A Complete Bibliography of the Publications of Jonathan Michael Borwein

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 August 2024 Version 1.93

Abstract

This bibliography records publications of Jonathan Michael Borwein.

Title word cross-reference

#11418 [BB09l]. #13553 [Bor81a].

(a, b) ← ((a + 3b)/4, (\sqrt{ab} + b)/2) [BBxxb]. (a, b) ← (\frac{a+3b}{4}, \frac{\sqrt{ab}+b}{2}) [BB89b]. (G) [BBL99]. 1/π [BB87b, BB88d, BB93d]. 24 [BB16o, CKM^{+}16]. $25 [BB93g]. $27.95 [BB91d]. $30.00 [Coh15]. $44.95 [BC96]. $45 [Zei05]. $45.00 [Sha05]. $49 [Zei05]. $49.00 [Ban10, Sha05]. $49.95 [Ber88]. 5 [Ade13, ZS12]. $59.50 [Bor06o]. 6 [ZZ14]. $65 [Odl11]. $69.95 [Bai91]. 8 [BB16o, Via16]. $99.00 [Bor09b]. [na + b] [Bor91n]. [nα + γ] [BB93e]. ^{∗} [BFG03]. b [BBG04b]. R [DL02]. C^{1} [BKW02, BFL02]. W [BL16]. D_{4} [Sol95]. DAD [BLN94b]. E_{6} [Sol95]. E_{8} [Sol95]. ℓ_{0} [BL11]. ℓ_{1} [XWQ14]. ϵ [LS00, YS00]. G [BBL97c]. k [BBB96b, BBB96c, BBB97d]. L [BB15c, BB07c]. L^{1} [BZ94b, BZ97, Hon85]. l^{∞} [Hon85]. l^{p} [Bor97g, Bor98g].

L_{1} [BL93b, BV97]. L_{1}(Ω, μ) [BF93d]. L_{1/2} [WSL16]. L_{p} [BTBT88, BBL10]. n [BB84d]. P [Alt20, BLS^{+}16, BLS^{+}17, BLS^{+}18]. π [AW97, ABBS12, Bai88, BBC^{+}11a, BBC^{+}12b, BBC^{+}12c, Bai16b, BBMW16, BB83, BB84b, BB84c, Bor85b, BB86b, BB86c, BB89a, BG96a, BB96d, BG97b, Borxx, BB11-31, Bor14p, Bor16n, Gan14, GG07, Gui08, Nim15, TK97, Wei15]. π^{2} [BBMW11, BBMW13]. q [LL01, PP11, War03]. R^{n} [BBW96]. \sqrt{5} log ϕ [Ade14b]. t [AX20]. θ(z, q) [HGB93]. × [BFG03]. ε [Bor82c]. W [Bor16l, Bor16m]. Weak^{∗} [BF95b]. x_{n}: = M(x_{n−1}, x_{n−2}, ⋯, x_{n−k}) [Bor94a]. xy + yz + zx [BC98a, BC00]. ζ(2n + 2) [BBB05, BBB06a]. ζ(4) [BB95f]. ζ(4n + 3) [AG99, BB97c, Bor97v, Bor97w, BB05f].

-analogue [PP11]. -ary [BBG04b]. -designs [AX20]. -elliptic [LL01]. -fold [BBB96b, BBB96c, BBB97d]. -function [BKW02]. -linear [DL02]. -regularized [XWQ14]. -Series [BB07c, BB15c]. -smooth [BFL02]. -Spheres [BLS^{+}17, BLS^{+}18, BLS^{+}16]. -subgradients [Bor82c]. -trinomial [War03]. -Variational [YS00, LS00].

0 [BC96, Bor06o]. 0-12-558630-2 [BC96]. 0-19-850763-1 [Bor06o]. 0-387-29570-4 [Bou06]. 0-387-87820-3 [Bor11-38]. 0-691-14247-5 [BO11b]. 0-89871 [Bor05g].

1 [BLN94a, Bor06o, Bor11-38, Bou06, Sha05]. 1-56881-136-5 [Sha05]. 1-56881-211-6 [Sha05]. 10 [Bai17e]. 100-Digit [Bor05g]. 100th [BB13v]. 11th [CGM95]. 12 [BB12-49]. 125th [AAB12]. 14th [IEE08]. 17th [IEE08]. 1880-2 [Bor09b]. 1983 [SBW84]. 1987 [AAB^{+}88]. 19th [Hd12].

2 [BC96]. 2000 [Tod03]. 2000j [BZ02a]. 2001 [BB12u, BB12n]. 2002 [KG04]. 2012 [BBL^{+}13]. 2013 [BB13v, BS14a]. 2014 [BBC^{+}14a]. 2016 [BBS17]. 2017 [Bai17e, BBB^{+}20, BE16]. 20th [BB12x, IEE08]. 21st [BB12p, BB12q, BB12-48, BBC^{+}14a, BBG03, Bor03z, Bor03-27, Bor03-28, Bor03-29, Bor04-27, Bor04w, Bor04x, Bor04y, Bor04z, Bor09r, Bor10a, HF05, Hoa05, R^{+}05, Zei05, BB04b]. 25 [Bai17a]. 2nd [Bou06].

3/14/15 [BB15t]. 38 [BZ02a, BZ02b].

4 [Bor81a]. 4N [Bor97q]. 4th [HY14].

5 [Sha05]. 51 [Bor81a]. 561-X [Bor05g]. 5th [BF06b].

60th [BBB^{+}13]. 6430-6435 [BSZ^{+}83].

7th [KG04].

8 [Zăl86]. 80th [Ano15]. 85h [Zăl86].

90d [BBB97a]. 978 [Bor11-38, Bou06]. 978-0-387-29570-1 [Bou06].

978-0-387-87820-1 [Bor11-38]. 978-0-691-1 [Bor09b]. 978-0-691-14247-0 [BO11b]. 978-0-88385-574-4 [Coh15]. 978-1-56881-271-7 [Odl11]. 978-1-56881-410-0 [Ban10].

= [lL09].

A. [BS14b]. AAECC [CGM95]. AAECC-11 [CGM95]. AARMS [Bor05d, Bor05e, Bor07a]. Abel [BB13a, Bor03o]. ability [BB11q]. Abraham [PR92]. Absence [BS11b, BS10b, BS10c, BS10d, Bor10i, Bor10j, Bor11r, Bor11s]. Absolute [BY84]. Abstract [BW79a, BW79b, BW81c, BW81b, BW82a, BW82b]. abundant [BB12-27, BB12e]. Academic [BC96]. academics [BBLZ16d]. Acceleration [BC18b]. Access [Bor04e, Bor04i, BB05e, Bor07d]. accuracy [Bor05g]. Accurate [BB14i, BB14h, BBLZ14b]. ACE [Bor05-28]. ACEnet [IEE08]. Action [BBC^{+}07b, Bor07m, Odl11, Lor09]. Activated [BBB^{+}96a]. Active [BL99]. Actually [Bor11g, BB12-36, BBWY11c, BBWY12c]. Acyclic [BW06]. Aczel [BB15d]. adaptive [FN15, LW18, LW19, NFB17a, NFB17b, QYX14, ZH06]. add [BB11f]. Addenda [BC15b]. Addendum [BZ02a]. Addition [BG95a]. Adjoint [Bor83a, BBWY11e, Zăl86]. admit [BV94a, BV96a, BV96b]. Adrian [Bou06, Tod03]. Advanced [Bai91, BL87, Ber88, BSZ^{+}83, BB85, Bor85a, BN86, Bor03b, Bor03c, Bor03a, Bor04f, Bor04g, Bor04h, Bor04e, Bor04d, Bor04a, Bor04b, Bor04c, Bor04i, Bor06d, Bor06b, Bor06c, IEE08, Sch85, SB87, SH87, SBW84, Bor06-28]. Advances [AHLC^{+}17a, AHLC^{+}17b, BBC10]. Advancing [KAA^{+}15]. Adventures [Bor15d, Bor16a, Bor17a, SZ20, Bor97v, Bor97w, Bor98q]. Advice [Bor03-30, BBLZ15d]. Advising [Bor03-30]. Aesthetics [Bor01a, Bor01b, Bor01c, Bor01d, Bor06e]. Affine [BGMS21, BW81a]. Affleck [SZ14]. AG [Bor10z, Bor10-27, Bor11-32]. Again [BB14c, BB15-31, BB13y, BB14w]. Age [Hol20, PR92, BB12-51, BB12-52, BB13-35, BB13-36]. ages [BB10g]. AGM [Ber88, Wim88, BB87d, BB88c, BB91c, Bor95c, BB98b, Bor03d, Bor03e, Bor03f, Bor04-30, Bor04-29, Bor04-28, BCF04, BC04a, Bre17, Bre20a, Lor08, Sol95, Ask88, Cas99]. Ago [bVP21]. agree [BB15m]. AI [BBLZ16a]. Aided [BB92b]. al [Gan17]. Alarm [BB12o]. Alexandria [SV14]. Alf [BSZ13]. Algebra [BB12p, BB12q, BB12-48, Bor11-29, CGM95]. Algebraic [BK05, Bor09-27, LY18, SV20, BBCP04, BB84d, BB87b, BLY13, CGM95]. algebras [KMY00]. Algorithm [AC18, Bai88, BB09j, Bor09c, BS11b, CZX21, Fin95, SV20, WSL16, Bai16b, BB93a, BB94a, BBL97a, BNCB99, BJCW13, BS10b, BLY13, BLY14, Jal24, JY12, Kom00, Kom02, Kom04, MP18, Pos13, QYX14, TK97, XSW12, XWQ14, ZL22]. Algorithms [BB95c, BB96b, Bor99w, Bor09p, Bor10c, Bor10d, Bor10r, BBC03, BBC^{+}11b, BB84b, BB86c, BL97, Bor98n, BL00b, Bor09-29, Bor11i, CGM95, Gui08, Gui16, Gui17, TK97]. Alignment [HMM20]. Alliance [BB13-44, BB13-43]. Almost [Moo18]. along [BB13-47]. Also [BB16r, BB16q]. Alternating [BB86a, Bor10c, Bor10d, Dil21, HNP10, BB93a, BB94a]. Alternative [Bor85c, BBG95b]. always [BB15m]. am [Bor11m, Bor11n]. America [Coh15, BB09n, Bor12u, Bor12v]. American [bVP21, BC15a, BC16]. Amir [BB15d]. among [BF94a, BF95a]. amongst [Bor94b]. AMS [BE16, AMM10, Bai16a, Jac09]. Anal [BZ02b]. Anal. [BZ02a]. analogue [PP11]. Analogues [Hir17, BBG93b, HGB93]. Analysis [Alt20, Ano15, ABMMY13, BBKL16, BBKL17, BBB^{+}20, Bor72, BBS89, BB92b, BLLN94, BI95, Bor96a, Bor99a, BMS99b, Bor99u, BL00a, Bor00v, BZ05, BM07d, Bor08i, Bor08j, Bor09z, BLY13, BLY14, Bor16j, Bor16k, Bor16l, BG16c, Bor16z, BLT17, Bor17b, Bou06, Roc20, Tod03, ABMMY14, BBLZ14q, BBMW17, Bor81b, BS86, BS87, Bor87k, Bor93p, Bor94h, Bor94i, Bor94j, Bor94k, BL94a, BZ94b, BS95, BI96, Bor97p, BS97a, BTZ97, BZ97, Bor98k, Bor98l, Bor98m, BZ99a, BZ99b, Bor99t, BL06, Bor06-30, BZ06, BM07c, Bor09l, Bor10p, Bor13-31, Bor13-34, Bor13-35, Bor13-33, BG15a, Bor15f, BLT15, BG15c, Bor16i, BLT16, BL16, BG18a, DLL05, IMR92, LY21, MTCB99, Bor92b]. analyst [Bor93b]. Analytic [Ber88, BB87d, Dil21, Wim88, Bor91e, Bor91f, Bor91g, Bor91j, Bor91k, Bor91i, Bor91l, Bor91m, Bor92e, Bor92f, Bor98k, BB98b, BZ99a, BZ99b]. analytical [BBB^{+}13]. Ancient [BB11h, BB12r, BB16d, JJ20, SV14]. Andrea [BB15u, BB16i]. Andrew [BE16]. Ann [Bor12a]. anniversary [AAB12, IEE08]. Answer [DP18]. Anthology [BC15a, BC16]. Anthony [BS14a]. antiderivative [BBB^{+}07]. antiproximal [Bor81a]. Antiproximinal [Bor81a, BJSM00, BJSM02]. Antisocial [BB15-31]. anxiety [BB12v, BB12i]. any [Ade11]. anyone [BWB97]. Apéry [Bor05f, AG99, BBB05, BBB06a, BB96c, BB97c, BB05f, BB05c, CS21]. Apéry-like [AG99, BBB05, BBB06a, BB96c, BB97c, BB05f, BB05c]. Apéry-type [Bor05f]. APICS [Bor89a]. APICS/FRASER [Bor89a]. appeal [Bor11v]. Appel [BB13e, BB13f]. Application [TB80, dPB21, BT14a, HYG09, Li15, LW18, LW19]. Applications [ABMMY13, ABMMY14, AI18, Bor96a, BL97, BL00b, Bor04-31, Bor07n, Bor09-28, Bor09-31, Bor12-31, Bor16u, Bor16v, Bor16w, Bor16x, Bor16y, CFG^{+}18, CG18, Geo05, HMM20, LY18, BBBG08, BB15c, BB16b, Bor79g, Bor86a, BP87, Bor87l, Bor88m, Bor88n, Bor89d, Bor90y, Bor90z, Bor90-27, Bor90-28, BZ94b, Bor94m, Bor95o, Bor95p, Bor95v, BZ95, BZ96, BZ97, BTZ99, BZ99c, BZ02a, BZ02b, BCFR04, Bor14e, BT14b, BT17, RZ15, BS87]. Applied [BB15n, HDG^{+}15, BLY13, BLY14, CGM95]. Approach [Alt20, BBC^{+}11a, BBC^{+}12b, Bor10h, Bor11p, GN16, Ade12, BTBT88, Bor77a, Bor79d, Bor93b, BMN98, BMW99a, BMW99b, BMW99c, BMN00, BMW01, BZ16]. Approaches [Bor09-30, Bor09-28, Bor09-31, Bor09-29, Bor09-27, BaO12, Bor90e, Bor90f]. Approaching [BC18b]. Approximate [BBW96]. Approximating [BB89a, BG97a, Bor85b]. Approximation [BM07d, Bor13g, Bor13h, Bor13i, Bor79g, Bor91h, Bor92d, Bor13e, Bor13f, Kom00, Kom02, Kom04, SBW84]. Approximations [BG96a, BG97b, BBB97c, BBB00b, BBB04b, BBB16, BBB97a, BB84d, BB84c, BS85, BB87a, Bor87f, BBB89]. Arabic [BS14b, BS14a]. arbitrary [Ade11, BBMW17, BBB96b, BBB96c, BBB97d, Bor14t, Bor15o, Bor15p, Bor15q]. arc [BBC08a]. Archimedes [Bor12o, Bor14p, Bor14s, Bor16o]. arcsin [BC07]. arctan [Nim15]. arguments [BV93b, BV94d]. arising [BB13g, BBCZ13, Cvi10]. Arithmetic [BB13q, BB84a, BLM96, BB97b, BLM97, BB00b, BB04a, Bor10-29, Bor11-33, Bor12b, BB11-29, BB15p, Bor87d, Bor88a, Bor88b, Bor88c, Bor88d, Bor88e, Bor88f, Bor89e, BBG93b, Bor10e, BNSW11, BB16t, Zah06]. Art [BB12u, BB12n, BBLZ13c]. Articles [BC15a, BC16]. ary [BBG04b]. Asen [Bor11-38]. aspects [BBBL97, BBBL98a, BBBL98b, Bor12b]. Asplund [Bor93a, BW07, Bor07b]. Assessment [MTCB98]. assets [BCM02, BCM03]. assisted [BB05a, BB08c, Bor93c, Bor93d, Bor06h, Bor07g, Bor08d, Bor08e, Bor08f, Bor08g, Bor09a, Bor09e, Bor09f, Bor09g, Bor09u, Bor12-33]. Associated [BCLM16, BCLM17, Liu01]. Association [BBLZ14p, Coh15, KG04]. Astonishing [BGMS21]. Astronomy [Fer91]. Asymptotes [BB93f]. Asymptotic [BBD97, BBD00, BBD04, BBD89, BBD16]. Asymptotics [BL92a, BSxx, Bor07i, Bor07j, BBC07c, BBC08b]. Atlantic [Bor04j, Bor04k]. attractors [BR16]. Aubin [Bor92b]. August [BF06b, BBS17, HY14, SBW84, BS16a]. Australia [BBB^{+}20, Bea13, BB13n, BB13-27, BB13-39, Bai17a, Bor10-30, Bor13d, Bor13a]. Australian [BB12k, BB13m]. Automated [BBK14]. Automatic [Bor87a]. Autour [Dev9x]. average [Zah06]. Averaged [BLT17, DLR20, BLT15, BLT16]. avoid [BBL^{+}16b]. Avoiding [Bor04-32]. Avriel [Bor90b]. away [BB11n, BB11d, BG16a, BG18b].

B [Ber88, Coh15]. Back [BBLZ13a, Bor11o, Bor11a, PD18]. Back-Testing [BBLZ13a]. Background [BB15t, BJL^{+}08]. Backing [Bor06f]. Backtest [BBS^{+}16a, BBLZ17, BBLZ14c, BBLZ14k, BBLZ14s, BBS^{+}15a, BBL^{+}16b, BBL16a, BBL16c]. backtested [BBLZ14a]. Bad [BB12t, BB12a]. baffle [Bor15a]. Baghdad [SV14]. Bailey [Gan17, Hoa05, Odl11, Sha05, Zei05, Bai16a, BJCW13, BCJW13, BE16, Fin95, PP11]. Baire [BS84a, BMW99a, BMW99b, BMW99c, BMW01]. Balkanica [Bor81a]. ball [BKW02, BS10a]. Banach [Bor81a, BB95a, BBC00a, BBC01, BBWY11a, BBWY11b, BBWY12a, BBWY12b, Bor78a, Bor82e, BS83, BS84b, BS86, Bor87m, BS87, BG87, BF89b, Bor91d, Bor92g, Bor92h, Bor92a, BL93a, Bor93f, Bor93g, BV94a, BV94b, Bor94h, Bor94i, Bor94j, Bor94k, BF94b, BN94, Bor95a, Bor95b, BZ95, BV96a, BV96b, BZ96, BFV97, BV97, BJ97, BTZ97, BJ98, BJSM00, BV00b, BV01, BG01, BJSM02, Bor02d, Bor02e, BG03a, BBL04, BM07c, Bor07x, BM07d, BE08, BG09, BGHV09, BV10a, Bor13e, Bor13f, Bor13g, Bor13h, Bor13i, BG15b, BG16b]. bang [BB14m]. barometer [BBLZ16b]. Barrow [BB09g, BB93g, Bor09b]. Bartle [BD03]. Barycentres [TB80]. Barzilai [IP17, IP18, RS02, AX20, AP16, AK22, BL17a, BL17b, CZX21, CPRZ20, DL02, DLL05, DF05, DHSZ06, DABY15, DK16, FN15, Fle05, GDT15, GS02, HNP10, HYG09, HD07, HLZ14, HL15a, HLZ15a, HLZ15b, HL15b, HLY16, HDL21, Jal24, JY12, JD13, KJR16, La 09, LLS11, LZ14, Li15, LW18, LW19, LY21, LL13, Mar91, MR96, MP18, MPB16, NWY09, NWY10, NFB17a, NFB17b, PT14, Pos13, PD18, QYX14, Ray93, Ray97, SI16, SD15, WM07, WSdSY15, WSL16, XH08, XSW12, XWQ14, XC11, YW12, ZH06, ZSQ10, ZL22, ZSZ16]. base [BB11-29]. base-10 [BB11-29]. Based [BB06a, BB08d, Bor06j, Bor06k, SI16, BCJW13, FN15, Jal24, JY12, LLS11]. Bases [Zhu91, Ade11, BBG95b]. Basic [BMS99b, BBLZ13b, BLY13, BLY14]. Battle [BB15i, BB15h]. Bauschke [Vir14]. BBP [AL10, Ade10, Ade11, Ade12, Ade13, Ade14a, Ade14b, Bor11i, Cha03, GG07, Lup02, Nim15, Wei15, ZS12, Zha13, ZZ14]. BBP-formulae [Cha03]. BBP-functions [Lup02]. BBP-Type [Ade14a, Ade14b, AL10, Ade10, Ade11, Ade12, Ade13, Nim15, Wei15, ZS12, Zha13, ZZ14]. Be [BB14f, BB12-45, BB14g, Bor15n, Bor16e]. beautiful [BB14-30, BB14-31, Bor15n, Bor16b, Bor16e]. Becomes [BB13-44, BB13-43]. been [BB11-29]. Behavior [ABT15, ABT16]. Behaviour [Bor99m, Bor99n, Bor00l, BDT16, BG16a, BG18b]. being [BB93g]. beispielorientierte [BD11]. Beliefs [BB09c]. Believing [BB12-33, BB12-32]. Bello [BS14b, BS14a]. bells [BB12o]. Benson [Yan94, Zho12]. Berggren [SV14]. Bernoulli [Dil20]. Bernstein [SZ14]. Bertinoro [ABD03]. beset [BB13-41]. Bessel [BBBG08, Bor07h, BS07, BBC08a, Bor08h, BS08, Cra04]. Bessel-integral [Cra04]. Best [Bor13e, Bor13f, Bor13g, Bor13h, Bor13i, BL91a, Mic03]. better [BB16e]. betting [BR14b]. Between [BB13-44, Goo20, BB13-43, BV93a, BV94c, BR14c, BR14a]. Beyond [Bor12o, Bor14p, Bor14s, Bor16o, ES01]. biconjugate [BV00a, BV02]. biconvex [Bor86b]. Big [BB14e, BB14m, BB15t]. bigger [BBWY11d, BBWY13]. Billion [BBB97c, BBB00b, BBB04b, BBB16, BBB97a, BBB89, BBxxc]. billionaire [BBLZ16d]. Binary [Ade14b, BCM20, Cha03, Ade10, Ade12, Ade13, BBCP04, BG95a]. Binational [IMR92]. Binomial [BBK00a, BBK00b, BBK01, BG05, Cos17, GG07]. Birthday [BBB^{+}13, BB13v, Ano15]. bit [Cra12]. blank [Bor11e]. Blaschke [BBFG00, BBFG01]. blessings [BB11e]. blue [Tre13]. Boltzmann [BH95]. Book [Abb00, Ask88, Bai91, BB09c, Ban10, Ber88, Bor90b, Bor92b, BB93g, BC96, Bor05g, Bor07c, Bor11-38, BS14b, Bou06, Cas99, Coh15, HF05, Hoa05, How14, Lor90, Lor09, Odl11, Rob06, Sha05, Wim88, Zei05, BBB97b, BBB00a, BBB04a, BB91d, Bor09b, BO11b, Tod03, Abb00, BBB03, Rob06]. Books [Bai91, Bou06]. Boole [BCM09]. boost [BBLZ13d]. Borchardt [Bor88f]. Bornemann [Bor05g]. Bornological [BFV93a, BFV93b, BF93c, BF95b]. Bornology [BGM18]. Borwein [Bai91, Ban10, Ber88, BO11a, Bou06, Coh15, Fin95, Hoa05, How14, KMT16, Odl11, PP11, Sha05, Tod03, Zei05, AG99, AX20, Ano15, Ano16, Ara07, Ara08, AP16, AK22, BCM20, BBB^{+}13, Bai16a, Bai16c, Bai16d, Bai17a, Bai17b, Bai17c, Bai17e, Bai20, Bai21, BL17a, BL17b, BWY10, Bej94, BJCW13, BCJW13, BS17, Bor03-32, Bor04n, Bor08s, BaHO20, BV24, BE16, Bre17, Bre20a, CW16, Cha16, CZX21, Cra04, Cra12, CPRZ20, Cvi10, DL02, DLL05, DF05, DHSZ06, DABY15, DK16, Dev9x, Dev17, Fab89, FN15, Fle05, FK00, GN16, GDT15, Geo05, GS02, Gui08, Gui16, Gui17, HNP10, HYG09, Hir17, HC09, HD07, HLZ14, HL15a, HLZ15a, HLZ15b, HL15b, HLY16, HDL21, IP17, IP18, Jal24, JY12, JD13, JN03, KMY00, Koh01, KJR16, KPS16, KPS17, La 09]. Borwein [LS00, LLS11, LZ14, Li15, LW18, LW19, LY21, Liu01, LL13, MW16, Mar91, Mer15, Mil90, Mil89, MW12, MR96, MP18, MPB16, MR11, NWY09, NWY10, NFB17a, NFB17b, Osb05, PT14, Pos13, PD18, QR07, QYX14, RP09, Ray93, Ray97, RS02, Rei02, SZ14, SI16, SD15, TK97, Tha02, Vir14, WM07, WSdSY15, War01, War03, WSL16, XH08, XSW12, XWQ14, XC11, Yan94, YS00, YW12, Zah06, Zăl86, ZH06, ZSQ10, ZL22, Zha10, Zho12, ZSZ16, Zhu91]. Borwein-based [JY12]. Borwein-Like [WSL16, DABY15, GDT15, Gui17, JD13]. Borwein-type [Gui16]. Borweins [AB15, AAW06, Bai88, Bai16b, Kom00, Kom02, Kom04, LL01, Liu00, XY12]. boson [BB13o]. bothered [BB12c]. bound [BMS13, BSM13, XH08]. Boundaries [Goo20, BS86, BS87]. Bounded [BBT00, BBL97c, BBL99, Bor86e]. Boundedness [BMV06, BF89a]. Bounds [BB06a, BB08d, BLL94, Bor06j, Bor06k, CPRZ20, PT20, BBL97c, BBL99, BC09]. Box [BBC07a, BCC10, BBC10, DF05, ZH06]. box-constrained [DF05]. Boys [BBLZ14m]. Bradley [AG99, PP11, Zha10]. Brailey [Bor15e]. brain [BB12v, BB12i]. brains [BB14d]. Brainstorming [Bor98c]. Brave [BBB^{+}96a]. breakthrough [BB13-42]. Breakthroughs [BB14e]. Bregman [BB95b, BB97a, BB00a, BB01b, BBC03, Bor02b, BRS11, BML18]. Bregman-Type [BML18]. Brézis [BBWY11a, BBWY12a]. BRICs [BB11v]. Brief [Bor77d, BC15a, BC16]. Bringing [Bor03g]. British [BBJC97]. Broadhurst [Cvi10, Zha10]. Brooks [Bai91]. Brooks/Cole [Bai91]. Brother [Bor08s, Bor12a]. Brothers [Bre20a]. Brouwer [Bor98d]. Browder [BBWY11a, BBWY12a]. Brown [BB13d]. Brun [PT20]. Bucks [BB14e]. Budget [Bor14a]. build [Bor13d, Bor13a]. bump [BFKL00, BFKL01, BFL02]. Bumps [Bor04l, Bor04m, Bor06r]. Burge [War01]. Burgers [NFB17b]. Burnaby [BBJC97]. burns [BB09k, BB11z]. Bytes [Bor05-40].

C2C [Bor07d, BM07b, BJL^{+}08]. CA [BC96]. calculate [BBxxc, BB12-31]. calculated [BB12i]. Calculation [BB84e, BS17, Bor89b, Bor97h, DS20, TK97]. Calculator [BL92e]. Calculus [WB87, BZ98, BZ99c, BZ02a, BZ02b]. California [Bai91]. Calkin [Odl11]. Callaghan [BBB^{+}20]. campaign [BB12-42, BB12-43]. Can [BB10a, BB14f, BB14g, SW21, BB12-45, BBLZ13h, Bor15n, Bor16e]. Canada [KG04, BB13-27, BD95, Bor03x, Bor03y, Bor04f, Bor04g, Bor04h, Bor04p, Bor05l, Bor05m, Bor05n, Bor05o, Bor05p, Bor06l, Bor06-28, R^{+}05]. Canadian [Ber88, KG04, Bor03h, CW16]. Cancellation [BO78]. cans [Bor10o]. can’t [BB14-28]. cap [BBLZ13d]. Carathéodory [TB80]. care [BC13]. Career [Bai16d]. CARMA [Bor13b, Bor16h, Bor09s, Bor10s, BR10, Bor11j, Bor11k, Bor11l, Bor11m, Bor11n, Bor12c, Bor12d, Bor14d]. Case [BBBR16, BBBR17, BCF04, BC04a, Bor05f, Bor11f, Gan17, BL94a, Bor04-30, Bor04-29, Bor10z, Bor10-27]. Cases [BBFG00, BBFG01]. CAT [BBS10]. Catalan [BBMW11, BBMW13, Bor10x, BBGW11, Bor11-30]. categorical [BMW99a, BMW99b, BMW99c, BMW01]. Cauchy [RS02]. Causa [Bor99o]. causation [BR14c, BR14a]. Cebysev [BK80]. CEIC [Bor00a, Bor02c, Bor02o, Bor04-33, Bor06-36]. Celebration [BBB^{+}20, BB15-28]. centenary [AAB^{+}88]. Central [BBK00a, BBK00b, BBK01]. centres [BK80]. Century [BBC^{+}14a, BBG03, Bor03z, Bor03-27, Bor03-28, Bor03-29, Bor04-27, Bor04w, Bor04x, Bor04y, Bor04z, Bor09r, Bor10a, KAA^{+}15, R^{+}05, BB12p, BB12q, BB12x, BB12-48, BB15-28, BB04b, Hoa05, Zei05, HF05]. Certain [BK05, BBS20, BB15q, BBS89, BBCM07b, BBS15b]. chain [BM96a, BM97d, BM98a]. Challenge [Bor05b, Bor05g, Bor09z]. Challenges [BB08b, BB13q, BB14a, BBC^{+}14a, BB01d, Bor01e, BB01e, Bor02s, Bor02t, Bor05h, Bor05-27, Bor08i, Bor08j, Hol20, BB11b, BB14t, Bor10p]. Champaign [AAB^{+}88]. Change [BB12-33, SW21, BB12-32]. changer [BB15b, BB15o]. changing [BB15v, Bor96k, Bor97x]. Chaos [BR12, BR13b]. Chapter [BM07c]. Character [Bor14e, Dil21, BB16b, BZB08]. Characterization [BW81c, BF94a, BF95a, BBP03, Bor84b, BRS11]. Characterizations [BW79a, BW79b, BW82a, BW82b, Bor94b, BFV97, DGLV20, BV10b, How14]. characterizes [BO78]. charlatanism [BBLZ14l, BBLZ14s]. Charnes [PR92]. chart [BBLZ14n]. Chasing [Bor03o]. Chebysev [Bor13i]. Chebyshev [Bor06u, Bor07y, Bor08t, Bor13e, Bor13f, Moo18]. Checkerboard [Bor13k, BH19, PHB14]. checking [BB11x]. cheque [Bor11e]. cherry [BBLZ14h]. children [BB12k]. CHIP [BBT98, BBT00]. Chiropractic [Bor11o, Bor11a]. Choi [HC09, Osb05, Tha02]. choice [Bor90c, Bor90d, Bor91b, Bor91c, Ray93]. Chronology [Bor04n]. ci [BB95e]. Circle [Bor94f, Bor90o, Bor90p]. City [IMR92]. Clarendon [BB93g]. Clarke [BF94a, BF95a, BW95a, BW97a, BM97e, BM97f, BW98b, BW00, BM00, BGV02, BW05b]. Class [BBBC07, BB93d, CFG^{+}18, BBC06, BG03b, Bor07e, LZ14]. classification [BFV94c]. Classifications [BFV94b, BFV94a, Bor94l, Bor95s]. Classroom [CDH^{+}21]. Clausen [BBK00a, BBK00b, BBK01]. Clearing [BR14c, BR14a]. Climate [BB12-33, BB12-32]. Climbing [BB11d]. Closed [BF95c, Bor10f, Bor10g, BC13, BBL97b, BS86, BS87, BFG03, BS16b]. closedness [BM09, BM10]. closure [BY12d, BY14b]. cm [Bai91, Ber88]. CMS [Bou06, Ano16]. co [IEE08]. co-hosted [IEE08]. Coast [BLM^{+}07, BJL^{+}08, BBJ12]. Coast-To-Coast [BBJ12, BLM^{+}07, BJL^{+}08]. Cobzas [Bor81a]. coderivatives [BBW96]. codes [CGM95]. coefficient [Cos17]. coefficients [BS17, BL05, BL08, War03]. coffee [BB13-40, BR13a]. coincide [BBWY11e]. coke [Bor10o]. Cold [BB15f, BB15g]. Cole [Bai91]. Collaboration [Bor03b, Bor03c, Bor03a, Bor04a, Bor04b, Bor04c, Bor04i, Bor05k, BLM^{+}07, BM07b, Bor09w, Bor09x, Bor11g, Bor11-37, Bor12n, BBJ12, Goo20]. Collaborative [Bor98e, Bor01f, Bor04e, Bor04d, Bor06d, Bor06b, Bor06c, Bor06g, Bor07f, IEE08, Bor16h]. Collaborator [Bor14i, Bor14j, Bor14k, Bor14l, Bor14m, Bor14n, Bor15h]. collapse [Bai17d]. collection [Bor97e]. college [BWB97]. collide [BB14-29]. Collins [BB95d, BB02]. color [BB13e, BB13f]. Colorful [BB13f, BB13e]. Coloring [AC18]. Columbia [BBJC97]. combat [BB12-29]. Combinatorial [ABT13b, ABT14b, BBBL97, BBBL98a, BBBL98b]. Combinatorics [GI01]. come [BB12g, BB13-47]. comes [Bor15b]. Coming [Bor07w, Bor07-32, Bor08n, Bor08o]. Commemorative [Bai17a]. Common [BLT17]. Communicating [BMPR02, BRR08, Ban10]. communications [Bor92c]. Community [Bor03p, BS05]. compact [BRLZ99, BLZ99, BRLZ00, BLZ01]. Compactly [BLM99, BLM00]. compactness [BF93c, BF95b]. Companion [HDG^{+}15, Bor09b]. Comparing [DLR20]. comparison [BGL93]. compendium [BBB96b, BBB96c, BBB97d]. Competition [Bor77d]. Complementarity [AI18, BD86, AR13, Bor84a, Bor85c, Bor87e, BD89, HLZ14, HLY16, KJR16, LLS11, LZ14, Li15]. complementary [BC09]. complete [BZ92]. completed [BB14j]. completely [SZ14]. Completeness [Bor83b, QR07]. Completion [ABT13a, ABT14a, CZX21, Bor13j, Bor14f, Bor14g, Bor15g, Bor16p]. Complex [BC04a, BMN98, BMN00, Bor04-29, Bor10-27]. Complex-Parameter [BC04a]. Complexity [BB84e, BB87d, BB88e, BBxxa, BB17, BB98b, Ber88, Wim88]. complicated [Bor14z, Bor16-27]. component [LY21]. composite [HL15a]. Composition [KMZ^{+}03]. compositions [BM97d]. Compound [BB93f]. Comprehensive [BS14a, BS14b]. Compressed [BB13g, BL17a, BL17b, Bor09c, Bor10h, Bor11p, QYX14]. compression [LY21]. compressive [XWQ14]. Computation [Bai88, BB08a, BBMW11, BB12y, BBC14b, BBC^{+}14a, BB15b, BB15a, BB16a, BB16b, BBMW17, BB18, BB84a, BB97b, Bor99g, Bor99h, Bor99i, Bor99j, Bor99k, Bor99l, Bor99v, BB00b, Bor00b, Bor00c, Bor00d, Bor00e, Bor00f, Bor00g, Bor00h, Bor00i, Bor00t, Bor01i, Bor01j, Bor01k, Bor03b, Bor03c, Bor03a, Bor04a, Bor04b, Bor04c, BB04a, Bor05-41, BH06, Bor07h, Bor07t, Bor07u, Bor08h, Bor09h, Bor09i, Bor09t, Bor10l, Bor10m, Bor11t, Bor11x, Bor11y, Bor11z, Bor11f, Bor11-27, Bor11-28, Bor12e, Bor12f, BMS13, BSM13, Bor14h, Bor14i, Bor14j, Bor14k, Bor14l, Bor14m, Bor14n, Bor15h, Bor16n, GI01, LLT18, MTCB99, BBP97, BB10d, BB11j, BBB12, BBMW13, BB15c, BB15o, Bai16b, BBMW16, BB16l, Bor90q, Bor90r, Bor90s, Bor90t, Bor90u, Bor90v, Bor90w, Bor90x, Bor93h]. computation [Bor93i, Bor94n, BMN98, BBxxc, BMN00, Bor10q, BB16t, IP17, IP18]. Computational [BB09a, BBB^{+}13, BBL^{+}13, BBBR16, BBBR17, Ber88, BB87d, BLLN94, BBC98, BS99d, BBC00b, Bor00s, Bor02j, Bor02k, Bor03l, BBG03, Bor05h, Bor05-38, Bor05-39, CC20a, GN16, Gan17, Hol20, SBB13, Wim88, Zei05, BB09e, BB17, Bor93p, BB98b, BS99b, BS00, BBG04a, BB10l, BLLN95]. Compute [BBB97c, BBB00b, BBB04b, BBB16, BBB97a, BBB89]. computed [MTCB98]. Computer [BB05a, BB08c, BBKL16, BBKL17, Bor92i, BB92b, Bor93c, Bor93d, Bor06h, Bor07g, Bor08d, Bor08e, Bor08f, Bor09d, Bor11-29, Bor14i, Bor14j, Bor14k, Bor14l, Bor14m, Bor14n, Bor15h, Dev20, BB11o, BB12j, BB12-44, BB12-36, BB13-35, BB13-36, Bor91e, Bor91f, Bor91g, Bor91j, Bor91k, Bor91i, Bor91l, Bor91m, Bor92e, Bor92f, Bor08c, BD09]. Computer-assisted [BB05a, BB08c, Bor06h, Bor07g, Bor08d, Bor08e, Bor08f]. computers [BB12s, BB12m, BB16e, BB16s]. Computing [BBLZ13a, BBS16b, Bor98h, BB01d, Bor01e, BB01e, Bor02s, Bor02t, Bor03h, Bor04f, Bor04g, Bor04h, Bor05-28, Cal16, IEE08, JWDS^{+}14, Bor92k, Bor92l, Bor92m, Bor98q, Bor03x, Bor03y, Bor06-28, BS11c, BS12a, Bor05g]. Conant [Bai16a, BE16]. concave [Bor86b]. Concavity [SZ81, Bor90b]. concept [BRS11]. Conditions [BTZ95, BBY12, BBY14, LY18, Bor82b, BZ88, BL91d, BTZ98]. Cone [BW81a, BW05a, BW81d, BB84f, BS89, BBL04, BG09]. Cone-convex [BW81a, BW81d]. Cone-monotone [BW05a, BBL04, BG09]. Cones [Bor77c, Bor78a, Bor86d, Bor87c, Bor87b, EB08, BO76, Bor78c, Bor80a, BM09, BM10, Zhu91]. Conference [Ano15, Bai17a, Bea13, HY14, lL09, AAB^{+}88, ABD03, BF06b, KG04, RZ15]. Confidence [BBLZ14g]. confirm [BB14m]. conflicted [BBLZ15d]. conformation [BT14b, BT14a, BT17]. confusing [BB10b]. confusion [BR14c, BR14a]. Congress [Bor05i]. congruence [Cos17]. Conical [BB98a, BB99b, BBL97c, BBL99]. Conjecture [DP18, Osb05, BB14j, BB14s, BBBG94, BBBG95, BW95b, BBBG96, BW97b, BMS13, BSM13, Cvi10, HC09, RP09, Tha02, War01, War03, Zah06, Zha10]. conjectured [ABBS12, BB11-31]. Conjugate [BPT84, SI16, BB96a, BB99c, BBWY11d, BBWY13, BV09, DK16, MP18, WSdSY15, XSW12]. Conjugates [BH06]. conjugation [BH09]. Consequence [Bor79b, Bor81e]. Consequences [Bor87c, Bor86d, Bor87b]. conspiracy [BB16g, BB16h]. Constant [BCM20, BBC09, BBMW11, Bor95q, Bor95r, Bor10x, Bor11-30, PT20, BBC97a, BBC9x, BBMW13, BBT85, BVW01, BVW03, BBGW11, Cra12]. Constant-Length [BCM20]. constants [Ade10, Ade12, BBP97, BB12-40, BBMW17, BBGPxx, GG07, Mer15]. Constrained [AI18, BTZ95, CPRZ20, AK22, BTZ98, DF05, XH08, XC11, ZH06]. constraint [BW79b, BW82a, BW82b, BW86]. constraints [Bor77a, BW81a, PD18, ZL22]. constructed [BB11i, BB12-34, BB12-35]. Constructible [BV04]. Construction [BBWY11b, BBWY12b, GG07, AX20, BGW97, BGW98]. Constructions [BV12, Com18, How14, BV10b]. Constructive [BK04]. contained [Ara07, Ara08]. containing [BV97]. Continuations [Dil21]. continue [BB15z, BB15y]. Continued [Bor03d, Bor03e, Bor03f, Bor04-30, Bor04-29, Bor04-28, BCP05, BvdPSZ14, BCLM16, BHL16b, BHL16a, BCLM17, BBGPxx, BL05, Bor05j, Bor06i, BL08, Bor10z, Bor10-27, Bor11-32, BHL17]. continues [BBxxc, Bor15c]. Continuity [BGM18, Bor82a, Bor87a, BV00a, BV02, BW05a, BY12e, BY13c]. Continuous [BB95a, BTZ95, BB96a, BB99c, BBW07, BTZ98, BW99, BW01]. continuously [BFKL00, BFKL01]. contraction [Bor83b]. Contractive [RZ18]. Control [ANR18, BB15l, BZ94a, LY18, BZ94b, BZ97, NFB17b]. controls [BB15w]. conundrums [Tre13]. converge [Bor98d]. Convergence [BB93b, BBT85, BL91a, BL93a, BL93c, BV95a, BBP95, BBP98, BV9x, BY06, BST13, BLT15, BLT16, BLT17, Gil18, Lor08, Mar91, AB12, AB13, AK22, BB93a, BB90a, Bor88j, BF89c, BL91c, BV93a, BV93b, BV94c, BV94d, BH94a, BH94b, BV95b, BV95c, BV95d, BV96c, Bor09-29, BLY13, BLY14, BST15, DL02, HL15b]. Convergent [Bai88, AL10, Bai16b, BB83, Bor94a, TK97]. converges [Bor94a]. converging [BB86c]. converse [BW98a]. Convex [ABMMY13, BB95c, BB96b, BBL97a, BW79a, Bor80b, Bor81b, BT84, Bor87c, Bor90e, Bor90f, Bor90c, Bor90d, Bor91b, Bor91c, Bor93e, BV94b, BFV94b, Bor95a, Bor95b, Bor96a, BV97, Bor99a, BRLZ99, BLZ99, BV9x, BL00a, BV00b, BRLZ00, BV01, BLZ01, Bor01o, Bor02b, BL06, Bor06s, Bor08u, Bor09-28, Bor09-32, BV10b, Bor10k, Bor10-33, Bor11q, BV12, Bor13p, BG15a, Bor15f, BL15, Bor16i, Bor16j, Bor16k, BG16c, Bor16u, Bor16v, Bor16w, Bor16x, Bor16y, BG18a, BML18, CFG^{+}18, Moo18, Roc20, AB12, AB13, ABMMY14, BBS10, BBL94, BBL97b, BBL97c, BBL99, Bor79e, BW79b, Bor79a, Bor80e, BW81a, BW81c, Bor81c, BW81d, Bor81d, BW81b, Bor82a, BW82a, BW82b, BPT84, Bor84e, BT85, Bor86e, Bor86a, Bor86b, Bor87a, Bor87k, BP87, Bor88l, Bor89i]. convex [Bor90g, Bor90h, Bor90a, Bor90-40, Bor90-41, Bor90-42, Bor90-43, Bor91d, Bor91h, BFK91, Bor91r, Bor91s, Bor91t, Bor91u, BZ91, Bor92d, Bor92g, Bor92h, BL92c, BL92d, Bor92a, BBT92, BL93a, BF93a, Bor93f, Bor93g, BFV94c, BFV94a, Bor94g, BLN94a, Bor94l, BN94, BL94a, BF95c, BV95a, BV95b, Bor95m, Bor95n, Bor95s, BV96c, BLN96, BFV97, BMN98, BZ98, BLM99, BV00a, BMN00, BLM00, Bor00r, Bor01p, Bor01q, Bor01r, BV02, BV04, Bor05-33, Bor05-34, Bor05-35, Bor05-36, Bor05-37, BMV06, Bor06-33, Bor06-30, Bor06-34, Bor06-35, BZ06, BM09, BGHV09, BM10, BBY12, BY12a, Bor12p, BLY13, BLY14, BBY14, Bor14o, BY14a, Bor15i, BG15b, BG15c, Bor15r, BG16b, Gil18, NWY09, PD18, YW12, ZL22, Zhu91, CFG^{+}18, Bou06, How14, Tod03]. convex-concave [Bor86b]. Convexity [BBFG00, Bor07-28, Bor07-29, Bor07-30, Bor07-31, BS11b, BS15a, BB11a, BBC00a, BB00a, BBC01, BB01b, BO76, Bor77a, BO78, Bor78c, BBFG01, Bor07-27, BS10b, BS10c, BS10d, Bor10i, Bor10j, Bor11r, Bor11s, BY12d, BY14b]. convolutions [BBEM10]. Copulas [Bor13k, PHB12, BH19, PHB14]. copyright [BB10b]. correcting [CGM95]. correlation [BR14c, BR14a]. cosmic [BB09d, BB11d]. could [BB12j]. Counter [Bor17b]. Counter-examples [Bor17b]. counterexamples [BV10b, How14]. Counterpart [BB88c, BB91c]. Counterparts [BB15s, BBLZ15b]. counting [BB11e, BB93g]. country [Bor13d, Bor13a]. courses [BB12-44]. crackers [Bor11o, Bor11a]. Crandall [BB12-38, BB15c]. Crash [BB15x, BBLZ15f]. creation [BB09h, BB13-41]. Creationism [BB10c]. creationists [BB13t]. Creativity [Bor09o, Bor12n, Coh15]. Crime [BB15l, BB15w]. Criminology [BB13h]. crisis [BB12-41, BB12-53]. critical [BKW02]. cross [DS20]. Crossing [Goo20]. Crucible [Bor09d, Bor08c, BD09]. Cubic [BB84b, BB88c, HGB93, Hir17, AB15, BB86b, BB90b, BB91c, BBG94c, Bor95c, LL01, Liu00, MP18, XY12]. cultures [Sel16]. Cup [BR14b].

Curiosity [BB12g]. curve [Bor90e, Bor90f]. CUSCOS [Bor89c, Bor89d, Bor90y, Bor90z, Bor90-27, Bor90-28, Bor91a]. Cusps [Bor04l, Bor04m, Bor06r]. Cutters [DLR20]. Cyclic [BT13a, BT13b, BBB^{+}07, BBL94, BBL97a, BBL97b, BL08, BLY13, BLY14, BT14c, BT15, DHSZ06, HLY16, XSW12, ZH06]. cyclotomic [HC09].

D [BB09g, BB93g, How14, Odl11, Bor05-46]. D-DRIVE [Bor05-46]. DALBAR [BBLZ14j]. Damping [BC18b]. Danger [BB11c, BB13i]. dangerous [BB12l]. Dark [BBLZ14m]. Data [BB14i, BB15h, BB15i, BTZ95, Bor09c, CZX21, BB12-50, BB14h, BBLZ14b, BTZ98, LY21, PHBH13]. dating [BB12d]. David [Hoa05, Sha05, Zei05, Bor03-32, Bor04n, BE16]. Day [BB13y, BB15-28, BB16k, Bor07v, Bor08l, Bor08m, Bor10u, Bor11x, Bor11y, Bor11z, Bor12u, Bor12v, Bor12w, Bor13s, Bor14u, Bor16n, Bor16c, BB14v, Bai17e, BB14c, BB14w, BB15t, BB16j, BB14u]. Days [Bor11d, Bor16n, Bor11h]. DC [Coh15, AMM10]. Deafening [Sol15]. death [BB11c]. December [Bea13, BBJC97, BBL^{+}13]. Decimal [Bai88, Gan14, BB11e, Bai16b, BBGPxx]. decision [Bea13]. decisions [BBLZ13b, BB13-39]. decline [BB11k, BB11y]. Decomposition [BL92b, Bor04o, CZX21, BW07]. Decompositions [Bor06t, BWY10, Bor79e, Bor07b]. decreasing [BL93a]. Dedekind [BG96a, BG97b, BG97a, BB98c, BB98d]. Dedekind-zeta [BB98c, BB98d]. Default [BBL^{+}13, SBB13, BBS13a, BB13-35, BB13-36, Bor13-30, Bor15m]. degree [Ade11, Ade13]. déjà [Tre13]. della [BB95d]. delta [BG15b, BG16b]. demand [JY12]. demonstrate [BBLZ14k]. demonstration [BBS^{+}15a]. Demyanov [DP18]. Denial [BB13-44, BB10c, BB12-28, BB13-43]. Denominators [BZ87]. dense [BB96a, BB99c, BBWY11c, BBWY11e, BBWY12c, BY12f]. Densities [BSWZ11, BSWZ12, BSV15, BSV16, Bor14t]. Density [Hon85, BS16b]. Department [Bor03i]. Derivative [BLLN94, Bor94g, BLN94a, BLN95, Bor95m, Bor95n, BLLN95, BLN96]. Derivatives [BB18, BFV93a, BD16a, BD18, AL10, BB16a, BFV93b, BF93c, Bor94m, BMW95, BF95b, Bor95v, BMW97]. Deriving [BB14-32]. Descent [Bor09c, SD15, RS02]. design [BBL16a, BBL16c]. designs [AX20]. Desperately [BB15j, BB15k]. detected [BB16f]. Determination [BBB06a, Las18, BBB05, BM97f, BM00, BT14b, BT14a, BT17]. Determinations [BB98c, BB98d]. determined [BB97c, BB05f]. developed [BB11q]. Developments [BB99a, BB01a]. devices [Bor00w]. dian [BB95e]. Dictionary [Bai91, BB88b, BB99d, BB99e, BS14a, Lor90, BB91a, BB02, BB90d, BWB97, BS14b]. Did [BB13j, BB09f, BB12g]. didn’t [BWB97]. Diego [BC96]. dies [Ano16, BB12-38, Bai16c, Bai20]. Diewert [Bor90b]. Difference [Bor11q, BB11a]. different [PHBH13, Zha13]. Differentiability [BBS10, Bor90g, Bor90h, Bor90i, Bor90k, Bor90j, Bor90l, Bor91d, Bor92a, Bor02d, Bor02e, BBL04, BV09, Bor76a, Bor82a, Bor86e, Bor86b, BFG87, BP87, Bor90a, Bor90-40, Bor90-41, Bor90-42, Bor90-43, Bor91r, Bor91s, Bor91t, Bor91u, Bor92g, Bor92h, BF93a, Bor93f, Bor93g, BF93b, BN94, BW05a, BMV06]. Differentiable [BW95a, BM95, Bor95d, BW97a, BFKL00, BFKL01]. Differential [BM97b, LY18, MR96]. Digit [Bor05g, Ade10, BB12-29, BBG04b, Bor11i]. digit-extraction [Bor11i]. Digital [Bor02f, BS03, Bor03-35, Bor05i, BRR08, BB09k, BB11z, Ban10, BM06, Bor06-36]. Digitally [BBB^{+}96a, Bor08g, Bor09a, Bor09e, Bor09f, Bor09g, Bor12-33, Bor09u]. Digitally-assisted [Bor08g, Bor09a, Bor09e, Bor09f, Bor09g, Bor12-33, Bor09u]. Digitized [BB05e]. Digitizing [Bor02g]. Digits [Bai88, BBMW11, BBC^{+}12c, BB13c, BB14i, BBBR16, BBBR17, BBB97c, BBB00b, BBB04b, Bor09z, BBB16, Gan17, BBB97a, BBMW13, BB13b, BB14h, BBLZ14b, BB14x, Bai16b, BBMW16, BB16p, BBMW17, BBB89, BBxxc]. Dilemmas [GS08]. dilogarithmic [Cvi10]. Dimension [BB09o, CKM^{+}16, Via16]. Dimensional [ANR18, BCC10, HDL21, AAW06, BW81c, BW86, Bor88f, Bor91h, Bor92d, Bor92n, Bor92j, BF94a, Bor94b, Bor94g, BF95a, Bor95m, Bor95n, Bor97f, Bor97m, DS20]. Dimensions [BB86a, WB87, BB16o, Bej94, BL91d, BFL02, Bor14t, BSV15, Bor15o, Bor15p, Bor15q, BS16b, BSV16]. dinosaurs [BB13j]. Diophantine [Kom00, Kom02, Kom04]. Dirac [BH94a, BH94b]. Direct [Bor79a, Bor81c, LLC^{+}95, FN15]. direction [ZL22]. Directionally [BS84a]. directions [BBLZ14d]. Directory [BMP05]. Dirichlet [BB15c, Bor01g, Bor02h, Bor02i, BC02, BC03, BC04b, Bor07e]. Dirk [Bor05g]. disciplinary [Bor16h]. Discourse [BS03]. Discover [BB09j]. discovered [BB11r, BB11-29, Bor95c, Bor97q, Bor97v, Bor97w]. Discovering [Bor91e, Bor91f, Bor91g, Bor91j, Bor91k, Bor91i, Bor91l, Bor91m, Bor92e, Bor92f]. Discovery [BB11l, BB11m, BBKL16, BBKL17, Bor02l, Bor02m, Bor03j, Bor03l, BBG03, Bor04p, Bor05l, Bor05m, Bor05n, Bor05o, Bor05p, Bor05a, Bor06l, Bor07g, Bor07l, Bor07k, Bor08g, Bor12-33, Bor16f, R^{+}05, Ade12, BB08c, BB10f, BBG04a, Bor06h, Bor08d, Bor08e, Bor08f, Bor09a, Bor09e, Bor09f, Bor09g, Bor09u, Zei05]. Discussion [BBGP95a]. disease [BB13-28]. distance [BB00a, BB01b, BF94b]. Distances [BML18]. Distinct [BW95a, BW97a, BBT98, BBT00]. Distributed [Bor99b, Bor99c, LY21]. Distribution [TB00, BG94a, BG94b]. distributions [BCM02, BCM03]. Ditor [BO11a, GN16, Mil90, Mil89, MW12]. divergence [Lor08]. Dizionario [BB95d]. DNA [BB12-50]. Do [BBLZ13d, BB13s, BBLZ14a, BB14c, BBLZ15b, BB15s, BBSL20, BB12-30, BB12-51, BB13r, BB13y, BB14w, BBLZ15d, Bor94n, BB15w]. Doctor [BB12b]. dodgy [BB12a]. Does [BB12v, BB12w, BB15l, BB15m, BB11-28, BB12-50, BB13o, BBLZ16b, BBLZ16a]. doesn’t [Bor07q, Bor07p]. Doing [Bor96b, Bor97a, Bor97b, Bor97c, Bor97e, Bor97d, Bor99e, Bor99f, Bor99d, BS99c, Bor00n, Bor98r, BS99a, BB11g]. dollar [BB14e]. domain [BY12d, BY14b]. done [BB12-45]. Don’t [Bor13c, BB11f]. Dontchev [Bor11-38]. double [BB12h, BZB08, Mer15]. Doubly [BLN94b]. Doubt [BB11t]. Douglas [AB12, ABT13a, AB13, ABT13b, ABT13c, ABT14a, ABT14b, ABT15, ABT16, AC18, BS10b, BS10c, BS10d, Bor10i, Bor10j, BS11b, Bor11r, Bor11s, BT13a, BT13b, Bor13j, Bor13r, BT14c, Bor14f, Bor14g, BT15, Bor15g, Bor15r, BG16a, BLS^{+}16, BLS^{+}17, BLS^{+}18, BG18b, Gil18]. dragons [BB13j]. Dreams [Bor02p]. drive [Bor13c, Bor05-46]. Driven [BC18b]. drug [BB13h]. Dual [BV93a, BV94c, BTBT88, BMN98, BMN00, BS10a]. Duality [BMCL18, BL91b, BF98, BF01, Bor09-28, BC10, BL15, DGLV20, Roc20, Ara07, Ara08, Bor80d, Bor80e, Bor83a, BK83, Bor83f, Bor86a, BL91d, BL92c, BLN94a, Bor94o, BLN96, BBY12, BBY14, Zăl86]. Dubious [BB14h, BBLZ14b, BB14i, BV24]. due [Koh01]. dull [Bor11b, Bor11c]. Dumb [BB13-38, BB13-37]. dummies [BB12-46]. dunce [BB13-39]. during [BBLZ14i, SBW84]. Dykstra [BB94a]. Dynamic [Bor02r, JJ20, KMZ^{+}03, BNCB99, BS95, BS97a, LLC^{+}95]. Dynamical [BC18b]. Dynamics [Bor04-30, Bor04-29, Bor04-28, BK05, BL05, Bor05j, Bor06i, BL08, BLS^{+}17, BLS^{+}18, BBB12, BBCM07b, Bor10q].

E. [BB13-45, BB13-46]. E2995 [ANO^{+}83, AJ86]. E2996 [NOL86]. E2997 [AJB86]. E2998 [KJ86]. E2999 [SZUM86]. E3000 [ANO^{+}83, EWM86]. E3159 [DNG^{+}86, DBCB88]. E3160 [NJS88]. E3161 [GC88]. E3162 [Mon89]. E3163 [KC89]. E3164 [DAK88, DNG^{+}86]. E3325 [Rud89]. E3335 [KWK^{+}90a, KWK^{+}90b, KWK^{+}90c]. E3384 [Stu90]. E3388 [CJKB92]. earlier [BB11-29]. early [BBLZ14r]. Earth [Bor13t, BB12-31, BB12-51, BB12-52, BB15q, BB12d, BB12g]. East [Bor05k]. Easy [Gui08]. Eberhard [Bor06o]. ecological [Bea13]. Economics [SZ81, BB13-31, BBLZ15e]. economies [BB12-41]. economists [BB09f]. economy [BB12p, BB12q, BB12-48]. Edelstein [BGMS21]. edge [BBLZ15a]. Edited [Bor06o, Coh15]. Edition [BBB03, Bou06]. editor [Zăl86, Bor11b, CW16, Cha16]. Editors [BM97a]. Education [AD20, Goo20, Hol20, JB21, lL09, PL20, BB10j, BB13w, Hd12]. Educational [BaHO20]. Educators [Goo20]. Effective [BB06a, BB08d, Bor06j, Bor06k, Bor07h, BBC07c, Bor07i, Bor07j, Bor08h, BBC08b]. effects [BBLZ14s]. efficiency [Bor80a, BZ91, BZ93, JN03, Zhu91]. Efficient [BCJW13, Bor77c, BJCW13, Bor83e, HLZ15a, Yan94, Zho12]. eigenvalue [AR13, GDT15, JD13]. eigenvalues [Bor84c]. Einführung [BD11]. Einstein [BB09d, BB16f]. Eisenstein [Liu01, XY12]. Ekeland [Bej94, Bor88g, Bor88h, Bor88i, Bor90m, Bor90n, LS00, YS00]. elastic [HYG09]. election [BBLZ14i]. Electron [BBSZ87, BBSZ88]. Electronic [Bor01n, Bor01m, Bor02n, Bor03p, BS97b, Bor97o]. Elementary [AJB86, ANO^{+}83, AJ86, BB84a, BB97b, BB00b, BB04a, CJKB92, DAK88, DNG^{+}86, DBCB88, EWM86, GC88, KJ86, KC89, KWK^{+}90a, KWK^{+}90b, KWK^{+}90c, Mon89, NJS88, NOL86, Rud89, SZUM86, Stu90, BB16t]. Ellipses [BLS^{+}17, BLS^{+}18, BLS^{+}16]. Elliptic [BBBG08, BB84c, BZ87, Borxx, Bor10x, Bor11-30, BBG95b, BZ92, BBGW11, LL01, PT14]. else [BWB97]. Emerging [BC98b, BC99]. Emmy [BB12x]. Empirical [Alt20, BBC^{+}11a, BBC^{+}12b, Bor97h]. Empirically [BB97c, BB05f]. Encourage [BB15l, BB15w]. Encyclopaedia [Sel16]. Encyclopedia [BC96]. End [Bor09d, BBLZ13d, Bor03x, Bor03y]. Energy [BB14l, BB14q, BB15g, BB16i, BBSZ87, BBSZ88, BB12-27, BB12e, BB14k, BB14p, BB15f, BB15u]. engaged [BB16g, BB16h]. Engineering [PL20, BBC^{+}11b]. engineers [BWB97]. Engines [Bor04p, Bor05l, Bor05m, Bor05n, Bor05o, Bor05p, Bor06l, R^{+}05]. enhancement [BM07a]. ENIAC [Bor12o, Bor14p, Bor14s, Bor16o]. Enlargements [BBY11, BBY13]. enough [BB14-28]. entire [Bor02g, BS10a]. Entropic [BL11]. entropies [BGL93, BH94a, BH94b]. Entropy [BL93c, BLL94, BLN94b, BLLN94, Bor97l, Bor01o, Bor05-33, Bor06-33, Bor08p, Bor09v, Bor10k, Bor10v, Bor10w, Bor12q, Bor13k, Bor13l, Bor13p, BHP14, BZ20b, Bor90c, Bor90d, BL91a, Bor91b, Bor91c, BL91b, Bor91h, Bor92d, Bor92n, Bor92j, Bor93e, Bor93k, BL93b, Bor94g, BLN94a, BH95, Bor95m, Bor95n, BLLN95, BLN96, BCM03, Bor12p, BH19, PHB12, PHBH12, PHB14, BCM02]. entropy-like [BL91b]. Entropy-Type [Bor01o]. Entry [BS16a]. Environment [IEE08]. Environments [Bor04e, Bor04d, Bor04i, Bor06d, Bor06b, Bor06c, Bor06g]. Epi [Bor87m, BLM99, BLM00]. Epi-Lipschitz-like [Bor87m]. epi-Lipschitzian [BLM99, BLM00]. Epigraphical [BV95b, BV96c]. episode [BBLZ15g]. equality [Bor77a]. equation [BB13g, BBCZ13, NFB17b]. Equations [AI18, BM97b, BBB97c, BBB00b, BBB04b, BBB16, LY18, BBB97a, Bor86f, Bor87g, Bor87f, BB89a, BBB89, Bor93k, BG94a, BG94b, DLL05, MR96]. Equilibrium [AI18, CG18]. equipped [BBLZ13b]. Equipping [HDL21]. Equivalence [BMS99b, Zho12, Bor77b, Yan94]. Era [BRR08, BB12-27, BB12e, Ban10]. Erdélyi [FK00]. Erdos [BB13v, Cra12, Mer15]. Ergodic [BG16a, BG18b]. Error [BB08d, Bor06j, Bor06k, BB13l, BB13-31, BB13-32, BBL97c, BBL99, CGM95]. error-correcting [CGM95]. Especially [Bor94f]. essays [BR01]. Essential [BBC00a, BBC01, BB12p, BB12q, BB12-48, BWB97]. Essentially [BM95, Bor95d, BM97c, BM97d, BM96a, BM96b, BM98a, BM98b]. estimates [BL91a, BL93b]. Estimation [Bor91h, Bor92d, BTBT88, Bor90e, Bor90f, Bor90c, Bor90d, Bor91b, Bor91c, BBT92]. eta [BG96a, BG97b, BG97a]. ETFs [BBLZ14a]. Euclidean [Bor84b, La 09]. Euler [BBG93a, BBG94a, BB06a, BB08d, BBD89, Bor89f, Bor90-29, Bor90-30, Bor90-31, Bor90-32, Bor90-33, Bor90-34, Bor90-35, Bor90-36, Bor90-37, Bor90-38, Bor90-39, Bor91j, Bor91k, Bor91i, Bor91l, Bor91m, Bor92e, Bor92f, Bor92i, BBG94b, BBG95c, Bor95e, BG95b, BBB96b, BBB96c, BG96b, Bor96f, Bor96g, Bor96h, BBB97d, Bor97f, BBD97, Bor98f, BBD00, BBD04, BB05g, Bor06j, Bor06k, BBB06b, BZB08, BCM09, BBD16, Dil20]. Euler/Zagier [BBB96b, BBB96c, BBB97d, Bor97f]. Eulerian [BBB15]. Evaluation [BBSL17a, BBSL17b, BB18, BBSL18, BZ87, BG95b, BG96b, Bor97f, BD16a, BBG93a, BBG94a, BB16a, BZ92, BBG94b, BBG95c, Bor95e, BBC08a, BZB08, BD18]. Evaluations [BBB96b, BBB96c, BBB97d, BG05, BBBG08, BS11a, BBSW11, BS12b, BBSW12]. even [BKW02]. ever [Mic03]. Every [BBWY11c, BBWY12c]. everybody [BB11n, BB11d, BB11-30, BB11g]. Everything [SW21, BB09g]. everywhere [Bor12l]. evolution [BB12w, BB12-50, BB13h]. Ewing [Jac09]. Exact [Bor99g, Bor99h, Bor00b, Bor00c, Bor00d, Bor00e, Bor00f, Bor00g, Bor00h, Bor00i, Bor01i, Bor01j, Bor01k]. example [Bor92c, BD11]. example-oriented [BD11]. Examples [BFV94b, BFV94c, BFV94a, Bor17b, Com18, DLR20, BB05b, Bor87m, Bor93p, Bor94l, Bor95s, BZ98, Bou06, Tod03]. Excel [BB13-31]. excluding [BBG04b]. Excursion [Bor87d, Bor88a, Bor88b, Bor88c, Bor88d, Bor88e]. Exercise [BB12r, BB16d]. Existence [BF89b, CG18, Bor82d, Bor83e, Bor84c, Bor88k, BL93b]. exp [BBC08a]. Expansion [Gan14, BB83]. Expansions [BBD97, BBD00, BBD04, BB07c, BBCP04, BBD89, BG95a, BBGPxx, BBD16]. expansive [BS10a]. Expectations [BBCR13, Bor12g, Bor12h, BR16]. Experience [Bor07d]. experiences [Bor08q, Bor12t]. experiencing [KMT16]. Experiment [BBG03, Bor03z, Bor03-27, Bor03-28, Bor03-29, Bor04u, BB04b, Bor05-31, Bor05-29, Bor05-30, BB08h, Bor10a, HF05, Zei05, Hoa05, Sha05]. Experimental [BBG93a, BBG94a, BB99a, BB01a, BBB05, BB05b, BBB06a, BBC^{+}07b, BB09e, BB09a, BB10d, BBBZ10b, BBBZ10a, BBL^{+}13, BB14a, BB15n, BB15o, BB16c, Bai17b, Bai17c, BB18, Bor94c, Bor94d, Bor94e, Bor94p, Bor94r, Bor94q, Bor95e, BBGP95a, Bor95f, Bor95g, Bor95h, Bor95i, Bor95j, Bor95k, Bor95l, Bor95w, Bor96c, Bor99g, Bor99h, Bor99i, Bor99j, Bor99k, Bor99l, BBGPxx, Bor00b, Bor00c, Bor00d, Bor00e, Bor00f, Bor00g, Bor00h, Bor00i, Bor00j, Bor00k, Bor01h, Bor01i, Bor01j, Bor01k, Bor02j, Bor02k, Bor02a, Bor02l, Bor02m, Bor04t, Bor05a, Bor05q, Bor05r, Bor05-38, Bor05-39, Bor05-41, Bor06m, Bor06n, Bor07l, Bor07k, Bor07m, Bor07n, Bor07r, Bor07s, Bor07t, Bor07u, BB08g, BB10l, BaO12, Bor14h, Bor16g, Bor16f, CDH^{+}21, CC20b, JB21, dPB21, AMM10, BBKW06, BB14t]. experimental [BB16a, Bor93p, Bor93q, BBGP95c, BBGP96, BC98b, BC99, Bor08c, Bor08b, BD09, BD11, Lor09, Odl11]. experimentally [ABBS12, Bor93j, BB11-31]. Experimentation [BB12y, Bor92i, BBGP95b, Bor03k, Bor03l, Bor03m, Bor03n, BBG04a, Bor04q, Bor04r, Bor04s, Bor09h, Bor09i, Bor10l, Bor10m, Bor11t, Bor12-33, Bor12i, Bor13m, Bor13n, BB11j, BB12z, Bor09u, Sha05, Zei05]. Experimentelle [BD11]. experimenting [KMT16]. Experiments [BBG03, BBG06, CS21]. Explainer [BR12, BR13b, BR14a, Tre13]. Explicit [BB06b, BB84d, BB87a, BL92d, BBG94b, BBG95c, BB86b, BS10a]. Exploration [BB12y, BB16n, BB16m]. Exploratory [BB11j, BB12z, Bor09h, Bor09i, Bor09u, Bor10l, Bor10m, Bor11t, Bor12-33, Bor12i, Bor13m, Bor13n, Bor14i, Bor14j, Bor14k, Bor14l, Bor14m, Bor14n, Bor15h]. Exploring [Bor01l, KMT16]. Exponential [BB94b, BG03b]. exposes [BBLZ15g]. exposing [Bor78b]. Expressions [BSW82, BBK14]. Extended [NWY09, NWY10, BBC14b]. Extending [SV20]. Extension [Cos17, La 09, Bor82e, DABY15, Mil90]. Extensions [Bor10x, Bor11-30, Bor88g, Bor88h, Bor88i, Bor94b, BGV02, BMV06, BBGW11]. extraction [Ade10, Bor11i]. Extraordinary [Bai16d]. extraterrestrial [BB11g]. Extremal [PR92]. Extreme [Bor06m, Bor06n, GDT15, JD13].

F [Ban10]. FAA [BBS17]. FAAAS [BBS17]. Face [Cal16, Bor96k, Bor97x]. Facial [BW81d]. Facilitating [BBS16b]. facilities [JY12]. fact [BB12f, BB15m]. factor [BBLZ14h]. Factorial [BC18a]. factorization [HNP10, HLZ15a, HLZ15b, LL13]. facts [BB11k]. fail [BW98a]. failing [BB12k]. failure [Bor92n, Bor93k]. fall [PD18]. fall-back [PD18]. fallacy [BBLZ14n]. False [dPB21]. Familiar [BB88e, BBxxa]. family [BB15e, Bor79c, Bor80e]. famine [BB12-27]. FAMS [BBS17]. Fan [BZ86]. FAQs [BBLZ14c]. far [BB11n, BB11d]. Fared [BB15r, BBLZ14f]. Farkas [Bor79d, Bor83d]. Fast [BB84a, BZ92, BLN95, BB97b, BB00b, BB04a, BD16a, BH95, BB16t, BD18]. FAustMS [BBS17]. Favourite [Bor07-28, Bor07-29, Bor07-30, Bor08u]. FBAS [BBS17]. Feasibility [ABT13a, ABT14a, ABT15, BB95c, BB96b, BT13b, Bor16u, Bor16v, Bor16w, Bor16x, Bor16y, ABT16, Bor12p, BT15, Bor15r]. Feasible [JD13, LLS11, ZL22]. Feast [BB12-27]. Featured [Bor06o]. February [ABD03, Bai17e]. Federated [BMP05]. Fedspeak [BBLZ14d]. Fee [Rei02]. feeling [BB09d]. Fenchel [BK83, BL91d, BH06, BH09]. Fenchel-duality [BK83]. Fermi [BB10e, BB15j, BB15k, BB15-29, BB15-30, BH94a, BH94b]. few [BB12b]. Feynman [BB98c, BB98d]. Fiasco [BB15x, BB13-29, BB13-30, BBLZ15f]. Fibonacci [Ade14a]. fiction [BB12f]. fiddling [BB09k, BB11z]. field [Cvi10]. Fields [Bor02p, BSZ13, Bor03o, Bor14b]. Fifty [Bor09j, Bor09k, Bor10n, Bor12j, Bor12k]. filter [AP16, ZSQ10]. Final [Bor06p, Bor09-27]. Finance [Ano15, BZ20b, BBLZ13g, BBLZ15e, BBLZ14p]. Financial [BBS^{+}16a, BBSL20, dPB21, BBLZ13b, BBLZ14l, BBLZ14s, BBLZ14q, BBLZ16d, Cam16]. Financially [BB14n, BBLZ14e]. Finding [BBG95a, BB06b, BBG04b, Bor07o, BB15d]. fine [BB14-29]. fine-tuned [BB14-29]. Finite [WB87, Bej94, BW81c, Bor88l, Bor89i, BL92c, BL92d, BL93b, La 09]. firmly [BRS11]. First [BC18b, Bor92g, Bor92h, Bor93f, Bor93g, Bor06q, BB12-41, BZ92]. first-world [BB12-41]. Fisher [BLN94a, BLN96]. Fitting [BBLZ13a]. Fitzpatrick [BBB^{+}07, BBW07, BBWY11c, BBWY11e, BBWY12c, Bor06a, Bor14o, Bor15i, BD15, BD16b]. Five [Bor07d, Bor15d]. Fixed [BBC^{+}11b, BB91b, BLT17, Bor84a, Bor92k, Bor92l, Bor92m, BLT15, BLT16]. Fixed-point [BBC^{+}11b, Bor84a]. Flash [BB15x, BBLZ15f, BBLZ14m]. fold [BBB96b, BBB96c, BBB97d]. Folkmar [Bor05g]. foolishness [BB13l]. Forecasters [BBSL17b, BBSL18, Swe17, BBSL17a]. Forecasts [BBSL20, BBLZ14q]. Forensic [BB12r, BB16d]. forever [BB11u, BB12-34, BB12-35, BB13-47]. form [BS16b]. Formal [Ade13, BB14j]. former [Ano16]. Forms [BBBC07, HMM20, Bor10f, Bor10g, BC13, La 09]. Formula [AW97, Ade14b, BBLZ13f, BG87, Borxx, Bor16b]. Formulae [BB96c, Bor99w, AG99, BB97c, BBG04b, BB05f, BB05c, Cha03, ZS12, Zha13, ZZ14]. Formulas [Ade14a, BB06b, AL10, Ade10, Ade11, Ade12, Ade13, ABBS12, BB11-31, GG07, Nim15, Wei15]. forthcoming [Cam16]. Foster [BSW82]. Fostering [Goo20]. found [BB16p]. Foundation [RZ15]. Four [Bor02c, Bor02q, Bor06r, Bor06s, Bor06t, Bor06u, BSW13, BB13e, Bor88f, BB13f]. four-color [BB13e, BB13f]. four-dimensional [Bor88f]. Four-Step [BSW13]. FPV [BEY11, BY13a, BY14c]. frack [BB14z, BB14-27]. Fractal [Bor10o, Bor11u, BBCR13, Bor12g, Bor12h]. Fractals [Bor12l, BR10]. Fraction [Bor03d, Bor03e, Bor03f, BCF04, BC04a, BBGPxx, BL05, BL08, Bor10z, Bor10-27, Bor11-32, Lor08]. Fractional [Bor76a]. Fractions [Bor04-30, Bor04-29, Bor04-28, BvdPSZ14, BCLM16, BHL16b, BHL16a, BCLM17, BZ92, BCP05, Bor05j, Bor06i, BHL17]. frame [FN15]. frame-based [FN15]. Framework [Roc20]. franc [BBLZ15g]. France [CGM95]. Frank [BB13k]. Frankowska [Bor92b]. Fraser [BBJC97, Bor89a]. Fraud [BB13l, BB90c, BB92a, BB11s, BB11f, BB13-33]. Fréchet [BV10a, BF93a]. Fredholm [Bor92n, Bor93k]. French [Dev9x]. frequency [BBLZ14o]. Friedman [BB13d]. Fritz [Bor76b]. FRSNSW [BBS17]. Function [BZ87, BB96c, BBC98, BBC00b, Bor03-33, Bor04-31, BK05, Bor08k, BBL10, BL11, BD16a, Hir17, AL10, AB15, BB15c, BS17, Bor91n, BZ92, BB93e, BLN95, BG96a, BG97b, BG97a, BKW02, BB05c, BC09, BS10a, Bor14o, Bor15i, BR16, Bor16l, Bor16m, BD18, HGB93, Liu00, NWY09, SZ14]. Functional [Bor72, BG94a, BG94b, Bor98k, BZ99a, BZ99b, LLC^{+}95]. Functionals [BB93b, Bor78b, BK01]. Functions [BB84a, BB88e, BFV93a, BV94b, BFV94b, BW95a, BB97b, BV9x, BBxxa, BB00b, Bor02b, BB04a, Bor07g, Bor07h, Bor07k, Bor08h, Bor08u, Bor09m, Bor11q, Bor11-29, BV12, BD15, BML18, Com18, Dil21, EB08, GI01, Las18, LPB01, SBW84, AB15, AAW06, BBS10, BBEM10, BB11a, BBB15, BBB^{+}07, BB95b, BB97a, BBC00a, BBC01, BBW07, BBWY11d, BBWY13, BBP03, BBG95b, BFG87, BP87, Bor90g, Bor90h, Bor90i, Bor90k, Bor90j, Bor90l, Bor90y, Bor90z, Bor90-27, Bor90-28, Bor90a, Bor90-40, Bor90-41, Bor90-42, Bor90-43, BB91b, Bor91a, Bor91r, Bor91s, Bor91t, Bor91u, BL92b, Bor92g, Bor92h, BF93a, Bor93e, Bor93f, Bor93g, BFV93b, BFV94c, BFV94a, BG94a, BG94b, BF94b, Bor94l, BN94, BV95a, BM95, BV95b, BMW95, Bor95d, Bor95s, BM96a, BV96c, BM96b, BFV97]. functions [BV97, BW97a, BM97c, BM97d, BM97e, BMW97, BM98a, BF98, BW98b, BM98b, Bor98o, BLM99, BRLZ99, BLZ99, BW99, BV00a, BLM00, BV00b, BRLZ00, BW00, BV01, BLZ01, BF01, BVW01, BW01, BV02, Bor02d, Bor02e, BGV02, BW03, BVW03, BBL04, BW05a, BW05b, BMV06, Bor06h, BBC08a, BV09, BG09, BGHV09, BV10b, BV10a, Bor11-38, BY12a, Bor12t, BY14a, BG15b, BB16t, BDT16, BS16b, BL16, BD16b, BG16b, How14, HL15a, LL01, Liu01, Lup02, SZ14, XY12]. fund [BBLZ14h]. Fundamental [BB05g, Bor13d, Bor13a]. Funding [Bor07o, BB10i, BB13u, BB13-39]. funds [BBLZ15a, BBLZ16c]. Further [BV93b, BV94d, Mil90]. Fusion [BB14k, BB14l, BB15g, BB15f]. Future [BB99a, BB05a, BB16c, BB16n, Bor05s, Bor07a, Bor08i, Bor08j, Bor10p, Bor12m, Bor15j, BB01a, BB12-39, BB12h, BB16m, BD95, Bor95t, Bor95u, Bor98c, Bor09l, Cam16].

Gâteaux [BF93a, BF93b]. game [BB12d, BB15b, BB15o]. games [BB12s, BB12m]. Gamma [BZ87, BK05, Bor12r, BC18a, BBB15, BZ92, BC09, BB15c]. gap [BB14s, BBY12, BBY14, Bor14o, Bor15i]. Garvan [Hir17]. gas [BB12-27, BB12e]. Gateway [Bor04j, Bor04k]. Gauss [Bor87d, Bor88a, Bor88b, Bor88c, Bor88d, Bor88e, Borxx, Cos17, TK97]. Gaussian [Cha03]. Gems [AMM10]. General [BB06b, AB15, BBWY11a, BBWY12a, Bor85c, BV00b, BV01, Bor07x]. Generalisation [BLS^{+}16]. Generalisations [Bor17b]. Generalization [Mil89, YS00, AB15, Bor97g, Bor98g, LS00]. Generalizations [TB80]. Generalized [Bor84a, Bor99m, Bor99n, BMW99a, BMW99b, BMW99c, Bor00l, Bor00m, BMW01, Bor10x, Bor11-30, BS11d, BS11e, BHL16b, BHL16a, BHL17, LPB01, RP09, SZ81, BFG87, Bor94b, BBGW11, Cha03, War01, War03, ZSZ16, Bor90b]. generated [SZ14]. Generating [Bor07g, Bor07k, Bor91n, BB93e, Bor06h, PHBH12]. Generation [PHBH13, BB16l, BJCW13]. generator [BCJW13]. generators [BB13x]. Generic [Bor86e, Bor99m, Bor99n, Bor00l, Bor00m, Bor86b, BF93b, BW98b, BW00, BK01]. generically [BW98a]. genius [Bor91p, Bor91q, BB91d, Bor11e]. geologic [BB10g]. Geometric [BB84a, BLM96, BB97b, BLM97, BB00b, BB04a, Bor87d, Bor88a, Bor88b, Bor88c, Bor88d, Bor88e, Bor88f, Bor89e, BBG93b, BB16t, IP17, IP18]. Geometry [Bor09-27, Bor11u, Bor80a]. German [BB96d, BD11]. get [BB09f, BB14-28]. Getting [BB13m]. Girgensohn [Odl11, Sha05, Zei05, Rei02, SZ14]. Giuga [BBBG94, BBBG95, BW95b, BBBG96, BW97b, BMS13, BSM13]. give [BB14e]. Glenn [BE16]. Global [AB12, AB13, ABT15, ABT16, BB12-28, BB10c, BB12c, NFB17a]. globalization [GS02]. Glum [BB13n]. glummer [BB13n]. go [Bor15a]. goals [BB10h, Bor13c]. God [BB12-32, BB12-33]. goes [BB11u, Bor05k]. Going [Bor12x]. Goldbach [Bor05c, BB05d, BB06c, Bor10b, Bor10-31]. Golden [Ade14a]. Good [Bor00j, Bor00k]. googol [Cra12]. googol-th [Cra12]. Got [Bor15t]. Gowers [Bor09b]. Gradient [BB88a, CZX21, SI16, SD15, BFKL00, BFKL01, BFL02, DL02, DLL05, DK16, GS02, Li15, LL13, Mar91, MP18, NFB17a, NFB17b, QYX14, Ray93, Ray97, WSdSY15, XH08, XSW12, XWQ14, YW12]. Gradients [Bor99m, Bor99n, Bor00l, Bor00m]. Grading [Swe17]. Graph [AC18]. graphics [BJCW13]. Graphs [BB93b, Ber88, BFG03]. Graves [BD03]. Gravitational [BB14m, BB16f]. great [BB11k, Bor13d, Bor13a]. Greatest [BB11l, BB11m, BB10f]. greco [Bor08a]. Greek [BS14b, BS14a, Bor90o, Bor90p, Bor94f, Bor08a, SV14]. Green [Bor09b, BB12-27, BB12e]. Grid [Bor03b, Bor03c, Bor03a, Bor04e, Bor04a, Bor04b, Bor04c, Bor04i, Bor05-28, Bor07d]. ground [BB12-30]. Groups [Bor16j, Bor16k, BG16c, BG15a, Bor15f, BG15c, Bor16i, BG18a]. Grove [Bai91]. guarantee [Cam16]. Guessing [Sei01]. Guide [Bor02j, Bor02k, Bor06o]. Guided [Bor92i]. Gun [BB15l, BB15w]. guru [BBLZ14q]. Gurus [BBSL20].

H [Bor92b, Hoa05, Odl11]. H. [MR11]. Haar [BF95c, Bor95a, Bor95b]. Hadamard [BF93d]. Hahn [Bor82e]. Haifa [IMR92, RZ15]. Half [WSL16]. Halloween [BBLZ13e]. Hand [BB12-29]. Hand-to-hand [BB12-29]. Handbook [Sch15]. handheld [Bor00w]. Handling [Bor03p]. happen [BB13-33]. Hard [Bor01e, Bor02s, Bor02t, BB14d, BBL^{+}16b, XC11]. hard-wired [BB14d]. Hardback [Ban10]. hardcover [BC96, Bor09b]. HarperCollins [BB91a]. hating [BO11b]. Hausdorff [BK80]. having [BF93a]. Hawking [BB10e]. headlines [BB12a]. heart [BB12-30]. Heats [BB15g, BB15f]. hedge [BBLZ14h, BBLZ15a, BBLZ16c]. Heisenberg [BBCM07a]. held [IMR92, SBW84]. Helen [Coh15]. Helly [Bor77b, Bor79b, Bor81e]. Here [Bor05s]. Hermitian [Bor84c]. Hersh [BO11b]. Hessian [BC18b]. heuristic [BH95, BLN95, JY12]. Heyting [Bor98d]. Hidden [BC18b]. Hide [BB13s, BB13r]. Higgs [BB13o, BB13p]. High [BB08a, BB08e, BB08b, BB11b, BBB12, BB13q, BBLZ15c, BB15p, BB90c, BL92e, BB92a, Bor98h, Bor05t, Bor05u, Bor05v, Bor05w, Bor05-47, Bor05-48, Bor05-49, Bor05-50, Bor05-51, Bor05-52, Bor06z, Bor06v, Bor06w, Bor06x, Bor06y, Bor06-37, Bor06-38, Bor06-39, Bor07f, BB09o, Bor10q, IEE08, BB09b, BBLZ14o, BB87a, BWB97, Bor03x, Bor03y, Bor05g, DS20]. high-accuracy [Bor05g]. high-dimensional [DS20]. high-end [Bor03x, Bor03y]. high-frequency [BBLZ14o]. High-Performance [IEE08]. High-Precision [BB08a, BB08e, BB08b, BB13q, BB11b, BBB12, BB15p, DS20]. Highend [Bor03h]. Higher [BCC10, AL10, BB84b, BSV15, BSV16]. Higher-Dimensional [BCC10]. Highly [BB08e, BB09b, Mic03]. Hilbert [BBEM10, BBL94, BBL97b, Bor05x, Bor08k, Bor09m, Bor10c, Bor10d]. History [Bor77d, BJL^{+}08, Bor11x, Bor11y, Bor11z, BC15a, Bor16n, Sel16, BB16l, Bor90q, Bor90r, Bor90s, Bor90t, Bor90u, Bor90v, Bor90w, Bor90x, Bor93h, Bor93i, BC16]. hits [BB13m]. Hölder [BLT15, BGW97, BGW98, BW03, BLT16]. holdout [BB13t]. Holistic [Alt20]. Homo [Thé16]. Homotopy [BO11a]. Honor [SV14, Ano15, BBB^{+}13]. Honoring [PR92]. Honoris [Bor99o]. honour [Bor17b]. Honours [BZ11]. Hope [BB14k, BB14l]. hoping [Bor01f]. Hopkins [BS14a]. hosted [IEE08]. Hot [BB12c]. HPC [Bor04p, Bor05l, Bor05m, Bor05n, Bor05o, Bor05p, Bor06l, R^{+}05]. HPCS [IEE08, IEE08]. HPCS06 [BB06a]. hull [BBL97c, BBL99]. Human [Cal16, WG17, BB15e]. humans [BB13j, BB16e, BB16s]. hundred [BBxxc]. Hurwitz [BB15c]. Hybrid [Bor11f, Bor11-27, Bor11-28]. Hype [BB13r, BB13s, BB14l, BB14k]. hyperbolic [BB98c, BB98d]. Hypergeometric [BBBC07, BBG93b, BCP05]. Hypergeometry [CDS20]. hypertangent [BS89]. hypotheses [BF89a]. Hypothesis [Alt20].

IBM [BB11o, Bor11e]. ICERM [BBL^{+}13, BB14a, BBC^{+}14a]. ICMI [lL09, Hd12]. ICMS [HY14]. I’d [Bor11e]. Ideas [JJ20]. Identities [BBB06a, BB05g, Bor07l, BV24, Dil20, BBG95a, BBB05, Bor85b, Bor93j, BBG94c, Bor97q, Bor97v, Bor97w, BBB06b, Liu00, Liu01, XY12]. Identity [BB88c, Bor87l, BB91c, KMY00, Liu00, PP11]. If [Bor11e, BB11-29, BB13y, BB14c, BB14w]. II [BB15a, BB15j, BB15k, BS87, BL92d, BLLN95, BLZ99, BLZ01, Bor03c, Bor03n, Bor04-27, BC04a, Bor05-31, Bor06n, Bor06s, Bor07j, Bor08f, Bor10d, Bor10r, Bor10-27, BM10, BBSW11, Bor11-35, Bor12f, BBSW12, Bor12-31, Bor13h, Bor13-35, Bor15h, Kom02, Pea07, War03]. III [Bor06t, Bor13i]. Illinois [AAB^{+}88]. illus [BB93g]. image [HYG09, WM07]. images [Jal24]. Imaging [Sch15]. Impact [BS99c, Bor00n, BB15b, BB15o, BS99a]. implementation [BJCW13, BCJW13]. Implication [Bor05-38, Bor05-39]. Implications [Bor04t, Bor05q, Bor05r, Bor08b, BB05b]. Implicit [LY18, Bor11-38]. important [BB11-27]. Improved [PT20, BMS13]. Improvement [TK97, FK00]. Imre [Bor09b]. IMU [Bor00a]. Inaccessible [BBMW11, BBMW13, BBMW16]. Inaugural [Bor93n]. inborn [BB11q]. Inclusions [BD15, BD16b]. incomplete [BB15c, BC09]. Inconsistent [BT13b, BT15]. independent [AK22]. Index [BBLZ14g, BBLZ14a]. indexed [OBB^{+}96]. Indian [BB91d, BB11h, BB12r, BB16d, Bor91p, Bor91q]. indicates [BB10h]. indicator [BBLZ13e]. indicators [BBLZ14r]. Indigenous [AD20, PL20]. individual [BBLZ13b, BBLZ14e, BBLZ14j, BBLZ15b, BBG04b]. Individuals [BB15s]. induced [Bor93j]. Inequalities [BSW82, BB93f, Bor99y, Bor99z, Bor09m, BF98, Bor98p, BF01, BG03b]. Inequality [Bor84d, BBFG00, Bor05x, Bor08k, SW21, Bor77a, Bor86c, Bor93b, Bor97g, Bor98g, BBFG01, Mer15]. Inexact [HD07]. inferred [BCM02, BCM03]. Infimal [BBEM10]. infimum [BBWY11e, BY12f]. Infinite [ANR18, Bor92j, BPB99, Bor79a, Bor81c, BK83, Bor83c, Bor83f, BW86, Bor91h, BL91d, Bor92d, Bor92n, Bor93k, Bor94g, Bor95m, Bor95n, BFL02, Bor11v, RZ15]. Infinite-Dimensional [ANR18, BW86]. Infinity [BB91d, Bor15b, Bor16d]. inflationary [BB14m]. Information [BLLN94, PL20, Bor94g, BLN94a, BLN95, Bor95m, Bor95n, BLLN95, BLN96]. informatique [Bor00o]. inhomogeneous [Kom00, Kom02, Kom04]. Initial [Goo20]. Initiatives [Bor00a, Bor01n, Bor01m, Bor02n, Bor03h]. Innovation [Bor09o, Bor12n]. Innumeracy [BB11p]. Insight [Bor99i, Bor99j, Bor99k, Bor99l, Bor07t, Bor07u]. Insignificance [Alt20]. inspired [GG07]. inspiring [KMT16]. Institute [SBW84]. institutional [Bor16h]. Instruments [MTB16]. Insult [BB12-33, BB12-32]. Integer [BB09j, BC96, Bor02a, BC07, Bor09p, Bor09q, Bor10r, SV20, BB93e, BL97, BL00b]. integrability [BM97f, BM00]. Integral [BB06b, BZ87, BBBG08, Bor84b, BB95f, BY12a, BY14a, Cra04, Cvi10]. Integrals [BBC06, BBBC07, Borxx, BCC10, Bor10x, Bor10-28, Bor10-29, Bor11f, Bor11-30, Bor11-27, Bor11-28, BS11d, BS11e, Bor11-34, Bor11-35, Bor11-33, Bor12r, Bor12-32, BSW13, BBC07a, BBCM07a, BBC10, BB10d, BB12-29, BBB15, BBB08, BB84c, BZ92, Bor00r, Bor01p, Bor01q, Bor01r, BBM01, BB01c, BBM02, Bor07e, BBC08a, BBGW11, BNSW11, BS13, DS20]. integrands [BY12a, BY14a]. integrate [Bor94n]. Integration [BB08e, BB08b, BB09o, BB09b, BB11b]. integrity [BB10c, BB12-28]. Interactive [Bor98j, Bor99p, Bor09-27, BWB97]. Interdisciplinarity [Bor07p, Bor12n]. Interdisciplinary [Bor07q]. Interference [HMM20]. interior [BG01, BG03a]. interiors [BL92c]. International [Bor03p, Bor09r, HY14, IEE08, ABD03, BB12o, BB13-27, BF06b, CGM95, Ano15, BBLZ14p, Bor01n, Bor01m, Bor02n]. Internet [Bor01l, CZX21]. interpolation [Bor98o, DS20]. intersection [BBL97c, BBL99]. Interview [Ano15, BB15u, BB16i, Bor12w]. intriguing [Bor93o, BB95f]. Intrinsic [Kru18]. Introduction [BZ20a, BC21, Bor97l, Bor02o, Bor07r, Bor07s, Bor07t, Bor07u, Bor09s, Bor09q, BR10, Bor11k, Bor11l, Bor13g, BvdPSZ14, Bor20, Bre20b, BL20, Bor08c, BD09, Bor10s, BD11, BS11c, BS12a]. invariance [BLZ99, BLZ01]. invariants [BB98c, BB98d]. invented [BB11r]. inventor [BB12-38]. Inverse [Bor97h, Bor08p, Bor09t, Bor09v, Bor10k, Bor10v, Bor10w, Bor12q, Bor13l, Bor13p, AL10, BBC^{+}11b, Bor92k, Bor92l, Bor92m, Bor12p, BT14b, BT14a, BT17]. investigation [BBGPxx]. Investing [BB14n, BBLZ13c, BBLZ13h, BBLZ14g]. Investment [BBLZ13a]. investments [BBLZ15g]. Investor [Bor14c]. investors [BBLZ13b, BBLZ14e, BBLZ14j, BBLZ15b, BBLZ15d]. Involving [BSW82, Bor93o, Mer15, XY12]. ISBN [Ban10, BC96, Bor05g, Bor06o, Bor09b, Bor11-38, BO11b, Bou06, Coh15, Odl11, Sha05]. ISBN-10 [BO11b, Bou06]. ISBN-13 [BO11b, Bou06]. Ising [BB06b, BBC06, BBBC07]. Ising-Class [BBBC07]. Islamic [SV14]. Isometry [BGMS21]. Israel [Bor90b, RZ15]. Issue [AHLC^{+}17b, BC21, AAB12]. Issues [BL99, Bor00t, Bor03p]. Italian [Bor08a]. Italy [ABD03]. iterated [BR16]. Iteration [BB89b, BBxxb, BT13a, Gil18, AB12, AB13, BB86b, BB90b, BBG93b, Bor94a, BT14c]. Iterations [Bor89g, Bor89h, BB93f, BLT17, BB91b, BRS92, Bor93j, BS10c, BS10d, Bor10i, Bor10j, Bor11r, Bor11s, Bor13r, BLT15, BLT16]. Iterative [Bor92k, Bor92l, Bor92m, WSL16, XC11]. IV [Bor06u].

J [BB13k, Ban10, Bor92b, BC96, Odl11]. J-P [Bor92b]. J. [Cra04, MR11, SV14, Zăl86]. Jacobi [BB88c, BB91c]. Jacobian [HGB93]. Jameson [BBL97c, BBL99]. January [AMM10, BBLZ16b]. Japanese [BB13t, BBLZ16a]. Jauregui [ABBS12]. Jeopardy [BB13a]. jiao [lL09]. JMB [Bor14c]. John [BB93g, BO11b, BS14a, IEE08, BB09g, Bor76b, Jac09]. John-Steiner [BO11b]. Joint [BB00a, BB01b]. Jon [How14, Ano15, Bai17e, Bor07-27]. Jonathan [Bou06, Hoa05, How14, MW16, Sha05, Tod03, Zei05, Ano16, Bai91, BBB^{+}13, Bai16c, Bai16d, Bai17a, Bai17b, Bai17c, BBB^{+}20, Bai21, Ber88, BBS17, BaHO20, BE16, Bre17, CW16, Cha16, Dev17, KMT16]. journal [BB11w, BS97b]. journalists [BB12-30]. journey [KMT16]. Joy [Bor98i]. JSrg [Bor05g]. July [CGM95, BBC^{+}14a]. June [AAB^{+}88, Bor09b]. Just [BB12g, BB14x].

Kadec [Bor88j, BF89c, BV93a, BV94c]. Kanigel [BB91d]. Karl [BBLZ14d]. Keeps [Sei01]. Kelly [BBLZ13f]. Kenneth [BB13e, BB13f]. Kentucky [BB13-41]. Kepler [BB14j]. Kernels [BLN94b]. Kestelman [BO11a]. Khinchine [Bor95q, Bor95r]. Khintchine [BBC97a, BBC9x]. killer [BB12a]. Kills [BB15-31]. kind [BZ92]. kinds [BN84]. Klee [BV93a, BV94c]. Knew [BB91d, Bor15b, Bor16d]. knots [BB98c, BB98d]. Know [BB14c, BB13y, BB14w]. Knowledge [BS05, ABD03, BF06b]. Known [CDH^{+}21]. Konjagin [Bor13h]. Korea [HY14]. Kós [FK00]. Krasnosel’ski [BRS92, Bor77b]. Kurt [BBC14c, BBC14d, BBC15].

L [Bai16a, Bor11-38, SV14]. L. [BSW82]. Laboratories [Bor99b, Bor99c]. Labs [BL99, Bor99p]. ladder [BB11d]. lagging [BB13-27]. Lagrange [BMCL18, Bor80b, Bor81d, BZ16]. Lagrangean [Bor79b]. Lagrangeans [Bor80d]. Lagrangian [Bor81e]. Laguerre [BBC07c, Bor07i, Bor07j, BBC08b]. Lambert [Bor16l, Bor16m, BL16]. Large [BBKL16, BBKL17, JWDS^{+}14, BBLZ13d, BBK14, DF05, LW18, LW19, Ray97, WM07, XH08, ZSZ16]. Large-Scale [JWDS^{+}14, DF05, LW18, LW19, WM07, XH08, ZSZ16]. largest [Bor10-30]. last [BB13t]. Later [BB13s, BB13r, BD95]. Latest [BB10h, BBLZ14j, BB12o, BB12-50]. Latin [BS14b, BS14a]. Lattice [BBCZ13, BLL94, BB94b, BBP95, BGM^{+}13, BB13g, BBT85, BBS89, BL92d, BBP98, BBS13b, BBS14b]. Lattices [BBSZ87, BS83, BY84, BS84b, BBSZ88]. Lau [Bor13h]. Launch [Bor03-31]. Laurie [Bor05g]. Law [BB12-39, BB12h, BB15z, BB15y, Bor15l]. Lawrence [Bor07c]. laws [BB10b]. Leader [Bor09b]. leadership [BB12-44]. leads [BB13w]. Learned [Dev20]. Learning [Bor05-42, Bor05-43, Bor05-44, MTB16]. Lecture [Bor06q, Bor06p, Bor09-27]. Lectures [Bor06r, Bor06s, Bor06t, Bor06u, Bor09-30, Bor09-28, Bor09-31, Bor09-29, Bor09-27, Bor13-31, Bor15r]. Legacies [BaHO20]. Legacy [BBB^{+}20, Dev17, BBC14c, BBC14d, BBC15]. Legendre [BB95b, BB97a, BBC00a, BBC01, Bor87d, Bor88a, Bor88b, Bor88c, Bor88d, Bor88e, BV00b, BV01, BV10a, BY12a, BY14a, TK97]. Legendre-type [BY12a, BY14a]. Leibniz [BWB97]. lemma [Bor79d, Bor83d]. Length [BCM20]. LENR [BB15f, BB15g, BB15u, BB16i]. lesson [BB13t]. Lessons [BBLZ15f, BB15x, KMZ^{+}03]. let [Bor13c]. Letter [Bor11b, CW16, Cha16, Zăl86]. Level [BB93b, BS99d, Bor00s, Bor11g, Bor11-37, BS99b, BS00]. Levi [Bai16a]. Lewis [Bou06, Tod03]. Lexicographic [Bor80c]. Library [Bor02f, Bor03-35]. Life [BB12-37, BB13f, BBB^{+}20, BB91d, Bor93m, Bor03q, Bor03r, Bor03s, Bor03t, Bor03u, Bor03v, Bor03w, Bor05y, Bor05z, Bor06-27, Bor07v, Bor08l, Bor08m, Bor10t, Bor10u, Bor11w, Bor11x, Bor11y, Bor11z, Bor12o, Bor13o, Bor14p, Bor14s, Bor14q, Bor14r, Bor16n, Bor16o, BB11g, BB12g, BB13e, Bor91p, Bor91q, BM06, Bor08a, Bor15b]. light [Fab89]. Like [BBP95, WSL16, AG99, BBB05, BBB06a, Bor87m, BL91b, BB96c, BB97c, Bor97v, Bor97w, BBP98, BB05f, BB05c, Bor07-27, Bor15d, DABY15, GDT15, Gui17, JD13]. likely [BB16g, BB16h]. Liljedahl [Coh15]. Limit [BS17, BF93c, BF95b]. Limiting [Bor79b, BZ98, Bor80d, Bor81e]. Limits [CS21, WG17, BBS13b, BBS14b]. line [BW03, IP17, IP18, YW12]. Linear [BB93b, Bor72, BD86, BBS14a, HMM20, BB95a, BB96a, BBL97c, BB99c, BBL99, BBT00, BBW07, BWY10, BBWY11e, Bor84a, BFG87, BD89, Bor93b, BM09, BM10, BY12b, BY13b, DL02, DLL05, DABY15, HLZ14, HLY16, KJR16, LLS11, LZ14, Li15, ZL22]. Linearly [CPRZ20, DGLV20]. lines [Bor79h]. link [BB15e]. links [BB98c, BB98d]. Lipschitz [BB11a, Bor87m, Bor90g, Bor90h, Bor90i, Bor90j, Bor90y, Bor90z, Bor90-27, Bor90-28, Bor90a, Bor90-40, Bor90-41, Bor90-42, Bor90-43, Bor91d, Bor91a, Bor91r, Bor91s, Bor91t, Bor91u, Bor92a, BFV93a, BFV93b, BV95a, BM95, BMW95, Bor95d, BM96a, BM96b, BFV97, BM97c, BM97d, BM97e, BMW97, BM98a, BW98b, BM98b, Bor98o, BW00, BVW01, BFL02, BGV02, BW03, BVW03, BW05b]. Lipschitz-constant [BVW01, BVW03]. Lipschitzian [BBEM10, BS84a, BLM99, BLM00]. Lists [Bor05h, Bor05-27]. literacy [BB13n, BB13m]. Literate [BB14n, BBLZ14e]. Literature [BB05e, BM07a, Bor02g]. little [Bor91i]. Littlewood [HC09]. live [BB13j]. Local [BF89a, BVW01, BVW03, QR07, BB96a, BB99c, Bor79g, JN03]. Locally [BFV93a, BFV93b, BB11a, BFV97, QR07]. locating [JY12]. Log [BB84e, BS11a, BBSW11, Bor11f, BS11d, BS11e, BS12b, BBSW12, Bor12r, BBB15, BS13]. log-gamma [BBB15]. Log-sine [BS11a, BBSW11, Bor11f, BS11d, BS11e, BS12b, BBSW12, BS13]. Logarithmic [BB93f]. Logarithms [BCLM16, BHL16b, BHL16a, BCLM17, BHL17, Cha03]. Logsine [Bor11-34, Bor11-35]. Long [Bor04p, Bor05l, Bor05m, Bor05n, Bor05o, Bor05p, Bor06l, R^{+}05, BBLZ14j, Bor03x, Bor03y, Bor06-28]. long-range [Bor06-28]. long-term [BBLZ14j]. lose [BBLZ15d]. losing [BBLZ15a]. Love [Dev20]. Loving [BO11b]. Low [BB14p, BB14q, BB13m]. Lowell [Bor77d]. Lower [CPRZ20, Bor90k, Bor90l, BMS13, BSM13, BLZ99, BLZ01]. LRP [Bor05-28]. lsc [Bor90a]. Ltd [Ban10]. Ludens [Thé16]. Luke [Odl11]. lun [lL09].

M [Ano16, Bai16d, BBB^{+}20, Ban10, Ber88, BBS17, BaHO20, Cha16, Hoa05, How14, KMT16, MW16, Odl11, Tod03, Zăl86]. MA [Odl11]. Machin [BBG04b]. Machin-type [BBG04b]. machines [BB10a]. Maclaurin [BB06a, BB08d, Bor06j, Bor06k]. Maclaurin-Based [BB06a, BB08d, Bor06j, Bor06k]. mad [BB13u]. Madelung [BBT85, BBP95, BBP98]. Madelung-Like [BBP95, BBP98]. Magic [BB11e]. Mahler [BCM20, Bor91j, Bor91k, Bor91i, Bor91l, Bor91m, Bor92e, Bor92f, Bor92i, BS11a, BBSW11, Bor11f, Bor11-34, Bor11-35, BS12b, BBSW12, BS13, BBC14c, BBC14d, BBC15]. Main [BB97d]. Make [BB12h, BB12a, BBLZ13b]. Making [BBGP95c, BBGP96, Bea13]. Malaises [BB93g]. Man [BB91d, Bor15b, Bor16d]. Manage [Bor12n]. Management [Bor06v, Bor06w, Bor06x, Bor06y, Bor07f, PR92, ABD03, BF06b]. Managing [BS03, Bor09o]. manifolds [BB98c, BB98d]. Mann [BRS92]. Many [BB16r, BB15q, BB16q, BR84]. MAPLE [Bor89f, Bor90-29, Bor90-30, Bor90-31, Bor90-32, Bor90-33, Bor90-34, Bor90-35, Bor90-36, Bor90-37, Bor90-38, Bor90-39, Bor92c, Bor92i, BL92e, BM97b, Bor06z, BS11c]. mapping [BB98a, BB99b, BM97f, BM00]. Mappings [RZ18, BB95a, BS83, BS84a, BS84b, Bor86b, Bor91d, Bor92a, BM97e, BM09, BM10, Bor11-38]. Maps [Bor09-29, GLR18, BZ88]. March [IMR92, BB13v]. Market [BB15r, BBSL17b, BBSL18, BBLZ14a, BBLZ14d, BBLZ14f, BBLZ14q, BBLZ14i, BBLZ15c, BBLZ16a, BBSL17a]. Markets [BBS^{+}16a]. Mars [BB12g, BB12-37]. Martians [BB12-37]. Marvels [Bor02p]. Massachusetts [BB13w]. Master [Zei05]. mate [BB12-36]. matematica [BB95d]. Math [Bor81a, Bor98j, Bor99d, BL99, Bor01e, Bor01l, Bor01m, Bor02s, Bor02t, Bor03g, Bor06m, Bor06n, Bor07w, Bor08n, Bor08o, BZ11, KMZ^{+}03, BB10j, BB11q, BB11x, BB12v, BB12-44, BB12-49, BB13w, BWB97, Bor98r, Bor14b].

Math. [Zăl86]. MathBrowser [Bor97k]. mathematic [BBLZ14g]. Mathematica [BS12a]. Mathematical [BB08a, BBBZ10b, BB11l, BB11m, BBB12, BB14r, BBC^{+}14a, BB15i, Bai16d, Ber88, BB93g, BBGP95b, Bor96d, Bor97i, Bor97j, Bor98e, Bor98a, BB01d, Bor01e, BB01e, Bor01n, Bor02j, Bor02k, Bor02n, Bor02p, Bor02s, Bor02t, Bor03o, Bor03p, BS03, Bor05-42, Bor05-43, Bor05-44, BLM^{+}07, BM07b, BM07a, Bor11-31, Bor13m, Bor13n, BS14a, BC15a, BC16, CW16, Coh15, bVP21, Sch15, SV14, TB80, BBBZ10a, BB10f, BB11w, BB14m, BB15b, BB15h, BB15o, BB15p, BB15v, BBMW17, Bor91p, Bor91q, Bor95t, Bor95u, Bor97x, BBGPxx, Bor02g, Bor08q, Bor09e, Bor09f, Bor09g, Bor10q, BS11c, BS12a, BS14b, RZ15, ABD03, BF06b, HY14, Bor14c]. Mathematician [BB12-38, Bai17b, Bai17c, Bai21, Bor98h, Bor01a, Bor01b, Bor01c, Bor01d, Bor02j, Bor02k, Bor02l, Bor02m, Bor05a, Bor16f, Bor06e, Bor15b, CKR15]. Mathematician/physicist/inventor [BB12-38]. Mathematicians [BB16r, Bor03-30, BMP05, Dev20, Goo20, BB16q, Coh15]. Mathematics [ABMMY13, AD20, BB99a, BB05a, BBC^{+}07b, BB09a, BBBZ10b, BB11s, BB12r, BBL^{+}13, BB14a, BB15n, BB16d, BB16c, BB02, Bor92i, Bor93c, Bor93d, Bor93q, Bor94f, Bor94p, Bor94r, Bor94q, BBGP95a, BD95, BBB^{+}96a, Bor96b, Bor96e, Bor96i, Bor96k, Bor97a, Bor97b, Bor97c, BB97d, Bor97n, Bor98j, Bor99e, Bor99f, Bor99b, Bor99c, Bor99g, Bor99h, Bor99i, Bor99j, Bor99k, Bor99l, Bor99p, BS99c, BS99d, Bor00b, Bor00c, Bor00d, Bor00e, Bor00f, Bor00g, Bor00h, Bor00i, Bor00j, Bor00k, Bor00n, Bor00s, Bor01h, Bor01i, Bor01j, Bor01k, Bor02f, Bor02a, BMPR02, Bor02r, Bor03j, Bor03k, Bor03l, Bor03m, Bor03n, BBG03, Bor03z, Bor03-27, Bor03-28, Bor03-29, Bor03-35, Bor04j, Bor04k, Bor04q, Bor04r, Bor04s, Bor04t, Bor04u, BB04b, Bor05h, Bor05q, Bor05r, Bor05t, Bor05u, Bor05v]. Mathematics [Bor05w, BS05, Bor05-27, Bor05-31, Bor05-29, Bor05-30, Bor05-38, Bor05-39, Bor05-47, Bor05-48, Bor05-49, Bor05-50, Bor05-51, Bor05-52, BBG06, Bor06z, Bor06v, Bor06w, Bor06x, Bor06y, Bor06-29, Bor06-37, Bor06-38, Bor06-39, Bor07f, Bor07m, Bor07n, Bor07r, Bor07s, Bor07t, Bor07u, BRR08, BB08h, Bor08r, BB08g, Bor09r, Bor10a, Bor12i, Bor12s, BJ12, Bor12z, Bor15j, Bor16g, Bor16-28, Bou06, CDH^{+}21, CC20b, DD15, ES01, Fer91, GS08, Goo20, HDG^{+}15, Hol20, JB21, KAA^{+}15, lL09, Lor09, MTB16, Odl11, PL20, Zei05, dPB21, AMM10, ABMMY14, BB01a, BB05b, BBKW06, BB09e, BBBZ10a, BB11r, BB11i, BB12z, BB12x, BB12c, BB12-39, BB12h, BBB^{+}13, BB14t, BBLZ14l, BBLZ14s, BB14o, BB14-30, BB14-31, BB15b, BB15o, BB15-27, BB16f, BB16e, BB16s, BB17, Ban10, BB88b, BB91a, BB99d, BB99e]. mathematics [Bor90o, Bor90p, Bor94c, Bor94d, Bor94e, Bor95f, Bor95g, Bor95h, Bor95i, Bor95j, Bor95k, Bor95l, BBGP95c, Bor95w, Bor96c, BBGP96, BBC^{+}96, Bor97e, Bor97d, Bor97h, BWB97, BBJC97, BBC^{+}97b, BC98b, BC99, BS99a, BS99b, BS00, Bor01f, BBG04a, Bor05-45, BF06a, Bor08c, Bor08b, BD09, Bor09y, BB10l, BD11, Bor11b, Bor11c, Bor14x, Bor14y, Bor16b, Hd12, KMT16, Sha05, Bor06o, Bor09b, BO11b, HF05, Hoa05, Sha05, Zei05]. Mathematics/Ouvrages [Bou06]. Mathematik [BD11, Fal96]. Mathématiques [Bor00o]. Mathemématiques [Bou06]. MathResource [BWB97, Bor97k]. Maths [Bor09u, Bor12m, BB11g, BB11f, BB12i, BB12k, BB13p, BB14e, Bor11e]. matrices [Bor84c, BR84]. Matrix [ABT13a, ABT14a, BRxx, Bor13j, Bor14f, Bor14g, Bor15g, Bor16p, HNP10, HLZ15a, HLZ15b, IP17, IP18, LL13]. Matter [BB09c, BB12-30, Bor10f, Bor10g]. Matters [Bor09u, BB14-30, BB14-31, Bor97v, Bor97w]. Mattingly [Bai16a, BE16]. mature [Bor94n]. Max [BB14r]. Maximal [BB96a, BB99c, Bor06-30, Bor09j, Bor09k, Bor12y, BD15, BD16b, EB08, Bor82c, BW98b, BW00, BVW01, BVW03, BW05b, Bor06-32, Bor07x, Bor10n, Bor12j, Bor12k]. Maximality [Bor06-31, Bor06-32, Bor07x, BY12b, BY13b]. Maximally [BY12f, BML18, BBWY11b, BBWY11c, BBWY11e, BBWY12b, BBWY12c, BY12b, BY12d, BY12e, BY13b, BY13a, BY13c, BY14b, BY14c]. Maximization [Bor94g, Bor04-32, Bor13-32, BZ20b, BBM99, BBM00]. Maximizations [Bor77c]. maximize [Bor09n]. Maximizing [Bor99q, Bor99r, Bor99s, Bor00p, Bor00q, Bor04v, Bor05-32]. Maximum [BL93c, BLN94a, BLLN94, Bor95m, Bor95n, BLLN95, BLN96, Bor97l, Bor01o, BCM02, Bor05-33, Bor06-33, Bor08p, Bor09v, Bor10v, Bor10w, Bor12p, Bor12q, Bor13p, BHP14, PHBH12, Bor92n, Bor93k, BL93b, BCM03, BH19, PHB12, PHB14]. May [BW95a, BBS17, IEE08, KG04, RZ15, BW97a, BW98a, Bor15d]. Maybe [BB12-37]. Me [Bor04n, Bor11j, Bor11k, Bor11l, Bor11m, Bor11n, Bor14d, Bor12c, Bor12d, Bor15e]. Mean [BB84a, BB89b, Bor89g, Bor89h, BB93f, BB97b, Bor99y, Bor99z, Bor99-27, BBxxb, BB00b, Bor00u, BB04a, BBS14a, BB11-28, BB13o, Bor87d, Bor88a, Bor88b, Bor88c, Bor88d, Bor88e, Bor88f, Bor89e, BB90b, BBG93b, Bor94a, BW98a, Bor98p, BB16t, IP17, IP18]. Mean-Value [Bor99-27, Bor00u]. Meaning [DD15]. Means [BB87c, Bor93j, BLM96, BLM97]. Measures [BCM20, Bor11f, Bor11-27, Bor11-28, Bor11-34, Bor11-35, Bor12-32, BS11a, BBSW11, BS12b, BBSW12, BS13]. medal [Bor89a, Bor14b]. media [BB12f, BB15-27, Bor12-28]. medical [Bor14a, HYG09]. Medicine [Sel16]. Medieval [SV14]. Meet [Bor14b]. meeting [Bai17e]. Meetings [Bor11-29, BL16]. meets [Bor05k]. Melbourne [BR14b]. Memorial [IEE08, SBW84, Bor06a]. Memoriam [TSB13]. memory [BSZ13]. Merchants [BB11t]. Mersenne [Cha03]. mesh [AK22]. mesh-independent [AK22]. mess [BB13-29, BB13-30]. Meters [BB13-38, Bor12-27, BB12l, BB12-46, BB13-37]. Method [ABT15, BL17a, BL17b, BT13b, BLS^{+}17, BLS^{+}18, HDL21, AR13, AX20, ABT16, AP16, AK22, BBL94, BB95b, BB97a, BBL97b, BS17, BH95, BT15, BLS^{+}16, DL02, DLL05, DHSZ06, DK16, FN15, Fle05, GS02, HYG09, HD07, HLZ14, HL15a, HLZ15a, HLZ15b, HLY16, IP17, IP18, Jal24, KJR16, La 09, LLS11, LZ14, Li15, LL13, Mar91, MR96, MPB16, NWY09, NWY10, NFB17a, NFB17b, PT14, PD18, Ray93, Ray97, RS02, WM07, WSdSY15, XH08, YW12, ZH06, ZSQ10, ZSZ16]. Methodology [BBGP95b]. Methods [ABT13a, ABT13b, ABT14a, ABT14b, BB88a, BL93c, Bor97l, Bor00t, Bor01o, BZ02a, Bor02b, Bor05-33, Bor06-33, Bor08p, Bor09q, Bor09v, Bor09-27, Bor10k, Bor10v, Bor10w, Bor12q, Bor13k, Bor13l, Bor13p, BST13, DLR20, PR92, Sch15, ABT13c, BB05b, BB10g, Bor92k, Bor92l, Bor92m, Bor94g, BLN95, Bor95m, Bor95n, Bor98k, BZ06, Bor12p, Bor13j, BZ13, Bor14f, Bor14g, BT14b, BT14a, Bor15g, BST15, Bor15r, Bor16p, BT17, DF05, GDT15, HNP10, HL15b, JD13, PHBH12]. Metric [BBT98, BGM18, BK80, BZ95, BZ96]. Metrical [HMM20]. Michel [Bor17b]. mid [BBLZ14i]. mid-term [BBLZ14i]. Might [CDH^{+}21, Bor07-27]. million [BB14e]. millions [BB15q]. mine [BB12h]. ming [lL09]. Minimal [Bor89c, Bor89d, Bor90y, Bor90z, Bor90-27, Bor90-28, Bor91a, BFK91, Bor95o, Bor95p, BF89a, BM97e, BK04]. Minimality [Bor87c, Bor82b, Bor86d, Bor87b, BM97f, BM00]. minimax [BZ86, Bor14z, Bor16-27]. Minimization [BLL94, BLN94b, Bor09-30, Bor09-28, Bor09-31, Bor09-29, Bor09-27, BL91b, Bor92j, BV09, NWY10, Ray97, XWQ14]. minimizing [HL15a, NWY09]. minimum [Bor79c, Bor80e]. Miraculous [Fin95]. miscalculate [BB11c]. Missing [Bor09c, BB15e]. Misuse [BB09h]. mixed [BH19]. MKM [ABD03, BF06b]. modal [Bor96e]. model [Bor16h, Cam16, ZSZ16]. Modelling [Bor13q, BHP14, PHB13, PHB14, Bea13]. Models [JJ20, BL92d, Cam16]. Modern [Bor09z, BB12-34, BB12-35, BB15b, BB15o, BS11c, BS12a]. Moderne [Fal96]. Modified [LL13, BS17, Jal24, XSW12]. MODSIM [Bea13]. Modular [BBB97c, BBB00b, BBB04b, BBB16, BBB97a, Bor85b, Bor86f, Bor87g, Bor87f, BB89a, BBB89, BBG94c, Liu00]. moduli [Zha13]. modulo [ZS12, ZZ14]. Moll [Odl11]. moment [Bor90e, Bor90f, BL91c, BGL93, BH94a, BH94b, BL94a, BH95]. Moments [BS07, BS08, Bor10x, BBGW11, Bor11-30, Bor14t, BS16a, TB00, BBBG08, BH19]. Mono [Ber88]. Mono- [Ber88]. Monochrome [Bor79h]. monoids [Bor15f, Bor16i]. Monotone [AHLC^{+}17a, AHLC^{+}17b, BBWY11d, BBWY13, Bor72, Bor02b, Bor04o, Bor05-34, Bor05-35, Bor05-36, Bor05-37, BW06, Bor06s, Bor06t, Bor06-34, Bor06-35, Bor06-31, Bor09-29, BBY11, BEY11, BY12c, BBY13, BD15, BML18, EB08, LLT18, Sim18, BB95a, BBC03, BBW07, BWY10, BBWY11b, BBWY11c, BBWY11e, BBWY12b, BBWY12c, Bor86b, BF89a, BFK91, Bor98n, Bor02d, Bor02e, BBL04, BW05a, Bor06-32, BW07, Bor07b, Bor07x, BE08, BG09, Bor12j, Bor12k, BY12f, BY12b, BY12d, BY12e, BY13b, BY13a, BY13c, BY14b, BY14c, BY15, BD16b, HLZ15a, SZ14]. Monotonicity [Bor09j, Bor09k, Bor12y, BBS15b, BBS20, BBB^{+}07, BB96a, BB99c, BBWY11e, Bor82c, Bor06-30, Bor10n, BRS11, Bor12j, Bor12k]. Month [bVP21]. Monthly [BB07a, BB12-47, BB09l, BB09m, BB10k, BC15a, BC16, BC18a, bVP21]. Montreal [KG04]. Moore [BB12-39, BB12h, BB15z, BB15y, Bor15l]. morass [BB10b]. Mordecai [Bor90b]. Mordell [BBC14b, BB15a, BB16a, BB16b, BB18, Bor12e, Bor12f, Bor12r]. Morozov [BMCL18]. Mosco [BB90a, BB93b, Bor88j, BF89c, BV93a, BV94c]. most [Bor16b]. Motivation [Bor09-30]. motive [BB09d]. Movements [BB13-44, BB13-43]. movies [Bor15b]. MR [Bor81a]. MR0716121 [Zăl86]. MR0991866 [BBB97a]. MRI [Jal24]. much [BBLZ15d]. Multi [Bor96e, Bor97m, BBM01, BBM02, Bor97f, Bor16h]. Multi-dimensional [Bor97m, Bor97f]. multi-disciplinary [Bor16h]. multi-institutional [Bor16h]. Multi-modal [Bor96e]. Multi-variable [BBM01, BBM02]. Multidimensional [Bor96f, Bor96g, Bor96h, BH06, BTBT88, Bor97q]. Multifunctional [Bor98k, BZ99a, BZ99b]. Multifunctions [Sim18, BF94a, Bor94b, BF95a, Bor95o, Bor95p, BMS97, BMS99a]. Multimedia [BMPR02]. Multimodal [Bor97n]. multiobjective [MPB16]. Multiple [BBBL98c, BBBL99, BBK00a, Bor10y, BZ11, BBBL97, BBBL98a, BBBL98b, BBK00b, BBK01, BBBL01, BC10, BDT16, JY12]. multiple-zeta [BC10]. Multiplier [Bor80b, Bor81d]. multipliers [Bor80c, BZ16]. Multivalued [Bor77a, Bor79d]. Multivariable [Bor00r, Bor01p, Bor01q, Bor01r]. Multivariate [HYG09, BL92b]. Museum [BB13-41]. Music [Bor12s]. Musicians [BB16r, BB16q]. My [Bor08q, Bor12t, Bor07-28, Bor07-29, Bor07-30, Bor08u, Bor12a]. Mysteries [Bor11-31]. mysterious [BB11-27]. myth [BBLZ13e].

N [BC96, Odl11]. National [Bor05k]. NATO [SBW84]. natural [RP09]. Nature [BB09c, BB15v]. Nearest [BG15b, BG16b, Bor88k, BF89b]. Nearly [Moo18]. Necessary [Bor82b, BTZ95, BTZ98, LY18, BZ88]. need [BB12-30]. needed [BB14o]. needs [Bor13d, Bor13a]. negative [BBWY11e, BY12f, LL13]. negative-infimum [BBWY11e]. Nested [BdB91]. Network [Bor99b, Bor99c]. Networking [Bor98e, BB15-31]. networks [BB12v]. Neumann [BB93a]. Nevanlinna [Bor03o]. Neverending [BvdPSZ14]. Newcastle [Bai17a]. Newfoundland [IEE08, SBW84]. Newly [BB12i]. news [BB12t, BB12a]. Newton [BWB97, CDH^{+}21]. Next [Bor02c, Bor02q, BB16l]. NI [BE08]. Nielsen [BS15b]. Nikodým [GLR18]. NJ [Bor09b]. NMR [BMN98, BMN00]. No [BB13r, BB13s, BM97a, BB13i, BKW02, Cam16, Zăl86, BB12-34, BB12-35]. no. [BZ02a]. Nobel [Bor14b]. Noether [BB12x]. noise [Jal24]. Non [ANR18, Bor72, Bor05-33, Bor06-33, Bor13p, Bor16u, Bor16v, Bor16w, Bor16x, Bor16y, Gil18, AB12, AB13, BBWY11b, BBWY12b, BZ94a, BE08, BS10a, Bor15r, LL13, Sel16, BM07c]. Non- [Bor05-33, Bor06-33]. Non-Convex [Bor16u, Bor16v, Bor16w, Bor16x, Bor16y, Bor13p, Gil18, AB12, AB13]. non-expansive [BS10a]. Non-Linear [Bor72]. non-negative [LL13]. non-reflexive [BBWY11b, BBWY12b, BZ94a, BE08]. Non-Smooth [ANR18, BM07c]. non-Western [Sel16]. nonattaining [BK01]. Nonconvex [ABT15, BC18b, Bor10k, Bor13r, ABT16, BJ97, BZ98, BJ98, Bor12p]. nondifferentiability [BG09]. Nonexpansive [BS83, BS84b, Bor09-29, BRS11]. Nonisolated [AI18]. Nonlinear [BBC09, Bor99a, BL00a, BZ02a, BZ02b, Bou06, Dil20, Tod03, BL06, IMR92, ZL22]. nonlocal [PT14]. Nonmonotone [BL17a, BL17b, GS02, QYX14, XWQ14, AP16, IP17, IP18, Li15, NFB17a, NFB17b, YW12, ZSQ10]. nonnegative [HNP10, HLZ15a, HLZ15b, WM07]. Nonnormality [BB12-40]. nonreflexive [BL93a, BV93b, BV94d, BZ94b, BZ97]. nonsense [BB12-42, BB12-43, BB13d, BB13-34]. Nonsmooth [BC18b, Bor94h, Bor94i, Bor94j, Bor94k, BM07d, CFG^{+}18, WB87, Bor98k, BZ99a, BZ99b, XWQ14, YW12]. noon [BBLZ15c]. Norm [Bor86a, BST13, BST15, Ara07, Ara08, BFG03]. Normal [BB13x, BB14c, BB13y, BB14w, BCJW13, BG87]. Normality [BBC^{+}11a, BBC^{+}12b, BBC^{+}12c, BBC^{+}12a, BN84]. Normed [BFV94b, BFG87, BRS92, BFV94c, BFV94a, Bor94l, Bor95s, BLM99, BLM00]. Norms [BBL10, BY84, BV93a, BV94c, BJSM00, BJSM02, BGV02]. notation [BB11e]. Note [BMCL18, BB86a, BM97a, Bor76b, Bor80d, Bor82d, Bor82c, Bor83d, BF94b, Rei02, Tha02]. Notes [Bor06-36, HC09]. notion [JN03]. Notions [Bor87c, BG01, BG03a, Bor86d, Bor87b]. novel [Ade12]. NSW [Bai17a]. Nuclear [BB14q, BB14p]. Null [BM96b, BM98b, BF95c, Bor95a, Bor95b]. Number [Ber88, BB87d, GI01, KG04, Wim88, BB11-27, BB13-42, BB13-47, BB16p, BJCW13, BCJW13, BB93d, BB96d, BB98b, BSZ13]. Numbered [Bor11d, Bor11h]. Numbers [Ade14a, ABBB13, BB88e, BBD97, BBxxa, BBD00, BBD04, Bor09t, Bor13-28, Bor13-29, Bor16-29, Bor16-30, Bor16-31, BBCP04, BB11e, BB12t, BB12a, BB13x, BB14d, BB14x, BCJW13, BBD89, BB90d, Bor11i, Bor13u, Bor13v, Bor13w, Bor13x, Bor13y, Bor13z, Bor13-27, Bor14x, Bor14y, BBD16, Bor16q, Bor16r, Bor16s, RP09, Bai91, Lor90]. Numeracy [BB09i, BB12-41, BB12-53]. Numerical [AX20, BB08e, BB08b, BB12-42, BB12-43, BS99d, BS99b, BS00, Bor00s, Bor09z, BB11b, Bor05g, MR96]. numerique [Bor00o]. Nurturing [Bor03-30].

O [BB13-45, BB13-46]. Obituary [BBS17]. objectives [Bor91h, Bor92d]. Objects [Bor06s, Bor91e, Bor91f, Bor91g, Bor91j, Bor91k, Bor91i, Bor91l, Bor91m, Bor92e, Bor92f, Bor05-34, Bor05-35, Bor05-36, Bor05-37, Bor06-34, Bor06-35]. Observations [BB92b]. odd [BS16b]. odds [BR14b]. Odyssey [BB12u, BB12n]. OEIS [Bor15d, Bor16a, Bor17a]. Official [Bor03-31]. often [Bor15a]. oil [BB12-27, BB12e]. Old [BB14-32, BB12-31, BB12d, BB15q, BB15z, BB15y, Bor15l]. Olver [BB13k]. once [BB13-47, BB15-28]. One [BBB97c, BBB00b, Bor03-33, BBB04b, BBB16, BBB97a, BBB89, BF94a, Bor94b, BF95a, BCFR04]. one-dimensional [BF94a, Bor94b, BF95a]. Online [BBS^{+}15a, BS97b, Bor97o, BBLZ14k, Bor01f]. only [BB13-39]. ontological [BB15b, BB15o]. Ontology [DD15, BB15b, BB15o]. Open [Bor88k, Bor03-34, Pea07, BBS13a, BB13-35, BB13-36, BB98a, BB99b]. openness [Bor87a, BZ88]. Oper. [Zăl86]. Operator [BY12c, BBWY11c, BBWY12c, BY12b, BY12d, BY13b, BY14b, BY15, BG16a, BG18b, KMY00]. Operators [AHLC^{+}17a, AHLC^{+}17b, Bor72, Bor04o, BW06, Bor06t, Bor06-31, BBY11, BML18, EB08, LLT18, BB96a, BB99c, BBW07, BBWY11b, BBWY11d, BBWY12b, BBWY13, Bor82a, BPT84, Bor84e, Bor86e, Bor86b, BF89a, BFK91, Bor92n, BT92, Bor98n, BRLZ99, BLZ99, BRLZ00, BLZ01, Bor05-34, Bor05-35, Bor05-36, Bor05-37, Bor06s, Bor06-34, Bor06-35, Bor06-32, BW07, Bor07b, Bor07x, BE08, BRS11, BEY11, Bor12j, Bor12k, BY12f, BY12e, BBY13, BY13a, BY13c, BY14c, RZ15]. Opinion [BBS13a, BB15m]. Opportunities [BB13q, BB14a, BBC^{+}14a, BB14t, Hol20]. Optimal [NFB17b, Pos13]. Optimality [BW79a, LY18, BW79b, BW81c, BW82a, BW82b]. Optimisation [Bor17b, BM07c, JN03]. Optimization [Ano15, ANR18, ABT13b, ABT14b, AHLC^{+}17a, AHLC^{+}17b, BBLZ13a, BC18b, Bor74, Bor78a, BTZ95, Bor99a, BL00a, Bor02b, Bor12-30, Bor12-31, Bor16l, Bor16m, Bou06, CFG^{+}18, CPRZ20, DGLV20, IMR92, SZ81, SI16, Tod03, AP16, AK22, BBL97c, BBL99, BBC03, Bor77a, Bor81b, BN84, BZ91, BZ93, BL94b, BTZ98, BL06, BL16, DHSZ06, LW18, LW19, MP18, MPB16, NFB17a, WSdSY15, XH08, XSW12, YW12, ZH06, ZSQ10, Zho12, ZSZ16, IMR92]. option [BCM02, BCM03]. Order [BC18b, BD86, Bor87e, EB08, BB84b, BB84d, Bor86e, BB87a, BD89, Bor92g, Bor92h, Bor93f, Bor93g, BF93b, BN94]. order-bounded [Bor86e]. orderings [Bor74]. Organic [Bor96i, BBJC97, BJ12, BBC^{+}96, Bor97e, BBC^{+}97b, BBJC97]. oriented [BD11]. Origami [AD20]. origin [BDT16, BG16a, BG18b]. originating [Bor05j, Bor06i]. Origins [BS14b, BS14a]. OSCAR [IEE08]. oscillatory [BB10d]. Other [Bor00j, Bor00k, Bor05-42, Bor05-43, Bor05-44, Bor16n, GS08, BBMW17, Bor92n, Bor93k, BFV97, Bor05-45, BL16, Tre13]. Our [BB11o, BB14d, BB14r]. out-of-sample [BBLZ14s]. Outlook [BB99a, BB01a]. outperform [BBLZ14a]. Ouvrages [Bou06]. Over-Fitting [BBLZ13a]. Overfitting [BBS^{+}16a, BBLZ17, BBLZ14c, BBLZ14k, BBLZ14s, BBS^{+}15a, BBL^{+}16b, BBL16a, BBL16c]. overlords [BB11o]. Overseas [BB15s, BBLZ15b]. Overview [Bor09-30]. Oxford [BB93g, Bor06o, BO11b, Bor06o]. Oz [Bor11m, Bor11n].

P [Bor92b]. PA [Bor05g]. PACBB [ZH06]. Pacific [Bai91]. packing [BB14j, BB16o, CKM^{+}16, Via16]. pages [Bou06, Sha05]. pain [BB12v, BB12i]. Pairs [Kru18]. Paleo [BB12r, BB16d]. Paleo-Mathematics [BB12r, BB16d]. Pamphlet [BBB03]. Paper [Bor14v, Bor14w, Bor81a, Zăl86]. Papers [BB14p, BB14q, Bor11b, Bor11c, Cam16, KG04]. Paradox [Bor04-32, BB15j, BB15-29, BB10e, BB15k, BB15-30]. Parallel [BB08e, Bor00t, DS20, BB09b, BJCW13]. Parameter [BCF04, BC04a, ZSZ16]. parameters [LLC^{+}95]. Parametric [BBB06b, Geo05]. Parbelos [CDS20]. Pareto [AR13, Bor80a, Bor83e]. Pari [Bor92c]. Paris [CGM95, Bai17e]. Part [AHLC^{+}17a, BLLN94, BB93e, Bor16b, BB15j, BB15k, BL92d, BLN94a, BLZ99, Bor03m, Bor03n, Bor08e, Bor08f, Bor12e, Bor12f, Bor12-30, Bor12-31, Bor13-34, Bor13-35]. Partial [DP18, Bor74, MR96]. Partially [Bor86b, Bor88l, BL92c, BL92d, BL93b, Bor97p, Bor98l, Bor98m, BTZ99, Bor99t, Bor99u, Bor00v]. Partially-finite [Bor88l, BL93b]. partitions [RP09]. Parts [Bor15h]. pass [BB12j, BB12-36]. Past [Bor07a, Cam16, Bor08r]. Patents [BB14q, BB14p]. pathological [BBWY11b, BBWY12b]. Paths [Bor03l, BBG03, Zei05, BBG04a]. pattern [BB16p]. Paul [BB13v]. PDE [AK22]. PDE-constrained [AK22]. peer [Bai17d]. perfect [Bor80d]. Performance [Bor98h, Bor05t, Bor05u, Bor05v, Bor05w, Bor05-47, Bor05-48, Bor05-49, Bor05-50, Bor05-51, Bor05-52, Bor06z, Bor06v, Bor06w, Bor06x, Bor06y, Bor06-37, Bor06-38, Bor06-39, Bor07f, IEE08, BBLZ13d, BBLZ14h, BBLZ14j, BBLZ14s, Cam16, MTCB98]. Person [BB12j]. personal [Bor14c, Mic03]. Perspective [Bor98h, Com18, BB12k, Bor14c]. Perth [Bea13]. perturbation [BCFR04]. perturbations [BZ94a, BZ94b, BZ97]. Perturbed [DGLV20, BV09]. Peter [Bai91, Ber88, Coh15, Bai20, Bor08s]. Peters [Ban10, Odl11, Sha05, Zei05]. Phelps [BBWY11c, BBWY11e, BBWY12c, TSB13]. Philadelphia [Bor05g]. Philosophical [Bor05q, Bor05r, Bor05-38, Bor05-39]. Philosophy [Bor04t, GS08, BB14o, BB14-29, Bor08b]. physicist [BB12-38]. Physics [BB08a, BBC09, BBBZ10b, BB15i, Fer91, GI01, BBBZ10a, BB12x, BBB12, BB15h, BB15p, Bor10q]. PI [Bor90q, Bor90r, Bor90s, Bor90t, Bor90u, Bor90v, Bor90w, Bor90x, AH01, BB11u, BB13c, BB13y, BB13z, BB14f, BB14g, BB14c, BB14v, BB14b, BB15t, BB15-28, BB16k, BB16l, BBBR16, Bai17d, BBBR17, BBB97b, BBB00a, BBB03, BBB04a, BB87d, Bor89e, BBD89, Bor89f, Bor90-29, Bor90-30, Bor90-31, Bor90-32, Bor90-33, Bor90-34, Bor90-35, Bor90-36, Bor90-37, Bor90-38, Bor90-39, Bor91i, Bor93h, Bor93i, BG97a, BBD97, BBB97c, Bor97r, Bor97s, Bor97t, Bor97u, Bor97y, Bor98i, BB98b, Bor98b, Bor99v, Bor99-28, Bor99-29, BBxxc, BBD00, BBB00b, Bor03q, Bor03r, Bor03s, Bor03t, Bor03u, Bor03v, Bor03w, BBD04, BBB04b, Bor05y, Bor05z, Bor06-27, Bor07v, Bor08l, Bor08m, Bor10t, Bor10u, Bor11v, Bor11w, Bor11x, Bor11y, Bor11z, Bor11d, Bor11h, Bor12o, Bor12u, Bor12v, Bor12w, Bor13o, Bor13s, Bor13t]. Pi [Bor14s, Bor14q, Bor14r, Bor14u, Bor14-27, Bor15k, BC15b, BC15a, Bor16n, Bor16o, Bor16c, BBD16, BC16, Bor16b, BBB16, Bre17, Bre20a, Fin95, Gan17, Gui16, Sei01, AL10, BBBP96, BBBP97, BBB97a, BBC^{+}12a, BB13b, BB14w, BB14-28, BB84d, Bor86f, BB87a, Bor87g, Bor87f, Bor89b, BBB89, BB01f, Bor08a, BB14c, BB14w, BB14u, BB15-27, BB16j, BBMW17, Ber88, BB88f, Bor91o, BB96d, BM06, Bor12x, BB16u, Abb00, Ask88, BBB03, BB93g, Cas99, Rob06, Wim88]. picking [BBLZ14h]. PIIGS [BB11v]. pillar [BB12x]. Pioneer [BB16i, BB15u]. PISA [BB13-27]. pitfalls [Bor94c, Bor94d, Bor94e, Bor95f, Bor95g, Bor95h, Bor95i, Bor95j, Bor95k, Bor95l, Bor96c]. Plagiarism [BB13-28]. Plan [Bor04p, Bor05l, Bor05m, Bor05n, Bor05o, Bor05p, Bor06l, R^{+}05, Bor03x, Bor03y, Bor06-28]. plane [Bor79h, BNSW10]. Planet [Bor13t, BB12-51, BB12-52, Bor06f]. plates [BB91d]. Plausible [Bor93c, Bor93d, BBG03, Bor03z, Bor03-27, Bor03-28, Bor03-29, BB04b, Bor04-27, Bor04w, Bor04x, Bor04y, Bor04z, Bor06-29, Bor10a, HF05, Hoa05, Zei05, BB11x]. playing [BB12s, BB12m]. Please [BB13-29, BB13-30]. Pleasure [Bor02l, Bor02m, Bor05a, Bor16f]. Plouffe [BC96, Fin95]. Point [BB88a, BLT17, BBC^{+}11b, Bor84a, BB91b, BLT15, BLT16, HD07]. Points [Bor77c, Bor84d, BB12-48, Bor83e, Bor86c, Bor88k, BF89b, Bor92k, Bor92l, Bor92m, BF93a, BW95a, BW97a, BKW02, BY12e, BY13c, BG15b, BG16b]. Poisson [BB13g, BBCZ13, BBKL16, BBKL17, TB00]. Pol [BB07c]. policy [BB09i]. Political [BB10i]. politicians [BB12-51, BB12-52]. politics [BB12b, BB12-45, BB13u, Bor13c]. polyhedra [Bor00r, Bor01p, Bor01q, Bor01r, BBM01, BBM02]. polylogarithm [Ade12]. polylogarithmic [BBP97, Bor97m, GG07]. Polylogarithms [BBBL98c, BBBL99, Bor14e, BB16b, Bor97q, BBBL01, BS15b]. polynomial [BH95]. Polynomials [BCM20, BBKL16, BBKL17, Dil20, HC09]. Pools [BBLZ14m]. Poor [BB12-44, BBLZ14j]. Poor-quality [BB12-44]. Poorten [BSZ13]. Popper [BBLZ14d]. Portfolio [Bor09o, Bor12n, BBLZ13d, BBL16a, BBL16c]. positive [DABY15]. Possible [Bor07w, Bor07-32, Bor08n, Bor08o, BBxxc]. Possibly [AI18]. postcards [Bor10o]. powers [BC07]. Pp [Ban10, Bai91, Ber88, BB91d, BB93g, BC96, Bor05g, Bor06o, Bor09b, BO11b, Coh15, Odl11, Zei05]. Practical [BL91d]. Practice [BBS16b, BJL^{+}08]. precedent [BB14b]. Precision [BB08a, BB08e, BB08b, BB13q, BB90c, BL92e, BB92a, BB09o, BB09b, BB11b, BBB12, BB15p, Bor10q, DS20]. Preconditioned [MR96]. predict [BBLZ16a]. predicted [BB16f]. prediction [BB14m]. Preface [AAB12, AHLC^{+}17a, AHLC^{+}17b, BMST18a, BMST18b]. Prefer [BB15t, BC15b, BC15a, Bor15k, BC16]. Preiss [Bej94, Dev9x, Fab89, Geo05, KPS16, KPS17, LS00, QR07, YS00]. Preisses’ [Bor89c]. Preparation [PL20, BB12-49]. Prepared [BB15-28]. prescribed [BMW95, BMW97, BW03, BH19]. Presence [Bor99e, Bor99f, Bor99d, Bor16z, Bor13-34, Bor13-35, Bor13-33, BZ13, BLT15, BLT16]. Present [Bor07a]. Presentation [Bor05e, Bor89a]. President [Ano16]. presidential [BB12-42, BB12-43]. Press [BB93g, BC96, Bor06o, Bor09b, BO11b, BS14a]. Previously [BBMW11, BBMW13, BBMW16]. Price [Bai91, Ber88]. prices [BCM02, BCM03]. primality [BBBG94, BBBG95, BW95b, BBBG96, BW97b, BMS13, BSM13]. prime [BB14s, BB16p]. primes [Cha03]. Princeton [Bor09b, BO11b, HDG^{+}15]. principal [LY21]. Principle [BMCL18, Bor03-33, Bor04-31, BHP14, Geo05, YS00, Bor83b, BB84f, Bor86g, BP87, Bor87h, Bor87i, Bor87j, Bor90m, Bor90n, BCM03, BCFR04, Fab89, KPS16, KPS17, LS00, QR07, BCM02]. Principles [BBS16b, BMS99b, Bor06r, Bor06s, Bor06t, Bor06u, Bor09-31, Bej94, BTZ99, BV09]. Prize [BB14e, Bor03o, BB13a, Bai16a, Bor14b, BE16]. Prizes [Bor03o]. Probability [BBLZ13a, BBLZ17, BCM02, BCM03, BB09h, BB12w]. Problem [ABT15, BB07b, BB07a, BB08f, BB09l, BB10k, BB12-47, BD86, Bor13e, Bor13f, Bor13i, WSL16, ABT16, BB16o, BW81d, BD89, BGL93, CKM^{+}16, GDT15, LLS11, PT14, Pos13, Ray97, Via16, Vir14, Zho12]. Problems [AJB86, ABT13a, ABT13b, ABT14a, ABT14b, AC18, AI18, ANO^{+}83, AJ86, BB09m, BB95c, BB96b, BL87, BC18b, BSZ^{+}83, BB85, Bor85a, BN86, BB87c, Bor93l, BB93c, BLN94b, BTZ95, Bor96j, BDT96, BBS^{+}97, BPB99, Bor05b, Bor08p, Bor09c, Bor09v, Bor09-30, Bor09-28, Bor09-31, Bor09-29, Bor09-27, Bor10k, Bor10v, Bor10w, Bor12q, BT13b, Bor13l, Bor13p, Bor16u, Bor16v, Bor16w, Bor16x, Bor16y, BLT17, BKL^{+}93, CJKB92, CG18, CPRZ20, DAK88, DNG^{+}86, DBCB88, DGLV20, EWM86, GRM^{+}97, GC88, KJ86, KC89, KWK^{+}90a, KWK^{+}90b, KWK^{+}90c, LPB01, Mon89, NJS88, NOL86, RSP^{+}93, Rud89, Sch85, SB87, SH87, SZUM86, Stu90, TB00, UVW^{+}21, AR13, AP16, BBKW06, BBC^{+}11b, BTBT88, Bor84a, Bor85c, Bor88k, BL91c, BL91b, Bor92k, Bor92l, Bor92m, BZ94a, BH94a, BH94b]. problems [BL94a, BZ94b, BH95, BZ97, BTZ98, Bor12p, Bor13j, Bor14f, Bor14g, BT14b, BT14a, BT15, Bor15g, Bor15r, BT17, HD07, HLZ14, HLY16, JD13, KJR16, LZ14, Li15, LW18, LW19, MPB16, NWY10, Pea07, PD18, WSdSY15, YW12]. Proceedings [Bor96i, lL09, AAB^{+}88, BBJC97, IMR92, HY14, ABD03, BF06b, CGM95, RZ15]. process [Bor83a, Zăl86]. processes [Bor86a, MTCB98]. processing [BJCW13]. Produce [BBSL20]. Product [BPB99, BB83]. productive [Mic03]. products [RZ15]. Professor [MW16]. Program [BW79a, BW79b, BW81c, BW81b, BW82a, BW82b, BWB97]. programmed [BB11c]. Programmes [Goo20]. Programming [Bor01o, Bor05-33, Bor06-33, BL15, TB80, Bor76a, Bor79a, BW81a, Bor81c, BW81d, Bor83c, Bor83f, BW86, Bor87k, Bor88l, Bor89i, Bor90e, Bor90f, Bor90c, Bor90d, Bor91b, Bor91c, BL92c, BL92d, BBT92, Bor93e, BL93b, Bor94g, Bor95m, Bor95n, BBY12, BBY14, DF05, ZL22]. Programs [CFG^{+}18, Bor79c, Bor80e, BK83, Bor91h, Bor92d]. Progress [BB08b, BB11b, Bor12y, BY12c, BY15]. progressions [Zah06]. Projected [BL17a, BL17b, DF05, LZ14, WM07, HNP10, HLZ15a, HLZ15b, HL15b, HLY16, PD18, ZH06]. Projection [BB95c, BB96b, Bor98n, Bor99w, Bor09v, Bor10c, Bor10d, Bor10k, Bor10v, Bor10w, Bor12q, Bor13p, BST13, DLR20, BB93a, BB94a, BBL97a, BLY13, BLY14, BST15]. projections [BBL94, BB95b, BB97a, BBL97b]. promises [Bor94c, Bor94d, Bor94e, Bor95f, Bor95g, Bor95h, Bor95i, Bor95j, Bor95k, Bor95l, Bor96c]. Promoting [BB12-27]. Proof [Bor02l, Bor02m, Bor05a, Bor07g, Bor07l, Bor07k, BS07, Bor08g, BS08, BB11-31, Bor12-33, Bor16f, Cvi10, GS08, Hd12, Ara07, Ara08, BB08c, BB14j, BB15v, Bor77b, Bor94a, Bor06h, Bor08d, Bor08e, Bor08f, Bor09a, Bor09e, Bor09f, Bor09g, Bor09u, BY12f, Bor14z, Bor16-27, lL09]. Proofs [CS21, Ade13, Gui08, Gui16]. Proper [Bor77c, JN03, Yan94, Zhu91]. properly [Zho12]. Properties [Bor00m, Com18, CPRZ20, BBEM10, BBT98, BBT00, Bor82a, Bor90g, Bor90h, Bor90i, Bor90k, Bor90j, Bor90l, Bor90a, Bor90-40, Bor90-41, Bor90-42, Bor90-43, Bor91d, Bor91r, Bor91s, Bor91t, Bor91u, Bor92a, BB01c, BNSW11, Mar91]. Property [GLR18, HDL21, Las18, AK22, BBL97c, BBL99, Bor82e, Bor88j, BF89c, BJ97, BJ98]. Prophets [BB15r, BBLZ14f, BBLZ15c]. propose [BBLZ14o]. Proposed [BB08f, BB11w]. Prospects [BB05a, Bor09w, Bor09x]. protein [BT14b, BT14a, BT17]. Prototype [BMP05]. Proving [lL09, Hd12]. prox [BBEM10]. prox-regular [BBEM10]. Proximal [BS86, BS87, BI95, BI96, BG87, BGW97, BGW98]. Proximality [Bor06u, Bor07y, Bor08t]. Pseudo [BBLZ14l, BBLZ14s, BCJW13]. Pseudo-mathematics [BBLZ14l, BBLZ14s]. pseudo-random [BCJW13]. pseudoconvex [QR07]. pseudorandom [BB13x]. PSLQ [BB09j, SV20]. psychology [BB13d, Bor09y]. Public [BB14n, Bor03g, BB09i, BB11p, BB15m, Bor12-28]. Publication [Bor97i, Bor98a, BS97b]. Publishing [Bor99x, Bor96d, Bor97j, Bor97o]. pursue [BB10h]. Putnam [Bor77d]. puzzles [Bor15a].

QC [KG04]. QCQP [PD18]. QPQC [Pos13]. Quadratic [Bor89g, Bor89h, BY06, HLZ15b, HDL21, Bor82b, DF05, La 09, NWY09]. quadratically [BB86c]. Quadrature [BB06a, BB08d, Bor06j, Bor06k, Bor06m, Bor06n, BY06]. qualification [BW79b, BW82a, BW82b, BW86]. quality [BB12-44]. Quantitative [Ano15, BBLZ14p, Koh01]. Quantum [CC20a, Cvi10]. Quartically [Bai88, Bai16b, TK97]. Quasi [BL92c]. quasiconvex [BBP03]. Quasidense [Sim18]. quest [BBBP96, BBBP97, BBxxc]. question [BB14z, BB14-27, MR11]. Questions [Bor03-34]. Quick [BB11x]. Quinn [BBC09].

R [Bor11-38, Odl11, TSB13]. Rachford [AB12, ABT13a, AB13, ABT13b, ABT13c, ABT14a, ABT14b, ABT15, ABT16, AC18, BS10b, BS10c, BS10d, Bor10i, Bor10j, BS11b, Bor11r, Bor11s, BT13a, BT13b, Bor13j, Bor13r, BT14c, Bor14f, Bor14g, BT15, Bor15g, Bor15r, BG16a, BLS^{+}16, BLS^{+}17, BLS^{+}18, BG18b, Gil18]. radicals [BdB91]. radiometric [BB10g]. Radon [GLR18]. Rainfall [Bor13k, BHP14, Bor13q, PHBH12, PHBH13, PHB13, PHB14]. Ramanujan [BB96d, AB15, AAB12, AAB^{+}88, BBB97a, BBG95b, BR01, Bor85b, Bor86f, BB87a, Bor87g, Bor87f, BB87b, Bor87l, BB88d, BB88f, BB89a, Bor89f, BBB89, Bor90-29, Bor90-30, Bor90-31, Bor90-32, Bor90-33, Bor90-34, Bor90-35, Bor90-36, Bor90-37, Bor90-38, Bor90-39, Bor91j, Bor91k, Bor91i, Bor91l, Bor91m, Bor91o, Bor91p, Bor91q, Bor92e, Bor92f, Bor92i, BB93d, Bor93m, BBG94c, BB96d, BBB97c, BBB00b, BB01f, Bor03d, Bor03e, Bor03f, Bor04-30, Bor04-29, Bor04-28, BCF04, BC04a, BBB04b, BL05, Bor05j, Bor06i, BL08, Bor10x, Bor10z, Bor10-27, Bor11-29, BBGW11, Bor11-30, Bor11-32, Bor12x, BBB16, Bor16d, BB16u, Liu00, Lor08, BB91d]. Ramanujan-type [BB87a, BB88d, BL08]. Ramble [Bor10-28, Bor10-29, Bor11-33]. Rand [BBC09]. Random [BB13c, BNSW10, Bor10-28, Bor10-29, Bor11-33, BSW13, CC20b, Gan14, BB13b, BB13-40, BB95b, BB97a, BJCW13, BCJW13, BL05, Bor10e, BSWZ11, BNSW11, Bor12b, BSWZ12, BR13a, BSV15, BS16b, BSV16, BS16a]. Randomness [BBBR16, BBBR17, Gan17]. Range [Bor04p, Bor05l, Bor05m, Bor05n, Bor05o, Bor05p, Bor06l, R^{+}05, BW81c, BFKL00, BFKL01, BFL02, Bor03x, Bor03y, Bor06-28]. Ranking [BBSL17b, BBSL18, BBSL17a]. rapid [BBP97]. rapidly [AL10, BB83]. Rate [BLT17, BLY13, BLY14, BLT15, BLT16, HL15b]. rating [BB11w]. Ratio [Ade14a]. Rational [BZ87, BB87b, BZ92, BB98c, BB98d]. Reactions [BB14q, BB14p]. Real [ABBB13, Bai91, BCF04, Bor13-28, Bor13-29, Lor90, BB13j, BFG87, BB90d, BB91b, Bor04-30, Bor10z, Bor14x, Bor14y, Bor16q, Bor16r, Bor16s]. Real-Parameter [BCF04]. Realistic [BST13, BST15]. Reality [Bor05-40, BB12u, BB12n, BB13p]. Really [BB14i, BB11-28, BB14h, BBLZ14b]. rearrangement [BLZ99, BLZ01]. Reasoning [Bor93c, Bor93d, BBG03, Bor03z, Bor03-27, Bor03-28, Bor03-29, BB04b, Bor04-27, Bor04w, Bor04x, Bor04y, Bor04z, Bor06-29, Bor10a, HF05, Hoa05, Zei05]. Receive [BE16, Bai16a]. recipients [BB14e]. Reconstruction [Bor09-27, Bor92n, Bor93k, BLN94a, BLN95, BLLN95, BLN96, LLC^{+}95, MTCB98]. reconstructions [MTCB99]. Recurrence [BS08, BBCM07b]. Recurrences [BBS14a]. Recursion [BS07]. Recursions [BB06b]. Reduced [BB84e]. reduction [BW81d, Jal24]. Refined [BBFG00, BBFG01, War03]. Reflection [BST13, BT14b, BT14a, Bor16p, BT17, BST15, Bor15r]. Reflections [BB09c]. Reflexive [BV94b, BBWY11b, BBWY12b, Bor93a, BZ94a, BTZ97, BE08, BV10a, Bor13g, Bor13h, Bor13i]. reflexivity [BB90a]. refute [BB12w]. region [ZSZ16]. regional [JY12]. registration [HYG09]. Regular [Bor84d, BBEM10, Bor86c]. regularity [BBL97c, BB98a, BBT98, BB99b, BBL99, BBT00, BZ88, BF94b, BZ95, BZ96, BLT15, BLT16]. Regularization [BL11, HLZ15b, ZL22]. regularizations [BV95a]. Regularized [WSL16, MTCB99, XWQ14]. Regularizing [BW81b]. Regulatory [BB15x, BBLZ15f]. Reich [Koh01]. Reinhart [BB13-31, BB13-32]. Related [Bor02b, BHL16b, BHL16a, BS84b, BB95f, BB01c, BSZ13, BHL17]. relating [BW95b, BW97b]. Relation [Bor09p, Bor09q, Bor10r, BL97, BL00b, BY12b, BY13b]. Relations [BB09j, Bor80b, Bor02a, BS15b, SV20, BWY10, BBWY11e, Bor81b, Bor81d, Bor87a, BBCM07b]. relationships [BL91b, BV93a, BV94c]. relative [BB09i, BB13i, BB13-34, BL92c, BG01, BG03a]. Relaxed [DLR20, RS02]. Reliability [BB13-32]. Reliable [BBSL20, BB10g, BB14x]. religious [BB09d]. Remark [Gil18, Osb05]. remarkable [BB11y, BB90b, BB01c]. Remarks [BG16c, BEO77, Bor81a, BG15c]. remembrance [Bai17e]. Remote [BLM^{+}07, BM07b, Bor09w, Bor09x, BBJ12]. Renaissance [Bai21]. renorming [BF93d, BV95c, BV95d]. replace [BB16s]. replication [Gui17]. Reply [Gan17]. Report [BBC^{+}14a, Bai17e, JWDS^{+}14, BBLZ14j, BBL^{+}13]. reported [BB14x]. reporting [BB12f]. reports [Mic03]. representation [BMS97, BMS99a]. representations [BC98a, BC00]. Representative [EB08]. Reproducibility [BBL^{+}13, BBS16b, BBBR16, BBBR17, Gan17, JWDS^{+}14, BB13-32, JWDS^{+}14]. Reproducible [BB13-35, BB13-36, BBL^{+}13, SBB13, BBLZ15e, Bor13-30, Bor15m]. Res [Zăl86]. Research [BB13s, Bor09o, Bor12n, Cam16, PR92, SBB13, BB09d, BB10h, BB13l, BB13r, BBLZ15e, Bor95t, Bor95u, Bor97x, Bor07q, Bor13d, Bor13a, Bor13c, Bor14a, Bor16h, RZ15]. researchers [BWB97]. Researching [Bor11g, Bor11-37]. Reseñas [Bou06]. Resolution [BBC09]. Resources [Bor98j]. Respect [Bor77c, Bor74]. Response [BaO12]. restoration [WM07]. Result [Mil89, BB11x, FK00, Mil90]. Results [ABT13b, ABT14b, BL93c, BLLN94, Bor96f, Bor96g, Bor96h, Bor07-28, Bor07-29, Bor07-30, Bor07-31, BB14-32, CG18, ABT13c, BB13n, BB13-42, BB13-47, BB14s, BLLN95, BW95b, BBB96b, BBB96c, BBB97d, BW97b, BK01, Bor07-27, Bor12j, Bor12k, BY12d, BY14b, Hon85]. retires [Jac09]. retraction [Bor15c]. Retro [BM07a]. Retro-enhancement [BM07a]. Retrospective [Bor08s]. Reuben [BO11b]. Review [Abb00, Ask88, Bai91, BB09c, BBLZ14m, Ban10, Ber88, Bor90b, Bor92b, BB93g, BC96, Bor05g, Bor06o, Bor11-38, BS14a, Cas99, Coh15, HF05, Hoa05, How14, Lor90, Lor09, Odl11, Rob06, Sha05, Wim88, Bai17d, BB91d, Bor09b, BO11b, BS14b, Tod03]. Reviews [Bou06, Zei05]. Reviews/Reseñas [Bou06]. Revisited [BLM96, BLM97, Bor08s, AAB^{+}88, BCM09, BY12f, KPS16]. Revivals [Bor96j]. Revolution [R^{+}05]. Richard [BB12-38]. Riemann [BS17, BB96c, BBC98, BBC00b, BB05c, Bor07g, BBS15b, BBS20]. Riemannian [IP17, IP18]. Risk [Roc20, BB09i, BB11c, BB11p, BB13i, BB13-34, Cam16]. risky [BBLZ15g]. Road [CC20a]. Robert [BB91d, TSB13]. robot [BBLZ16a]. Robust [ANR18, CFG^{+}18, DGLV20]. Rocha [Ban10]. Rock [Bor14v, Bor14w]. Rockafellar [Ano15, Bor11-38, BBB^{+}07]. Rodrigues [Ban10]. Rogoff [BB13-31, BB13-32]. Roland [Sha05, Zei05]. Role [Bor02l, Bor02m, Bor05a, Bor16f]. Rome [BB09k, BB11z]. Romney [Bor12a]. root [BB13p]. Roots [BB12r, BB16d, BB11h, BR84, BS14a, BS14b]. Rossi [BB15u, BB16i]. Rotund [BGV02]. rotundity [BL94b]. routes [Ade11]. Rule [BY06, BM96a, BM98a]. Rules [CPRZ20, BB12-30, BBLZ14o, BM97d]. Ryabova [DP18].

S [Bou06, Tod03]. S. [Bor91p, Bor91q, Bor93m, Bor81a]. Sad [BB10j]. saddle [HD07]. Salamin [Borxx]. salt [BF06a]. Same [BW95a, BB96a, BB99c, BW97a]. sample [BBLZ14s, KJR16]. sampler [BG15a, BG18a]. San [BC96]. Sandwich [Bor80b, BT92, Bor98o, Bor81d]. sandwiched [BF98, BF01]. Sank [Bor11-36, BBS12]. Santalo [BBFG00, BBFG01]. Sapiens [Thé16]. Satire [Bor07c]. say [BB12-50]. Scale [JWDS^{+}14, DF05, LW18, LW19, Ray97, WM07, XH08, ZSZ16]. scales [PHBH13]. scaling [WSdSY15]. scary [BBLZ14n]. sceptics [BB12d]. Schaible [Bor90b]. Scheme [BT13a, BT14c]. Schemes [BB08d, Bor06j, Bor06k]. scholars [Mic03]. School [BB12k, BWB97]. Science [BB13-38, BB13-44, BB15l, BBBR16, BBBR17, Bor95t, Bor95u, Gan17, PR92, RZ15, Sel16, SBB13, BB10i, BB10j, BB11i, BB12f, BB12-39, BB12-34, BB12-35, BB12-44, BB13n, BB13m, BB13w, BB13u, BB13-29, BB13-30, BB13-35, BB13-36, BB13-33, BB13-37, BB13-39, BB13-43, BB14o, BB14-29, BB15w, BBC^{+}11b, Bor96k, Bor97x, Bor98r, Bor14a, Bor15c, BB09c]. Sciences [Bor98e, Bor07o, Bor13m, Bor13n, SV14]. Scientific [BB13s, BB13-33, BB13-34, BBS16b, Bor04i, BB09d, BB10c, BB11s, BB11f, BB12-28, BB12-30, BB13l, BB13r, BB15m]. scientist [BB09h]. Scientists [BB12-45, BB12b, BB15q, BB16g, BB16h, BWB97]. SCIHTBB [XC11]. Scissors [Bor14v, Bor14w]. score [BB12-45]. scores [BB12o, BB13-27]. Scribner [BB91d]. search [FN15, IP17, IP18, YW12]. Searching [BB96c, BB05c]. Seasonal [BHP14, Bor13q, PHB13, PHB14]. SEC [BBLZ14o]. Second [BN94, EB08, ABD03, Bor92g, Bor92h, Bor93f, Bor93g, BF93b]. second-order [BF93b]. Security [BB15s, BBLZ15b]. Seeing [Bor12z, Bor13-28, Bor13-29, Bor13u, Bor13v, Bor13w, Bor13x, Bor13y, Bor13z, Bor13-27, Bor14x, Bor14y, Bor16q, Bor16r, Bor16s]. Seeking [BB15k, BB15j]. select [BBGPxx]. selected [BB12z, BB10l]. Selection [Bor12-30, Bor12-31]. Self [Gui17, Ara07, Ara08]. self-contained [Ara07, Ara08]. Self-replication [Gui17]. sell [BB12e]. Semi [Bor83f, Bor89i, Bor79a, Bor81c, Bor83c, BLY13]. semi-algebraic [BLY13]. Semi-finite [Bor89i]. Semi-infinite [Bor83f, Bor79a, Bor81c, Bor83c]. Semialgebraic [CFG^{+}18, BLY14]. semicontinuity [BLZ99, BLZ01]. Semicontinuous [BTZ95, Bor90g, Bor90h, Bor90k, Bor90l, Bor90-40, Bor90-41, Bor90-42, Bor90-43, Bor91d, Bor91r, Bor91s, Bor91t, Bor91u, Bor92a, BT92, BTZ98]. Semigroups [Bor16j, Bor16k, BG15a, BG18a]. Seminar [BBLZ14p, BLM^{+}07, Bor07d, BM07b, BJL^{+}08, BBJ12]. Semiotic [BB09k, BB11z]. Semismooth [Las18]. sense [BBGP95c, BBGP96, JN03]. Sensing [BL17a, BL17b, Bor09c, Bor10h, Bor11p, QYX14, XWQ14]. Sensitivity [BTZ97]. Seoul [HY14]. Separable [BM97f, BM00, Bor95a, Bor95b, Bor02d, Bor02e, BBL04, PD18]. separably [BK83]. separably-infinite [BK83]. separate [BB00a, BB01b]. separation [BB84f, BJ97, BJ98]. September [Bai17a, BBB^{+}20, SBW84]. Sequence [BSxx, BL92a]. sequences [BL93a, Bor98d, Bor15d, BC96]. Sequential [BV9x, BF93c, BF95b]. Sequentially [BV94b, Bor93a]. Ser. [BZ02a]. Series [Ber88, BB86a, BB90c, BB92a, Bor01g, Bor05f, BB07c, BBJ12, BB15c, BB87b, BB88d, BB93d, Bor93o, BB95f, Bor02h, Bor02i, BC02, BC03, BC04b, BCP05, BG05, Bor07e, Liu01, Nim15, XY12]. Serious [Bor07c, BB13i]. Serving [Zei05, BBB03]. Session [AMM10, Bea13]. Set [BBS13a, BB13-35, BB13-36, Bor13-30, Bor15m, BZ88, BV95c, BV95d, Zho12, Bor92b]. set-valued [BZ88, Zho12, Bor92b]. Sets [BB14a, BB93b, BT84, Bor06u, Kru18, Moo18, RZ18, BBCR13, BB93a, BB94a, BBL94, BBL97a, BBL97b, Bor81a, BT85, BS86, Bor87m, BS87, BFK91, BL93a, BV94a, BF94b, BF95c, Bor95a, Bor95b, BV96a, BV96b, BM96b, BM98b, BLM99, BLM00, BV04, Bor07y, Bor08t, Bor12g, Bor12h, BLY13, BLY14]. Setting [BBL^{+}13, Bor07z, Gil18, SBB13]. Seven [Bor13-31]. Several [BB86a, Wei15]. Shafrir [Koh01]. Shannon [BH95]. shape [SZ14]. Share [BW95a, BW97a]. Short [BM97b, Bor10-29, Bor11f, Bor11-34, Bor11-35, Bor11-33, Bor15o, Bor15p, Bor15q, SZ20, BSWZ11, BNSW11, Bor12b, BSWZ12, BS13, Bor14t, BSV15, Bor15n, BSV16, Bor16e]. show [BB13-27]. Shrum [Bor93n]. Shu [BB95e, lL09]. SIAM [Bor05g, BB08f, Bor09z]. Siegfried [Bor90b]. signal [Bor90e, Bor90f]. significance [BB14x]. Significance* [Alt20]. Silence [Sol15]. Silicon [Zei05]. Simon [BC96, BBJC97, Bor06a]. Simple [AW97, BW86, BLS^{+}16, ZSZ16]. simplification [BBK14]. Simpsons [BB13z]. Simulate [Bor13k]. simulated [PHBH12, PHBH13]. Simulation [BHP14, Bor13q, PHB13, PHB14]. Sinc [BBL10, Bor11-36, BBS12, BB14-32, BBB08, Bor00r, Bor01p, Bor01q, Bor01r, BBM01, BB01c, BBM02]. sine [BS11a, BBSW11, Bor11f, BS11d, BS11e, BS12b, BBSW12, BS13]. Single [Bor04-31, BZ88]. single-valued [BZ88]. Singly [CPRZ20]. singular [BB91d]. Sinh [BY06]. Six [BBJ12]. Size [BB88a, SI16, KJR16, LW18, LW19, LY21, XC11]. Skepticism [BB13-44, BB13-43]. skews [BR14b]. sky [BB93g, Tre13]. Slice [BB93b, BV93a, BV93b, BV94c, BV94d]. Slices [Bor04l, Bor04m, Bor06r]. Sloane [BC96]. sloppy [BB13-33]. Small [BZ87, HMM20, BFK91, BZ92]. Smart [BB12l, BB12-46, BB13-37, BB13-38, Bor12-27]. SMC [Bou06]. smell [BB13-40, BR13a]. Smooth [ANR18, BI95, Bor99u, Bor00v, YS00, Bor86g, BP87, Bor87h, Bor87i, Bor87j, Bor90m, Bor90n, BF93d, Bor94h, Bor94i, Bor94j, Bor94k, BZ95, BM96a, BM96b, BI96, BZ96, BM97c, BM97d, Bor97p, BM98a, BM98b, Bor98l, Bor98m, BTZ99, Bor99t, BFL02, LS00, LLS11, BM07c]. Smoothing [HLY16, Li15]. smoothness [BBC00a, BBC01]. soaring [BB15w]. sobering [BBLZ14q]. SOC [ZL22]. Social [BB15s, Bor15c, BBLZ15b]. socially [BB11i, BB12-34, BB12-35]. Society [BB16c, Ber88, BB11k, CW16]. softcover [Bor05g, Bor06o]. Software [Bai91, HY14, Bor08q]. Sokal [BB13d]. Solution [BB07b, BB07a, BB09l, BB10k, BB12-47, BBS14a, Bor11-38, BB12-53, MR96, Zho12]. Solutions [AJB86, AI18, ANO^{+}83, AJ86, BB09m, BL87, BSW82, BSZ^{+}83, BB85, Bor85a, BN86, Bor93l, BB93c, Bor96j, BDT96, BBS^{+}97, BPB99, BKL^{+}93, CJKB92, DAK88, DNG^{+}86, DBCB88, EWM86, GRM^{+}97, GC88, KJ86, KC89, KWK^{+}90a, KWK^{+}90b, KWK^{+}90c, LPB01, Mon89, NJS88, NOL86, RSP^{+}93, Rud89, Sch85, SB87, SH87, SZUM86, Stu90, TB00, UVW^{+}21, BZ95, BZ96, Yan94]. solved [BB16o]. Solving [AC18, BB95c, BB96b, CPRZ20, AR13, AP16, Bor92k, Bor92l, Bor92m, LW18, LW19]. Some [BEO77, Bor81a, BSW82, Bor85b, BB92b, Bor93o, BBG94c, BB94b, Bor98p, BMS99b, Bor99y, Bor99z, Bor99-27, Bor00u, BK01, BB01c, Bor03-30, Bor07-27, Bor07-28, Bor07-29, Bor07-30, Bor08u, BNSW11, BY12d, BY14b, BG15c, BG16c, Liu01, Lup02, TB80, BB95f, Bor96f, Bor96g, Bor96h, Bor05j, Bor06i, BB11-31, Gui08, Liu00]. sorry [BB13i]. SOS [CFG^{+}18]. SOS-Convex [CFG^{+}18]. Soul [BB15i, BB15h]. sound [BB12o]. Source [Abb00, BBB03, Rob06, BBB97b, BBB00a, BBB04a]. sourcebook [BB16l]. sources [Cam16]. South [HY14]. Space [BB12u, BB12n, BB16m, BB16n, BGM18, Bor78a, BM07d, Bor10c, Bor10d, Bor13e, Bor13f, WG17, BB17, BBL94, BB95a, BBL97b, BBWY11a, BBWY12a, Bor84b, BS86, BFG87, Bor87m, BS87, BG87, BZ94a, BF94b, Bor02d, Bor02e, BBL04, BM07c, Bor07x, Bor13g, Bor13h, Bor13i]. Spaces [BV94b, BFV94b, BI95, BBS10, BBEM10, BBC00a, BBC01, BBWY11b, BBWY12b, Bor81a, BS84a, BF89b, Bor91d, Bor92g, Bor92h, BRS92, Bor92a, Bor93a, BL93a, BV93b, Bor93f, Bor93g, BV94a, BFV94c, BFV94a, BV94d, Bor94h, Bor94i, Bor94j, Bor94k, Bor94l, BN94, BZ94b, Bor95a, Bor95b, Bor95s, BZ95, BV96a, BV96b, BI96, BZ96, BFV97, BV97, BJ97, BTZ97, BZ97, BJ98, BLM99, BJSM00, BLM00, BV00b, BV01, BG01, BJSM02, BG03a, BE08, BG09, BGHV09, BV10a, BG15b, BG16b, La 09, QR07]. Sparsity [XC11]. Spatio [CZX21]. Spatio-Temporal [CZX21]. Special [AMM10, AHLC^{+}17b, BC21, BBBL98c, BBBL99, BBFG00, BBBL01, Bor11-29, BS11d, BS11e, GI01, AAB12, Bor83c, Bor83f, BBFG01, Bor12t, BL16]. SPECT [BNCB99, BS95, BS97a, Bor02r, LLC^{+}95]. spectra [BMN98, BMN00]. Spectral [Bor87k, BBT92, BLLN94, CPRZ20, BTBT88, Bor90c, Bor90d, Bor91b, Bor91c, BRLZ99, BLZ99, BRLZ00, BLZ01]. spent [Bor10-30]. Sphere [BB16o, BB14j, BKW02, CKM^{+}16, Via16]. Spheres [BLS^{+}17, BLS^{+}18, BLS^{+}16]. spherical [AX20]. spin [BBCM07a]. Spline [SBW84]. sports [BB13h]. Springer [Bor11-38, Tod03]. Square [BB12r, BB16d, BB11h, BRxx]. Squares [Bor01g, Bor02h, Bor02i, BC02, BC03, BC04b]. Srinivasa [BB96d, Bor12x]. St [IEE08]. Stability [AI18, Bor84d, Bor86c, BM09, BM10, BW81a, BS95, BS97a, MTCB99]. stabilized [LY21]. Stable [DGLV20]. Stage [Bor07z]. Stan [Bor05g]. Standing [JWDS^{+}14]. Starshape [BEO76, BEO77, Bor78c]. state [BB10j]. Static [BBSZ87, BBSZ88]. Statistical [Alt20, BSW82]. Statistically [Gan14]. Statistics [BB09a, BB15l, BB09e, BB11f, BB15w]. staunch [BW05b]. steepest [RS02]. Steiner [BO11b]. Step [BB88a, BSW13, SI16, Bor10e, LW18, LW19, LY21, SD15, XC11]. step-size [LY21]. Stephen [BB10e]. steplength [Pos13, Ray93, XSW12]. stepsize [DABY15, MP18]. Still [BB14c, Bor01e, Bor02s, Bor02t, BB13y, BB14w, BB14-28]. Stochastic [BLN94b, SD15, HLZ14, HLY16, KJR16, LLS11, LZ14, Li15, LY21]. Stock [BBL16a, BBL16c, BBLZ14i, BBLZ16a]. stocks [BBLZ13d]. Stoneham [BB12-40]. Stop [Dev20, BB12f]. Story [Bor94f, Bor09z, Bor90o, Bor90p]. Strange [BB90c, BB92a]. Strategies [BBLZ13a, BBC98, BBC00b]. Strategy [dPB21]. STRAW [BB11v]. Street [BB97d]. strict [BBC00a, BBC01]. strictly [BM95, Bor95d, NWY09, PD18]. Strogatz [BBC09]. Strong [BBL97c, BBL99, BL94b, BBT98, BBT00, Bor80e, Bor12x]. strongly [Bor78b]. Structure [BY12e, BY13c, BB16b]. Students [PL20, BWB97]. Studies [SV14, BWB97]. Study [BBBR16, BBBR17, Ber88, BB87d, Bor05f, Bor11f, Bor11-27, Bor11-28, Gan17, lL09, SBW84, Wim88, BB98b, Bor05g, Hd12]. Stuff [Bor00j, Bor00k]. Stupid [BB13-39]. Style [Bor11-29]. Subderivatives [Bor88m, Bor88n, BMW95, BZ95, BZ96, BGW97, BMW97, BGW98]. Subdifferentiability [BW99, BW01, Fab89, BP87]. Subdifferential [BW95a, Las18, BW97a, BM97e, BM97f, BZ98, BZ99c, BM00, BZ02a, BZ02b, BS10a]. Subdifferentials [BFG03, BBEM10, BW98b, BMW99a, BMW99b, BMW99c, BW00, BMW01, BVW01, BGV02, BW03, BVW03, BW05b]. Subgradient [BMS97, BMS99a, Bor09c, Bor10h, Bor11p]. Subgradients [Bor84e, Bor82d, Bor82c, BFG87, Bor91a, BF94a, Bor94b, BF95a, BBW96]. Subject [CPRZ20]. Subspace [XH08, LL13]. Substance [DD15]. Substitutions [BCM20]. success [Cam16]. sufficiency [Bor76b]. sufficient [Bor82b, BZ88]. suggest [Cam16]. Sum [BB18, BY13a, BY14c, BB16a, BBB06b, BY12b, BY13b]. Summary [BB06a, BC04b]. summation [BCM09]. Sums [BB94b, BBP95, BG95b, Bor96f, Bor96g, Bor96h, BBK00a, Bor01g, BB05g, Bor06-31, Bor12r, BGM^{+}13, BBS20, BBG93a, BBG94a, BB13g, BBCZ13, BBC14b, BB15a, BB16b, BBB08, BBT85, BBS89, BBG94b, BBG95c, Bor95e, BBB96b, BBB96c, BG96b, BBB97d, Bor97f, Bor97m, BBP98, Bor98f, BBK00b, BBK01, Bor02h, Bor02i, BC02, BC03, BC04b, Bor06-32, Bor07x, BZB08, Bor12e, Bor12f, BBS13b, BBS14b, BBS15b, GG07]. sunlight [BB13-40, BR13a]. Super [BZ91, BZ93]. supercomputers [BBG95a]. superrelaxation [Pos13]. Supplement [BBB03]. support [BV94a, BV96a, BV96b]. supportability [Bor79g]. Supportless [BT84, BT85]. Supremacy [CC20a]. Surmise [DD15, Bor02g]. Surprise [Bor99q, Bor99r, Bor99s, BBM99, Bor00p, Bor00q, BBM00, Bor04v, Bor04-32, Bor05-32, Bor09-27, Bor13-32, Bor09n]. Surprising [BBB08]. Survey [BL93c, BV9x, Bor90a, Bor90-40, Bor90-41, Bor90-42, Bor90-43, Bor91r, Bor91s, Bor91t, Bor91u, Bor94l, Bor95s, BW95b, BV95c, BV95d, BW97b, BZ99c, BZ02a, BZ02b]. Surveys [SV14, BR01]. SVM [SD15]. Swedroe [Swe17]. Swiss [BBLZ15g]. Sylvester [Bor79f]. Symbolic [Ade11, Bor98h, Bor00t, Bor05-41, BH06, Bor09t, BH09, GI01, LLT18, BBK14, Bor97h, Bor98q]. Symbolically [BB96c, Bor97q, Bor97v, Bor97w, BB05c]. Symbols [Bor09t]. symmetric [DABY15, JD13]. Symmetry [Bor16z, BBS20, Bor13-34, Bor13-35, Bor13-33, BZ13]. Symposium [IEE08, CGM95]. symptom [BB13-28]. system [BB11w]. Systems [ANR18, BC18b, Bor84d, LY18, PR92, Bea13, Bor86c, Bor92n, Bor93b, Bor93k, BS95, BS97a, BR16, DABY15].

tails [BCP05, BC10]. tales [BBLZ13f]. Talk [Bor93n, Bor07v, Bor08l, Bor08m, Bor10u, Bor11x, Bor11y, Bor11z, Bor11-29, Bor16n, Bor16t, Bor89a]. Talking [BB12-48, Bor97r, Bor97s, Bor97t, Bor97u, Bor98b, Bor99-28, Bor10-30, Bor12-28]. talks [BB14e]. Tangency [Bor99w]. Tangent [BO76, Bor78c, Bor78a, AL10, BB84f]. Tangential [BS85]. Tanh [BY06]. Taylor [Nim15]. teach [BB10a, BBLZ13h]. Teacher [Goo20, Mic03]. teachers [BB12-49, BWB97]. Teaching [AD20, Bor11g, Bor11-37]. Technical [Bor16t]. Technion [IMR92]. Techniques [BZ05, Bor94n, BZ99a, BZ99b, GS02]. technological [BB12-44]. Technologies [JJ20, PL20]. Technology [Bor98e, Bor99e, Bor99f, Bor99d, BS99c, Bor00n, Bor07f, Sel16, BS99a]. Tegmark [BB14r]. Telco [Bor10-30]. telelearning [Bor00w]. Telstra [Bor10-30]. Temporal [CZX21]. Ten [BBKW06, Bor05b, Bor09-30, Bor09-28, Bor09-31, Bor09-29, Bor09-27]. Tensor [CZX21]. tentative [BB12-34, BB12-35]. term [BBLZ14j, BBLZ14i]. Termination [HDL21]. Terms [BC18b]. ternary [Ade10]. Terry [Ano15]. Tertiary [Bor11g, Bor11-37]. test [BB12o, BB12-36, BB13-27, BB12j]. Testing [Alt20, BBLZ13a, BBLZ14r, BB13h]. tests [BB11x]. Texas [BB13-29, BB13-30]. textbook [BB13-29, BB13-30]. Texts [Ber88]. th [BB84d, Cra12]. their [BBLZ15a, Bor88m, Bor88n, Bor89d, Bor95o, Bor95p, Bor14e, RZ15]. themselves [BB10a]. Theorem [BBWY11a, Bor80b, GN16, TB80, dPB21, Ara07, Ara08, BB13e, BBWY12a, BO11a, Bor79f, Bor80e, Bor81e, Bor81d, BZ86, Bor88g, Bor88h, Bor88i, Bor89c, Bor90m, Bor90n, BW98a, BD03, Bor14z, Bor16-27, Dev9x, Koh01, MW12, OBB^{+}96, Rei02, BB13f, Bor79b, Bor13h]. théorème [Dev9x]. Theorems [Bor99-27, Bor00u, Bor12-30, Bor12-31, Bor14h, Bor14i, Bor14j, Bor14k, Bor14l, Bor14m, Bor14n, Bor15h, Bor16-28, BB98a, BB99b, BS17, Bor77b, Bor79a, Bor81c, Bor85c, Bor87m, BT92, BG95a, Bor98o, BY13a, BY14c]. Theoretical [BaO12]. Theories [BB09g, BBG95b]. Theory [AHLC^{+}17a, AHLC^{+}17b, BB15i, Ber88, BB87d, BZ02a, BM07d, Bor09d, Bor12e, Bor12f, BR12, BY12c, Bor12-30, BR13b, Bor16u, Bor16v, Bor16w, Bor16x, Bor16y, Bou06, DLR20, GI01, HMM20, SBW84, Tod03, Wim88, BBC10, BB13-42, BB13-47, BBC14b, BB15a, BB15h, Bor84a, BL92c, Bor94m, Bor95v, BB98b, BM07c, BY12e, BSZ13, BY13c, BY15, Cvi10, KG04, BS86]. Théra [Bor17b]. there [BB15e, BB12-53, Bor14a]. Theta [Hir17, AB15, AAW06, Bor87l, HGB93, LL01, Liu00, XY12]. Theta-Function [Hir17]. Things [Bor13-28, Bor13-29, BB11f, Bor12z, Bor13u, Bor13v, Bor13w, Bor13x, Bor13y, Bor13z, Bor13-27, Bor14x, Bor14y, Bor16q, Bor16r, Bor16s]. think [BB12-51]. Thinking [BaO12, BB12-52, BB93g, Bor94o]. Third [BBB03]. Thirty [BB05d, BB06c, Bor10-31]. Thirty-two [BB05d, BB06c]. Thompson [Bor07-27]. thousand [BB12-29]. thousand-digit [BB12-29]. threaten [BB12-44]. threatens [BB12-41, BB13-29, BB13-30]. threats [BB10i]. Three [Bor93p, Bor97v, Bor97w, Bor98q, Bor03-34, Bor07-31, BSW13, BB14e, BB93d]. Three-Step [BSW13]. Thresholding [WSL16, XC11]. Tilting [BB14y]. Time [WG17, BB17, PHBH13]. time-scales [PHBH13]. times [BBLZ16c, Bor05b]. Timothy [Bor09b]. Tipsy [BB13-40, BR13a]. TMA [BZ02b]. Together [JWDS^{+}14]. tomographic [MTCB99]. tomography [MTCB98]. Tony [Bor15d]. Tool [AD20, BBLZ14k, BWB97]. Tools [Bor00v, BMPR02, Bor05-42, Bor05-43, Bor05-44, Bor06d, Bor11g, Bor11-37, MTB16, BB15b, BB15o, BBS^{+}15a, BC98b, BC99, Bor05-45]. topics [BS84b]. Topological [BG16c, BG15c]. topology [Pea07]. Tornheim [BBC14b, BB15a, BB16a, BB16b, BB18, Bor12r, BBB15, Bor12e, Bor12f, BD18, Dil21]. Total [Jal24]. tottering [BB13-40, BR13a]. Tough [BBLZ16c]. towers [BBLZ13g]. Tractable [CFG^{+}18]. Trademarked [BB14f, BB14g]. trademarking [BB14b]. trading [BBLZ14o]. Traffic [CZX21]. train [Bor15c]. transform [War01]. transitivity [Hon85]. Transversality [Kru18]. Treasury [Fer91]. treated [Bor84a]. tree [BB15e]. trenches [BS97b, Bor97o, Bor06-36]. Tribute [BB13f, BB13e]. trigger [BB12v]. triggers [BB12i]. Trigonometric [BB94b, LPB01]. trilogarithm [Ade10]. trinomial [War03]. Triple [BG95b, CZX21, BG96b]. Troubles [BB13-41]. troubling [BB14b]. trust [ZSZ16]. trust-region [ZSZ16]. trustworthiness [Fab89]. Tsallis [ABBS12]. tuned [BB14-29]. Turing [BB12-36, BB12j]. turn [Bor11e]. Turns [BB15k, BB15-30, BB15j, BB15-29]. Tutorial [BM97b, Bor92j]. twenty [BBxxc]. twenty-two [BBxxc]. Two [BB13-42, BBLZ13f, BB88a, Bor79f, BN84, BB05g, Bor10-31, Bor10-30, Bor15r, HDL21, AAW06, BBLZ13g, BB13-47, BB93a, BB94a, BS97b, Bor97o, BBxxc, BB05d, Bor06-32, BB06c, Bor07x, Cam16]. two-dimensional [AAW06]. Two-Point [BB88a]. Type [Ade14a, Ade14b, Bor01o, BML18, AL10, Ade10, Ade11, Ade12, Ade13, BB96a, BB99c, BBWY11c, BBWY11e, BBWY12c, BB87a, BB88d, Bor91h, Bor92d, BB93d, Bor93e, BH94a, BH94b, BV00b, BV01, BBG04b, Bor05f, BE08, BL08, BEY11, BY12a, BY12f, BY13a, BY14a, BY14c, Gui16, HLZ14, HL15a, Nim15, Wei15, ZS12, Zha13, ZZ14]. typical [BW99, BW01]. Tyrrell [Bor11-38].

U.S. [BB12-42, BB12-43, BB12-51, BB12-52]. UK [BF06b, BB13-27]. ultraproducts [BS15a]. Unbounded [RZ18]. Uncertain [DGLV20, BB12c]. unconsciously [BB10h]. Unconstrained [SI16, AP16, DHSZ06, MP18, NWY10, NFB17a, Ray97, WSdSY15, XSW12]. uncovers [Cam16]. Underdetermined [BL94a, BGL93]. Undergraduate [BS99d, Bor00s, BS99b, BS00]. underscores [BBLZ14j]. Understand [BB15s, BBLZ15b]. Understanding [WG17]. uneven [BB12-49]. Unexpected [BB16p]. Unholy [BB13-44, BB13-43]. unified [Bor77a]. Uniform [BGM18, BH94a, BH94b, BC09, Bor10-29, Bor11-33, BV95b, BV96c, BSWZ11, BSWZ12, BSV15, BSV16]. Uniformly [BGHV09, BV12]. Union [Bor01n, Bor01m, Bor02n]. units [BJCW13]. Universe [Bor11-31, BB14-29, BB14r]. University [AAB^{+}88, BB93g, BBJC97, Bor06o, Bor09b, BO11b, BS14a, IEE08, KG04, SBW84, BWB97]. Unknown [Bor02j, Bor02k]. Unleashed [AH01]. unlimited [ES01]. Unscientific [BB09n]. Unsolved [BB87c]. unsymmetric [DLL05]. untitled [Bor08v, Bor10-32, Bor12-29, Bor15s]. Update [BB15g, BB15f, SD15]. Upon [BB14c, BB13y, BB14w]. Upper [CPRZ20, Las18]. Urbana [AAB^{+}88]. Urbana-Champaign [AAB^{+}88]. US$29.95 [BO11b]. US$57.00 [Bor05g]. USA [Bor05g, BB13-27]. uscos [BFK91, BK04]. Use [Bor12-30, Bor12-31, Bor00w]. used [BB10g]. useful [Bor85b]. User [Bor06o]. uses [BWB97]. Using [Bai88, BLLN94, BHP14, Bai16b, BFG87, Bor91h, Bor92d, BZ92, Bor94g, BLN94a, BLN95, Bor95m, Bor95n, BLLN95, BLN96, BRS11, LY21, PHB14]. Utility [Roc20].

V [BSW82, Odl11]. Value [Bor99-27, Bor00u, BW98a, Bor98p, Bor99y, Bor99z]. valued [BBP03, BZ88, Zho12, Bor92b]. Values [BZ87, BB96c, BBBL98c, BBBL99, BBK00a, BK05, Bor10y, BZ11, BS11d, BS11e, BS17, BBBL97, BBBL98a, BBBL98b, BBK00b, BBK01, BBBL01, BB05c, BC10]. Vanderwerff [How14]. variable [BBM01, BBM02, KJR16]. Variant [YS00, LS00]. variants [Bor79f]. variation [Jal24]. Variational [Ano15, BZ94b, BZ97, BMS99b, Bor99u, Bor00v, Bor03-33, Bor04-31, BZ05, Bor06r, Bor06s, Bor06t, Bor06u, BZ06, Bor07n, Bor08i, Bor08j, Bor09-30, Bor09-28, Bor09-31, Bor09-29, Bor09-27, Bor13-34, Bor13-35, Bor13-33, BZ13, Bor16z, Geo05, YS00, Bor86g, BP87, Bor87h, Bor87i, Bor87j, Bor90m, Bor90n, Bor97p, Bor98l, Bor98m, BTZ99, Bor99t, BCFR04, Bor09l, Bor10p, Bor13-31, BZ16, Fab89, KPS16, KPS17, LS00, QR07]. Variations [Bor05c, BB05d, Bor10b, Bor10-31, BB06c]. various [BBP97, Bor92g, Bor92h, Bor93f, Bor93g]. vector [BBP03, BY84, BN84, BZ91, BZ93, JN03, KPS17]. vector-valued [BBP03]. Vectors [BSxx, BL92a]. Vera [BO11b]. Verifiable [BZ88]. version [BWB97, Koh01]. versus [BB12u, BB12n]. vertex [KMY00]. very [BB83, Bor14z, Bor16-27]. via [BMCL18, Bor87k, BBT92, BG96a, BG97b, BFV97, BCM02, BCM03, Bor06-30, BBC08a, CZX21, EB08, NFB17b, TB80]. victorious [BB11o]. victory [BB11-28]. view [BB17]. Views [DD15, BS97b, Bor97o, Bor98c]. viii [Bai91]. violence [BB11y]. viral [Bor15a]. Virtual [Bor95t, Bor95u, Bor96k, Bor97x, Bor98r]. Viscosity [Bor94m, Bor95v, BZ95, BZ96]. viscous [NFB17b]. Visibility [BEO76, BEO77]. vision [Bor94n]. Visit [BV24]. Visual [Bor14h, Bor14i, Bor14j, Bor14k, Bor14l, Bor14m, Bor14n, Bor15h, Bor16-28, JJ20]. Visualisation [Bor05-42, Bor05-43, Bor05-44]. Visualization [BBB^{+}20, Bor05-45, Bor14-27]. visualizing [BWB97]. vita [Bor08a]. Vol [BM97a]. volume [Bor06a]. volumes [Bor00r, Bor01p, Bor01q, Bor01r, BBM01, BBM02]. vs [BB13n, BB15h, BB15i]. vu [Tre13].

W [BB13k]. Wadsworth [Bai91]. Wagon [Bor05g]. wait [BB13-47]. Waldvogel [Bor05g]. Walk [BSW13, CC20b, SZ20, BNSW11, Bor15n, Bor16e]. Walking [ABBB13, Bor13-28, Bor13-29, Bor16-29, Bor16-30, Bor16-31, Bor13u, Bor13v, Bor13w, Bor13x, Bor13y, Bor13z, Bor13-27, Bor14x, Bor14y, Bor16q, Bor16r, Bor16s]. Walks [Bor10-28, Bor10-29, Bor11f, Bor11-27, Bor11-28, Bor11-34, Bor11-35, Bor11-33, Bor12-32, Bor10e, BNSW10, BSWZ11, Bor12b, BSWZ12, BS13, Bor14t, BSV15, Bor15o, Bor15p, Bor15q, BS16b, BSV16, BS16a]. Walter [Bor90b]. warming [BB10c, BB12-28, BB12c]. warning [BBLZ14r]. Washington [AMM10, Coh15]. Watson [BB11o, BB11-28, Bor11e]. waves [BB14m, BB16f]. Way [BB12-34, BB12-35, BB13s, BB87c, Bor15t, BB13r, BB13w, Bor11o, Bor11a]. Ways [Bor94o]. Weak [Bor78a, Bor79g, BF93c, BF93d, BFG03]. weaker [BBLZ14i]. Web [Bor97a, Bor97b, Bor97c, Bor99x, Bor97e, Bor97d, Bor97j, Bor98r, BBB^{+}96a, Bor96b, Bor96d, Bor97i, Bor98a]. weeks [Bor10-30]. Welcome [Bor02r]. Well [BB15s, BBLZ16b, BBLZ16a]. Wellesley [Odl11]. were [BB12-52, BB13j]. West [Bor05k]. Western [BB11k, Sel16]. WestGrid [Bor01m, Bor03-31]. Where [BB11-30, BB11g, BB15-29, BB15-30, BBLZ16d]. whether [BB11x]. which [BF93a]. while [BB09k, BB11z]. Who [BB91d, Bor15b, Bor15t, BWB97, Bor16d]. whose [BFG03, BS10a]. Wide [BBB^{+}96a]. Wiersma [BWY10, MR11]. Wightwick [Bai16a, BE16]. Wigner [BBS13b, BBS14b]. Wijsman [BV93a, BV94c]. wild [Bor02g]. Wiley [Ber88]. Will [BB15z, BB16s, BB15y]. William [Bor77d]. Wilson [BB13-45, BB13-46]. windmills [BB14y]. winners [Bor14b]. Winter [BM97a]. wired [BB14d]. wireless [Bor00w]. wishing [Bor01f]. within [ABMMY13, ABMMY14]. without [Bor76a, BW79b, BW82a, BW82b, Bor84a, BBY11, BBY13]. Witt [BL92a, BSxx]. Witten [Bor12e, Bor12f, BBC14b, BB15a, BBB15, BB16a, BB16b, BB18, Bor05x, Bor08k, Bor09m, Bor12r, BDT16, BD16a]. Wittgenstein [BBLZ13h]. Wokingham [BF06b]. Wonderful [Bor93m, Bor91p, Bor91q]. word [BB12d]. Words [BS14a, BS14b]. work [BBLZ16b, Bor02o, Bor04-33, Bor06-36]. Working [Bor01a, Bor01b, Bor01c, Bor01d, Bor06e]. works [BB12z, Bor07q, Bor07p, BR14b]. Workshop [BBL^{+}13, BBC^{+}14a, BBJC97, IMR92, RZ15, BB14a]. Workspaces [Bor98j]. World [Bor03-35, BMP05, Fer91, BB12-41, BBB^{+}96a]. Worrying [Dev20]. Would [BB12-36]. wreck [Bor15c]. writings [BB10l]. wrong [BB09f, BB13m, BB13-45, BB13-46]. WSN [LY21]. WWII [BB13t].

X [Bor05g, BB91d, Zei05]. xii [BB93g, BC96, Bou06, Odl11]. xil [Bor05g]. XSEDE [JWDS^{+}14]. xue [BB95e, lL09, lL09]. xv [Ber88]. xviii [Coh15]. xxii [Bor06o, Bor09b].

year [BBLZ13d, BB15-28, BBxxc]. Years [Bor02c, Bor02q, Bor07d, Bor09j, Bor09k, BBJ12, bVP21, BBLZ14i, BB15q, BB15z, BB15y, BD95, Bor08r, Bor10n, Bor12j, Bor12k, Bor15l]. Yes [BB12-53, BB13-33]. York [Ber88, BB91d, BB93g, Tod03]. Young [Bor97g, Bor98g]. you’re [BB13i]. yourself [BB12-31]. yu [lL09].

Zagier [BBB96b, BBB96c, BBB97d, Bor97f]. Zahl [BB96d]. Zang [Bor90b]. Zeidler [Bor06o]. zero [BB11-27, BBY12, BBY14, BB15d]. ZETA [Bor97q, BB96c, BBC98, BBK00a, BBC00b, Bor05x, Bor07g, Bor08k, Bor09m, Bor10y, BZ11, BD16a, Dil21, BB15c, BBB15, BS17, BBBL97, BBBL98a, BBBL98b, BB98c, BB98d, BBK00b, BBK01, BB05c, Bor06h, BC10, BDT16, BD18]. Zeta-Function [Bor08k, BS17]. Zhai [Coh15]. zheng [lL09]. Zoo [BV24].

References

Andrews:1988:RRP

[AAB^{+}88] George E. Andrews, Richard A. Askey, Bruce C. Berndt, K. G. Ramanathan, Robert A. Rankin, et al., editors. Ramanujan revisited: proceedings of the centenary conference, University of Illinois at Urbana-Champaign, June 1–5, 1987. Academic Press, New York, NY, USA, 1988. ISBN 0-12-058560-X. LCCN QA1 .R26 1987.

Alladi:2012:PRA

[AAB12] Krishnaswami Alladi, George E. Andrews, and Jonathan M. Borwein. Preface to Ramanujan’s 125th anniversary special issue. The Ramanujan Journal, 29(1–3):1–2, December 2012. CODEN RAJOF9. ISSN 1382-4090 (print), 1572-9303 (elec-

            tronic). URL http://link.springer.com/article/10.1007/
            s11139-012-9448-9.

Alaca:2006:TDT

[AAW06] Ayşe Alaca, Şaban Alaca, and Kenneth S. Williams. On the twodimensional theta functions of the Borweins. Acta Arithmetica, 124(2):177–195, 2006. CODEN AARIA9. ISSN 0065-1036 (print), 1730-6264 (electronic).

AragonArtacho:2012:GCN

[AB12] F. J. Aragón Artacho and J. M. Borwein. Global convergence of a non-convex Douglas–Rachford iteration. ArXiv e-prints, March

2012. URL http://adsabs.harvard.edu/abs/2012arXiv1203.
2392A;http://docserver.carma.newcastle.edu.au/1061/.

AragonArtacho:2013:GCN

[AB13] Francisco J. Aragón Artacho and Jonathan M. Borwein. Global convergence of a non-convex Douglas–Rachford iteration. Journal of Global Optimization, 57(3):753–769, 2013. CODEN JGOPEO. ISSN 0925-5001 (print), 1573-2916 (electronic). URL

http://arxiv.org/abs/1203.2392;http://docserver.carma.
newcastle.edu.au/1061/.

Adiga:2015:RGT

[AB15] Chandrashekar Adiga and Nasser Abdo Saeed Bulkhali. On Ramanujan’s general theta function and a generalization of the Borweins’ cubic theta functions. Asian-Eur. J. Math., 8(1):1550002, 16, 2015. ISSN 1793-5571 (print), 1793-7183 (electronic).

Abbott:2000:BRP

[Abb00] Steve Abbott. Book review: Pi: A Source Book. Mathematical Gazette, 84(501):547, November 2000. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.

jstor.org/stable/3620801.

AragonArtacho:2013:WRN

[ABBB13] Francisco J. Aragón Artacho, David H. Bailey, Jonathan M. Borwein, and Peter B. Borwein. Walking on real numbers. The Mathematical Intelligencer, 35(1):42–60, March 2013. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic). URL

http://docserver.carma.newcastle.edu.au/1475/;http://
gigapan.com/gigapans/106803;http://link.springer.com/
article/10.1007/s00283-012-9340-x;http://www.davidhbailey.
com/dhbpapers/tools-walk.pdf.

Amdeberhan:2012:FEC

[ABBS12] Tewodros Amdeberhan, David Borwein, Jonathan M. Borwein, and Armin Straub. On formulas for π experimentally conjectured by Jauregui–Tsallis. Journal of Mathematical Physics, 53(7):073708, July 2012. CODEN JMAPAQ.

ISSN 0022-2488. URL http://adsabs.harvard.edu/abs/
2012JMP....53g3708A;http://docserver.carma.newcastle.
edu.au/1378/.

Asperti:2003:MKM

[ABD03] Andrea Asperti, Bruno Buchberger, and James Harold Davenport, editors. Mathematical knowledge management: second international conference, MKM 2003, Bertinoro, Italy, February 16– 18, 2003: proceedings, volume 2594 of Lecture notes in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 3-540-00568-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.95

.I565 2003. URL http://www.loc.gov/catdir/enhancements/
fy0817/2003042408-d.html.

AragonArtacho:2013:ACA

[ABMMY13] F. J. Aragón Artacho, J. M. Borwein, V. Martín-Márquez, and L. Yao. Applications of convex analysis within mathematics. ArXiv e-prints, February 2013. URL http://adsabs.

harvard.edu/abs/2013arXiv1302.1978A;http://docserver.
carma.newcastle.edu.au/1513/.

AragonArtacho:2014:ACA

[ABMMY14] Francisco J. Aragón Artacho, Jonathan M. Borwein, Victoria Martín-Márquez, and Liangjin Yao. Applications of convex analysis within mathematics. Mathematical Programming, 148 (1–2):49–88, December 2014. CODEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic). URL http://arxiv.org/

abs/1302.1978;http://docserver.carma.newcastle.edu.au/

1513/. Special volume of Serdica Mathematical Journal in honour of Asen Dontchev.

AragonArtacho:2013:DRF

[ABT13a] F. J. Aragón Artacho, J. M. Borwein, and M. K. Tam. Douglas– Rachford feasibility methods for matrix completion problems. ArXiv e-prints, August 2013. URL http://adsabs.harvard.

edu/abs/2013arXiv1308.4243A.

AragonArtacho:2013:RRDa

[ABT13b] F. J. Aragón Artacho, J. M. Borwein, and M. K. Tam. Recent results on Douglas–Rachford methods for combinatorial optimization problems. ArXiv e-prints, May 2013. URL http://adsabs.

harvard.edu/abs/2013arXiv1305.2657A;http://docserver.
carma.newcastle.edu.au/1510/.

AragonArtacho:2013:RRDb

[ABT13c] Francisco J. Aragón Artacho, Jonathan M. Borwein, and Matthew K. Tam. Recent results on Douglas–Rachford methods. Serdica. Mathematical Journal. Serdika. Matematichesko Spisanie, 39(3–4):313–330, 2013. ISSN 1310-6600.

AragonArtacho:2014:DRF

[ABT14a] Francisco J. Aragón Artacho, Jonathan M. Borwein, and Matthew K. Tam. Douglas–Rachford feasibility methods for matrix completion problems. The ANZIAM Journal, 55(4):299–326, April 2014. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735

(electronic). URL http://arxiv.org/abs/1308.4243;http://
journals.cambridge.org/action/displayAbstract?fromPage=
online&aid=9347336;https://www.cambridge.org/core/journals/
anziam-journal/article/douglasrachford-feasibility-methods-for-matrix-completion-problems/
0DCB430BF7CF1187A7A1DB5B9C3C2BCC.

AragonArtacho:2014:RRD

[ABT14b] Francisco J. Aragón Artacho, Jonathan M. Borwein, and Matthew K. Tam. Recent results on Douglas–Rachford methods for combinatorial optimization problems. Journal of Optimization Theory and Applications, 163(1):1–30, October 2014. CODEN JOTABN. ISSN 0022-3239 (print), 1573-2878 (electronic). URL

http://arxiv.org/abs/1305.2657;http://docserver.carma.
newcastle.edu.au/1510/.

AragonArtacho:2015:GBD

[ABT15] F. J. Aragón Artacho, J. M. Borwein, and M. K. Tam. Global behavior of the Douglas–Rachford method for a nonconvex feasibility problem. ArXiv e-prints, June 2015. URL http://adsabs.

harvard.edu/abs/2015arXiv150609026A;http://docserver.
carma.newcastle.edu.au/1701/.

AragonArtacho:2016:GBD

[ABT16] Francisco J. Aragón Artacho, Jonathan M. Borwein, and Matthew K. Tam. Global behavior of the Douglas–Rachford method for a nonconvex feasibility problem. Journal of Global Optimization, 65(2):309–327, June 2016. CODEN JGOPEO. ISSN 0925-5001 (print), 1573-2916 (electronic). URL http:

//docserver.carma.newcastle.edu.au/1701/.

Artacho:2018:SGC

[AC18] Francisco J. Aragón Artacho and Rubén Campoy. Solving graph coloring problems with the Douglas–Rachford algorithm. Set-Valued and Variational Analysis, 26(2):277–304, June 2018.

CODEN ???? ISSN 1877-0533 (print), 1877-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s11228-017-0461-4.

Assis:2020:OTT

[AD20] Michael Assis and Michael Donovan. Origami as a teaching tool for indigenous mathematics education. In Bailey et al. [BBB^{+}20], pages 171–188. ISBN 3-030-36567-0 (print), 3-030-36568-9 (ebook). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Adegoke:2010:NBT

[Ade10] Kunle Adegoke. New binary and ternary digit extraction (BBPtype) formulas for trilogarithm constants. New York J. Math., 16: 361–367, 2010. ISSN 1076-9803. URL http://nyjm.albany.edu:

8000/j/2010/16_361.html.

Adegoke:2011:SRB

[Ade11] Kunle Adegoke. Symbolic routes to BBP-type formulas of any degree in arbitrary bases. Appl. Math. Inf. Sci., 5(2):264–275, 2011. ISSN 1935-0090 (print), 2325-0399 (electronic).

Adegoke:2012:NAD

[Ade12] Kunle Adegoke. A novel approach to the discovery of binary BBP-type formulas for polylogarithm constants. Integers, 12(3): 345–371, 2012. CODEN INTEHN. ISSN 1867-0652 (print), 1867-0660 (electronic).

Adegoke:2013:FPD

[Ade13] Kunle Adegoke. Formal proofs of degree 5 binary BBP-type formulas. Functiones et Approximatio Commentarii Mathematici, 48(part 1):19–27, 2013. ISBN 83-232-2533-8. ISSN 0208-6573 (print), 2080-9433 (electronic).

Adegoke:2014:GRF

[Ade14a] Kunle Adegoke. The Golden Ratio, Fibonacci numbers and BBPtype formulas. Fibonacci Quarterly, 52(2):129–138, May 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.

ca/Abstracts/52-2/adegoke.pdf.

Adegoke:2014:NBB

[Ade14b] Kunle Adegoke. A new binary BBP-type formula for \sqrt{5} log ϕ. Fibonacci Quarterly, 52(4):357–359, November 2014. CODEN FIBQAU. ISSN 0015-0517. URL http://www.fq.math.ca/

Abstracts/52-4/adegoke.pdf;http://www.fq.math.ca/Papers/
52-4/adegoke4282014.pdf.

Almkvist:1999:BBA

[AG99] Gert Almkvist and Andrew Granville. Borwein and Bradley’s Apéry-like formulae for ζ(4n + 3). Experimental Mathematics, 8 (2):197–203, 1999. CODEN ???? ISSN 1058-6458 (print), 1944-

950X (electronic). URL http://projecteuclid.org/euclid.
em/1047477060.

Arndt:2001:PU

[AH01] Jörg Arndt and Christoph Haenel. Pi — Unleashed. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-66572-2 (paperback), 3-642-56735-5 (ebook). xii + 270 pp. LCCN QA484.A7513 2001. US$. Includes CD-ROM. Translated from the German by Catriona and David Lischka.

AragonArtacho:2017:AMO

[AHLC^{+}17a] Francisco Javier Aragón Artacho, Rene Henrion, Marco Antonio Lopez-Cerda, Claudia Sagastizabal, and Jonathan M. Borwein. Advances in monotone operators theory and optimization (part 2) preface. Set-Valued and Variational Analysis, 25(4, 2):637–638, December 2017. ISSN 1877-0533 (print), 1877-0541 (electronic).

AragonArtacho:2017:SIA

[AHLC^{+}17b] Francisco Javier Aragón Artacho, Rene Henrion, Marco Antonio Lopez-Cerda, Claudia Sagastizabal, and Jonathan M. Borwein. Special issue: Advances in monotone operators theory and optimization preface. Set-Valued and Variational Analysis, 25(3 (SI)):463–465, September 2017. ISSN 1877-0533 (print), 1877-0541 (electronic).

Arutyunov:2018:SPN

[AI18] A. V. Arutyunov and A. F. Izmailov. Stability of possibly nonisolated solutions of constrained equations, with applications to complementarity and equilibrium problems. Set-Valued and Variational Analysis, 26(2):327–352, June 2018. CODEN ???? ISSN 1877-0533 (print), 1877-0541 (electronic). URL http://link.

springer.com/article/10.1007/s11228-017-0459-y.

Asic:1986:PSS

[AJ86] Miroslav D. Asic and A. A. Jagers. Problems and solutions: Solutions of elementary problems: E2995. American Mathematical Monthly, 93(5):401–402, May 1986. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [ANO^{+}83].

Adler:1986:PSS

[AJB86] Irving Adler, A. A. Jagers, and J. M. Borwein. Problems and solutions: Solutions of elementary problems: E2997. American Mathematical Monthly, 93(4):304, April 1986. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [ANO^{+}83].

Azmi:2022:CMI

[AK22] Behzad Azmi and Karl Kunisch. On the convergence and meshindependent property of the Barzilai–Borwein method for PDEconstrained optimization. IMA Journal of Numerical Analysis, 42(4):2984–3021, October 2022. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic). URL http://academic.

oup.com/imajna/article/42/4/2984/6340562.

Adegoke:2010:HDI

[AL10] Kunle Adegoke and Olawanle Layeni. The higher derivatives of the inverse tangent function and rapidly convergent BBP-type formulas for pi. Appl. Math. E-Notes, 10:70–75, 2010. ISSN 1607-2510.

Altman:2020:HAE

[Alt20] Morris Altman. A holistic approach to empirical analysis: The insignificance of P, hypothesis testing and statistical significance*. In Bailey et al. [BBB^{+}20], pages 233–253. ISBN 3-030-36567-0 (print), 3-030-36568-9 (e-book). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Amdeberhan:2010:GEM

[AMM10] Tewodros Amdeberhan, Luis A. Medina, and Victor H. Moll, editors. Gems in experimental mathematics: AMS Special Session, Experimental Mathematics, January 5, 2009, Washington, DC, volume 517 of Contemporary Mathematics. American Mathematical Society, Providence, RI, USA, 2010. ISBN 0-8218-4869-0, 0-8218-8196-5 (e-book). LCCN QA164 .A475 2009.

Asic:1983:PSE

[ANO^{+}83] Miroslav D. Asic, Phil Novinger, Daniel Oberlin, Irving Adler, Clark Kimberling, J. D. Shallit, and P. Erdős. Problems and solutions: Elementary problems: E2995–E3000. American Mathematical Monthly, 90(5):334–335, May 1983. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [AJB86, AJ86, EWM86, KJ86, NOL86, SZUM86].

Anonymous:2015:IJB

[Ano15] Anonymous. Interview with Jon Borwein at the International Conference on Variational Analysis, Optimization and Quantitative Finance in honor of Terry Rockafellar’s 80th Birthday. Web

video, May 18–22, 2015. URL http://terryfest2015.xlim.fr/
?page_id=482.

Anonymous:2016:JMB

[Ano16] Anonymous. Jonathan M. Borwein, former CMS President, dies at 65. ACM Communications in Computer Algebra, 50(3):121– 122, September 2016. CODEN ???? ISSN 1932-2232 (print), 1932-2240 (electronic).

Apkarian:2018:NSO

[ANR18] Pierre Apkarian, Dominikus Noll, and Laleh Ravanbod. Nonsmooth optimization for robust control of infinite-dimensional systems. Set-Valued and Variational Analysis, 26(2):405–429, June 2018. CODEN ???? ISSN 1877-0533 (print), 1877-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s11228-017-0453-4.

Arzani:2016:NNF

[AP16] F. Arzani and M. Reza Peyghami. A new nonmonotone filter Barzilai–Borwein method for solving unconstrained optimization problems. International Journal of Computer Mathematics, 93 (3):596–608, 2016. CODEN IJCMAT. ISSN 0020-7160.

Adly:2013:NMS

[AR13] Samir Adly and Hadia Rammal. A new method for solving Pareto eigenvalue complementarity problems. Computational optimization and applications, 55(3):703–731, July 2013. CODEN CPPPEF. ISSN 0926-6003 (print), 1573-2894 (elec-

tronic). URL http://link.springer.com/article/10.1007/

s10589-013-9534-y. Dedicated to Jonathan Borwein in honor of his 60th birthday.

AragonArtacho:2007:NSC

[Ara07] Francisco J. Aragón Artacho. A new and self-contained proof of Borwein’s norm duality theorem. Set-Valued Analysis, 15(3): 307–315, 2007. CODEN SVANEG. ISSN 0927-6947 (print), 1572-

932x (electronic). URL http://docserver.carma.newcastle.
edu.au/914.

AragonArtacho:2008:NSC

[Ara08] Francisco J. Aragón Artacho. A new and self-contained proof of Borwein’s norm duality theorem. Report, Department of Statistics and Operations Research, University of Alicante, 03071 Alicante, Spain, May 29, 2008. 13 pp. URL http://docserver.

carma.newcastle.edu.au/914.

Askey:1988:BRP

[Ask88] Richard Askey. Book review: Pi and the AGM. American Mathematical Monthly, 95(9):895–897, November 1988. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

https://www.jstor.org/stable/2322925.

Adamchik:1997:SF

[AW97] Victor Adamchik and Stan Wagon. A simple formula for π. American Mathematical Monthly, 104(9):852–855, November 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://www.maa.org/pubs/monthly_nov97_toc.

html. The authors employ Mathematica to extend earlier work of Bailey, Borwein, and Plouffe, [BBP97], done in 1995, but only just published, that discovered an amazing formula for π as is a power series in 16^{−k}, enabling any base-16 digit of π to be computed without knowledge of any prior digits. In this paper, Mathematica is used to find several simpler formulas having powers of 4^{−k}. They also note that it has been proven that their methods cannot be used to exhibit similar formulas in powers of 10^{−k}.

An:2020:NCS

[AX20] Congpei An and Yuchen Xiao. Numerical construction of spherical t-designs by Barzilai–Borwein method. Applied Numerical Mathematics: Transactions of IMACS, 150(??):295–302, April 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0168927419302831.

Borwein:2020:ELJ

[BaHO20] Naomi Simone Borwein and Judy anne Heather Osborn. On the educational legacies of Jonathan M. Borwein. In Bailey et al. [BBB^{+}20], pages 103–131. ISBN 3-030-36567-0 (print), 3-030-36568-9 (e-book). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Bailey:1988:CDD

[Bai88] David H. Bailey. The computation of π to 29,360,000 decimal digits using Borweins’ quartically convergent algorithm. Mathematics of Computation, 50(181):283–296, January 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

https://www.jstor.org/stable/2007932.

Bailey:1991:BRD

[Bai91] David H. Bailey. Book review: Jonathan Borwein and Peter Borwein, A Dictionary of Real Numbers, Wadsworth & Brooks/ Cole Advanced Books & Software, Pacific Grove, California, 1990, viii + 424 pp., 281 cm. Price $69.95. Mathematics of Computation, 56(193):403–404, January 1991. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

https://www.jstor.org/stable/2008568.

Bailey:2016:BBM

[Bai16a] David H. Bailey. Bailey, Borwein, Mattingly and Wightwick to receive the Levi L. Conant Prize from AMS. Web document., De-

cember 2, 2016. URL http://experimentalmath.org/2016/12/
bailey-borwein-mattingly-and-wightwick-to-receive-the-levi-l-conant-prize-from-ams/
;http://www.ams.org/news?news_id=3232.

Bailey:2016:CDD

[Bai16b] David H. Bailey. The computation of π to 29,360,000 decimal digits using Borweins’ quartically convergent algorithm (1988). In Bailey and Borwein [BB16l], pages 109–124. ISBN 3-319-32375-X, 3-319-32377-6 (e-book). LCCN QA251. URL

http://docserver.carma.newcastle.edu.au/1716/;http://
lib.myilibrary.com?id=941862.

Bailey:2016:JBD

[Bai16c] David H. Bailey. Jonathan Borwein dies at 65. Math Drudge,

August 2, 2016. URL https://experimentalmath.info/blog/
2016/08/jonathan-borwein-dies-at-65/.

Bailey:2016:JMB

[Bai16d] David H. Bailey. Jonathan M. Borwein’s extraordinary mathematical career. Canadian Mathematical Society Notes, 48(6):14– 15, December 2016. ISSN 0045-5164. URL http://cms.math.ca/

notes/v48/n6/Notesv48n6.pdf;http:www.jonborwein.org/jmbpapers/.

Bailey:2017:JBC

[Bai17a] David H. Bailey. Jonathan Borwein commemorative conference[: 25–29 September 2017 in Newcastle, NSW, Australia].

Web blog, June 6, 2017. URL http://mathscholar.org/
jonathan-borwein-commemorative-conference.

Bailey:2017:JBEa

[Bai17b] David H. Bailey. Jonathan Borwein: Experimental mathematician. Report, Department of Computer Science, University of California, Davis, Davis, CA 95616, USA, January

10, 2017. URL http://www.davidhbailey.com/dhbpapers/
dhb-jmb-em.pdf;http://www.tandfonline.com/doi/abs/10.
1080/10586458.2017.1284625.

Bailey:2017:JBEb

[Bai17c] David H. Bailey. Jonathan Borwein: Experimental mathematician. Experimental Mathematics, 26(2):125–129, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2017.1284625.

Bailey:2017:PCP

[Bai17d] David H. Bailey. Pi and the collapse of peer review. Math-Scholar Web blog., July 20, 2017. URL https://mathscholar.

org/2017/07/pi-and-the-collapse-of-peer-review/.

Bailey:2017:RJB

[Bai17e] David H. Bailey. Report on the Jon Borwein remembrance day meeting in Paris [10 February 2017]. Web blog.,

February 15, 2017. URL http://jonborwein.org/2017/02/
report-on-the-jon-borwein-remembrance-day-meeting-in-paris/.

Bailey:2020:PBD

[Bai20] David H. Bailey. Peter Borwein dies at 67. MathScholar Web

blog., August 29, 2020. URL https://mathscholar.org/2020/
08/peter-borwein-dies-at-67/.

Bailey:2021:JBR

[Bai21] David H. Bailey. Jonathan Borwein: Renaissance mathematician. American Mathematical Monthly, 128(9):773–779, 2021. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bankov:2010:BRC

[Ban10] Kiril Bankov. Book review: Communicating mathematics in the digital era, by J. M. Borwein, E. M. Rocha, and J. F. Rodrigues. Pp. 325. $49.00 (Hardback), 2008. ISBN 978-1-56881-410-0 (A K Peters, Ltd.). Mathematical Gazette, 94(531):557–559, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.jstor.org/stable/25759758.

Borwein:2012:REA

[BaO12] Jonathan M. Borwein and Judy anne Osborn. Response to “Experimental Approaches to Theoretical Thinking … ”. In Hanna and de Villiers [Hd12], page ?? ISBN 94-007-2128-5, 94-007-2129-3 (e-book). LCCN QA9.54 .P766 2012. URL

http://www.springerlink.com/content/978-94-007-2129-6.

Borwein:1983:VRC

[BB83] J. M. Borwein and P. B. Borwein. A very rapidly convergent product expansion for π. BIT (Nordisk tidskrift for informationsbehandling), 23(4):538–540, December 1983. CO-DEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-

9125 (electronic). URL http://docserver.carma.newcastle.
edu.au/1634/;http://link.springer.com/article/10.1007/
BF01933626;http://www.springerlink.com/openurl.asp?genre=
article&issn=0006-3835&volume=23&issue=4&spage=538.

Borwein:1984:AGM

[BB84a] J. M. Borwein and P. B. Borwein. The arithmetic–geometric mean and fast computation of elementary functions. SIAM Review, 26(3):351–366, July 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https://www.jstor.org/

stable/2031275.

Borwein:1984:CHO

[BB84b] J. M. Borwein and P. B. Borwein. Cubic and higher order algorithms for π. Canadian mathematical bulletin = Bulletin canadien de mathématiques, 27(4):436–443, December 1984. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL

http://docserver.carma.newcastle.edu.au/1622/.

Borwein:1984:EIA

[BB84c] J. M. Borwein and P. B. Borwein. Elliptic integrals and approximations to π. Typescript, with 84-01 added by hand on cover page., January 1984.

Borwein:1984:EAO

[BB84d] J. M. Borwein and P. B. Borwein. Explicit algebraic n th order approximations to pi. In Singh et al. [SBW84], pages 247–256. ISBN 94-009-6466-8, 94-009-6468-4. ISSN 1389-2185.

LCCN ???? URL http://link.springer.com/chapter/10.
1007/978-94-009-6466-2_12.

Borwein:1984:RCC

[BB84e] J. M. Borwein and P. B. Borwein. Reduced complexity calculation of log. Technical Report DALTR 84-01, Department of Mathematics, Dalhousie University, Halifax, NS, Canada, January 1984. 17 pp.

Borwein:1984:TCS

[BB84f] J. M. Borwein and P. B. Borwein. A tangent cone separation principle. Report DALTR 84-03, Department of Mathematics, Dalhousie University, Halifax, NS, Canada, January 1984.

Borwein:1985:PSA

[BB85] Jon Borwein and Anatole Beck. Problems and solutions: Advanced problems: 6491–6492. American Mathematical Monthly, 92(3):217, March 1985. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BL87, BN86].

Borwein:1986:NAS

[BB86a] David Borwein and Jonathan M. Borwein. A note on alternating series in several dimensions. American Mathematical Monthly, 93(7):531–539, August/September 1986. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://docserver.carma.newcastle.edu.au/1610/.

Borwein:1986:ECI

[BB86b] J. M. Borwein and P. B. Borwein. An explicit cubic iteration for π. BIT (Nordisk tidskrift for informationsbehandling), 26(1): 123–126, March 1986. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic). URL http://docserver.

carma.newcastle.edu.au/1609/;http://link.springer.com/
article/10.1007/BF01939368;http://www.springerlink.com/
openurl.asp?genre=article&issn=0006-3835&volume=26&issue=
1&spage=123.

Borwein:1986:MQC

[BB86c] J. M. Borwein and P. B. Borwein. More quadratically converging algorithms for π. Mathematics of Computation, 46(173):247–253, January 1986. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-

6842 (electronic). URL http://docserver.carma.newcastle.
edu.au/1614/.

Borwein:1987:ERT

[BB87a] J. M. Borwein and P. B. Borwein. Explicit Ramanujan-type approximations to pi of high order. Proceedings of the Indian Academy of Sciences — Mathematical Sciences, 97(1–3):53–59, 1987. ISSN 0253-4142 (print), 0973-7685 (electronic). URL http:

//link.springer.com/article/10.1007/BF02837813. Centen-

nial issue.

Borwein:1987:RRA

[BB87b] J. M. Borwein and P. B. Borwein. Ramanujan’s rational and algebraic series for 1/π. J. Indian Math. Soc. (N.S.), 51:147–160 (1988), 1987. CODEN JIMTA2. ISSN 0019-5839. Ramanujan Centennial issue.

Borwein:1987:UPW

[BB87c] J. M. Borwein and P. B. Borwein. Unsolved problems: The way of all means. American Mathematical Monthly, 94(6):519–522, June/July 1987. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL https://www.jstor.org/stable/
2322842.

Borwein:1987:PAS

[BB87d] Jonathan M. Borwein and Peter B. Borwein. Pi and the AGM: a Study in Analytic Number Theory and Computational Complexity. Canadian Mathematical Society series of monographs and advanced texts = Monographies et études de la Société mathématique du Canada. Wiley, New York, NY, USA, 1987. ISBN 0-471-83138-7, 0-471-31515-X (paperback). xv + 414 pp. LCCN QA241 .B774 1987.

Barzilai:1988:TPS

[BB88a] Jonathan Barzilai and Jonathan M. Borwein. Two-point step size gradient methods. IMA Journal of Numerical Analysis, 8(1):141– 148, 1988. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Borowski:1988:DM

[BB88b] E. J. (Ephraim J.) Borowski and Jonathan M. Borwein. Dictionary of mathematics. Collins, London, UK, 1988. ISBN 0-00-434347-6 (paperback). xi + 659 pp. LCCN QA5. With contributions by J. F. Bowers.

Borwein:1988:CCJ

[BB88c] J. M. Borwein and P. B. Borwein. A cubic counterpart of Jacobi’s identity and the AGM. Report, Department of Mathematics, Statistics and Computing Science, Dalhousie University, Halifax, NS B3H 3J5, Canada, December 31, 1988. 20 pp.

Borwein:1988:MRT

[BB88d] J. M. Borwein and P. B. Borwein. More Ramanujan-type series for 1/π. In Andrews et al. [AAB^{+}88], pages 359–374. ISBN 0-12-058560-X. LCCN QA1 .R26 1987.

Borwein:1988:CFF

[BB88e] J. M. Borwein and P. B. Borwein. On the complexity of familiar functions and numbers. SIAM Review, 30(4):589–601, December 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-

7200 (electronic). URL http://docserver.carma.newcastle.
edu.au/1593/;https://www.jstor.org/stable/2030559.

Borwein:1988:RP

[BB88f] Jonathan M. Borwein and Peter B. Borwein. Ramanujan and Pi. Scientific American, 258(2):112–117, February 1988. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://docserver.carma.newcastle.edu.au/1379/;http://
www.nature.com/scientificamerican/journal/v258/n2/pdf/
scientificamerican0288-112.pdf. Japanese edition (April

1988), Russian edition (April 1988), German edition (May 1988). Reprinted in [Bor91o, BB96d, BB01f, BB16l].

Borwein:1989:ARM

[BB89a] J. M. Borwein and P. B. Borwein. Approximating π with Ramanujan’s modular equations. Rocky Mountain Journal of Mathematics, 19(1):93–102, 1989. CODEN RMJMAE. ISSN 0035-7596 (print), 1945-3795 (electronic). Constructive Function Theory— 86 Conference (Edmonton, AB, 1986).

Borwein:1989:MI

[BB89b] J. M. Borwein and P. B. Borwein. On the mean iteration (a, b) ← (\frac{a+3b}{4}, \frac{\sqrt{ab}+b}{2}). Mathematics of Computation, 53(187):311–326, July 1989. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-

6842 (electronic). URL http://docserver.carma.newcastle.
edu.au/1586/.

Beer:1990:MCR

[BB90a] Gerald Beer and Jonathan M. Borwein. Mosco convergence and reflexivity. Proceedings of the American Mathematical Society, 109(2):427–436, 1990. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL http://docserver.carma.

newcastle.edu.au/1581/.

Borwein:1990:RCM

[BB90b] J. M. Borwein and P. B. Borwein. A remarkable cubic mean iteration. Lecture Notes in Mathematics, 1435:27–31, 1990. CO-DEN LNMAA2. ISBN 3-540-52768-0 (print), 3-540-47139-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0087894/.

Borwein:1990:SSH

[BB90c] J. M. Borwein and P. B. Borwein. Strange series and high precision fraud. Preliminary report, Department of Mathematics, Statistics and Computing Science, Dalhousie University, Halifax, NS B3H 3J5, Canada, September 4, 1990. 28 pp.

Borwein:1990:DRN

[BB90d] Jonathan Borwein and Peter B. Borwein. A dictionary of real numbers. Wadsworth, Pacific Grove, CA, USA, 1990. ISBN 0-534-12840-8. viii + 424 pp. LCCN QA47.B625 1990.

Borowski:1991:HDM

[BB91a] E. J. Borowski and J. M. Borwein. The HarperCollins dictionary of mathematics. HarperPerennial, New York, NY, USA, 1991. ISBN 0-06-461019-5, 0-00-434347-6 (paperback). xii + 659 pp. LCCN QA5 B67 1989. URL http://docserver.carma. newcastle.edu.au/1738/. With the assistance of J. F. Bowers, A. Robertson and M. McQuillan, Revised reprint of the 1988 original [BB88b].

Borwein:1991:FPI

[BB91b] David Borwein and Jonathan Borwein. Fixed point iterations for real functions. Journal of Mathematical Analysis and Applications, 157(1):112–126, 1991. CODEN JMANAK. ISSN 0022-247X (print), 1096-0813 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022247X9190139Q.

Borwein:1991:CCJ

[BB91c] J. M. Borwein and P. B. Borwein. A cubic counterpart of Jacobi’s identity and the AGM. Transactions of the American Mathematical Society, 323(2):691–701, February 1991. CO-DEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
1578/;https://www.jstor.org/stable/2001551.

Borwein:1991:SGB

[BB91d] J. M. Borwein and P. B. Borwein. A singular genius: [book review] The Man Who Knew Infinity. A Life of the Indian Genius Ramanujan. Robert Kanigel, Scribner, New York, 1991. x + 438 pp. + plates, $27.95. Science, 253(5017):334–335, July 19, 1991. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL https://www.jstor.org/stable/2879085;http:
//www.ncbi.nlm.nih.gov/pubmed/17794700.

Borwein:1992:SSH

[BB92a] J. M. Borwein and P. B. Borwein. Strange series and high precision fraud. American Mathematical Monthly, 99(7):622–640, August/September 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also subsequent work [BV24].

Borwein:1992:SOC

[BB92b] Jonathan Borwein and Peter B. Borwein. Some observations on computer aided analysis. Notices of the American Mathematical Society, 39(8):825–829, October 1992. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://www.ams.

org/journals/notices/199210/199210FullIssue.pdf.

Bauschke:1993:CNA

[BB93a] H. H. Bauschke and J. M. Borwein. On the convergence of von Neumann’s alternating projection algorithm for two sets. Set-Valued Analysis, 1(2):185–212, 1993. CODEN SVANEG. ISSN 0927-6947 (print), 1572-932X (electronic). URL http:

//docserver.carma.newcastle.edu.au/1555/;http://link.
springer.com/article/10.1007/BF01027691.

Beer:1993:MSC

[BB93b] Gerald Beer and Jonathan M. Borwein. Mosco and slice convergence of level sets and graphs of linear functionals. Journal of Mathematical Analysis and Applications, 175(1):53–67, 1993. CODEN JMANAK. ISSN 0022-247X (print), 1096-0813

(electronic). URL http://docserver.carma.newcastle.edu.
au/1562/;http://www.sciencedirect.com/science/article/
pii/S0022247X83711510.

Borwein:1993:PSPb

[BB93c] David Borwein and Jonathan Borwein. Problems and solutions: Problems: 10335. American Mathematical Monthly, 100(8):796– 798, October 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Borwein:1993:CNT

[BB93d] J. M. Borwein and P. B. Borwein. Class number three Ramanujan type series for 1/π. Journal of Computational and Applied Mathematics, 46(1–2):281–290, June 14, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://docserver.carma.newcastle.edu.au/1560/;http://
www.sciencedirect.com/science/article/pii/037704279390302R.

Computational complex analysis.

Borwein:1993:GFI

[BB93e] J. M. Borwein and P. B. Borwein. On the generating function of the integer part: [nα + γ]. Journal of Number Theory, 43(3):293– 318, 1993. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658

(electronic). URL http://docserver.carma.newcastle.edu.
au/1564/;http://www.sciencedirect.com/science/article/
pii/S0022314X83710231.

Borwein:1993:ICM

[BB93f] Jonathan M. Borwein and Peter B. Borwein. Inequalities for compound mean iterations with logarithmic asymptotes. Journal of Mathematical Analysis and Applications, 177(2):572–582, 1993. CODEN JMANAK. ISSN 0022-247X (print), 1096-0813

(electronic). URL http://docserver.carma.newcastle.edu.
au/1553/;http://www.sciencedirect.com/science/article/
pii/S0022247X83712783.

Borwein:1993:MMB

[BB93g] Jonathan M. Borwein and Peter B. Borwein. Mathematical malaises: Book review: Pi in the sky: counting, thinking and being, John D. Barrow. Clarendon (Oxford University Press), New York, 1992. xii, 317 pp., illus. $25. Science, 259 (5103):1928–1930, March 26, 1993. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

ncbi.nlm.nih.gov/pubmed/17836254;https://www.science.
org/doi/10.1126/science.259.5103.1928.a.

Bauschke:1994:DAP

[BB94a] H. H. Bauschke and J. M. Borwein. Dykstra’s alternating projection algorithm for two sets. Journal of Approximation Theory, 79(3):418–443, 1994. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic). URL http://docserver.

carma.newcastle.edu.au/1543/;http://www.sciencedirect.
com/science/article/pii/S0021904584711361.

Borwein:1994:STE

[BB94b] David Borwein and Jonathan M. Borwein. On some trigonometric and exponential lattice sums. Journal of Mathematical Analysis and Applications, 188(1):209–218, 1994.

CODEN JMANAK. ISSN 0022-247X (print), 1096-0813

(electronic). URL http://docserver.carma.newcastle.edu.
au/1542/;http://www.sciencedirect.com/science/article/
pii/S0022247X84714223.

Bauschke:1995:CLM

[BB95a] Heinz H. Bauschke and Jonathan M. Borwein. Continuous linear monotone mappings in Banach space. CECM Research Report 95:049, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, August 29, 1995. 48

pp. URL http://docserver.carma.newcastle.edu.au/114/
;https://web.archive.org/web/20110313015539/.

Bauschke:1995:LFM

[BB95b] Heinz H. Bauschke and Jonathan M. Borwein. Legendre functions and the method of random Bregman projections. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, March 2, 1995. 48 pp. URL http://docserver.

carma.newcastle.edu.au/96. Published in [BB97a].

Bauschke:1995:PAS

[BB95c] Heinz H. Bauschke and Jonathan M. Borwein. On projection algorithms for solving convex feasibility problems. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, April

28, 1995. 99 pp. URL http://docserver.carma.newcastle.
edu.au/100/2/95_034-Bauschke-Borwein.pdf. Published in

[BB96b].

Borowski:1995:DCD

[BB95d] Ephraim J. Borowski and Jonathan M. Borwein. Dizionario Collins della matematica. Dizionari Gremese. Gremese Editore, Roma, Italy, 1995. ISBN 88-7605-813-3. 423 pp. LCCN ???? Translation to Italian by Andrea Stracca of [BB91a].

Borowski:1995:SXC

[BB95e] Ephraim J. Borowski and Jonathan M. Borwein. Shu xue ci dian, volume 10. Mao tou ying chu ban, Taibei, 1995. ISBN 957-0337-14-1. 879 pp. LCCN ???? Chinese edition of [BB91a].

Borwein:1995:IIS

[BB95f] David Borwein and Jonathan M. Borwein. On an intriguing integral and some series related to ζ(4). Proceedings of the American Mathematical Society, 123(4):1191–1198, 1995. CODEN PAM-YAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL

http://docserver.carma.newcastle.edu.au/1538/.

Bauschke:1996:MMD

[BB96a] Heinz H. Bauschke and Jonathan M. Borwein. Maximal monotonicity of dense type, local maximal monotonicity, and monotonicity of the conjugate are all the same for continuous linear operators. Report, Department of Combinatorics & Optimization, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada, November 20, 1996. 23 pp. URL http://docserver.carma.

newcastle.edu.au/180/. Published in [BB99c].

Bauschke:1996:PAS

[BB96b] Heinz H. Bauschke and Jonathan M. Borwein. On projection algorithms for solving convex feasibility problems. SIAM Review, 38(3):367–426, September 1996. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//docserver.carma.newcastle.edu.au/100/;http://epubs.
siam.org/sam-bin/dbq/article/25171.

Borwein:1996:SSA

[BB96c] Jonathan Borwein and David Bradley. Searching symbolically for Apéry-like formulae for values of the Riemann zeta function. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 30(2):2–7, June 1996. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
170/;http://doi.acm.org/10.1145/235699.235700.

Borwein:1996:SRZ

[BB96d] Jonathan M. Borwein and Peter B. Borwein. Srinivasa Ramanujan und die Zahl Pi. (German) [Srinivasa Ramanujan and the number π]. In Moderne Mathematik [Fal96], pages 60–68. ISBN 3-8274-0025-2. LCCN ???? URL http://docserver.

carma.newcastle.edu.au/1379/;https://web.archive.org/
web/20170227063951/. Reprint of [BB88f].

Bauschke:1997:LFM

[BB97a] Heinz H. Bauschke and Jonathan M. Borwein. Legendre functions and the method of random Bregman projections. Journal of Convex Analysis, 4(1):27–67, 1997. ISSN 0944-6532 (print), 2363-

6394 (electronic). URL http://docserver.carma.newcastle.
edu.au/96/;http://www.heldermann-verlag.de/jca/jca04/
jca04002.pdf.

Borwein:1997:AGMa

[BB97b] J. M. Borwein and P. B. Borwein. The arithmetic–geometric mean and fast computation of elementary functions. In Berggren et al. [BBB97b], pages 537–552. ISBN 0-387-94924-0, 1-4757-2736-4 (e-book), 1-4757-2738-0 (print), 3-540-94924-0. LCCN QA484

.P5 1997. URL http://link.springer.com/chapter/10.1007/
978-1-4757-2736-4_56. Reprint of [BB84a].

Borwein:1997:EDA

[BB97c] Jonathan Borwein and David Bradley. Empirically determined Apéry-like formulae for ζ(4n + 3). Experimental Mathematics, 6(3):181–194, 1997. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://docserver.carma.

newcastle.edu.au/163/;http://projecteuclid.org/euclid.
em/1047920419.

Borwein:1997:MMS

[BB97d] Jonathan M. Borwein and Peter B. Borwein. Mathematics on main street. Board–Faculty Association Dinner, Simon Fraser University, Burnaby, BC, Canada., April 24, 1997.

Bauschke:1998:COM

[BB98a] Heinz H. Bauschke and Jonathan M. Borwein. Conical open mapping theorems and regularity. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, November 10,

1998. 14 pp. URL http://docserver.carma.newcastle.edu.

au/208/. Published in [BB99b].

Borwein:1998:PAS

[BB98b] Jonathan M. Borwein and Peter B. Borwein. Pi and the AGM: A study in analytic number theory and computational complexity. Canadian Mathematical Society Series of Monographs and Advanced Texts, 4. Wiley, New York, NY, USA, 1998. ISBN 0-471-83138-7, 0-471-31515-X (paperback). xvi + 414 pp. Reprint of the 1987 original.

Borwein:1998:DRDa

[BB98c] Jonathan M. Borwein and David J. Broadhurst. Determinations of rational Dedekind-zeta invariants of hyperbolic manifolds and Feynman knots and links. ArXiv High Energy Physics — Theory e-prints, November 1998. URL http://adsabs.

harvard.edu/abs/1998hep.th...11173B;http://arxiv.org/
abs/hep-th/9811173;http://docserver.carma.newcastle.edu.
au/213/.

Borwein:1998:DRDb

[BB98d] Jonathan M. Borwein and David J. Broadhurst. Determinations of rational Dedekind-zeta invariants of hyperbolic manifolds and Feynman knots and links. CECM Preprint CECM-98-120, OUT-4102-76, hep-th/9811173, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, November 1998. 53

pp. URL http://docserver.carma.newcastle.edu.au/213/.

arXiv preprint hep-th/9811173.

Bailey:1999:EMR

[BB99a] David H. Bailey and Jonathan M. Borwein. Experimental mathematics: Recent developments and future outlook. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada,

1999. 18 pp. URL http://docserver.carma.newcastle.edu.
au/250.

Bauschke:1999:COM

[BB99b] Heinz H. Bauschke and Jonathan M. Borwein. Conical open mapping theorems and regularity. In J. Giles and B. Ninness, editors, Proceedings of the Centre for Mathematics and its Applications (Australian National University) National Symposium on Functional Analysis, Optimization and Applications, March 1998, volume 36, pages 1–10. ????, ????, 1999. URL http:

//docserver.carma.newcastle.edu.au/208/.

Bauschke:1999:MMD

[BB99c] Heinz H. Bauschke and Jonathan M. Borwein. Maximal monotonicity of dense type, local maximal monotonicity, and monotonicity of the conjugate are all the same for continuous linear operators. Pacific Journal of Mathematics, 189(1):1–20, 1999. CO-DEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic).

URL http://docserver.carma.newcastle.edu.au/180/.

Borowski:1999:DMa

[BB99d] E. J. (Ephraim J.) Borowski and Jonathan M. Borwein. Dictionary of mathematics. Unwin Hyman, London, UK, 1999. ISBN 0-261-67198-7. 660 pp. LCCN ????

Borowski:1999:DMb

[BB99e] E. J. (Ephraim J.) Borowski and Jonathan M. Borwein. Dictionary of mathematics. HarperCollins, London, UK, 1999. ISBN 0-00-434347-6 (paperback). 672 pp. LCCN ????

Borwein:19xx:CFF

[BBxxa] J. M. Borwein and P. B. Borwein. On the complexity of familiar functions and numbers. Report, Department of Mathematics, Statistics and Computing Science, Dalhousie University, Halifax, NS B3H 3J5, Canada, 19xx. 30 pp.

Borwein:19xx:MI

[BBxxb] J. M. Borwein and P. B. Borwein. On the mean iteration (a, b) ← ((a + 3b)/4, (\sqrt{ab} + b)/2). Report, Department of Mathematics, Statistics and Computing Science, Dalhousie University, Halifax, NS B3H 3J5, Canada, 19xx. 39 pp.

Borwein:19xx:PCT

[BBxxc] Jonathan M. Borwein and Peter B. Borwein. Pi and its computation: a twenty-two hundred year quest continues: why it is now possible to calculate a billion digits of pi. Report, Dalhousie University, Halifax, NS, Canada, 19xx. 21 pp. Undated typeset manuscript intended for Scientific American, but never published. The latest reference is to a 1987 book.

Bauschke:2000:JSC

[BB00a] Heinz H. Bauschke and Jonathan M. Borwein. Joint and separate convexity of the Bregman distance. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, 2000. 17 pp.

URL http://docserver.carma.newcastle.edu.au/227/.

Borwein:2000:AGM

[BB00b] J. M. Borwein and P. B. Borwein. The arithmetic–geometric mean and fast computation of elementary functions. In Berggren et al. [BBB00a], pages 537–552. ISBN 0-387-98946-3 (hardcover).

LCCN QA484 .P5 2000. URL http://link.springer.com/
chapter/10.1007/978-1-4757-3240-5_56. Reprint of [BB84a].

Bailey:2001:EMR

[BB01a] David H. Bailey and Jonathan M. Borwein. Experimental mathematics: recent developments and future outlook. In Engquist and Schmid [ES01], pages 51–66. ISBN 3-540-66913-2. LCCN QA7

.M32423 2001. URL http://docserver.carma.newcastle.edu.
au/250/.

Bauschke:2001:JSC

[BB01b] Heinz H. Bauschke and Jonathan M. Borwein. Joint and separate convexity of the Bregman distance. In Dan Butnariu, Yair Censor, and Simeon Reich, editors, Inherently Parallel Algorithms in Feasibility and Optimization and their Applications (Haifa, 2000), volume 8 of Studies in Computational Mathematics, pages 23–36. North-Holland, Amsterdam, The Netherlands, 2001. ISSN 1570-579X (print), 2212-1145 (electronic). URL http://docserver.

carma.newcastle.edu.au/227/;http://www.sciencedirect.
com/science/article/pii/S1570579X01800045.

Borwein:2001:SRP

[BB01c] David Borwein and Jonathan M. Borwein. Some remarkable properties of sinc and related integrals. The Ramanujan Journal, 5(1): 73–89, 2001. CODEN RAJOF9. ISSN 1382-4090 (print), 1572-

9303 (electronic). URL http://docserver.carma.newcastle.
edu.au/249/.

Borwein:2001:CMCa

[BB01d] Jonathan M. Borwein and Peter B. Borwein. Challenges in mathematical computing. Computing in Science and Engineering, 3(3):48–53, May/June 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL http://

dlib.computer.org/cs/books/cs2001/pdf/c3048.pdf;http:
//docserver.carma.newcastle.edu.au/135/;http://www.computer.
org/cse/cs1999c3048abs.htm.

Borwein:2001:CMCc

[BB01e] Jonathan M. Borwein and Peter B. Borwein. Challenges in mathematical computing. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, February 19, 2001. 12

pp. URL http://docserver.carma.newcastle.edu.au/135/2/
01_160-Borwein-Borwein-ocr.pdf.

Borwein:2001:RP

[BB01f] Jonathan M. Borwein and Peter B. Borwein. Ramanujan and pi. In Berndt and Rankin [BR01], pages 187–199. ISBN 0-8218-2624-7. ISSN 0899-2428. LCCN QA29.R3 R29 2001. URL

http://docserver.carma.newcastle.edu.au/1379/. Reprint

of [BB88f].

Borowski:2002:MCD

[BB02] Ephraim J. Borowski and Jonathan M. Borwein. Mathematics: Collins dictionary. Collins, London, UK, second edition, 2002. ISBN 0-00-710295-X. 633 pp. LCCN ????

Borwein:2004:AGMa

[BB04a] J. M. Borwein and P. B. Borwein. The arithmetic–geometric mean and fast computation of elementary functions. In Berggren et al. [BBB04a], pages 537–552. ISBN 0-387-

20571-3. URL http://link.springer.com/chapter/10.1007/
978-1-4757-4217-6_56. Reprint of [BB84a].

Borwein:2004:MEP

[BB04b] Jonathan M. Borwein and David H. Bailey. Mathematics by Experiment: Plausible Reasoning in the 21st Century. A. K. Peters, Ltd., Wellesley, MA, USA, 2004. ISBN 1-56881-211-

6. x + 288 pp. LCCN QA76.95 .B67 2003. US$45.00. URL

http://docserver.carma.newcastle.edu.au/272/.

Bailey:2005:FPC

[BB05a] D. H. Bailey and J. M. Borwein. Future prospects for computerassisted mathematics. Canadian Mathematical Society Notes, 37 (8):2–6, December 2005. ISSN 1193-9273 (print), 1496-4295 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
304/;https://cms.math.ca/notes/v37/n8/Notesv37n8.pdf.

Bailey:2005:EME

[BB05b] David H. Bailey and Jonathan M. Borwein. Experimental mathematics: examples, methods and implications. Notices of the American Mathematical Society, 52(5):502–514, May 2005. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
1241/;http://www.ams.org/notices/200505/fea-borwein.pdf.

Borwein:2005:SSA

[BB05c] J. M. Borwein and David M. Bradley. Searching symbolically for Apéry-like formulae for values of the Riemann zeta function. ArXiv Mathematics e-prints, May 2005. URL http://adsabs.

harvard.edu/abs/2005math......5093B;http://docserver.
carma.newcastle.edu.au/170/.

Borwein:2005:TTG

[BB05d] J. M. Borwein and David M. Bradley. Thirty-two Goldbach variations. ArXiv Mathematics e-prints, February 1,

2005. URL http://adsabs.harvard.edu/abs/2005math.....
.2034B;http://arxiv.org/abs/math/0502034;http://docserver.
carma.newcastle.edu.au/301/.

Borwein:2005:ADL

[BB05e] Jonathan M. Borwein and John Ball. Access to the digitized literature. MSRI Workshop on Digitizing Mathematics, April 15– 17, Berkeley, CA, USA., April 15, 2005.

Borwein:2005:EDA

[BB05f] Jonathan M. Borwein and David M. Bradley. Empirically determined Apéry-like formulae for ζ(4n + 3). ArXiv Mathematics

e-prints, May 2005. URL http://adsabs.harvard.edu/abs/
2005math......5124B;http://docserver.carma.newcastle.
edu.au/163/.

Borwein:2005:TFI

[BB05g] Jonathan M. Borwein and David M. Bradley. On two fundamental identities for Euler sums. arxiv.org, ??(??):??, February

2005. URL http://arxiv.org/abs/math.NT/0502034/;http:
//docserver.carma.newcastle.edu.au/279/.

Bailey:2006:EBE

[BB06a] D. H. Bailey and J. M. Borwein. Effective bounds in Euler– Maclaurin-based quadrature (summary for HPCS06). In 20th International Symposium on High-Performance Computing in an Advanced Collaborative Environment (HPCS’06), page 34. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, May 2006. ISSN 1550-5243 (print),

2378-2099 (electronic). URL http://www.davidhbailey.com/
dhbpapers/hpcs06.pdf.

Bailey:2006:FGE

[BB06b] David H. Bailey and Jonathan M. Borwein. Finding general explicit formulas for Ising integral recursions. Report, Lawrence Berkeley National Laboratory and Faculty of Computer Science, Dalhousie University, Berkeley, CA 94720, USA and Halifax, NS, B3H 2W5, Canada, October 2, 2006. URL http:

//www.davidhbailey.com/dhbpapers/ising-int.pdf.

Borwein:2006:TTG

[BB06c] Jonathan M. Borwein and David M. Bradley. Thirty-two Goldbach variations. International Journal of Number Theory, 2(1): 65–103, March 2006. ISSN 1793-0421 (print), 1793-7310 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
301/;https://www.worldscientific.com/doi/10.1142/S1793042106000383.

Bailey:2007:SPb

[BB07a] David H. Bailey and Jonathan M. Borwein. Solution to monthly problem 11277. Report, Lawrence Berkeley National Laboratory and Faculty of Computer Science, Dalhousie University, Berkeley, CA 94720, USA and Halifax, NS, B3H 2W5, Canada, March 23,

2007. 3 pp. URL http://www.davidhbailey.com/dhbpapers/
amm-11277.pdf.

Bailey:2007:SPa

[BB07b] David H. Bailey and Jonathan M. Borwein. Solution to problem 11275. Report, Lawrence Berkeley National Laboratory and Faculty of Computer Science, Dalhousie University, Berkeley, CA 94720, USA and Halifax, NS, B3H 2W5, Canada, February 24,

2007. 2 pp. URL http://www.davidhbailey.com/dhbpapers/
amm-11275.pdf.

Borwein:2007:VPE

[BB07c] David Borwein and Jonathan Borwein. van der Pol expansions of L-series. Canadian mathematical bulletin = Bulletin canadien de mathématiques, 50(1):11–23, March 2007. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL http://

docserver.carma.newcastle.edu.au/283/.

Bailey:2008:HPC

[BB08a] D. H. Bailey and J. M. Borwein. High-precision computation and mathematical physics. In Proceedings of XII Advanced Computing and Analysis Techniques in Physics Research, page 14.

????, ????, 2008. URL http://adsabs.harvard.edu/abs/
2008acat.confE..14B;http://docserver.carma.newcastle.
edu.au/386/.

Bailey:2008:HPN

[BB08b] D. H. Bailey and J. M. Borwein. High-precision numerical integration: Progress and challenges. In MICA conference May 2008, pages 1–19. ????, ????, 2008. URL http://docserver.carma.

newcastle.edu.au/1444/.

Bailey:2008:CAD

[BB08c] David H. Bailey and Jonathan M. Borwein. Computer-assisted discovery and proof. In Tapas in experimental mathematics, volume 457 of Contemp. Math., pages 21–52. American Mathematical Society, Providence, RI, USA, 2008. URL http://docserver.

carma.newcastle.edu.au/338/;http://www.ams.org/books/
conm/457/8901.

Bailey:2008:EEB

[BB08d] David H. Bailey and Jonathan M. Borwein. Effective error bounds for Euler–Maclaurin-based quadrature schemes. In IEEE [IEE08], pages 1–7. ISBN 0-7695-2582-2. ISSN 1550-5243 (print), 2378-2099 (electronic). LCCN QA76.88 .I5858 2006.

Bailey:2008:HPH

[BB08e] David H. Bailey and Jonathan M. Borwein. Highly parallel, highprecision numerical integration. Report, Lawrence Berkeley National Laboratory and Faculty of Computer Science, Dalhousie University, Berkeley, CA 94720, USA and Halifax, NS, B3H 2W5, Canada, April 1, 2008. 18 pp. URL http://www.davidhbailey.

com/dhbpapers/quadparallel.pdf.

Bailey:2008:PSP

[BB08f] David H. Bailey and Jonathan M. Borwein. Proposed SIAM problem. Report, Lawrence Berkeley National Laboratory and Faculty of Computer Science, Dalhousie University, Berkeley, CA 94720, USA and Halifax, NS, B3H 2W5, Canada, August 12,

2008. 1–2 pp. URL http://www.davidhbailey.com/dhbpapers/
siam-bessel.pdf.

Borwein:2008:WEM

[BB08g] J. M. Borwein and D. H. Bailey. What is experimental mathematics. In Giovanni Treccani, editor, Enciclopedia della Scienza e della Tecnica, page ?? ????, Roma, Italia, 2008. Reprint of [BB04b, Chapter 1].

Borwein:2008:ME

[BB08h] Jonathan Borwein and David Bailey. Mathematics by Experiment. A. K. Peters, Ltd., Wellesley, MA, USA, second edition, 2008. ISBN 1-56881-442-9, 1-4398-6536-1 (PDF e-book). xii + 377 pp. LCCN QA76.95 .B67 2008. URL http://docserver. carma.newcastle.edu.au/272/. Plausible reasoning in the 21st Century.

Bailey:2009:EMCb

[BB09a] D. H. Bailey and J. M. Borwein. Experimental mathematics and computational statistics. In Online Encyclopedia of Computational Statistics. A Wiley Interdisciplinary Review, page ?? Wiley, New York, NY, USA, July 2009. ISBN 0-470-97388-9 (hardcover). LCCN GE45.S73 E56 2012. URL http://ca.wiley.

com/WileyCDA/Section/id-390445.html;http://docserver.
carma.newcastle.edu.au/1227/;http://www.wiley.com/WileyCDA/

WileyTitle/productCd-0470973889.html. Six volumes (lix +

3150 pages).

Bailey:2009:HPH

[BB09b] David Bailey and Jonathan Borwein. Highly parallel, high precision integration. International Journal of Computational Science and Engineering, 4(??):??, ???? 2009. ISSN 1742-7185 (print), 1742-7193 (electronic). URL http://docserver.carma.

newcastle.edu.au/294/.

Bailey:2009:BRW

[BB09c] David H. Bailey and Jonathan M. Borwein. Book review: Why Beliefs Matter: Reflections on the Nature of Science. Math

Drudge, October 12, 2009. URL https://experimentalmath.
info/blog/2012/04/book-review-why-beliefs-matter-reflections-on-the-nature-of-science/.

Bailey:2009:ECR

[BB09d] David H. Bailey and Jonathan M. Borwein. Einstein on the ‘cosmic religious feeling’ as motive for scientific research. Math

Drudge, June 10, 2009. URL https://experimentalmath.info/
blog/2009/06/einstein/.

Bailey:2009:EMCa

[BB09e] David H. Bailey and Jonathan M. Borwein. Experimental mathematics and computational statistics. Wiley Interdisciplinary Reviews: Computational Statistics, 1(1):12–24, July–August 2009. CODEN ???? ISSN 1939-0068 (print), 1939-5108 (electronic).

URL http://docserver.carma.newcastle.edu.au/1227/.

Bailey:2009:HDE

[BB09f] David H. Bailey and Jonathan M. Borwein. How did the economists get it so wrong. Math Drudge, September 12,

2009. URL https://experimentalmath.info/blog/2009/09/
how-did-the-economists-get-it-so-wrong/.

Bailey:2009:JDB

[BB09g] David H. Bailey and Jonathan M. Borwein. John D. Barrow’s New Theories of Everything. Math Drudge, October 12,

2009. URL https://experimentalmath.info/blog/2009/10/
john-d-barrows-new-theories-of-everything/.

Bailey:2009:MPC

[BB09h] David H. Bailey and Jonathan M. Borwein. Misuse of probability by ‘creation scientists’. Math Drudge, August 13,

2009. URL https://experimentalmath.info/blog/2009/08/
misuse-of-probability-by-creation-scientists-and-others/.

Bailey:2009:NRR

[BB09i] David H. Bailey and Jonathan M. Borwein. Numeracy, relative risk and public policy. Math Drudge, June 25,

2009. URL https://experimentalmath.info/blog/2009/07/
innumeracy-and-assessment-of-relative-risk/.

Bailey:2009:PAD

[BB09j] David H. Bailey and Jonathan M. Borwein. PSLQ: an algorithm to discover to integer relations. Computeralgebra Rundbrief, 45(??):8–11, October 2009. ISSN 0933-5994. URL http://

www.fachgruppe-computeralgebra.de/cms/tiki-index.php?
page=Rundbrief.

Bailey:2009:SFW

[BB09k] David H. Bailey and Jonathan M. Borwein. Semiotic fiddling while a digital Rome burns. Math Drudge, June 23, 2009. URL

https://experimentalmath.info/blog/2009/06/50/.

Bailey:2009:SMPa

[BB09l] David H. Bailey and Jonathan M. Borwein. Solution to Monthly problem #11418. Report, Lawrence Berkeley National Laboratory and Faculty of Computer Science, Dalhousie University and School of Mathematical and Physical Sciences, University of Newcastle, Berkeley, CA 94720, USA and Halifax, NS, B3H 2W5, Canada and Callaghan, NSW 2308, Australia, July 6, 2009. URL

http://www.davidhbailey.com/dhbpapers/amm-11418.pdf.

Bailey:2009:SMPb

[BB09m] David H. Bailey and Jonathan M. Borwein. Solutions to Monthly problems 11456 and 11457. Report, Lawrence Berkeley National Laboratory and Faculty of Computer Science, Dalhousie University, Berkeley, CA 94720, USA and Halifax, NS, B3H 2W5, Canada, October 10, 2009. 2 pp. URL http://www.

davidhbailey.com/dhbpapers/amm-11456.pdf.

Bailey:2009:UA

[BB09n] David H. Bailey and Jonathan M. Borwein. Unscientific America. Math Drudge, August 1, 2009. URL https://

experimentalmath.info/blog/2009/08/unscientific-america/.

Borwein:2009:HPH

[BB09o] Jonathan M. Borwein and D. H. Bailey. High precision, high dimension integration. Third International Workshop on High Dimensional Approximation, University of New South Wales, Sydney, Australia, February 16–20. Awarded most-entertaining presentation prize., February 20, 2009.

Bailey:2010:CMT

[BB10a] David H. Bailey and Jonathan M. Borwein. Can machines teach themselves. Math Drudge, October 10, 2010. URL https://

experimentalmath.info/blog/2010/10/can-machines-teach-themselves-2/.

Bailey:2010:CMC

[BB10b] David H. Bailey and Jonathan M. Borwein. The confusing morass of copyright laws. Math Drudge, January 28,

2010. URL https://experimentalmath.info/blog/2010/01/
the-confusing-morass-of-copyright-laws/.

Bailey:2010:CGW

[BB10c] David H. Bailey and Jonathan M. Borwein. Creationism, global warming denial, and scientific integrity. Math Drudge, March 7,

2010. URL https://experimentalmath.info/blog/2010/03/
creationism-global-warming-denial-and-scientific-integrity/.

Bailey:2010:ECO

[BB10d] David H. Bailey and Jonathan M. Borwein. Experimental computation with oscillatory integrals. In Amdeberhan et al. [AMM10], pages 25–40. ISBN 0-8218-4869-0, 0-8218-8196-5 (e-book). LCCN

QA164 .A475 2009. URL http://docserver.carma.newcastle.
edu.au/391/.

Bailey:2010:FPS

[BB10e] David H. Bailey and Jonathan M. Borwein. Fermi’s Paradox and Stephen Hawking. Math Drudge, April 30, 2010. URL https://

experimentalmath.info/blog/2010/04/fermis-paradox-and-stephen-hawking/.

Bailey:2010:GMD

[BB10f] David H. Bailey and Jonathan M. Borwein. The greatest mathematical discovery. Math Drudge, February 6,

2010. URL https://experimentalmath.info/blog/2010/02/
the-greatest-mathematical-discovery/.

Bailey:2010:HRR

[BB10g] David H. Bailey and Jonathan M. Borwein. How reliable are the radiometric methods used for geologic ages. Math Drudge, May

14, 2010. URL https://experimentalmath.info/blog/2010/
05/how-reliable-are-the-radiometric-methods-used-for-geologic-ages/.

Bailey:2010:LRI

[BB10h] David H. Bailey and Jonathan M. Borwein. Latest research indicates we unconsciously pursue goals. Math Drudge, July 4,

2010. URL https://experimentalmath.info/blog/2010/07/
latest-research-indicates-we-unconsciously-pursue-goals/.

Bailey:2010:PTS

[BB10i] David H. Bailey and Jonathan M. Borwein. Political threats to science funding. Math Drudge, September 12, 2010. URL https://

experimentalmath.info/blog/2010/09/political-threats-to-science-funding/.

Bailey:2010:SSM

[BB10j] David H. Bailey and Jonathan M. Borwein. Sad state of math and science education. Math Drudge, January 7,

2010. URL https://experimentalmath.info/blog/2010/01/
sad-state-of-math-and-science-education/.

Bailey:2010:SMP

[BB10k] David H. Bailey and Jonathan M. Borwein. Solution to Monthly problem 11515. Report, Lawrence Berkeley National Laboratory and School of Mathematical and Physical Sciences, University of Newcastle, Berkeley, CA 94720, USA and Callaghan, NSW 2308, Australia, June 15, 2010. 2 pp. URL http://www.davidhbailey.

com/dhbpapers/amm-11515.pdf.

Borwein:2010:ECM

[BB10l] Jonathan Borwein and Peter B. Borwein. Experimental and computational mathematics: selected writings. Perfectly Scientific Press, Portland, OR, USA, 2010. ISBN 1-935638-05-X. viii + 297 pp.

Bacak:2011:DCL

[BB11a] Miroslav Bačák and Jonathan M. Borwein. On difference convexity of locally Lipschitz functions. Optimization, 60(8–9):961–978, 2011. CODEN OPTZDQ. ISSN 0233-1934, 0323-3898. URL

http://docserver.carma.newcastle.edu.au/1446/.

Bailey:2011:HPN

[BB11b] D. H. Bailey and J. M. Borwein. High-precision numerical integration: Progress and challenges. Journal of Symbolic Computation, 46(7):741–754, July 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855x (electronic). URL http://docserver.

carma.newcastle.edu.au/1444/;http://www.sciencedirect.
com/science/article/pii/S0747717110001409. Special Issue

in Honour of Keith Geddes on his 60th Birthday.

Bailey:2011:DDW

[BB11c] David Bailey and Jonathan Borwein. Danger of death: are we programmed to miscalculate risk? The Conversation, ??(??):

??, December 8, 2011. URL https://theconversation.com/
danger-of-death-are-we-programmed-to-miscalculate-risk-4598.

Bailey:2011:HFEb

[BB11d] David Bailey and Jonathan Borwein. How far away is everybody? Climbing the cosmic ladder. The Conversation, ??(??):

??, October 24, 2011. URL https://theconversation.com/
how-far-away-is-everybody-climbing-the-cosmic-distance-ladder-3548.

Bailey:2011:MNC

[BB11e] David Bailey and Jonathan Borwein. Magic numbers: counting the blessings of decimal notation. The Conversation, ??

(??):??, August 23, 2011. URL https://theconversation.com/
magic-numbers-the-beauty-of-decimal-notation-2538.

Bailey:2011:WTD

[BB11f] David Bailey and Jonathan Borwein. When things don’t add up: statistics, maths and scientific fraud. The Conversation, ??(??):

??, November 13, 2011. URL https://theconversation.com/
when-things-dont-add-up-statistics-maths-and-scientific-fraud-4185.

Part of series on The State Of Science.

Bailey:2011:WED

[BB11g] David Bailey and Jonathan Borwein. Where is everybody? Doing the maths on extraterrestrial life. The Conversation, ??(??):

??, September 15, 2011. URL https://theconversation.com/
where-is-everybody-doing-the-maths-on-extraterrestrial-life-3390.

Bailey:2011:AIS

[BB11h] David H. Bailey and Jonathan M. Borwein. Ancient Indian square roots. Math Drudge, June 14, 2011. URL https://

experimentalmath.info/blog/2011/06/ancient-indian-square-roots/.

Bailey:2011:SMS

[BB11i] David H. Bailey and Jonathan M. Borwein. Are science and mathematics socially constructed. Math Drudge, May 29,

2011. URL https://experimentalmath.info/blog/2011/05/
are-science-and-mathematics-socially-constructed/.

Bailey:2011:EEC

[BB11j] David H. Bailey and Jonathan M. Borwein. Exploratory experimentation and computation. Notices of the American Mathematical Society, 58(10):1410–1419, November 2011. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL

http://docserver.carma.newcastle.edu.au/1396/;https://
www.ams.org/notices/201110/rtx111001410p.pdf. Chinese

translation in Mathematical Advances in Translation (Chinese Academy of Science), June 2012.

Bailey:2011:GDW

[BB11k] David H. Bailey and Jonathan M. Borwein. The great decline of Western society: What are the facts. Math Drudge, December 23,

2011. URL https://experimentalmath.info/blog/2011/12/
the-great-decline-of-western-society-what-are-the-facts/.

Bailey:2011:GMDa

[BB11l] David H. Bailey and Jonathan M. Borwein. The greatest mathematical discovery? Report, Lawrence Berkeley National Laboratory and Centre for Computer Assisted RMA, University of Newcastle, Berkeley, CA 94720, USA and Callaghan, NSW 2308, Australia, May 8, 2011. 10 pp. URL http://www.davidhbailey.

com/dhbpapers/decimal.pdf.

Bailey:2011:GMDb

[BB11m] David H. Bailey and Jonathan M. Borwein. The greatest mathematical discovery. Math Drudge, May 9, 2011. URL https://

experimentalmath.info/blog/2011/05/the-greatest-mathematical-discovery-2/.

Bailey:2011:HFEa

[BB11n] David H. Bailey and Jonathan M. Borwein. How far away is everybody. Math Drudge, September 29, 2011. URL https://

experimentalmath.info/blog/2011/09/how-far-away-is-everybody/.

Bailey:2011:IWV

[BB11o] David H. Bailey and Jonathan M. Borwein. IBM’s ‘Watson’ victorious: Our new computer overlords. Math Drudge, February

17, 2011. URL https://experimentalmath.info/blog/2011/
02/ibms-watson-victorious-our-new-computer-overlord/.

Bailey:2011:IPR

[BB11p] David H. Bailey and Jonathan M. Borwein. Innumeracy and public risk. Math Drudge, December 4, 2011. URL https://

experimentalmath.info/blog/2011/12/innumeracy-and-public-risk/.

Bailey:2011:MAI

[BB11q] David H. Bailey and Jonathan M. Borwein. Is math ability inborn or developed. Math Drudge, August 14, 2011. URL https://

experimentalmath.info/blog/2011/08/is-math-ability-inborn-or-developed/.

Bailey:2011:MID

[BB11r] David H. Bailey and Jonathan M. Borwein. Is mathematics invented or discovered. Math Drudge, August 7,

2011. URL https://experimentalmath.info/blog/2011/08/
is-mathematics-invented-or-discovered/.

Bailey:2011:MSF

[BB11s] David H. Bailey and Jonathan M. Borwein. Mathematics and scientific fraud. Math Drudge, November 3, 2011. URL https://

experimentalmath.info/blog/2011/11/mathematics-and-scientific-fraud/.

Bailey:2011:MD

[BB11t] David H. Bailey and Jonathan M. Borwein. Merchants of doubt. Math Drudge, July 17, 2011. URL https://experimentalmath.

info/blog/2011/07/merchants-of-doubt/.

Bailey:2011:PGF

[BB11u] David H. Bailey and Jonathan M. Borwein. Pi goes on forever. Math Drudge, March 9, 2011. URL https://experimentalmath.

info/blog/2011/03/pi-goes-on-forever/.

Bailey:2011:PBS

[BB11v] David H. Bailey and Jonathan M. Borwein. PIIGS, BRICs and STRAW. Math Drudge, July 23, 2011. URL https://

experimentalmath.info/blog/2011/07/piigs-brics-and-straw/.

Bailey:2011:PMJ

[BB11w] David H. Bailey and Jonathan M. Borwein. Proposed mathematical journal rating system. Math Drudge, November 20,

2011. URL https://experimentalmath.info/blog/2011/11/
proposed-mathematical-journal-rating-system/.

Bailey:2011:QTC

[BB11x] David H. Bailey and Jonathan M. Borwein. Quick tests for checking whether a new math result is plausible. Math Drudge, June 8,

2011. URL https://experimentalmath.info/blog/2011/06/
quick-tests-for-checking-whether-a-new-math-result-is-plausible/.

Bailey:2011:RDV

[BB11y] David H. Bailey and Jonathan M. Borwein. The remarkable decline of violence. Math Drudge, December 15, 2011. URL https:

//experimentalmath.info/blog/2011/12/the-remarkable-decline-of-violence/.

Bailey:2011:SFW

[BB11z] David H. Bailey and Jonathan M. Borwein. Semiotic fiddling while a digital Rome burns. Math Drudge, May 29,

2011. URL https://experimentalmath.info/blog/2011/05/
semiotic-fiddling-while-a-digital-rome-burns/.

Bailey:2011:MIN

[BB11-27] David H. Bailey and Jonathan M. Borwein. That mysterious but important number zero. Math Drudge, November 21,

2011. URL https://experimentalmath.info/blog/2011/11/
that-mysterious-but-important-number-zero/.

Bailey:2011:WDW

[BB11-28] David H. Bailey and Jonathan M. Borwein. What does Watson’s victory really mean. Math Drudge, February 20,

2011. URL https://experimentalmath.info/blog/2011/02/
what-does-watsons-victory-really-mean/.

Bailey:2011:WIB

[BB11-29] David H. Bailey and Jonathan M. Borwein. What if base-10 arithmetic had been discovered earlier. Math Drudge, July 27,

2011. URL https://experimentalmath.info/blog/2011/07/
what-if-base-10-arithmetic-had-been-discovered-earlier/.

Bailey:2011:WE

[BB11-30] David H. Bailey and Jonathan M. Borwein. Where is everybody. Math Drudge, September 9, 2011. URL https://

experimentalmath.info/blog/2011/09/where-is-everybody/.

Borwein:2011:PSE

[BB11-31] D. Borwein and Jonathan M. Borwein. Proof of some experimentally conjectured formulas for π. Preprint, Department of Mathematics, University of Western Ontario and Centre for Computerassisted Research Mathematics and its Applications (CARMA), School of Mathematical and Physical Sciences, University of Newcastle, London, ON, Canada and Callaghan, NSW 2308, Australia, December 4, 2011.

Bailey:2012:BNM

[BB12a] David Bailey and Jonathan Borwein. Bad numbers make for killer headlines – and dodgy news. The Conversation, ??

(??):??, June 26, 2012. URL https://theconversation.com/
bad-numbers-make-for-killer-headlines-and-dodgy-news-7894.

Bailey:2012:DDW

[BB12b] David Bailey and Jonathan Borwein. Doctor, doctor: why so few scientists in politics? The Conversation, ??(??):??,

February 29, 2012. URL https://theconversation.com/
doctor-doctor-why-so-few-scientists-in-top-government-jobs-5561.

Bailey:2012:HBU

[BB12c] David Bailey and Jonathan Borwein. Hot and bothered: the uncertain mathematics of global warming. The Conversation, ??(??):

??, February 17, 2012. URL https://theconversation.com/
hot-and-bothered-the-uncertain-mathematics-of-global-warming-5369.

Bailey:2012:HOEb

[BB12d] David Bailey and Jonathan Borwein. How old is Earth? A word to sceptics on the dating game. The Conversation, ??

(??):??, March 22, 2012. URL https://theconversation.com/
how-old-is-earth-a-word-to-sceptics-on-the-dating-game-5971.

Bailey:2012:HSG

[BB12e] David Bailey and Jonathan Borwein. How to sell green energy in an era of abundant gas and oil. The Conversation, ??

(??):??, June 15, 2012. URL https://theconversation.com/
how-to-sell-green-energy-in-an-era-of-abundant-gas-and-oil-7668.

Bailey:2012:HSM

[BB12f] David Bailey and Jonathan Borwein. How to stop the media reporting science fiction as fact. The Conversation, ??(??):

??, October 23, 2012. URL https://theconversation.com/
how-to-stop-the-media-reporting-science-fiction-as-fact-10252.

Bailey:2012:JCD

[BB12g] David Bailey and Jonathan Borwein. Just out of Curiosity, did life on Earth come from Mars? The Conversation, ??(??):

??, December 3, 2012. URL https://theconversation.com/
just-out-of-curiosity-did-life-on-earth-come-from-mars-11109.

Bailey:2012:MMD

[BB12h] David Bailey and Jonathan Borwein. Make mine a double: Moore’s Law and the future of mathematics. The Conversation, ??

(??):??, January 28, 2012. URL https://theconversation.com/
make-mine-a-double-moores-law-and-the-future-of-mathematics-4957.

Condensed and revised version of [BB12-39].

Bailey:2012:NCM

[BB12i] David Bailey and Jonathan Borwein. Newly calculated: maths anxiety triggers pain in the brain. The Conversation, ??(??):

??, November 2, 2012. URL https://theconversation.com/
newly-calculated-maths-anxiety-triggers-pain-in-the-brain-10453.

Bailey:2012:PCC

[BB12j] David Bailey and Jonathan Borwein. Person or computer: could you pass the Turing Test? The Conversation, ??

(??):??, May 3, 2012. URL https://theconversation.com/
person-or-computer-could-you-pass-the-turing-test-6769.

Bailey:2012:SMF

[BB12k] David Bailey and Jonathan Borwein. School maths is failing children — a US and Australian perspective. The Conversation,

??(??):??, July 25, 2012. URL https://theconversation.com/
school-maths-is-failing-children-a-us-and-australian-perspective-8397.

Bailey:2012:SMA

[BB12l] David Bailey and Jonathan Borwein. Smart meters are about as dangerous as …. The Conversation, ??(??):??,

October 23, 2012. URL https://theconversation.com/
smart-meters-are-about-as-dangerous-as-9413.

Bailey:2012:CPGb

[BB12m] David A. Bailey and Jonathan M. Borwein. Are computers playing games with us? Huffington Post, ??(??):??, April 2, 2012.

Bailey:2012:SOAb

[BB12n] David A. Bailey and Jonathan M. Borwein. 2001: A Space Odyssey: Art versus 2012 reality. Huffington Post, ??(??):

??, May 9, 2012. URL http://www.huffingtonpost.com/
david-h-bailey/2001-a-space-odyssey-art-_b_1501750.html.

Bailey:2012:ABS

[BB12o] David H. Bailey and Jonathan M. Borwein. Alarm bells sound over latest international test scores. Math Drudge, December 12,

2012. URL https://experimentalmath.info/blog/2012/12/
alarm-bells-sound-over-latest-international-test-scores/.

Bailey:2012:AECa

[BB12p] David H. Bailey and Jonathan M. Borwein. Algebra is essential in a 21st century economy. Math Drudge, July 31,

2012. URL https://experimentalmath.info/blog/2012/07/
algebra-is-essential-in-a-21st-century-economy/.

Bailey:2012:AECb

[BB12q] David H. Bailey and Jonathan M. Borwein. Algebra is essential in a 21st century economy. Huffington Post, ??(??):

??, August 1, 2012. URL http://www.huffingtonpost.com/
jonathan-m-borwein/.

Bailey:2012:AIS

[BB12r] David H. Bailey and Jonathan M. Borwein. Ancient Indian square roots: An exercise in forensic paleo-mathematics. American Mathematical Monthly, 119(8):646–657, October 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://docserver.carma.newcastle.edu.
au/1385/;https://www.jstor.org/stable/10.4169;https://
www.jstor.org/stable/pdfplus/10.4169/amer.math.monthly.
119.08.646.pdf.

Bailey:2012:CPGa

[BB12s] David H. Bailey and Jonathan M. Borwein. Are computers playing games with us. Math Drudge, March 27,

2012. URL https://experimentalmath.info/blog/2012/03/
are-computers-playing-games-with-us/.

Bailey:2012:BNB

[BB12t] David H. Bailey and Jonathan M. Borwein. Bad numbers are bad news. Math Drudge, June 24, 2012. URL https://

experimentalmath.info/blog/2012/06/bad-numbers-are-bad-news/.

Bailey:2012:SOAa

[BB12u] David H. Bailey and Jonathan M. Borwein. 2001: A Space Odyssey: Art versus 2012 reality. Math Drudge, May 31,

2012. URL https://experimentalmath.info/blog/2012/05/
2001-a-space-odyssey-art-versus-2012-reality/.

Bailey:2012:DMA

[BB12v] David H. Bailey and Jonathan M. Borwein. Does math anxiety trigger pain networks in the brain. Math Drudge, October 31,

2012. URL https://experimentalmath.info/blog/2012/10/
does-math-anxiety-trigger-pain-networks-in-the-brain/.

Bailey:2012:DPR

[BB12w] David H. Bailey and Jonathan M. Borwein. Does probability refute evolution. Math Drudge, January 22, 2012. URL https://

experimentalmath.info/blog/2012/01/does-probability-refute-evolution/.

Bailey:2012:ENP

[BB12x] David H. Bailey and Jonathan M. Borwein. Emmy Noether: pillar of 20th century mathematics and physics. Math Drudge, March

27, 2012. URL https://experimentalmath.info/blog/2012/
03/emmy-noether-pillar-of-20th-century-mathematics-and-physics/.

Bailey:2012:EEC

[BB12y] David H. Bailey and Jonathan M. Borwein. Exploration, experimentation and computation. International Advances in Mathematics, 31(1):1–14, January 2012. ISSN 1003-3092. URL http://

www.davidhbailey.com/dhbpapers/expexp-chinese.pdf. Chi-

nese translation of [BB11j].

Bailey:2012:EEM

[BB12z] David H. Bailey and Jonathan M. Borwein. Exploratory experimentation in mathematics: selected works. Perfectly Scientific Press, Portland, OR, USA, 2012. ISBN 1-935638-24-6. 364 pp. LCCN QA76.95 .B34 2012.

Bailey:2012:FFP

[BB12-27] David H. Bailey and Jonathan M. Borwein. Feast or famine? Promoting green energy in an era of abundant gas and oil. Math

Drudge, June 11, 2012. URL https://experimentalmath.info/
blog/2012/06/feast-or-famine-promoting-green-energy-in-an-era-of-abundant-gas-and-oil/.

Bailey:2012:GWD

[BB12-28] David H. Bailey and Jonathan M. Borwein. Global warming denial and scientific integrity. Math Drudge, February 10,

2012. URL https://experimentalmath.info/blog/2012/02/
global-warming-denial-and-scientific-integrity/.

Bailey:2012:HHC

[BB12-29] David H. Bailey and Jonathan M. Borwein. Hand-to-hand combat with thousand-digit integrals. Journal of Computational Science, 3(3):77–86, 2012. ISSN 1877-7503 (print), 1877-7511

(electronic). URL http://docserver.carma.newcastle.edu.
au/1393/;http://www.sciencedirect.com/science/article/

pii/S1877750310000773. Scientific Computation Methods and Applications.

Bailey:2012:HMD

[BB12-30] David H. Bailey and Jonathan M. Borwein. The heart of the matter: do scientific journalists need ground rules. Math Drudge, Oc-

tober 19, 2012. URL https://experimentalmath.info/blog/
2012/10/do-scientific-journalists-need-ground-rules/.

Bailey:2012:HOEa

[BB12-31] David H. Bailey and Jonathan M. Borwein. How old is the earth? Calculate it for yourself. Math Drudge, March 3,

2012. URL https://experimentalmath.info/blog/2012/03/
how-old-is-the-earth-calculate-it-for-yourself/.

Bailey:2012:BCCa

[BB12-32] David H. Bailey and Jonathan M. Borwein. Is believing in climate change ‘an insult to God’. Math Drudge, December 6,

2012. URL https://experimentalmath.info/blog/2012/12/
is-believing-in-climate-change-an-insult-to-god/.

Bailey:2012:BCCb

[BB12-33] David H. Bailey and Jonathan M. Borwein. Is believing in climate change “an insult to God”? Huffington Post, ??(??):

??, December 7, 2012. URL http://www.huffingtonpost.com/
david-h-bailey/climate-change-religion_b_2254016.html.

Bailey:2012:MSFa

[BB12-34] David H. Bailey and Jonathan M. Borwein. Is modern science ‘forever tentative’ and ‘socially constructed’? No way. Math Drudge,

May 31, 2012. URL https://experimentalmath.info/blog/
2012/05/is-modern-science-forever-tentative-and-socially-constructed-no-way/.

Bailey:2012:MSFb

[BB12-35] David H. Bailey and Jonathan M. Borwein. Is modern science ‘forever tentative’ and ‘socially constructed?’ No way! Huffington Post, ??(??):??, June 7, 2012. URL http://www.

huffingtonpost.com/jonathan-m-borwein/.

Bailey:2012:YMA

[BB12-36] David H. Bailey and Jonathan M. Borwein. Is your mate actually a computer? Would you pass the “Turing test”? Math Drudge,

April 27, 2012. URL https://experimentalmath.info/blog/
2012/04/will-computers-soon-pass-the-turing-test/.

Bailey:2012:LMM

[BB12-37] David H. Bailey and Jonathan M. Borwein. Life on Mars!? Maybe we are all Martians. Math Drudge, December 3,

2012. URL https://experimentalmath.info/blog/2012/12/
life-on-mars-maybe-we-are-all-martians/.

Bailey:2012:MPI

[BB12-38] David H. Bailey and Jonathan M. Borwein. Mathematician/ physicist/inventor Richard Crandall dies at 64. Math Drudge, De-

cember 21, 2012. URL https://experimentalmath.info/blog/
2012/12/mathematicianphysicistinventor-richard-crandall-dies-at-64/.

Bailey:2012:MLF

[BB12-39] David H. Bailey and Jonathan M. Borwein. Moore’s Law and the future of science and mathematics. Math Drudge, January 2,

2012. URL https://experimentalmath.info/blog/2012/01/
moores-law-and-the-future-of-science-and-mathematics/.

Condensed and revised version appears in [BB12h].

Bailey:2012:NSC

[BB12-40] David H. Bailey and Jonathan M. Borwein. Nonnormality of Stoneham constants. The Ramanujan Journal, 29(1–3):409–422, 2012. CODEN RAJOF9. ISSN 1382-4090 (print), 1572-9303

(electronic). URL http://docserver.carma.newcastle.edu.
au/1386/.

Bailey:2012:NCT

[BB12-41] David H. Bailey and Jonathan M. Borwein. Numeracy crisis threatens first-world economies. Math Drudge, March 4,

2012. URL https://experimentalmath.info/blog/2012/03/
numeracy-crisis-threatens-first-world-economies/.

Bailey:2012:NNUa

[BB12-42] David H. Bailey and Jonathan M. Borwein. Numerical nonsense in the U.S. presidential campaign. Math Drudge, August 17,

2012. URL https://experimentalmath.info/blog/2012/08/
numerical-nonsense-in-the-u-s-presidential-campaign/.

Bailey:2012:NNUb

[BB12-43] David H. Bailey and Jonathan M. Borwein. Numerical nonsense in the U.S. presidential campaign. Huffington Post, ??(??):

??, August 18, 2012. URL http://www.huffingtonpost.com/
jonathan-m-borwein/.

Bailey:2012:PQM

[BB12-44] David H. Bailey and Jonathan M. Borwein. Poor-quality math and computer science courses threaten technological leadership. Math Drudge, January 10, 2012. URL https://

experimentalmath.info/blog/2012/01/poor-quality-math-and-computer-science-courses-threaten-technological-leadership/.

Bailey:2012:SPW

[BB12-45] David H. Bailey and Jonathan M. Borwein. Scientists in politics: What is the score, and what can be done. Math Drudge, February

16, 2012. URL https://experimentalmath.info/blog/2012/
02/scientists-in-politics-what-is-the-score-and-what-can-be-done/.

Bailey:2012:SMD

[BB12-46] David H. Bailey and Jonathan M. Borwein. Smart meters for dummies. Math Drudge, September 6, 2012. URL https://

experimentalmath.info/blog/2012/09/smart-meters-for-dummies/.

Bailey:2012:SMP

[BB12-47] David H. Bailey and Jonathan M. Borwein. Solution to monthly problem 11650. Report, Lawrence Berkeley National Laboratory and Centre for Computer Assisted Research Mathematics and its Applications (CARMA), University of Newcastle and King Abdulaziz University, Berkeley, CA 94720, USA and Callaghan, NSW 2308, Australia and Jeddah 80200, Saudi Arabia., June 14,

2012. 2 pp. URL http://www.davidhbailey.com/dhbpapers/
amm-11650.pdf.

Bailey:2012:TPA

[BB12-48] David H. Bailey and Jonathan M. Borwein. Talking points for ‘algebra is essential in a 21st century economy’. Math Drudge,

August 9, 2012. URL https://experimentalmath.info/blog/
2012/08/talking-points-for-algebra-is-essential-in-a-21st-century-economy/.

Bailey:2012:UPK

[BB12-49] David H. Bailey and Jonathan M. Borwein. The uneven preparation of K–12 math teachers. Math Drudge, July 19,

2012. URL https://experimentalmath.info/blog/2012/07/
the-uneven-preparation-of-k-12-math-teachers/.

Bailey:2012:WDL

[BB12-50] David H. Bailey and Jonathan M. Borwein. What does the latest DNA data say about evolution. Math Drudge, May 1,

2012. URL https://experimentalmath.info/blog/2012/05/
what-does-the-latest-dna-data-say-about-evolution/.

Bailey:2012:WED

[BB12-51] David H. Bailey and Jonathan M. Borwein. What on earth do they think? U.S. politicians on the age of the planet. Math Drudge,

November 19, 2012. URL https://experimentalmath.info/
blog/2012/11/what-on-earth-do-they-think-u-s-politicians-on-the-age-of-the-planet/.

Bailey:2012:WEW

[BB12-52] David H. Bailey and Jonathan M. Borwein. What on earth were they thinking? U.S. politicians on the age of the planet. Huffington Post, ??(??):??, December 4, 2012. URL http://www.

huffingtonpost.com/jonathan-m-borwein/.

Borwein:2012:YTN

[BB12-53] Jonathan M. Borwein and David H. Bailey. Yes, there’s a numeracy crisis — so what’s the solution? The Conversation, ??

(??):??, April 12, 2012. URL https://theconversation.com/
yes-theres-a-numeracy-crisis-so-whats-the-solution-6386.

Bailey:2013:APJ

[BB13a] David H. Bailey and Jonathan M. Borwein. The Abel Prize on Jeopardy. Math Drudge, May 10, 2013. URL https://

experimentalmath.info/blog/2013/05/the-abel-prize-on-jeopardy/.

Bailey:2013:DPRa

[BB13b] David H. Bailey and Jonathan M. Borwein. Are the digits of pi random. Math Drudge, April 15, 2013. URL https://

experimentalmath.info/blog/2013/04/are-the-digits-of-pi-random/.

Bailey:2013:DPRb

[BB13c] David H. Bailey and Jonathan M. Borwein. Are the digits of pi random? Huffington Post, ??(??):??, April 16, 2013. URL http:

//www.huffingtonpost.com/david-h-bailey/are-the-digits-of-pi-random_
b_3085725.html.

Bailey:2013:BSF

[BB13d] David H. Bailey and Jonathan M. Borwein. Brown, Sokal and Friedman on nonsense in psychology. Math Drudge, July 18,

2013. URL https://experimentalmath.info/blog/2013/07/
sokal-on-nonsense-in-psychology/.

Bailey:2013:CLFa

[BB13e] David H. Bailey and Jonathan M. Borwein. The colorful life of the four-color theorem: A tribute to Kenneth Appel. Math Drudge,

May 7, 2013. URL https://experimentalmath.info/blog/
2013/05/the-colorful-life-of-the-four-color-theorem-a-tribute-to-kenneth-appel/.

Bailey:2013:CLFb

[BB13f] David H. Bailey and Jonathan M. Borwein. The colorful life of the Four-Color Theorem: A tribute to Kenneth Appel. Huffington Post, ??(??):??, May 8, 2013. URL http://www.

huffingtonpost.com/david-h-bailey/kenneth-appel-four-color-theorem_
b_3233775.html.

Bailey:2013:CLS

[BB13g] David H. Bailey and Jonathan M. Borwein. Compressed lattice sums arising from the Poisson equation. Boundary Value Problems, 75(1):1–18, 2013. ISSN 1687-2770. URL http:

//boundaryvalueproblems.springeropen.com/articles/10.
1186/1687-2770-2013-75;http://docserver.carma.newcastle.
edu.au/1476/;http://www.boundaryvalueproblems.com/content/

2013/1/75. Special volume in honour of Hari Srivastava.

Bailey:2013:CSD

[BB13h] David H. Bailey and Jonathan M. Borwein. Criminology, sports drug testing and evolution. Math Drudge, April 1,

2013. URL https://experimentalmath.info/blog/2013/04/
criminology-sports-drug-testing-and-evolution/.

Bailey:2013:DYS

[BB13i] David H. Bailey and Jonathan M. Borwein. Danger, you’re at serious risk of … no, sorry, it’s all relative. The Conversation, ??(??):

??, February 15, 2013. URL https://theconversation.com/
danger-youre-at-serious-risk-of-no-sorry-its-all-relative-12218.

Bailey:2013:DDL

[BB13j] David H. Bailey and Jonathan M. Borwein. Did dinosaurs live with humans? Were dragons real. Math Drudge, August 27,

2013. URL https://experimentalmath.info/blog/2013/08/
did-dinosaurs-live-with-humans-were-dragons-real/.

Bailey:2013:FWJ

[BB13k] David H. Bailey and Jonathan M. Borwein. Frank W. J. Olver (1924–2013). Math Drudge, May 12, 2013. URL https://

experimentalmath.info/blog/2013/05/frank-w-j-olver-1924-2013/.

Bailey:2013:FFE

[BB13l] David H. Bailey and Jonathan M. Borwein. Fraud, foolishness and error in scientific research. Math Drudge, May 4,

2013. URL https://experimentalmath.info/blog/2013/05/
fraud-foolishness-and-error-in-scientific-research/.

Bailey:2013:GIW

[BB13m] David H. Bailey and Jonathan M. Borwein. Getting it wrong: Australian science literacy hits new low. Math Drudge, July 17,

2013. URL https://experimentalmath.info/blog/2013/07/
australian-science-literacy-hits-new-low/.

Bailey:2013:GGA

[BB13n] David H. Bailey and Jonathan M. Borwein. Glum and glummer: Australia vs US on science literacy results. The Conversation,

??(??):??, July 19, 2013. URL https://theconversation.com/
glum-and-glummer-australia-vs-us-on-science-literacy-results-16222.

Bailey:2013:HBW

[BB13o] David H. Bailey and Jonathan M. Borwein. The Higgs boson: What does it mean. Math Drudge, March 18,

2013. URL https://experimentalmath.info/blog/2013/03/
the-higgs-boson-what-does-it-mean/.

Bailey:2013:HHL

[BB13p] David H. Bailey and Jonathan M. Borwein. A Higgs, the Higgs … is maths at the root of reality? The Conversation, ??

(??):??, March 21, 2013. URL https://theconversation.com/
a-higgs-the-higgs-is-maths-at-the-root-of-reality-12943.

Bailey:2013:HPA

[BB13q] David H. Bailey and Jonathan M. Borwein. High-precision arithmetic: Opportunities and challenges. Preprint., August 2013.

Bailey:2013:HNHa

[BB13r] David H. Bailey and Jonathan M. Borwein. Hype now, hide later: No way to do scientific research. Math Drudge, May 26,

2013. URL https://experimentalmath.info/blog/2013/05/
hype-now-hide-later-no-way-to-do-scientific-research/.

Bailey:2013:HNHb

[BB13s] David H. Bailey and Jonathan M. Borwein. Hype now, hide later: No way to do scientific research. Huffington Post, ??

(??):??, May 28, 2013. URL http://www.huffingtonpost.com/
david-h-bailey/scientific-research_b_3340682.html.

Bailey:2013:LJW

[BB13t] David H. Bailey and Jonathan M. Borwein. The last Japanese WWII holdout: A lesson for creationists. Math Drudge, April 1,

2013. URL https://experimentalmath.info/blog/2013/04/
the-last-japanese-wwii-holdout-a-lesson-for-creationists/.

Bailey:2013:MPS

[BB13u] David H. Bailey and Jonathan M. Borwein. The mad politics of science funding. Math Drudge, May 8, 2013. URL https://

experimentalmath.info/blog/2013/05/the-politics-of-science-funding/.

Bailey:2013:MBP

[BB13v] David H. Bailey and Jonathan M. Borwein. March 26, 2013: The 100th birthday of Paul Erdős. Math Drudge, March 25,

2013. URL https://experimentalmath.info/blog/2013/03/
march-26-2013-the-100th-birthday-of-paul-erdos/.

Bailey:2013:MLW

[BB13w] David H. Bailey and Jonathan M. Borwein. Massachusetts leads the way in science and math education. Math Drudge, September

3, 2013. URL https://experimentalmath.info/blog/2013/
09/massachusetts-leads-the-way-in-science-and-math-education/.

Bailey:2013:NNP

[BB13x] David H. Bailey and Jonathan M. Borwein. Normal numbers and pseudorandom generators. In Bailey et al. [BBB^{+}13], pages 1–

18. ISBN 1-4614-7620-8, 1-4614-7621-6 (e-book). ISSN 2194-1009.

LCCN QA241. URL http://docserver.carma.newcastle.edu.
au/1448/.

Bailey:2013:PDU

[BB13y] David H. Bailey and Jonathan M. Borwein. Pi day is upon us again and we still do not know if pi is normal. Report, Lawrence Berkeley National Laboratory and Centre for Computer Assisted Research Mathematics and its Applications (CARMA), University of Newcastle, Berkeley, CA 94720, USA and Callaghan, NSW 2308, Australia, May 29, 2013. 20 pp. URL http://www.carma.

newcastle.edu.au/jon/pi-monthly.pdf.

Bailey:2013:PS

[BB13z] David H. Bailey and Jonathan M. Borwein. Pi in the Simpsons. Math Drudge, November 18, 2013. URL https://

experimentalmath.info/blog/2013/11/pi-in-the-simpsons/.

Bailey:2013:PIT

[BB13-27] David H. Bailey and Jonathan M. Borwein. PISA international test scores show Australia, Canada, UK, USA lagging. Math

Drudge, December 6, 2013. URL https://experimentalmath.
info/blog/2013/12/pisa-international-test-scores-show-australia-canada-uk-usa-lagging/.

Bailey:2013:PSD

[BB13-28] David H. Bailey and Jonathan M. Borwein. Plagiarism is a symptom not a disease. Math Drudge, May 31,

2013. URL https://experimentalmath.info/blog/2013/05/
plagiarism-is-a-symptom-not-a-disease/.

Bailey:2013:PMTa

[BB13-29] David H. Bailey and Jonathan M. Borwein. Please mess with Texas: Texas textbook fiasco threatens US science. Math Drudge,

September 15, 2013. URL https://experimentalmath.info/
blog/2013/09/texas-textbook-fiasco-threatens-us-science/.

Bailey:2013:PMTb

[BB13-30] David H. Bailey and Jonathan M. Borwein. Please mess with Texas: Texas textbook fiasco threatens US science. Huffington Post, ??(??):??, September 18, 2013. URL http://www.

huffingtonpost.com/david-h-bailey/texas-textbooks-science_
b_3935203.html.

Bailey:2013:RRE

[BB13-31] David H. Bailey and Jonathan M. Borwein. The Reinhart–Rogoff error or how not to Excel at economics. The Conversation, ??

(??):??, April 23, 2013. URL https://theconversation.edu.
au/profiles/jon-borwein-101.

Bailey:2013:RRR

[BB13-32] David H. Bailey and Jonathan M. Borwein. Reliability, reproducibility and the Reinhart–Rogoff error. Math Drudge, April 18,

2013. URL https://experimentalmath.info/blog/2013/04/
reliability-reproducibility-and-the-reinhart-rogoff-error/.

Bailey:2013:SFS

[BB13-33] David H. Bailey and Jonathan M. Borwein. Scientific fraud, sloppy science yes, they happen. The Conversation, ??(??):

??, May 6, 2013. URL https://theconversation.com/
scientific-fraud-sloppy-science-yes-they-happen-13948.

Bailey:2013:SNR

[BB13-34] David H. Bailey and Jonathan M. Borwein. Scientific nonsense and relative risk. Math Drudge, February 13, 2013. URL https:

//experimentalmath.info/blog/2013/02/scientific-nonsense-and-relative-risk/.

Bailey:2013:SDOa

[BB13-35] David H. Bailey and Jonathan M. Borwein. Set the default to ‘open’: Reproducible science in the computer age. Math Drudge,

January 20, 2013. URL https://experimentalmath.info/
blog/2013/01/set-the-default-to-open-reproducible-science-in-the-computer-age/.

Bailey:2013:SDOb

[BB13-36] David H. Bailey and Jonathan M. Borwein. Set the default to “open”: Reproducible science in the computer age. Huffington Post, ??(??):??, February 7, 2013. URL http://www.

huffingtonpost.com/david-h-bailey/set-the-default-to-open-r_
b_2635850.html.

Bailey:2013:SMDa

[BB13-37] David H. Bailey and Jonathan M. Borwein. Smart meters, dumb science. Math Drudge, February 26, 2013. URL https://

experimentalmath.info/blog/2013/02/smart-meters-dumb-science/.

Bailey:2013:SMDb

[BB13-38] David H. Bailey and Jonathan M. Borwein. Smart meters, dumb science. The Conversation, ??(??):??, February 28, 2013. URL

http://www.huffingtonpost.com/david-h-bailey/smart-meters-dumb-science_
b_2768405.html.

Bailey:2013:SSF

[BB13-39] David H. Bailey and Jonathan M. Borwein. Stupid science funding decisions? Australia’s not the only dunce. The Conversation,

??(??):??, May 12, 2013. URL https://theconversation.com/
stupid-science-funding-decisions-australias-not-the-only-dunce-14087.

Bailey:2013:TTS

[BB13-40] David H. Bailey and Jonathan M. Borwein. Tipsy tottering, sunlight and the smell of coffee: it’s all random. Math Drudge, De-

cember 18, 2013. URL https://experimentalmath.info/blog/
2013/12/tipsy-tottering-sunlight-and-the-smell-of-coffee-its-all-random/.

Bailey:2013:TBK

[BB13-41] David H. Bailey and Jonathan M. Borwein. Troubles beset Kentucky’s Creation Museum. Math Drudge, August 27,

2013. URL https://experimentalmath.info/blog/2013/08/
troubles-beset-kentuckys-creation-museum/.

Bailey:2013:TBR

[BB13-42] David H. Bailey and Jonathan M. Borwein. Two breakthrough results in number theory. Math Drudge, May 24,

2013. URL https://experimentalmath.info/blog/2013/05/
two-breakthrough-results-in-number-theory/.

Bailey:2013:WSBa

[BB13-43] David H. Bailey and Jonathan M. Borwein. When skepticism becomes denial: The unholy alliance between science denial movements. Math Drudge, November 3, 2013. URL https://

experimentalmath.info/blog/2013/11/when-skepticism-becomes-denial-the-unholy-alliance-between-science-denial-movements/.

Bailey:2013:WSBb

[BB13-44] David H. Bailey and Jonathan M. Borwein. When skepticism becomes denial: The unholy alliance between science denial movements. Huffington Post, ??(??):??, November 5, 2013. URL http:

//www.huffingtonpost.com/david-h-bailey/when-skepticism-becomes-d_
b_4215286.html.

Bailey:2013:WWWa

[BB13-45] David H. Bailey and Jonathan M. Borwein. Why E. O. Wilson is wrong. Math Drudge, April 17, 2013. URL https://

experimentalmath.info/blog/2013/04/why-e-o-wilson-is-wrong/.

Bailey:2013:WWWb

[BB13-46] David H. Bailey and Jonathan M. Borwein. Why E. O. Wilson is wrong. Huffington Post, ??(??):??, April 17, 2013. URL http://

www.huffingtonpost.com/david-h-bailey/why-eo-wilson-is-wrong_
b_3103122.html.

Bailey:2013:YWF

[BB13-47] David H. Bailey and Jonathan M. Borwein. You wait forever for number theory results … then two come along at once. The Conversation, ??(??):??, June 3, 2013. URL https:

//theconversation.edu.au/profiles/jon-borwein-101.

Bailey:2014:IWS

[BB14a] D. H. Bailey and J. M. Borwein. ICERM Workshop sets out opportunities and challenges in experimental mathematics. SIAM News, 47(9):4, 8, ???? 2014. ISSN 0036-1437.

URL http://www.davidhbailey.com/dhbpapers/icerm-summ.
pdf;https://sinews.siam.org/Portals/Sinews2/Issue%20Pdfs/
sn_November2014.pdf.

Bailey:2014:RTP

[BB14b] D. H. Bailey and J. M. Borwein. The recent trademarking of pi: a troubling precedent. Notices of the American Mathematical Society, 61(10):121–124, November 2014. CODEN AMNOAN.

ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

docserver.carma.newcastle.edu.au/id/eprint/1501;http:
//www.ams.org/notices/201410/rnoti-p1224.pdf.

Bailey:2014:PDUa

[BB14c] David H. Bailey and Jonathan Borwein. Pi Day is upon us again and we still do not know if pi is normal. American Mathematical Monthly, 121(3):191–206, March 2014. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.121.03.191.pdf. Chinese translation

in Mathematical Advances in Translation (Chinese Academy of Science) [metadata unknown].

Bailey:2014:OBH

[BB14d] David H. Bailey and Jonathan M. Borwein. Are our brains hardwired for numbers. Math Drudge, January 14, 2014. URL https:

//experimentalmath.info/blog/2014/01/are-our-brains-hard-wired-for-numbers/.

Bailey:2014:BBB

[BB14e] David H. Bailey and Jonathan M. Borwein. Big bucks for big breakthroughs: Prize recipients give three million dollar maths talks. Math Drudge, November 11, 2014. URL https://

experimentalmath.info/blog/2014/11/breakthrough-prize-recipients-give-math-seminar-talks/.

Bailey:2014:CPP

[BB14f] David H. Bailey and Jonathan M. Borwein. Can pi be trademarked? Huffington Post, ??(??):??, June 20, 2014. URL http://

www.huffingtonpost.com/david-h-bailey/can-pi-be-trademarked_
b_5513392.html.

Bailey:2014:CPT

[BB14g] David H. Bailey and Jonathan M. Borwein. Can pi be trademarked. Math Drudge, June 11, 2014. URL https://

experimentalmath.info/blog/2014/06/can-pi-be-trademarked-2/.

Bailey:2014:DDDa

[BB14h] David H. Bailey and Jonathan M. Borwein. Dubious digits: Is this data really that accurate. Math Drudge, November 8,

2014. URL https://experimentalmath.info/blog/2014/11/
dubious-digits-is-this-data-really-that-accurate/.

Bailey:2014:DDDc

[BB14i] David H. Bailey and Jonathan M. Borwein. Dubious digits: Is this data really that accurate? Huffington Post, ??(??):??,

November 10, 2014. URL http://www.huffingtonpost.com/
david-h-bailey/dubious-digits-is-this-da_b_6129864.html.

Bailey:2014:FPC

[BB14j] David H. Bailey and Jonathan M. Borwein. Formal proof completed for Kepler’s conjecture on sphere packing. Math Drudge,

August 16, 2014. URL https://experimentalmath.info/blog/
2014/08/formal-proof-completed-for-keplers-conjecture-on-sphere-packing/.

Bailey:2014:FEHa

[BB14k] David H. Bailey and Jonathan M. Borwein. Fusion energy: Hope or hype. Math Drudge, October 18, 2014. URL https://

experimentalmath.info/blog/2014/10/fusion-energy-hope-or-hype/.

Bailey:2014:FEHb

[BB14l] David H. Bailey and Jonathan M. Borwein. Fusion energy: Hope or hype? Huffington Post, ??(??):??, Oc-

tober 23, 2014. URL http://www.huffingtonpost.com/
david-h-bailey/fusion-energy-hope-or-hype_b_6031968.
html.

Bailey:2014:GWC

[BB14m] David H. Bailey and Jonathan M. Borwein. Gravitational waves confirm mathematical prediction of inflationary big bang. Math

Drudge, March 18, 2014. URL https://experimentalmath.
info/blog/2014/03/gravitational-waves-confirm-mathematical-prediction-of-inflationary-big-bang/.

Bailey:2014:HFLb

[BB14n] David H. Bailey and Jonathan M. Borwein. How financially literate is the investing public? Huffington Post, ??(??):

??, July 29, 2014. URL http://www.huffingtonpost.com/
david-h-bailey/how-financially-literate-investing-public_
b_5625649.html.

Bailey:2014:PNM

[BB14o] David H. Bailey and Jonathan M. Borwein. Is philosophy needed in mathematics and science. Math Drudge, May 13,

2014. URL https://experimentalmath.info/blog/2014/03/
is-philosophy-needed-in-mathematics-and-science/.

Bailey:2014:LENa

[BB14p] David H. Bailey and Jonathan M. Borwein. Low energy nuclear reactions: Papers and patents. Math Drudge, November

19, 2014. URL https://experimentalmath.info/blog/2014/
11/low-energy-nuclear-reactions-papers-and-patents/.

Bailey:2014:LENb

[BB14q] David H. Bailey and Jonathan M. Borwein. Low energy nuclear reactions: Papers and patents. Huffington Post, ??(??):??,

November 28, 2014. URL http://www.huffingtonpost.com/
david-h-bailey/low-energy-nuclear-reacti_b_6189772.html.

Bailey:2014:MTO

[BB14r] David H. Bailey and Jonathan M. Borwein. Max Tegmark’s Our Mathematical Universe. Math Drudge, January 31,

2014. URL https://experimentalmath.info/blog/2014/01/
max-tegmarks-our-mathematical-universe/.

Bailey:2014:NRP

[BB14s] David H. Bailey and Jonathan M. Borwein. New results on the prime gap conjecture. Math Drudge, December 11,

2014. URL https://experimentalmath.info/blog/2014/12/
new-results-on-the-prime-gap-conjecture/.

Bailey:2014:OCEa

[BB14t] David H. Bailey and Jonathan M. Borwein. Opportunities and challenges in experimental mathematics. Math Drudge, August

26, 2014. URL https://experimentalmath.info/blog/2014/
08/opportunities-and-challenges-in-experimental-mathematics/.

Bailey:2014:PDa

[BB14u] David H. Bailey and Jonathan M. Borwein. Pi day 3.14 (14). Math Drudge, February 23, 2014. URL https://experimentalmath.

info/blog/2014/02/pi-day-3-14-14/.

Bailey:2014:PDb

[BB14v] David H. Bailey and Jonathan M. Borwein. Pi day 3.14 (14). Huffington Post, ??(??):??, March 15, 2014. URL http://www.

huffingtonpost.com/david-h-bailey/pi-day-314-14_b_4851011.
html.

Bailey:2014:PDUb

[BB14w] David H. Bailey and Jonathan M. Borwein. Pi Day is upon us again, and we still do not know if pi is normal. Mathematical Advances of Translation, 33(??):222–235, ???? 2014. Chinese translation of [BB14c].

Bailey:2014:SDJ

[BB14x] David H. Bailey and Jonathan M. Borwein. The significance of digits: just how reliable are reported numbers? The Conversation, ??(??):??, November 27, 2014. URL https://theconversation.

com/the-significance-of-digits-just-how-reliable-are-reported-numbers-34295.

Bailey:2014:TW

[BB14y] David H. Bailey and Jonathan M. Borwein. Tilting at windmills. Math Drudge, September 17, 2014. URL https://

experimentalmath.info/blog/2014/09/tilting-at-windmills/.

Bailey:2014:FFQ

[BB14z] David H. Bailey and Jonathan M. Borwein. To frack or not to frack: That is the question. Huffington Post, ??(??):??,

September 9, 2014. URL http://www.huffingtonpost.com/
david-h-bailey/to-frack-or-not-to-frack-that-is-the-question_
b_5781634.html.

Bailey:2014:FFT

[BB14-27] David H. Bailey and Jonathan M. Borwein. To frack or not to frack: That’s not the question. Math Drudge, September 7,

2014. URL https://experimentalmath.info/blog/2014/09/
to-frack-or-not-to-frack-thats-not-the-question/.

Bailey:2014:WCG

[BB14-28] David H. Bailey and Jonathan M. Borwein. We still can’t get enough pi … but why? The Conversation, ??(??):

??, March 13, 2014. URL https://theconversation.com/
we-still-cant-get-enough-pi-but-why-23960.

Bailey:2014:WSP

[BB14-29] David H. Bailey and Jonathan M. Borwein. When science and philosophy collide in a ‘fine-tuned’ universe. The Conversation, ??(??):??, April 3, 2014.

Bailey:2014:WMBa

[BB14-30] David H. Bailey and Jonathan M. Borwein. Why mathematics is beautiful and why that matters. Math Drudge, February 14,

2014. URL https://experimentalmath.info/blog/2014/02/
why-mathematics-is-beautiful-and-why-it-matters/.

Bailey:2014:WMBb

[BB14-31] David H. Bailey and Jonathan M. Borwein. Why mathematics is beautiful and why that matters. Huffington Post, ??(??):

??, February 18, 2014. URL http://www.huffingtonpost.com/
david-h-bailey/why-mathematics-matters_b_4794617.html.

Borwein:2014:DNS

[BB14-32] David Borwein and Jonathan Borwein. Deriving new sinc results from old. American Mathematical Monthly, 121(8):700–705, October 2014. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://docserver.carma.newcastle.edu.
au/1502/;https://www.jstor.org/stable/10.4169/amer.math.
monthly.121.08.700.

Bailey:2015:CTM

[BB15a] D. H. Bailey and J. M. Borwein. Computation and theory of Mordell–Tornheim–Witten sums II. Journal of Approximation Theory, 197(??):115–140, September 2015. CODEN JAX-TAZ. ISSN 0021-9045 (print), 1096-0430 (electronic). URL

http://docserver.carma.newcastle.edu.au/1494/;http://
www.carma.newcastle.edu.au/jon/MTW2.pdf;http://www.sciencedirect.
com/science/article/pii/S0021904514001841. Special Issue

Dedicated to Dick Askey on the occasion of his 80th birthday.

Bailey:2015:COG

[BB15b] D. H. Bailey and J. M. Borwein. Computation as an ontological game changer: The impact of modern mathematical computation tools on the ontology of mathematics. In Davis and Davis [DD15], pages 25–67. ISBN 3-319-21472-1, 3-319-21473-X (e-book). LCCN

QA8.4 .M38 2015. URL http://link.springer.com/chapter/
10.1007/978-3-319-21473-3_3.

Bailey:2015:CCI

[BB15c] D. H. Bailey and J. M. Borwein. Crandall’s computation of the incomplete Gamma function and the Hurwitz zeta function, with applications to Dirichlet L-series. Applied Mathematics and Computation, 268(??):462–477, October 1, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0096300315008292.

Bailey:2015:AAF

[BB15d] David H. Bailey and Jonathan M. Borwein. Amir Aczel’s Finding Zero. Math Drudge, January 13, 2015. URL https://

experimentalmath.info/blog/2015/01/amir-aczels-finding-zero/.

Bailey:2015:TML

[BB15e] David H. Bailey and Jonathan M. Borwein. Are there ‘missing links’ in the human family tree. Math Drudge, June 2,

2015. URL https://experimentalmath.info/blog/2015/06/
are-there-missing-links-in-the-human-family-tree/.

Bailey:2015:CFHa

[BB15f] David H. Bailey and Jonathan M. Borwein. Cold fusion heats up: Fusion energy and LENR update. Math Drudge, August 27,

2015. URL https://experimentalmath.info/blog/2015/08/
cold-fusion-heats-up-fusion-energy-and-lenr-update/.

Bailey:2015:CFHb

[BB15g] David H. Bailey and Jonathan M. Borwein. Cold fusion heats up: Fusion energy and LENR update. Huffington Post, ??(??):

??, August 27, 2015. URL http://www.huffingtonpost.com/
david-h-bailey/post_10010_b_8052326.html.

Bailey:2015:DVTa

[BB15h] David H. Bailey and Jonathan M. Borwein. Data vs. theory: the mathematical battle for the soul of physics. Math Drudge, De-

cember 27, 2015. URL https://experimentalmath.info/blog/
2015/12/data-vs-theory-the-mathematical-battle-for-the-soul-of-physics/.

Bailey:2015:DVTb

[BB15i] David H. Bailey and Jonathan M. Borwein. Data vs theory: The mathematical battle for the soul of physics. Huffington Post, ?? (??):??, December 30, 2015. URL http://www.huffingtonpost.

com/david-h-bailey/data-vs-theory-the-mathem_b_8886292.
html.

Bailey:2015:DSFa

[BB15j] David H. Bailey and Jonathan M. Borwein. Desperately seeking ET: Fermi’s paradox turns 65 (Part II). Math Drudge, April 16,

2015. URL https://experimentalmath.info/blog/2015/04/
desperately-seeking-et-fermis-paradox-turns-65-part-ii/.

Bailey:2015:DSFb

[BB15k] David H. Bailey and Jonathan M. Borwein. Desperately seeking ET: Fermi’s Paradox turns 65 (Part II). Huffington Post, ??

(??):??, April 17, 2015. URL http://www.huffingtonpost.com/
david-h-bailey/where-is-et-fermis-parado_b_7014044.html.

See Part I [?] and response [Sol15].

Bailey:2015:DGC

[BB15l] David H. Bailey and Jonathan M. Borwein. Does gun control encourage crime? The science of crime statistics. Huffington Post,

??(??):??, August 3, 2015. URL http://www.huffingtonpost.
com/david-h-bailey/does-gun-control-encourage-crime_
b_7917684.html.

Bailey:2015:DPO

[BB15m] David H. Bailey and Jonathan M. Borwein. Does public opinion always agree with scientific fact. Math Drudge, January 30,

2015. URL https://experimentalmath.info/blog/2015/01/
does-public-opinion-always-agree-with-scientific-fact/.

Bailey:2015:EAM

[BB15n] David H. Bailey and Jonathan M. Borwein. Experimental applied mathematics. In Higham et al. [HDG^{+}15], pages 925–933. ISBN 0-691-15039-7 (hardcover). LCCN QA155 .P75 2015. URL http:

//docserver.carma.newcastle.edu.au/1460/. Entry VIII.6.

Bailey:2015:ECO

[BB15o] David H. Bailey and Jonathan M. Borwein. Experimental computation as an ontological game changer: the impact of modern mathematical computation tools on the ontology of mathematics. In Mathematics, substance and surmise, pages 25–67. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK /

etc., 2015. URL http://www.davidhbailey.com/dhbpapers/
ontology.pdf.

Bailey:2015:HPA

[BB15p] David H. Bailey and Jonathan M. Borwein. High-precision arithmetic in mathematical physics. Mathematics, 3(2):337–367, June

2015. ISSN 2227-7390. URL http://www.mdpi.com/2227-7390/
3/2/337.

Bailey:2015:HCS

[BB15q] David H. Bailey and Jonathan M. Borwein. How certain are scientists that the earth is many millions of years old. Math Drudge,

June 2, 2015. URL https://experimentalmath.info/blog/
2015/06/how-certain-are-scientists-that-the-earth-is-many-millions-of-years-old/.

Bailey:2015:HMP

[BB15r] David H. Bailey and Jonathan M. Borwein. How have 2014 market prophets fared? Huffington Post, ??(??):??, De-

cember 9, 2015. URL http://www.huffingtonpost.com/
david-h-bailey/how-have-2014-market-prophets-fared_b_
6273434.html.

Bailey:2015:HWD

[BB15s] David H. Bailey and Jonathan M. Borwein. How well do individuals understand social security (and its overseas counterparts)? Huffington Post, ??(??):??, June 25, 2015. URL http://www.

huffingtonpost.com/david-h-bailey/how-well-do-individuals-u_
b_7664706.html.

Bailey:2015:PPB

[BB15t] David H. Bailey and Jonathan M. Borwein. I prefer pi: Background for Big Pi Day (3/14/15). Math Drudge, March 12,

2015. URL https://experimentalmath.info/blog/2015/03/
i-prefer-pi-background-for-big-pi-day-31415/.

Bailey:2015:IAR

[BB15u] David H. Bailey and Jonathan M. Borwein. Interview with Andrea Rossi, LENR energy pioneer. Math Drudge, October 5,

2015. URL https://experimentalmath.info/blog/2015/10/
interview-with-andrea-rossi-lenr-energy-pioneer/.

Bailey:2015:NMP

[BB15v] David H. Bailey and Jonathan M. Borwein. Is the nature of mathematical proof changing. Math Drudge, February 6,

2015. URL https://experimentalmath.info/blog/2015/02/
is-the-nature-of-mathematical-proof-changing/.

Bailey:2015:UCS

[BB15w] David H. Bailey and Jonathan M. Borwein. Is US crime soaring? Do gun controls encourage crime? The science of crime statistics. Math Drudge, July 31, 2015. URL https://experimentalmath.

info/blog/2015/07/is-us-crime-soaring-do-gun-controls-encourage-crime-the-science-of-crime-statistics/.

Bailey:2015:LFCb

[BB15x] David H. Bailey and Jonathan M. Borwein. Lessons from the “flash crash” regulatory fiasco. Huffington Post, ??(??):

??, April 27, 2015. URL http://www.huffingtonpost.com/
david-h-bailey/lessons-from-the-flash-cr_b_7148898.html.

Bailey:2015:MLYb

[BB15y] David H. Bailey and Jonathan M. Borwein. Moore’s Law is 50 years old but will it continue? The Conversation, ??

(??):??, July 20, 2015. URL https://theconversation.com/
moores-law-is-50-years-old-but-will-it-continue-44511.

Bailey:2015:MLYa

[BB15z] David H. Bailey and Jonathan M. Borwein. Moore’s Law is 50 years old: Will it continue? Math Drudge, July 14,

2015. URL https://experimentalmath.info/blog/2015/07/
moores-law-is-50-years-old-will-it-continue/.

Bailey:2015:MMP

[BB15-27] David H. Bailey and Jonathan M. Borwein. More mathematics (and Pi) in the media. Math Drudge, December 3,

2015. URL https://experimentalmath.info/blog/2015/12/
more-mathematics-and-pi-in-the-media/.

Bailey:2015:PPD

[BB15-28] David H. Bailey and Jonathan M. Borwein. Prepared for pi day? This year it’s a once in a century celebration. The Conversation, ??(??):??, March 13, 2015. URL https://theconversation.

com/prepared-for-pi-day-this-year-its-a-once-in-a-century-celebration-38576.

Bailey:2015:WFPa

[BB15-29] David H. Bailey and Jonathan M. Borwein. Where is ET? Fermi’s paradox turns 65. Math Drudge, April 6,

2015. URL https://experimentalmath.info/blog/2015/04/
where-is-et-fermis-paradox-turns-2015/.

Bailey:2015:WFPb

[BB15-30] David H. Bailey and Jonathan M. Borwein. Where is ET? Fermi’s Paradox turns 65. Huffington Post, ??(??):

??, April 10, 2015. URL http://www.huffingtonpost.com/
david-h-bailey/where-is-et-fermis-parado_b_7014044.html.

See also Part II [?].

Borwein:2015:ANK

[BB15-31] Naomi Borwein and Jonathan M. Borwein. Antisocial networking kills, again. Math Drudge, October 7, 2015. URL https://

experimentalmath.info/blog/2015/10/antisocial-networking-kills-again/.

Bailey:2016:CEE

[BB16a] D. H. Bailey and J. M. Borwein. Computation and experimental evaluation of Mordell–Tornheim–Witten sum derivatives. Technical report, University of California at Davis and Centre for Computer Assisted Research Mathematics and its Applications (CARMA), University of Newcastle, Davis, CA, USA and Callaghan, NSW 2308, Australia, August 16,

2016. 9 pp. URL http://www.davidhbailey.com/dhbpapers/
omega-numerics.pdf.

Bailey:2016:CSC

[BB16b] D. H. Bailey and J. M. Borwein. Computation and structure of character polylogarithms with applications to character Mordell–Tornheim–Witten sums. Mathematics of Computation, 85(297):295–324, 2016. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

docserver.carma.newcastle.edu.au/1492/;http://www.ams.
org/journals/mcom/2016-85-297/S0025-5718-2015-02974-3;http:
//www.ams.org/journals/mcom/2016-85-297/S0025-5718-2015-02974-3/
S0025-5718-2015-02974-3.pdf;http://www.ams.org/mathscinet/
search/author.html?authorName=Borwein%2C%20J.%20M.;http:
//www.ams.org/mathscinet/search/author.html?mrauthid=
29355;http://www.carma.newcastle.edu.au/jon/MTWIII.pdf.

Bailey:2016:EMS

[BB16c] D. H. Bailey and J. M. Borwein. Experimental mathematics in the society of the future. In 7th European Mathematics Congress (EMC) on Mathematics in the Society of the Future, pages 7–25. European Mathematical Society, Zürich, Switzerland, 2016. URL

http://docserver.carma.newcastle.edu.au/1675/.

Bailey:2016:AIS

[BB16d] David H. Bailey and Jonathan M. Borwein. Ancient Indian square roots: An exercise in forensic paleo-mathematics. In Encyclopaedia of the History of Science, Technology, and Medicine in non-Western cultures [Sel16], pages 305–310. ISBN 94-007-7746-9 (set), 94-007-7747-7 (e-book), 94-007-7748-5 (bundle).

LCCN ???? URL http://docserver.carma.newcastle.edu.

au/1385/. Redacted from [BB12r].

Bailey:2016:HCB

[BB16e] David H. Bailey and Jonathan M. Borwein. Are humans or computers better at mathematics? Blog posting, November 27, 2016. This article was co-authored with Jonathan M. Borwein before his death on 2 August 2016. A condensed version of this article appeared in [BB16s].

Bailey:2016:GWD

[BB16f] David H. Bailey and Jonathan M. Borwein. Gravitational waves detected, as predicted by Einstein’s mathematics. Math Drudge,

February 11, 2016. URL https://experimentalmath.info/
blog/2016/02/gravitational-waves-detected-as-predicted-by-einsteins-mathematics/.

Bailey:2016:HLIa

[BB16g] David H. Bailey and Jonathan M. Borwein. How likely is it that scientists are engaged in a conspiracy. Math Drudge, January 30,

2016. URL https://experimentalmath.info/blog/2016/01/
how-likely-is-it-that-scientists-are-engaged-in-a-conspiracy/.

Bailey:2016:HLIb

[BB16h] David H. Bailey and Jonathan M. Borwein. How likely is it that scientists are engaged in a conspiracy? Huffington Post, ??(??):

??, February 2, 2016. URL http://www.huffingtonpost.com/
david-h-bailey/how-likely-is-it-that-sci_b_9121342.html.

Bailey:2016:IAR

[BB16i] David H. Bailey and Jonathan M. Borwein. Interview with Andrea Rossi, LENR energy pioneer. Huffington Post, ??(??):

??, October 7, 2016. URL http://www.huffingtonpost.com/
david-h-bailey/interview-with-andrea-ros_b_8248624.html.

Bailey:2016:PDa

[BB16j] David H. Bailey and Jonathan M. Borwein. Pi Day 2016. Math

Drudge, March 10, 2016. URL https://experimentalmath.
info/blog/2016/03/pi-day-2016/.

Bailey:2016:PDb

[BB16k] David H. Bailey and Jonathan M. Borwein. Pi day 2016. Huffington Post, ??(??):??, March 14, 2016. URL http://www.

huffingtonpost.com/david-h-bailey/pi-day-2016_b_9432600.
html.

Bailey:2016:PNG

[BB16l] David H. Bailey and Jonathan M. Borwein, editors. Pi: the next generation: a sourcebook on the recent history of Pi and its computation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-319-32375-X, 3-319-32377-6 (e-book). xiv + 507 pp. LCCN

QA251. URL http://docserver.carma.newcastle.edu.au/
1716/;http://lib.myilibrary.com?id=941862.

Bailey:2016:SEFa

[BB16m] David H. Bailey and Jonathan M. Borwein. Space exploration: The future is now. Math Drudge, April 14,

2016. URL https://experimentalmath.info/blog/2016/04/
space-exploration-the-future-is-now/.

Bailey:2016:SEFb

[BB16n] David H. Bailey and Jonathan M. Borwein. Space exploration: The future is now. Huffington Post, ??(??):

??, April 15, 2016. URL http://www.huffingtonpost.com/
david-h-bailey/space-exploration-the-fut_b_9697226.html.

Bailey:2016:SPP

[BB16o] David H. Bailey and Jonathan M. Borwein. Sphere packing problem solved in 8 and 24 dimensions. Blog posting, April 1,

2016. URL https://experimentalmath.info/blog/2016/04/
sphere-packing-problem-solved-in-8-and-24-dimensions/.

See research papers [Via16, CKM^{+}16].

Bailey:2016:UPF

[BB16p] David H. Bailey and Jonathan M. Borwein. Unexpected pattern found in prime number digits. Math Drudge, March 15,

2016. URL https://experimentalmath.info/blog/2016/03/
unexpected-pattern-found-in-prime-numbers-digits/.

Bailey:2016:WMMa

[BB16q] David H. Bailey and Jonathan M. Borwein. Why are so many mathematicians also musicians? Math Drudge, April 30,

2016. URL https://experimentalmath.info/blog/2016/04/
why-are-so-many-mathematicians-also-musicians/.

Bailey:2016:WMMb

[BB16r] David H. Bailey and Jonathan M. Borwein. Why are so many mathematicians also musicians? Huffington Post, ??

(??):??, May 3, 2016. URL http://www.huffingtonpost.com/
david-h-bailey/why-are-so-many-mathemati_b_9814796.html.

Bailey:2016:WCR

[BB16s] David H. Bailey and Jonathan M. Borwein. Will computers replace humans in mathematics? The Conversation, ??(??):??, June

2, 2016. URL https://theconversation.edu.au/profiles/
jon-borwein-101.

Borwein:2016:AGM

[BB16t] J. M. Borwein and P. B. Borwein. The arithmetic–geometric mean and fast computation of elementary functions (1984). In Bailey and Borwein [BB16l], pages 79–96. ISBN 3-319-32375-X, 3-319-32377-6 (e-book). LCCN QA251. URL http://link.springer.

com/chapter/10.1007/978-3-319-32377-0_4.

Borwein:2016:RP

[BB16u] Jonathan M. Borwein and Peter B. Borwein. Ramanujan and Pi. In Bailey and Borwein [BB16l], pages 165–174. ISBN 3-319-32375-X, 3-319-32377-6 (e-book). LCCN QA251. URL

http://docserver.carma.newcastle.edu.au/1379/. Reprint

of [BB88f].

Bailey:2017:CMV

[BB17] David H. Bailey and Jonathan M. Borwein. A computational mathematics view of space, time and complexity. In Space, Time, and Frontiers of Human Understanding [WG17], pages 403–416. ISBN 3-319-44417-4, 3-319-44418-2 (e-book). LCCN BD161. URL

http://docserver.carma.newcastle.edu.au/1696/.

Bailey:2018:CEE

[BB18] D. H. Bailey and J. M. Borwein. Computation and experimental evaluation of Mordell–Tornheim–Witten sum derivatives. Experimental Mathematics, 27(3):370–376, 2018. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://www.

davidhbailey.com/dhbpapers/omega-numerics.pdf;http://
www.tandfonline.com/doi/full/10.1080/10586458.2017.1295687.

Borwein:1989:RME

[BBB89] J. M. Borwein, P. B. Borwein, and D. H. Bailey. Ramanujan, modular equations, and approximations to pi, or How to compute one billion digits of pi. American Mathematical Monthly, 96(3):201–219, March 1989. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//docserver.carma.newcastle.edu.au/1587/;https://www.
jstor.org/stable/2325206;https://web.archive.org/web/

20170328083101. Reprinted in [BBB97a].

Borwein:1996:DAM

[BBB^{+}96a] Jonathan M. Borwein, Peter B. Borwein, Stephen Braham, Robert Corless, and Loki Jörgenson. Digitally activated mathematics for a brave new World Wide Web. Education, Research and Perspectives, 23(2):28–47, December 1996. ISSN 0311-2543 (print), 1446-0017 (electronic). URL http:

//docserver.carma.newcastle.edu.au/1533/;http://nova.
newcastle.edu.au/vital/access/manager/Repository/uon:
13132;https://search.informit.com.au/documentSummary%
3Bdn=970909729%3Bres=IELAPA. Special issue on multimedia

technologies and education.

Borwein:1996:EFEa

[BBB96b] Jonathan M. Borwein, David M. Bradley, and David J. Broadhurst. Evaluations of k-fold Euler/Zagier sums: a compendium of results for arbitrary k. Technical Report CECM-96-067, OUT-4102-63, hep-th/9611004, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, September 2, 1996. 21 pp. URL

http://docserver.carma.newcastle.edu.au/161.

Borwein:1996:EFEb

[BBB96c] Jonathan M. Borwein, David M. Bradley, and David J. Broadhurst. Evaluations of k-fold Euler/Zagier sums: a compendium of results for arbitrary k. ArXiv High Energy Physics — Theory e-prints, November 1996. URL http://adsabs.

harvard.edu/abs/1996hep.th...11004B;http://arxiv.org/
abs/hep-th/9611004.

Bailey:1997:RME

[BBB97a] D. H. Bailey, J. M. Borwein, and P. B. Borwein. Ramanujan, modular equations, and approximations to pi or How to compute one billion digits of pi [MR0991866 (90d:11143)]. In Borwein et al. [BBJC97], pages 35–71. ISBN 0-8218-0668-8. ISSN 0731-1036. LCCN QA1 .O67 1995. URL

http://docserver.carma.newcastle.edu.au/1587/;http://
www.cecm.sfu.ca/organics/papers/borwein/index.html;https:
//web.archive.org/web/20170328083101/. Reprint of [BBB89].

Berggren:1997:PSB

[BBB97b] Lennart Berggren, Jonathan M. Borwein, and Peter B. Borwein, editors. Pi, a source book. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 0-387-94924-0, 1-4757-2736-4 (e-book), 1-4757-2738-0 (print), 3-540-94924-0. xix + 716 pp. LCCN QA484 .P5 1997.

Borwein:1997:RME

[BBB97c] J. M. Borwein, P. B. Borwein, and D. H. Bailey. Ramanujan, modular equations, and approximations to pi or how to compute one billion digits of pi. In Berggren et al. [BBB97b], page ?? ISBN 0-387-94924-0, 1-4757-2736-4 (e-book), 1-4757-2738-0 (print), 3-540-94924-0. LCCN QA484 .P5 1997. URL http://docserver.

carma.newcastle.edu.au/1587/;http://link.springer.com/
chapter/10.1007/978-1-4757-2736-4_64;http://www.cecm.
sfu.ca/organics/papers/borwein/index.html.

Borwein:1997:EFE

[BBB97d] Jonathan M. Borwein, David M. Bradley, and David J. Broadhurst. Evaluations of k-fold Euler/Zagier sums: a compendium of results for arbitrary k. Electronic Journal of Combinatorics, 4 (2):Research Paper 5, 1997. ISSN 1077-8926 (print), 1097-1440

(electronic). URL http://www.combinatorics.org/Volume_4/

Abstracts/v4i2r5.html. The Wilf Festschrift (Philadelphia, PA,

Berggren:2000:PSB

[BBB00a] Lennart Berggren, Jonathan Borwein, and Peter Borwein, editors. Pi: a source book. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2000. ISBN 0-387-98946-3 (hardcover). xx + 736 pp. LCCN QA484 .P5 2000.

Borwein:2000:RME

[BBB00b] J. M. Borwein, P. B. Borwein, and D. H. Bailey. Ramanujan, modular equations, and approximations to pi or how to compute one billion digits of pi. In Berggren et al. [BBB00a], pages 623–641. ISBN 0-387-98946-3 (hardcover). LCCN

QA484 .P5 2000. URL http://docserver.carma.newcastle.
edu.au/1587/;http://link.springer.com/chapter/10.1007/
978-1-4757-2736-4_64;http://link.springer.com/chapter/
10.1007/978-1-4757-3240-5_64.

Berggren:2003:PP

[BBB03] J. Lennart Berggren, Jonathan Borwein, and Peter Borwein. A pamphlet on pi serving as a supplement for the third edition of Pi: A Source Book. CECM Preprint 2003:210, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, July 17, 2003.

65 pp. URL http://docserver.carma.newcastle.edu.au/33/.

Berggren:2004:PSB

[BBB04a] Lennart Berggren, Jonathan Borwein, and Peter Borwein, editors. Pi: a source book. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2004. ISBN 0-387-20571-3. xx + 797 pp.

Borwein:2004:RME

[BBB04b] J. M. Borwein, P. B. Borwein, and D. H. Bailey. Ramanujan, modular equations, and approximations to pi or how to compute one billion digits of pi. In Berggren et al. [BBB04a], pages 623–641. ISBN 0-387-20571-3. URL http://docserver.

carma.newcastle.edu.au/1587/;http://link.springer.com/
chapter/10.1007/978-1-4757-4217-6_64.

Bailey:2005:EDA

[BBB05] David H. Bailey, Jonathan M. Borwein, and David M. Bradley. Experimental determination of Apéry-like identities for ζ(2n + 2). ArXiv Mathematics e-prints, May

2005. URL http://adsabs.harvard.edu/abs/2005math...
...5270B;http://docserver.carma.newcastle.edu.au/163/
;http://docserver.carma.newcastle.edu.au/295/.

Bailey:2006:EDA

[BBB06a] David H. Bailey, Jonathan M. Borwein, and David M. Bradley. Experimental determination of Apéry-like identities for ζ(2n + 2). Experimental Mathematics, 15(3):281–289, 2006. CODEN ???? ISSN 1058-6458 (print), 1944-950X (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
295/;http://projecteuclid.org/euclid.em/1175789759.

Borwein:2006:PES

[BBB06b] David Borwein, Jonathan M. Borwein, and David M. Bradley. Parametric Euler sum identities. Journal of Mathematical Analysis and Applications, 316(1):328–338, April 2006. CO-DEN JMANAK. ISSN 0022-247X (print), 1096-0813 (elec-

tronic). URL http://adsabs.harvard.edu/abs/2006JMAA.
.316..328B;http://docserver.carma.newcastle.edu.au/1236/
;http://www.sciencedirect.com/science/article/pii/S0022247X05003811.

Bartz:2007:FFC

[BBB^{+}07] Sedi Bartz, Heinz H. Bauschke, Jonathan M. Borwein, Simeon Reich, and Xianfu Wang. Fitzpatrick functions, cyclic monotonicity and Rockafellar’s antiderivative. Nonlinear Analysis, Theory, Methods and Applications, 66(5):1198–1223, 2007. CODEN NOANDD. ISSN 0362-546x (print), 1873-5215

(electronic). URL http://docserver.carma.newcastle.edu.
au/1232/;http://www.sciencedirect.com/science/article/
pii/S0362546X06000435.

Baillie:2008:SSS

[BBB08] Robert Baillie, David Borwein, and Jonathan M. Borwein. Surprising sinc sums and integrals. American Mathematical Monthly, 115(10):888–901, December 2008. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://docserver.carma.newcastle.edu.au/332/;https://
www.jstor.org/stable/27642636.

Bailey:2012:HPC

[BBB12] David H. Bailey, Roberto Barrio, and Jonathan M. Borwein. Highprecision computation: Mathematical physics and dynamics. Applied Mathematics and Computation, 218(20):10106–10121, June 15, 2012. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://docserver.carma.newcastle.edu.
au/775/;http://www.sciencedirect.com/science/article/
pii/S0096300312003505.

Bailey:2013:CAM

[BBB^{+}13] David H. Bailey, Heinz H. Bauschke, Peter Borwein, Frank Garvan, Michel Théra, Jon D. Vanderwerff, and Henry Wolkowicz, editors. Computational and analytical mathematics: in honor of Jonathan Borwein’s 60th Birthday, volume 50 of Springer proceedings in mathematics and statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-7620-8, 1-4614-7621-6 (e-book). ISSN 2194-1009. xv + 701 pp. LCCN QA241. URL http://public.

eblib.com/choice/publicfullrecord.aspx?p=1466708;http:
//swb.eblib.com/patron/FullRecord.aspx?p=1466708;http:
//www.myilibrary.com?id=547562.

Bailey:2015:ELG

[BBB15] David H. Bailey, David Borwein, and Jonathan M. Borwein. On Eulerian log-gamma integrals and Tornheim–Witten zeta functions. The Ramanujan Journal, 36(1–2):43–68, February 2015. ISSN 1382-4090 (print), 1572-9303 (electronic). URL http:

//docserver.carma.newcastle.edu.au/1377/.

Borwein:2016:RME

[BBB16] J. M. Borwein, P. B. Borwein, and D. H. Bailey. Ramanujan, modular equations, and approximations to pi or how to compute one billion digits of pi. In Bailey and Borwein [BB16l], pages 175– 195. ISBN 3-319-32375-X, 3-319-32377-6 (e-book). LCCN QA251.

URL http://docserver.carma.newcastle.edu.au/1587/.

Bailey:2020:AVC

[BBB^{+}20] David H. Bailey, Naomi Simone Borwein, Richard P. Brent, Regina S. Burachik, Judy anne Heather Osborn, Brailey Sims, and Qiji J. Zhu, editors. From Analysis to Visualization: A Celebration of the Life and Legacy of Jonathan M. Borwein, Callaghan, Australia, September 2017, volume 313 of Springer Proceedings in Mathematics & Statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2020. ISBN 3-030-36567-0 (print), 3-030-36568-9 (e-book). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Bailey:2007:HFI

[BBBC07] D. H. Bailey, D. Borwein, J. M. Borwein, and R. E. Crandall. Hypergeometric forms for Ising-class integrals. Experimental Mathematics, 16(3):257–276, 2007. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://docserver.carma.

newcastle.edu.au/326/;http://projecteuclid.org/euclid.
em/1204928528.

Borwein:1994:GCP

[BBBG94] David Borwein, Jonathan M. Borwein, Peter B. Borwein, and Roland Girgensohn. Giuga’s conjecture on primality. Centre de Récherche Mathématiques XXV Anniversary Conference, Montreal, Quebec., October 7, 1994. URL http://docserver.carma.

newcastle.edu.au/88/.

Borwein:1995:GCP

[BBBG95] D. Borwein, J. M. Borwein, P. B. Borwein, and R. Girgensohn. Giuga’s conjecture on primality. Report, Department of Mathematics & Statistics, Simon Fraser University, Burnaby, BC V5A 156, Canada, June 9, 1995. 16 pp. URL http://docserver.

carma.newcastle.edu.au/88. Published in [BBBG96].

Borwein:1996:GCP

[BBBG96] D. Borwein, J. M. Borwein, P. B. Borwein, and R. Girgensohn. Giuga’s conjecture on primality. American Mathematical Monthly, 103(1):40–50, 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://docserver.

carma.newcastle.edu.au/88/.

Bailey:2008:EIE

[BBBG08] David H. Bailey, Jonathan M. Borwein, David Broadhurst, and M. L. Glasser. Elliptic integral evaluations of Bessel moments and applications. Journal of Physics A: Mathematical and Theoretical, 41(20):205203, May 2008. CODEN JPAMB5. ISSN 1751-8113 (print), 1751-8121 (electronic). URL http://adsabs.harvard.

edu/abs/2008JPhA...41t5203B;http://arxiv.org/abs/0801.
0891;http://iopscience.iop.org/article/10.1088/1751-8113/
41/20/205203;http://stacks.iop.org/1751-8121/41/i=20/
a=205203.

Borwein:1997:CAM

[BBBL97] Jonathan M. Borwein, David M. Bradley, David J. Broadhurst, and Petr Lisoněk. Combinatorial aspects of multiple zeta values. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, July 31, 1997. 12 pp. URL http://docserver.

carma.newcastle.edu.au/201/.

Borwein:1998:CAMa

[BBBL98a] Jonathan M. Borwein, David M. Bradley, David J. Broadhurst, and Petr Lisoněk. Combinatorial aspects of multiple zeta values. ArXiv Mathematics e-prints, December

1998. URL http://adsabs.harvard.edu/abs/1998math.....
12020B;http://docserver.carma.newcastle.edu.au/201/.

Borwein:1998:CAMb

[BBBL98b] Jonathan M. Borwein, David M. Bradley, David J. Broadhurst, and Petr Lisoněk. Combinatorial aspects of multiple zeta values. Electronic Journal of Combinatorics, 5:Research Paper 38, 1998. ISSN 1077-8926 (print), 1097-1440 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
201/;http://www.combinatorics.org/Volume_5/Abstracts/
v5i1r38.html.

Borwein:1998:SVM

[BBBL98c] Jonathan M. Borwein, David M. Bradley, David J. Broadhurst, and Petr Lisoněk. Special values of multiple polylogarithms. CECM Research Report 98-106, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, May 14, 1998. 45 pp.

URL http://docserver.carma.newcastle.edu.au/200.

Borwein:1999:SVM

[BBBL99] Jonathan M. Borwein, David M. Bradley, David J. Broadhurst, and Petr Lisoněk. Special values of multiple polylogarithms. ArXiv Mathematics e-prints, October 1999. URL http://

adsabs.harvard.edu/abs/1999math.....10045B;http://docserver.
carma.newcastle.edu.au/200/.

Borwein:2001:SVM

[BBBL01] Jonathan M. Borwein, David M. Bradley, David J. Broadhurst, and Petr Lisoněk. Special values of multiple polylogarithms. Transactions of the American Mathematical Society, 353(3):907– 941, March 2001. CODEN TAMTAM. ISSN 0002-9947 (print),

1088-6850 (electronic). URL https://www.jstor.org/stable/
221840.

Bailey:1996:QP

[BBBP96] David H. Bailey, Jonathan M. Borwein, Peter B. Borwein, and Simon Plouffe. The quest for pi. Technical report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, June

25, 1996. 16 pp. URL http://docserver.carma.newcastle.

edu.au/164/. Published in [BBBP97].

Bailey:1997:QP

[BBBP97] David H. Bailey, Jonathan M. Borwein, Peter B. Borwein, and Simon Plouffe. The quest for pi. The Mathematical Intelligencer, 19(1):50–57, 1997. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic). URL http://docserver.carma.

newcastle.edu.au/164/.

Bailey:2016:RCS

[BBBR16] David H. Bailey, Jonathan M. Borwein, Richard P. Brent, and Mohsen Reisi. Reproducibility in computational science: A case study: Randomness of the digits of pi. Experimental Mathematics, 22(??):1–8, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://docserver.carma.

newcastle.edu.au/1695/;http://www.tandfonline.com/doi/
full/10.1080/10586458.2016.1163755. See [Gan14].

Bailey:2017:RCS

[BBBR17] David H. Bailey, Jonathan M. Borwein, Richard P. Brent, and Mohsen Reisi. Reproducibility in computational science: A case study: Randomness of the digits of pi. Experimental Mathematics, 26(3):298–305, 2017. CODEN ???? ISSN 1058-6458 (print),

1944-950X (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/10586458.2016.1163755. See reply [Gan17].

Bailey:2010:EMMb

[BBBZ10a] David H. Bailey, Jonathan M. Borwein, David Broadhurst, and Wadim Zudilin. Experimental mathematics and mathematical physics. In Gems in experimental mathematics, volume 517 of Contemp. Math., pages 41–58. American Mathematical Society, Providence, RI, USA, 2010. URL http://docserver.carma.

newcastle.edu.au/1215/.

Bailey:2010:EMMa

[BBBZ10b] David H. Bailey, Jonathan M. Borwein, David J. Broadhurst, and Wadim Zudilin. Experimental mathematics and mathematical physics. ArXiv e-prints, May 2010. URL http://adsabs.

harvard.edu/abs/2010arXiv1005.0414B;http://docserver.
carma.newcastle.edu.au/1215/.

Borwein:1996:WOM

[BBC^{+}96] Jonathan M. Borwein, Peter B. Borwein, Robert M. Corless, Loki Jörgenson, and Nathalie Sinclair. What is organic mathematics? Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, April 24, 1996. 20 pp. URL http://docserver.

carma.newcastle.edu.au/157/. Published in [BBC^{+}97b].

Bailey:1997:KC

[BBC97a] David H. Bailey, Jonathan M. Borwein, and Richard E. Crandall. On the Khintchine constant. Mathematics of Computation, 66 (217):417–431, 1997. CODEN MCMPAF. ISSN 0025-5718 (print),

1088-6842 (electronic). URL http://adsabs.harvard.edu/abs/
1997MaCom..66..417B;http://docserver.carma.newcastle.
edu.au/102/.

Borwein:1997:WOM

[BBC^{+}97b] Jonathan M. Borwein, Peter B. Borwein, Robert M. Corless, Loki Jörgenson, and Nathalie Sinclair. What is organic mathematics? In Borwein et al. [BBJC97], pages 1– 18. ISBN 0-8218-0668-8. ISSN 0731-1036. LCCN QA1 .O67

1995. URL http://docserver.carma.newcastle.edu.au/157/
;http://www.cecm.sfu.ca/organics.

Borwein:1998:CSR

[BBC98] Jonathan M. Borwein, David M. Bradley, and Richard E. Crandall. Computational strategies for the Riemann zeta function. Report CECM-98-118, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, October 30, 1998. 68 pp.

URL http://docserver.carma.newcastle.edu.au/211;http:
//people.reed.edu/~crandall/papers/attach01.pdf. Pub-

lished in [BBC00b].

Bailey:199x:KC

[BBC9x] David H. Bailey, Jonathan M. Borwein, and Richard E. Crandall. On the Khintchine constant. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, 199x. 19 pp. Published in [BBC97a].

Bauschke:2000:ESE

[BBC00a] Heinz H. Bauschke, Jonathan M. Borwein, and Patrick L. Combettes. Essential smoothness, essential strict convexity, and Legendre functions in Banach spaces. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, 2000. 32 pp. Published in [BBC01].

Borwein:2000:CSR

[BBC00b] Jonathan M. Borwein, David M. Bradley, and Richard E. Crandall. Computational strategies for the Riemann zeta function.

Journal of Computational and Applied Mathematics, 121(1–2): 247–296, September 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://people.reed.edu/

~crandall/papers/attach01.pdf. Numerical analysis in the

20th century, Vol. I, Approximation theory.

Bauschke:2001:ESE

[BBC01] Heinz H. Bauschke, Jonathan M. Borwein, and Patrick L. Combettes. Essential smoothness, essential strict convexity, and Legendre functions in Banach spaces. Communications in Contemporary Mathematics, 3(4):615–647, 2001. ISSN 0219-1997 (print), 1793-6683 (electronic). URL http://docserver.carma.

newcastle.edu.au/232/.

Bauschke:2003:BMO

[BBC03] Heinz H. Bauschke, Jonathan M. Borwein, and Patrick L. Combettes. Bregman monotone optimization algorithms. SIAM Journal on Control and Optimization, 42(2):596–636, 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL

http://docserver.carma.newcastle.edu.au/127/.

Bailey:2006:IIC

[BBC06] D. H. Bailey, J. M. Borwein, and R. E. Crandall. Integrals of the Ising class. Journal of Physics A (Mathematical and General), 39(40):12271–12302, October 2006. CO-DEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (elec-

tronic). URL http://adsabs.harvard.edu/abs/2006JPhA...
39..001B;http://docserver.carma.newcastle.edu.au/324/
;http://iopscience.iop.org/article/10.1088/0305-4470/
39/40/001;http://stacks.iop.org/0305-4470/39/i=40/a=001.

Bailey:2007:BI

[BBC07a] D. H. Bailey, J. M. Borwein, and R. E. Crandall. Box integrals. Journal of Computational and Applied Mathematics, 206(1):196–208, September 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://adsabs.

harvard.edu/abs/2007JCoAM.206..196B;http://docserver.
carma.newcastle.edu.au/320/;http://www.sciencedirect.
com/science/article/pii/S0377042706004250.

Bailey:2007:EMA

[BBC^{+}07b] David H. Bailey, Jonathan M. Borwein, Neil J. Calkin, Roland Girgensohn, D. Russell Luke, and Victor Moll. Experimental Mathematics in Action. A. K. Peters, Ltd., Wellesley, MA, USA, 2007. ISBN 1-56881-271-X. xii + 322 pp. LCCN QA8.7 .E97 2007.

URL http://docserver.carma.newcastle.edu.au/1733/.

Borwein:2007:ELAa

[BBC07c] David Borwein, Jonathan Borwein, and Richard Crandall. Effective Laguerre asymptotics. Report, Department of Computer Science, Dalhousie University, Halifax, NS, Canada B3H 4R2,

2007. URL http://docserver.carma.newcastle.edu.au/334/
;http://locutus.cs.dal.ca:8088/archive/00000334/;http:
//web.archive.org/web/20070208032521/h.

Borwein:2008:EBF

[BBC08a] David Borwein, Jonathan M. Borwein, and O-Yeat Chan. The evaluation of Bessel functions via exp–arc integrals. Journal of Mathematical Analysis and Applications, 341(1): 478–500, May 2008. CODEN JMANAK. ISSN 0022-247X (print), 1096-0813 (electronic). URL http://adsabs.

harvard.edu/abs/2008JMAA..341..478B;http://docserver.
carma.newcastle.edu.au/1231/;http://www.sciencedirect.
com/science/article/pii/S0022247X07012346.

Borwein:2008:ELA

[BBC08b] David Borwein, Jonathan M. Borwein, and Richard E. Crandall. Effective Laguerre asymptotics. SIAM Journal on Numerical Analysis, 46(6):3285–3312, 2008. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic). URL http://docserver.

carma.newcastle.edu.au/334/.

Bailey:2009:RQR

[BBC09] D. H. Bailey, J. M. Borwein, and R. E. Crandall. Resolution of the Quinn–Rand–Strogatz constant of nonlinear physics. Experimental Mathematics, 18(1):107–116, 2009. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://docserver.

carma.newcastle.edu.au/345/;http://projecteuclid.org/
euclid.em/1243430534.

Bailey:2010:ATB

[BBC10] D. H. Bailey, J. M. Borwein, and R. E. Crandall. Advances in the theory of box integrals. Mathematics of Computation, 79(271):1839–1866, July 2010. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

http://docserver.carma.newcastle.edu.au/389/;https://
www.jstor.org/stable/20779125.

Bailey:2011:EAN

[BBC^{+}11a] David H. Bailey, Jonathan M. Borwein, Cristian S. Calude, Michael J. Dinneen, Monica Dumitrescu, and Alex Yee. An empirical approach to the normality of π. Report CDMTCS-413, Centre for Discrete Mathematics and Theoretical Computer Science, University of Auckland, Auckland, New Zealand, November 2011.

i + 31 pp. URL http://docserver.carma.newcastle.edu.
au/1382/;http://www.cs.auckland.ac.nz/research/groups/
CDMTCS/researchreports/?download&paper_file=413.

Bauschke:2011:FPA

[BBC^{+}11b] Heinz H. Bauschke, Regina S. Burachik, Patrick Louis Combettes, Veit Elser, D. Russell Luke, and Henry Wolkowicz, editors. Fixedpoint algorithms for inverse problems in science and engineering, volume 49 of Springer optimization and its applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 1-4419-9568-4, 1-4419-9569-2 (e-book). ISSN 1931-6828 (print), 1931-6836 (electronic). xi + 402 pp. LCCN

QA378.5 .F59 2011. URL http://0-dx.doi.org.fama.us.es/
10.1007/978-1-4419-9569-8.

Bailey:2012:NP

[BBC^{+}12a] D. H. Bailey, J. M. Borwein, C. S. Calude, M. J. Dinneen, M. Dumitrescu, and A. Yee. Normality and pi. Preprint.,

February 2012. URL http://www.carma.newcastle.edu.au/
jon/normality-long.pdf.

Bailey:2012:EAN

[BBC^{+}12b] David H. Bailey, Jonathan M. Borwein, Cristian S. Calude, Michael J. Dinneen, Monica Dumitrescu, and Alex Yee. An empirical approach to the normality of π. Experimental Mathematics, 21(4):375–384, 2012. CODEN ???? ISSN 1058-6458 (print), 1944-

950X (electronic). URL http://docserver.carma.newcastle.
edu.au/1382/.

Bailey:2012:ND

[BBC^{+}12c] David H. Bailey, Jonathan M. Borwein, Cristian S. Calude, Michael J. Dinneen, Monica Dumitrescu, and Alex Yee. Normality and the digits of π. In David H. Bailey and Jonathan M. Borwein, editors, Exploratory Experimentation in Mathematics: Selected Works. Perfectly Scientific Press, Portland, OR, USA, Febru-

ary 3, 2012. URL http://www.davidhbailey.com/dhbpapers/
normality-digits-pi.pdf.

Bailey:2014:OCC

[BBC^{+}14a] David H. Bailey, Jonathan M. Borwein, Olga Caprotti, Ursula Martin, Bruno Salvy, and Michela Taufer. Opportunities and challenges in 21st century mathematical computation: ICERM workshop report, July 10–14, 2014. Technical report, ????, ????,

August 28, 2014. 18 pp. URL http://www.davidhbailey.com/

dhbpapers/ICERM-2014.pdf. Preprint in collaboration with the workshop participants.

Bailey:2014:CTE

[BBC14b] David H. Bailey, Jonathan M. Borwein, and Richard E. Crandall. Computation and theory of extended Mordell– Tornheim–Witten sums. Mathematics of Computation, 83 (288):1795–1821, July 2014. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL http://

docserver.carma.newcastle.edu.au/1459/;http://www.ams.
org/journals/mcom/2014-83-288/S0025-5718-2014-02768-3;http:
//www.ams.org/journals/mcom/2014-83-288/S0025-5718-2014-02768-3/
S0025-5718-2014-02768-3.pdf.

Borwein:2014:LKMa

[BBC14c] Jonathan M. Borwein, Yann Bugeaud, and Michael Coons. The legacy of Kurt Mahler. European Mathematical Society. Newsletter, 91:19–23, March 2014. ISSN 1027-488X. URL http://www.

ems-ph.org/journals/journal.php?jrn=news.

Borwein:2014:LKMb

[BBC14d] Jonathan M. Borwein, Yann Bugeaud, and Michael Coons. The legacy of Kurt Mahler. Australian Mathematical Society Gazette, 41(1):11–21, March 2014. ISSN 0311-0729 (print), 1326-2297

(electronic). URL http://www.austms.org.au/Publ/Gazette/
2014/Mar14/Mahler.pdf.

Borwein:2015:LKM

[BBC15] Jonathan M. Borwein, Yann Bugeaud, and Michael Coons. The legacy of Kurt Mahler. Notices of the American Mathematical Society, 62(5):526–531, May 2015. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://www.ams.

org/notices/201505/rnoti-p526.pdf.

Bailey:2007:HSI

[BBCM07a] D. Bailey, J. Borwein, R. Crandall, and D. Manna. Heisenberg spin integrals. Preprint., August 2007.

Borwein:2007:DCR

[BBCM07b] D. Borwein, J. Borwein, R. Crandall, and R. Mayer. On the dynamics of certain recurrence relations. The Ramanujan Journal, 13(1–3):63–101, 2007. CODEN RAJOF9. ISSN 1382-4090 (print), 1572-9303 (electronic). URL http://docserver.

carma.newcastle.edu.au/253/;http://link.springer.com/
article/10.1007/s11139-006-0243-3.

Bailey:2004:BEA

[BBCP04] David H. Bailey, Jonathan M. Borwein, Richard E. Crandall, and Carl Pomerance. On the binary expansions of algebraic numbers. Journal de Théorie des Nombres de Bordeaux, 16(3): 487–518, 2004. ISSN 1246-7405 (print), 2118-8572 (electronic).

URL http://docserver.carma.newcastle.edu.au/36/;http:
//jtnb.cedram.org/item?id=JTNB_2004__16_3_487_0.

Bailey:2013:EFS

[BBCR13] David H. Bailey, Jonathan M. Borwein, Richard E. Crandall, and Michael G. Rose. Expectations on fractal sets. Applied Mathematics and Computation, 220(??):695–721, September 1, 2013. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649

(electronic). URL http://docserver.carma.newcastle.edu.
au/1454/;http://www.sciencedirect.com/science/article/
pii/S0096300313007273.

Bailey:2013:LSA

[BBCZ13] David H. Bailey, Jonathan M. Borwein, Richard E. Crandall, and I. John Zucker. Lattice sums arising from the Poisson equation. Journal of Physics A: Mathematical and Theoretical, 46(11): 115201, March 2013. CODEN JPAMB5. ISSN 1751-8113 (print),

1751-8121 (electronic). URL http://adsabs.harvard.edu/abs/
2013JPhA...46k5201B;http://docserver.carma.newcastle.
edu.au/1458/;http://iopscience.iop.org/article/10.1088/
1751-8113/46/11/115201;http://stacks.iop.org/1751-8121/
46/i=11/a=115201;http://www.davidhbailey.com/dhbpapers/
PoissonLattice.pdf.

Borwein:1989:PEN

[BBD89] J. M. Borwein, P. B. Borwein, and K. Dilcher. Pi, Euler numbers, and asymptotic expansions. American Mathematical Monthly, 96(8):681–687, 1989. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://docserver.

carma.newcastle.edu.au/1583/.

Borwein:1997:PEN

[BBD97] J. M. Borwein, P. B. Borwein, and K. Dilcher. Pi, Euler numbers, and asymptotic expansions. In Berggren et al. [BBB97b], pages 642–648. ISBN 0-387-94924-0, 1-4757-2736-4 (e-book), 1-4757-2738-0 (print), 3-540-94924-0. LCCN

QA484 .P5 1997. URL http://docserver.carma.newcastle.
edu.au/1583/;http://link.springer.com/chapter/10.1007/
978-1-4757-2736-4_65.

Borwein:2000:PEN

[BBD00] J. M. Borwein, P. B. Borwein, and K. Dilcher. Pi, Euler numbers, and asymptotic expansions. In Berggren et al. [BBB00a], pages 642–648. ISBN 0-387-98946-3 (hardcover). LCCN

QA484 .P5 2000. URL http://docserver.carma.newcastle.
edu.au/1583/;http://link.springer.com/chapter/10.1007/
978-1-4757-3240-5_65.

Borwein:2004:PEN

[BBD04] J. M. Borwein, P. B. Borwein, and K. Dilcher. Pi, Euler numbers, and asymptotic expansions. In Berggren et al. [BBB04a], pages 642–648. ISBN 0-387-20571-3. URL http://docserver.

carma.newcastle.edu.au/1583/;http://link.springer.com/
chapter/10.1007/978-1-4757-4217-6_65. Reprint of [BBD89].

Borwein:2016:PEN

[BBD16] J. M. Borwein, P. B. Borwein, and K. Dilcher. Pi, Euler numbers, and asymptotic expansions. In Bailey and Borwein [BB16l], pages 197–205. ISBN 3-319-32375-X, 3-319-32377-6 (e-book).

LCCN QA251. URL http://docserver.carma.newcastle.edu.
au/1583/.

Bacak:2010:ICL

[BBEM10] Miroslav Bačák, Jonathan M. Borwein, Andrew Eberhard, and Boris S. Mordukhovich. Infimal convolutions and Lipschitzian properties of subdifferentials for prox-regular functions in Hilbert spaces. Journal of Convex Analysis, 17(3–4):737–763, 2010. ISSN 0944-6532 (print), 2363-6394 (electronic). URL

http://docserver.carma.newcastle.edu.au/1224/;http://
www.heldermann.de/JCA/JCA17/JCA173/jca17049.htm. Special

volume in honour of Hedi Attouch.

Borwein:2000:RCS

[BBFG00] David Borwein, Jonathan Borwein, Greg Fee, and Roland Girgensohn. Refined convexity and special cases of the Blaschke–Santalo inequality. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, April 3, 2000. 8 pp. URL http:

//docserver.carma.newcastle.edu.au/222.

Borwein:2001:RCS

[BBFG01] D. Borwein, J. Borwein, G. Fee, and R. Girgensohn. Refined convexity and special cases of the Blaschke–Santalo inequality. Mathematical Inequalities & Applications, 4(4):631– 638, 2001. ISSN 1331-4343 (print), 1848-9966 (electronic). URL

http://docserver.carma.newcastle.edu.au/222/.

Bailey:1993:EEE

[BBG93a] David H. Bailey, Jonathan M. Borwein, and Roland Girgensohn. Experimental evaluation of Euler sums. Technical Report RNR-93-014, Department of Mathematics & Statistics, Simon Fraser University, Burnaby, BC V5A 156, Canada, October 20, 1993. 24

pp. URL http://docserver.carma.newcastle.edu.au/60.

Borwein:1993:HAA

[BBG93b] J. Borwein, P. Borwein, and F. Garvan. Hypergeometric analogues of the arithmetic–geometric mean iteration. Constructive Approximation, 9(4):509–523, 1993. ISSN 0176-4276 (print), 1432-

0940 (electronic). URL http://docserver.carma.newcastle.
edu.au/1556/;http://link.springer.com/article/10.1007/
BF01204654.

Bailey:1994:EEEb

[BBG94a] David H. Bailey, Jonathan M. Borwein, and Roland Girgensohn. Experimental evaluation of Euler sums. Experimental Mathematics, 3(1):17–30, 1994. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://docserver.carma.

newcastle.edu.au/60/;http://projecteuclid.org/euclid.
em/1062621000.

Borwein:1994:EEEa

[BBG94b] David Borwein, Jonathan M. Borwein, and Roland Girgensohn. Explicit evaluation of Euler sums. Report, Department of Mathematics and Statistics, Simon Fraser University, Burnaby, BC V5A 1S6, Canada, March 18, 1994. 21 pp. URL http://docserver.

carma.newcastle.edu.au/58.

Borwein:1994:SCM

[BBG94c] J. M. Borwein, P. B. Borwein, and F. G. Garvan. Some cubic modular identities of Ramanujan. Transactions of the American Mathematical Society, 343(1):35–47, May 1994. CO-DEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
1547/;https://www.jstor.org/stable/2154520.

Bailey:1995:FNI

[BBG95a] D. H. Bailey, J. M. Borwein, and R. Girgensohn. Finding new identities with supercomputers. In ????, editor, SIAM Conference on Parallel Computing, San Francisco, February 1995, page ?? Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1995.

Berndt:1995:RTE

[BBG95b] Bruce C. Berndt, S. Bhargharva, and Frank G. Garvan. Ramanujan’s theories of elliptic functions to alternative bases. Transactions of the American Mathematical Society, 347(11):4163– 4124, November 1995. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic). URL http://www.ams.org/

journals/tran/1995-347-11/S0002-9947-1995-1311903-0/
S0002-9947-1995-1311903-0.pdf.

Borwein:1995:EEEa

[BBG95c] David Borwein, Jonathan M. Borwein, and Roland Girgensohn. Explicit evaluation of Euler sums. Proceedings of the Edinburgh Mathematical Society (2), 38(2):277–294, June 1995. CODEN PRMSA3. ISSN 0013-0915 (print), 1464-3839 (electronic). URL

http://docserver.carma.newcastle.edu.au/58/.

Borwein:2003:MEP

[BBG03] Jonathan M. Borwein, David H. Bailey, and Roland Girgensohn. Mathematics by experiment: Plausible reasoning in the 21st century and experiments in mathematics: Computational paths to discovery. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, August 28, 2003.

iv + 67 pp. URL https://citeseerx.ist.psu.edu/viewdoc/
download?doi=10.1.1.187.5170&rep=rep1&type=pdf.

Borwein:2004:EMCa

[BBG04a] Jonathan M. Borwein, David H. Bailey, and Roland Girgensohn. Experimentation in mathematics: computational paths to discovery. A. K. Peters, Ltd., Wellesley, MA, USA, 2004. ISBN 1-56881-136-5. x + 357 pp. LCCN QA12 .B67 2004.

Borwein:2004:FEA

[BBG04b] Jonathan M. Borwein, David Borwein, and William F. Galway. Finding and excluding b-ary Machin-type individual digit formulae. Canadian Journal of Mathematics = Journal canadien de mathématiques, 56(5):897–925, ???? 2004. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL

http://docserver.carma.newcastle.edu.au/47/. This paper

established the result that there are no degree-1 BBP-type formulas for π, except when the base is 2 (or an integer power thereof).

Borwein:2006:EM

[BBG06] Jonathan M. Borwein, David H. Bailey, and Roland Girgensohn. Experiments in Mathematics. A. K. Peters, Ltd., Wellesley, MA, USA, 2006. ISBN 1-56881-283-3, 1-56881-211-6, 1-56881-136-5. 14 + vi + 294 + viii + 363 pp. LCCN QA76.95 .E9 2006. Combined interactive CD version.

Borwein:1995:EMD

[BBGP95a] J. Borwein, P. Borwein, R. Girgensohn, and S. Parnes. Experimental mathematics: a discussion. Report, Department of Mathematics & Statistics, Simon Fraser University, Burnaby, BC V5A 156, Canada, June 14, 1995. URL http://docserver.carma.

newcastle.edu.au/98.

Borwein:1995:MEMa

[BBGP95b] J. M. Borwein, P. B. Borwein, R. Girgensohn, and S. Parnes. Mathematical experimentation and methodology. Unpublished report., January 1995.

Borwein:1995:MSE

[BBGP95c] Jonathan M. Borwein, Peter B. Borwein, Roland Girgensohn, and S. Parnes. Making sense of experimental mathematics. CECM preprint 95:032, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, 1995. URL http://docserver.carma.

newcastle.edu.au/98/.

Borwein:1996:MSE

[BBGP96] Jonathan Borwein, Peter B. Borwein, Roland Girgensohn, and S. Parnes. Making sense of experimental mathematics. The Mathematical Intelligencer, 18(4):12–18, Fall 1996. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic). URL

http://docserver.carma.newcastle.edu.au/98/.

Borwein:19xx:EMI

[BBGPxx] Jonathan M. Borwein, Peter B. Borwein, Roland Girgensohn, and S. Parnes. Experimental mathematical investigation of decimal and continued fraction expansions of select constants. Unpublished., 19xx.

Borwein:2011:MRGa

[BBGW11] David Borwein, Jonathan M. Borwein, M. Larry Glasser, and James G. Wan. Moments of Ramanujan’s generalized elliptic integrals and extensions of Catalan’s constant. Journal of Mathematical Analysis and Applications, 384(2):478–496, 2011. CODEN JMANAK. ISSN 0022-247X (print), 1096-0813

(electronic). URL http://docserver.carma.newcastle.edu.
au/1443/;http://www.sciencedirect.com/science/article/
pii/S0022247X1100549X.

Borwein:2012:RCS

[BBJ12] Jonathan Borwein, Peter B. Borwein, and Veselin Jungić. Remote collaboration: Six years of the coast-to-coast seminar series. Science Communication, 34(3):419–428, June 2012. ISSN 1075-5470 (print), 1552-8545 (electronic). URL http://docserver.carma.

newcastle.edu.au/1387/;http://scx.sagepub.com/content/
34/3/419.abstract.

Borwein:1997:OMP

[BBJC97] Jonathan M. Borwein, Peter B. Borwein, Loki Jörgenson, and Rob Corless, editors. Organic mathematics: proceedings of the Organic Mathematics Workshop, December 12–14, 1995, Simon Fraser University, Burnaby, British Columbia, volume 20 of CMS conference proceedings. American Mathematical Society, Providence, RI, USA, 1997. ISBN 0-8218-0668-8. ISSN 0731-1036. LCCN QA1 .O67 1995.

Borwein:2000:CBSa

[BBK00a] Jonathan M. Borwein, David J. Broadhurst, and Joel Kamnitzer. Central binomial sums, multiple Clausen values and zeta values. ArXiv High Energy Physics — Theory e-prints, April

2000. URL http://adsabs.harvard.edu/abs/2000hep.th...
.4153B;http://docserver.carma.newcastle.edu.au/243/.

Borwein:2000:CBSb

[BBK00b] Jonathan Michael Borwein, David J. Broadhurst, and Joel Kamnitzer. Central binomial sums, multiple Clausen values, and zeta values. Report, Department of Mathematics, Simon Fraser University, Burnaby, BC V5A 1S6, Canada, November 4, 2000. 17 pp. Published in [BBK01].

Borwein:2001:CBS

[BBK01] Jonathan Michael Borwein, David J. Broadhurst, and Joel Kamnitzer. Central binomial sums, multiple Clausen values, and zeta values. Experimental Mathematics, 10(1):25–34, 2001. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://arxiv.org/abs/hep-th/0004153;http://docserver.
carma.newcastle.edu.au/243/;http://projecteuclid.org/
euclid.em/999188418.

Bailey:2014:ASL

[BBK14] David H. Bailey, Jonathan M. Borwein, and Alexander D. Kaiser. Automated simplification of large symbolic expressions. Journal of Symbolic Computation, 60(??):120–136, 2014. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855x

(electronic). URL http://docserver.carma.newcastle.edu.
au/1456/;http://www.sciencedirect.com/science/article/
pii/S074771711300117X.

Bailey:2016:CDA

[BBKL16] David H. Bailey, Jonathan M. Borwein, Jason Kimberley, and Watson Ladd. Computer discovery and analysis of large Poisson polynomials. Experimental Mathematics, ??(??):??, 2016. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://docserver.carma.newcastle.edu.
au/1689/;http://www.tandfonline.com/doi/full/10.1080/
10586458.2016.1180565.

Bailey:2017:CDA

[BBKL17] David H. Bailey, Jonathan M. Borwein, Jason S. Kimberley, and Watson Ladd. Computer discovery and analysis of large Poisson polynomials. Experimental Mathematics, 26(3):349–363, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2016.1180565.

Bailey:2006:TPE

[BBKW06] David H. Bailey, Jonathan M. Borwein, Vishaal Kapoor, and Eric W. Weisstein. Ten problems in experimental mathematics. American Mathematical Monthly, 113(6):481–509, June/July 2006. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL http://docserver.carma.newcastle.edu.
au/1237/;https://www.jstor.org/stable/27641975.

Bauschke:1994:MCP

[BBL94] H. H. Bauschke, J. M. Borwein, and A. S. Lewis. On the method of cyclic projections for convex sets in Hilbert space. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, February 10, 1994. 51 pp. URL http:

//docserver.carma.newcastle.edu.au/77.

Bauschke:1997:CSC

[BBL97a] H. H. Bauschke, J. M. Borwein, and A. S. Lewis. Convex sets and the cyclic projection algorithm. In Yair Censor and Simeon Reich, editors, Recent Developments in Optimization Theory and Nonlinear Analysis, volume 204 of Contemporary Mathematics, pages 1– 38. American Mathematical Society, Providence, RI, USA, 1997.

Bauschke:1997:MCP

[BBL97b] Heinz H. Bauschke, Jonathan M. Borwein, and Adrian S. Lewis. The method of cyclic projections for closed convex sets in Hilbert space. In Recent developments in optimization theory and nonlinear analysis (Jerusalem, 1995), volume 204 of Contemp. Math., pages 1–38. American Mathematical Society, Providence, RI, USA, 1997.

Bauschke:1997:SCH

[BBL97c] Heinz H. Bauschke, Jonathan M. Borwein, and Wu Li. Strong conical hull intersection property, bounded linear regularity, Jameson’s property (G), and error bounds in convex optimization. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, September 26, 1997. 30 pp. URL http:

//docserver.carma.newcastle.edu.au/195.

Bauschke:1999:SCH

[BBL99] Heinz H. Bauschke, Jonathan M. Borwein, and Wu Li. Strong conical hull intersection property, bounded linear regularity, Jameson’s property (G), and error bounds in convex optimization. Mathematical Programming, 86(1, Ser. A):135–160, 1999. CO-DEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

URL http://docserver.carma.newcastle.edu.au/195/.

Borwein:2004:DCM

[BBL04] Jonathan M. Borwein, James V. Burke, and Adrian S. Lewis. Differentiability of cone-monotone functions on separable Banach space. Proceedings of the American Mathematical Society, 132 (4):1067–1076, April 2004. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL http://docserver.

carma.newcastle.edu.au/123/1/;http://docserver.carma.
newcastle.edu.au/1242/;https://www.jstor.org/stable/4097135.

Borwein:2010:NSF

[BBL10] David Borwein, Jonathan M. Borwein, and Isaac E. Leonard. L_{p} norms and the sinc function. American Mathematical Monthly, 117(6):528–539, June 2010. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/pdfplus/10.4169/000298910X492817.pdf.

Bailey:2013:SDRa

[BBL^{+}13] D. H. Bailey, J. Borwein, R. J. LeVeque, W. Rider, W. Stein, and V. Stodden. Setting the default to reproducible. Report on the ICERM workshop: Reproducibility in computational and experimental mathematics, December 10–14, 2012. Preprint in collaboration with the workshop participants., January 2013. URL

http://www.carma.newcastle.edu.au/jon/icerm12.pdf.

Bailey:2016:SPDa

[BBL16a] D. Bailey, J. Borwein, and M. López de Prado. Stock portfolio design and backtest overfitting. Journal of Investment Management, ??(??):??, ???? 2016. ISSN 1545-9144 (print), 1545-9152 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
1684/;http://www.davidhbailey.com/dhbpapers/stockfund.
pdf.

Bailey:2016:HHI

[BBL^{+}16b] D. Bailey, J. Borwein, M. López de Prado, A. Salehipour, and Qiji Zhu. How hard is it to avoid backtest overfitting? Submitted to Applied Mathematical Finance, July 2016., 2016.

Bailey:2016:SPDb

[BBL16c] David H. Bailey, Jonathan M. Borwein, and Marcos López de Prado. Stock portfolio design and backtest overfitting. SSRN Electronic Journal, 2016.

Bailey:2013:CPF

[BBLZ13a] D. H. Bailey, J. M. Borwein, M. López de Prado, and Qiji Zhu. Computing the probability of over-fitting in the back-testing and optimization of investment strategies. Preprint., Septem-

ber 2013. URL http://papers.ssrn.com/sol3/papers.cfm?
abstract_id=2308659.

Bailey:2013:IIE

[BBLZ13b] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. Are individual investors equipped to make basic financial decisions. Mathematical Investor, December 30,

2013. URL http://www.financial-math.org/blog/2013/12/
are-individual-investors-equipped-to-make-basic-financial-decisions/.

Bailey:2013:AI

[BBLZ13c] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. The art of investing. Mathematical Investor,

November 15, 2013. URL http://www.financial-math.org/
blog/2013/11/the-art-of-investing/.

Bailey:2013:DLC

[BBLZ13d] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. Do large cap stocks boost portfolio performance at year’s end. Mathematical Investor, November 28,

2013. URL http://www.financial-math.org/blog/2013/11/
do-large-cap-stocks-boost-portfolio-performance-at-years-end/.

Bailey:2013:MHI

[BBLZ13e] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. The myth of the Halloween indicator. Mathematical Investor, November 15, 2013. URL http://www.

financial-math.org/blog/2013/11/the-myth-of-the-halloween-indicator/.

Bailey:2013:TTK

[BBLZ13f] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. Two tales of the Kelly formula. Mathematical Investor, October 27, 2013. URL http://www.financial-math.

org/blog/2013/10/two-tales-of-the-kelly-formula/.

Bailey:2013:TTF

[BBLZ13g] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. The two towers of finance. Mathematical In-

vestor, October 21, 2013. URL http://www.financial-math.
org/blog/2013/10/the-two-towers-of-finance/.

Bailey:2013:WWC

[BBLZ13h] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. What Wittgenstein can teach all of us about investing. Mathematical Investor, December 28,

2013. URL http://www.financial-math.org/blog/2013/12/
what-wittgenstein-can-teach-all-of-us-about-investing/.

Bailey:2014:DNB

[BBLZ14a] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. Do new backtested index ETFs outperform the market. Mathematical Investor, April 23,

2014. URL http://www.financial-math.org/blog/2014/04/
do-new-backtested-index-etfs-outperform-the-market/.

Bailey:2014:DDDb

[BBLZ14b] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. Dubious digits: Is this data really that accurate. Mathematical Investor, November 9,

2014. URL http://www.financial-math.org/blog/2014/11/
dubious-digits-is-this-data-really-that-accurate/.

Bailey:2014:FBO

[BBLZ14c] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. FAQs on backtest overfitting. Mathematical In-

vestor, April 13, 2014. URL http://www.financial-math.org/
blog/2014/04/faqs-on-backtest-overfitting/.

Bailey:2014:FKP

[BBLZ14d] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. Fedspeak, Karl Popper and market directions. Mathematical Investor, January 19, 2014. URL http://www.

financial-math.org/blog/2014/01/fedspeak-karl-popper-and-market-directions/.

Bailey:2014:HFLa

[BBLZ14e] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. How financially literate are individual investors. Mathematical Investor, July 19, 2014. URL http://www.

financial-math.org/blog/2014/07/how-financially-literate-are-individual-investors/.

Bailey:2014:HMP

[BBLZ14f] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. How have 2014 market prophets fared. Mathematical Investor, December 2, 2014. URL http://www.

financial-math.org/blog/2014/12/how-have-2014-market-prophets-fared/.

Bailey:2014:IIC

[BBLZ14g] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. Index investing: ‘confidence in the mathematics’. Mathematical Investor, November 4, 2014. URL http://www.

financial-math.org/blog/2014/11/index-investing-confidence-in-the-mathematics/.

Bailey:2014:CPF

[BBLZ14h] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. Is ‘cherry picking’ a factor in hedge fund performance. Mathematical Investor, September 25,

2014. URL http://www.financial-math.org/blog/2014/09/
is-cherry-picking-a-factor-in-hedge-fund-performance/.

Bailey:2014:SMW

[BBLZ14i] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. Is the stock market weaker during mid-term election years. Mathematical Investor, April 16,

2014. URL http://www.financial-math.org/blog/2014/04/
is-the-stock-market-weaker-during-mid-term-election-years/.

Bailey:2014:LDR

[BBLZ14j] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. Latest DALBAR report underscores poor long-term performance of individual investors. Mathematical Investor, May

9, 2014. URL http://www.financial-math.org/blog/2014/
05/latest-dalbar-report-underscores-poor-long-term-performance-of-individual-investors/.

Bailey:2014:NOT

[BBLZ14k] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. New online tool to demonstrate backtest overfitting. Mathematical Investor, August 28,

2014. URL http://www.financial-math.org/blog/2014/08/
new-online-tool-to-demonstrate-backtest-overfitting/.

Bailey:2014:PMFa

[BBLZ14l] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. Pseudo-mathematics and financial charlatanism. Mathematical Investor, April 4, 2014. URL https://www.

financial-math.org/blog/2014/04/pseudo-mathematics-and-financial-charlatanism/.

Bailey:2014:RDP

[BBLZ14m] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. Review of Dark Pools and Flash Boys. Mathematical

Investor, April 18, 2014. URL https://www.financial-math.
org/blog/2014/04/review-of-dark-pools-and-flash-boys/.

Bailey:2014:SCF

[BBLZ14n] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. The ‘scary chart’ fallacy. Mathematical Investor,

February 16, 2014. URL http://www.financial-math.org/
blog/2014/02/the-scary-chart-fallacy/.

Bailey:2014:SPN

[BBLZ14o] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. Sec to propose new rules for high-frequency trading. Mathematical Investor, June 7, 2014. URL http://www.

financial-math.org/blog/2014/06/sec-to-propose-new-rules-for-high-frequency-trading/.

Bailey:2014:SIA

[BBLZ14p] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. Seminar at the International Association for Quantitative Finance. Mathematical Investor, January 14,

2014. URL http://www.financial-math.org/blog/2014/01/
seminar-at-the-international-association-for-quantitative-finance/.

Bailey:2014:SAF

[BBLZ14q] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. A sobering analysis of financial gurus’ market forecasts. Mathematical Investor, March 30,

2014. URL http://www.financial-math.org/blog/2014/03/
a-sobering-analysis-of-financial-gurus-forecasts/.

Bailey:2014:TEW

[BBLZ14r] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. Testing early warning indicators. Mathematical

Investor, January 5, 2014. URL http://www.financial-math.
org/blog/2014/01/testing-early-warning-indicators/.

Bailey:2014:PMFb

[BBLZ14s] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Qiji Jim Zhu. Pseudo-mathematics and financial charlatanism: the effects of backtest overfitting on outof-sample performance. Notices of the American Mathematical Society, 61(5):458–471, May 2014. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

docserver.carma.newcastle.edu.au/1506/;http://www.ams.
org/notices/201405/rnoti-p458.pdf.

Bailey:2015:HFL

[BBLZ15a] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. Are hedge funds losing their edge. Mathematical Investor, September 16, 2015. URL http://www.financial-math.

org/blog/2015/09/are-hedge-funds-losing-their-edge/.

Bailey:2015:DII

[BBLZ15b] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. Do individual investors understand Social Security and its overseas counterparts. Mathematical Investor, June 16,

2015. URL http://www.financial-math.org/blog/2015/06/
do-individual-investors-understand-social-security-and-its-overseas-counterparts/.

Bailey:2015:HNM

[BBLZ15c] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. High noon for 2015 market prophets. Mathematical Investor, December 24, 2015. URL http://www.

financial-math.org/blog/2015/12/high-noon-for-2015-market-prophets/.

Bailey:2015:HMD

[BBLZ15d] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. How much do investors lose from conflicted advice. Mathematical Investor, February 28, 2015. URL http://www.

financial-math.org/blog/2015/02/how-much-do-investors-lose-from-conflicted-advice/.

Bailey:2015:RFE

[BBLZ15e] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. Is research in finance and economics reproducible. Mathematical Investor, October 10,

2015. URL http://www.financial-math.org/blog/2015/10/
is-research-in-finance-and-economics-reproducible/.

Bailey:2015:LFCa

[BBLZ15f] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. Lessons from the ‘Flash Crash’ regulatory fiasco. Mathematical Investor, April 26, 2015. URL http://www.

financial-math.org/blog/2015/04/lessons-from-the-flash-crash-regulatory-fiasco/.

Bailey:2015:SFE

[BBLZ15g] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. Swiss franc episode exposes risky investments. Mathematical Investor, January 19, 2015. URL http://www.

financial-math.org/blog/2015/01/swiss-franc-episode-exposes-risky-investments/.

Bailey:2016:HWDb

[BBLZ16a] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. How well does a ‘robot AI’ predict the Japanese stock market. Mathematical Investor, February 19,

2016. URL http://www.financial-math.org/blog/2016/02/
how-well-does-a-robot-ai-predict-the-japanese-stock-market/.

Bailey:2016:HWDa

[BBLZ16b] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. How well does the ‘January barometer’ work. Mathematical Investor, January 21, 2016. URL http://www.

financial-math.org/blog/2016/01/how-well-does-the-january-barometer-work/.

Bailey:2016:TTH

[BBLZ16c] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. Tough times for hedge funds. Mathematical In-

vestor, April 22, 2016. URL http://www.financial-math.org/
blog/2016/04/tough-times-for-hedge-funds/.

Bailey:2016:WBF

[BBLZ16d] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Jim Zhu. Where are the billionaire financial academics. Mathematical Investor, July 18, 2016. URL http://www.

financial-math.org/blog/2016/07/where-are-the-billionaire-financial-academics/.

Bailey:2017:PBO

[BBLZ17] David H. Bailey, Jonathan M. Borwein, Marcos López de Prado, and Qiji Jim Zhu. The probability of backtest overfitting. Journal of Computational Finance, 20(4):39–69, April 2017. ISSN 1460-1559 (print), 1755-2850 (electronic). URL http://docserver.

carma.newcastle.edu.au/1710/;http://ssrn.com/abstract=
2326253;https://www.risk.net/journal-of-computational-finance/
2471206/the-probability-of-backtest-overfitting.

Borwein:1999:SM

[BBM99] D. Borwein, J. M. Borwein, and P. Maréchal. Surprise maximization. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, July 25, 1999. 16 pp. URL

http://docserver.carma.newcastle.edu.au/209/. Published

in [BBM00].

Borwein:2000:SM

[BBM00] D. Borwein, J. M. Borwein, and P. Maréchal. Surprise maximization. American Mathematical Monthly, 107(6):517–527, June/ July 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://docserver.carma.newcastle.
edu.au/209/.

Borwein:2001:MVS

[BBM01] David Borwein, Jonathan M. Borwein, and Bernard A. Mares, Jr. Multi-variable sinc integrals and volumes of polyhedra. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, February 9, 2001. 19 pp. URL http://docserver.

carma.newcastle.edu.au/56/.

Borwein:2002:MVS

[BBM02] David Borwein, Jonathan M. Borwein, and Bernard A. Mares, Jr. Multi-variable sinc integrals and volumes of polyhedra. The Ramanujan Journal, 6(2):189–208, 2002. CODEN RA-JOF9. ISSN 1382-4090 (print), 1572-9303 (electronic). URL

http://docserver.carma.newcastle.edu.au/56/.

Bailey:2011:CPI

[BBMW11] David H. Bailey, Jonathan M. Borwein, Andrew Mattingly, and Glenn Wightwick. The computation of previously inaccessible digits of π^{2} and Catalan’s constant. Report, Lawrence Berkeley National Laboratory; Centre for Computer Assisted Research Mathematics and its Applications (CARMA), University of Newcastle; IBM Australia, Berkeley, CA, USA; Callaghan, NSW 2308, Australia; St. Leonards, NSW 2065, Australia; Pyrmont, NSW 2009, Australia, April 11, 2011. 18 pp. URL http:

//crd.lbl.gov/~dhbailey/dhbpapers/bbp-bluegene.pdf.

Bailey:2013:CPI

[BBMW13] David H. Bailey, Jonathan M. Borwein, Andrew Mattingly, and Glenn Wightwick. The computation of previously inaccessible digits of π^{2} and Catalan’s constant. Notices of the American Mathematical Society, 60(7):844–854, August 2013. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL

http://docserver.carma.newcastle.edu.au/1436/;http://
www.ams.org/notices/201307/rnoti-p844.pdf.

Bailey:2016:CPI

[BBMW16] David H. Bailey, Jonathan M. Borwein, Andrew Mattingly, and Glenn Wightwick. The computation of previously inaccessible digits of π. In Bailey and Borwein [BB16l], pages 327–339. ISBN 3-319-32375-X, 3-319-32377-6 (e-book). LCCN

QA251. URL http://docserver.carma.newcastle.edu.au/
1716/;http://lib.myilibrary.com?id=941862.

Bailey:2017:CAA

[BBMW17] David H. Bailey, Jonathan M. Borwein, Andrew Mattingly, and Glenn Wightwick. Computation and analysis of arbitrary digits of Pi and other mathematical constants. Slides for Levi Conant Prize lecture at Worcester Polytechnic Institute, Worcester, MA 01609-2280 USA., September 15, 2017. URL http:

//www.davidhbailey.com/dhbtalks/dhb-conant.pdf.

Borwein:1995:CML

[BBP95] David Borwein, Jonathan M. Borwein, and Christopher Pinner. Convergence of Madelung-like lattice sums. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, June

1995. 46 pp. URL http://docserver.carma.newcastle.edu.
au/104/2/95_040-Borwein-Borwein-Pinner.pdf.

Bailey:1997:RCV

[BBP97] David Bailey, Peter B. Borwein, and Simon Plouffe. On the rapid computation of various polylogarithmic constants. Mathematics of Computation, 66(218):903–913, April 1997. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842

(electronic). URL http://docserver.carma.newcastle.edu.
au/121/;http://www.ams.org/journals/mcom/1997-66-218/
S0025-5718-97-00856-9/S0025-5718-97-00856-9.pdf;https:
//www.jstor.org/stable/2153905;https://en.wikipedia.org/
wiki/Bailey%E2%80%93Borwein%E2%80%93Plouffe_formula. See

[AW97].

Borwein:1998:CML

[BBP98] David Borwein, Jonathan M. Borwein, and Christopher Pinner. Convergence of Madelung-like lattice sums. Transactions of the American Mathematical Society, 350(8):3131–3167, August 1998. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-

6850 (electronic). URL http://docserver.carma.newcastle.
edu.au/104/;https://www.jstor.org/stable/117772.

Benoist:2003:CQV

[BBP03] Joël Benoist, Jonathan M. Borwein, and Nicolae Popovici. A characterization of quasiconvex vector-valued functions. Proceedings of the American Mathematical Society, 131(4):1109–1113, April 2003. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-

6826 (electronic). URL http://docserver.carma.newcastle.
edu.au/145/;https://www.jstor.org/stable/1194189.

Borwein:1989:ACL

[BBS89] D. Borwein, J. M. Borwein, and R. Shail. Analysis of certain lattice sums. Journal of Mathematical Analysis and Applications, 143(1):126–137, 1989. CODEN JMANAK. ISSN 0022-247X (print), 1096-0813 (electronic). URL http://docserver.

carma.newcastle.edu.au/1589/;http://www.sciencedirect.
com/science/article/pii/0022247X89900322.

Borwein:1997:PSS

[BBS^{+}97] David Borwein, Jonathan Borwein, Heinz-Jürgen Seiffert, Robert D. Brown, and Pawel Szeptycki. Problems and solutions: Solutions: 10335. American Mathematical Monthly, 104 (1):72–74, 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Bacak:2010:DCF

[BBS10] M. Bacák, J. Borwein, and B. Sims. Differentiability of convex functions in CAT(0) spaces. Preprint., August 2010.

Borwein:2012:SS

[BBS12] David Borwein, Jonathan M. Borwein, and Armin Straub. A sinc that sank. American Mathematical Monthly, 119(7):535–549, August 2012. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://docserver.carma.newcastle.
edu.au/1391/;https://www.jstor.org/stable/pdfplus/10.
4169/amer.math.monthly.119.07.535.pdf.

Bailey:2013:OSD

[BBS13a] D. H. Bailey, J. M. Borwein, and Victoria Stodden. Opinion: Set the default to “open”. Notices of the American Mathematical Society, 60(6):679–680, June/July 2013. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL

http://www.ams.org/notices/201306/rnoti-p679.pdf.

Borwein:2013:LSW

[BBS13b] D. Borwein, J. M. Borwein, and A. Straub. On lattice sums and Wigner limits. ArXiv e-prints, October

2013. URL http://adsabs.harvard.edu/abs/2013arXiv1310.
1423B;http://docserver.carma.newcastle.edu.au/1509/.

Borwein:2014:SLM

[BBS14a] David Borwein, Jonathan M. Borwein, and Brailey Sims. On the solution of linear mean recurrences. American Mathematical Monthly, 121(6):486–498, June/July 2014. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://docserver.carma.newcastle.edu.au/1505/;https://
www.jstor.org/stable/10.4169.

Borwein:2014:LSW

[BBS14b] David Borwein, Jonathan M. Borwein, and Armin Straub. On lattice sums and Wigner limits. Journal of Mathematical Analysis and Applications, 414(2):489–513, 2014. CODEN JMANAK. ISSN 0022-247X (print), 1096-0813 (electronic).

URL http://arxiv.org/abs/1310.1423;http://docserver.
carma.newcastle.edu.au/1509/;http://www.sciencedirect.
com/science/article/pii/S0022247X14000134.

Bailey:2015:OTD

[BBS^{+}15a] David H. Bailey, Jonathan M. Borwein, Amir Salehipour, Marcos López de Prado, and Qiji Zhu. Online tools for demonstration of backtest overfitting. Technical report, Lawrence Berkeley National Laboratory and four others, Berkeley, CA 94720, USA and others, November 29, 2015. 13 pp. URL http:

//ssrn.com/abstract=2597421.

Borwein:2015:MCR

[BBS15b] D. Borwein, J. M. Borwein, and B. Sims. Monotonicity of certain Riemann sums. Submitted, Mathematics Magazine, Jan-

uary 2015., 2015. URL http://docserver.carma.newcastle.
edu.au/1692/.

Bailey:2016:BOF

[BBS^{+}16a] David H. Bailey, Jonathan M. Borwein, Amir Salehipour, Marcos López de Prado, and Qiji Zhu. Backtest overfitting in financial markets. Automated Trader Magazine, 39(Q1):52–

57, ???? 2016. URL http://docserver.carma.newcastle.
edu.au/1713/;http://www.automatedtrader.net/articles/
strategies/155469/backtest-overfitting-in-financial-markets.

Bailey:2016:FRS

[BBS16b] David H. Bailey, Jonathan M. Borwein, and Victoria Stodden. Facilitating reproducibility in scientific computing: Principles and practice. In Harald Atmanspacher and Sabine Maasen, editors, Reproducibility: Principles, Problems, Practices, and Prospects, pages 205–231. Wiley, New York, NY, USA, 2016. ISBN 1-118-86506-5. This book was awarded the 2017 Prose Award (“Honorable Mention”) in the category “Textbook / Best in Physical Sciences and Mathematics”; see https://

proseawards.com/winners/2017-award-winners/ and http:
//experimentalmath.org/2017/02/reproducibility-principles-problems-practices-and-prospects/.

Borwein:2017:OJM

[BBS17] Judith Borwein, Naomi Borwein, and Brailey Sims. Obituary: Jonathan M. Borwein FAAAS, FBAS, FAustMS, FAA, FAMS, FRSNSW 20 May 1951 to 2 August 2016. Australian Mathematical Society Gazette, 44(5):289–293, November 2017. ISSN 0311-0729 (print), 1326-2297 (electronic). URL http://www.austms.

org.au/Publ/Gazette/2017/Nov17/ObitBorwein.pdf.

Borwein:2020:SMC

[BBS20] David Borwein, Jonathan M. Borwein, and Brailey Sims. Symmetry and the monotonicity of certain Riemann sums. In Bailey et al. [BBB^{+}20], pages 7–20. ISBN 3-030-36567-0 (print), 3-030-36568-9 (e-book). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Bailey:2017:ERMa

[BBSL17a] David H. Bailey, Jonathan M. Borwein, Amir Salehipour, and Marcos López de Prado. Evaluation and ranking of market forecasters. Preprint, Department of Computer Science, University of California, Davis, Davis, CA 95616, USA, March 1, 2017. 22 pp.

URL http://www.davidhbailey.com/dhbpapers/forecaster.
pdf;https://papers.ssrn.com/sol3/papers.cfm?abstract_
id=2944853.

Bailey:2017:ERMb

[BBSL17b] David H. Bailey, Jonathan M. Borwein, Amir Salehipour, and Marcos López de Prado. Evaluation and ranking of market forecasters. SSRN Electronic Journal, 2017.

Bailey:2018:ERM

[BBSL18] David H. Bailey, Jonathan M. Borwein, Amir Salehipour, and Marcos López de Prado. Evaluation and ranking of market forecasters. Journal of Investment Management, 16(2):47–64, April 2018. ISSN 1545-9144 (print), 1545-9152 (electronic).

Bailey:2020:DFG

[BBSL20] David H. Bailey, Jonathan M. Borwein, Amir Salehipour, and Marcos López de Prado. Do financial gurus produce reliable forecasts? In Bailey et al. [BBB^{+}20], pages 255–274. ISBN 3-030-36567-0 (print), 3-030-36568-9 (e-book). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Borwein:2011:LSEb

[BBSW11] David Borwein, Jonathan M. Borwein, Armin Straub, and James Wan. Log-sine evaluations of Mahler measures, II. ArXiv e-

prints, March 2011. URL http://adsabs.harvard.edu/abs/
2011arXiv1103.3035B;http://docserver.carma.newcastle.
edu.au/1435/.

Borwein:2012:LSEb

[BBSW12] David Borwein, Jonathan M. Borwein, Armin Straub, and James Wan. Log-sine evaluations of Mahler measures, II. Integers, 12 (6):1179–1212, 2012. CODEN INTEHN. ISSN 1867-0652 (print),

1867-0660 (electronic). URL http://arxiv.org/abs/1103.
3035;http://docserver.carma.newcastle.edu.au/1435/.

Borwein:1987:ESE

[BBSZ87] D. Borwein, J. M. Borwein, R. Shail, and I. J. Zucker. Energy of static electron lattices. Report, University of Western Ontario and Department of Mathematics, Statistics and Computing Science, Dalhousie University and University of Surrey and University of Surrey, London, ON N6A 3K7, Canada and Halifax, NS B3H 3J5, Canada and Guildford, UK GU2 5XH and Guildford, UK GU2 5XH, September 11, 1987. 20 pp.

Borwein:1988:ESE

[BBSZ88] D. Borwein, J. M. Borwein, R. Shail, and I. J. Zucker. Energy of static electron lattices. Journal of Physics A (Mathematical and General), 21(7):1519–1531, April 1988. CO-DEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (elec-

tronic). URL http://adsabs.harvard.edu/abs/1988JPhA...
21.1519B;http://docserver.carma.newcastle.edu.au/1592/
;http://stacks.iop.org/0305-4470/21/1519;http://stacks.
iop.org/0305-4470/21/i=7/a=015.

Borwein:1985:CLS

[BBT85] David Borwein, Jonathan M. Borwein, and Keith F. Taylor. Convergence of lattice sums and Madelung’s constant. Journal of Mathematical Physics, 26(11):2999–3009, November 1985. CODEN JMAPAQ. ISSN 0022-2488. URL http://adsabs.

harvard.edu/abs/1985JMP....26.2999B;http://docserver.
carma.newcastle.edu.au/1618/.

Borwein:1992:SEC

[BBT92] J. M. Borwein, A. Ben Tal, and M. Teboulle. Spectral estimation via convex programming. In Phillips and Rousseau [PR92], pages 275–290. ISBN 1-4613-6599-6 (print), 1-4615-3600-6 (ebook). LCCN HD30.23.

Bauschke:1998:MRS

[BBT98] Heinz H. Bauschke, Jonathan M. Borwein, and Paul Tseng. Metric regularity, strong CHIP, and CHIP are distinct properties. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, July 20, 1998. 17 pp. URL http://docserver.

carma.newcastle.edu.au/204.

Bauschke:2000:BLR

[BBT00] Heinz H. Bauschke, Jonathan M. Borwein, and Paul Tseng. Bounded linear regularity, strong CHIP, and CHIP are distinct properties. Journal of Convex Analysis, 7(2):395–412, 2000. ISSN 0944-6532 (print), 2363-6394 (electronic). URL http://www.

heldermann.de/JCA/JCA07/JCA072/jca07020.htm.

Borwein:1996:ASC

[BBW96] D. Borwein, J. M. Borwein, and Xianfu Wang. Approximate subgradients and coderivatives in R^{n}. Set-Valued Analysis, 4(4): 375–398, 1996. CODEN SVANEG. ISSN 0927-6947 (print), 1572-

932x (electronic). URL http://docserver.carma.newcastle.
edu.au/152/;http://link.springer.com/article/10.1007/
BF00436112.

Bauschke:2007:FFC

[BBW07] Heinz H. Bauschke, Jonathan M. Borwein, and Xianfu Wang. Fitzpatrick functions and continuous linear monotone operators. SIAM Journal on Optimization, 18(3):789–809, 2007. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic). URL

http://docserver.carma.newcastle.edu.au/319/.

Bauschke:2011:BBT

[BBWY11a] H. H. Bauschke, J. M. Borwein, X. Wang, and L. Yao. The Brézis–Browder theorem in a general Banach space. ArXiv e-

prints, October 2011. URL http://adsabs.harvard.edu/abs/
2011arXiv1110.5706B.

Bauschke:2011:CPM

[BBWY11b] H. H. Bauschke, J. M. Borwein, X. Wang, and L. Yao. Construction of pathological maximally monotone operators on non-reflexive Banach spaces. ArXiv e-prints, August

2011. URL http://adsabs.harvard.edu/abs/2011arXiv1108.
1463B;http://docserver.carma.newcastle.edu.au/1090/.

Bauschke:2011:EMM

[BBWY11c] H. H. Bauschke, J. M. Borwein, X. Wang, and L. Yao. Every maximally monotone operator of Fitzpatrick–Phelps type is actually of dense type. ArXiv e-prints, April 2011. URL http://adsabs.

harvard.edu/abs/2011arXiv1104.0750B;http://docserver.
carma.newcastle.edu.au/1442/.

Bauschke:2011:MOB

[BBWY11d] H. H. Bauschke, J. M. Borwein, X. Wang, and L. Yao. Monotone operators and “bigger conjugate” functions. ArXiv e-

prints, August 2011. URL http://adsabs.harvard.edu/abs/
2011arXiv1108.2578B;http://docserver.carma.newcastle.
edu.au/1477/.

Bauschke:2011:MML

[BBWY11e] Heinz H. Bauschke, Jonathan M. Borwein, Xianfu Wang, and Liangjin Yao. For maximally monotone linear relations, dense type, negative-infimum type, and Fitzpatrick–Phelps type all coincide with monotonicity of the adjoint. ArXiv e-prints, page 15,

March 21, 2011. URL http://adsabs.harvard.edu/abs/
2011arXiv1103.6239B;http://arxiv.org/abs/1103.6239.

Bauschke:2012:BBT

[BBWY12a] Heinz H. Bauschke, Jonathan M. Borwein, Xianfu Wang, and Liangjin Yao. The Brézis–Browder theorem in a general Banach space. Journal of Functional Analysis, 262(12):4948–4971, 2012. CODEN JFUAAW. ISSN 0022-1236 (print), 1096-0783

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022123612001462.

Bauschke:2012:CPM

[BBWY12b] Heinz H. Bauschke, Jonathan M. Borwein, Xianfu Wang, and Liangjin Yao. Construction of pathological maximally monotone operators on non-reflexive Banach spaces. Set-Valued and Variational Analysis, 20(3):387–415, 2012. ISSN 1877-0533 (print),

1877-0541 (electronic). URL http://arxiv.org/abs/1108.
1463;http://docserver.carma.newcastle.edu.au/1090/.

Bauschke:2012:EMM

[BBWY12c] Heinz H. Bauschke, Jonathan M. Borwein, Xianfu Wang, and Liangjin Yao. Every maximally monotone operator of Fitzpatrick– Phelps type is actually of dense type. Optimization Letters, 6(8):1875–1881, 2012. ISSN 1862-4472 (print), 1862-4480

(electronic). URL http://arxiv.org/abs/1104.0750;http://
docserver.carma.newcastle.edu.au/1442/.

Bauschke:2013:MOB

[BBWY13] Heinz H. Bauschke, Jonathan M. Borwein, Xianfu Wang, and Liangjin Yao. Monotone operators and “bigger conjugate” functions. Journal of Convex Analysis, 20(1):143–155, 2013. ISSN 0944-6532 (print), 2363-6394 (electronic). URL

http://docserver.carma.newcastle.edu.au/1477/;http://
www.heldermann.de/JCA/JCA20/JCA201/jca20009.htm.

Borwein:2011:MOE

[BBY11] J. M. Borwein, R. Burachik, and L. Yao. Monotone operators without enlargements. ArXiv e-prints, October 11,

2011. URL http://adsabs.harvard.edu/abs/2011arXiv1110.
3102B;http://arxiv.org/abs/1110.3102;http://docserver.
carma.newcastle.edu.au/1447/.

Borwein:2012:CZD

[BBY12] Jonathan M. Borwein, Regina S. Burachik, and Liangjin Yao. Conditions for zero duality gap in convex programming. ArXiv e-prints, page 10, November 2012. URL http://adsabs.

harvard.edu/abs/2012arXiv1211.4953B;http://docserver.
carma.newcastle.edu.au/1465/.

Borwein:2013:MOE

[BBY13] Jonathan M. Borwein, Regina S. Burachik, and Liangjin Yao. Monotone operators without enlargements. In Bailey et al. [BBB^{+}13], pages 79–103. ISBN 1-4614-7620-8, 1-4614-7621-6 (ebook). ISSN 2194-1009. LCCN QA241. URL http://docserver.

carma.newcastle.edu.au/1447/.

Borwein:2014:CZD

[BBY14] Jonathan M. Borwein, Regina S. Burachik, and Liangjin Yao. Conditions for zero duality gap in convex programming. Journal of Nonlinear and Convex Analysis, 15(1):167– 190, 2014. ISSN 1345-4773 (print), 1880-5221 (electronic).

URL http://arxiv.org/abs/1211.49533;http://docserver.
carma.newcastle.edu.au/1465/.

Borwein:1996:BRE

[BC96] Jonathan M. Borwein and Rob M. Corless. Book review: The Encyclopedia of Integer Sequences, by N. J. A. Sloane and Simon Plouffe. Academic Press, San Diego, CA, 1995. $44.95. xii + 587 pp., hardcover. ISBN 0-12-558630-2. SIAM Review, 38(2): 333–337, June 1996. CODEN SIREAD. ISSN 0036-1445 (print),

1095-7200 (electronic). URL https://www.jstor.org/stable/
2132886.

Borwein:1998:R

[BC98a] Jonathan Borwein and Kwok-Kwong Stephen Choi. On the representations of xy+yz +zx. Report, Department of Mathematics, Simon Fraser University, Burnaby, BC V5A 1S6, Canada, 1998.

8 pp. URL http://docserver.carma.newcastle.edu.au/212.

Borwein:1998:ETE

[BC98b] Jonathan M. Borwein and Robert M. Corless. Emerging tools for experimental mathematics. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, November 27, 1998.

28 pp. URL http://docserver.carma.newcastle.edu.au/203.

Published in [BC99].

Borwein:1999:ETE

[BC99] Jonathan M. Borwein and Robert M. Corless. Emerging tools for experimental mathematics. American Mathematical Monthly, 106(10):889–909, 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://docserver.carma.newcastle.edu.au/203/;https://
www.jstor.org/stable/2589743.

Borwein:2000:R

[BC00] Jonathan Borwein and Kwok-Kwong Stephen Choi. On the representations of xy + yz + zx. Experimental Mathematics, 9(1): 153–158, 2000. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://docserver.carma.newcastle.edu.
au/212/;http://projecteuclid.org/euclid.em/1046889597.

Borwein:2002:DSSc

[BC02] Jonathan Michael Borwein and Kwok-Kwong Stephen Choi. On Dirichlet series for sums of squares. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, February 5,

2002. 29 pp. URL http://docserver.carma.newcastle.edu.

au/142/. Published in [BC03].

Borwein:2003:DSS

[BC03] Jonathan Michael Borwein and Kwok-Kwong Stephen Choi. On Dirichlet series for sums of squares. The Ramanujan Journal, 7(1–3):95–127, 2003. CODEN RAJOF9. ISSN 1382-4090 (print), 1572-9303 (electronic). URL http://docserver.carma. newcastle.edu.au/142/. Rankin memorial issues.

Borwein:2004:RAFb

[BC04a] J. Borwein and R. Crandall. On the Ramanujan AGM fraction, II: The complex-parameter case. Experimental Mathematics, 13(3): 287–295, 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://docserver.carma.newcastle.edu.
au/29/;http://projecteuclid.org/euclid.em/1103749837.

Borwein:2004:DSS

[BC04b] Jonathan Michael Borwein and Kwok-Kwong Stephen Choi. Dirichlet series for squares of sums of squares: a summary. In Kisilevsky and Goren [KG04], pages 27–34. ISBN 0-8218-3331-6. LCCN QA241 .N8645 2004.

Borwein:2007:IPA

[BC07] Jonathan M. Borwein and Marc Chamberland. Integer powers of arcsin. International Journal of Mathematics and Mathematical Sciences, pages Art. ID 19381, 10, June 2007. ISSN 0161-1712 (print), 1687-0425 (electronic). URL http://docserver.carma.

newcastle.edu.au/287/.

Borwein:2009:UBC

[BC09] Jonathan M. Borwein and O-Yeat Chan. Uniform bounds for the complementary incomplete gamma function. Mathematical Inequalities & Applications, 12(1):115–121, 2009. ISSN 1331-4343 (print), 1848-9966 (electronic). URL http://docserver.carma.

newcastle.edu.au/335/.

Borwein:2010:DTM

[BC10] Jonathan M. Borwein and O-Yeat Chan. Duality in tails of multiple-zeta values. International Journal of Number Theory, 6(3):501–514, May 2010. ISSN 1793-0421 (print), 1793-7310

(electronic). URL http://docserver.carma.newcastle.edu.
au/1218/;https://www.worldscientific.com/doi/10.1142/
S1793042110003058.

Borwein:2013:CFW

[BC13] Jonathan M. Borwein and Richard E. Crandall. Closed forms: what they are and why we care. Notices of the American Mathematical Society, 60(1):50–65, January 2013. CO-DEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
767/;http://www.ams.org/notices/201301/rnoti-p50.pdf.

Borwein:2015:PPB

[BC15a]  Jonathan M. Borwein and Scott T. Chapman.  I pre-

fer pi: a brief history and anthology of articles in the American Mathematical Monthly. American Mathematical Monthly, 122(3):195–216, March 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//docserver.carma.newcastle.edu.au/1615/;https://www.
jstor.org/stable/10.4169;https://www.jstor.org/stable/
10.4169/amer.math.monthly.122.03.195. See addenda [BC15b].

Borwein:2015:PPA

[BC15b] Jonathan M. Borwein and Scott T. Chapman. I prefer pi: Addenda. American Mathematical Monthly, 122(8):800, October 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL https://www.jstor.org/stable/10.4169/
amer.math.monthly.122.8.800. See [BC15a]. The addenda cor-

rect formulas (4) and (12), references [34] and [74], and add a new reference.

Borwein:2016:PPB

[BC16] Jonathan M. Borwein and Scott T. Chapman. I prefer pi: a brief history and anthology of articles in the American Mathematical Monthly. In Bailey and Borwein [BB16l], pages 475–499. ISBN 3-319-32375-X, 3-319-32377-6 (e-book). LCCN QA251. URL http:

//docserver.carma.newcastle.edu.au/1615/.

Borwein:2018:GFM

[BC18a] Jonathan M. Borwein and Robert M. Corless. Gamma and factorial in the Monthly. American Mathematical Monthly, 125(5): 400–424, May 2018. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

BoAa:2018:ANN

[BC18b] Radu Ioan Boţ and Ernö Robert Csetnek. Approaching nonsmooth nonconvex optimization problems through first order dynamical systems with hidden acceleration and Hessian driven damping terms. Set-Valued and Variational Analysis, 26(2):227–245, June 2018. CODEN ???? ISSN 1877-0533 (print), 1877-0541 (electronic). URL http://link.

springer.com/article/10.1007/s11228-017-0411-1;https:
//link.springer.com/content/pdf/10.1007/s11228-017-0411-1.
pdf.

Bailey:2021:ISI

[BC21] David H. Bailey and Scott Chapman. Introduction to the special issue. American Mathematical Monthly, 128(9):772, 2021. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Borwein:2010:HDB

[BCC10] Jonathan M. Borwein, O-Yeat Chan, and R. E. Crandall. Higher-dimensional box integrals. Experimental Mathematics, 19(4):431–446, 2010. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://docserver.

carma.newcastle.edu.au/1223/;http://projecteuclid.org/
euclid.em/1317758103.

Borwein:2004:RAFa

[BCF04] J. Borwein, R. Crandall, and G. Fee. On the Ramanujan AGM fraction, I: The real-parameter case. Experimental Mathematics, 13(3):275–285, 2004. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). URL http://docserver.carma.

newcastle.edu.au/27/;http://projecteuclid.org/euclid.
em/1103749836.

Borwein:2004:OPV

[BCFR04] Jonathan Borwein, Lixin Cheng, Marián Fabian, and Julian P. Revalski. A one perturbation variational principle and applications. Set-Valued Analysis, 12(1–2):49–60, 2004. CODEN SVANEG. ISSN 0927-6947 (print), 1572-932x (electronic). URL

http://docserver.carma.newcastle.edu.au/14/.

Beliakov:2013:EIBb

[BCJW13] G. Beliakov, D. Creighton, M. Johnstone, and T. Wilkin. Efficient implementation of Bailey and Borwein pseudo-random number generator based on normal numbers. Computer Physics Communications, 184(8):1999–2004, 2013. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0010465513001276.

Borwein:2016:CLA

[BCLM16] Jonathan M. Borwein, Neil J. Calkin, Scott B. Lindstrom, and Andrew Mattingly. Continued logarithms and associated continued fractions. Experimental Mathematics, 26(4):412–429, September 2016. ISSN 1058-6458 (print), 1944-950X (electronic). URL

http://docserver.carma.newcastle.edu.au/1687/.

Borwein:2017:CLA

[BCLM17] Jonathan M. Borwein, Neil J. Calkin, Scott B. Lindstrom, and Andrew Mattingly. Continued logarithms and associated continued fractions. Experimental Mathematics, 26(4):412–429, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2016.1195307.

Borwein:2002:PDA

[BCM02] J. Borwein, R. Choksi, and P. Maréchal. Probability distributions of assets inferred from option prices via the Principle of Maximum Entropy. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, January 7, 2002. 19

pp. URL http://docserver.carma.newcastle.edu.au/10/2/
02_3A176-Borwein-Choksi-Marechal.pdf.

Borwein:2003:PDA

[BCM03] J. Borwein, R. Choksi, and P. Maréchal. Probability distributions of assets inferred from option prices via the principle of maximum entropy. SIAM Journal on Optimization, 14(2):464– 478, 2003. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-

7189 (electronic). URL http://docserver.carma.newcastle.
edu.au/10/.

Borwein:2009:EBS

[BCM09] Jonathan M. Borwein, Neil J. Calkin, and Dante Manna. Euler– Boole summation revisited. American Mathematical Monthly, 116 (5):387–412, 2009. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Baake:2020:BCL

[BCM20] Michael Baake, Michael Coons, and Neil Mañibo. Binary constant-length substitutions and Mahler measures of Borwein polynomials. In Bailey et al. [BBB^{+}20], pages 303–322. ISBN 3-030-36567-0 (print), 3-030-36568-9 (e-book). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Borwein:2005:CFT

[BCP05] Jonathan Michael Borwein, Kwok-Kwong Stephen Choi, and Wilfried Pigulla. Continued fractions of tails of hypergeometric series. American Mathematical Monthly, 112(6):493–501, June/ July 2005. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://docserver.carma.newcastle.
edu.au/39/;https://www.jstor.org/stable/30037519.

Borwein:1986:LOC

[BD86] J. M. Borwein and M. A. H. Dempster. The linear order complementarity problem. Report DALTR 86-3, Department of Mathematics, Statistics & Computing Science, Dalhousie University and Balliol College, Oxford University, Halifax, NS, Canada and Oxford, England OX1 3BJ, July 1986. 47 pp.

Borwein:1989:LOC

[BD89] J. M. Borwein and M. A. H. Dempster. The linear order complementarity problem. Mathematics of Operations Research, 14(3):534–558, 1989. CODEN MOREDQ. ISSN 0364-765x (print), 1526-5471 (electronic). URL http://docserver.carma.

newcastle.edu.au/1585/.

Borwein:1995:MCF

[BD95] Jonathan M. Borwein and Kenneth R. Davidson. Mathematics in Canada: the future of mathematics in Canada 50 years later. In Canadian Mathematical Society. 1945–1995, volume 1 of CMS Publications, pages 231–268 (249–268 françiai). Canadian Mathematical Society, Ottawa, ON, Canada, 1995. Dual English–French text.

Borwein:2003:BGT

[BD03] J. M. Borwein and A. L. Dontchev. On the Bartle–Graves theorem. Proceedings of the American Mathematical Society, 131(8):2553–2560, 2003. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL http://docserver.carma.

newcastle.edu.au/128/.

Borwein:2009:CCI

[BD09] Jonathan M. Borwein and Keith J. Devlin. The computer as crucible: an introduction to experimental mathematics. A. K. Peters, Ltd., Wellesley, MA, USA, 2009. ISBN 1-56881-343-0. xi + 158 pp. LCCN QA8.7 .B67 2009. URL http://

docserver.carma.newcastle.edu.au/1730/;http://www.loc.
gov/catdir/toc/fy0904/2008022180.html.

Borwein:2011:EMB

[BD11] Jonathan Borwein and Keith Devlin. Experimentelle Mathematik: Eine beispielorientierte Einführung. (German) [Experimental mathematics: an example-oriented introduction]. Spektrum Akademischer Verlag, Heidelberg, Germany, 2011. ISBN 3-8274-2661-8, 3-8274-2662-6 (e-book). xii + 158 pp. LCCN

QA8.7 .B67 2011. URL http://link.springer.com/book/10.

1007/978-3-8274-2662-8. German translation by Roland Girgensohn of [BD09], with updates.

Borwein:2015:MMI

[BD15] Jonathan M. Borwein and Joydeep Dutta. Maximal monotone inclusions and Fitzpatrick functions. Journal of Optimization Theory and Applications, ??(??):1–28, 2015. CODEN JOTABN. ISSN 0022-3239 (print), 1573-2878 (electronic). URL http:

//docserver.carma.newcastle.edu.au/1715/;http://link.
springer.com/article/10.1007/s10957-015-0813-x. Special

issue of JOTA on Nondifferentiable Optimization and Nonsmooth Analysis, dedicated to Vladimir Demyanov.

Borwein:2016:DFE

[BD16a] J. M. Borwein and Karl Dilcher. Derivatives and fast evaluation of the Witten zeta function. Submitted to the Ramanujan Journal, April 2016., 2016.

Borwein:2016:MMI

[BD16b] J. M. Borwein and J. Dutta. Maximal monotone inclusions and Fitzpatrick functions. Journal of Optimization Theory and Applications, 171(3):757–784, December 2016. CODEN JOTABN. ISSN 0022-3239 (print), 1573-2878 (electronic).

Borwein:2018:DFE

[BD18] Jonathan M. Borwein and Karl Dilcher. Derivatives and fast evaluation of the Tornheim zeta function. The Ramanujan Journal, 45(2):413–432, February 2018. ISSN 1382-4090 (print), 1572-9303 (electronic).

Borwein:1991:NR

[BdB91] Jonathan M. Borwein and G. de Barra. Nested radicals. American Mathematical Monthly, 98(8):735–739, 1991. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

docserver.carma.newcastle.edu.au/1580/.

Borwein:1996:PSS

[BDT96] Jonathan M. Borwein, Donald A. Darling, and D. B. Tyler. Problems and solutions: Solutions: 10281. American Mathematical Monthly, 103(2):181–183, 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Borwein:2016:BOM

[BDT16] J. M. Borwein, K. Dilcher, and H. Tomkins. The behaviour at the origin of multiple Witten zeta functions. In preparation., July 2016.

Borwein:2008:BST

[BE08] J. M. Borwein and A. C. Eberhard. Banach spaces of type (NI) and monotone operators on non-reflexive spaces. Preprint., June 2008.

Breen:2016:DBJ

[BE16] Mike Breen and Annette Emerson. David Bailey, Jonathan Borwein, Andrew Mattingly, and Glenn Wightwick to receive 2017 AMS Conant Prize. AMS press release., December 2, 2016. URL

http://www.ams.org/news?news_id=3232.

Bean:2013:MDM

[Bea13] Nigel Bean, editor. Modelling for decision making in ecological systems [… a session … at the MODSIM 2011 Conference in Perth, Australia, 12–16 December 2011], volume 249.2013 of Ecological modelling Special issue. Elsevier, Amsterdam, The Netherlands, 2013. LCCN ????

Bejancu:1994:EBP

[Bej94] Aurelian Bejancu. On the Ekeland and Borwein–Preiss principles in finite dimensions. An. Ştiinţ. Univ. Al. I. Cuza Iaşi Secţ. I a Mat., 40(1):63–67, 1994. CODEN AUZMAV. ISSN 0041-9109.

Borwein:1976:VS

[BEO76] J. Borwein, M. Edelstein, and R. O’Brien. Visibility and starshape. Journal of the London Mathematical Society (series 2), S2-14(2):313–318, November 1976. CODEN JLMSAK. ISSN 0024-6107 (print), 1469-7750 (electronic). URL http:

//docserver.carma.newcastle.edu.au/1670/;http://jlms.
oxfordjournals.org/content/s2-14/2/313.full.pdf+html.

Borwein:1977:SRV

[BEO77] J. Borwein, M. Edelstein, and R. O’Brien. Some remarks on visibility and starshape. Journal of the London Mathematical Society (series 2), S2-15(2):342–344, April 1977. CODEN JLM-SAK. ISSN 0024-6107 (print), 1469-7750 (electronic). URL http:

//docserver.carma.newcastle.edu.au/1666/;http://jlms.
oxfordjournals.org/content/s2-15/2/342.full.pdf+html.

Berndt:1988:BRJ

[Ber88] Bruce C. Berndt. Book review: Jonathan M. Borwein and Peter B. Borwein, Pi and the AGM — A Study of Analytic Number Theory and Computational Complexity, Canadian Mathematical Society Series of Mono- graphs and Advanced Texts, Wiley, New York, 1987, xv + 414 pp., 24 cm. Price $49.95. Mathematics of Computation, 50(181):352–354, January 1988. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

https://www.jstor.org/stable/2007942.

Borwein:2011:MOT

[BEY11] Jonathan M. Borwein, Andrew Eberhard, and Liangjin Yao. Monotone operators of type FPV. Preprint., August 2011.

Borwein:1989:LBM

[BF89a] Jon Borwein and Simon Fitzpatrick. Local boundedness of monotone operators under minimal hypotheses. Bulletin of the Australian Mathematical Society, 39(3):439–441, 1989. CODEN ALNBAB. ISSN 0004-9727 (print), 1755-1633 (electronic). URL

http://docserver.carma.newcastle.edu.au/1590/.

Borwein:1989:ENP

[BF89b] Jonathan M. Borwein and Simon Fitzpatrick. Existence of nearest points in Banach spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 41(4):702–720, 1989. CO-DEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

URL http://docserver.carma.newcastle.edu.au/1582/.

Borwein:1989:MCK

[BF89c] Jonathan M. Borwein and Simon Fitzpatrick. Mosco convergence and the Kadec property. Proceedings of the American Mathematical Society, 106(3):843–851, 1989. CODEN PAM-YAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL

http://docserver.carma.newcastle.edu.au/1584/.

Borwein:1993:CFH

[BF93a] J. M. Borwein and M. Fabián. On convex functions having points of Gâteaux differentiability which are not points of Fréchet differentiability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 45(6):1121–1134, 1993. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL

http://docserver.carma.newcastle.edu.au/1552/.

Borwein:1993:GSO

[BF93b] J. M. Borwein and M. Fabián. On generic second-order Gâteaux differentiability. Nonlinear Analysis, Theory, Methods and Applications, 20(12):1373–1382, 1993. CODEN NOANDD. ISSN 0362-546x (print), 1873-5215 (electronic). URL http://docserver.

carma.newcastle.edu.au/1563/;http://www.sciencedirect.
com/science/article/pii/0362546X9390166P.

Borwein:1993:WSC

[BF93c] Jonathan Borwein and Simon Fitzpatrick. Weak sequential compactness and bornological limit derivatives. Report, Department of Mathematics and Statistics, Simon Fraser University, Burnaby, BC V5A 1S6, Canada, September 7, 1993. 10

pp. URL http://docserver.carma.newcastle.edu.au/58/2/
93_001-Borwein-Borwein-Girgensohn.pdf.

Borwein:1993:WHS

[BF93d] Jonathan M. Borwein and Simon Fitzpatrick. A weak Hadamard smooth renorming of L_{1}(Ω, μ). Canadian mathematical bulletin = Bulletin canadien de mathématiques, 36(4):407–413, 1993. CO-DEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

URL http://cms.math.ca/10.4153/CMB-1993-055-5;http://
docserver.carma.newcastle.edu.au/1554/.

Borwein:1994:CCS

[BF94a] Jonathan Borwein and Simon Fitzpatrick. Characterization of Clarke subgradients among one-dimensional multifunctions. Report, Department of Mathematics, Simon Fraser University, Burnaby, BC V5A 1S6, Canada, February 2, 1994. 14 pp. URL

http://docserver.carma.newcastle.edu.au/23/. Published

in [BF95a].

Borwein:1994:NRS

[BF94b] Jonathan M. Borwein and Marián Fabián. A note on regularity of sets and of distance functions in Banach space. Journal of Mathematical Analysis and Applications, 182(2):566–570, 1994. CO-DEN JMANAK. ISSN 0022-247X (print), 1096-0813 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022247X84711048.

Borwein:1995:CCS

[BF95a] J. M. Borwein and S. Fitzpatrick. Characterization of Clarke subgradients among one-dimensional multifunctions. In Proceedings of the Optimization Miniconference II, pages 5–12. University of Ballarat, Ballarat, VI, Australia, 1995. URL http:

//docserver.carma.newcastle.edu.au/23/.

Borwein:1995:SCB

[BF95b] Jonathan Borwein and Simon Fitzpatrick. Weak^{∗} sequential compactness and bornological limit derivatives. Journal of Convex Analysis, 2(1–2):59–67, 1995. ISSN 0944-6532 (print), 2363-6394

(electronic). URL http://www.heldermann.de/JCA/jcacon.
htmvol.2_no.1+2/j4_49.ps.gz. Dedicated to R. T. Rockafel-

lar on his 60th birthday.

Borwein:1995:CCH

[BF95c] Jonathan M. Borwein and Simon Fitzpatrick. Closed convex Haar null sets. CECM Research Report 95:052, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, August 1995. 9

pp. URL http://docserver.carma.newcastle.edu.au/117/.

Borwein:1998:DIS

[BF98] J. M. Borwein and S. P. Fitzpatrick. Duality inequalities and sandwiched functions. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, July 8, 1998. 15 pp. URL

http://docserver.carma.newcastle.edu.au/204.

Borwein:2001:DIS

[BF01] J. M. Borwein and S. P. Fitzpatrick. Duality inequalities and sandwiched functions. Nonlinear Analysis, Theory, Methods and Applications, 46(3):365–380, 2001. CODEN NOANDD. ISSN 0362-546x (print), 1873-5215 (electronic). URL http://docserver.

carma.newcastle.edu.au/204/;http://www.sciencedirect.
com/science/article/pii/S0362546X00001292.

Borwein:2006:MS

[BF06a] Jonathan Borwein and Helaman Ferguson. The mathematics of salt. Preprint., December 2006.

Borwein:2006:MKM

[BF06b] Jonathan M. Borwein and William Michael Farmer, editors. Mathematical knowledge management: 5th international conference, MKM 2006, Wokingham, UK, August 11– 12, 2006: proceedings, volume 4108 of Lecture notes in computer science; Lecture notes in artificial intelligence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 3-540-37104-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.95

.M56 2006. URL http://www.loc.gov/catdir/enhancements/
fy0825/2006930246-d.html;http://www.loc.gov/catdir/enhancements/
fy1402/2006930246-t.html.

Borwein:1987:DRF

[BFG87] J. M. Borwein, S. P. Fitzpatrick, and J. R. Giles. The differentiability of real functions on normed linear space using generalized subgradients. Journal of Mathematical Analysis and Applications, 128(2):512–534, 1987. CODEN JMANAK. ISSN 0022-247X (print), 1096-0813 (electronic). URL http://docserver.

carma.newcastle.edu.au/1598/;http://www.sciencedirect.
com/science/article/pii/0022247X87902034.

Borwein:2003:SWG

[BFG03] Jonathan Borwein, Simon Fitzpatrick, and Roland Girgensohn. Subdifferentials whose graphs are not norm × weak^{∗} closed. Canadian mathematical bulletin = Bulletin canadien de mathématiques, 46(4):538–545, December 2003. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

URL http://cms.math.ca/10.4153/CMB-2003-051-5;http://
docserver.carma.newcastle.edu.au/149/.

Borwein:1991:MCU

[BFK91] Jonathan Borwein, Simon Fitzpatrick, and Petar Kenderov. Minimal convex uscos and monotone operators on small sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 43(3):461–476, 1991. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL http://docserver.carma.

newcastle.edu.au/1573/.

Borwein:2000:RGC

[BFKL00] J. M. Borwein, M. Fabian, I. Kortezov, and P. D. Loewen. The range of the gradient of a continuously differentiable bump. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, October 12, 2000. 21 pp. URL http://docserver.

carma.newcastle.edu.au/231/. Published in [BFKL01].

Borwein:2001:RGC

[BFKL01] J. M. Borwein, M. Fabian, I. Kortezov, and P. D. Loewen. The range of the gradient of a continuously differentiable bump. Journal of Nonlinear and Convex Analysis, 2(1):1–19, 2001. ISSN 1345-4773 (print), 1880-5221 (electronic). URL http://

docserver.carma.newcastle.edu.au/231/;http://www.ybook.

co.jp/online2/opjnca/vol2/p1.html. Special issue for Profes-

sor Ky Fan.

Borwein:2002:RGL

[BFL02] J. M. Borwein, M. Fabian, and P. D. Loewen. The range of the gradient of a Lipschitz C^{1}-smooth bump in infinite dimensions. Israel Journal of Mathematics, 132(1):239–251, December 2002. CODEN ISJMAP. ISSN 0021-2172 (print), 1565-

8511 (electronic). URL http://docserver.carma.newcastle.
edu.au/144/;http://link.springer.com/article/10.1007/
BF02784514.

Borwein:1993:LLFa

[BFV93a] J. M. Borwein, M. Fabian, and J. Vanderwerff. Locally Lipschitz functions and bornological derivatives. ArXiv Mathematics e-prints, page 13, February 1993. URL http://adsabs.

harvard.edu/abs/1993math......2211B;http://docserver.
carma.newcastle.edu.au/69/.

Borwein:1993:LLFb

[BFV93b] Jonathan M. Borwein, Marian Fabian, and Jon Vanderwerff. Locally Lipschitz functions and bornological derivatives. CECM Research Report 93:012, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, December 15, 1993. 13 pp. URL

http://docserver.carma.newcastle.edu.au/69/.

Borwein:1994:ECFc

[BFV94a] Jon Borwein, Simon Fitzpatrick, and Jon Vanderwerff. Examples of convex functions and classifications of normed spaces. Journal of Convex Analysis, 1(1):61–73, 1994. ISSN 0944-6532 (print), 2363-6394 (electronic). URL http://docserver.carma.

newcastle.edu.au/64/;http://www.heldermann-verlag.de/
jca/jca01/jca01004.pdf.

Borwein:1994:ECFa

[BFV94b] Jon Borwein, Simon Fitzpatrick, and Jon Vanderwerff. Examples of convex functions and classifications of normed spaces. Report, Department of Mathematics & Statistics, Simon Fraser University, Burnaby, BC V5A 156, Canada, March 1, 1994. 16 pp. URL

http://docserver.carma.newcastle.edu.au/64. Published in

[BFV94a].

Borwein:1994:ECFb

[BFV94c] Jonathan M. Borwein., Simon Fitzpatrick, and Jon Vanderwerff. Examples of convex functions and classification of normed spaces. Published in [BFV94a]. Plenary talk, VII Colloque Franco–Allemand d’Optimisation, Dijon, France., June 27, 1994.

URL http://docserver.carma.newcastle.edu.au/64/.

Borwein:1997:CBS

[BFV97] Jon Borwein, Marian Fabian, and Jon Vanderwerff. Characterizations of Banach spaces via convex and other locally Lipschitz functions. Acta Mathematica Vietnamica, 22(1):53–69, 1997. ISSN 0251-4184 (print), 2315-4144 (electronic). URL http:

//docserver.carma.newcastle.edu.au/179/. Special edition in

honour of Huang Tuy’s seventieth birthday.

Borwein:1987:PNF

[BG87] J. M. Borwein and J. R. Giles. The proximal normal formula in Banach space. Transactions of the American Mathematical Society, 302(1):371–381, 1987. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic). URL http://docserver.

carma.newcastle.edu.au/1602/.

Borwein:1994:FEDa

[BG94a] Jonathan M. Borwein and Roland Girgensohn. Functional equations and distribution functions. Report, Department of Mathematics & Statistics, Simon Fraser University, Burnaby, BC V5A 156, Canada, July 15, 1994. 10 pp. URL http://docserver. carma.newcastle.edu.au/90. Dedicated to Professor János Aczél’s on the occasion of his 70th birthday.

Borwein:1994:FEDb

[BG94b] Jonathan M. Borwein and Roland Girgensohn. Functional equations and distribution functions. Results in Mathematics. Resultate der Mathematik, 26(3–4):229–237, 1994. ISSN 0378-

6218. URL http://docserver.carma.newcastle.edu.au/90/.

For János Aczél’s 70th birthday.

Borwein:1995:ATB

[BG95a] Jonathan M. Borwein and Roland Girgensohn. Addition theorems and binary expansions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 47(2):262–273, 1995. CO-DEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

URL http://docserver.carma.newcastle.edu.au/1537/.

Borwein:1995:ETE

[BG95b] Jonathan M. Borwein and Roland Girgensohn. Evaluation of triple Euler sums. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, 1995. 27 pp.

URL http://docserver.carma.newcastle.edu.au/118/2/95_
053-Borwein-Girgensohn.pdf.

Borwein:1996:ADE

[BG96a] J. M. Borwein and F. G. Garvan. Approximations to π via the Dedekind eta function. Report, Department of Mathematics, Simon Fraser University, Burnaby, BC V5A 1S6, Canada, March

27, 1996. 28 pp. URL http://docserver.carma.newcastle.
edu.au/158/.

Borwein:1996:ETE

[BG96b] Jonathan M. Borwein and Roland Girgensohn. Evaluation of triple Euler sums. Electronic Journal of Combinatorics, 3(1): Research Paper 23, 1996. ISSN 1077-8926 (print), 1097-1440 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
118/;http://www.combinatorics.org/Volume_3/Abstracts/
v3i1r23.html.

Borwein:1997:APD

[BG97a] J. M. Borwein and F. Garvan. Approximating pi with the Dedekind eta function. In ????, editor, Organic Mathematics Proceedings, 12 April 1996, page ?? ????, ????, 1997. ISSN 0731-

1036. URL http://www.cecm.sfu.ca/organics.

Borwein:1997:ADE

[BG97b] J. M. Borwein and F. G. Garvan. Approximations to π via the Dedekind eta function. In Borwein et al. [BBJC97], pages 89–115. ISBN 0-8218-0668-8. ISSN 0731-1036. LCCN QA1 .O67 1995.

URL http://docserver.carma.newcastle.edu.au/158/.

Borwein:2001:NRI

[BG01] J. Borwein and R. Goebel. Notions of relative interior in Banach spaces. Technical report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, October 10, 2001. 17 pp. URL

http://docserver.carma.newcastle.edu.au/138/. Published

in [BG03a].

Borwein:2003:NRI

[BG03a] J. Borwein and R. Goebel. Notions of relative interior in Banach spaces. Journal of Mathematical Sciences (New York), 115(4):2542–2553, 2003. CODEN JMTSEW. ISSN 1072-3374 (print), 1573-8795 (electronic). URL http://docserver.

carma.newcastle.edu.au/138/;http://link.springer.com/
article/10.1023/A%3A1022988116044. Special issue on opti-

mization and related topics, 1.

Borwein:2003:CEI

[BG03b] Jonathan Borwein and Roland Girgensohn. A class of exponential inequalities. Mathematical Inequalities & Applications, 6(3):397– 411, 2003. ISSN 1331-4343 (print), 1848-9966 (electronic). URL

http://docserver.carma.newcastle.edu.au/148/.

Borwein:2005:EBS

[BG05] Jonathan M. Borwein and Roland Girgensohn. Evaluations of binomial series. Aequationes Mathematicae, 70(1–2):25–36, 2005. CODEN AEMABN. ISSN 0001-9054 (print), 1420-8903

(electronic). URL http://docserver.carma.newcastle.edu.
au/131/.

Borwein:2009:NCM

[BG09] Jonathan Borwein and Rafal Goebel. On the nondifferentiability of cone-monotone functions in Banach spaces. In Optimization, volume 32 of Springer Optim. Appl., pages 3–14. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009.

Borwein:2015:CAGa

[BG15a] J. M. Borwein and O. Giladi. Convex analysis in groups and semigroups: a sampler. ArXiv e-prints, October 2015. URL http://

adsabs.harvard.edu/abs/2015arXiv151004480B;http://docserver.
carma.newcastle.edu.au/1693/.

Borwein:2015:NPD

[BG15b] J. M. Borwein and O. Giladi. Nearest points and delta convex functions in Banach spaces. ArXiv e-prints, October 2015.

URL http://adsabs.harvard.edu/abs/2015arXiv151004471B.

Preprint of [BG16b].

Borwein:2015:SRC

[BG15c] Jonathan M. Borwein and Ohad Giladi. Some remarks on convex analysis in topological groups. ArXiv e-prints, October 2015. URL

http://adsabs.harvard.edu/abs/2015arXiv151004487B;http:
//docserver.carma.newcastle.edu.au/1697/.

Borwein:2016:EBD

[BG16a] J. M. Borwein and Ohad Giladi. Ergodic behaviour of a Douglas– Rachford operator away from the origin. Submitted to the Journal of Global Optimization, May 2016., 2016.

Borwein:2016:NPD

[BG16b] Jonathan M. Borwein and Ohad Giladi. Nearest points and delta convex functions in Banach spaces. Bulletin of the Australian Mathematical Society, 93(2):283–294, April 2016. CODEN ALNBAB. ISSN 0004-9727 (print), 1755-1633 (electronic). URL

http://journals.cambridge.org/action/displayAbstract?
fromPage=online&aid=10230840.

Borwein:2016:SRC

[BG16c] Jonathan M. Borwein and Ohad Giladi. Some remarks on convex analysis in topological groups. Journal of Convex Analysis, 23(2):313–332, ???? 2016. ISSN 0944-6532 (print), 2363-6394 (electronic). URL http://docserver.

carma.newcastle.edu.au/1697/;http://www.heldermann.de/
JCA/JCA23/JCA232/jca23012.htm.

Borwein:2018:CAG

[BG18a] Jonathan M. Borwein and Ohad Giladi. Convex analysis in groups and semigroups: a sampler. Mathematical Programming, 168(1–2 (Series B)):11–53, 2018. CODEN MHPGA4. ISSN 0025-5610.

Borwein:2018:EBD

[BG18b] Jonathan M. Borwein and Ohad Giladi. Ergodic behaviour of a Douglas–Rachford operator away from the origin. Journal of Nonlinear and Convex Analysis, 19(8):1395–1407, 2018. ISSN 1345-4773 (print), 1880-5221 (electronic).

Borwein:2009:UCF

[BGHV09] J. Borwein, A. J. Guirao, P. Hájek, and J. Vanderwerff. Uniformly convex functions on Banach spaces. Proceedings of the American Mathematical Society, 137(3):1081–1091, 2009. CO-DEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic).

URL http://docserver.carma.newcastle.edu.au/340/.

Borwein:1993:CEU

[BGL93] J. M. Borwein, R. K. Goodrich, and M. A. Limber. A comparison of entropies in the underdetermined moment problem. Preprint., May 1993.

Borwein:2013:LST

[BGM^{+}13] J. M. Borwein, M. L. Glasser, R. C. McPhedran, J. G. Wan, and I. J. Zucker. Lattice Sums: Then and Now, volume 150 of Encyclopedia of Mathematics and its Applications. Cambridge University Press, Cambridge, UK, 2013. ISBN 1-107-03990-8. xx +

368 pp. URL http://ebooks.cambridge.org/ebook.jsf?bid=

CBO9781139626804. With a foreword by Helaman Ferguson and Claire Ferguson.

Beer:2018:UCN

[BGM18] Gerald Beer, M. Isabel Garrido, and Ana S. Meroño. Uniform continuity and a new bornology for a metric space. Set-Valued and Variational Analysis, 26(1):49–65, March 2018. CODEN ???? ISSN 1877-0533 (print), 1877-0541 (electronic). URL http://

link.springer.com/article/10.1007/s11228-017-0429-4.

Bauschke:2021:EAA

[BGMS21] Heinz H. Bauschke, Sylvain Gretchko, Walaa M. Moursi, and Matthew Saurette. Edelstein’s astonishing affine isometry. American Mathematical Monthly, 128(9):796–809, 2021. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Borwein:2002:RNC

[BGV02] Jon Borwein, John Giles, and Jon Vanderwerff. Rotund norms, Clarke subdifferentials and extensions of Lipschitz functions. Nonlinear Analysis, Theory, Methods and Applications, 48(2):287– 301, 2002. CODEN NOANDD. ISSN 0362-546x (print), 1873-5215

(electronic). URL http://docserver.carma.newcastle.edu.
au/219/;http://www.sciencedirect.com/science/article/
pii/S0362546X00001875.

Borwein:1997:CHP

[BGW97] J. M. Borwein, R. Girgensohn, and Xianfu Wang. On the construction of Hölder and proximal subderivatives. Report, Department of Mathematics, Simon Fraser University, Burnaby, BC V5A 1S6, Canada, March 14, 1997. 13 pp. URL http://docserver.

carma.newcastle.edu.au/186/. Published in [BGW98].

Borwein:1998:CHP

[BGW98] J. M. Borwein, R. Girgensohn, and Xianfu Wang. On the construction of Hölder and proximal subderivatives. Canadian mathematical bulletin = Bulletin canadien de mathématiques, 41(4): 497–507, 1998. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-

4287 (electronic). URL http://docserver.carma.newcastle.
edu.au/186/.

Borwein:1994:UCMa

[BH94a] J. M. Borwein and W. Huang. Uniform convergence for moment problems with Fermi–Dirac type entropies. Report, Department of Mathematics & Statistics, Simon Fraser University, Burnaby, BC V5A 156, Canada, January 7, 1994. 11 pp. URL http:

//docserver.carma.newcastle.edu.au/70/.

Borwein:1994:UCMb

[BH94b] J. M. Borwein and W. Huang. Uniform convergence for moment problems with Fermi–Dirac type entropies. ZOR — Mathematical Methods of Operations Research, 40(3):239–252, 1994. CODEN ZMRREP. ISSN 0340-9422. URL http://docserver.carma.

newcastle.edu.au/70/;http://link.springer.com/article/
10.1007/BF01432968.

Borwein:1995:FHMa

[BH95] J. M. Borwein and W. Z. Huang. A fast heuristic method for polynomial moment problems with Boltzmann–Shannon entropy. SIAM Journal on Optimization, 5(1):68–99, 1995. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic). URL

http://docserver.carma.newcastle.edu.au/1539/.

Borwein:2006:SCM

[BH06] Jonathan M. Borwein and Chris H. Hamilton. Symbolic computation of multidimensional Fenchel conjugates. In Proceedings of the 2006 International Symposium on Symbolic and Algebraic Computation: ISSAC 2006, ISSAC ’06, pages 23–30. ACM Press, New York, NY 10036, USA, 2006. ISBN 1-59593-276-3. URL

http://doi.acm.org/10.1145/1145768.1145780.

Borwein:2009:SFC

[BH09] Jonathan M. Borwein and Chris H. Hamilton. Symbolic Fenchel conjugation. Mathematical Programming, 116(1–2):17–35, 2009. CODEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/

1479/. Special issue in honor of Alfred Auslender.

Borwein:2019:CCM

[BH19] Jonathan Borwein and Phil Howlett. Checkerboard copulas of maximum entropy with prescribed mixed moments. Journal of the Australian Mathematical Society, 107(3):302–318, 2019. CODEN JAUMAX. ISSN 1446-7887.

Borwein:2016:GCLb

[BHL16a] J. M. Borwein, K. Hare, and J. Lynch. Generalized continued logarithms and related continued fractions. Submitted to Math of Computation, June 2016., 2016.

Borwein:2016:GCLa

[BHL16b] J. M. Borwein, K. G. Hare, and J. G. Lynch. Generalized continued logarithms and related continued fractions. ArXiv

e-prints, June 2016. URL http://adsabs.harvard.edu/abs/
2016arXiv160606984B.

Borwein:2017:GCL

[BHL17] Jonathan M. Borwein, Kevin G. Hare, and Jason G. Lynch. Generalized continued logarithms and related continued fractions. Journal of Integer Sequences, 20(5):Art. 17.5.7, 51, 2017. ISSN 1530-7638.

Borwein:2014:MSS

[BHP14] Jonathan Borwein, Phil Howlett, and Julia Piantadosi. Modelling and simulation of seasonal rainfall using the principle of maximum entropy. Entropy, 16(2):747–769, February 2014. CODEN ENTRFG. ISSN 1099-4300. URL http://adsabs.

harvard.edu/abs/2014Entrp..16..747B;http://docserver.
carma.newcastle.edu.au/1507/;http://www.mdpi.com/1099-4300/
16/.

Borwein:1995:PAS

[BI95] Jonathan M. Borwein and Alexander Ioffe. Proximal analysis in smooth spaces. Report, Department of Mathematics & Statistics, Simon Fraser University, Burnaby, BC V5A 156, Canada, Febru-

ary 21, 1995. 30 pp. URL http://docserver.carma.newcastle.
edu.au/61.

Borwein:1996:PAS

[BI96] Jonathan M. Borwein and Alexander Ioffe. Proximal analysis in smooth spaces. Set-Valued Analysis, 4(1):1–24, 1996. CODEN SVANEG. ISSN 0927-6947 (print), 1572-932x (electronic). URL

http://docserver.carma.newcastle.edu.au/61/.

Borwein:1997:NSP

[BJ97] Jonathan M. Borwein and Alejandro Jofré. A nonconvex separation property in Banach spaces. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, November 13, 1997.

11 pp. URL http://docserver.carma.newcastle.edu.au/197.

Published in [BJ98].

Borwein:1998:NSP

[BJ98] Jonathan M. Borwein and Alejandro Jofré. A nonconvex separation property in Banach spaces. Zeitschrift für Operations Research. Mathematical Methods of Operations Research, 48(2): 169–179, 1998. CODEN ZMMRFZ. ISSN 1432-2994 (print), 1432-

5217 (electronic). URL http://docserver.carma.newcastle.

edu.au/197/. Set-valued optimization.

Borwein:2012:OMT

[BJ12] Jonathan Borwein and Veselin Jungić. Organic mathematics: then and now. Notices of the American Mathematical Society, 59(3):416–419, March 2012. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL http://

docserver.carma.newcastle.edu.au/1397/;http://www.ams.
org/notices/201203/rtx120300416p.pdf.

Beliakov:2013:EIBa

[BJCW13] Gleb Beliakov, Michael Johnstone, Doug Creighton, and Tim Wilkin. An efficient implementation of Bailey and Borwein’s algorithm for parallel random number generation on graphics processing units. Computing: Archiv für Informatik und Numerik, 95(4):309–326, April 2013. CODEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (electronic). URL http://link.

springer.com/article/10.1007/s00607-012-0234-8. See also

[BCJW13].

Borwein:2008:CCC

[BJL^{+}08] J. Borwein, V. Jungić, D. Langstroth, M. Macklem, and S. Wilson. Coast-to-coast (C2C) seminar: Background, history, and practice. In Borwein et al. [BRR08], chapter 9, pages 127–140. ISBN 1-56881-410-0. LCCN QA76.95 .C59 2008. URL http:

//www.crcnetbase.com/doi/pdfplus/10.1201/b10587-12.

Borwein:2000:ANB

[BJSM00] J. M. Borwein, M. Jiménez-Sevilla, and J. P. Moreno. Antiproximinal norms in Banach spaces. Technical report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, Octo-

ber 2000. 13 pp. URL http://docserver.carma.newcastle.
edu.au/233/.

Borwein:2002:ANB

[BJSM02] J. M. Borwein, M. Jiménez-Sevilla, and J. P. Moreno. Antiproximinal norms in Banach spaces. Journal of Approximation Theory, 114(1):57–69, 2002. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic). URL http://docserver.

carma.newcastle.edu.au/233/;http://www.sciencedirect.
com/science/article/pii/S0021904501936366.

Borwein:1980:HMC

[BK80] J. Borwein and L. Keener. The Hausdorff metric and Čebyšev centres. Journal of Approximation Theory, 28(4):366–376, 1980. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430

(electronic). URL http://docserver.carma.newcastle.edu.
au/1652/;http://www.sciencedirect.com/science/article/
pii/0021904580900714.

Borwein:1983:FDS

[BK83] J. M. Borwein and K. O. Kortanek. Fenchel-duality and separably-infinite programs. Optimization, 14(1):37–48., 1983. CODEN OPTZDQ. ISSN 0323-3898.

Borwein:2001:SGR

[BK01] Jonathan M. Borwein and Ivaylo Kortezov. Some generic results on nonattaining functionals. Set-Valued Analysis, 9(1–2): 35–47, 2001. CODEN SVANEG. ISSN 0927-6947 (print), 1572-

932x (electronic). URL http://docserver.carma.newcastle.

edu.au/221/. Wellposedness in optimization and related topics (Gargnano, 1999).

Borwein:2004:CMU

[BK04] J. Borwein and I. Kortezov. Constructive minimal uscos. Comptes rendus de l’Académie bulgare des sciences, 57(12):9–12, 2004. CO-DEN DBANAD. ISSN 1310-1331 (print), 2367-5535 (electronic).

URL http://docserver.carma.newcastle.edu.au/273/.

Borwein:2005:ADC

[BK05] J. M. Borwein and K. Karamanos. Algebraic dynamics of certain gamma function values. In Andrew Eberhard, Nicolas Hadjisavvas, and Dinh The Luc, editors, Generalized Convexity, Generalized Monotonicity and Applications, volume 77 of Nonconvex Optim. Appl., pages 3–21. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN

0-387-23638-4. URL http://docserver.carma.newcastle.
edu.au/256/;http://link.springer.com/chapter/10.1007/
0-387-23639-2_1.

Byerly:1993:PSP

[BKL^{+}93] Robert E. Byerly, Murray S. Klamkin, A. Liu, Stephen M. Gagola, Jr., L. E. Mattics, Raphael M. Robinson, M. Al-Ahmar, Donald E. Knuth, and Jonathan M. Borwein. Problems and solutions: Problems: 10274–10281. American Mathematical Monthly, 100(1):75– 77, 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Borwein:2002:FES

[BKW02] J. M. Borwein, I. Kortezov, and H. Wiersma. A C^{1}-function that is even on a sphere and has no critical points in the ball. Journal of Nonlinear and Convex Analysis, 3(1):1–16, 2002. ISSN 1345-4773 (print), 1880-5221 (electronic). URL http://docserver.carma.

newcastle.edu.au/140/;http://www.ybook.co.jp/online2/
jncav3.html.

Beck:1987:PSS

[BL87] Anatole Beck and O. P. Lossers. Problems and solutions: Solutions of advanced problems: 6492. American Mathematical Monthly, 94(1):82–83, January 1987. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BB85].

Borwein:1991:CBE

[BL91a] Jonathan M. Borwein and Adrian S. Lewis. Convergence of best entropy estimates. SIAM Journal on Optimization, 1(2):191– 205, 1991. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-

7189 (electronic). URL http://docserver.carma.newcastle.
edu.au/1576/.

Borwein:1991:DRE

[BL91b] Jonathan M. Borwein and Adrian S. Lewis. Duality relationships for entropy-like minimization problems. SIAM Journal on Control and Optimization, 29(2):325–338, 1991. CODEN SJ-CODE. ISSN 0363-0129 (print), 1095-7138 (electronic). URL

http://docserver.carma.newcastle.edu.au/1577/.

Borwein:1991:CMP

[BL91c] Jonathan M. Borwein and Adrian S. Lewis. On the convergence of moment problems. Transactions of the American Mathematical Society, 325(1):249–271, May 1991. CO-DEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
1579/;https://www.jstor.org/stable/2001670.

Borwein:1991:PCF

[BL91d] Jonathan M. Borwein and Adrian S. Lewis. Practical conditions for Fenchel duality in infinite dimensions. In J-B. Baillon and M. Théra, editors, Fixed point theory and applications (Marseille, 1989), volume 252 of Pitman Lecture Notes in Mathematics, pages 83–89. Longman Scientific and Technical, Harlow, Essex, UK,

Borwein:1992:ASW

[BL92a] Jonathan Borwein and Shi Tuo Lou. Asymptotics of a sequence of Witt vectors. Journal of Approximation Theory, 69(3):326– 337, 1992. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430

(electronic). URL http://docserver.carma.newcastle.edu.
au/1565/;http://www.sciencedirect.com/science/article/
pii/002190459290006A.

Borwein:1992:DMF

[BL92b] Jonathan M. Borwein and Adrian S. Lewis. Decomposition of multivariate functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 44(3):463–482, 1992. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL

http://docserver.carma.newcastle.edu.au/1569/.

Borwein:1992:PFCa

[BL92c] Jonathan M. Borwein and Adrian S. Lewis. Partially finite convex programming. I. Quasi relative interiors and duality theory. Mathematical Programming, 57(1–3):15–48, 1992. CODEN MH-PGA4. ISSN 0025-5610 (print), 1436-4646 (electronic). URL

http://docserver.carma.newcastle.edu.au/1571/;http://
link.springer.com/article/10.1007/BF01581072.

Borwein:1992:PFCb

[BL92d] Jonathan M. Borwein and Adrian S. Lewis. Partially finite convex programming, Part II: Explicit lattice models. Mathematical Programming, 57(1–3):49–83, 1992. CODEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic). URL http://docserver.

carma.newcastle.edu.au/1572/;http://link.springer.com/
article/10.1007/BF01581073.

Borwein:1992:MHP

[BL92e] Jonathan M. Borwein and Mark A. Limber. Maple as a high precision calculator. Maple Technical Newsletter, 0(8):39–44, Fall

1992. ISSN 1061-5733. URL http://www.can.nl/Systems_and_
Packages/Per_Purpose/General/Maple/mtn/mtn8.html.

Borwein:1993:CDS

[BL93a] Jonathan M. Borwein and Adrian S. Lewis. Convergence of decreasing sequences of convex sets in nonreflexive Banach spaces. Set-Valued Analysis, 1(4):355–363, 1993. CODEN SVANEG. ISSN 0927-6947 (print), 1572-932x (electronic). URL http:

//docserver.carma.newcastle.edu.au/1551/.

Borwein:1993:PFP

[BL93b] Jonathan M. Borwein and Adrian S. Lewis. Partially-finite programming in L_{1} and the existence of maximum entropy estimates. SIAM Journal on Optimization, 3(2):248–267, 1993. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic). URL

http://docserver.carma.newcastle.edu.au/1558/.

Borwein:1993:SCR

[BL93c] Jonathan M. Borwein and Adrian S. Lewis. A survey of convergence results for maximum entropy methods. In A. Mohammad-Djafari and G. Demoments, editors, Maximum entropy and Bayesian methods (Paris, 1992), volume 53 of Fund. Theories Phys., pages 39–48. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1993. URL http://

link.springer.com/chapter/10.1007/978-94-017-2217-9_5.

Borwein:1994:UMP

[BL94a] J. M. Borwein and M. A. Limber. Underdetermined moment problems: a case for convex analysis. SIAG/OPT, ??(5):9–13, Fall 1994. Invited talk at 1994 SIAM Conference on Optimization.

Borwein:1994:SRO

[BL94b] Jonathan M. Borwein and Adrian S. Lewis. Strong rotundity and optimization. SIAM Journal on Optimization, 4(1):146– 158, 1994. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-

7189 (electronic). URL http://docserver.carma.newcastle.
edu.au/1550/.

Borwein:1997:AIR

[BL97] Jonathan M. Borwein and Petr Lisoněk. Applications of integer relation algorithms. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, November 18, 1997.

URL http://docserver.carma.newcastle.edu.au/198. Pub-

lished in [BL00b].

Borwein:1999:IAM

[BL99] Jonathan M. Borwein and June Lester. Issues for active math and math labs. Issues for next generation telelearning systems, Telelearning 1999, Montreal, QC, Canada, November 6–9., November 9, 1999.

Borwein:2000:CAN

[BL00a] Jonathan M. Borwein and Adrian S. Lewis. Convex Analysis and Nonlinear Optimization. CMS Books in Mathematics/Ouvrages de Mathématiques de la SMC, 3. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 0-387-98940-4. x + 273 pp. Theory and examples.

Borwein:2000:AIR

[BL00b] Jonathan M. Borwein and Petr Lisoněk. Applications of integer relation algorithms. Discrete Mathematics, 217(1–3):65–82, 2000. CODEN DSMHA4. ISSN 0012-365x (print), 1872-681x (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
198/;http://www.sciencedirect.com/science/article/pii/

S0012365X99002563. Formal power series and algebraic combinatorics (Vienna, 1997).

Borwein:2005:DCF

[BL05] Jonathan M. Borwein and D. Russell Luke. Dynamics of a continued fraction of Ramanujan with random coefficients. Abstract and Applied Analysis, 2005(5):449–467, ???? 2005. ISSN 1085-3375 (print), 1687-0409 (electronic). URL http://

adsabs.harvard.edu/abs/2005AbApA2005..449B;http://www.
hindawi.com/journals/aaa/2005/162798/abs/.

Borwein:2006:CAN

[BL06] Jonathan M. Borwein and Adrian S. Lewis. Convex analysis and nonlinear optimization, volume 3 of CMS Books in Mathematics/ Ouvrages de Mathématiques de la SMC. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2006. ISBN 0-387-29570-4. xii + 310 pp. Theory and examples.

Borwein:2008:DRT

[BL08] Jonathan M. Borwein and D. Russell Luke. Dynamics of a Ramanujan-type continued fraction with cyclic coefficients. The Ramanujan Journal, 16(3):285–304, 2008. CODEN RAJOF9. ISSN 1382-4090 (print), 1572-9303 (electronic). URL http:

//docserver.carma.newcastle.edu.au/1480/.

Borwein:2011:ERF

[BL11] Jonathan M. Borwein and D. Russell Luke. Entropic regularization of the ℓ_{0} function. In Bauschke et al. [BBC^{+}11b], pages 65– 92. ISBN 1-4419-9568-4, 1-4419-9569-2 (e-book). ISSN 1931-6828 (print), 1931-6836 (electronic). LCCN QA378.5 .F59 2011. URL

http://docserver.carma.newcastle.edu.au/1451/;http://
link.springer.com/chapter/10.1007/978-1-4419-9569-8_5.

Borwein:2015:DCP

[BL15] Jonathan M. Borwein and D. Russell Luke. Duality and convex programming. In Scherzer [Sch15], pages 257–304. ISBN 1-4939-0789-1 (set), 1-4939-0790-5 (e-book). LCCN RC78.7.D53

H358 2015. URL http://docserver.carma.newcastle.edu.au/
1217/.

Borwein:2016:MLO

[BL16] Jonathan M. Borwein and Scott B. Lindstrom. Meetings with Lambert W and other special functions in optimization and analysis. Pure and Applied Functional Analysis, 1(3):361–396, ???? 2016. ISSN 2189-3756 (print), 2189-3764 (electronic). URL

http://www.ybook.co.jp/online2/oppafa/vol1/p361.html.

Bao:2017:NPBa

[BL17a] Yanyan Bao and Hongwei Liu. Nonmonotone projected Barzilai– Borwein method for compressed sensing. In 2017 2nd International Conference on Image, Vision and Computing (ICIVC), pages 756–760. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2017.

Bao:2017:NPBb

[BL17b] Yanyan Bao and Hongwei Liu. Nonmonotone projected Barzilai– Borwein method for compressed sensing. In 2017 International Conference on Smart Grid and Electrical Automation (ICSGEA), pages 264–268. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2017.

Burachik:2020:I

[BL20] Regina S. Burachik and Guoyin Li. Introduction. In Bailey et al. [BBB^{+}20], pages 3–5. ISBN 3-030-36567-0 (print), 3-030-36568-9 (e-book). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Borwein:1994:EML

[BLL94] Jonathan M. Borwein, Adrian S. Lewis, and Mark A. Limber. Entropy minimization with lattice bounds. Journal of Approximation Theory, 79(1):1–16, 1994. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic). URL http://docserver.

carma.newcastle.edu.au/1545/;http://www.sciencedirect.
com/science/article/pii/S0021904584711105.

Borwein:1994:MES

[BLLN94] J. M. Borwein, A. S. Lewis, M. N. Limber, and D. Noll. Maximum entropy spectral analysis using derivative information part 2: Computational results. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, January 10, 1994. 19 pp.

URL http://docserver.carma.newcastle.edu.au/71.

Borwein:1995:MER

[BLLN95] J. M. Borwein, A. S. Lewis, M. N. Limber, and D. Noll. Maximum entropy reconstruction using derivative information. II. Computational results. Numerische Mathematik, 69(3):243–256, 1995. CO-DEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic).

URL http://docserver.carma.newcastle.edu.au/71/;http:
//link.springer.com/article/10.1007/s002110050090.

Borwein:1996:AGM

[BLM96] Jonathan A. Borwein, Petr Lisoněk, and John A. Macdonald. Arithmetic–geometric means revisited. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, 1996. 8 pp.

URL http://docserver.carma.newcastle.edu.au/4/.

Borwein:1997:AGMb

[BLM97] Jonathan A. Borwein, Petr Lisoněk, and John A. Macdonald. Arithmetic–geometric means revisited. Maple Technical Newsletter, 4(1):20–27, Winter 1997. ISSN 1061-5733. URL http:

//docserver.carma.newcastle.edu.au/4/. Special issue on

Maple in the mathematical sciences.

Borwein:1999:CEL

[BLM99] Jonathan Borwein, Yves Lucet, and Boris Mordukhovich. Compactly epi-Lipschitzian convex sets and functions in normed spaces. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, June 9, 1999. 20 pp. URL http://docserver.

carma.newcastle.edu.au/210/. Published in [BLM00].

Borwein:2000:CEL

[BLM00] Jonathan Borwein, Yves Lucet, and Boris Mordukhovich. Compactly epi-Lipschitzian convex sets and functions in normed spaces. Journal of Convex Analysis, 7(2):375–393, 2000. ISSN 0944-6532 (print), 2363-6394 (electronic). URL http://

docserver.carma.newcastle.edu.au/210/;http://www.heldermann.
de/JCA/JCA07/JCA072/jca07019.htm.

Borwein:2007:CCS

[BLM^{+}07] Jonathan M. Borwein, David Langstroth, Mason Macklem, Scott Wilson, and V. Jungić. The coast-to-coast seminar and remote mathematical collaboration. In 21st International Symposium on High Performance Computing Systems and Applications. HPCS 2007, Saskatoon, SK, Canada, 13–16 May 2007, pages 23–29. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, May 2007.

Borwein:1994:MER

[BLN94a] J. M. Borwein, A. S. Lewis, and D. Noll. Maximum entropy reconstruction using derivative information. Part 1: Fisher information and convex duality. Report, Department of Combinatorics and Optimization University of Waterloo, Waterloo, ON N2L 3G1, Canada, December 14, 1994. 31 pp. URL

http://docserver.carma.newcastle.edu.au/82.

Borwein:1994:EMPa

[BLN94b] Jonathan M. Borwein, Adrian S. Lewis, and Roger D. Nussbaum. Entropy minimization, DAD problems, and doubly stochastic kernels. Journal of Functional Analysis, 123(2):264–307, 1994. CODEN JFUAAW. ISSN 0022-1236 (print), 1096-0783

(electronic). URL http://docserver.carma.newcastle.edu.
au/1546/;http://www.sciencedirect.com/science/article/
pii/S0022123684710895.

Borwein:1995:FHMb

[BLN95] J. M. Borwein, M. N. Limber, and D. Noll. Fast heuristic methods for function reconstruction using derivative information. Applicable Analysis, 58(3–4):241–261, 1995. CODEN APANCC. ISSN

0003-6811. URL http://docserver.carma.newcastle.edu.au/
1536/.

Borwein:1996:MER

[BLN96] J. M. Borwein, A. S. Lewis, and D. Noll. Maximum entropy reconstruction using derivative information. I. Fisher information and convex duality. Mathematics of Operations Research, 21(2):442– 468, 1996. CODEN MOREDQ. ISSN 0364-765x (print), 1526-

5471 (electronic). URL http://docserver.carma.newcastle.
edu.au/82/.

Borwein:2016:GDR

[BLS^{+}16] J. Borwein, S. Lindstrom, A. Schneider, B. Sims, and M. Skerritt. Generalisation of the Douglas–Rachford method from simple spheres to ellipses and p-spheres. Preprint., June 2016.

Borwein:2017:DDR

[BLS^{+}17] Jonathan M. Borwein, Scott B. Lindstrom, Brailey Sims, Anna Schneider, and Matthew P. Skerritt. Dynamics of the Douglas– Rachford method for ellipses and p-spheres. Set-Valued and Variational Analysis, ??(??):1–19, October 2017. ISSN 1877-0533 (print), 1877-0541 (electronic).

Borwein:2018:DDR

[BLS^{+}18] Jonathan M. Borwein, Scott B. Lindstrom, Brailey Sims, Anna Schneider, and Matthew P. Skerritt. Dynamics of the Douglas– Rachford method for ellipses and p-spheres. Set-Valued and Variational Analysis, 26(2):385–403, June 2018. CODEN ???? ISSN 1877-0533 (print), 1877-0541 (electronic). URL http://link.

springer.com/article/10.1007/s11228-017-0457-0.

Borwein:2015:CRA

[BLT15] J. M. Borwein, G. Li, and M. K. Tam. Convergence rate analysis for averaged fixed point iterations in the presence of hölder regularity. ArXiv e-prints, October 2015. URL http://adsabs.

harvard.edu/abs/2015arXiv151006823B;http://docserver.
carma.newcastle.edu.au/1690/.

Borwein:2016:CRA

[BLT16] Jonathan M. Borwein, Gouyin Li, and Matthew Tam. Convergence rate analysis for averaged fixed point iterations in the presence of Hölder regularity. Revised SIAM Optimization, April

2016., 2016. URL http://arxiv.org/abs/1510.06823;http:
//docserver.carma.newcastle.edu.au/1690/.

Borwein:2017:CRA

[BLT17] Jonathan M. Borwein, Guoyin Li, and Matthew K. Tam. Convergence rate analysis for averaged fixed point iterations in common fixed point problems. SIAM Journal on Optimization, 27(1):1– 33, January 2017. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

Borwein:2013:ACR

[BLY13] J. M. Borwein, G. Li, and L. Yao. Analysis of the convergence rate for the cyclic projection algorithm applied to basic semi-algebraic convex sets. ArXiv e-prints, April 2013. URL http://adsabs.

harvard.edu/abs/2013arXiv1304.7965B;http://docserver.
carma.newcastle.edu.au/1508/.

Borwein:2014:ACR

[BLY14] Jonathan M. Borwein, Guoyin Li, and Liangjin Yao. Analysis of the convergence rate for the cyclic projection algorithm applied to basic semialgebraic convex sets. SIAM Journal on Optimization, 24(1):498–527, 2014. CODEN SJOPE8. ISSN 1052-6234 (print),

1095-7189 (electronic). URL http://arxiv.org/abs/1304.
7965;http://docserver.carma.newcastle.edu.au/1508/.

Borwein:1999:CSFb

[BLZ99] Jonathan M. Borwein, Adrian S. Lewis, and Qiji J. Zhu. Convex spectral functions of compact operators, Part II: Lower semicontinuity and rearrangement invariance. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, 1999. 18 pp.

URL http://docserver.carma.newcastle.edu.au/251/.

Borwein:2001:CSF

[BLZ01] Jonathan M. Borwein, Adrian S. Lewis, and Qiji J. Zhu. Convex spectral functions of compact operators. II. Lower semicontinuity and rearrangement invariance. In Optimization and related topics (Ballarat/Melbourne, 1999), volume 47 of Appl. Optim., pages 179–196. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. URL http://docserver.

carma.newcastle.edu.au/251/.

Borwein:1995:ESD

[BM95] Jonathan M. Borwein and Warren B. Moors. Essentially strictly differentiable Lipschitz functions. Technical report, Department of Mathematics & Statistics, Simon Fraser University, Burnaby, BC V5A 156, Canada, March 5, 1995. 50 pp. URL

http://docserver.carma.newcastle.edu.au/95. West Coast

Optimization Meeting, University of Washington, Seattle, WA, USA.

Borwein:1996:CRE

[BM96a] Jonathan M. Borwein and Warren B. Moors. A chain rule for essentially smooth Lipschitz functions. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, January 16,

1996. URL http://docserver.carma.newcastle.edu.au/151/.

Published in [BM98a].

Borwein:1996:NSE

[BM96b] Jonathan M. Borwein and Warren B. Moors. Null sets and essentially smooth Lipschitz functions. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, June 4, 1996. 22

pp. URL http://docserver.carma.newcastle.edu.au/162/.

Published in [BM98b].

Borwein:1997:NEW

[BM97a] Jon Borwein and Michael Monagan. A note from the Editors: Winter 1997, vol. 4, no. 1. Maple Technical Newsletter, 4(1):1, Winter 1997. ISSN 1061-5733.

Borwein:1997:DEM

[BM97b] Jonathan Borwein and Michael Monagan. Differential equations in Maple: a short tutorial. Maple Technical Newsletter, 4(1):55– 58, Winter 1997. ISSN 1061-5733.

Borwein:1997:ESLa

[BM97c] Jonathan M. Borwein and Warren B. Moors. Essentially smooth Lipschitz functions. Journal of Functional Analysis, 149(2):305– 351, 1997. CODEN JFUAAW. ISSN 0022-1236 (print), 1096-0783

(electronic). URL http://docserver.carma.newcastle.edu.
au/95/;http://www.sciencedirect.com/science/article/pii/
S0022123697931013.

Borwein:1997:ESLb

[BM97d] Jonathan M. Borwein and Warren B. Moors. Essentially smooth Lipschitz functions: compositions and chain rules. In F. Cucker and M. Shub, editors, Foundations of computational mathematics (Rio de Janeiro, 1997), pages 16–22. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997.

Borwein:1997:LFM

[BM97e] Jonathan M. Borwein and Warren B. Moors. Lipschitz functions with minimal Clarke subdifferential mappings. In B. M. Glover and D. Ralph, editors, Proceedings of the Optimization Miniconference III, University of Melbourne, July 1996, pages 5– 12. University of Ballarat, Ballarat, VI, Australia, 1997. URL

http://docserver.carma.newcastle.edu.au/178/.

Borwein:1997:SDIb

[BM97f] Jonathan M. Borwein and Warren B. Moors. Separable determination of integrability and minimality of the Clarke subdifferential mapping. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, September 30, 1997. 10 pp.

URL http://docserver.carma.newcastle.edu.au/196. Pub-

lished in [BM00].

Borwein:1998:CRE

[BM98a] Jonathan M. Borwein and Warren B. Moors. A chain rule for essentially smooth Lipschitz functions. SIAM Journal on Optimization, 8(2):300–308, 1998. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic). URL http://docserver.carma.

newcastle.edu.au/151/.

Borwein:1998:NSE

[BM98b] Jonathan M. Borwein and Warren B. Moors. Null sets and essentially smooth Lipschitz functions. SIAM Journal on Optimization, 8(2):309–323, 1998. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic). URL http://docserver.carma.

newcastle.edu.au/162/.

Borwein:2000:SDI

[BM00] Jonathan M. Borwein and Warren B. Moors. Separable determination of integrability and minimality of the Clarke subdifferential mapping. Proceedings of the American Mathematical Society, 128(1):215–221, January 2000. CODEN PAM-YAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL

http://docserver.carma.newcastle.edu.au/196/;https://
www.jstor.org/stable/119403.

Borwein:2006:DLP

[BM06] Jonathan M. Borwein and Mason S. Macklem. The (digital) life of Pi. Australian Mathematical Society Gazette, 33(4):243– 248, September 2006. ISSN 0311-0729 (print), 1326-2297 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
267/;http://www.austms.org.au/Publ/Gazette/2006/Sep06/
pi.pdf.

Borwein:2007:REM

[BM07a] J. M. Borwein and Mason Macklem. Retro-enhancement of mathematical literature. Preprint., March 2007. URL http:

//docserver.carma.newcastle.edu.au/339/.

Borwein:2007:CSR

[BM07b] Jonathan M. Borwein and M. Macklem. The C2C seminar and remote mathematical collaboration. Presentation to e+Calculus Conference Lisbon, Portugal., February 8, 2007.

Borwein:2007:CNS

[BM07c] Jonathan M. Borwein and Warren B. Moors. Chapter 50: Nonsmooth analysis, optimisation theory and Banach space theory.

In Elliott Pearl, editor, Open Problems in Topology II, pages 549–559. Elsevier, Amsterdam, The Netherlands, 2007. ISBN 0-

444-52208-5. URL http://www.sciencedirect.com/science/
article/pii/B9780444522085500508.

Borwein:2007:NAA

[BM07d] Jonathan M. Borwein and Warren B. Moors. Nonsmooth analysis, approximation theory and Banach space theory. In Pearl [Pea07], pages 557–566. ISBN 0-444-52208-5, 0-08-047529-9 (ebook). LCCN QA611 .O562 2007.

Borwein:2009:SCC

[BM09] Jonathan M. Borwein and Warren B. Moors. Stability of closedness of convex cones under linear mappings. Journal of Convex Analysis, 16(3–4):699–705, 2009. ISSN 0944-6532 (print), 2363-6394 (electronic). URL http://docserver.

carma.newcastle.edu.au/1229/;http://www.heldermann.de/
JCA/JCA16/JCA163/jca16041.htm.

Borwein:2010:SCC

[BM10] Jonathan M. Borwein and Warren B. Moors. Stability of closedness of convex cones under linear mappings II. Journal of Nonlinear Analysis and Optimization: Theory & Applications, 1(1):1–7, 2010. ISSN 1906-9685. URL

http://docserver.carma.newcastle.edu.au/1226/;http://
www.math.sci.nu.ac.th/ojs234/index.php/jnao/article/view/
5.

Bonnefond:2018:NMP

[BMCL18] Xavier Bonnefond, Pierre Maréchal, Walter Cedric, and Simo Tao Lee. A note on the Morozov principle via Lagrange duality. Set-Valued and Variational Analysis, 26(2):265–275, June 2018. CODEN ???? ISSN 1877-0533 (print), 1877-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s11228-018-0470-y.

Burachik:2018:BTD

[BML18] Regina S. Burachik and Juan E. Martínez-Legaz. On Bregmantype distances for convex functions and maximally monotone operators. Set-Valued and Variational Analysis, 26(2):369–384, June 2018. CODEN ???? ISSN 1877-0533 (print), 1877-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s11228-017-0443-6.

Borwein:1998:CDA

[BMN98] Jonathan M. Borwein, Pierre Maréchal, and David Naugler. A convex dual approach to the computation of NMR complex spectra. Report, Department of Mathematics, Simon Fraser University, Burnaby, BC V5A 1S6, Canada, November 10, 1998. 14 pp.

URL http://docserver.carma.newcastle.edu.au/207/. Pub-

lished in [BMN00].

Borwein:2000:CDA

[BMN00] Jonathan M. Borwein, Pierre Maréchal, and David Naugler. A convex dual approach to the computation of NMR complex spectra. Math. Methods Oper. Res., 51(1):91–102, 2000. ISSN 1432-2994 (print), 1432-5217 (electronic). URL http://docserver.

carma.newcastle.edu.au/207/.

Borwein:2005:PFW

[BMP05] Jonathan M. Borwein, Mason Macklem, and Jaehyun Paek. A prototype for the federated world directory of mathematicians. In ????, editor, SIAM Conference on Computational Science and Engineering, Orlando, February 2005, page ?? Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2005.

Borwein:2002:MTC

[BMPR02] Jonathan M. Borwein, Maridee Morales, Konrad Polthier, and Jose Francisco Rodrigues, editors. Multimedia Tools for Communicating Mathematics, volume 8 of Mathematics and visualization. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 3-540-42450-4 (hardcover). LCCN

QA20.C65 M85 2002. URL http://www.loc.gov/catdir/
enhancements/fy0817/2002283245-d.html;http://www.loc.
gov/catdir/toc/fy032/2002283245.html.

Borwein:1997:SRM

[BMS97] Jonathan M. Borwein, Warren B. Moors, and Yongheng Shao. Subgradient representation of multifunctions. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, Febru-

ary 21, 1997. 14 pp. URL http://docserver.carma.newcastle.

edu.au/184/. Published in [BMS99a].

Borwein:1999:SRM

[BMS99a] Jonathan M. Borwein, Warren B. Moors, and Yongheng Shao. Subgradient representation of multifunctions. Journal of the Australian Mathematical Society. Series B, 40(3):

301–313, January 1999. CODEN JAMMDU. ISSN 0334-

2700. URL http://docserver.carma.newcastle.edu.au/184/
;http://journals.cambridge.org/action/displayAbstract?
fromPage=online&aid=3964380;https://www.cambridge.org/
core/journals/anziam-journal/article/subgradient-representation-of-multifunctions/
A035439B7CEE8FBA04EAD6EB6888CDDB. Volume in honour of B.

Craven and B. Mond.

Borwein:1999:ESB

[BMS99b] Jonathan M. Borwein, Boris S. Mordukhovich, and Yongheng Shao. On the equivalence of some basic principles in variational analysis. Journal of Mathematical Analysis and Applications, 229(1):228–257, 1999. CODEN JMANAK. ISSN 0022-247X (print), 1096-0813 (electronic). URL http://docserver.

carma.newcastle.edu.au/192/;http://www.sciencedirect.
com/science/article/pii/S0022247X98961571.

Borwein:2013:CIL

[BMS13] Jonathan Borwein, Christopher Maitland, and Matthew Skerritt. Computation of an improved lower bound to Giuga’s primality conjecture. Integers, 13:Paper No. A67, 14, 2013. CODEN INTEHN. ISSN 1553-1732. URL http://docserver.carma.

newcastle.edu.au/1511/.

Bauschke:2018:Pa

[BMST18a] Heinz H. Bauschke, Boris S. Mordukhovich, Claudia S. Sagastizábal, and Michel A. Théra. Preface. Set-Valued and Variational Analysis, 26(1):1–4, March 2018. CODEN ???? ISSN 1877-0533 (print), 1877-0541 (electronic). URL http://link.

springer.com/article/10.1007/s11228-018-0473-8;https:
//link.springer.com/content/pdf/10.1007/s11228-018-0473-8.
pdf.

Bauschke:2018:Pb

[BMST18b] Heinz H. Bauschke, Boris S. Mordukhovich, Claudia S. Sagastizábal, and Michel A. Théra. Preface. Set-Valued and Variational Analysis, 26(2):205–206, June 2018. CODEN ???? ISSN 1877-0533 (print), 1877-0541 (electronic). URL http://link.

springer.com/article/10.1007/s11228-018-0484-5;https:
//link.springer.com/content/pdf/10.1007/s11228-018-0484-5.
pdf.

Borwein:2006:BDE

[BMV06] Jonathan Borwein, Vicente Montesinos, and Jon Vanderwerff. Boundedness, differentiability and extensions of convex functions.

Journal of Convex Analysis, 13(3–4):587–602, 2006. ISSN 0944-6532 (print), 2363-6394 (electronic). URL http://docserver.

carma.newcastle.edu.au/288/;http://www.heldermann.de/
JCA/JCA13/JCA133/jca13046.htm.

Borwein:1995:LFP

[BMW95] Jonathan M. Borwein, Warren B. Moors, and Xianfu Wang. Lipschitz functions with prescribed derivatives and subderivatives. Report, Department of Mathematics & Statistics, Simon Fraser University, Burnaby, BC V5A 156, Canada, July 25, 1995. 16 pp.

URL http://docserver.carma.newcastle.edu.au/92.

Borwein:1997:LFP

[BMW97] Jonathan M. Borwein, Warren B. Moors, and Xianfu Wang. Lipschitz functions with prescribed derivatives and subderivatives. Nonlinear Analysis, Theory, Methods and Applications, 29(1):53– 63, 1997. CODEN NOANDD. ISSN 0362-546x (print), 1873-5215

(electronic). URL http://docserver.carma.newcastle.edu.
au/92/;http://www.sciencedirect.com/science/article/pii/
S0362546X96000508.

Borwein:1999:GSBa

[BMW99a] Jonathan M. Borwein, Warren B. Moors, and Xianfu Wang. Generalized subdifferentials: a Baire categorical approach. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, March 8, 1999. 19 pp. Published in [BMW01].

Borwein:1999:GSBb

[BMW99b] Jonathan M. Borwein, Warren B. Moors, and Xianfu Wang. Generalized subdifferentials: a Baire categorical approach. Technical report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, April 15, 1999. 6 pp. Published in [BMW99c].

Borwein:1999:GSBc

[BMW99c] Jonathan M. Borwein, Warren B. Moors, and Xianfu Wang. Generalized subdifferentials: a Baire categorical approach. C. R. Math. Acad. Sci. Soc. R. Can., 21(4):132–138, 1999. ISSN 0706-

1994. URL http://docserver.carma.newcastle.edu.au/234/.

Borwein:2001:GSB

[BMW01] Jonathan M. Borwein, Warren B. Moors, and Xianfu Wang. Generalized subdifferentials: a Baire categorical approach. Transactions of the American Mathematical Society, 353(10):3875–3893, 2001. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850

(electronic). URL http://docserver.carma.newcastle.edu.
au/234/.

Borwein:1984:TKN

[BN84] J. M. Borwein and J. W. Nieuwenhuis. Two kinds of normality in vector optimization. Mathematical Programming, 28(2):185– 191, 1984. CODEN MHPGA4. ISSN 0025-5610 (print), 1436-

4646 (electronic). URL http://docserver.carma.newcastle.
edu.au/1626/;http://link.springer.com/article/10.1007/
BF02612358.

Borwein:1986:PSS

[BN86] Jon Borwein and William A. Newcomb. Problems and solutions: Solutions of advanced problems: 6491. American Mathematical Monthly, 93(10):822–823, 1986. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Borwein:1994:SOD

[BN94] Jonathan M. Borwein and Dominikus Noll. Second order differentiability of convex functions in Banach spaces. Transactions of the American Mathematical Society, 342(1):43–81, March 1994. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850

(electronic). URL http://docserver.carma.newcastle.edu.
au/1549/;https://www.jstor.org/stable/2154684.

Bauschke:1999:EAD

[BNCB99] Heinz H. Bauschke, Dominikus Noll, Anna Celler, and Jonathan M. Borwein. An EM algorithm for dynamic SPECT. IEEE Transactions on Medical Imaging, 18(3):252–261, March 1999. CO-DEN ITMID4. ISSN 0278-0062 (print), 1558-254x (electronic).

URL http://docserver.carma.newcastle.edu.au/187;http:
//www.ncbi.nlm.nih.gov/pubmed/10363703.

Borwein:2010:RWP

[BNSW10] Jonathan M. Borwein, Dirk Nuyens, Armin Straub, and James Wan. Random walks in the plane. In 22nd International Conference on Formal Power Series and Algebraic Combinatorics (FP-SAC 2010), Discrete Math. Theor. Comput. Sci. Proc., AN, pages 191–202. Assoc. Discrete Math. Theor. Comput. Sci., Nancy, France, 2010.

Borwein:2011:SAP

[BNSW11] Jonathan M. Borwein, Dirk Nuyens, Armin Straub, and James Wan. Some arithmetic properties of short random walk integrals.

The Ramanujan Journal, 26(1):109–132, 2011. CODEN RAJOF9. ISSN 1382-4090 (print), 1572-9303 (electronic). URL http://

docserver.carma.newcastle.edu.au/1440/.

Borwein:1976:TCC

[BO76] J. Borwein and R. O’Brien. Tangent cones and convexity. Canadian mathematical bulletin = Bulletin canadien de mathématiques, 19(3):257–261, 1976. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL http:

//docserver.carma.newcastle.edu.au/1517/.

Borwein:1978:CCC

[BO78] J. M. Borwein and R. C. O’Brien. Cancellation characterizes convexity. Nanta Mathematica, 11(1):100–102, 1978. CODEN NAMAB2. ISSN 0077-2739. URL http://docserver.carma.

newcastle.edu.au/1659/.

Bingham:2011:HKB

[BO11a] N. H. Bingham and A. J. Ostaszewski. Homotopy and the Kestelman–Borwein–Ditor theorem. Canadian mathematical bulletin = Bulletin canadien de mathématiques, 54(1):12–20, 2011. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:2011:LHM

[BO11b] Jonathan M. Borwein and Judy-Anne Osborn. Loving and hating mathematics, by Reuben Hersh and Vera John-Steiner. Princeton and Oxford: Princeton University Press, 2010, 428 pp., US$29.95. ISBN-10: 0-691-14247-5, ISBN-13: 978-0-691-14247-0 [book review of Cno. 2731560]. The Mathematical Intelligencer, 33(4): 63–69, December 2011. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic). URL http://link.springer.

com/article/10.1007/s00283-011-9260-1.

Borwein:1972:MON

[Bor72] Jonathan Michael Borwein. Monotone operators and non-linear functional analysis. M.Sc. thesis, Oxford University, Oxford, UK, 1972.

Borwein:1974:ORP

[Bor74] Jonathan Michael Borwein. Optimization with respect to partial orderings. D.Phil. thesis, Oxford University, Oxford, UK,

1974. 200 pp. URL http://search.proquest.com/pqdtglobal/
docview/1767375969/.

Borwein:1976:FPD

[Bor76a] J. M. Borwein. Fractional programming without differentiability. Mathematical Programming, 11(1):283–290, 1976. CODEN MH-PGA4. ISSN 0025-5610 (print), 1436-4646 (electronic). URL

http://docserver.carma.newcastle.edu.au/1671/;http://
link.springer.com/article/10.1007/BF01580396.

Borwein:1976:NFJ

[Bor76b] J. M. Borwein. A note on Fritz John sufficiency. Bulletin of the Australian Mathematical Society, 15(2):293–296, October 1976. CODEN ALNBAB. ISSN 0004-9727 (print), 1755-1633 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
1515/;http://journals.cambridge.org/action/displayAbstract?
fromPage=online&aid=4812456.

Borwein:1977:MCO

[Bor77a] J. Borwein. Multivalued convexity and optimization: a unified approach to inequality and equality constraints. Mathematical Programming, 13(2):183–199, 1977. CODEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic). URL http://docserver.

carma.newcastle.edu.au/1669/;http://link.springer.com/
article/10.1007/BF01584336.

Borwein:1977:PEH

[Bor77b] J. Borwein. A proof of the equivalence of Helly’s and Krasnoselski’s theorems. Canadian mathematical bulletin = Bulletin canadien de mathématiques, 20(1):35–37, March 1977. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL

http://docserver.carma.newcastle.edu.au/1518/.

Borwein:1977:PEP

[Bor77c] J. Borwein. Proper efficient points for maximizations with respect to cones. SIAM Journal on Control and Optimization, 15 (1):57–63, January 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http://docserver.carma.

newcastle.edu.au/1667/.

Borwein:1977:WLP

[Bor77d] J. M. Borwein. The William Lowell Putnam competition. A brief history. Unpublished notes., January 26, 1977.

Borwein:1978:WTC

[Bor78a] J. Borwein. Weak tangent cones and optimization in a Banach space. SIAM Journal on Control and Optimization, 16(3):512– 522, May 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-

7138 (electronic). URL http://docserver.carma.newcastle.
edu.au/1664/.

Borwein:1978:SEF

[Bor78b] J. M. Borwein. On strongly exposing functionals. Proceedings of the American Mathematical Society, 69(1):46–48, 1978. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
1665/;https://www.jstor.org/stable/2043185.

Borwein:1978:TCS

[Bor78c] J. M. Borwein. Tangent cones, starshape and convexity. International Journal of Mathematics and Mathematical Sciences, 1(4): 459–477, 1978. ISSN 0161-1712 (print), 1687-0425 (electronic).

URL http://docserver.carma.newcastle.edu.au/1663/.

Borwein:1979:DTS

[Bor79a] J. M. Borwein. Direct theorems in semi-infinite convex programming. Report 11, Department of Mathematics, Dalhousie University, Halifax, NS, Canada, April 1979. 33 pp.

Borwein:1979:LLC

[Bor79b] J. M. Borwein. The limiting lagrangean as a consequence of Helly’s Theorem. Report 7, Department of Mathematics, Dalhousie University, Halifax, NS, Canada, March 1979.

Borwein:1979:MFP

[Bor79c] J. M. Borwein. The minimum of a family of programs. In Third Symposium on Operations Research (Univ. Mannheim, Mannheim, 1978), Section I, volume 31 of Operations Res. Verfahren, pages 99–111. Hain, 1979.

Borwein:1979:MAF

[Bor79d] J. M. Borwein. A multivalued approach to the Farkas lemma. Mathematical Programming Study, 10:42–47, 1979. CO-DEN MPSTDF. ISSN 0303-3929. URL http://docserver.

carma.newcastle.edu.au/1660/;http://link.springer.com/
chapter/10.1007/BFb0120842. Point-to-set maps and mathe-

matical programming.

Borwein:1979:CD

[Bor79e] J. M. Borwein. On convex decompositions. Accepted for publication in Nanta Mathematica, but the journal ceased publication in 1979 before the paper appeared., 1979. URL http:

//docserver.carma.newcastle.edu.au/1672/.

Borwein:1979:TVS

[Bor79f] J. M. Borwein. Two variants of Sylvester’s theorem. Atlantic Mathematics Bulletin, 3(??):11–13, ???? 1979. ISSN 0705-9078.

URL http://docserver.carma.newcastle.edu.au/1658/.

Borwein:1979:WLS

[Bor79g] J. M. Borwein. Weak local supportability and applications to approximation. Pacific Journal of Mathematics, 82(2):323– 338, 1979. CODEN PJMAAI. ISSN 0030-8730 (print), 1945-

5844 (electronic). URL http://projecteuclid.org/euclid.
pjm/1102784876.

Borwein:1979:MLP

[Bor79h] Jonathan M. Borwein. Monochrome lines in the plane. Mathematics Magazine, 52(1):41–45, 1979. CODEN MAMGA8. ISSN 0025-570X (print), 1930-0980 (electronic). URL http://docserver.

carma.newcastle.edu.au/1662/.

Borwein:1980:GPE

[Bor80a] J. M. Borwein. The geometry of Pareto efficiency over cones. Optimization, 11(2):235–248, 1980. CODEN OPTZDQ. ISSN 0323-

3898. URL http://www.tandfonline.com/doi/abs/10.1080/
02331938008842650.

Borwein:1980:LMT

[Bor80b] J. M. Borwein. A Lagrange multiplier theorem and a sandwich theorem for convex relations. Research Report 80-1, Department of Mathematics, Dalhousie University, Halifax, NS, Canada, January 15, 1980. 26 pp.

Borwein:1980:LM

[Bor80c]  J. M. Borwein.  Lexicographic multipliers.  Journal of

Mathematical Analysis and Applications, 78(1):309–327, 1980. CODEN JMANAK. ISSN 0022-247X (print), 1096-0813

(electronic). URL http://docserver.carma.newcastle.edu.
au/1653/;http://www.sciencedirect.com/science/article/
pii/0022247X80902280.

Borwein:1980:NPD

[Bor80d] J. M. Borwein. A note on perfect duality and limiting Lagrangeans. Mathematical Programming, 18(1):330–337, 1980. CODEN MHPGA4. ISSN 0025-5610 (print), 1436-

4646 (electronic). URL http://docserver.carma.newcastle.
edu.au/1654/;http://link.springer.com/article/10.1007/
BF01588327.

Borwein:1980:SDT

[Bor80e] J. M. Borwein. A strong duality theorem for the minimum of a family of convex programs. Journal of Optimization Theory and Applications, 31(4):453–472, 1980. CODEN JOTABN. ISSN 0022-3239 (print), 1573-2878 (electronic). URL http:

//docserver.carma.newcastle.edu.au/1657/;http://link.
springer.com/article/10.1007/BF00934472.

Borwein:1981:SRP

[Bor81a] J. Borwein. Some remarks on a paper of Ş. Cobzaş on antiproximal sets: “Antiproximinal sets in some Banach spaces” [Math. Balkanica 4 (1974), 79–82; MR 51 #13553]. Bulletin of the Calcutta Mathematical Society, 73(1):5–8, 1981. CODEN BCMSA5. ISSN

0008-0659. URL http://docserver.carma.newcastle.edu.au/
1648/.

Borwein:1981:CRO

[Bor81b] J. M. Borwein. Convex relations in optimization and analysis. In Generalized Convexity in Optimization and Economics, pages 335–377. Academic Press, New York, NY, USA, 1981.

Borwein:1981:DTS

[Bor81c] J. M. Borwein. Direct theorems in semi-infinite convex programming. Mathematical Programming, 21(1):301–318, 1981. CODEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic). URL

http://link.springer.com/article/10.1007/BF01584251.

Borwein:1981:LMT

[Bor81d] J. M. Borwein. A Lagrange multiplier theorem and a sandwich theorem for convex relations. Mathematica Scandinavica, 48(2): 189–204, 1981. CODEN MTSCAN. ISSN 0025-5521 (print), 1903-

1807 (electronic). URL http://docserver.carma.newcastle.
edu.au/1646/;http://www.mscand.dk/article/view/11911/
9927.

Borwein:1981:LLC

[Bor81e] J. M. Borwein. The limiting Lagrangian as a consequence of Helly’s theorem. Journal of Optimization Theory and Applications, 33(4):497–513, 1981. CODEN JOTABN. ISSN 0022-3239 (print), 1573-2878 (electronic). URL http://docserver.

carma.newcastle.edu.au/1647/;http://link.springer.com/
article/10.1007/BF00935755.

Borwein:1982:CDP

[Bor82a] J. M. Borwein. Continuity and differentiability properties of convex operators. Proceedings of the London Mathematical Society. Third Series, 44(3):420–444, 1982. CODEN PLMTAL. ISSN 0024-6115 (print), 1460-244x (electronic). URL http:

//docserver.carma.newcastle.edu.au/1640/.

Borwein:1982:NSC

[Bor82b] J. M. Borwein. Necessary and sufficient conditions for quadratic minimality. Numerical Functional Analysis and Optimization, 5(2):127–140, 1982. CODEN NFADOL. ISSN 0163-0563 (print), 1532-2467 (electronic). URL http://docserver.carma.

newcastle.edu.au/1639/.

Borwein:1982:NSM

[Bor82c] J. M. Borwein. A note on ε-subgradients and maximal monotonicity. Pacific Journal of Mathematics, 103(2):307–314, 1982. CO-DEN PJMAAI. ISSN 0030-8730 (print), 1945-5844 (electronic).

URL http://projecteuclid.org/euclid.pjm/1102723964.

Borwein:1982:NES

[Bor82d] J. M. Borwein. A note on the existence of subgradients. Mathematical Programming, 24(2):225–228, 1982. CODEN MH-PGA4. ISSN 0025-5610 (print), 1436-4646 (electronic). URL

http://docserver.carma.newcastle.edu.au/1641/;http://
link.springer.com/article/10.1007/BF01585105.

Borwein:1982:HBE

[Bor82e] Jonathan M. Borwein. On the Hahn–Banach extension property. Proceedings of the American Mathematical Society, 86(1):42–46, 1982. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic).

Borwein:1983:APD

[Bor83a] J. M. Borwein. Adjoint process duality. Mathematics of Operations Research, 8(3):403–434, 1983. CODEN MOREDQ. ISSN 0364-765x (print), 1526-5471 (electronic). URL http://

docserver.carma.newcastle.edu.au/1633/. See letter [Zăl86].

Borwein:1983:CCP

[Bor83b] J. M. Borwein. Completeness and the contraction principle. Proceedings of the American Mathematical Society, 87(2):246–250, 1983. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826

(electronic). URL http://docserver.carma.newcastle.edu.
au/1637/.

Borwein:1983:HSS

[Bor83c] J. M. Borwein. How special is semi-infinite programming? In Semi-infinite Programming, pages 139–172. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983.

Borwein:1983:NFL

[Bor83d] J. M. Borwein. A note on the Farkas lemma. Utilitas Mathematica, 24:235–241, 1983. CODEN UTMADA. ISSN 0315-3681. URL

http://docserver.carma.newcastle.edu.au/1631/.

Borwein:1983:EPE

[Bor83e] Jonathan M. Borwein. On the existence of Pareto efficient points. Mathematics of Operations Research, 8(1):64–73, February 1983. CODEN MOREDQ. ISSN 0364-765x (print), 1526-

5471 (electronic). URL http://docserver.carma.newcastle.
edu.au/1636/;https://www.jstor.org/stable/3689411.

Borwein:1983:SIP

[Bor83f] Jonathan M. Borwein. Semi-infinite programming duality: how special is it? In Anthony V. Fiacco and Kenneth O. Kortanek, editors, Semi-Infinite Programming and Applications: an International Symposium, Austin, Texas, September 8 -10, 1981, volume 215 of Lecture Notes in Economics and Mathematical Systems, pages 10–36. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. ISBN 0-387-12304-0 (paperback: New York), 3-540-12304-0 (paperback: Berlin). LCCN QA402.5 .S428 1983.

Borwein:1984:GLC

[Bor84a] J. M. Borwein. Generalized linear complementarity problems treated without fixed-point theory. Journal of Optimization Theory and Applications, 43(3):343–356, 1984. CODEN JOTABN. ISSN 0022-3239 (print), 1573-2878 (electronic). URL http:

//link.springer.com/article/10.1007/BF00934459.

Borwein:1984:ICE

[Bor84b] J. M. Borwein. An integral characterization of Euclidean space. Bulletin of the Australian Mathematical Society, 29(3):357–364, 1984. CODEN ALNBAB. ISSN 0004-9727 (print), 1755-1633

(electronic). URL http://docserver.carma.newcastle.edu.
au/1623/.

Borwein:1984:EEH

[Bor84c] J. M. Borwein. On the existence of eigenvalues for Hermitian matrices. In ????, editor, APICS Proceedings, pages 1–5. ????, ????, 1984.

Borwein:1984:SRP

[Bor84d] J. M. Borwein. Stability and regular points of inequality systems. Report DALTR 84-02, Department of Mathematics, Dalhousie University and Carnegie-Mellon University, Halifax, NS, Canada and Pittsburgh, PA, USA, 1984. 67 pp.

Borwein:1984:SCO

[Bor84e] J. M. Borwein. Subgradients of convex operators. Optimization, 15(2):179–191, 1984. CODEN OPTZDQ. ISSN 0323-3898.

Borwein:1985:PSS

[Bor85a] J. Borwein. Problems and solutions: Solutions of advanced problems: 6430. American Mathematical Monthly, 92(2):148–149, February 1985. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BSZ^{+}83].

Borwein:1985:SMI

[Bor85b] Jon Borwein. Some modular identities of Ramanujan useful in approximating π. Proceedings of the American Mathematical Society, 95(3):365–371, November 1985. CODEN PAM-YAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL

http://docserver.carma.newcastle.edu.au/1616/;https://
www.jstor.org/stable/2045804.

Borwein:1985:ATG

[Bor85c] Jonathan M. Borwein. Alternative theorems for general complementarity problems. In Infinite programming (Cambridge, 1984), volume 259 of Lecture Notes in Econom. and Math. Systems, pages 194–203. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985.

Borwein:1986:NDC

[Bor86a] J. M. Borwein. Norm duality for convex processes and applications. Journal of Optimization Theory and Applications, 48(1): 53–64, 1986. CODEN JOTABN. ISSN 0022-3239 (print), 1573-

2878 (electronic). URL http://docserver.carma.newcastle.
edu.au/1613/;http://link.springer.com/article/10.1007/

BF00938589. Fourth symposium on nonlinear programming with data perturbations.

Borwein:1986:PMO

[Bor86b] J. M. Borwein. Partially monotone operators and the generic differentiability of convex-concave and biconvex mappings. Israel Journal of Mathematics, 54(1):42–50, 1986. CODEN ISJMAP. ISSN 0021-2172 (print), 1565-8511 (electronic). URL http:

//docserver.carma.newcastle.edu.au/1605/;http://link.
springer.com/article/10.1007/BF02764875.

Borwein:1986:SRP

[Bor86c] J. M. Borwein. Stability and regular points of inequality systems. Journal of Optimization Theory and Applications, 48(1): 9–52, 1986. CODEN JOTABN. ISSN 0022-3239 (print), 1573-

2878 (electronic). URL http://docserver.carma.newcastle.
edu.au/1612/;http://link.springer.com/article/10.1007/

BF00938588. Fourth symposium on nonlinear programming with data perturbations.

Borwein:1986:CMN

[Bor86d] Jonathan M. Borwein. Cones, minimality notions and consequences. International conference on Vector Optimization, Darmstadt, West Germany., August 5, 1986.

Borwein:1986:GDO

[Bor86e] Jonathan M. Borwein. Generic differentiability of order-bounded convex operators. Journal of the Australian Mathematical Society. Series B, 28(1):22–29, July 1986. CODEN JAM-

MDU. ISSN 0334-2700. URL http://docserver.carma.
newcastle.edu.au/1607/;http://journals.cambridge.org/
action/displayAbstract?fromPage=online&aid=3973752;https:
//www.cambridge.org/core/journals/anziam-journal/article/
generic-differentiability-of-orderbounded-convex-oparators/
EBC10922AB270DAAF5784308C7E3157A.

Borwein:1986:RME

[Bor86f] Jonathan M. Borwein. Ramanujan, modular equations and pi. Colloquium, St. Francis Xavier University, Antigonish, NS, Canada., November 24, 1986. URL http://docserver.carma.

newcastle.edu.au/1587/.

Borwein:1986:SVP

[Bor86g] Jonathan M. Borwein. A smooth variational principle. Computer Science/Optimization Seminar, University of Waterloo, Waterloo, ON, Canada., 1986.

Borwein:1987:ACO

[Bor87a] J. M. Borwein. Automatic continuity and openness of convex relations. Proceedings of the American Mathematical Society, 99(1):49–55, January 1987. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL http:

//docserver.carma.newcastle.edu.au/1604/;https://www.
jstor.org/stable/2046269.

Borwein:1987:CMN

[Bor87b] J. M. Borwein. Cones, minimality notions and consequences. In Recent Advances and Historic Development of Vector Optimization, pages 62–85. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987.

Borwein:1987:CCM

[Bor87c] J. M. Borwein. Convex cones, minimality notions, and consequences. In Recent Advances and Historical Development of Vector Optimization, page ?? Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1987. URL http://

link.springer.com/chapter/10.1007/978-3-642-46618-2_3.

Borwein:1987:AGM

[Bor87d] Jonathan M. Borwein. The arithmetic–geometric mean of Gauss and Legendre: An excursion. Canadian Mathematical Society, Coxeter–James Lecture, Vancouver, BC, Canada., December 15, 1987.

Borwein:1987:OC

[Bor87e] Jonathan M. Borwein. Order complementarity. Colloquium, University of Western Ontario, London, ON, Canada., March 24, 1987.

Borwein:1987:RMEb

[Bor87f] Jonathan M. Borwein. Ramanujan, modular equations and approximations to pi. Ramanujan Centenary Meeting, University of Illinois, Champaign-Urbana, IL, USA., June 1, 1987. URL

http://docserver.carma.newcastle.edu.au/1587/;https://
web.archive.org/web/20170328083101/h.

Borwein:1987:RMEa

[Bor87g] Jonathan M. Borwein. Ramanujan, modular equations and pi. Colloquium, Concordia., March 19, 1987. URL http:

//docserver.carma.newcastle.edu.au/1587/.

Borwein:1987:SVPb

[Bor87h] Jonathan M. Borwein. A smooth variational principle. AMS Winter meetings, San Antonio, TX, USA., 1987.

Borwein:1987:SVPc

[Bor87i] Jonathan M. Borwein. A smooth variational principle. Analysis/Applied Math seminar, University of Toronto, Toronto, ON, Canada., March 26, 1987.

Borwein:1987:SVPd

[Bor87j] Jonathan M. Borwein. A smooth variational principle. Franco– Quebec Conference on Non-linear Analysis, Perpignan, France., June 25, 1987.

Borwein:1987:SAC

[Bor87k] Jonathan M. Borwein. Spectral analysis via convex programming. Charnes’ 70th birthday conference, IC2, University of Texas at Austin, Austin, TX, USA., October 15, 1987.

Borwein:1987:TIR

[Bor87l] Jonathan M. Borwein. A theta identity of Ramanujan’s and applications. AMS Summer Research Institute, Bowdoin College, Brunswick, ME, USA., July 21, 1987.

Borwein:1987:ELL

[Bor87m] Jonathan Michael Borwein. Epi-Lipschitz-like sets in Banach space: theorems and examples. Nonlinear Analysis, Theory, Methods and Applications, 11(10):1207–1217, 1987. CODEN NOANDD. ISSN 0362-546x (print), 1873-5215

(electronic). URL http://docserver.carma.newcastle.edu.
au/1599/;http://www.sciencedirect.com/science/article/
pii/0362546X87900083.

Borwein:1988:AGMa

[Bor88a] Jonathan M. Borwein. The arithmetic–geometric mean of Gauss and Legendre: An excursion. Distinguished Lecturer Series, University of Delaware, Newark, DE, USA., May 13, 1988.

Borwein:1988:AGMb

[Bor88b] Jonathan M. Borwein. The arithmetic–geometric mean of Gauss and Legendre: An excursion. Colloquium, University of Newcastle, Newcastle, NSW, Australia., June 14, 1988.

Borwein:1988:AGMc

[Bor88c] Jonathan M. Borwein. The arithmetic–geometric mean of Gauss and Legendre: An excursion. Colloquium, University of New England, Armidale, NSW, Australia., June 27, 1988.

Borwein:1988:AGMd

[Bor88d] Jonathan M. Borwein. The arithmetic–geometric mean of Gauss and Legendre: An excursion. Colloquium, Auckland University, Auckland, New Zealand., July 27, 1988.

Borwein:1988:AGMe

[Bor88e] Jonathan M. Borwein. The arithmetic–geometric mean of Gauss and Legendre: An excursion. Colloquium, Macquarie University, Sydney, NSW, Australia., September 12, 1988.

Borwein:1988:BFD

[Bor88f] Jonathan M. Borwein. Borchardt’s four-dimensional arithmetic– geometric mean. Seminar, Macquarie University, Sydney, NSW, Australia., September 14, 1988.

Borwein:1988:ETEa

[Bor88g] Jonathan M. Borwein. Ekeland’s theorem and its extensions. Distinguished Lecturer Series, University of Delaware, Newark, DE, USA., May 12, 1988.

Borwein:1988:ETEb

[Bor88h] Jonathan M. Borwein. Ekeland’s theorem and its extensions. Colloquium, University of New England, Armidale, NSW, Australia., June 29, 1988.

Borwein:1988:ETEc

[Bor88i] Jonathan M. Borwein. Ekeland’s theorem and its extensions. Colloquium, Melbourne University, Melbourne, VIC, Australia., August 1, 1988.

Borwein:1988:MCK

[Bor88j] Jonathan M. Borwein. Mosco convergence and the Kadec property. Workshop on Functional Analysis and Optimization, Australian National University, Canberra, ACT, Australia., August

24, 1988. URL http://docserver.carma.newcastle.edu.au/

Borwein:1988:OPE

[Bor88k] Jonathan M. Borwein. Open problems on the existence of nearest points. Workshop on Functional Analysis and Optimization, Australian National University, Canberra, ACT, Australia., August 9, 1988.

Borwein:1988:PFC

[Bor88l] Jonathan M. Borwein. Partially-finite convex programming. AMS Winter Meetings, Atlanta, GA, USA., January 6, 1988.

Borwein:1988:STAa

[Bor88m] Jonathan M. Borwein. Subderivatives and their applications. Conference on Functional Analysis and Optimization, Australian National University, Canberra, ACT, Australia., August 17, 1988.

Borwein:1988:STAb

[Bor88n] Jonathan M. Borwein. Subderivatives and their applications. Joint Colloquium, University of New South Wales and Sydney University, Sydney, NSW, Australia., September 9, 1988.

Borwein:1989:AFM

[Bor89a] Jonathan M. Borwein. APICS/FRASER medal presentation talk. Dalhousie University, Halifax, NS, Canada., May 12, 1989.

Borwein:1989:CPH

[Bor89b] Jonathan M. Borwein. The calculation of pi. How, why, what? Nova Scotia Institute of Science, Halifax, NS, Canada., February 8, 1989.

Borwein:1989:MCP

[Bor89c] Jonathan M. Borwein. Minimal CUSCOS and Preisses’ theorem. Miniconference on Optimization Theory, University of Pau, France., June 12, 1989.

Borwein:1989:MCT

[Bor89d] Jonathan M. Borwein. Minimal CUSCOS and their applications. Plenary talk, Conference on Fixed Point Theory, CIRM, Marseille, France., June 9, 1989.

Borwein:1989:PAG

[Bor89e] Jonathan M. Borwein. Pi and the arithmetic–geometric mean. Colloquium, Rutger’s University, New Brunswick, NJ, USA., April 14, 1989.

Borwein:1989:PER

[Bor89f] Jonathan M. Borwein. Pi, Euler, Ramanujan, and MAPLE. Colloquium, Department of Computer Science, University of Manitoba, Winnipeg, MB, Canada., November 9, 1989.

Borwein:1989:QMIa

[Bor89g] Jonathan M. Borwein. Quadratic mean iterations. Carleton University/Université d’Ottawa joint Colloquium, Carleton University, Ottawa, ON, Canada., March 4, 1989.

Borwein:1989:QMIb

[Bor89h] Jonathan M. Borwein. Quadratic mean iterations. Seminar, Rutger’s University, New Brunswick, NJ, USA., April 12, 1989.

Borwein:1989:SFC

[Bor89i] Jonathan M. Borwein. Semi-finite convex programming. ORSA/ TIMS National Meeting, New York (presented by A. Lewis)., October 17, 1989.

Borwein:1990:SDPa

[Bor90a] J. M. Borwein. A survey of differentiability properties of convex, Lipschitz and lsc functions. In Gustave Choquet et al., editors, Séminaire d’initiation à l’analyze, page ?? Publ. Math. Univ. Pierre et Marie Curie, Paris, France, 1990. URL https://

carma.newcastle.edu.au/jon/Preprints/Books/CUP/CUPold/
convsurvey.pdf.

Borwein:1990:BRG

[Bor90b] Jonathan Borwein. Book review: Generalized Concavity (Mordecai Avriel, Walter E. Diewert, Siegfried Schaible, and Israel Zang). SIAM Review, 32(4):689–690, December 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Borwein:1990:CPCa

[Bor90c] Jonathan M. Borwein. Convex programming and the choice of entropy in spectral estimation. Seminar, Department of Combinatorics and Optimization, University of Waterloo, Waterloo, ON, Canada., November 21, 1990.

Borwein:1990:CPCb

[Bor90d] Jonathan M. Borwein. Convex programming and the choice of entropy in spectral estimation. Seminar, Department of Combinatorics and Optimization, Waterloo, Waterloo, ON, Canada., November 21, 1990.

Borwein:1990:CPAa

[Bor90e] Jonathan M. Borwein. Convex programming approaches to moment, curve, and signal estimation. Miniconference on Optimization Theory, Dalhousie University, Halifax, NS, Canada., August 22, 1990.

Borwein:1990:CPAb

[Bor90f] Jonathan M. Borwein. Convex programming approaches to moment, curve, and signal estimation. Miniconference on Optimization Theory, Dalhousie University, Halifax, NS, Canada., August 22, 1990.

Borwein:1990:DPCa

[Bor90g] Jonathan M. Borwein. Differentiability properties of convex, Lipschitz and semicontinuous functions. Ontario Math Meetings #88, Brock University, St. Catharines, ON, Canada., April 21, 1990.

Borwein:1990:DPCb

[Bor90h] Jonathan M. Borwein. Differentiability properties of convex, Lipschitz and semicontinuous functions. Ontario Math Meetings #88, Brock University, St. Catharines, ON, Canada., April 21, 1990.

Borwein:1990:DPLa

[Bor90i] Jonathan M. Borwein. Differentiability properties of Lipschitz functions. Nonlinear Analysis Seminar #1, Technion, Haifa, Israel., May 15, 1990.

Borwein:1990:DPLc

[Bor90j] Jonathan M. Borwein. Differentiability properties of Lipschitz functions. Nonlinear Analysis Seminar #1, Technion, Haifa, Israel., May 15, 1990.

Borwein:1990:DPLb

[Bor90k] Jonathan M. Borwein. Differentiability properties of lower semicontinuous functions. Nonlinear Analysis Seminar #2, Technion, Haifa, Israel., May 21, 1990.

Borwein:1990:DPLd

[Bor90l] Jonathan M. Borwein. Differentiability properties of lower semicontinuous functions. Nonlinear Analysis Seminar #2, Technion, Haifa, Israel., May 21, 1990.

Borwein:1990:ETSa

[Bor90m] Jonathan M. Borwein. Ekeland’s theorem and the smooth variational principle. Conference on Topological Methods, Brock University, St. Catharines, ON, Canada., April 20, 1990.

Borwein:1990:ETSb

[Bor90n] Jonathan M. Borwein. Ekeland’s theorem and the smooth variational principle. Conference on Topological Methods, Brock University, St. Catharines, ON, Canada., April 20, 1990.

Borwein:1990:GMSa

[Bor90o] Jonathan M. Borwein. Greek mathematics and the story of the circle. Junior High presentation, Dalhousie University, Halifax, NS, Canada., December 6, 1990.

Borwein:1990:GMSb

[Bor90p] Jonathan M. Borwein. Greek mathematics and the story of the circle. Junior High presentation, Dalhousie University, Halifax, NS, Canada., December 6, 1990.

Borwein:1990:HCPa

[Bor90q] Jonathan M. Borwein. The history of the computation of pi. APICS Lecture, University of Prince Edward Island, Charlottetown, PE C1A 4P3, Canada., March 16, 1990.

Borwein:1990:HCPb

[Bor90r] Jonathan M. Borwein. The history of the computation of pi. APICS Lecture, St. Francis Xavier University, Antigonish, NS, Canada., March 24, 1990.

Borwein:1990:HCPc

[Bor90s] Jonathan M. Borwein. The history of the computation of pi. APICS Lecture, Memorial University, St John’s, NL, Canada., March 31, 1990.

Borwein:1990:HCPd

[Bor90t] Jonathan M. Borwein. The history of the computation of pi. APICS Lecture, Université de Moncton, Moncton, NB, Canada., April 5, 1990.

Borwein:1990:HCPe

[Bor90u] Jonathan M. Borwein. The history of the computation of pi. APICS Lecture, University of Prince Edward Island, Charlottetown, PE C1A 4P3, Canada., March 16, 1990.

Borwein:1990:HCPf

[Bor90v] Jonathan M. Borwein. The history of the computation of pi. APICS Lecture, St. Francis Xavier University, Antigonish, NS, Canada., March 24, 1990.

Borwein:1990:HCPg

[Bor90w] Jonathan M. Borwein. The history of the computation of pi. APICS Lecture, Memorial University, St John’s, NL, Canada., March 31, 1990.

Borwein:1990:HCPh

[Bor90x] Jonathan M. Borwein. The history of the computation of pi. APICS Lecture, Université de Moncton, Moncton, NB, Canada., April 5, 1990.

Borwein:1990:MCAa

[Bor90y] Jonathan M. Borwein. Minimal CUSCOS and applications to Lipschitz functions. AMS Winter Meetings, Louisville, KY, USA., January 19, 1990.

Borwein:1990:MCAb

[Bor90z] Jonathan M. Borwein. Minimal CUSCOS and applications to Lipschitz functions. Nonlinear Analysis Seminar #3, Technion, Haifa, Israel., May 25, 1990.

Borwein:1990:MCAc

[Bor90-27] Jonathan M. Borwein. Minimal CUSCOS and applications to Lipschitz functions. AMS Winter Meetings, Louisville, KY, USA., January 19, 1990.

Borwein:1990:MCAd

[Bor90-28] Jonathan M. Borwein. Minimal CUSCOS and applications to Lipschitz functions. Nonlinear Analysis Seminar #3, Technion, Haifa, Israel., May 25, 1990.

Borwein:1990:PERa

[Bor90-29] Jonathan M. Borwein. Pi, Euler, Ramanujan, and MAPLE. APICS Lecture, Mount St Vincent University, Halifax, NS, Canada., January 22, 1990.

Borwein:1990:PERb

[Bor90-30] Jonathan M. Borwein. Pi, Euler, Ramanujan, and MAPLE. APICS Lecture, Mount Allison University, Sackville, NB, Canada., January 26, 1990.

Borwein:1990:PERc

[Bor90-31] Jonathan M. Borwein. Pi, Euler, Ramanujan, and MAPLE. APICS Lecture, University College of Cape Breton, Sydney, NS B1P 6L2, Canada., March 15, 1990.

Borwein:1990:PERe

[Bor90-32] Jonathan M. Borwein. Pi, Euler, Ramanujan, and MAPLE. APICS Lecture, University of New Brunswick, Moncton, NB, Canada., April 6, 1990.

Borwein:1990:PERf

[Bor90-33] Jonathan M. Borwein. Pi, Euler, Ramanujan, and MAPLE. Seminar, Technion, Haifa, Israel., June 15, 1990.

Borwein:1990:PERg

[Bor90-34] Jonathan M. Borwein. Pi, Euler, Ramanujan, and MAPLE. APICS Lecture, Mount St Vincent University, Halifax, NS, Canada., January 22, 1990.

Borwein:1990:PERh

[Bor90-35] Jonathan M. Borwein. Pi, Euler, Ramanujan, and MAPLE. APICS Lecture, Mount Allison University, Sackville, NB, Canada., January 26, 1990.

Borwein:1990:PERi

[Bor90-36] Jonathan M. Borwein. Pi, Euler, Ramanujan, and MAPLE. APICS Lecture, University College of Cape Breton, Sydney, NS B1P 6L2, Canada., March 15, 1990.

Borwein:1990:PERj

[Bor90-37] Jonathan M. Borwein. Pi, Euler, Ramanujan, and MAPLE. APICS Lecture, Acadia University, Wolfville, NS B4P 2R6, Canada., March 23, 1990.

Borwein:1990:PERk

[Bor90-38] Jonathan M. Borwein. Pi, Euler, Ramanujan, and MAPLE. APICS Lecture, University of New Brunswick, Moncton, NB, Canada., April 6, 1990.

Borwein:1990:PERl

[Bor90-39] Jonathan M. Borwein. Pi, Euler, Ramanujan, and MAPLE. Seminar, Technion, Haifa, Israel., June 15, 1990.

Borwein:1990:SDPb

[Bor90-40] Jonathan M. Borwein. A survey of differentiability properties of convex, Lipschitz and semicontinuous functions. Seminar, Technion Israel., May 24, 1990.

Borwein:1990:SDPc

[Bor90-41] Jonathan M. Borwein. A survey of differentiability properties of convex, Lipschitz and semicontinuous functions. Seminar, Ben Gurion University, Israel., May 28, 1990.

Borwein:1990:SDPd

[Bor90-42] Jonathan M. Borwein. A survey of differentiability properties of convex, Lipschitz and semicontinuous functions. Seminar, Technion, Haifa, Israel., May 24, 1990.

Borwein:1990:SDPe

[Bor90-43] Jonathan M. Borwein. A survey of differentiability properties of convex, Lipschitz and semicontinuous functions. Seminar, Ben Gurion University, Israel., May 28, 1990.

Borwein:1991:MCS

[Bor91a] J. M. Borwein. Minimal CUSCOS and subgradients of Lipschitz functions. In Michael A. Théra and Jean-Bernard Baillon, editors, Fixed point theory and applications: proceedings of the International Conference on Fixed Point Theory and Applications, held at CIRM (Centre International de Rencontres Mathématiques) located in the campus of the University of Marseille-Luminy, June 5–9, 1989, volume 252 of Pitman Res. Notes Math. Ser., pages 57– 81. Longman Scientific and Technical, Harlow, Essex, UK, 1991. ISBN 0-582-08063-0 (Harlow), 0-470-21759-6 (New York). LCCN QA329.9 .F58 1991.

Borwein:1991:CPCa

[Bor91b] Jonathan M. Borwein. Convex programming and the choice of entropy in spectral estimation. First Plenary talk, Journees d’Optimization, Université de Limoges, Limoges, France., May 15, 1991.

Borwein:1991:CPCb

[Bor91c] Jonathan M. Borwein. Convex programming and the choice of entropy in spectral estimation. Special session on Dynamic Optimization, CMS Summer Meeting, Université de Sherbrooke, Sherbrooke, QC, Canada., May 29, 1991.

Borwein:1991:DPC

[Bor91d] Jonathan M. Borwein. Differentiability properties of convex, of Lipschitz, and of semicontinuous mappings on Banach spaces. In Séminaire d’Initiation à l’Analyse, volume 104(19) of Publ. Math. Univ. Pierre et Marie Curie, pages 1–11. Université de Paris VI, Paris, France, 1991.

Borwein:1991:DAOa

[Bor91e] Jonathan M. Borwein. Discovering analytic objects by computer. Miniconference on Symbolic computation, Dalhousie University, Halifax, NS, Canada., January 25, 1991.

Borwein:1991:DAOb

[Bor91f] Jonathan M. Borwein. Discovering analytic objects by computer. Miniconference on Symbolic computation, Dalhousie University, Halifax, NS, Canada., January 25, 1991.

Borwein:1991:DAOc

[Bor91g] Jonathan M. Borwein. Discovering analytic objects by computer. Colloquium, Department of Mathematics, Guelph University, Guelph, ON, Canada., November 12, 1991.

Borwein:1991:EAU

[Bor91h] Jonathan M. Borwein. Estimation and approximation using infinite dimensional convex programs with entropy type objectives. Special session on Constrained Approximation, AMS Regional Meeting, University of North Dakota, Fargo, ND, USA., October 26, 1991.

Borwein:1991:EMRc

[Bor91i] Jonathan M. Borwein. Euler, Mahler, Ramanujan and a little pi: Discovering analytic objects by computer. One of two invited talks at the Festkolloquium for Dr. A. Peyerimhoff ’s 65th birthday, Ulm, Germany., April 25, 1991.

Borwein:1991:EMRa

[Bor91j] Jonathan M. Borwein. Euler, Mahler, Ramanujan: Discovering analytic objects by computer. Colloquium Pure Mathematics Department, Waterloo, Waterloo, ON, Canada., March 13, 1991.

Borwein:1991:EMRb

[Bor91k] Jonathan M. Borwein. Euler, Mahler, Ramanujan: Discovering analytic objects by computer. Number Theory Seminar, Université de Limoges, Limoges, France., April 23, 1991.

Borwein:1991:EMRd

[Bor91l] Jonathan M. Borwein. Euler, Mahler, Ramanujan: Discovering analytic objects by computer. Seminar Project Algorithms Group, INRIA, Paris., May 21, 1991.

Borwein:1991:EMRe

[Bor91m] Jonathan M. Borwein. Euler, Mahler, Ramanujan: Discovering analytic objects by computer. Colloquium, Department of Mathematics, Simon Fraser University, Burnaby, BC, Canada., June 28, 1991.

Borwein:1991:GFB

[Bor91n] Jonathan M. Borwein. On the generating function of [na + b]. International Conference on Functional Equations, Acadia University, Wolfville, NS B4P 2R6, Canada., June 5, 1991. URL

http://docserver.carma.newcastle.edu.au/1564/.

Borwein:1991:RP

[Bor91o] Jonathan M. Borwein. Ramanujan and Pi. In Ferris [Fer91], pages 647–659. ISBN 0-316-28129-8. LCCN QC71 .W67 1991. URL

http://docserver.carma.newcastle.edu.au/1379/;https://
web.archive.org/web/20170227063951/. With a foreword by

Clifton Fadiman, general editor.

Borwein:1991:RWLa

[Bor91p] Jonathan M. Borwein. Ramanujan: the wonderful life of the Indian mathematical genius S. Ramanujan (1887–1920). Seminar, Faculty of Science, Simon Fraser University, Burnaby, BC, Canada., June 27, 1991.

Borwein:1991:RWLb

[Bor91q] Jonathan M. Borwein. Ramanujan: the wonderful life of the Indian mathematical genius S. Ramanujan (1887–1920). Colloquium, Combinatorics and Optimization, University of Waterloo, Waterloo, ON, Canada., July 26, 1991.

Borwein:1991:SDPa

[Bor91r] Jonathan M. Borwein. A survey of differentiability properties of convex, Lipschitz and semicontinuous functions. Colloquium, Universität Stuttgart, Stuttgart, Germany., April 29, 1991.

Borwein:1991:SDPb

[Bor91s] Jonathan M. Borwein. A survey of differentiability properties of convex, Lipschitz and semicontinuous functions. Optimization Seminar, Université de Limoges, Limoges, France., May 17, 1991.

Borwein:1991:SDPc

[Bor91t] Jonathan M. Borwein. A survey of differentiability properties of convex, Lipschitz and semicontinuous functions. Analysis Seminar, Université de Paris VI, Paris, France., May 23, 1991.

Borwein:1991:SDPd

[Bor91u] Jonathan M. Borwein. A survey of differentiability properties of convex, Lipschitz and semicontinuous functions. Analysis Seminar, York University, Toronto, ON, Canada., October 9, 1991.

Borwein:1992:SDP

[Bor92a] J. M. Borwein. Differentiability properties of convex, of Lipschitz, and of semicontinuous mappings on Banach spaces. In Ioffe et al. [IMR92], pages 39–52. ISBN 0-582-08065-7 (paperback), 0-470-21943-2. ISSN 0269-3674. LCCN QA402.5 .O6424 1991.

Borwein:1992:BRS

[Bor92b] Jonathan Borwein. Book review: Set-valued analysis, J-P. Aubin and H. Frankowska. Bulletin of the American Mathematical Society (new series), 26(1):157–160, 1992. CODEN BAMOAD. ISSN 0273-0979 (print), 1088-9485 (electronic).

Borwein:1992:CEM

[Bor92c] Jonathan M. Borwein. A communications example: Maple and Pari. Annual Maple Retreat, Sparrow Lake, ON, Canada., June 15, 1992.

Borwein:1992:EAU

[Bor92d] Jonathan M. Borwein. Estimation and approximation using infinite dimensional convex programs with entropy type objectives. Colloquium, Industrial and Organizational Engineering, University of Michigan, Ann Arbor, MI, USA., February 19, 1992.

Borwein:1992:EMRa

[Bor92e] Jonathan M. Borwein. Euler, Mahler, Ramanujan: Discovering analytic objects by computer. Colloquium, Department of Mathematics, York University, Toronto, ON, Canada., February 6, 1992.

Borwein:1992:EMRb

[Bor92f] Jonathan M. Borwein. Euler, Mahler, Ramanujan: Discovering analytic objects by computer. Seminar, Department of Mathematics, University of Michigan, Ann Arbor, MI, USA., February 20, 1992.

Borwein:1992:FSOa

[Bor92g] Jonathan M. Borwein. First and second order differentiability of convex functions on various Banach spaces. Variational Analysis and Related Topics, University of California at Davis, Davis, CA, USA., May 16, 1992.

Borwein:1992:FSOb

[Bor92h] Jonathan M. Borwein. First and second order differentiability of convex functions on various Banach spaces. Variational Analysis and Related Topics, First World Congress of Nonlinear Analysts, Tampa, FL, USA., August 20, 1992.

Borwein:1992:GCE

[Bor92i] Jonathan M. Borwein. Guided computer experimentation in mathematics: Euler, Mahler, Ramanujan and Maple. Harry H. Gehman Lecture, MAA/OMM Meeting, Queen’s University, Kingston, ON, Canada., May 2, 1992.

Borwein:1992:IDE

[Bor92j] Jonathan M. Borwein. Infinite dimensional entropy minimization: a tutorial. 14th Symposium on Mathematical Programming with Data Perturbations, George Washington University, Washington, DC, USA., May 21, 1992.

Borwein:1992:IMSa

[Bor92k] Jonathan M. Borwein. Iterative methods for solving inverse problems and computing fixed points. Colloquium, Department of Mathematics, Statistics and Computing Science, Dalhousie University, Halifax, NS, Canada., April 9, 1992.

Borwein:1992:IMSb

[Bor92l] Jonathan M. Borwein. Iterative methods for solving inverse problems and computing fixed points. Colloquium, Department of Pure Mathematics, University of Western Ontario, London, ON, Canada., April 23, 1992.

Borwein:1992:IMSc

[Bor92m] Jonathan M. Borwein. Iterative methods for solving inverse problems and computing fixed points. Third FrancoLatin American Conference on Applied Mathematics, Santiago, Chile., September 4, 1992.

Borwein:1992:FME

[Bor92n] Jonathan M. Borwein. On the failure of ‘maximum entropy’ reconstruction for Fredholm operators and other infinite dimensional systems. 14th Symposium on Mathematical Programming with Data Perturbations, George Washington University, Washington, DC, USA, 10. June 15th., May 22, 1992.

Borwein:1993:ASS

[Bor93a] J. M. Borwein. Asplund spaces are sequentially reflexive. Accepted for publication in the Canadian Journal of Mathematics, but withdrawn and merged with another paper. Jon Borwein recorded that as publication number 121, but because the article numbers changed with each update of his CV, that number has long been incorrect., 1993.

Borwein:1993:AAL

[Bor93b] Jonathan M. Borwein. An analyst’s approach to linear inequality systems. Seminar, Department of Mathematics, University of Colorado, Boulder, CO, USA., February 12, 1993.

Borwein:1993:CAMa

[Bor93c] Jonathan M. Borwein. Computer assisted ‘mathematics and plausible reasoning’. Kempner Colloquium, Department of Mathematics, University of Colorado, Boulder, CO, USA., February 15, 1993.

Borwein:1993:CAMb

[Bor93d] Jonathan M. Borwein. Computer assisted ‘mathematics and plausible reasoning’. Colloquium, Department of Mathematics, Pennsylvania State University, State College, PA, USA., April 8, 1993.

Borwein:1993:CPE

[Bor93e] Jonathan M. Borwein. Convex programming and entropy type functions. Plenary Lecture, XVIII Symposium on Operations Research, University of Cologne., September 2, 1993.

Borwein:1993:FSOa

[Bor93f] Jonathan M. Borwein. First and second order differentiability of convex functions on various Banach spaces. Colloquium, University of Western Ontario, London, ON, Canada., February 2, 1993.

Borwein:1993:FSOb

[Bor93g] Jonathan M. Borwein. First and second order differentiability of convex functions on various Banach spaces. Regional Functional Analysis Conference, Miami University, Oxford, OH, USA., May 1, 1993.

Borwein:1993:HCPa

[Bor93h] Jonathan M. Borwein. A history of the computation of pi. Undergraduate Colloquium, University of Western Ontario, London, ON, Canada., February 3, 1993.

Borwein:1993:HCPb

[Bor93i] Jonathan M. Borwein. A history of the computation of pi. Colloquium, University of Vermont, Burlington, VT, USA., March 25, 1993.

Borwein:1993:MIE

[Bor93j] Jonathan M. Borwein. Means, iterations and experimentally induced identities. MAA-CMS Invited Lecture, Joint AMS/MAA/ CMS Summer Meetings, University of British Columbia, Vancouver, BC, Canada., August 15, 1993.

Borwein:1993:FME

[Bor93k] Jonathan M. Borwein. On the failure of maximum entropy reconstruction for Fredholm equations and other infinite systems. Mathematical Programming, 61(2, Ser. A):251–261, 1993. CO-DEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic).

URL http://docserver.carma.newcastle.edu.au/1559/.

Borwein:1993:PSPa

[Bor93l] Jonathan M. Borwein. Problems and solutions: Problems: 10281. American Mathematical Monthly, 100(1):76–77, January 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Borwein:1993:RWL

[Bor93m] Jonathan M. Borwein. S. Ramanujan: a wonderful life? South Asian Colloquium of the Pacific Northwest, Harbour Centre, Simon Fraser University, Burnaby, BC, Canada., October 16, 1993.

Borwein:1993:SIT

[Bor93n] Jonathan M. Borwein. Shrum inaugural talk. Harbour Centre, Simon Fraser University, Burnaby, BC, Canada., October 19, 1993.

Borwein:1993:SIS

[Bor93o] Jonathan M. Borwein. Some intriguing series involving (4). Tutte Seminar, Department of Combinatorics and Optimization, University of Waterloo, Waterloo, ON, Canada., May 7, 1993.

Borwein:1993:TEE

[Bor93p] Jonathan M. Borwein. Three examples of experimental computational analysis? Pacific Northwest Numerical Analysis Seminar, University of Washington, Seattle, WA, USA., October 9, 1993.

Borwein:1993:WEM

[Bor93q] Jonathan M. Borwein. What is experimental mathematics? Applied Mathematics Colloquium, University of British Columbia, Vancouver, BC, Canada., September 27, 1993.

Borwein:1994:CMI

[Bor94a] Jonathan Borwein. A convergent mean iteration: a proof that x_{n}: = M(x_{n−1}, x_{n−2}, ⋯, x_{n−k}) converges. Aequationes Mathematicae, 47(1):115–118, 1994. CODEN AEMABN. ISSN 0001-9054 (print), 1420-8903 (electronic). URL http://docserver.

carma.newcastle.edu.au/1541/.

Borwein:1994:CGS

[Bor94b] Jonathan M. Borwein. Characterizations of generalized subgradients amongst one-dimensional multifunctions: and extensions. CMS Winter Meeting, Special Session on Nonsmooth Analysis Meridien Hotel, McGill University, Montreal, QC, Canada., December 11, 1994.

Borwein:1994:EMPb

[Bor94c] Jonathan M. Borwein. Experimental mathematics, promises and pitfalls. Maple Summer Workshop and Symposium., August 11, 1994.

Borwein:1994:EMPc

[Bor94d] Jonathan M. Borwein. Experimental mathematics, promises and pitfalls. Colloquium, Department of Mathematics, Indiana University, Bloomington, IN, USA., November 18, 1994.

Borwein:1994:EMPd

[Bor94e] Jonathan M. Borwein. Experimental mathematics, promises and pitfalls. Colloquium, Department of Mathematics and Statistics, University of Calgary, Calgary, AB, Canada., November 24, 1994.

Borwein:1994:GME

[Bor94f] Jonathan M. Borwein. Greek mathematics and especially the story of the circle. High School Science Evening, Simon Fraser University, Burnaby, BC, Canada., April 14, 1994.

Borwein:1994:MEM

[Bor94g] Jonathan M. Borwein. Maximization entropy methods (using derivative information) and infinite dimensional convex programming. XV International Mathematical Programming Symposium, Ann Arbor, MI, USA., August 18, 1994.

Borwein:1994:NASa

[Bor94h] Jonathan M. Borwein. Nonsmooth analysis in smooth Banach spaces. Colloquium, Department of Mathematics, University of Washington, Seattle, WA, USA., January 18, 1994.

Borwein:1994:NASb

[Bor94i] Jonathan M. Borwein. Nonsmooth analysis in smooth Banach spaces. Analysis Seminar, University of California, Santa Barbara, Santa Barbara, CA, USA., March 4, 1994.

Borwein:1994:NASc

[Bor94j] Jonathan M. Borwein. Nonsmooth analysis in smooth Banach spaces. Colloquium, University of Victoria, Victoria, BC, Canada., April 11, 1994.

Borwein:1994:NASd

[Bor94k] Jonathan M. Borwein. Nonsmooth analysis in smooth Banach spaces. Colloquium, University of Limoges, Limoges, France., July 5, 1994.

Borwein:1994:SEC

[Bor94l] Jonathan M. Borwein. A survey of examples of convex functions and classifications of normed spaces. Report, Department of Mathematics & Statistics, Simon Fraser University, Burnaby, BC V5A 156, Canada, December 1, 1994.

13 pp. URL http://docserver.carma.newcastle.edu.au/
92;http://docserver.carma.newcastle.edu.au/93. Revised

22 January 1995. Published in [Bor95s].

Borwein:1994:VDT

[Bor94m] Jonathan M. Borwein. Viscosity derivatives: theory and applications. XV International Mathematical Programming Symposium, Ann Arbor, MI, USA., August 18, 1994.

Borwein:1994:VHD

[Bor94n] Jonathan M. Borwein. The vision: how do we integrate … mature computation techniques. Maple Summer Workshop and Symposium., August 11, 1994.

Borwein:1994:WTA

[Bor94o] Jonathan M. Borwein. Ways of thinking about duality. Student Session, XV International Mathematical Programming Symposium, Ann Arbor, MI, USA., August 16, 1994.

Borwein:1994:WEMa

[Bor94p] Jonathan M. Borwein. What is experimental mathematics? Colloquium, University of California, Santa Barbara, Santa Barbara, CA, USA., March 3, 1994.

Borwein:1994:WEMc

[Bor94q] Jonathan M. Borwein. What is experimental mathematics? Algorithms Seminar, Samedi de Recherche, University of Ottawa, Ottawa, ON, Canada. 8. June 27th., April 23, 1994.

Borwein:1994:WEMb

[Bor94r] Jonathan M. Borwein. What’s experimental mathematics? Talk to Grade 12 Students Spring Break, Simon Fraser University, Burnaby, BC, Canada., March 22, 1994.

Borwein:1995:CHNa

[Bor95a] Jonathan M. Borwein. Convex Haar null sets in separable Banach spaces. Lecture at Honoris Causa ceremony for R. T. Rockafellar, Université de Montpelier II, Montpelier, France., October 24, 1995.

Borwein:1995:CHNb

[Bor95b] Jonathan M. Borwein. Convex Haar null sets in separable Banach spaces. Functional Analysis Seminar, Department of Mathematics and Statistics, University of Saskatchewan, Saskatoon, SK, Canada., November 8, 1995.

Borwein:1995:CAD

[Bor95c] Jonathan M. Borwein. The cubic AGM discovered. Specialist Colloquium Lecture, University of Utrecht, Utrecht, The Netherlands., October 26, 1995.

Borwein:1995:MEMc

[Bor95d] Jonathan M. Borwein. Essentially strictly differentiable Lipschitz functions. Seminar, University of Newcastle, Newcastle, NSW, Australia., July 17, 1995.

Borwein:1995:EEEb

[Bor95e] Jonathan M. Borwein. Experimental evaluation of Euler sums. Halberstam retirement conference, Urbana, IL, USA, May 16–

21, 1995., May 17, 1995. URL http://docserver.carma.
newcastle.edu.au/60/.

Borwein:1995:EMPa

[Bor95f] Jonathan M. Borwein. Experimental mathematics, promises and pitfalls. Colloquium, Department of Mathematics and Computing Science, University of Northern British Columbia, Prince George, BC, Canada., April 7, 1995.

Borwein:1995:EMPb

[Bor95g] Jonathan M. Borwein. Experimental mathematics, promises and pitfalls. Principal Lecture, Australian Mathematical Society Meeting, University of Tasmania, Hobart, TAS, Australia., July 5, 1995.

Borwein:1995:EMPc

[Bor95h] Jonathan M. Borwein. Experimental mathematics, promises and pitfalls. University Public Lecture, University of Newcastle, Newcastle, NSW, Australia., July 19, 1995.

Borwein:1995:EMPd

[Bor95i] Jonathan M. Borwein. Experimental mathematics, promises and pitfalls. Colloquium, Department of Mathematics and Statistics, University of Western Australia, Crawley, WA 6009, Australia., July 27, 1995.

Borwein:1995:EMPe

[Bor95j] Jonathan M. Borwein. Experimental mathematics, promises and pitfalls. Colloquium, Department of Mathematics and Statistics, Murdoch University. Perth, WA, Australia., August 7, 1995.

Borwein:1995:EMPf

[Bor95k] Jonathan M. Borwein. Experimental mathematics: promises and pitfalls. General Colloquium Lecture, University of Utrecht, Utrecht, The Netherlands., October 26, 1995.

Borwein:1995:EMPg

[Bor95l] Jonathan M. Borwein. Experimental mathematics, promises and pitfalls. Colloquium, Department of Mathematics and Statistics, University of Saskatchewan, Saskatoon, SK, Canada., November 9, 1995.

Borwein:1995:MEMb

[Bor95m] Jonathan M. Borwein. Maximum entropy methods (using derivative information) and infinite dimensional convex programming.

Principal Lecture, Optimization Miniconference, University of NSW, Sydney, NSW, Australia. 11. July 17th., July 11, 1995.

Borwein:1995:MEMd

[Bor95n] Jonathan M. Borwein. Maximum entropy methods (using derivative information) and infinite dimensional convex programming. Pure Mathematics Seminar, University of Western Australia, Crawley, WA 6009, Australia., August 1, 1995.

Borwein:1995:MMTa

[Bor95o] Jonathan M. Borwein. Minimal multifunctions and their applications. Special Session on multivalued nonlinear dynamics, AMS Winter Meeting, Hilton Hotel, San Francisco, CA, USA., January 7, 1995.

Borwein:1995:MMTb

[Bor95p] Jonathan M. Borwein. Minimal multifunctions and their applications. Workshop on Nonsmooth Analysis and Applications, University of California at Santa Barbara, Santa Barbara, CA, USA, April 1–2., April 1, 1995.

Borwein:1995:KCa

[Bor95q] Jonathan M. Borwein. On Khinchine’s constant. Seminar, University of Newcastle, Newcastle, NSW, Australia., July 18, 1995.

Borwein:1995:KCb

[Bor95r] Jonathan M. Borwein. On Khinchine’s constant. Colloquium, Department of Mathematics and Statistics, University of Calgary, Calgary, AB, Canada., December 7, 1995.

Borwein:1995:SEC

[Bor95s] Jonathan M. Borwein. A survey of examples of convex functions and classifications of normed spaces. In Roland Durier and Chistian Michelot, editors, Recent Developments in Optimization: Seventh French–German Conference on Optimization (Dijon 1994), volume 429 of Lecture Notes in Economic and Mathematical Systems, pages 60–71. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-642-46823-3. LCCN QA402.5. URL http://docserver.carma.

newcastle.edu.au/93/.

Borwein:1995:VSFa

[Bor95t] Jonathan M. Borwein. Virtual science: the future of mathematical research. President’s Lecture Series, Simon Fraser University, Burnaby, BC, Canada., February 22, 1995.

Borwein:1995:VSFb

[Bor95u] Jonathan M. Borwein. Virtual science: the future of mathematical research. Science I, University of British Columbia, Vancouver, BC, Canada., March 2, 1995.

Borwein:1995:VDT

[Bor95v] Jonathan M. Borwein. Viscosity derivatives: theory and applications. Analysis Seminar, University of Auckland, New Zealand., June 28, 1995.

Borwein:1995:WEM

[Bor95w] Jonathan M. Borwein. What is experimental mathematics? Principal Lecture, Workshop on Experimental Mathematics, CARMA, Technical University of Denmark, Lyngby, Denmark., October 5, 1995.

Borwein:1996:CAA

[Bor96a] Jonathan M. Borwein. Convex analysis and applications. AMS Mathfest, University of Washington, Seattle, WA, USA., August 10, 1996.

Borwein:1996:DMW

[Bor96b] Jonathan M. Borwein. Doing mathematics on the Web. Colloquium, Department of Mathematics, University of British Columbia, Vancouver, BC, Canada., November 15, 1996.

Borwein:1996:EMP

[Bor96c] Jonathan M. Borwein. Experimental mathematics, promises and pitfalls. Colloquium & MAA Visiting Lecture, Department of Mathematics, Western Washington University, Bellingham, WA 98225, USA., February 6, 1996.

Borwein:1996:MPW

[Bor96d] Jonathan M. Borwein. Mathematical publishing on the Web. 10th Pacific North West Numerical Analysis Seminar, Vancouver, BC, Canada., September 21, 1996. URL http://docserver.carma.

newcastle.edu.au/190/.

Borwein:1996:MMM

[Bor96e] Jonathan M. Borwein. Multi-modal mathematics. First Annual TeleLearning Meeting and Conference (as part of Plenary — Theme 5: Post Secondary Education), Montreal, QC, Canada, November 5–7., November 5, 1996.

Borwein:1996:MESa

[Bor96f] Jonathan M. Borwein. Multidimensional Euler sums: some recent results. Combinatorics and Graph Theory Conference (in honour of Herbert Wilf’s 65th birthday), June 13–15, University of Pennsylvania, Philadelphia, PA 19104, USA., June 13, 1996.

Borwein:1996:MESb

[Bor96g] Jonathan M. Borwein. Multidimensional Euler sums: some recent results. CECM Conference on Analysis and its Computational Applications, Simon Fraser University, Burnaby, BC, Canada, August 14–15., August 14, 1996.

Borwein:1996:MESc

[Bor96h] Jonathan M. Borwein. Multidimensional Euler sums: some recent results. Fifth Canadian Number Theory Association Meeting, Carleton University, Ottawa, ON, Canada, August 17–22., August 21, 1996.

Borwein:1996:OMP

[Bor96i] Jonathan M. Borwein. The organic mathematics proceedings. Colloquium, University of Manitoba, Winnipeg, MB, Canada., March 29, 1996.

Borwein:1996:PSR

[Bor96j] Jonathan M. Borwein. Problems and solutions: Revivals: 10281. American Mathematical Monthly, 103(10):911–912, 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-

0972 (electronic). URL http://links.jstor.org/sici?sici=
0002-9890(199612)103:10<911:1>2.0.CO%3B2-5&origin=MSN.

Borwein:1996:VSC

[Bor96k] Jonathan M. Borwein. Virtual science: the changing face of mathematics. National Council of Teachers of Mathematics, Canadian Regional Meeting, Vancouver, BC, Canada, August 22–23., August 23, 1996.

Borwein:1997:DMWa

[Bor97a] Jonathan M. Borwein. Doing mathematics on the web. Colloquium, Department of Mathematics and Statistics, Simon Fraser University, Burnaby, BC, Canada., January 15, 1997.

Borwein:1997:DMWb

[Bor97b] Jonathan M. Borwein. Doing mathematics on the web. Colloquium, Department of Mathematics, Stats and CS, Dalhousie University, Halifax, NS, Canada., February 20, 1997.

Borwein:1997:DMWc

[Bor97c] Jonathan M. Borwein. Doing mathematics on the web. Colloquium, Science Faculty, Malaspina University College, Nanaimo, BC, Canada., March 26, 1997.

Borwein:1997:DMWe

[Bor97d] Jonathan M. Borwein. Doing mathematics on the web. 1997 Elizabeth Laird Lecture, University of Winnipeg, Winnipeg, MB, Canada., October 6, 1997.

Borwein:1997:DMWd

[Bor97e] Jonathan M. Borwein. Doing mathematics on the web: the organic mathematics collection. Two lectures, Canada–USA Mathcamps, Babson College, Wellesley, MA, USA., August 6, 1997.

Borwein:1997:EMD

[Bor97f] Jonathan M. Borwein. Evaluation of multi-dimensional Euler/ Zagier sums. AMS Special Session on Algebraic and Elementary Number Theory, Corvallis, OR, USA, April 19–20., April 19, 1997.

Borwein:1997:GYI

[Bor97g] Jonathan M. Borwein. A generalization of Young’s l^{p} inequality. Report, Department of Mathematics, Simon Fraser University, Burnaby, BC V5A 1S6, Canada, June 24, 1997. 6 pp. URL

http://docserver.carma.newcastle.edu.au/189/.

Borwein:1997:ISC

[Bor97h] Jonathan M. Borwein. Inverse symbolic calculation: empirical mathematics. CRM Workshop on Computer Algebra and Statistics, Montreal, QC, Canada, September 21–27., September 24, 1997.

Borwein:1997:MPWa

[Bor97i] Jonathan M. Borwein. Mathematical publication on the Web. Lecture slides, Department of Mathematics, Simon Fraser University, Burnaby, BC V5A 1S6, Canada, June 1, 1997. 20 (printed

4-up) pp. URL http://docserver.carma.newcastle.edu.au/

190. CAMS–Fields Mini–Colloquium on Technology and Mathematical Education, Toronto, ON, Canada.

Borwein:1997:MPWb

[Bor97j] Jonathan M. Borwein. Mathematical publishing on the web. Colloquium, School of Mathematical Sciences, Lakehead University, Thunder Bay, ON P7B 5E1, Canada., September 22, 1997. URL

http://docserver.carma.newcastle.edu.au/190/.

Borwein:1997:MMa

[Bor97k] Jonathan M. Borwein. The MathResource and the MathBrowser. 13 presentations at NECC, Seattle, WA, USA., June 29–30, 1997.

Borwein:1997:MEM

[Bor97l] Jonathan M. Borwein. Maximum entropy methods an introduction. VHHSC Medical Imaging Group Open House, Vancouver Hospital and Health Science Centre, Vancouver, BC, Canada., March 4, 1997.

Borwein:1997:MDP

[Bor97m] Jonathan M. Borwein. Multi-dimensional polylogarithmic sums. CRM Workshop on Experimental Mathematics and Combinatorics, Montreal, QC, Canada, May 19–23., May 20, 1997.

Borwein:1997:MMb

[Bor97n] Jonathan M. Borwein. Multimodal mathematics. Software Demonstrations, Telelearning NCE, Second Annual Conference, Toronto, ON, Canada., November 5–6, 1997.

Borwein:1997:OPT

[Bor97o] Jonathan M. Borwein. Online publishing: two views from the electronic trenches. Scholarly Communication in the Next Millennium, Simon Fraser University (Harbour Centre), Burnaby, BC, Canada, March 5–8., March 7, 1997.

Borwein:1997:PSV

[Bor97p] Jonathan M. Borwein. Partially smooth variational analysis. AMS Special Session on Optimization and Variational Analysis, Wayne State University, Detroit, MI, USA, May 2–4., May 2, 1997.

Borwein:1997:SDIa

[Bor97q] Jonathan M. Borwein. Symbolically discovered identities for zeta(4n+3) and multidimensional polylogarithms. Penn State Number Theory Conference, July 31–Aug 3., July 31, 1997.

Borwein:1997:TAPa

[Bor97r] Jonathan M. Borwein. Talking about pi. Mathematics and Statistics Department Colloquium, Western Michigan University, Kalamazoo, MI, USA., May 5, 1997.

Borwein:1997:TAPb

[Bor97s] Jonathan M. Borwein. Talking about pi. Two lectures, Canada– USA Mathcamps, Babson College, Wellesley, MA, USA., August 4, 1997.

Borwein:1997:TAPc

[Bor97t] Jonathan M. Borwein. Talking about pi. Colloquium, School of Mathematical Sciences, Lakehead University, Thunder Bay, ON P7B 5E1, Canada., September 22, 1997.

Borwein:1997:TAPd

[Bor97u] Jonathan M. Borwein. Talking about pi. Undergraduate Colloquium, University of Western Ontario, London, ON, Canada., November 3, 1997.

Borwein:1997:TASa

[Bor97v] Jonathan M. Borwein. Three adventures: Symbolically discovered identities for ζ(4n + 3) and like matters. Plenary talk, Formal Power Series and Algebraic Combinatorics, 9, Vienna, Austria, July 14–18., July 14, 1997.

Borwein:1997:TASb

[Bor97w] Jonathan M. Borwein. Three adventures: Symbolically discovered identities for ζ(4n + 3) and like matters. Joint CS/C&O Colloquium, University of Waterloo, Waterloo, ON, Canada., October 9, 1997.

Borwein:1997:VSC

[Bor97x] Jonathan M. Borwein. Virtual science: the changing face of mathematical research. Three lectures, Canada–USA Mathcamps, Babson College, Wellesley, MA, USA., August 5, 1997.

Borwein:1997:WP

[Bor97y] Jonathan M. Borwein. Why pi? Colloquium, Department of Mathematics and Statistics, University of Winnipeg, Winnipeg, MB, Canada., October 7, 1997.

Borwein:1998:MPW

[Bor98a] Jonathan Borwein. Mathematical publication on the Web. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 32(1):4–7, March 1998. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
190/;http://doi.acm.org/10.1145/294833.294836.

Borwein:1998:TAP

[Bor98b] Jonathan Borwein. Talking about pi. The original URL is no longer found, but the archive URL worked on 26-Apr-2011., January 20, 1998.

Borwein:1998:BVF

[Bor98c] Jonathan M. Borwein. Brainstorming: views of the future. Presentation, First Workshop of the IMU Committee on Electronic Information and Communication, (Nov 13–14), Konrad-Zuse-Zentrum für Informationstechnik, Berlin, Germany., November 13, 1998.

Borwein:1998:BHS

[Bor98d] Jonathan M. Borwein. Brouwer–Heyting sequences converge. The Mathematical Intelligencer, 20(1):14–15, 1998. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Borwein:1998:CNT

[Bor98e] Jonathan M. Borwein. Collaborative networking technology in the mathematical sciences. MITACS/Canada–China Opening, Asia– Pacific Centre, University of British Columbia, Vancouver, BC, Canada., November 18, 1998.

Borwein:1998:ES

[Bor98f] Jonathan M. Borwein. Euler sums. CECM98 Analysis Day, Simon Fraser University, Burnaby, BC, Canada., June 29, 1998.

Borwein:1998:GYI

[Bor98g] Jonathan M. Borwein. A generalization of Young’s l^{p} inequality. Mathematical Inequalities & Applications, 1(1):131–136, 1998. ISSN 1331-4343 (print), 1848-9966 (electronic). URL http:

//docserver.carma.newcastle.edu.au/189/.

Borwein:1998:HPS

[Bor98h] Jonathan M. Borwein. High performance symbolic computing: A mathematician’s perspective. Plenary Lecture, NESRC-MSRI Workshop on Parallel Symbolic Computation (Oct. 1–3), Berkeley, CA, USA., October 1, 1998.

Borwein:1998:JP

[Bor98i] Jonathan M. Borwein. The joy of pi. Joint presentation and book signing with D. Blatner and L. Berggren, University of Washington Bookstore, Seattle, WA, USA., April 22, 1998.

Borwein:1998:MRI

[Bor98j] Jonathan M. Borwein. Math resources: Interactive mathematics workspaces. Eleventh International Conference on Technology in Collegiate Mathematics, New Orleans, LA, USA (Nov 20–22)., November 21, 1998.

Borwein:1998:MFA

[Bor98k] Jonathan M. Borwein. Multifunctional and functional analytic methods in nonsmooth analysis. Four Lectures, NATO Advanced Study Institute on Analyse non linéaire, équations différentielles et contrôle, Université de Montréal, Montréal, QC, Canada, July 27–Aug 7., August 3–7, 1998.

Borwein:1998:PSVa

[Bor98l] Jonathan M. Borwein. Partially smooth variational analysis. Workshop talk, CMA National Symposium on Functional Analysis, Optimization and Applications, University of Newcastle, Newcastle, NSW, Australia, March 9–21., March 9, 1998.

Borwein:1998:PSVb

[Bor98m] Jonathan M. Borwein. Partially smooth variational analysis. Spring 1998 West Coast Optimization Meeting, Harbour Centre, Simon Fraser University, Burnaby, BC, Canada., April 24, 1998.

Borwein:1998:PAM

[Bor98n] Jonathan M. Borwein. Projection algorithms and monotone operators. Plenary lecture in conjunction with CMA National Symposium on Functional Analysis, Optimization and Applications, University of Newcastle (CIDACS and Mathematics), Newcastle, NSW, Australia, March 9–21., March 20, 1998.

Borwein:1998:SIT

[Bor98o] Jonathan M. Borwein. Sandwich (interpolation) theorems for Lipschitz functions. Workshop talk, CMA National Symposium on Functional Analysis, Optimization and Applications, University of Newcastle, Newcastle, NSW, Australia, March 9–21., March 13, 1998.

Borwein:1998:SNM

[Bor98p] Jonathan M. Borwein. Some new mean value inequalities. Analysis Seminar, Dalhousie University, Halifax, NS, Canada., June 10, 1998.

Borwein:1998:TAS

[Bor98q] Jonathan M. Borwein. Three adventures in symbolic computing. The Macquarie Mathematics Colloquium and Number Theory Seminar., March 31, 1998.

Borwein:1998:VSD

[Bor98r] Jonathan M. Borwein. Virtual science: doing math on the web. Public lecture in conjunction with CMA National Symposium on Functional Analysis, Optimization and Applications, University of Newcastle, Newcastle, NSW, Australia, March 9–21., March 16, 1998.

Borwein:1999:CAN

[Bor99a] Jonathan M. Borwein. Convex analysis and nonlinear optimization. Mini-course (9 hours), 5th International Conference on Approximation and Optimization in the Caribbean, Guadeloupe, March 28 April 2, 1999., March 28, 1999.

Borwein:1999:DNMa

[Bor99b] Jonathan M. Borwein. Distributed network mathematics laboratories. TL-NCE Project Leaders Meeting, Toronto, ON, Canada., June 13, 1999.

Borwein:1999:DNMb

[Bor99c] Jonathan M. Borwein. Distributed network mathematics laboratories. MITACS Day, CECM, Simon Fraser University, Burnaby, BC, Canada., August 3, 1999.

Borwein:1999:DMPc

[Bor99d] Jonathan M. Borwein. Doing math in the presence of technology. Colloquium, Department of Mathematics and Statistics, Miami University of Ohio (1999 Buckingham Fellow in Residence)., October 14, 1999.

Borwein:1999:DMPa

[Bor99e] Jonathan M. Borwein. The doing of mathematics in the presence of technology. Canadian Mathematics Education Study Group (CMESG), First Plenary, Brock University, St. Catharines, ON, Canada, June 4–8., June 4, 1999.

Borwein:1999:DMPb

[Bor99f] Jonathan M. Borwein. The doing of mathematics in the presence of technology. Session on Electronic Information and Communication, Joint Australian–American Math Society Meetings, Melbourne, VIC, Australia, July 12–15., July 13, 1999.

Borwein:1999:EMEa

[Bor99g] Jonathan M. Borwein. Experimental mathematics and exact computation. Plenary Lecture, International Symposium on Symbolic and Algebraic Commutation (ISSAC), Vancouver, BC, Canada, July 29–31, 1999., July 29, 1999.

Borwein:1999:EMEb

[Bor99h] Jonathan M. Borwein. Experimental mathematics and exact computation. Colloquium, Physics Department, University of Bologna, Bologna, Italy., September 17, 1999.

Borwein:1999:EMIa

[Bor99i] Jonathan M. Borwein. Experimental mathematics: Insight from computation. MAA Invited Address, Combined Mathematics Meetings, San Antonio, TX, USA, January 12–16., January 16, 1999.

Borwein:1999:EMIb

[Bor99j] Jonathan M. Borwein. Experimental mathematics: Insight from computation. Lecture II, Institute of Advanced Research in Mathematics (IAS), Technion, Haifa, Israel., January 25, 1999.

Borwein:1999:EMIc

[Bor99k] Jonathan M. Borwein. Experimental mathematics: Insight from computation. 2 hour Invited Address, MAA Pacific Northwest Section Meeting, Willamette University, Salem, OR, USA, March 12–13, 1999., February 8, 1999.

Borwein:1999:EMId

[Bor99l] Jonathan M. Borwein. Experimental mathematics: Insight from computation. Twenty-Seventh Annual Fall Conference: Twenty-Seventh Annual Fall Conference: “Experimental Mathematics”, Miami University, October 15–16., October 15, 1999.

Borwein:1999:GBGa

[Bor99m] Jonathan M. Borwein. Generic behaviour of generalized gradients. Special Session on Nonlinear Analysis, Canadian Mathematical Society Summer Meeting, Memorial University, St John’s, NL, Canada., May 29, 1999.

Borwein:1999:GBGb

[Bor99n] Jonathan M. Borwein. Generic behaviour of generalized gradients. Session on Nonlinear Dynamics and Optimization, Joint Australian-American Math Society Meetings, Melbourne, VIC, Australia, July 12–15., July 13, 1999.

Borwein:1999:HC

[Bor99o] Jonathan M. Borwein. Honoris causa. Acceptance speech, University of Limoges, Limoges, France., September 22, 1999.

Borwein:1999:IML

[Bor99p] Jonathan M. Borwein. Interactive mathematics labs. CECM-MITACS Day Presentation, Simon Fraser University, Burnaby, BC, Canada., November 12, 1999.

Borwein:1999:MSa

[Bor99q] Jonathan M. Borwein. Maximizing surprise. Session on Operations Research, Joint Australian-American Math Society Meetings, Melbourne, VIC, Australia, July 12–15., July 13, 1999.

Borwein:1999:MSb

[Bor99r] Jonathan M. Borwein. Maximizing surprise. Colloque: Analyse et Applications., September 23, 1999.

Borwein:1999:MSc

[Bor99s] Jonathan M. Borwein. Maximizing surprise. Colloquium, Pure Mathematics Department, University of Western Ontario, London, ON, Canada., October 1, 1999.

Borwein:1999:PSVb

[Bor99t] Jonathan M. Borwein. Partially smooth variational analysis. Nonlinear analysis seminar, Technion, Haifa, Israel., January 24, 1999.

Borwein:1999:PSVc

[Bor99u] Jonathan M. Borwein. Partially smooth variational analysis. Seventh Conference on Well-posedness and Stability of Optimization Problems, Gargnano, Italy, September 13–18., September 14, 1999.

Borwein:1999:PC

[Bor99v] Jonathan M. Borwein. Pi and its computation. Twenty-Seventh Annual Fall Conference: Twenty-Seventh Annual Fall Conference: “Experimental Mathematics”, Miami University, October 15–16., October 16, 1999.

Borwein:1999:PAT

[Bor99w] Jonathan M. Borwein. Projection algorithms & tangency formulae. Lecture III, Institute of Advanced Research in Mathematics (IAS), Technion, Haifa, Israel., January 28, 1999.

Borwein:1999:PW

[Bor99x] Jonathan M. Borwein. Publishing on the web. Burnaby Rotary Club, Burnaby, BC, Canada., February 8, 1999.

Borwein:1999:SNMa

[Bor99y] Jonathan M. Borwein. Some new mean–value inequalities. Lecture I, Institute of Advanced Research in Mathematics (IAS), Technion, Haifa, Israel., January 21, 1999.

Borwein:1999:SNMb

[Bor99z] Jonathan M. Borwein. Some new mean–value inequalities. Winter 1998 West Coast Optimization Meeting, University of Washington, Seattle, WA, USA, Feb 5–6., February 6, 1999.

Borwein:1999:SNMc

[Bor99-27] Jonathan M. Borwein. Some new mean-value theorems. Sixth Australian Optimization Day, Ballarat, VIC, Australia., July 16, 1999.

Borwein:1999:TAP

[Bor99-28] Jonathan M. Borwein. Talking about pi. Technion Mathclub Lecture, Technion, Haifa, Israel., January 27, 1999.

Borwein:1999:WP

[Bor99-29] Jonathan M. Borwein. Why pi? Dinner Address, MAA Pacific Northwest Section Meeting, Willamette University, Salem, OR, USA, March 12–13, 1999., March 13, 1999.

Borwein:19xx:EIG

[Borxx] Jonathan M. Borwein. Elliptic integrals and the Gauss–Salamin formula for π. Unpublished hand-written lecture slides at Dalhousie University. The slides are undated, and the 10 references on the final slides include a reference to E. Salamin (1976) and to P. Backmann’s 1977 book A History of Pi., 19xx.

Borwein:2000:CII

[Bor00a] Jonathan M. Borwein. CEIC–IMU initiatives. CMS special session on Mathematics on the Internet, II (MOTI-2), CMS Year 2000 Summer Meeting, Hamilton, ON, Canada, June 10–13., June 10, 2000.

Borwein:2000:EMEa

[Bor00b] Jonathan M. Borwein. Experimental mathematics and exact computation. Washington State Meeting on Exact Algorithmics, Pullman, WA, USA., April 8, 2000.

Borwein:2000:EMEb

[Bor00c] Jonathan M. Borwein. Experimental mathematics and exact computation. Colloquium, Mathematics Department, Temple University, Philadelphia, PA, USA., April 12, 2000.

Borwein:2000:EMEc

[Bor00d] Jonathan M. Borwein. Experimental mathematics and exact computation. Colloquium as Thirteenth Annual Donald H. Clanton Visiting Mathematician, Furman University, Greenville, SC 29613, USA., April 13, 2000.

Borwein:2000:EMEd

[Bor00e] Jonathan M. Borwein. Experimental mathematics and exact computation. Colloquium, University of Western Australia, Crawley, WA 6009, Australia., April 19, 2000.

Borwein:2000:EMEe

[Bor00f] Jonathan M. Borwein. Experimental mathematics and exact computation. Colloquium at GSF-Forschungszentrum Inst. für Biomathematik und Biometriei, University of Munich, München, Germany., October 4, 2000.

Borwein:2000:EMEf

[Bor00g] Jonathan M. Borwein. Experimental mathematics and exact computation. Ernst Schrödinger Lecture, Schrödinger Institute, University of Vienna, Vienna, Austria., October 5, 2000.

Borwein:2000:EMEg

[Bor00h] Jonathan M. Borwein. Experimental mathematics and exact computation. Colloquium, University of Coimbra, Coimbra, Portugal., November 27, 2000.

Borwein:2000:EMEh

[Bor00i] Jonathan M. Borwein. Experimental mathematics and exact computation. Colloquium, University of Lisbon, Lisbon, Portugal., November 28, 2000.

Borwein:2000:EMOa

[Bor00j] Jonathan M. Borwein. Experimental mathematics and other good stuff. Four Hour Lecture Series, Canada–US Mathcamps, University of British Columbia, Vancouver, BC, Canada., July 13–14,

Borwein:2000:EMOb

[Bor00k] Jonathan M. Borwein. Experimental mathematics and other good stuff. Science One Presentation, University of British Columbia, Vancouver, BC, Canada., October 17, 2000.

Borwein:2000:GBG

[Bor00l] Jonathan M. Borwein. The generic behaviour of generalized gradients. Third World Congress of Nonlinear Analysts, Special session on “Variational Analysis and Optimization”, July 19–25, 2000, Catania, Italy., July 19, 2000.

Borwein:2000:GPG

[Bor00m] Jonathan M. Borwein. Generic properties of generalized gradients. Colloque, Université des Antilles et de la Guyane, Guadeloupe., May 31, 2000.

Borwein:2000:ITD

[Bor00n] Jonathan M. Borwein. The impact of technology on the doing of mathematics. Public Lecture as Donald H. Clanton Visiting Mathematician, Furman University, Greenville, SC 29613, USA.,

April 13, 2000. URL http://docserver.carma.newcastle.edu.
au/248/.

Borwein:2000:MNI

[Bor00o] Jonathan M. Borwein. Mathématiques numerique et informatique. Conférence, 5ieme Colloque de l’IREM (Institut de recherche sur l’enseignement des mathématiques) Antilles-Guyane, Guadeloupe., June 2, 2000.

Borwein:2000:MSa

[Bor00p] Jonathan M. Borwein. Maximizing surprise. Colloque, Université des Antilles et de la Guyane, Guadeloupe., May 24, 2000.

Borwein:2000:MSb

[Bor00q] Jonathan M. Borwein. Maximizing surprise. Colloquium, Mathematics Department, Michigan State University, East Lansing, MI, USA., November 2, 2000.

Borwein:2000:MSI

[Bor00r] Jonathan M. Borwein. Multivariable sinc integrals and volumes of convex polyhedra. Special Session on Classical and Computational Analysis, Canadian Mathematical Society Winter Meeting, Vancouver, BC, Canada., December 10, 2000.

Borwein:2000:NCMb

[Bor00s] Jonathan M. Borwein. Numerical and computational mathematics at the undergraduate level. Plenary lecture, Pacific Northwest Sectional MAA Meeting, University of British Columbia, Vancouver, BC, Canada, 16–17 June., June 17, 2000. URL

http://docserver.carma.newcastle.edu.au/246/.

Borwein:2000:PSC

[Bor00t] Jonathan M. Borwein. Parallel symbolic computation: Methods and issues. Haifa-Technion Workshop on ‘Inherently parallel algorithms in optimization and feasibility and their applications’, March 14., March 13–16, 2000.

Borwein:2000:SNM

[Bor00u] Jonathan M. Borwein. Some new mean-value theorems. Colloque, Université des Antilles et de la Guyane, Guadeloupe., May 29, 2000.

Borwein:2000:TPS

[Bor00v] Jonathan M. Borwein. Tools for (partially) smooth variational analysis. Third World Congress of Nonlinear Analysts, Plenary Lecture, July 19–25, 2000, Catania, Italy., July 25, 2000.

Borwein:2000:UWH

[Bor00w] Jonathan M. Borwein. The use of wireless and handheld devices in telelearning. Panel, Telelearning Annual Meeting, Toronto, ON, Canada., November 5, 2000.

Borwein:2001:AWMa

[Bor01a] Jonathan M. Borwein. Aesthetics for the working mathematician. Report, Simon Fraser University, Burnaby, BC V5A 1S6, Canada, April 18, 2001. 22 pp. URL http://docserver.carma. newcastle.edu.au/150/. Public Lecture at Queen’s University Symposium on Beauty and the Mathematical Beast, April 18–19, Kingston, ON, Canada.

Borwein:2001:AWMb

[Bor01b] Jonathan M. Borwein. Aesthetics for the working mathematician. Special Mathematics Seminar, University of New South Wales, Sydney, NSW, Australia., August 20, 2001. URL http:

//docserver.carma.newcastle.edu.au/150/.

Borwein:2001:AWMc

[Bor01c] Jonathan M. Borwein. Aesthetics for the working mathematician. Mathematics Colloquium, Macquarie University, Sydney, NSW, Australia., August 21, 2001. URL http://docserver.carma.

newcastle.edu.au/150/.

Borwein:2001:AWMd

[Bor01d] Jonathan M. Borwein. Aesthetics for the working mathematician. Special Session on History of Mathematics, CMS Winter Meeting, Toronto, ON, Canada, December 8–10, 2001., December 9, 2001.

URL http://docserver.carma.newcastle.edu.au/150/.

Borwein:2001:CMCb

[Bor01e] Jonathan M. Borwein. Challenges in mathematical computing — why math is still hard. MAA Seaway Sectional Meeting, after dinner lecture, Brock University, St. Catharines, ON, Canada, November 2–3, 2001., November 2, 2001.

Borwein:2001:COM

[Bor01f] Jonathan M. Borwein. Collaborative online mathematics: wishing and hoping. Plenary lecture, Fields Institute Workshop on Online Mathematics, November 15–17, 2001., November 15, 2001.

Borwein:2001:DSS

[Bor01g] Jonathan M. Borwein. Dirichlet series of squares of sums of squares. ALGO Seminar, INRIA — Rocquencourt, France., October 22, 2001.

Borwein:2001:EM

[Bor01h] Jonathan M. Borwein. Experimental mathematics. Australian Mathematical Society Gazette, 28(??):77–80, ???? 2001. ISSN 0311-0729 (print), 1326-2297 (electronic). URL http:

//docserver.carma.newcastle.edu.au/1527/.

Borwein:2001:EMEa

[Bor01i] Jonathan M. Borwein. Experimental mathematics and exact computation. Number Theory Seminar, Macquarie University, Sydney, NSW, Australia., August 21, 2001.

Borwein:2001:EMEb

[Bor01j] Jonathan M. Borwein. Experimental mathematics and exact computation. Distinguished Visitor Colloquium, INRIA — Rocquencourt, France., October 23, 2001.

Borwein:2001:EMEc

[Bor01k] Jonathan M. Borwein. Experimental mathematics and exact computation. Colloquium, University of Limoges, France., October 24, 2001.

Borwein:2001:EMI

[Bor01l] Jonathan M. Borwein. Exploring math on the Internet. Esso-CMS-PIMS Math Camp, (9.00–12.00), Simon Fraser University, Burnaby, BC, Canada, June 25–29., June 28, 2001.

Borwein:2001:IMU

[Bor01m] Jonathan M. Borwein. The International Math Union’s electronic initiatives — and WestGrid. CECM01 Summer Conference, Analysis, Computation and Communication Simon Fraser, July 27– 28., July 27, 2001.

Borwein:2001:FIW

[Bor01n] Jonathan M. Borwein. The International Mathematical Union’s electronic initiatives. First International Workshop on Mathematical Knowledge Management, Sept 24–26, RISC Linz, Austria., September 25, 2001.

Borwein:2001:MET

[Bor01o] Jonathan M. Borwein. Maximum entropy-type methods and convex programming. Workshop on New Approaches to the Phase Problem, Lawrence Berkeley National Laboratory, Berkeley, CA, USA, May 17–19 (replaced by poster)., May 17, 2001.

Borwein:2001:MSIa

[Bor01p] Jonathan M. Borwein. Multivariable sinc integrals and volumes of convex polyhedra. Special Session on Series and Integrals, Combined Mathematics Meetings, New Orleans, LA, USA, January 9–13., January 10, 2001.

Borwein:2001:MSIb

[Bor01q] Jonathan M. Borwein. Multivariable sinc integrals and volumes of convex polyhedra. Analysis Seminar, University of Newcastle, Newcastle, NSW, Australia., August 16, 2001.

Borwein:2001:MSIc

[Bor01r] Jonathan M. Borwein. Multivariable sinc integrals and volumes of convex polyhedra. ALGO Seminar, INRIA — Rocquencourt, France., October 22, 2001.

Borwein:2002:EMI

[Bor02a] J. M. Borwein. Experimental mathematics and integer relations. European Research Consortium for Informatics and Mathematics, ERCIM News, 50(??):30–31, July 2002. URL https://www.

ercim.eu/publication/Ercim_News/enw50/borwein.html. Sec-

tion Special Theme ERCIMathematics.

Borwein:2002:BMO

[Bor02b] Jonathan M. Borwein. Bregman monotone optimization methods and related convex functions. Plenary Lecture, IV Brazilian Workshop on Continuous Optimization, IMPA, Rio de Janeiro, Brazil, July 15–20, 2002., July 16, 2002.

Borwein:2002:CNF

[Bor02c] Jonathan M. Borwein. The CEIC: The next four years. West Coast Optimization Fall Meeting, University of Washington, Seattle, WA, USA., November 2, 2002.

Borwein:2002:DMFa

[Bor02d] Jonathan M. Borwein. Differentiability of monotone functions on separable Banach space. Spring 2002 West Coast Optimization Meeting, Burnaby Mountain Campus, Simon Fraser University, Burnaby, BC, Canada., May 4, 2002.

Borwein:2002:DMFb

[Bor02e] Jonathan M. Borwein. Differentiability of monotone functions on separable Banach space. Nonlinear Analysis Seminar, University of Pau, 64012 Pau, France., November 13, 2002.

Borwein:2002:DLM

[Bor02f] Jonathan M. Borwein. The digital library of mathematics. Presentation at ICM Satellite Meeting on Electronic Information and Communication in Mathematics, Beijing, August 29–31, 2002., August 31, 2002.

Borwein:2002:DEM

[Bor02g] Jonathan M. Borwein. Digitizing the entire mathematical literature: what wild surmise! CMS Special Session on History of Mathematics, Ottawa, December 8–10, 2002. (Also presented to CISTI Board, December 6th.)., December 9, 2002.

Borwein:2002:DSSa

[Bor02h] Jonathan M. Borwein. Dirichlet series for squares of sums of squares. Plenary Lecture at Seventh Canadian Number Theory Association Conference at CRM, May 19–25, 2002., May 22, 2002.

Borwein:2002:DSSb

[Bor02i] Jonathan M. Borwein. Dirichlet series for squares of sums of squares. Discrete Mathematics Seminar, University of Calgary, Calgary, AB, Canada., October 18, 2002.

Borwein:2002:EMCa

[Bor02j] Jonathan M. Borwein. The experimental mathematician: A computational guide to the mathematical unknown. Plenary Lecture at The First International Congress of Mathematical Software, Beijing, China, August 17–19, 2002., August 17–19, 2002.

Borwein:2002:EMCb

[Bor02k] Jonathan M. Borwein. The experimental mathematician: A computational guide to the mathematical unknown. Numerical Analysis Seminar, University of Pau, 64012 Pau, France., November 14, 2002.

Borwein:2002:EMPa

[Bor02l] Jonathan M. Borwein. The experimental mathematician: The pleasure of discovery and the role of proof. Plenary Lecture at 25th Anniversary Meeting of the Canadian Math Educators Study Group (CMESG), Queen’s University, Kingston, ON, May 25–29, 2002., May 26, 2002. URL http://docserver.carma.

newcastle.edu.au/264/.

Borwein:2002:EMPb

[Bor02m] Jonathan M. Borwein. The experimental mathematician: The pleasure of discovery and the role of proof. Response and Discussion, 25th Anniversary Meeting of the Canadian Math Educators Study Group (CMESG), Queen’s University, Kingston, ON, Canda, May 25–29, 2002., May 27, 2002. URL http:

//docserver.carma.newcastle.edu.au/264/.

Borwein:2002:IMU

[Bor02n] Jonathan M. Borwein. The International Mathematical Union’s electronic initiatives. Workshop on Managing digital information in mathematics: From journals to the gray literature. during the Fourth Annual CEIC Meeting, Vancouver Wosk Centre, Vancouver, BC, Canada, February 15–17, 2001., February 16, 2002.

Borwein:2002:IWC

[Bor02o] Jonathan M. Borwein. Introduction to the work of the CEIC. Electronic Information Afternoon at the ICM, Beijing, August 20–27, 2002., August 26, 2002.

Borwein:2002:MMF

[Bor02p] Jonathan M. Borwein. Mathematical marvels: Fields of dreams. Simon Fraser series A Passion For Excellence, on the Nobel and like Prizes., March 26, 2002.

Borwein:2002:NFY

[Bor02q] Jonathan M. Borwein. The next four years. Invited Lecture at ICM Satellite Meeting on Electronic Information and Communication in Mathematics, Beijing, China, August 29–31, 2002., August 29, 2002.

Borwein:2002:WMD

[Bor02r] Jonathan M. Borwein. Welcome to the mathematics of dynamic SPECT. Workshop on Exploring the Frontiers of Dynamic SPECT, Wall Institute, University of British Columbia, Vancouver, BC, Canada, September 20–23, 2002., September 21, 2002.

Borwein:2002:WMSa

[Bor02s] Jonathan M. Borwein. Why math is (still) hard: Challenges for mathematical computing. Distinguished Visitors Colloquium, Wayne State, Detroit, MI, USA., March 20, 2002.

Borwein:2002:WMSb

[Bor02t] Jonathan M. Borwein. Why math is (still) hard: Challenges for mathematical computing. Colloquium, Centre de Recherches Mathématiques, Montreal, QC, Canada., April 22, 2002.

Borwein:2003:ACGc

[Bor03a] Jonathan M. Borwein. Advanced collaboration and grid computation. Presentation to the SFU Board of Governors, Burnaby, BC, Canada., November 27, 2003.

Borwein:2003:ACGa

[Bor03b] Jonathan M. Borwein. Advanced collaboration and grid computation, I. ICIAM 2003 Mini-symposium, International Congress on Industrial and Applied Mathematics, Sydney, NSW, Australia., July 9, 2003.

Borwein:2003:ACGb

[Bor03c] Jonathan M. Borwein. Advanced collaboration and grid computation, II. ICIAM 2003 Mini-symposium, International Congress on Industrial and Applied Mathematics, Sydney, NSW, Australia., July 9, 2003.

Borwein:2003:ACFa

[Bor03d] Jonathan M. Borwein. The AGM continued fraction of Ramanujan. CECM Day 2003, Simon Fraser University, Burnaby, BC, Canada., July 31, 2003.

Borwein:2003:ACFb

[Bor03e] Jonathan M. Borwein. The AGM continued fraction of Ramanujan. First Plenary Lecture, First Congress of the Mathematical Society of South East Europe (MASSEÉ), Borovets, Bulgaria., September 16, 2003.

Borwein:2003:ACFc

[Bor03f] Jonathan M. Borwein. The AGM continued fraction of Ramanujan. Colloquium, Reed College, OR, USA., October 14, 2003.

Borwein:2003:BMP

[Bor03g] Jonathan M. Borwein. Bringing math to the public. Panel Moderator, CMS National School Math Forum, May 16–18, Montreal, QC, Canada., May 18, 2003.

Borwein:2003:CHC

[Bor03h] Jonathan M. Borwein. Canadian Highend Computing Initiatives. NRC-CISTI Presentation, CISTI Advisory Board, Fredericton, NB, Canada., April 30, 2003.

Borwein:2003:DSF

[Bor03i] Jonathan M. Borwein. The department. Simon Fraser University, Burnaby, BC, Canada., February 26, 2003.

Borwein:2003:DM

[Bor03j] Jonathan M. Borwein. Discovery in mathematics. Workshop on Special Functions in the Digital Age, Simon Fraser University, January 23–24, 2003, Burnaby, BC, Canada., January 24, 2003.

Borwein:2003:EM

[Bor03k] Jonathan M. Borwein. Experimentation in mathematics. Dalhousie University, Faculty of Computing Science Colloquium, 2003, Halifax, NS, Canada., April 28, 2003.

Borwein:2003:EMC

[Bor03l] Jonathan M. Borwein. Experimentation in mathematics: Computational paths to discovery. Colloquium, University of Adelaide, Adelaide, SA, Australia., June 27, 2003.

Borwein:2003:EMPa

[Bor03m] Jonathan M. Borwein. Experimentation in mathematics: Part I. Frontiers of Mathematics, Lecture Series, Texas A&M University, College Station, TX, USA, March 22–27, 2003., March 25, 2003.

Borwein:2003:EMPb

[Bor03n] Jonathan M. Borwein. Experimentation in mathematics: Part II. Frontiers of Mathematics, Lecture Series, Texas A&M University, College Station, TX, USA, March 22–27, 2003., March 26, 2003.

Borwein:2003:FNA

[Bor03o] Jonathan M. Borwein. The Fields, Nevanlinna and Abel Prizes: Chasing the mathematical prize. In lecture series ”Recognizing Excellence. The Nobels and Other Prizes”, Series, SFU Harbour Centre, Burnaby, BC, Canada, 2003., March 13, 2003.

Borwein:2003:HEI

[Bor03p] Jonathan M. Borwein. Handling electronic issues in the international mathematical community. ICIAM 2003 Minisymposium, International Congress on Industrial and Applied Mathematics, Sydney, NSW, Australia., July 7, 2003.

Borwein:2003:LPa

[Bor03q] Jonathan M. Borwein. The life of pi. Pi Day Open House, Simon Fraser University, Burnaby, BC, Canada., March 14, 2003.

Borwein:2003:LPb

[Bor03r] Jonathan M. Borwein. The life of pi. Frontiers of Mathematics, Lecture Series, Texas A&M University, March 22–27, 2003., March 24, 2003.

Borwein:2003:LPc

[Bor03s] Jonathan M. Borwein. The life of pi. Colloquium, University of South Australia, Adelaide, SA, Australia., July 1, 2003.

Borwein:2003:LPd

[Bor03t] Jonathan M. Borwein. The life of pi. Colloquium, Royal Melbourne Institute of Technology, Melbourne, VIC, Australia., July 3, 2003.

Borwein:2003:LPe

[Bor03u] Jonathan M. Borwein. The life of pi. Colloquium, University of Newcastle, Newcastle, NSW, Australia., July 15, 2003.

Borwein:2003:LPf

[Bor03v] Jonathan M. Borwein. The life of pi. Colloquium, Dalhousie University, Halifax, NS, Canada., October 10, 2003.

Borwein:2003:LPg

[Bor03w] Jonathan M. Borwein. The life of pi. Colloquium, University of Regina, Regina, SK, Canada., October 30, 2003.

Borwein:2003:LRPa

[Bor03x] Jonathan M. Borwein. The long range plan for high-end computing in Canada. Vancouver ‘town hall‘ meeting, Simon Fraser University, Burnaby, BC, Canada., March 3, 2003.

Borwein:2003:LRPb

[Bor03y] Jonathan M. Borwein. The long range plan for high-end computing in Canada. Victoria ‘town hall’ meeting, University of Victoria, Victoria, BC, Canada., March 10, 2003.

Borwein:2003:MEPa

[Bor03z] Jonathan M. Borwein. Mathematics by experiment: Plausible reasoning in the 21st century. Colloquium, University of Adelaide, SA, Australia., June 25, 2003. URL http://docserver.carma.

newcastle.edu.au/272/.

Borwein:2003:MEPb

[Bor03-27] Jonathan M. Borwein. Mathematics by experiment: Plausible reasoning in the 21st century. 98 lecture slides from the Royal Society Lecture Series. Simon Fraser University, Burnaby, BC, Canada., October 21, 2003. URL http://www.lacim.uqam.ca/

~plouffe/articles/rsc-talk.

Borwein:2003:MEPc

[Bor03-28] Jonathan M. Borwein. Mathematics by experiment: Plausible reasoning in the 21st century. Colloquium, University of Northern British Columbia, Prince George, BC, Canada., October 23, 2003.

URL http://docserver.carma.newcastle.edu.au/272/.

Borwein:2003:MEPd

[Bor03-29] Jonathan M. Borwein. Mathematics by experiment: Plausible reasoning in the 21st century. Colloquium, University of Saskatchewan, Saskatoon, SK, Canada., October 30, 2003. URL

http://docserver.carma.newcastle.edu.au/272/.

Borwein:2003:NNM

[Bor03-30] Jonathan M. Borwein. Nurturing new mathematicians: Some advice on advising. Presentation at Panel on Supervision, Project NexTMAC, CMS 2003 Summer Meeting, University of Alberta, Edmonton, Alberta., June 13, 2003.

Borwein:2003:OWL

[Bor03-31] Jonathan M. Borwein. Official WestGrid launch. Vancouver MC, NewMIC, Edmonton, Calgary, Ottawa., May 8, 2003.

Borwein:2003:DB

[Bor03-32] Jonathan M. Borwein. On David Borwein. In Michalos [Mic03], chapter 14, page ?? ISBN 0-920354-53-X. LCCN LA2321 .B47 2003.

Borwein:2003:OFV

[Bor03-33] Jonathan M. Borwein. A one function variational principle. Eighth Conference on Well-posedness and Stability of Optimization Problems, Marseilles, France, September 8–12., September 10, 2003.

Borwein:2003:TOQ

[Bor03-34] Jonathan M. Borwein. Three open questions. Special Session in Honour of Petar Kenderov’s 60th Birthday, First Congress of the Mathematical Society of South East Europe (MASSEÉ), Bulgaria., September 17, 2003.

Borwein:2003:WDM

[Bor03-35] Jonathan M. Borwein. The world digital mathematics library. Special Session, First Congress of the Mathematical Society of South East Europe (MASSEÉ), Bulgaria., September 16, 2003.

Borwein:2004:ACGa

[Bor04a] Jonathan M. Borwein. Advanced collaboration and grid computation. Plenary Lecture, North American Knowledge Management Meeting, Phoenix, AZ, USA., January 6, 2004.

Borwein:2004:ACGb

[Bor04b] Jonathan M. Borwein. Advanced collaboration and grid computation. Seminar, Dalhousie Faculty of Computer Science, Halifax, NS, Canada., January 29, 2004.

Borwein:2004:ACGc

[Bor04c] Jonathan M. Borwein. Advanced collaboration and grid computation. Informal AARMS Workshop, Dalhousie Faculty of Computer Science, Halifax, NS, Canada., March 12, 2004.

Borwein:2004:ACEb

[Bor04d] Jonathan M. Borwein. Advanced collaborative environments. Colloquium, St Francis Xavier University, Antigonish, NS, Canada., October 1, 2004.

Borwein:2004:ACEa

[Bor04e] Jonathan M. Borwein. Advanced collaborative environments and the access grid. 4th European Math Congress, Stockholm, Sweden (delivered by Alf van der Poorten)., June 28, 2004.

Borwein:2004:ACCa

[Bor04f] Jonathan M. Borwein. Advanced computing in Canada. Presentations on the Long Range Plan for Advanced Computing in Canada to the CMS Development Group and to the Board, CMS Summer Meeting, Halifax, NS, Canada., June 13, 2004.

Borwein:2004:ACCb

[Bor04g] Jonathan M. Borwein. Advanced computing in Canada. Presentations on the Long Range Plan for HPC in Canada, University of Saskatchewan, Saskatoon, SK, Canada., November 4, 2004.

Borwein:2004:ACCc

[Bor04h] Jonathan M. Borwein. Advanced computing in Canada. Presentations on the Long Range Plan for HPC in Canada, University of Saskatchewan, Saskatoon, SK, Canada., November 5, 2004.

Borwein:2004:ASC

[Bor04i] Jonathan M. Borwein. Advanced scientific collaboration environments and the access grid. Computer Science Colloquium, University of Saskatchewan, Saskatoon, SK, Canada., November 4, 2004.

Borwein:2004:AGMb

[Bor04j] Jonathan M. Borwein. The Atlantic gateway to mathematics. First AGATE-M Annual Conference, Mount Allison University, Sackville, NB, Canada, December 3–4., December 4, 2004.

Borwein:2004:AGMc

[Bor04k] Jonathan M. Borwein. The Atlantic gateway to mathematics. Presentation to Nova Scotia Provincial Mathematics Team Meeting., December 16, 2004.

Borwein:2004:BSCa

[Bor04l] Jonathan M. Borwein. Bumps, slices and cusps. Plenary Lecture on Nonsmooth Analysis, First Franco–Canadian Math Meeting, Toulouse, France, July 12–15., July 15, 2004.

Borwein:2004:BSCb

[Bor04m] Jonathan M. Borwein. Bumps, slices and cusps. Sixth Midwest Optimization Seminar, Plenary Talk, Wayne State University, Detroit, MI, USA, September 11., September 11, 2004.

Borwein:2004:DBM

[Bor04n] Jonathan M. Borwein. David Borwein and me: a chronology. CMS Summer Meeting, Halifax, NS, Canada., June 14, 2004.

Borwein:2004:DMO

[Bor04o] Jonathan M. Borwein. Decomposition of monotone operators. Workshop on Control, Set-Valued Analysis and Applications University of French Antilles and Guyana, April 5–8., April 5, 2004.

Borwein:2004:ED

[Bor04p] Jonathan M. Borwein. Engines of discovery: The long range plan for HPC in Canada. Seminar, Dalhousie Faculty of Computer Science, Halifax, NS, Canada., November 8, 2004.

Borwein:2004:EMa

[Bor04q] Jonathan M. Borwein. Experimentation in mathematics. Graduate Seminar, Mathematics Department, Dalhousie University, Halifax, NS, Canada., January 28, 2004.

Borwein:2004:EMb

[Bor04r] Jonathan M. Borwein. Experimentation in mathematics. Opening Lecture, Workshop on Experimental Mathematics, Oakland, CA, USA, March 29–30., March 29, 2004.

Borwein:2004:EMc

[Bor04s] Jonathan M. Borwein. Experimentation in mathematics. Plenary Lecture, East Coast Computer Algebra Day, Wilfred Laurier University, Waterloo, ON N2L 3C5, Canada., May 8, 2004.

Borwein:2004:IEM

[Bor04t] Jonathan M. Borwein. Implications of experimental mathematics for the philosophy of mathematics. CMS Winter Meeting, Session on History of Mathematics, McGill University, Montreal, QC, Canada., December 12, 2004. URL http://docserver.carma.

newcastle.edu.au/280/.

Borwein:2004:ME

[Bor04u] Jonathan M. Borwein. Mathematics by experiment. Opening Lecture, Discovery by Computer GERAD-DIMACS Workshop, Montreal, June 2–5., June 2, 2004.

Borwein:2004:MS

[Bor04v] Jonathan M. Borwein. Maximizing surprise. Informal AARMS Workshop, Dalhousie Faculty of Computer Science, Halifax, NS, Canada., March 11, 2004.

Borwein:2004:PRCb

[Bor04w] Jonathan M. Borwein. Plausible reasoning in the 21st century. Regular Lecture, ICME10, Copenhagen, Denmark, July 5–11., July 9, 2004.

Borwein:2004:PRCc

[Bor04x] Jonathan M. Borwein. Plausible reasoning in the 21st century. Colloquium, St Francis Xavier University, Antigonish, NS, Canada., October 1, 2004.

Borwein:2004:PRCd

[Bor04y] Jonathan M. Borwein. Plausible reasoning in the 21st century. Acadia Symposium on Modelling and Computation, Acadia University, Wolfville, NS B4P 2R6, Canada., October 4, 2004.

Borwein:2004:PRCe

[Bor04z] Jonathan M. Borwein. Plausible reasoning in the 21st century. Maritime Teachers Association, New Glasgow, NS, Canada., October 24, 2004.

Borwein:2004:PRCa

[Bor04-27] Jonathan M. Borwein. Plausible reasoning in the 21st century I & II. NSF Undergraduate Research Experience Lectures, Clemson University, Clemson, SC, USA., June 25, 2004.

Borwein:2004:RACc

[Bor04-28] Jonathan M. Borwein. Ramanujan’s AGM continued fractions and dynamics. Workshop on Analytic and Computational Number Theory, August 23–27, Dalhousie University, Halifax, NS, Canada., August 27, 2004.

Borwein:2004:RACb

[Bor04-29] Jonathan M. Borwein. Ramanujan’s AGM continued fractions and dynamics: the complex case. Analysis Seminar, Mathematics Department, Dalhousie University, Halifax, NS, Canada., March 10, 2004.

Borwein:2004:RACa

[Bor04-30] Jonathan M. Borwein. Ramanujan’s AGM continued fractions and dynamics: the real case. Colloquium, Mathematics Department, Dalhousie University, Halifax, NS, Canada., March 4, 2004.

Borwein:2004:SFV

[Bor04-31] Jonathan M. Borwein. A single function variational principle and applications. Large Scale Nonlinear and Semidefinite Programming Workshop, University of Waterloo, ON, Canada, May 12– 15., May 14, 2004.

Borwein:2004:SMA

[Bor04-32] Jonathan M. Borwein. Surprise maximization: Avoiding a paradox. Mathematics Colloquium, University of Saskatchewan, Saskatoon, SK, Canada., November 5, 2004.

Borwein:2004:WC

[Bor04-33] Jonathan M. Borwein. The work of the CEIC. Presentation to ICMI General Assembly, ICME10, Copenhagen, Denmark, July 5–11., July 9, 2004.

Borwein:2005:EMPa

[Bor05a] J. M. Borwein. The experimental mathematician: The pleasure of discovery and the role of proof. International Journal of Computers for Mathematical Learning, 10(2):75– 108, May 2005. ISSN 1382-3892 (print), 1573-1766 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
264/;http://link.springer.com/article/10.1007/s10758-005-5216-x;https:
//web.archive.org/web/20040330173752/. Counterpart pre-

sentation published in CMESG25 Proceedings, 2002, with lecture

slides at http://www.cecm.sfu.ca/personal/jborwein/proof.
pdf.

Borwein:2005:TTC

[Bor05b] Jonathan M. Borwein. (2 times) ten challenge problems. Third Clifford Lecture, Tulane University, New Orleans, LA, USA., April 1, 2005.

Borwein:2005:GV

[Bor05c] Jonathan M. Borwein. 32 Goldbach variations. Analysis Seminar, Dalhousie University, Halifax, NS, Canada., November 18, 2005.

Borwein:2005:A

[Bor05d] Jonathan M. Borwein. Aarms. Presentation, Department of Math and Stats, Memorial University, St John’s, NL, Canada., November 17, 2005.

Borwein:2005:AP

[Bor05e] Jonathan M. Borwein. “AARMS” presentation. Department of Math and Stats, University of New Brunswick, Fredericton, NB, Canada., November 1, 2005.

Borwein:2005:ATS

[Bor05f] Jonathan M. Borwein. Apéry-type series: a case study. Fourth Clifford Lecture, Tulane University, New Orleans, LA, USA., April 2, 2005.

Borwein:2005:BRS

[Bor05g] Jonathan M. Borwein. Book review: The SIAM 100-Digit challenge: a study in high-accuracy numerical computing, Folkmar Bornemann, Dirk Laurie, Stan Wagon, and JSrg Waldvogel, SIAM, Philadelphia, PA, USA 2004, xil + 306 pp. softcover, ISBN 0-89871 561-X, US$57.00. The Mathematical Intelligencer, 27(4): 40–48, 2005. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Borwein:2005:CLC

[Bor05h] Jonathan M. Borwein. Computational lists and challenges in mathematics? Applied and Computational Mathematics Seminar, Dalhousie University, Halifax, NS, Canada., October 28, 2005.

Borwein:2005:DC

[Bor05i] Jonathan M. Borwein. The digital congress. Site visit presentation, Canadian bid to hold ICM 2010 in Montreal, QC, Canada., April 20, 2005.

Borwein:2005:DSC

[Bor05j] Jonathan M. Borwein. Dynamics of some continued fractions originating with Ramanujan. Special Session on Dynamical Systems, Combined Mathematics Meetings, Atlanta, GA, USA., January 6, 2005.

Borwein:2005:EMW

[Bor05k] Jonathan M. Borwein. East meets West: Collaboration goes national. Delivered over the Access Grid to the opening of IRMACS at Simon Fraser University, Burnaby, BC, Canada., April 8, 2005.

Borwein:2005:EDLa

[Bor05l] Jonathan M. Borwein. Engines of discovery: the long range plan for HPC in Canada. Presentation to HPCS05, May 15–18, 2005., May 17, 2005.

Borwein:2005:EDLb

[Bor05m] Jonathan M. Borwein. Engines of discovery: the long range plan for HPC in Canada. Presentation to ACOA, Halifax, NS., May 27, 2005.

Borwein:2005:EDLc

[Bor05n] Jonathan M. Borwein. Engines of discovery: the long range plan for HPC in Canada. Presentation to NSERC Executive VP, Ottawa, ON, Canada., May 31, 2005.

Borwein:2005:EDLd

[Bor05o] Jonathan M. Borwein. Engines of discovery: the long range plan for HPC in Canada. Presentation to NRC President, Ottawa., May 27, 2005.

Borwein:2005:EDLe

[Bor05p] Jonathan M. Borwein. Engines of discovery: the long range plan for HPC in Canada. Presentation to Industry Canada, Ottawa, ON, Canada., August 15, 2005.

Borwein:2005:EMPb

[Bor05q] Jonathan M. Borwein. Experimental mathematics and its philosophical implications. Colloquium, Australian Mathematical Sciences Institute, Melbourne, VIC, Australia., October 5, 2005.

Borwein:2005:EMPc

[Bor05r] Jonathan M. Borwein. Experimental mathematics and its philosophical implications. Colloquium, Dias Abertos, University of Porto, Porto, Portugal., November 26, 2005.

Borwein:2005:FH

[Bor05s] Jonathan M. Borwein. The future is here? Presentation to National Educational Forum, Fields Institute, Toronto, ON M5T 3J1, Canada, May 6–8., May 6, 2005.

Borwein:2005:HPMa

[Bor05t] Jonathan M. Borwein. High performance mathematics. First Plenary, HPCS05, Guelph, ON, Canada, May 15–18, 2005., May 16, 2005.

Borwein:2005:HPMb

[Bor05u] Jonathan M. Borwein. High performance mathematics. Presentation to HPC@Dal, Dalhousie University, Halifax, NS, Canada., June 10, 2005.

Borwein:2005:HPMc

[Bor05v] Jonathan M. Borwein. High performance mathematics. Presentation to Media Light Paths project ‘kick off’, June 10, 2005., June 10, 2005.

Borwein:2005:HPMd

[Bor05w] Jonathan M. Borwein. High performance mathematics. Plenary Lecture, Ontario R&E Summit, Toronto, ON, Canada, June 13– 14, 2005., June 14, 2005.

Borwein:2005:HIW

[Bor05x] Jonathan M. Borwein. Hilbert’s inequality and Witten’s zeta. Trans Canada Computational Science Seminar., August 25, 2005.

Borwein:2005:LPa

[Bor05y] Jonathan M. Borwein. The life of pi. Colloquium: University of Melbourne, Melbourne, VIC, Australia., October 3, 2005.

Borwein:2005:LPb

[Bor05z] Jonathan M. Borwein. The life of pi. Colloquium: La Trobe University, Melbourne, VIC, Australia., October 4, 2005.

Borwein:2005:LCM

[Bor05-27] Jonathan M. Borwein. Lists and challenges in mathematics? Colloquium, Mathematics Department, Rutgers, the State University of New Jersey., November 10, 2005.

Borwein:2005:LGC

[Bor05-28] Jonathan M. Borwein. The LRP, grid computing and ACE’s. Dalhousie Senate Computing and Information Technology Planning Committee (SCITPC) Presentation, Halifax, NS, Canada., March 2, 2005.

Borwein:2005:MEa

[Bor05-29] Jonathan M. Borwein. Mathematics by experiment. Dalhousie Math Circles — for High Schools, Halifax, NS, Canada., March 3, 2005.

Borwein:2005:MEb

[Bor05-30] Jonathan M. Borwein. Mathematics by experiment, I. First Clifford Lecture, Tulane University, New Orleans, LA, USA., March 31, 2005.

Borwein:2005:MEI

[Bor05-31] Jonathan M. Borwein. Mathematics by experiment, II. Second Clifford Lecture, Tulane University, New Orleans, LA, USA., April 1, 2005.

Borwein:2005:MS

[Bor05-32] Jonathan M. Borwein. Maximizing surprise. Seminar, University of Aviero, Centre for Studies on Optimization and Control, Aveiro, Portugal., November 25, 2005.

Borwein:2005:MEM

[Bor05-33] Jonathan M. Borwein. Maximum entropy methods and (non-) convex programming. Special Session on Nonsmooth Analysis and Imaging, Combined Mathematics Meetings, Atlanta, GA, USA., January 7, 2005.

Borwein:2005:MOCa

[Bor05-34] Jonathan M. Borwein. Monotone operators as convex objects. 6th Stability Workshop, Borevetz, Bulgaria, September 5–9., September 8, 2005.

Borwein:2005:MOCb

[Bor05-35] Jonathan M. Borwein. Monotone operators as convex objects. Keynote talk, Fitzpatrick Memorial Meeting, Perth, WA, Australia, September 25th., September 25, 2005.

Borwein:2005:MOCc

[Bor05-36] Jonathan M. Borwein. Monotone operators as convex objects. Keynote Address, 6th Midwest Optimization Meeting, Kalamazoo, MI, USA., October 14, 2005.

Borwein:2005:MOCd

[Bor05-37] Jonathan M. Borwein. Monotone operators as convex objects. Colloquium, University of Lisbon, Lisbon, Portugal., November 23, 2005.

Borwein:2005:PIEa

[Bor05-38] Jonathan M. Borwein. Philosophical implication of experimental (computational) mathematics. Philosophy of Mathematics, Invited Presentation, MAA, Atlanta, GA, USA., January 7, 2005.

Borwein:2005:PIEb

[Bor05-39] Jonathan M. Borwein. Philosophical implication of experimental (computational) mathematics. Honours Seminar, Mathematics Department, Dalhousie University, Halifax, NS, Canada., January 26, 2005.

Borwein:2005:RB

[Bor05-40] Jonathan M. Borwein. Reality bytes. Joint German-Austrian-AMS Meeting, Mainz, Germany, June 16–19, 2005. (Given by Martin Groetschel)., June 16, 2005.

Borwein:2005:SEC

[Bor05-41] Jonathan M. Borwein. Symbolic and experimental computation. Shad Valley Afternoon in D-DRIVE., July 12, 2005.

Borwein:2005:VOMa

[Bor05-42] Jonathan M. Borwein. Visualisation and other mathematical learning tools. TransCanada Computational Science Seminar., September 13, 2005.

Borwein:2005:VOMb

[Bor05-43] Jonathan M. Borwein. Visualisation and other mathematical learning tools. Lecture to Teachers, Australian Math Society Meetings, Perth, WA, Australia, September 26–30th., September 28, 2005.

Borwein:2005:VOMc

[Bor05-44] Jonathan M. Borwein. Visualisation and other mathematical learning tools. Dalhousie Math Circles–for High Schools., November 15, 2005.

Borwein:2005:VOT

[Bor05-45] Jonathan M. Borwein. Visualization and other tools for mathematics. Colloquium, Statistics and Operations Research Department, University of North Carolina., November 9, 2005.

Borwein:2005:WDD

[Bor05-46] Jonathan M. Borwein. What is D-DRIVE? Exhibit and presentation, Dymaxion Exchange, Halifax, NS, Canada., April 27, 2005.

Borwein:2005:WHPa

[Bor05-47] Jonathan M. Borwein. What is high performance mathematics. Shad Valley Colloquium., July 19, 2005.

Borwein:2005:WHPb

[Bor05-48] Jonathan M. Borwein. What is high performance mathematics. First Plenary Lecture, Australian Math Society Meetings, Perth, WA, Australia, September 26–30th., September 26, 2005.

Borwein:2005:WHPc

[Bor05-49] Jonathan M. Borwein. What is high performance mathematics? Colloquium, Department of Math and Stats, Western Michigan University, Kalamazoo, MI, USA., October 13, 2005.

Borwein:2005:WHPd

[Bor05-50] Jonathan M. Borwein. What is high performance mathematics? Colloquium, Department of Math and Stats, University of New Brunswick, Fredericton., November 1, 2005.

Borwein:2005:WHPe

[Bor05-51] Jonathan M. Borwein. What is high performance mathematics. Colloquium, University of Aviero, Aveiro, Portugal., November 25, 2005.

Borwein:2005:WHPf

[Bor05-52] Jonathan M. Borwein. What is high performance mathematics? Colloquium, University of Lisbon, Lisbon, Portugal., November 28, 2005.

Borwein:2006:SFM

[Bor06a] Jonathan Borwein. Simon Fitzpatrick memorial volume. Journal of Convex Analysis, 13(3–4):463–476, 2006. ISSN 0944-6532 (print), 2363-6394 (electronic). URL http://www.heldermann.

de/JCA/JCA13/JCA133/jca13040.htm. Introduction to Fitz-

patrick memorial volume.

Borwein:2006:ACEb

[Bor06b] Jonathan M. Borwein. Advanced collaborative environments. Symposium on the Computer: the once and future medium for the social sciences and humanities, York University, Toronto, ON, Canada, May 30., May 30, 2006.

Borwein:2006:ACEc

[Bor06c] Jonathan M. Borwein. Advanced collaborative environments. Presentation to Canadian International Olympiad Team, Dalhousie University, Halifax, NS, Canada., June 28, 2006.

Borwein:2006:ACEa

[Bor06d] Jonathan M. Borwein. Advanced collaborative environments and tools. Workshop on Intelligent Computing, Dalhousie, March 6–7, 2006., March 6, 2006.

Borwein:2006:AWM

[Bor06e] Jonathan M. Borwein. Aesthetics for the working mathematician. In N. Sinclair and W. Higginson, editors, Mathematics and the aesthetic, volume 25 of CMS Books Math./Ouvrages Math. SMC, pages 21–40. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. URL http:

//docserver.carma.newcastle.edu.au/150/.

Borwein:2006:BP

[Bor06f] Jonathan M. Borwein. Backing up the planet. Interview on Quirks and Quarks, CBC Radio One., November 25, 2006.

Borwein:2006:CE

[Bor06g] Jonathan M. Borwein. Collaborative environments. Panel Discussion HPCS 06, Memorial University, St John’s, NL, Canada., May 17, 2006.

Borwein:2006:CAD

[Bor06h] Jonathan M. Borwein. Computer-assisted discovery and proof of generating functions for zeta functions. Atlantic Analysis Days, January 20–21, Dalhousie University, Halifax, NS, Canada., January 21, 2006.

Borwein:2006:DSC

[Bor06i] Jonathan M. Borwein. Dynamics of some continued fractions originating with Ramanujan. Number Theory Seminar, University of Waterloo, Waterloo, ON, Canada., January 31, 2006.

Borwein:2006:EEBa

[Bor06j] Jonathan M. Borwein. Effective error bounds for Euler– Maclaurin-based quadrature schemes. HPCS 06, Memorial University, St John’s, NL, Canada., May 16, 2006. URL http:

//docserver.carma.newcastle.edu.au/297/.

Borwein:2006:EEBb

[Bor06k] Jonathan M. Borwein. Effective error bounds for Euler– Maclaurin-based quadrature schemes. Computer Science Seminar, University of Saskatchewan, Saskatoon, SK, Canada., May 26,

2006. URL http://docserver.carma.newcastle.edu.au/297/.

Borwein:2006:ED

[Bor06l] Jonathan M. Borwein. Engines of discovery: the long range plan for HPC in Canada. Presentation to NSERC President and Senior Administrators, Ottawa., April 12, 2006.

Borwein:2006:EMEa

[Bor06m] Jonathan M. Borwein. Experimental math and extreme quadrature, I. Analysis Seminar, D-DRIVE, Dalhousie University, Halifax, NS, Canada., October 25, 2006.

Borwein:2006:EMEb

[Bor06n] Jonathan M. Borwein. Experimental math and extreme quadrature, II. Analysis Seminar, D-DRIVE, Dalhousie University, Halifax, NS, Canada., November 1, 2006.

Borwein:2006:FRO

[Bor06o] Jonathan M. Borwein. Featured review: Oxford Users’ Guide to Mathematics. Edited by Eberhard Zeidler. Oxford University Press, Oxford, 2004. $59.50. xxii + 1285 pp., softcover. ISBN 0-19-850763-1. SIAM Review, 48(3):585–594, September 2006. CODEN SIREAD. ISSN 0036-1445 (print), 1095-

7200 (electronic). URL http://epubs.siam.org/doi/abs/10.
1137/SIREAD000048000003000585000001;https://www.jstor.
org/stable/20453842.

Borwein:2006:FLb

[Bor06p] Jonathan M. Borwein. Final lecture. MAA Shortcourse on Experimental Mathematics, San Antonio, TX, USA., January 11,

Borwein:2006:FLa

[Bor06q] Jonathan M. Borwein. First lecture. MAA Shortcourse on Experimental Mathematics, San Antonio, TX, USA., January 10, 2006.

Borwein:2006:FLVa

[Bor06r] Jonathan M. Borwein. Four lectures on variational principles. I: Bumps, cusps and slices. Spring School on Analysis, Paseky, Czech Republic., April 24, 2006.

Borwein:2006:FLVb

[Bor06s] Jonathan M. Borwein. Four lectures on variational principles. II: Monotone operators as convex objects. Spring School on Analysis, Paseky, Czech Republic, April 25, 2006.

Borwein:2006:FLVc

[Bor06t] Jonathan M. Borwein. Four lectures on variational principles. III: Decompositions of monotone operators. Spring School on Analysis, Paseky, Czech Republic, April 26, 2006.

Borwein:2006:FLVd

[Bor06u] Jonathan M. Borwein. Four lectures on variational principles. IV: Chebyshev sets and proximality. Spring School on Analysis, Paseky, Czech Republic., April 28, 2006.

Borwein:2006:HPMb

[Bor06v] Jonathan M. Borwein. High performance mathematics and its management. Colloquium, Information Technology Laboratory, NIST, Washington, DC, USA., June 16, 2006.

Borwein:2006:HPMc

[Bor06w] Jonathan M. Borwein. High performance mathematics and its management. Colloquium, Department of Math and Stats, York University, Toronto, ON, Canada., November 27, 2006.

Borwein:2006:HPMd

[Bor06x] Jonathan M. Borwein. High performance mathematics and its management. Colloquium, AMSI Access Grid Conference, La Trobe University, Melbourne, VIC, Australia (given over Access Grid)., December 3, 2006.

Borwein:2006:HPMe

[Bor06y] Jonathan M. Borwein. High performance mathematics and its management. IMA Hot Topics Workshop, The Evolution of Mathematical Communication in the Age of Digital Libraries, December 8–9., December 8, 2006.

Borwein:2006:HPMa

[Bor06z] Jonathan M. Borwein. High performance mathematics in Maple. Seminar, MapleSoft, Waterloo, Waterloo, ON, Canada., March 24, 2006.

Borwein:2006:LP

[Bor06-27] Jonathan M. Borwein. Life of pi. Colloquium, Mona Campus, University of the West Indies, Kingston, Jamaica., March 13, 2006.

Borwein:2006:LRP

[Bor06-28] Jonathan M. Borwein. The long-range plan for advanced computing in Canada. Presentation to Petter Nicholson, President Council of Canadian Academies (CCA), Ottawa, ON, Canada., September 28, 2006.

Borwein:2006:MPR

[Bor06-29] Jonathan M. Borwein. Mathematics and plausible reasoning. ISM, Graduate Student Conference, Laval, QC, Canada. May 23–25, 2006., May 24, 2006.

Borwein:2006:MMC

[Bor06-30] Jonathan M. Borwein. Maximal monotonicity via convex analysis. Journal of Convex Analysis, 13(3–4):561–586, 2006. ISSN 0944-6532 (print), 2363-6394 (electronic). URL http://docserver.

carma.newcastle.edu.au/1235/;http://www.heldermann.de/
JCA/JCA13/JCA133/jca13045.htm.

Borwein:2006:MSM

[Bor06-31] Jonathan M. Borwein. Maximality of sums of monotone operators. Special Session on Functional Analysis., December 10, 2006.

Borwein:2006:MST

[Bor06-32] Jonathan M. Borwein. Maximality of sums of two maximal monotone operators. Proceedings of the American Mathematical Society, 134(10):2951–2955, 2006. CODEN PAM-YAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL

http://docserver.carma.newcastle.edu.au/1234/;https://
www.jstor.org/stable/4098152.

Borwein:2006:MEM

[Bor06-33] Jonathan M. Borwein. Maximum entropy methods and (non-) convex programming. Plenary talk Atlantic Optimization Days, Fredericton, NB, Canada, Oct 5–6., October 5, 2006.

Borwein:2006:MOCa

[Bor06-34] Jonathan M. Borwein. Monotone operators as convex objects. Optimization Seminar, University of Waterloo, Waterloo, ON, Canada., January 30, 2006.

Borwein:2006:MOCb

[Bor06-35] Jonathan M. Borwein. Monotone operators as convex objects. Plenary Lecture, Mathematics of Optimization and Decision, Guadeloupe, April 18–21, 2006., April 19, 2006.

Borwein:2006:NDT

[Bor06-36] Jonathan M. Borwein. Notes from, the digital trenches: the work of the CEIC. C2C Seminar, from D-DRIVE., October 10, 2006.

Borwein:2006:WHPa

[Bor06-37] Jonathan M. Borwein. What is high performance mathematics? Colloquium, University of Waterloo, Waterloo, ON, Canada., January 30, 2006.

Borwein:2006:WHPb

[Bor06-38] Jonathan M. Borwein. What is high performance mathematics? Mathematics Colloquium, Acadia University, Wolfville, NS B4P 2R6, Canada., March 3, 2006.

Borwein:2006:WHPc

[Bor06-39] Jonathan M. Borwein. What is high performance mathematics? Colloquium, Spring School on Analysis, Paseky, Czech Republic., April 26, 2006.

Borwein:2007:APP

[Bor07a] Jonathan M. Borwein. AARMS: Past, present and future. Colloquium, Department of Math and Stats, Dalhousie University, Halifax, NS, Canada., February 5, 2007.

Borwein:2007:ADMb

[Bor07b] Jonathan M. Borwein. Asplund decompositions of monotone operators. In Alain Pietrus and Michel H. Geoffroy, editors, CSVAA 2004—control set-valued analysis and applications, ESAIM: Proceedings, volume 17 of ESAIM Proc., pages 19–25. EDP Sci., Les

Ulis, France, 2007. URL http://docserver.carma.newcastle.
edu.au/299/.

Borwein:2007:BLS

[Bor07c] Jonathan M. Borwein. The book of Lawrence: a serious satire. Presentation, Halifax Unitarian Universalist Assembly, Halifax, NS, Canada., February 4, 2007.

Borwein:2007:CSF

[Bor07d] Jonathan M. Borwein. The C2C seminar: Five years of experience with the Access Grid. HPCS07, University of Manitoba, Winnipeg, MB, Canada., May 13–16, 2007.

Borwein:2007:CDS

[Bor07e] Jonathan M. Borwein. A class of Dirichlet series integrals. American Mathematical Monthly, 114(1):70–76, January 2007. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
262/;https://www.jstor.org/stable/27642121.

Borwein:2007:CTH

[Bor07f] Jonathan M. Borwein. Collaborative technology: High performance mathematics and its management. AARMS/ACENet/ MITAC Summer Workshop on High Performance Computing in the Mathematical Sciences, Acadia University, Wolfville, NS, July 9–12., July 12, 2007.

Borwein:2007:CAD

[Bor07g] Jonathan M. Borwein. Computer-assisted discovery and proof of generating functions for Riemann’s zeta. First Lecture, MAA Special Session on Experimental Mathematics., January 5, 2007.

Borwein:2007:ECB

[Bor07h] Jonathan M. Borwein. Effective computation of Bessel functions. 8th Bluenose Numerical Analysis Days, St Marys University, Halifax, NS, Canada., July 27, 2007.

Borwein:2007:ELAb

[Bor07i] Jonathan M. Borwein. Effective Laguerre asymptotics, I. Dalhousie Analysis Seminar, Dalhousie University, Halifax, NS, Canada., January 30, 2007.

Borwein:2007:ELAc

[Bor07j] Jonathan M. Borwein. Effective Laguerre asymptotics, II. Dalhousie Analysis Seminar, Dalhousie University, Halifax, NS, Canada., February 14, 2007.

Borwein:2007:EDPb

[Bor07k] Jonathan M. Borwein. Experimental discovery and proof of generating functions. Special session on Algorithmic Challenges in Polynomial and Linear Algebra CMS Winter Meeting, University of Western Ontario, London, ON, Canada., December 9, 2007.

Borwein:2007:EDPa

[Bor07l] Jonathan M. Borwein. Experimental discovery and proof of identities. Chaitin 60th Birthday Symposium, IBM Watson Centre, Yorktown, NY, USA., November 15, 2007.

Borwein:2007:EMA

[Bor07m] Jonathan M. Borwein. Experimental mathematics in action. Carleton College, Northfield, MN., July 16–20, 2007. URL http:

//docserver.carma.newcastle.edu.au/1733/.

Borwein:2007:EMV

[Bor07n] Jonathan M. Borwein. Experimental mathematics with variational applications. ICCOPT 2007, McMaster University, Hamilton, ON, Canada., August 11, 2007.

Borwein:2007:FFS

[Bor07o] Jonathan M. Borwein. Finding funding in the sciences. Interdisciplinary PhD Presentation, Dalhousie University, Halifax, NS, Canada., March 29, 2007.

Borwein:2007:IWW

[Bor07p] Jonathan M. Borwein. Interdisciplinarity: what works, what doesn’t. Colloquium, University of Newcastle, Newcastle, NSW, Australia., October 23, 2007.

Borwein:2007:IRW

[Bor07q] Jonathan M. Borwein. Interdisciplinary research: What works, what doesn’t. Keynote address, 2007 Faculty Research Day, Dalhousie Faculty of Management, Dalhousie University, Halifax, NS, Canada., June 12, 2007.

Borwein:2007:IEMa

[Bor07r] Jonathan M. Borwein. Introduction to experimental mathematics. Honours Seminar, Dalhousie Mathematics Department., September 19, 2007.

Borwein:2007:IEMb

[Bor07s] Jonathan M. Borwein. Introduction to experimental mathematics. Colloquium, Okanagan Community College, Kelowna, BC, Canada., September 26, 2007.

Borwein:2007:IEMc

[Bor07t] Jonathan M. Borwein. Introduction to experimental mathematics: Insight through computation. Interdisciplinary Colloquium, IR-MACS, Simon Fraser University, Burnaby, BC, Canada., September 28, 2007.

Borwein:2007:IEMd

[Bor07u] Jonathan M. Borwein. Introduction to experimental mathematics: Insight through computation. Applied Mathematics Colloquium, Cornell University, Ithaca, NY, USA., November 16, 2007.

Borwein:2007:LPT

[Bor07v] Jonathan M. Borwein. The life of pi — a talk for pi day. Analysis Seminar, Dalhousie University, Halifax, NS, Canada., March 14, 2007.

Borwein:2007:MWN

[Bor07w] Jonathan M. Borwein. Math: What’s new, what’s possible, what’s coming? Talks to Grade nine and ten students at AB Lucas Secondary School, London, ON, Canada., December 12, 2007.

Borwein:2007:MST

[Bor07x] Jonathan M. Borwein. Maximality of sums of two maximal monotone operators in general Banach space. Proceedings of the American Mathematical Society, 135(12):3917–3924, December 2007. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-

6826 (electronic). URL http://docserver.carma.newcastle.
edu.au/322/;https://www.jstor.org/stable/20535029.

Borwein:2007:PCS

[Bor07y] Jonathan M. Borwein. Proximality and Chebyshev sets. Optimization Letters, 1(1):21–32, 2007. ISSN 1862-4472 (print), 1862-

4480 (electronic). URL http://docserver.carma.newcastle.
edu.au/323/.

Borwein:2007:SS

[Bor07z] Jonathan M. Borwein. Setting the stage. Opening presentation to Workshop on Math Knowledge Management: Sustainability, Scalability, and Interoperability, Dalhousie University, Halifax, NS, Canada, April 26–28, 2007., April 26, 2007.

Borwein:2007:SCR

[Bor07-27] Jonathan M. Borwein. Some convexity results a Jon or a Thompson might like. 65th Birthday Colloquium lecture for Jon Thompson, (Inter-Campus Seminar Day), University of New Brunswick, Moncton, NB, Canada., June 8, 2007.

Borwein:2007:SMFa

[Bor07-28] Jonathan M. Borwein. Some of my favourite convexity results. OCANA Seminar, Department of Mathematics, Statistics and Physics, UBCO, Kelowna, BC., September 27, 2007.

Borwein:2007:SMFb

[Bor07-29] Jonathan M. Borwein. Some of my favourite convexity results. Math Colloquium, University of Newcastle, Newcastle, NSW, Australia., October 25, 2007.

Borwein:2007:SMFc

[Bor07-30] Jonathan M. Borwein. Some of my favourite convexity results. Applied Nonlinear Optimization Day, CORS and Ddrive, Dalhousie University, Halifax, NS, Canada., November 19, 2007.

Borwein:2007:TCR

[Bor07-31] Jonathan M. Borwein. Three convexity results. Second Annual AARMS Analysis Days, in honour of Tony Thompson, Dalhousie University, Halifax, NS, Canada, March 30–31., March 30, 2007.

Borwein:2007:WNW

[Bor07-32] Jonathan M. Borwein. What’s new, what’s possible, what’s coming? Dalhousie Open House, Dalhousie FCS, Halifax, NS, Canada., February 10, 2007.

Borwein:2008:VPG

[Bor08a] J. M. Borwein. La vita di pi greco. (Italian) [The life of Greek pi]. In ????, editor, Mathematics and Culture, La matematica: Problemi e teoremi, page ?? Guilio Einaudi Editori, Turino, Italy, 2008. ISBN ???? LCCN ???? URL http://www.carma.

newcastle.edu.au/~jb616/pi-2010.pdf.

Borwein:2008:IEM

[Bor08b] Jonathan Borwein. Implications of experimental mathematics for the philosophy of mathematics. In Gold and Simons [GS08], chapter 2, pages 33–59. ISBN 0-88385-567-4. LCCN QA8.4 .P755 2008.

URL http://docserver.carma.newcastle.edu.au/280/.

Borwein:2008:CCI

[Bor08c] Jonathan M. Borwein. The computer as crucible: an introduction to experimental mathematics. Physics Colloquium, University of Newcastle, Newcastle, NSW, Australia., March 6, 2008. URL

http://docserver.carma.newcastle.edu.au/1730/.

Borwein:2008:CADa

[Bor08d] Jonathan M. Borwein. Computer-assisted discovery and proof. Colloquium, Department of Mathematics and Statistics, University of Melbourne, Melbourne, VIC, Australia., September 30,

2008. URL http://docserver.carma.newcastle.edu.au/338/.

Borwein:2008:CADb

[Bor08e] Jonathan M. Borwein. Computer-assisted discovery and proof: Part I. Colloquium, Department of Mathematics, Australian National University, Canberra, ACT, Australia., November 13, 2008.

Borwein:2008:CADc

[Bor08f] Jonathan M. Borwein. Computer-assisted discovery and proof: Part II. Colloquium, Department of Mathematics, Australian National University, Canberra, ACT, Australia., November 14, 2008.

Borwein:2008:DAD

[Bor08g] Jonathan M. Borwein. Digitally-assisted discovery and proof. Invited lecture in Special Session on University Mathematics Education, Teaching and Learning, 7th NZ-AustMS Joint Meeting, ANZMC2008, Christchurch, New Zealand, December 8–12., De-

cember 11, 2008. URL http://docserver.carma.newcastle.
edu.au/390/.

Borwein:2008:ECB

[Bor08h] Jonathan M. Borwein. Effective computation of Bessel functions. SIAM-AMS Special Session on Special Functions, Combined Membership Meetings, San Diego, CA, USA, Jan 6–9, 2008., January 6, 2008.

Borwein:2008:FCVa

[Bor08i] Jonathan M. Borwein. Future challenges for variational analysis. Plenary Lecture (presented by Andrew Eberhard), Conference on Variational Analysis and Nonlinear Optimization, National Sun-Yat-Sen University, Kaohsiung, Taiwan, November 28–30.,

November 29, 2008. URL http://docserver.carma.newcastle.
edu.au/396/.

Borwein:2008:FCVb

[Bor08j] Jonathan M. Borwein. Future challenges for variational analysis. Plenary Lecture in Special Session on Nonlinear Optimization and Applications, 7th NZ-AustMS Joint Meeting, ANZMC2008, Christchurch, New Zealand, December 8–12., December 10, 2008.

URL http://docserver.carma.newcastle.edu.au/396/.

Borwein:2008:HIW

[Bor08k] Jonathan M. Borwein. Hilbert’s inequality and Witten’s zetafunction. American Mathematical Monthly, 115(2):125–137, February 2008. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL https://www.jstor.org/stable/
27642418.

Borwein:2008:LPTa

[Bor08l] Jonathan M. Borwein. Life of pi: a talk for pi day. AMSI Access Grid Colloquium, Australia., March 14, 2008.

Borwein:2008:LPTb

[Bor08m] Jonathan M. Borwein. Life of pi: a talk for pi day. IRMACS Access Grid Colloquium, Simon Fraser University, Burnaby, BC, Canada, Canada., March 15, 2008.

Borwein:2008:MWNa

[Bor08n] Jonathan M. Borwein. Math: What’s new, what’s possible, what’s coming? ICE-EM/ICE day from Victoria University, Melbourne VIC 3000, Australia, presented from Wollongong Access Grid Room., August 13, 2008.

Borwein:2008:MWNb

[Bor08o] Jonathan M. Borwein. Math: What’s new, what’s possible, what’s coming? Colloquium, Department of Mathematical Sciences, RMIT, Melbourne, VIC, Australia., October 1, 2008.

Borwein:2008:MEMb

[Bor08p] Jonathan M. Borwein. Maximum entropy methods for inverse problems. Colloquium, Department of Mathematical Sciences, RMIT, Melbourne, VIC, Australia., September 29, 2008.

Borwein:2008:MEMa

[Bor08q] Jonathan M. Borwein. My experiences with mathematical software. Seminar, Department of Mathematics, University of Newcastle, Newcastle, NSW, Australia., August 20, 2008.

Borwein:2008:PYM

[Bor08r] Jonathan M. Borwein. The past 60 years in mathematics. Colloquium, Department of Mathematics, University of Auckland, Auckland, New Zealand., December 4, 2008.

Borwein:2008:PBR

[Bor08s] Jonathan M. Borwein. Peter Borwein revisited: A brother’s retrospective. IRMACS Conference celebrating Peter Borwein’s 55th Birthday, Simon Fraser University, Burnaby, BY, Canada, May 12–16, 2008., May 14, 2008.

Borwein:2008:PCS

[Bor08t] Jonathan M. Borwein. Proximality and Chebyshev sets. Analysis Seminar, University of Newcastle, Newcastle, NSW, Australia.,

July 3, 2008. URL http://docserver.carma.newcastle.edu.
au/323/.

Borwein:2008:SMF

[Bor08u] Jonathan M. Borwein. Some of my favourite convex functions. NZIMA Plenary Lecture, 7th NZ-AustMS Joint Meeting, ANZMC2008, Christchurch, New Zealand, December 8–12., December 9, 2008.

Borwein:2008:U

[Bor08v] Jonathan M. Borwein. [untitled]. CEIC Presentation to IMU Executive, Hungarian Academy of Science, Budapest, Hungary., April 21, 2008.

Borwein:2009:DAD

[Bor09a] J. M. Borwein. Digitally-assisted discovery and proof. In lai Lin [lL09], pages I.3–I.11. ISBN 986-01-8210-8. LCCN QA109;

QA11.A1; QA9.54. URL http://docserver.carma.newcastle.

edu.au/390/. Plenary address.

Borwein:2009:PCM

[Bor09b] Jonathan M. Borwein. The Princeton companion to mathematics. Timothy Gowers, with June Barrow Green and Imre Leader. Princeton University Press, Princeton, NJ, 2008. $99.00. xxii + 1034 pp., hardcover. ISBN 978-0-691-1 1880-2. [book review of Cno. 2467561]. SIAM Review, 51(4):790–794, December 2009. CODEN SIREAD. ISSN 0036-1445 (print), 1095-

7200 (electronic). URL http://docserver.carma.newcastle.
edu.au/1228/;https://www.jstor.org/stable/25662346.

Borwein:2009:CSS

[Bor09c] Jonathan M. Borwein. Compressed sensing: A subgradient descent algorithm for missing data problems. CARMA Colloquium, Newcastle, NSW, Australia., November 19, 2009.

Borwein:2009:CCE

[Bor09d] Jonathan M. Borwein. The computer as crucible: The end of theory? Second Annual Rubinov Memorial Lecture, University of Ballarat, VIC, Australia., November 9, 2009.

Borwein:2009:DAMa

[Bor09e] Jonathan M. Borwein. Digitally-assisted mathematical discovery and proof. CARMA Colloquium, University of Newcastle, Newcastle, NSW, Australia., May 7, 2009.

Borwein:2009:DAMb

[Bor09f] Jonathan M. Borwein. Digitally-assisted mathematical discovery and proof. Plenary address, ICMI Study 19., May 11, 2009.

Borwein:2009:DAMc

[Bor09g] Jonathan M. Borwein. Digitally-assisted mathematical discovery and proof. Responses to Plenary address, ICMI Study 19., May 14, 2009.

Borwein:2009:EECa

[Bor09h] Jonathan M. Borwein. Exploratory experimentation and computation. Colloquium, Mathematics Dept, University of Victoria, BC, Canada., September 15, 2009. URL http://docserver.

carma.newcastle.edu.au/1396/.

Borwein:2009:EECb

[Bor09i] Jonathan M. Borwein. Exploratory experimentation and computation. Plenary lecture Fields–IRMACS Workshop on Discovery and Experiment in Number Theory, Simon Fraser University, Burnaby, BC, Canada and Toronto, ON, Canada, September 22–26., September 23, 2009. URL http://docserver.carma.

newcastle.edu.au/1396/.

Borwein:2009:FYMa

[Bor09j] Jonathan M. Borwein. Fifty years of maximal monotonicity. Plenary lecture for Optimization Theory and Methods Special Session, 53rd Annual Australian Mathematical Society Meetings, University of South Australia, Adelaide SA 5001, Australia, Sept 28-Oct 1., September 29, 2009. URL http://docserver.carma.

newcastle.edu.au/1221/.

Borwein:2009:FYMb

[Bor09k] Jonathan M. Borwein. Fifty years of maximal monotonicity. Session on Functional Analysis, University of Newcastle, Official Opening and Workshop Oct 30-Nov 1, October 31, 2009. URL

http://docserver.carma.newcastle.edu.au/1221/.

Borwein:2009:FVA

[Bor09l] Jonathan M. Borwein. The future of variational analysis. Keynote Lecture, Special session on Convex and nonlinear analysis, CMS Winter Meeting, Windsor, ON, Canada, Dec. 5–7., December 6, 2009.

Borwein:2009:HIW

[Bor09m] Jonathan M. Borwein. Hilbert inequalities and Witten zeta functions. AMSI-SIGopt Seminar. Delivered to RMIT and USA from Newcastle, NSW, Australia., August 20, 2009.

Borwein:2009:HMS

[Bor09n] Jonathan M. Borwein. How to maximize surprise. AMSI-SIGopt Seminar. Delivered to RMIT and USA from Newcastle, NSW, Australia., June 25, 2009.

Borwein:2009:ICM

[Bor09o] Jonathan M. Borwein. Innovation and creativity or managing a research portfolio? Keynote Address, Annual Research Day, University of Ballarat, Ballarat, VIC, Australia., November 11, 2009.

Borwein:2009:IRA

[Bor09p] Jonathan M. Borwein. Integer relation algorithms, I. CARMA Number Theory Seminar., November 29, 2009.

Borwein:2009:IRM

[Bor09q] Jonathan M. Borwein. Integer relation methods: an introduction. Special Session on Scientific Computation, First Pacific Rim Conference on Mathematics and Applications (PRIMA), University of New South Wales, Sydney, NSW, Australia, July 6–10, 2009., July 9, 2009.

Borwein:2009:IMC

[Bor09r] Jonathan M. Borwein. International mathematics in the 21st century. Minisymposium in honour of Phil Broadbridge’s retirement as Director of AMSI, University of Melbourne, Melbourne, VIC, Australia, June 15., June 15, 2009.

Borwein:2009:IC

[Bor09s] Jonathan M. Borwein. Introduction to CARMA. Presentation to students from Dungog High School in CARMA., August 11, 2009.

Borwein:2009:ISC

[Bor09t] Jonathan M. Borwein. Inverse symbolic computation: Symbols from numbers. Education Afternoon, 53rd Annual Australian Mathematical Society Meetings, University of South Australia, Adelaide SA 5001, Australia, Sept 28-Oct 1., September 29, 2009.

Borwein:2009:MME

[Bor09u] Jonathan M. Borwein. Maths matters: Exploratory experimentation: digitally-assisted discovery and proof. Australian Mathematical Society Gazette, 36(3):166–175, July 2009. ISSN 0311-0729 (print), 1326-2297 (electronic).

Borwein:2009:MEP

[Bor09v] Jonathan M. Borwein. Maximum entropy and projection methods for inverse problems. CARMA Colloquium, University of Newcastle, Newcastle, NSW, Australia, February 12, 2009.

Borwein:2009:PRCa

[Bor09w] Jonathan M. Borwein. Prospects for remote collaboration. IR-MACS Retreat on Remote Collaboration, April 24–25, Simon Fraser University, Burnaby, BC, Canada. Delivered from Newcastle, NSW, Australia., April 25, 2009.

Borwein:2009:PRCb

[Bor09x] Jonathan M. Borwein. Prospects for remote collaboration. First AMSI-SIGopt Seminar. Delivered to RMIT and USA from Newcastle, NSW, Australia., April 30, 2009.

Borwein:2009:PM

[Bor09y] Jonathan M. Borwein. The psychology of mathematics. Math Drudge, November 15, 2009. URL https://experimentalmath.

info/blog/2009/11/192/.

Borwein:2009:SDC

[Bor09z] Jonathan M. Borwein. The SIAM 100 Digits Challenge: a story of modern numerical analysis. CARMA Colloquium, University of Newcastle, Newcastle, NSW, Australia., May 28, 2009.

Borwein:2009:TLVe

[Bor09-27] Jonathan M. Borwein. Ten lectures on variational approaches to minimization problems: Algebraic reconstruction methods and interactive geometry (and a final lecture on surprise). IMA 2009 Summer Program for Graduate Students on The Mathematics of Inverse Problems, University of Delaware, Newark, DE, USA., July 3, 2009.

Borwein:2009:TLVb

[Bor09-28] Jonathan M. Borwein. Ten lectures on variational approaches to minimization problems: Convex duality and applications. IMA 2009 Summer Program for Graduate Students on The Mathematics of Inverse Problems, University of Delaware, Newark, DE, USA., June 30, 2009.

Borwein:2009:TLVd

[Bor09-29] Jonathan M. Borwein. Ten lectures on variational approaches to minimization problems: Monotone and nonexpansive maps: algorithms and convergence. IMA 2009 Summer Program for Graduate Students on The Mathematics of Inverse Problems, University of Delaware, Newark, DE, USA., July 2, 2009.

Borwein:2009:TLVa

[Bor09-30] Jonathan M. Borwein. Ten lectures on variational approaches to minimization problems: Motivation and overview. IMA 2009 Summer Program for Graduate Students on The Mathematics of Inverse Problems, University of Delaware, Newark, DE, USA., June 29, 2009.

Borwein:2009:TLVc

[Bor09-31] Jonathan M. Borwein. Ten lectures on variational approaches to minimization problems: Variational principles and applications. IMA 2009 Summer Program for Graduate Students on The Mathematics of Inverse Problems, University of Delaware, Newark, DE, USA., July 1, 2009.

Borwein:2009:WC

[Bor09-32] Jonathan M. Borwein. Why convex? Plenary Lecture, CMS Winter Meeting, Windsor, ON, Canada, December 5–7., December 5, 2009.

Borwein:2010:MEPa

[Bor10a] J. M. Borwein. Mathematics by experiment: Plausible reasoning in the 21st century. In ICME10 Proceedings, July 2010, page ??

????, ????, 2010. URL http://docserver.carma.newcastle.
edu.au/272/.

Borwein:2010:GV

[Bor10b] Jonathan M. Borwein. 32 Goldbach variations. CARMA Colloquium and Number Theory and Analysis Seminar, University of Newcastle, Newcastle, NSW, Australia., March 19, 2010.

Borwein:2010:APAa

[Bor10c] Jonathan M. Borwein. Alternating projection algorithms in Hilbert space, I. CARMA Analysis and Optimization Seminar, University of Newcastle, Newcastle, NSW, Australia., April 20, 2010.

Borwein:2010:APAb

[Bor10d] Jonathan M. Borwein. Alternating projection algorithms in Hilbert space, II. CARMA Analysis and Optimization Seminar, University of Newcastle, Newcastle, NSW, Australia., April 27, 2010.

Borwein:2010:ASR

[Bor10e] Jonathan M. Borwein. The arithmetic of 3 and 4 step random walks. Keynote Lecture, AMSI-CARMA Workshop on Exploratory Experimentation and Computation Theory, CARMA, July 7–9., July 9, 2010.

Borwein:2010:CFWa

[Bor10f] Jonathan M. Borwein. Closed forms: what they are and why they matter. Part I, CARMA Number Theory seminar., September 15,

2010. URL http://docserver.carma.newcastle.edu.au/767/.

Borwein:2010:CFWb

[Bor10g] Jonathan M. Borwein. Closed forms: what they are and why they matter. Part II, CARMA Number Theory seminar., September

22, 2010. URL http://docserver.carma.newcastle.edu.au/

Borwein:2010:CSS

[Bor10h] Jonathan M. Borwein. Compressed sensing: a subgradient approach. Special session on Optimization, 54th Australian Math Society Meetings, September 27–30, 2010., September 28, 2010.

Borwein:2010:DRIc

[Bor10i] Jonathan M. Borwein. Douglas–Rachford iterations in the absence of convexity. Clemson Research Experience for Undergraduates, presented from Newcastle, NSW, Australia., June 22, 2010.

Borwein:2010:DRId

[Bor10j] Jonathan M. Borwein. Douglas–Rachford iterations in the absence of convexity. Keynote Lecture, AMSI-CARMA Workshop on Applied Functional Analysis, CARMA, October 2–4., October 3, 2010.

Borwein:2010:EPM

[Bor10k] Jonathan M. Borwein. Entropy and projection methods for convex and nonconvex inverse problems. Prepared for Univ of South Australia Distinguished Lecture (54 slides)., March

29, 2010. URL http://citeseerx.ist.psu.edu/viewdoc/
download?doi=10.1.1.175.5066&rep=rep1&type=pdf.

Borwein:2010:EECa

[Bor10l] Jonathan M. Borwein. Exploratory experimentation and computation. First Plenary Lecture, 2010 German Math Society Meetings (joint with Mathematical Education), Munich, Germany, March 8–12., March 8, 2010. URL http://docserver.carma.

newcastle.edu.au/1396/.

Borwein:2010:EECb

[Bor10m] Jonathan M. Borwein. Exploratory experimentation and computation. Mathematics Department Colloquium, University of Adelaide, Adelaide, SA, Australia., April 16, 2010. URL http:

//docserver.carma.newcastle.edu.au/1396/.

Borwein:2010:FYM

[Bor10n] Jonathan M. Borwein. Fifty years of maximal monotonicity. Optimization Letters, 4(4):473–490, 2010. ISSN 1862-4472 (print), 1862-4480 (electronic). URL http://docserver.carma.

newcastle.edu.au/1221/.

Borwein:2010:FPC

[Bor10o] Jonathan M. Borwein. Fractal postcards and coke cans. Presentation to West Wallsend High School students in CARMA., September 24, 2010.

Borwein:2010:FCV

[Bor10p] Jonathan M. Borwein. Future challenges for variational analysis. In Variational analysis and generalized differentiation in optimization and control, volume 47 of Springer Optim. Appl., pages 95–107. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. URL http://docserver.

carma.newcastle.edu.au/396/.

Borwein:2010:HPC

[Bor10q] Jonathan M. Borwein. High precision computation in mathematical physics and dynamics. AMSI-SIGopt Seminar. Delivered to Australian National University, UWA, RMIT and USA from Newcastle, NSW, Australia., June 2, 2010.

Borwein:2010:IRA

[Bor10r] Jonathan M. Borwein. Integer relation algorithms, II. CARMA Number Theory Seminar., January 20, 2010.

Borwein:2010:IC

[Bor10s] Jonathan M. Borwein. An introduction to CARMA. University of Newcastle 2010 Teachers’ Visit Day, presented inside CARMA., June 28, 2010.

Borwein:2010:LP

[Bor10t] Jonathan M. Borwein. Life of pi. Public Lecture, 54th Australian Math Society Meetings, September 27–30, 2010., September 28, 2010.

Borwein:2010:LPT

[Bor10u] Jonathan M. Borwein. Life of pi: a talk for pi day. Newcastle Undergraduate Mathematics Club, Newcastle, NSW, Australia., March 15, 2010.

Borwein:2010:MEPb

[Bor10v] Jonathan M. Borwein. Maximum entropy and projection methods for inverse problems. Distinguished Lecture Series, Mathematics Dept, University of South Australia, Adelaide SA 5001, Australia., March 29, 2010.

Borwein:2010:MEPc

[Bor10w] Jonathan M. Borwein. Maximum entropy and projection methods for inverse problems. Plenary Lecture, Second South Pacific Conference on Mathematics, University of New Caledonia, Nouméa 98851, New Caledonia., August 31, 2010.

Borwein:2010:MRGa

[Bor10x] Jonathan M. Borwein. Moments of Ramanujan’s generalized elliptic integrals and extensions of Catalan’s constant. CARMA Colloquium, University of Newcastle, NSW, Australia., August

12, 2010. URL http://docserver.carma.newcastle.edu.au/
1443/.

Borwein:2010:MZV

[Bor10y] Jonathan M. Borwein. Multiple zeta values. CARMA Number Theory and Analysis Seminar, University of Newcastle, Newcastle, NSW, Australia., March 31, 2010.

Borwein:2010:RACa

[Bor10z] Jonathan M. Borwein. Ramanujan’s AG continued fraction, I: the real case. CARMA Colloquium., November 4, 2010.

Borwein:2010:RACb

[Bor10-27] Jonathan M. Borwein. Ramanujan’s AG continued fraction, II: the complex case. CARMA Number Theory Seminar., November 4, 2010.

Borwein:2010:RWR

[Bor10-28] Jonathan M. Borwein. Random walks and ramble integrals. Analysis Seminar, University of Adelaide, Adelaide, SA, Australia., April 16, 2010.

Borwein:2010:SWR

[Bor10-29] Jonathan M. Borwein. Short walks and ramble integrals: The arithmetic of uniform random walks. Plenary Lecture, 54th Australian Math Society Meetings, September 27–30, 2010., September 30, 2010.

Borwein:2010:TTT

[Bor10-30] Jonathan M. Borwein. Talking to Telstra: Two weeks spent with Australia’s largest Telco. Response to Australian Communications and Media Authority (ACMA) enquiry, Reconnecting the Customer, Sept 27, 2010., September 27, 2010.

Borwein:2010:TTG

[Bor10-31] Jonathan M. Borwein. Thirty two Goldbach variations. CARMA Workshop on Multi-zeta Values, University of Newcastle, Newcastle, NSW, Australia., October 20, 2010.

Borwein:2010:U

[Bor10-32] Jonathan M. Borwein. [untitled]. General interview on Newcastle Drive, ABC Radio Newcastle, NSW, Australia., March 30, 2010.

Borwein:2010:WC

[Bor10-33] Jonathan M. Borwein. Why convex? Colloquium, Universität der Bundeswehr München, Germany., March 10, 2010.

Borwein:2011:CCNb

[Bor11a] J. M. Borwein. Chiropractic: crackers now, and crackers way back when. The Conversation, ??(??):??, De-

cember 23, 2011.  URL https://theconversation.com/
chiropractic-crackers-now-and-crackers-way-back-when-4836.

Borwein:2011:LEW

[Bor11b] J. M. Borwein. Letter to the Editor: Why are mathematics papers so dull? SIAM News, 44(6):7, July/August 2011. ISSN 0036-1437.

URL http://www.siam.org/news/news.php?id=1903.

Borwein:2011:WMP

[Bor11c] J. M. Borwein. Why are mathematics papers so dull? SIAM News, 44(6):7, July 15, 2011. ISSN 0036-1437. URL

http://docserver.carma.newcastle.edu.au/1450/;http://
www.siam.org/news/news.php?id=1903.

Borwein:2011:PDNa

[Bor11d] Jonathan Borwein. Are pi’s days numbered? The Conversation,

??(??):??, May 4, 2011. URL https://theconversation.com/
are-pis-days-numbered-39.

Borwein:2011:IBC

[Bor11e] Jonathan Borwein. If I had a blank cheque I’d … turn IBM Watson into a maths genius. The Conversation, ??

(??):??, July 8, 2011. URL https://theconversation.com/
if-i-had-a-blank-cheque-id-turn-ibms-watson-into-a-maths-genius-1213.

Borwein:2011:MMS

[Bor11f] Jonathan Borwein. Mahler measures, short walks and log-sine integrals: A case study in hybrid computation. In Proceedings of the 2011 International Workshop on Symbolic–Numeric Computation, SNC ’11, page 1. ACM Press, New York, NY 10036, USA,

2011. ISBN 1-4503-0515-6. URL http://doi.acm.org/10.1145/
2331684.2331685.

Borwein:2011:ATR

[Bor11g] Jonathan M. Borwein. Actually: Teaching and researching at the tertiary level with collaboration tools. CARMA Colloquium., November 3, 2011.

Borwein:2011:PDNb

[Bor11h] Jonathan M. Borwein. Are pi’s days numbered? Interview with ABC Goldcoast, Australia., July 18, 2011.

Borwein:2011:BND

[Bor11i] Jonathan M. Borwein. BBP numbers and digit-extraction algorithms. CARMA Seminar., May 23, 2011.

Borwein:2011:CM

[Bor11j] Jonathan M. Borwein. CARMA and me. New Fellows Seminar, Australian Academy of Science, Shine Dome, Canberra, ACT, Australia., May 4, 2011.

Borwein:2011:CMIa

[Bor11k] Jonathan M. Borwein. CARMA and me: An introduction. CDSC-CARMA-CISRA (Canon Information Systems Research Australia) afternoon, CARMA., April 26, 2011.

Borwein:2011:CMIb

[Bor11l] Jonathan M. Borwein. CARMA and me: An introduction. First CARMA Retreat, Shortland Wetland., July 19, 2011.

Borwein:2011:CMWa

[Bor11m] Jonathan M. Borwein. CARMA and me: or why am i in Oz? JonFest 2011, IRMACS, Simon Fraser University, Burnaby, BC, Canada, 16–20 May., May 17, 2011.

Borwein:2011:CMWb

[Bor11n] Jonathan M. Borwein. CARMA and me: or why am i in Oz? Two presentations to 2011 Teachers’ Visit Day, University of Newcastle, NSW, Australia. July 8., June 30, 2011.

Borwein:2011:CCNa

[Bor11o] Jonathan M. Borwein. Chiropractic: crackers now, and crackers way back when. Math Drudge, December 23,

2011. URL https://experimentalmath.info/blog/2011/12/
chiropractic-crackers-now-and-crackers-way-back-when/.

Borwein:2011:CSS

[Bor11p] Jonathan M. Borwein. Compressed sensing: a subgradient approach. Special Session on Optimization, ANZIAM 2011., February 1, 2011.

Borwein:2011:DCF

[Bor11q] Jonathan M. Borwein. Difference convex functions. Workshop on Continuous Optimization, University of South Australia, Adelaide SA 5001, Australia., January 29, 2011.

Borwein:2011:DRIa

[Bor11r] Jonathan M. Borwein. Douglas–Rachford iterations in the absence of convexity. AMS Special Session on Douglas–Rachford Methods, Combined Membership Meetings, New Orleans, LA, USA., January 9, 2011.

Borwein:2011:DRIb

[Bor11s] Jonathan M. Borwein. Douglas–Rachford iterations in the absence of convexity. Plenary Lecture (given from Newcastle), 13th Midwest Optimization Meeting and Workshop on Large Scale Optimization and Applications, Fields Institute, Toronto, ON M5T 3J1, Canada, October 14–15., October 15, 2011.

Borwein:2011:EEC

[Bor11t] Jonathan M. Borwein. Exploratory experimentation and computation. AMS Special Session in Logic and Analysis, Combined Membership Meetings, New Orleans, LA, USA., January 7, 2011.

URL http://docserver.carma.newcastle.edu.au/1396/.

Borwein:2011:FG

[Bor11u] Jonathan M. Borwein. Fractal geometry. Presentation to Year 7 students form Wallsend with Michael Rose to the NSW MEGS program (Making Educational Goals Sustainable)., February 16, 2011.

Borwein:2011:IAP

[Bor11v] Jonathan M. Borwein. The infinite appeal of pi. ABC Science, ??(??):??, March 10, 2011. URL http://www.abc.net.au/

science/articles/2011/03/10/3158045.htm.

Borwein:2011:LP

[Bor11w] Jonathan M. Borwein. Life of pi. Graduate Colloquium, Mathematics Dept, University of Florida, Gainesville, FL, USA., January 12, 2011.

Borwein:2011:LPHa

[Bor11x] Jonathan M. Borwein. Life of pi: History and computation — a talk for pi day. University of Technology Sydney, NSW, Australia., March 14, 2011.

Borwein:2011:LPHb

[Bor11y] Jonathan M. Borwein. Life of pi: History and computation — a talk for pi day. Interviews with ABC Sydney, Goldcoast and Tasmania/Victoria, Australia., March 14, 2011.

Borwein:2011:LPHc

[Bor11z] Jonathan M. Borwein. Life of pi: History and computation — a talk for pi day. AMSI AGR Talk for PiDay in America., March 15, 2011.

Borwein:2011:MWIa

[Bor11-27] Jonathan M. Borwein. Measures, walks and integrals: a study in hybrid computation. First Plenary Lecture, Numeric-Symbolic Computation (2011 FCCAR Meetings), San Jose, CA, USA., June 7, 2011.

Borwein:2011:MWIb

[Bor11-28] Jonathan M. Borwein. Measures, walks and integrals: a study in hybrid computation. Keynote Lecture, AustMS Special Session on Computational Mathematics (presented by James Wang) 2011 AustMS Meetings, Wollongong, NSW, Australia, Sept 26– 29., September 27, 2011.

Borwein:2011:MCA

[Bor11-29] Jonathan M. Borwein. Meetings with computer algebra and special functions: A Ramanujan style talk. Plenary Lecture at Jon-Fest Down Under, CARMA–AMSI Workshop, Nov 29–Dec 1, Newcastle, NSW, Australia., November 29, 2011.

Borwein:2011:MRGb

[Bor11-30] Jonathan M. Borwein. Moments of Ramanujan’s generalized elliptic integrals and extensions of Catalan’s constant. Number Theory Seminar, Mathematics Dept, University of Florida., Jan-

uary 11, 2011. URL http://docserver.carma.newcastle.edu.
au/1443/.

Borwein:2011:MMU

[Bor11-31] Jonathan M. Borwein. Mysteries of the mathematical universe. World Science Festival Panel, New York. Panel shared with Devlin, de Sautoy, Singh., June 4, 2011.

Borwein:2011:RAC

[Bor11-32] Jonathan M. Borwein. Ramanujan’s AG continued fraction. Colloquium, Mathematics Dept, University of Florida, Gainesville, FL, USA., January 11, 2011.

Borwein:2011:SWR

[Bor11-33] Jonathan M. Borwein. Short walks and ramble integrals: The arithmetic of uniform random walks. AMS Special Session on Special Functions, Combined Membership Meetings, New Orleans, LA, USA., January 9, 2011.

Borwein:2011:SWMa

[Bor11-34] Jonathan M. Borwein. Short walks, Mahler measures and logsine integrals, I. CARMA Seminar., March 16, 2011.

Borwein:2011:SWMb

[Bor11-35] Jonathan M. Borwein. Short walks, Mahler measures and logsine integrals, II. CARMA Seminar., March 23, 2011.

Borwein:2011:SS

[Bor11-36] Jonathan M. Borwein. A sinc that sank. CARMA Analysis Seminar, Newcastle, NSW, Australia., September 6, 2011. URL

http://docserver.carma.newcastle.edu.au/1391/.

Borwein:2011:TRT

[Bor11-37] Jonathan M. Borwein. Teaching and researching at the tertiary level with collaboration tools. Plenary Lecture, ALTC Workshop, 2011 AustMS Meetings, Wollongong, NSW, Australia, Sept 29– 30., September 29, 2011.

Borwein:2011:BRI

[Bor11-38] Jonathan Michael Borwein. Book review: Implicit functions and solution mappings, by Asen L. Dontchev and R. Tyrrell Rockafellar, Springer, 2009, ISBN 10: 0-387-87820-3, ISBN 13: 978-0-387-87820-1. IEEE Control Systems Magazine, 31(1):74–77, 2011. CODEN ISMAD7. ISSN 1066-033X (print), 1941-000X (electronic).

Borwein:2012:ARM

[Bor12a] Jonathan M. Borwein. Ann Romney and my brother. Math

Drudge, August 31, 2012. URL https://experimentalmath.
info/blog/2012/08/ann-romney-and-my-brother/.

Borwein:2012:AAS

[Bor12b] Jonathan M. Borwein. Arithmetic aspects of short random walks. Experimental mathematics seminar, University of Melbourne, Melbourne, VIC, Australia., May 14, 2012.

Borwein:2012:CMa

[Bor12c] Jonathan M. Borwein. CARMA and me: 2012. Annual CARMA Retreat, Fort Scratchley Newcastle, NSW, Australia., August 18, 2012.

Borwein:2012:CMb

[Bor12d] Jonathan M. Borwein. CARMA and me: 2012. APEC-ITB Workshop, Bandung Indonesia, given from Newcastle, NSW, Australia., October 23, 2012.

Borwein:2012:CTMa

[Bor12e] Jonathan M. Borwein. Computation and theory of Mordell– Witten–Tornheim sums. Part I theory. CARMA Analysis seminar, Newcastle, NSW, Australia., May 1, 2012.

Borwein:2012:CTMb

[Bor12f] Jonathan M. Borwein. Computation and theory of Mordell– Witten–Tornheim sums. Part II computation. CARMA Analysis seminar, Newcastle, NSW, Australia. 9. May 14., May 8, 2012.

Borwein:2012:EFSa

[Bor12g] Jonathan M. Borwein. Expectations over fractal sets. Applied Mathematics Seminar, University of New South Wales, Sydney, NSW, Australia., August 2, 2012.

Borwein:2012:EFSb

[Bor12h] Jonathan M. Borwein. Expectations over fractal sets. SigmaOPT Colloquium, CARMA., August 12, 2012.

Borwein:2012:EEM

[Bor12i] Jonathan M. Borwein. Exploratory experimentation in mathematics. ICERM Workshop on Reproducibility in Computational and Experimental Mathematics, ICERM, Providence, RI, USA, December 10–14., December 10, 2012.

Borwein:2012:FYMa

[Bor12j] Jonathan M. Borwein. Fifty years of maximal monotonicity: recent results on maximal monotone operators. Conference Presentation to the Workshop on Infinite Products of Operators and Their Applications, Technion, Haifa, May 21–24., May 24, 2012.

Borwein:2012:FYMb

[Bor12k] Jonathan M. Borwein. Fifty years of maximal monotonicity: recent results on maximal monotone operators. Keynote presentation to special session on Variational Analysis 56th AustMS Meetings, Ballarat, VIC, Australia, Sept 23–27., September 24, 2012.

Borwein:2012:FE

[Bor12l] Jonathan M. Borwein. Fractals everywhere. Presentation to West Wallsend High School students in CARMA., September 7, 2012.

Borwein:2012:FM

[Bor12m] Jonathan M. Borwein. The future of maths. Interview with ABC Radio National Future tense (Sunday 11.30)., Novem-

ber 11, 2012. URL http://www.abc.net.au/radionational/
programs/futuretense/the-future-of-maths/4355778..

Borwein:2012:IIC

[Bor12n] Jonathan M. Borwein. Interdisciplinarity, innovation, collaboration and creativity or how to manage a research portfolio. CARMA Colloquium., September 13, 2012.

Borwein:2012:LPA

[Bor12o] Jonathan M. Borwein. The life of pi: From Archimedes to ENIAC and beyond. Report, Centre for Computer Assisted Research Mathematics and its Applications (CARMA), University of Newcastle, Callaghan, NSW 2308, Australia, June 19, 2012. 30

pp. URL http://carma.newcastle.edu.au/jon/pi-2012.pdf.

Prepared for Berggren Festschrift. Updated and revised version of [Bor08a].

Borwein:2012:MEF

[Bor12p] Jonathan M. Borwein. Maximum entropy and feasibility methods for convex and nonconvex inverse problems. Optimization, 61(1): 1–33, 2012. CODEN OPTZDQ. ISSN 0233-1934, 0323-3898. URL

http://docserver.carma.newcastle.edu.au/1439/.

Borwein:2012:MEP

[Bor12q] Jonathan M. Borwein. Maximum entropy and projection methods for inverse problems. Technion Mathematics Colloquium and Conference Presentation to the Workshop on Infinite Products of Operators and Their Applications, Technion, Haifa, May 21–24., May 21, 2012.

Borwein:2012:MTW

[Bor12r] Jonathan M. Borwein. Mordell–Tornheim–Witten sums and log gamma integrals. Special session on Number Theory 56th AustMS Meetings, Ballarat, VIC, Australia, Sept 23–27., September 25, 2012.

Borwein:2012:MM

[Bor12s] Jonathan M. Borwein. Music and mathematics. Concert including a visualisation and sonification of a random walk through Pi with Jon Drummond and Fran Aragon. Harold Lobb Concert Hall, Newcastle, NSW, Australia., October 25, 2012.

Borwein:2012:MES

[Bor12t] Jonathan M. Borwein. My experiences with special functions. Colloquium, University of Melbourne, Melbourne, VIC, Australia., May 15, 2012.

Borwein:2012:PDAa

[Bor12u] Jonathan M. Borwein. Pi day in America. Huffington Post, ??(??):

??, March 15, 2012. URL http://www.huffingtonpost.com/
jonathan-m-borwein/.

Borwein:2012:PDAb

[Bor12v] Jonathan M. Borwein. Pi day in America. Conference in Honour of Alf van der Poorten, CARMA, Newcastle, NSW, Australia., March 15, 2012.

Borwein:2012:PDI

[Bor12w] Jonathan M. Borwein. Pi day interview. ABC Radio Queensland, Australia., March 14, 2012.

Borwein:2012:RPS

[Bor12x] Jonathan M. Borwein. Ramanujan and Pi: Srinivasa Ramanujan: Going strong at 125. Notices of the American Mathematical Society, 59(11):1534–1537, December 2012. CODEN AM-NOAN. ISSN 0002-9920 (print), 1088-9477 (electronic). URL

http://docserver.carma.newcastle.edu.au/1379/;http://
www.ams.org/journals/notices/201211/rtx121101522p.pdf.

Borwein:2012:RPMb

[Bor12y] Jonathan M. Borwein. Recent progress on maximal monotonicity. ANZIAM 2012, Warrnambool, VIC, Australia., January 31, 2012.

Borwein:2012:STMb

[Bor12z] Jonathan M. Borwein. Seeing things in mathematics. CARMA Workshop on Effective Visualisation in the Mathematical Sciences (eViMS)., November 25, 2012.

Borwein:2012:SM

[Bor12-27] Jonathan M. Borwein. Smart meters. Interview with ABC Radio Canberra, ACT, Australia., November 8, 2012.

Borwein:2012:TMP

[Bor12-28] Jonathan M. Borwein. Talking to, with and for the media and the public. CARMA-MAPS Outreach afternoon., August 17, 2012.

Borwein:2012:U

[Bor12-29] Jonathan M. Borwein. [untitled]. Interview with Ginger Gorman at 666 ABC Canberra, ACT, Australia on Pi Walk., August 11, 2012.

Borwein:2012:USTa

[Bor12-30] Jonathan M. Borwein. The use of selection theorems in optimization. Part I theory. CARMA/SigmaOPT seminar, Newcastle, NSW, Australia., March 22, 2012.

Borwein:2012:USTb

[Bor12-31] Jonathan M. Borwein. The use of selection theorems in optimization. Part II applications. CARMA/SigmaOPT seminar, Newcastle, NSW, Australia., March 29, 2012.

Borwein:2012:WMI

[Bor12-32] Jonathan M. Borwein. Walks, measures and integrals. Conference in Honour of Alf van der Poorten, CARMA, Newcastle, NSW, Australia., March 16, 2012.

Borwein:2012:EED

[Bor12-33] Jonathan Michael Borwein. Exploratory experimentation: Digitally-assisted discovery and proof. In Hanna and de Villiers [Hd12], pages 69–96. ISBN 94-007-2128-5, 94-007-2129-3 (ebook). LCCN QA9.54 .P766 2012. URL http://docserver.

carma.newcastle.edu.au/393/;http://link.springer.com/
chapter/10.1007/978-94-007-2129-6_4.

Borwein:2013:ANFb

[Bor13a] J. M. Borwein. Australia needs fundamental research to build a great country. The Conversation, ??(??):??, November 11, 2013.

URL https://theconversation.com/australia-needs-fundamental-research-to-build-a-great-country-20031.

Borwein:2013:C

[Bor13b] J. M. Borwein. Carma. One of a small selected number of researchers featured in an issue with a focus on “Mathematics, Statistics, Numerical and Computational Mathematics,” International Innovation., July 2013. URL http://www.

international-innovation-northamerica.com.

Borwein:2013:DLP

[Bor13c] J. M. Borwein. Don’t let politics drive research goals. Newcastle Herald, page 11, November 25, 2013. URL http://

www.theherald.com.au/story/1931600/opinion;http://www.
theherald.com.au/story/1931600/opinion-dont-let-politics-drive-research-goals/
;https://theconversation.edu.au/profiles/jon-borwein-101.

Borwein:2013:ANFa

[Bor13d] Jonathan M. Borwein. Australia needs fundamental research to build a great country. Math Drudge, November 11,

2013. URL https://experimentalmath.info/blog/2013/11/
australia-needs-fundamental-research-to-build-a-great-country/.

Borwein:2013:BABa

[Bor13e] Jonathan M. Borwein. Best approximation in Banach space: the Chebyshev problem. Naresuan University, Workshop on Nonsmooth Variational Inequalities, Optimization Problems and Fixed Point Theory, April 24–26, Naresuan University, Phitsanulok, Thailand., April 26, 2013.

Borwein:2013:BABb

[Bor13f] Jonathan M. Borwein. Best approximation in Banach space: the Chebyshev problem. Colloquium, Department of Mathematics, Chiang Mai University, Chiang Mai, Thailand., April 29, 2013.

Borwein:2013:BARa

[Bor13g] Jonathan M. Borwein. Best approximation in (reflexive) Banach space, I: Introduction. OANTS-AMSI seminar, CARMA., March 25, 2013.

Borwein:2013:BARb

[Bor13h] Jonathan M. Borwein. Best approximation in (reflexive) Banach space, II: The Lau–Konjagin Theorem. OANTS-AMSI seminar, CARMA., April 8, 2013.

Borwein:2013:BARc

[Bor13i] Jonathan M. Borwein. Best approximation in (reflexive) Banach space, III: The Chebysev problem. OANTS-AMSI seminar, CARMA., April 15, 2013.

Borwein:2013:DRM

[Bor13j] Jonathan M. Borwein. Douglas–Rachford methods for matrix completion problems. CARMA/OCANA Seminar, University of Newcastle, Newcastle, NSW, Australia., October 23, 2013.

Borwein:2013:EMC

[Bor13k] Jonathan M. Borwein. Entropy methods and checkerboard copulas to simulate rainfall. Department of Economics, Chiang Mai University, Chiang Mai, Thailand., April 30, 2013.

Borwein:2013:EMI

[Bor13l] Jonathan M. Borwein. Entropy methods for inverse problems. Colloquium, Department of Economics, Chiang Mai University, Chiang Mai, Thailand., April 30, 2013.

Borwein:2013:EEMa

[Bor13m] Jonathan M. Borwein. Exploratory experimentation in the mathematical sciences. Colloquium, Department of Economics, Chiang Mai University, Chiang Mai, Thailand., April 29, 2013.

Borwein:2013:EEMb

[Bor13n] Jonathan M. Borwein. Exploratory experimentation in the mathematical sciences. National Mathematics Seminar, Bulgarian Academy of Science., June 5, 2013.

Borwein:2013:LP

[Bor13o] Jonathan M. Borwein. The life of pi. Maths Enrichment Session, CARMA., September 20, 2013.

Borwein:2013:MEP

[Bor13p] Jonathan M. Borwein. Maximum entropy and projection methods for convex and non-convex inverse problems. First Keynote Lecture, MaxEnt33, Canberra, ACT, Australia (Dec 15–20)., December 16, 2013.

Borwein:2013:MSS

[Bor13q] Jonathan M. Borwein. Modelling and simulation of seasonal rainfall. Recent Advances in OR, RMIT AGR–University of Newcastle AGR–Adelaide, July 17., July 17, 2013. URL http:

//docserver.carma.newcastle.edu.au/1453/.

Borwein:2013:NDR

[Bor13r] Jonathan M. Borwein. Nonconvex Douglas–Rachford iterations. Workshop on Nonsmooth Variational Inequalities, Optimization Problems and Fixed Point Theory, April 24–26, Naresuan University, Phitsanulok, Thailand., April 26, 2013.

Borwein:2013:PD

[Bor13s] Jonathan M. Borwein. Pi day. Interview with ABC radio South Australia., March 14, 2013.

Borwein:2013:PPE

[Bor13t] Jonathan M. Borwein. The pi of planet earth. Plenary Lecture, AMSI AGR Distinguished Lecture Series from CARMA., March 14, 2013.

Borwein:2013:STWc

[Bor13u] Jonathan M. Borwein. Seeing things by walking on numbers. Workshop on Nonsmooth Variational Inequalities, Optimization Problems and Fixed Point Theory, April 24–26, Naresuan University, Phitsanulok, Thailand., April 26, 2013.

Borwein:2013:STWd

[Bor13v] Jonathan M. Borwein. Seeing things by walking on numbers. Colloquium, Department of Mathematics, Chiang Mai University, Chiang Mai, Thailand., April 30, 2013.

Borwein:2013:STWe

[Bor13w] Jonathan M. Borwein. Seeing things by walking on numbers. First keynote, workshop on Topological Methods in Analysis and Optimization, Bulgarian Academy of Science, June 10–13., June 10, 2013.

Borwein:2013:STWf

[Bor13x] Jonathan M. Borwein. Seeing things by walking on numbers. National Youth Science Forum students visit to University of Newcastle, NSW, Australia, July 1–2, CARMA., July 2, 2013.

Borwein:2013:STWg

[Bor13y] Jonathan M. Borwein. Seeing things by walking on numbers. Maths Enrichment Session, CARMA., August 14, 2013.

Borwein:2013:STWh

[Bor13z] Jonathan M. Borwein. Seeing things by walking on numbers. Number Theory Session, 57th AustMS Meeting, University of Sydney, NSW, Australia., September 30, 2013.

Borwein:2013:STWi

[Bor13-27] Jonathan M. Borwein. Seeing things by walking on numbers. Google’s CS4HS programme, University of Newcastle, Newcastle, NSW, Australia., November 12, 2013.

Borwein:2013:STWa

[Bor13-28] Jonathan M. Borwein. Seeing things by walking on real numbers. Fellows’ Lecture, 2013 National Mathematics Summer School, Shine Dome, Canberra, ACT, Australia., January 17, 2013.

Borwein:2013:STWb

[Bor13-29] Jonathan M. Borwein. Seeing things by walking on real numbers. Plenary Lecture, Third South Pacific Optimization Meeting, Newcastle, NSW, Australia., February 11, 2013.

Borwein:2013:SDR

[Bor13-30] Jonathan M. Borwein. Set the default to ‘reproducible’. Session on reproducible computational science, MPE 2013, Melbourne, VIC, Australia, July 8–12., July 9, 2013.

Borwein:2013:SLV

[Bor13-31] Jonathan M. Borwein. Seven lectures on variational analysis. CIMPA–Unesco–India Research School on Generalized Nash Equilibrium Problems, Delhi University, New Delhi, India (Nov 25th-Dec 6)., November 25–30, 2013.

Borwein:2013:SM

[Bor13-32] Jonathan M. Borwein. Surprise maximization. Department of Economics, Chiang Mai University, Chiang Mai, Thailand., April

30, 2013. URL http://docserver.carma.newcastle.edu.au/
209/.

Borwein:2013:VAPc

[Bor13-33] Jonathan M. Borwein. Variational analysis in the presence of symmetry. Optimization of Planet Earth Session, 57th AustMS Meeting, University of Sydney, Sydney, NSW, Australia., October 1, 2013.

Borwein:2013:VAPa

[Bor13-34] Jonathan M. Borwein. Variational analysis in the presence of symmetry. Part I. OAANTS AGR Seminar., July 30, 2013.

Borwein:2013:VAPb

[Bor13-35] Jonathan M. Borwein. Variational analysis in the presence of symmetry. Part II. OAANTS AGR Seminar., August 6, 2013.

Borwein:2014:BTM

[Bor14a] J. M. Borwein. Budget 2014: there’s more to science than medical research. The Conversation, ??(??):??, May 13, 2014. URL

https://theconversation.com/budget-2014-theres-more-to-science-than-medical-research-26632.

Borwein:2014:MWF

[Bor14b] J. M. Borwein. Meet the winners of the Fields Medal: the ‘Nobel Prize of maths’. The Conversation, ??(??):

??, August 12, 2014. URL https://theconversation.com/
meet-the-winners-of-the-fields-medal-the-nobel-prize-of-maths-30411.

Borwein:2014:MIP

[Bor14c] Jonathan M. Borwein. The Mathematical Investor: A personal perspective by JMB. Mathematical Investor, January 7,

2014. URL https://www.financial-math.org/blog/2014/01/
the-mathematical-investor-a-personal-perspective-by-jmb/.

Borwein:2014:CM

[Bor14d] Jonathan M. Borwein. CARMA and me. Opening of CRM, University of Western Sydney, NSW, Australia., May 28, 2014.

Borwein:2014:CPT

[Bor14e] Jonathan M. Borwein. Character polylogarithms and their applications. First lecture, Number Theory Down Under II, Newcastle, October 24–25, 2014., October 24, 2014.

Borwein:2014:DRMa

[Bor14f] Jonathan M. Borwein. Douglas–Rachford methods for matrix completion problems. ANZIAM 2014, Rotorua, New Zealand., February 6, 2014.

Borwein:2014:DRMb

[Bor14g] Jonathan M. Borwein. Douglas–Rachford methods for matrix completion problems. ANZIAM 2014, Federation University meeting in honour of Vladimir Demyanov (given from Burnaby)., April 16, 2014.

Borwein:2014:ECVa

[Bor14h] Jonathan M. Borwein. Experimental computation and visual theorems. In Hong and Yap [HY14], pages 1–8. ISBN 3-662-44198-5, 3-662-44199-3 (e-book). LCCN QA76.95.

Borwein:2014:ECVb

[Bor14i] Jonathan M. Borwein. Exploratory computation and visual theorems: The computer as collaborator. Part I, CARMA Colloquium., May 21, 2014.

Borwein:2014:ECVc

[Bor14j] Jonathan M. Borwein. Exploratory computation and visual theorems: The computer as collaborator. Part II, CARMA Colloquium., May 29, 2014.

Borwein:2014:ECVd

[Bor14k] Jonathan M. Borwein. Exploratory computation and visual theorems: The computer as collaborator. Part III, CARMA Colloquium., June 14, 2014.

Borwein:2014:ECVe

[Bor14l] Jonathan M. Borwein. Exploratory computation and visual theorems: The computer as collaborator. First Plenary, ICERM Workshop on Challenges for 21st Century Experimental Mathematics, Brown University, Providence, RI, USA, July 21–25, 2014., July 21, 2014.

Borwein:2014:ECVf

[Bor14m] Jonathan M. Borwein. Exploratory computation and visual theorems: The computer as collaborator. Final Keynote, Fourth International Congress on Mathematical Software (and ICM Satellite Meeting), Hanyang University, Seoul, Souh Korea August 5–8, 2014., August 9, 2014.

Borwein:2014:ECVg

[Bor14n] Jonathan M. Borwein. Exploratory computation and visual theorems: The computer as collaborator. Plenary lecture Effective Visualisation in the Mathematical Sciences (EViMS2), Australian National University, November 21–23., November 22, 2014.

Borwein:2014:FFC

[Bor14o] Jonathan M. Borwein. The Fitzpatrick function as a convex gap function. Special session on Optimization, ANZMC8, Melbourne, VIC, Australia, December 8–12, 2014., December 8, 2014.

Borwein:2014:LAE

[Bor14p] Jonathan M. Borwein. The life of π: From Archimedes to ENIAC and beyond. In Sidoli and Van Brummelen [SV14], pages 531–561. ISBN 3-642-36735-6 (hardcover), 3-642-36736-4 (e-book). LCCN

QA21 .F76 2014. URL http://www.carma.newcastle.edu.au/

~jb616/pi-2010.pdf. Extended and updated version of [Bor08a].

Borwein:2014:LPb

[Bor14q] Jonathan M. Borwein. The life of pi. A Talk for Pi Day or other Days. Part I, History. Inaugural Möbius Lecture Series, Department of Mathematics, Baylor University, Waco, TX, USA., April 22, 2014.

Borwein:2014:LPc

[Bor14r] Jonathan M. Borwein. The life of pi. A Talk for Pi Day or other Days. Part II, Computation. Colloquium, Inaugural Möbius Lecture Series, Department of Mathematics, Baylor University, Waco, Texas., April 23, 2014.

Borwein:2014:LPA

[Bor14s] Jonathan M. Borwein. The life of pi: From Archimedes to ENIAC and beyond. In Sidoli and Van Brummelen [SV14], pages 531–561. ISBN 3-642-36735-6 (hardcover), 3-642-36736-4 (e-book). LCCN

QA21 .F76 2014. URL http://docserver.carma.newcastle.
edu.au/265/.

Borwein:2014:MDS

[Bor14t] Jonathan M. Borwein. Moments and densities of short walks in arbitrary dimensions. Plenary lecture New Directions in Fractal Geometry, Australian National University, November 23–27., November 25, 2014.

Borwein:2014:PD

[Bor14u] Jonathan M. Borwein. Pi day. Interview and call-in show with ABC radio South Australia., March 14, 2014.

Borwein:2014:RPSa

[Bor14v] Jonathan M. Borwein. Rock, paper, scissors. Interview on ABC Radio Newcastle, NSW, Australia., May 8, 2014.

Borwein:2014:RPSb

[Bor14w] Jonathan M. Borwein. Rock, paper, scissors. Interview on ABC Radio Riverena., May 12, 2014.

Borwein:2014:STMb

[Bor14x] Jonathan M. Borwein. Seeing things in mathematics by walking on real numbers. Workshop in honour of David Borwein’s 90th birthday, IRMACS, Burnaby, BC, Canada., April 16, 2014.

Borwein:2014:STMc

[Bor14y] Jonathan M. Borwein. Seeing things in mathematics by walking on real numbers. Inaugural Möbius Lecture Series, Colloquium, Department of Mathematics, Baylor University, Waco, Texas., April 24, 2014.

Borwein:2014:VCP

[Bor14z] Jonathan M. Borwein. A very complicated proof of the minimax theorem. CARMA Workshop on Optimisation and Risk, July 12, 2012., July 12, 2014. URL http://docserver.carma.

newcastle.edu.au/1499/.

Borwein:2014:VPP

[Bor14-27] Jonathan M. Borwein. The visualization of pi. Presentation at Destination Maitland: City of the Future. Conservatorium Campus, Maitland, NSW, Australia., April 10, 2014.

Borwein:2015:OPB

[Bor15a] J. M. Borwein. It’s often the puzzles that baffle that go viral. The Conversation, ??(??):??, April 16, 2015. URL https://

theconversation.com/its-often-the-puzzles-that-baffle-that-go-viral-40216.

Borwein:2015:MWK

[Bor15b] J. M. Borwein. The man who knew infinity: a mathematician’s life comes to the movies. The Conversation, ??(??):

??, November 17, 2015. URL https://theconversation.com/
the-man-who-knew-infinity-a-mathematicians-life-comes-to-the-movies-50777.

Borwein:2015:TWC

[Bor15c] J. M. Borwein. The ‘train wreck’ continues: another social science retraction. The Conversation, ??(??):??, May 29, 2015. URL

https://theconversation.com/the-train-wreck-continues-another-social-science-retraction-42404.

Borwein:2015:AOF

[Bor15d] Jonathan M. Borwein. Adventures with the OEIS: Five sequences Tony may like. Tony Guttmann: Seventy and counting, December 7–8, Newcastle, NSW, Australia., December 7, 2015.

Borwein:2015:BMY

[Bor15e] Jonathan M. Borwein. Brailey and me (and you). Dinner speech, CARMA Workshop in honour of Brailey Sims, August 21–23, Newcastle, NSW, Australia., August 22, 2015.

Borwein:2015:CAGb

[Bor15f] Jonathan M. Borwein. Convex analysis on groups and monoids. Plenary Variational Analysis and Optimisation Special Session, AustMS 59, Flinders University, Bedford Park SA 5042, Australia, 28 Sept–Oct 1, 2015., September 29, 2015.

Borwein:2015:DRM

[Bor15g] Jonathan M. Borwein. Douglas–Rachford methods for matrix completion problems. Federation University Workshop on Continuous Optimization in honour of Alex Rubinov (given from Newcastle)., April 16, 2015.

Borwein:2015:ECV

[Bor15h] Jonathan M. Borwein. Exploratory computation and visual theorems: The computer as collaborator. Parts I and II. Opening general lectures, AMSI Summer School Newcastle, January 5–26., January 8, 2015.

Borwein:2015:FFC

[Bor15i] Jonathan M. Borwein. The Fitzpatrick function as a convex gap function. Special session in memory of Simon Fitzpatrick, South Pacific Continuous Optimization Meeting, Adelaide, SA, Australia, February 8–12, 2015., February 10, 2015.

Borwein:2015:FM

[Bor15j] Jonathan M. Borwein. The future of mathematics: 1965 to 2065. In Kennedy et al. [KAA^{+}15], pages 313–329. ISBN 0-88385-588-7 (hardcover), 1-61444-522-2 (e-book). LCCN QA26 .C46 2015.

URL http://docserver.carma.newcastle.edu.au/1720/.

Borwein:2015:PP

[Bor15k] Jonathan M. Borwein. I prefer pi. CARMA Colloquium., March 12, 2015.

Borwein:2015:MLY

[Bor15l] Jonathan M. Borwein. Moore’s Law is 50 years old. Interview with ABC Radio Melbourne, VIC, Australia., July 22, 2015.

Borwein:2015:SDR

[Bor15m] Jonathan M. Borwein. Set the default to reproducible. CARMA workshop on Mathematical Aspects of Behavioural Economics and Finance, November 13–14, Newcastle, NSW, Australia., November 13, 2015.

Borwein:2015:SWC

[Bor15n] Jonathan M. Borwein. A short walk can be beautiful. Maths Education Special Session, AustMS 59, Flinders University, Bedford Park SA 5042, Australia, 28 Sept–Oct 1, 2015., October 1, 2015.

URL http://docserver.carma.newcastle.edu.au/1712/.

Borwein:2015:SWAa

[Bor15o] Jonathan M. Borwein. Short walks in arbitrary dimensions. Final lecture Workshop on Mathematics and Computation, CARMA 21–23 June, 2015., June 23, 2015.

Borwein:2015:SWAb

[Bor15p] Jonathan M. Borwein. Short walks in arbitrary dimensions. ASEMS Workshop on Stochastics and Special Functions, University of Melbourne, Melbourne, VIC, Australia, 13–14 August, 2015., August 13, 2015.

Borwein:2015:SWAc

[Bor15q] Jonathan M. Borwein. Short walks in arbitrary dimensions. Number Theory Down Under 3, CARMA, University of Newcastle, 18–19 September, 2015., September 18, 2015.

Borwein:2015:TLD

[Bor15r] Jonathan M. Borwein. Two lectures on Douglas–Rachford reflection methods for convex and non convex feasibility problems. Keynote lectures, RMIT Workshop on Optimisation, August 11, 2013., August 11, 2015.

Borwein:2015:U

[Bor15s] Jonathan M. Borwein. [untitled]. NZ Radio interview on the Melbourne Cup, Newstalk ZB., November 3, 2015.

Borwein:2015:WWH

[Bor15t] Jonathan M. Borwein. Who we are and how we got that way? In Casazza et al. [CKR15], pages 140–156. ISBN 0-88385-585-2 (print), 1-61444-521-4 (e-book). LCCN QA28 .I22 2015. URL

http://docserver.carma.newcastle.edu.au/777/.

Borwein:2016:AO

[Bor16a] J. M. Borwein. Adventures with the OEIS. The Ramanujan Journal, ??(??):??, ???? 2016. ISSN 1382-4090 (print), 1572-

9303 (electronic). URL http://docserver.carma.newcastle.
edu.au/1686/.

Borwein:2016:PPM

[Bor16b] J. M. Borwein. Pi and its part in the most beautiful formula in mathematics. The Conversation, ??(??):

??, March 13, 2016. URL https://theconversation.com/
pi-and-its-part-in-the-most-beautiful-formula-in-mathematics-56067.

Borwein:2016:PD

[Bor16c] J. M. Borwein. Pi day 2016. Rhodes Scholar Blog, ??(??):??,

April 8, 2016. URL http://www.rhodeshouse.ox.ac.uk/blog/
pi-day-2016.

Borwein:2016:RMW

[Bor16d] J. M. Borwein. Ramanujan – the man who knew infinity. Rhodes Scholar Blog, ??(??):??, February 23, 2016. URL http://www.

rhodeshouse.ox.ac.uk/blog/ramanujan-the-man-who-knew-infinity.

Borwein:2016:SWC

[Bor16e] J. M. Borwein. A short walk can be beautiful. Journal of Humanistic Mathematics, 6(1):5–28, January 2016. ISSN 2159-8118.

URL http://docserver.carma.newcastle.edu.au/1712/.

Borwein:2016:EMP

[Bor16f] Jonathan Borwein. The experimental mathematician: The pleasure of discovery and the role of proof. Canadian Mathematical Society Notes, 48(6):10–13, December 2016. ISSN 0045-

5164. URL http://cms.math.ca/notes/v48/n6/Notesv48n6.

pdf. Excerpts from a 2002 CMESG Plenary Lecture given by the author.

Borwein:2016:EM

[Bor16g] Jonathan Borwein. Experimental mathematics. In The Human Face of Computing [Cal16], pages 141–156. ISBN 1-78326-643-0 (hardcover), 1-78326-645-7 (e-book). LCCN QA76.9.C66. URL

http://docserver.carma.newcastle.edu.au/1527/.

Borwein:2016:CMM

[Bor16h] Jonathan M. Borwein. Carma: a model for multi-disciplinary and multi-institutional collaborative research. Colloquium, Florida ADRC (Alzheimer’s Disease Research Center), McKnight Brain Institute, University of Florida, Gainesville, FL, USA., May 17, 2016.

Borwein:2016:CAGb

[Bor16i] Jonathan M. Borwein. Convex analysis on groups and monoids. Keynote lecture, South Pacific Optimization Meeting V, University of New Caledonia, Nouméa 98851, New Caledonia., February 29, 2016.

Borwein:2016:CAGc

[Bor16j] Jonathan M. Borwein. Convex analysis on groups and semigroups. Plenary lecture, Fourth International Workshop on Functional Analysis, University of Cartagena, June 9–10., June 9, 2016.

Borwein:2016:CAGd

[Bor16k] Jonathan M. Borwein. Convex analysis on groups and semigroups. Plenary lecture, Mathematical Optimisation Down Under (MODU2016), Melbourne, Melbourne, VIC, Australia, July 18– 22. (Given from London, ON, Canada, on 21 July 2016)., July 20,

2016. URL https://www.carma.newcastle.edu.au/jon/cogs.
pdf.

Borwein:2016:LFA

[Bor16l] Jonathan M. Borwein. The Lambert W function in analysis and optimization. Keynote lecture, Fields Workshop on the 20th anniversary of the Lambert W function, July 25–28, Western University, London, ON, Canada., July 25, 2016.

Borwein:2016:LFOa

[Bor16m] Jonathan M. Borwein. The Lambert W function in optimization. Third Keynote lecture, OVA7, in honour of Michel Théra at 70, June 1–3, Alicante, Spain., June 3, 2016.

Borwein:2016:LHC

[Bor16n] Jonathan M. Borwein. The life of π: History and computation: A talk for pi day or other days. Nerenberg Memorial Lecture, University of Western Ontario, London, ON, Canada.,

April 5, 2016. URL https://www.carma.newcastle.edu.au/
jon/piday-16-sm.pdf.

Borwein:2016:LPA

[Bor16o] Jonathan M. Borwein. The life of pi: From Archimedes to ENIAC and beyond. In Bailey and Borwein [BB16l], pages 443–474. ISBN 3-319-32375-X, 3-319-32377-6 (e-book). LCCN QA251. URL

http://docserver.carma.newcastle.edu.au/265/.

Borwein:2016:RMM

[Bor16p] Jonathan M. Borwein. Reflection methods for matrix completion. Tutte Seminar, University of Waterloo, Waterloo, ON, Canada., May 6, 2016.

Borwein:2016:STWa

[Bor16q] Jonathan M. Borwein. Seeing things by walking on real numbers. Mathematics Colloquium, University of Florida, Gainesville, FL, USA., May 19, 2016.

Borwein:2016:STWb

[Bor16r] Jonathan M. Borwein. Seeing things by walking on real numbers. Mathematics Colloquium, Dalhousie University, Halifax, NS, Canada., June 30, 2016.

Borwein:2016:STWc

[Bor16s] Jonathan M. Borwein. Seeing things by walking on real numbers. Mathematics Seminar, University of Toronto, Toronto, ON, Canada., July 14, 2016.

Borwein:2016:TT

[Bor16t] Jonathan M. Borwein. Technical talk. 25th Owens Memorial Lecture, Wayne State University, Detroit, MI, USA., April 13, 2016.

Borwein:2016:TACa

[Bor16u] Jonathan M. Borwein. Theory and applications of convex and non-convex feasibility problems. Applied Mathematics Spring Lecture Series — University of Western Ontario, London, ON, Canada. Lecture 1., April 29, 2016.

Borwein:2016:TACb

[Bor16v] Jonathan M. Borwein. Theory and applications of convex and non-convex feasibility problems. Applied Mathematics Spring Lecture Series — University of Western Ontario, London, ON, Canada. Lecture 2., April 29, 2016.

Borwein:2016:TACc

[Bor16w] Jonathan M. Borwein. Theory and applications of convex and non-convex feasibility problems. Applied Mathematics Spring Lecture Series — University of Western Ontario, London, ON, Canada. Lecture 3., May 3, 2016.

Borwein:2016:TACd

[Bor16x] Jonathan M. Borwein. Theory and applications of convex and non-convex feasibility problems. Applied Mathematics Spring Lecture Series — University of Western Ontario, London, ON, Canada. Lecture 4., May 3, 2016.

Borwein:2016:TACe

[Bor16y] Jonathan M. Borwein. Theory and applications of convex and non-convex feasibility problems. Applied Mathematics Spring Lecture Series — University of Western Ontario, London, ON, Canada. Lecture 5., May 5, 2016.

Borwein:2016:VAP

[Bor16z] Jonathan M. Borwein. Variational analysis in the presence of symmetry. Informal Worshop on Nonlinear Optimization, Western, London, ON, Canada, June 24–25., June 25, 2016.

Borwein:2016:VCP

[Bor16-27] Jonathan M. Borwein. A very complicated proof of the minimax theorem. Minimax Theory and its Applications, 1(1):21–27, ???? 2016. ISSN 2199-1413 (print), 2199-1421 (electronic). URL

http://docserver.carma.newcastle.edu.au/1499/;http://
www.heldermann.de/MTA/MTA01/MTA011/mta01002.htm.

Borwein:2016:VTM

[Bor16-28] Jonathan M. Borwein. Visual theorems in mathematics. Fields Workshop on Computationally Assisted Mathematical Discovery and Experimental Mathematics, First Plenary Lecture, Western University, London, ON, Canada, May 12–15., May 12, 2016.

Borwein:2016:WNa

[Bor16-29] Jonathan M. Borwein. Walking on numbers. Technical Talk, Nerenberg Memorial Lecture, University of Western Ontario, London, ON, Canada., April 7, 2016.

Borwein:2016:WNb

[Bor16-30] Jonathan M. Borwein. Walking on numbers. 25th Owens Memorial Lecture, Wayne State University, Detroit, MI, USA., April 12, 2016.

Borwein:2016:WNc

[Bor16-31] Jonathan M. Borwein. Walking on numbers. Mathematics Department Colloquium, Western Michigan University, Kalamazoo, MI, USA., April 14, 2016.

Borwein:2017:AO

[Bor17a] Jonathan M. Borwein. Adventures with the OEIS. In George E. Andrews and Frank (Frank G.) Garvan, editors, Analytic Number Theory, Modular Forms and q-Hypergeometric Series: in honor of Krishna Alladi’s 60th birthday, University of Florida, Gainesville, March 2016, pages 123–138. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017. ISBN 3-319-68375-6 (print), 3-319-68376-4 (e-book). ISSN 2194-1009. LCCN QA241.

Borwein:2017:GEC

[Bor17b] Jonathan M. Borwein. Generalisations, examples, and counterexamples in analysis and optimisation: In honour of Michel Théra at 70. Set-Valued and Variational Analysis, 25(3): 467–479, September 2017. ISSN 1877-0533 (print), 1877-

0541 (electronic). URL http://docserver.carma.newcastle.
edu.au/1688/;http://link.springer.com/article/10.1007/

s11228-016-0379-2. Special issue for Michel Théra at 70. 7th International Seminar on Optimization and Variational Analysis (OVA), Universidad Alicante, Alicante, Spain, 1–3 June 2016.

Borwein:2020:I

[Bor20] Naomi Simone Borwein. Introduction. In Bailey et al. [BBB^{+}20], pages 101–102. ISBN 3-030-36567-0 (print), 3-030-36568-9 (ebook). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Bouza:2006:BRR

[Bou06] C. N. Bouza. Book reviews/reseñas: Convex Analysis and Nonlinear Optimization. Theory and Examples, 2nd edition, Jonathan Borwein and Lewis Adrian S. (2006), xii + 310 pages. CMS Books in Mathematics/Ouvrages de Mathemématiques de la SMC, ISBN-10 0-387-29570-4. ISBN-13 978-0-387-29570-1. Revista Investigación Operacional, 27(2):206–208, 2006. CO-DEN ???? ISSN 0257-4306 (print), 2224-5405 (electronic).

URL http://archives-web.univ-paris1.fr/rev-inv-ope/
fileadmin/rev-inv-ope/files/27206/IO_27206-BOOK_REVIEWS.
pdf.

Borwein:1987:SVPa

[BP87] Jonathan M. Borwein and David Preiss. A smooth variational principle with applications to subdifferentiability and to differentiability of convex functions. Transactions of the American Mathematical Society, 303(2):517–527, October 1987. CO-DEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
1597/;https://www.jstor.org/stable/2000681.

Borwein:1999:PSS

[BPB99] Jonathan M. Borwein, C. G. Pinner, and David Bradley. Problems and solutions: Solutions: An infinite product: 10605. American Mathematical Monthly, 106(2):173–174, 1999. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Borwein:1984:CCO

[BPT84] J. M. Borwein, J.-P. Penot, and M. Théra. Conjugate convex operators. Journal of Mathematical Analysis and Applications, 102(2):399–414, 1984. CODEN JMANAK. ISSN 0022-247X (print), 1096-0813 (electronic). URL http://docserver.

carma.newcastle.edu.au/1620/;http://www.sciencedirect.
com/science/article/pii/0022247X8490180X.

Borwein:1984:HMM

[BR84] J. M. Borwein and B. Richmond. How many matrices have roots? Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(2):286–299, 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL http:

//docserver.carma.newcastle.edu.au/1628/.

Borwein:19xx:WMS

[BRxx] J. M. Borwein and B. Richmond. When is a matrix a square? Research report 5, Department of Mathematics, Dalhousie University and Department of Combinatorics and Optimization, University of Waterloo, Halifax, NS, Canada and Waterloo, ON, Canada, 19xx. 22 pp.

Berndt:2001:RES

[BR01] Bruce C. Berndt and Robert A. (Robert Alexander) Rankin, editors. Ramanujan: essays and surveys, volume 22 of History of mathematics. American Mathematical Society, Providence, RI, USA, 2001. ISBN 0-8218-2624-7. ISSN 0899-2428. xvi + 347 pp. LCCN QA29.R3 R29 2001.

Borwein:2010:ICF

[BR10] Jonathan M. Borwein and Michael Rose. Introduction to CARMA and fractals. BOOST + on campus high school visit to University of Newcastle, Newcastle, NSW, Australia., October 19, 2010.

Borwein:2012:EWC

[BR12] Jonathan Borwein and Michael Rose. Explainer: what is chaos theory? The Conversation, ??(??):??, November 19, 2012. URL

https://theconversation.com/explainer-what-is-chaos-theory-10620.

Borwein:2013:TTS

[BR13a] J. M. Borwein and Michael Rose. Tipsy tottering, sunlight and the smell of coffee: its all random. The Conversation, ??(??):

??, December 19, 2013. URL https://theconversation.com/
tipsy-tottering-sunlight-and-the-smell-of-coffee-its-all-random-21598.

Borwein:2013:EWC

[BR13b] Jonathan Borwein and Michael Rose. Explainer: what is chaos theory. In The explainer: from déjà vu to why the sky is blue, and other conundrums [Tre13], page ?? ISBN 1-

4863-0050-2 (paperback). URL https://theconversation.com/
explainer-what-is-chaos-theory-10620.

Borwein:2014:ECC

[BR14a] J. M. Borwein and M. Rose. Explainer: clearing up confusion between correlation and causation. The Conversation, ??(??):??,

September 23, 2014. URL https://theconversation.edu.au/
profiles/jon-borwein-101.

Borwein:2014:HBW

[BR14b] J. M. Borwein and M. Rose. How betting works and why the Melbourne Cup skews the odds. The Conversation, ??(??):

??, October 30, 2014. URL https://theconversation.com/
how-betting-works-and-why-the-melbourne-cup-skews-the-odds-33357.

Borwein:2014:CCB

[BR14c] Jonathan Borwein and Michael Rose. Clearing up confusion between correlation and causation. The Conversation, ??(??):

??, September 22, 2014. URL https://theconversation.com/
clearing-up-confusion-between-correlation-and-causation-30761.

Borwein:2016:EAI

[BR16] Jonathan M. Borwein and Michael Rose. Expectations over attractors of iterated function systems. Submitted JMAA, June

2016., ???? 2016. URL http://docserver.carma.newcastle.
edu.au/1680/.

Brent:2017:JBP

[Bre17] Richard P. Brent. Jonathan Borwein, pi and the AGM. Talk slides, Australian National University and CARMA, University of Newcastle, Canberra, ACT and Newcastle, NSW, Australia, September 26, 2017. 76 pp. URL https:

//carma.newcastle.edu.au/meetings/jbcc/abstracts/pdf/
JBCC-Richard_Brent.pdf.

Brent:2020:BBP

[Bre20a] Richard P. Brent. The Borwein brothers, pi and the AGM. In Bailey et al. [BBB^{+}20], pages 323–347. ISBN 3-030-36567-0 (print), 3-030-36568-9 (e-book). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Brent:2020:I

[Bre20b] Richard P. Brent. Introduction. In Bailey et al. [BBB^{+}20], pages 299–302. ISBN 3-030-36567-0 (print), 3-030-36568-9 (e-book). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Borwein:1999:CSFa

[BRLZ99] Jonathan M. Borwein, John Read, Adrian S. Lewis, and Qiji Zhu. Convex spectral functions of compact operators. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, March 10, 1999. 27 pp. Published in [BRLZ00].

Borwein:2000:CSF

[BRLZ00] Jonathan M. Borwein, John Read, Adrian S. Lewis, and Qiji Zhu. Convex spectral functions of compact operators. Journal of Nonlinear and Convex Analysis, 1(1):17–35, 2000. ISSN 1345-4773 (print), 1880-5221 (electronic). URL http://www.ybook.co.jp/

online2/jncav1.html.

Borwein:2008:CMD

[BRR08] Jonathan M. Borwein, E. M. (Eugenio M.) Rocha, and José-Francisco Rodrigues, editors. Communicating Mathematics in the Digital Era. A. K. Peters, Ltd., Wellesley, MA, USA, 2008. ISBN 1-56881-410-0. xii + 325 pp. LCCN QA76.95 .C59 2008. URL

http://www.loc.gov/catdir/toc/fy0903/2008022183.html.

Borwein:1992:KMI

[BRS92] Jonathan Borwein, Simeon Reich, and Itai Shafrir. Krasnosel’ski– Mann iterations in normed spaces. Canadian mathematical bulletin = Bulletin canadien de mathématiques, 35(1):21–28, 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:2011:CBF

[BRS11] Jonathan M. Borwein, Simeon Reich, and Shoham Sabach. A characterization of Bregman firmly nonexpansive operators using a new monotonicity concept. Journal of Nonlinear and Convex Analysis, 12(1):161–184, 2011. ISSN 1345-4773 (print), 1880-

5221 (electronic). URL http://docserver.carma.newcastle.
edu.au/1445/.

Borwein:1983:NMB

[BS83] Jonathan Borwein and Brailey Sims. Nonexpansive mappings on Banach lattices. Comptes rendus de l’Académie des sciences. Canada, 5(1):21–26, 1983. ISSN 0706-1994. URL http:

//docserver.carma.newcastle.edu.au/1635/.

Borwein:1984:DLM

[BS84a] J. M. Borwein and H. M. Strójwas. Directionally Lipschitzian mappings on Baire spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 36(1):95–130, 1984. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL

http://docserver.carma.newcastle.edu.au/1627/.

Borwein:1984:NMB

[BS84b] Jon M. Borwein and Brailey Sims. Nonexpansive mappings on Banach lattices and related topics. Houston Journal of Mathematics, 10(3):339–356, 1984. CODEN HJMADZ. ISSN

0362-1588. URL https://www.math.uh.edu/~hjm/restricted/
archive/v010n3/0339BORWEIN.pdf.

Borwein:1985:TA

[BS85] J. M. Borwein and H. M. Strójwas. Tangential approximations. Nonlinear Analysis, Theory, Methods and Applications, 9(12):1347–1366, 1985. CODEN NOANDD. ISSN 0362-546x (print), 1873-5215 (electronic). URL http://docserver.

carma.newcastle.edu.au/1617/;http://www.sciencedirect.
com/science/article/pii/0362546X85900951.

Borwein:1986:PAB

[BS86] J. M. Borwein and H. M. Strójwas. Proximal analysis and boundaries of closed sets in Banach space. I. Theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 38(2):431– 452, 1986. CODEN CJMAAB. ISSN 0008-414X (print), 1496-

4279 (electronic). URL http://docserver.carma.newcastle.
edu.au/1611/.

Borwein:1987:PAB

[BS87] J. M. Borwein and H. M. Strójwas. Proximal analysis and boundaries of closed sets in Banach space. II. Applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 39(2):428–472, 1987. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL http://docserver.carma.

newcastle.edu.au/1600/.

Borwein:1989:HC

[BS89] J. M. Borwein and H. M. Strójwas. The hypertangent cone. Nonlinear Analysis, Theory, Methods and Applications, 13(2):125– 144, 1989. CODEN NOANDD. ISSN 0362-546x (print), 1873-5215

(electronic). URL http://docserver.carma.newcastle.edu.
au/1588/;http://www.sciencedirect.com/science/article/
pii/0362546X89900394.

Borwein:1995:SAD

[BS95] J. M. Borwein and W. Sun. The stability analysis of dynamic SPECT systems. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, 1995. 14 pp. URL http:

//docserver.carma.newcastle.edu.au/99.

Borwein:1997:SAD

[BS97a] J. M. Borwein and W. Sun. The stability analysis of dynamic SPECT systems. Numerische Mathematik, 77(3):283–298, 1997. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
99/;http://link.springer.com/article/10.1007/s002110050287.

Borwein:1997:OJP

[BS97b] Jonathan Borwein and Richard Smith. Online journal publication: two views from the electronic trenches. Canadian Journal of Communication, 22(??):135–152, ???? 1997. URL http:

//edie.cprost.sfu.ca/~scom.

Borwein:1999:ITDb

[BS99a] J. M. Borwein and T. Stanway. The impact of technology on the doing of mathematics. In John Grant McGloughlin, editor, Canadian Mathematics Education Study Group (CMESG) Proceedings 1999 Annual Meeting, pages 3–6. Memorial University Press, St. Johns, NL, Canada, 1999. URL http://docserver.

carma.newcastle.edu.au/248/.

Borwein:1999:NCMb

[BS99b] J. M. Borwein and T. Stanway. Numerical and computational mathematics (at the undergraduate level). Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, August 31, 1999. 22 pp. Published in [BS00].

Borwein:1999:ITDa

[BS99c] Jonathan M. Borwein and Terry Stanway. The impact of technology on the doing of mathematics. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, October 8, 1999.

7 pp. URL http://docserver.carma.newcastle.edu.au/248/.

Published in [BS99a].

Borwein:1999:NCMa

[BS99d] Jonathan M. Borwein and Terry Stanway. Numerical and computational mathematics at the undergraduate level. Technology in Mathematics Education (TMEST), Plenary, Brock University, St. Catharines, ON, Canada, June 3–4., June 4, 1999. URL

http://docserver.carma.newcastle.edu.au/246/.

Borwein:19xx:ASW

[BSxx] Jonathan Borwein and Lou Shitou. Asymptotics of a sequence of Witt vectors. Report, Department of Mathematics, Statistics and Computing Science, Dalhousie University, Halifax, NS B3H 3J5, Canada, 19xx. 13 pp.

Borwein:2000:NCMa

[BS00] J. M. Borwein and T. Stanway. Numerical and computational mathematics (at the undergraduate level). In Bruce Cload and Tom Jenkyns, editors, Proceedings of the Conference on Technology in Mathematics Education at the Secondary and Tertiary Levels, June 2–4, 1999 at Brock University, pages 20–37. Brock University Press, St. Catharines, ON, L2S 3A1 Canada, 2000.

URL http://docserver.carma.newcastle.edu.au/246/.

Borwein:2003:MDM

[BS03] J. M. Borwein and T. Stanway. Managing digital mathematical discourse. In Asperti et al. [ABD03], pages 45–55. ISBN 3-540-00568-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.95 .I565 2003. URL http://www.loc.gov/

catdir/enhancements/fy0817/2003042408-d.html.

Borwein:2005:KCM

[BS05] Jonathan Borwein and Terry Stanway. Knowledge and community in mathematics. The Mathematical Intelligencer, 27(2):7– 16, 2005. CODEN MAINDC. ISSN 0343-6993 (print), 1866-

7414 (electronic). URL http://docserver.carma.newcastle.
edu.au/274/.

Borwein:2007:PRB

[BS07] J. M. Borwein and B. Salvy. A proof of a recursion for Bessel moments. ArXiv e-prints, June 2007. URL http://adsabs.

harvard.edu/abs/2007arXiv0706.1409B;http://docserver.
carma.newcastle.edu.au/346/.

Borwein:2008:PRB

[BS08] Jonathan M. Borwein and Bruno Salvy. A proof of a recurrence for Bessel moments. Experimental Mathematics, 17(2):223– 230, 2008. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://docserver.carma.newcastle.edu.
au/346/;http://projecteuclid.org/euclid.em/1227118973.

Borwein:2010:ENE

[BS10a] Jonathan M. Borwein and Scott Sciffer. An explicit non-expansive function whose subdifferential is the entire dual ball. In Nonlinear analysis and optimization II. Optimization, volume 514 of Contemp. Math., pages 99–103. American Mathematical Society, Providence, RI, USA, 2010. URL http://docserver.carma.

newcastle.edu.au/385/.

Borwein:2010:DRA

[BS10b] Jonathan M. Borwein and Brailey Sims. Douglas–Rachford algorithm in the absence of convexity. Technical report, Centre for Computer-assisted Research Mathematics and its Applications (CARMA), School of Mathematical and Physical Sciences, University of Newcastle, Callaghan, NSW 2308, Australia, May 22,

2010. 20 pp. URL https://www.carma.newcastle.edu.au/jon/
dr.pdf.

Borwein:2010:DRIa

[BS10c] Jonathan M. Borwein and Brailey Sims. Douglas–Rachford iterations in the absence of convexity. CARMA Colloquium., Jan-

uary 28, 2010. URL https://www.carma.newcastle.edu.au/
jon/dr-talk.pdf.

Borwein:2010:DRIb

[BS10d] Jonathan M. Borwein and Brailey Sims. Douglas–Rachford iterations in the absence of convexity. ANZIAM-SigmaOpt Session, Queenstown, New Zealand, February 1–4., February 1, 2010. URL

https://www.carma.newcastle.edu.au/jon/dr-talk.pdf.

Borwein:2011:LSEa

[BS11a] J. M. Borwein and A. Straub. Log-sine evaluations of Mahler measures. ArXiv e-prints, March 2011. URL http://adsabs.

harvard.edu/abs/2011arXiv1103.3893Bhttps://arxiv.org/
abs/1103.3893;https://arxiv.org/pdf/1103.3893.pdf.

Borwein:2011:DRA

[BS11b] Jonathan M. Borwein and Brailey Sims. The Douglas–Rachford algorithm in the absence of convexity. In Bauschke et al. [BBC^{+}11b], pages 93–109. ISBN 1-4419-9568-4, 1-4419-9569-2 (e-book). ISSN 1931-6828 (print), 1931-6836 (electronic).

LCCN QA378.5 .F59 2011. URL http://link.springer.com/
chapter/10.1007/978-1-4419-9569-8_6.

Borwein:2011:IMM

[BS11c] Jonathan M. Borwein and Matthew P. Skerritt. An introduction to modern mathematical computing: with Maple. Springer Undergraduate Texts in Mathematics and Technology. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 1-4614-0121-6. xvi + 216 pp. LCCN QA76.6

.B665 2011. URL http://docserver.carma.newcastle.edu.
au/1727/.

Borwein:2011:SVGa

[BS11d]  Jonathan M. Borwein and Armin Straub.  Special val-

ues of generalized log-sine integrals. ArXiv e-prints, March

2011. URL http://adsabs.harvard.edu/abs/2011arXiv1103.
4298B;http://docserver.carma.newcastle.edu.au/1398/.

Borwein:2011:SVGb

[BS11e] Jonathan M. Borwein and Armin Straub. Special values of generalized log-sine integrals. In Proceedings of the 36th International Symposium on Symbolic and Algebraic Computation, ISSAC ’11, pages 43–50. ACM Press, New York, NY 10036, USA, 2011. ISBN 1-4503-0675-6. URL http://

docserver.carma.newcastle.edu.au/1398/;http://doi.acm.
org/10.1145/1993886.1993899. Awarded ISSAC 2011 Best

SIGSAM-ACM Student Paper Prize.

Borwein:2012:IMM

[BS12a] Jonathan M. Borwein and Matthew P. Skerritt. An introduction to modern mathematical computing: with Mathematica(R). Springer undergraduate texts in mathematics and technology. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 1-4614-4252-4 (print), 1-4614-4253-2 (e-book). ISSN 1867-5506 (print), 1867-5514 (electronic). LCCN

QA76.95 .B67 2012. URL http://public.eblib.com/choice/
publicfullrecord.aspx?p=1030664.

Borwein:2012:LSEa

[BS12b] Jonathan M. Borwein and Armin Straub. Log-sine evaluations of Mahler measures. Journal of the Australian Mathematical Society, 92(1):15–36, 2012. CODEN JAUMAX. ISSN 1446-7887 (print), 1446-8107 (electronic). URL http://docserver.carma.

newcastle.edu.au/1384/;http://journals.cambridge.org/
action/displayAbstract?fromPage=online&aid=8610681.

Borwein:2013:MMS

[BS13] Jonathan M. Borwein and Armin Straub. Mahler measures, short walks and log-sine integrals. Theoretical Computer Science, 479:4– 21, 2013. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294

(electronic). URL http://docserver.carma.newcastle.edu.
au/1383/;http://www.sciencedirect.com/science/article/

pii/S0304397512009462. Symbolic-Numerical Algorithms.

Borwein:2014:ROM

[BS14a] J. M. Borwein and Brailey Sims. Review of Origins of Mathematical Words: A Comprehensive Dictionary of Latin, Greek and Arabic Roots by Anthony lo Bello. The John Hopkins University Press, 2013. Australian Mathematical Society Gazette, 41(2):116– 118, May 2014. ISSN 0311-0729 (print), 1326-2297 (electronic).

URL http://www.austms.org.au/Publ/Gazette/2014/May14/
BkRev.pdf.

Borwein:2014:BRB

[BS14b] Jonathan M. Borwein and Brailey Sims. Book review of: A. Lo Bello, Origins of mathematical words. A comprehensive dictionary of Latin, Greek and Arabic roots. Australian Mathematical Society Gazette, 41(2):116–118, May 2014. ISSN 0311-0729 (print),

1326-2297 (electronic). URL http://www.austms.org.au/Publ/
Gazette/2014/May14/BkRev.pdf.

Borwein:2015:CU

[BS15a] Jonathan M. Borwein and Brailey Sims. Convexity in ultraproducts. Preprint., August 2015.

Borwein:2015:RNP

[BS15b] Jonathan M. Borwein and Armin Straub. Relations for Nielsen polylogarithms. Journal of Approximation Theory, 193(??):74–88, May 2015. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430

(electronic). URL http://docserver.carma.newcastle.edu.
au/1500/;http://www.sciencedirect.com/science/article/

pii/S0021904513001202. Special Issue Dedicated to Dick Askey on the occasion of his 80th birthday.

Borwein:2016:EAM

[BS16a] J. M. Borwein and A. Straub. Entry for August on “Moments of random walks”. In Mathematical Beauties 2016 Calendar, page ?? Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016.

Borwein:2016:CFD

[BS16b] Jonathan M. Borwein and Corwin W. Sinnamon. A closed form for the density functions of random walks in odd dimensions. Bulletin of the Australian Mathematical Society, 93(2): 330–339, April 2016. CODEN ALNBAB. ISSN 0004-9727 (print), 1755-1633 (electronic). URL http://docserver.carma.

newcastle.edu.au/1702/;http://journals.cambridge.org/
action/displayAbstract?fromPage=online&aid=10230818.

Belovas:2017:LTC

[BS17] Igoris Belovas and Leonidas Sakalauskas. Limit theorems for the coefficients of the modified Borwein method for the calculation of the Riemann zeta-function values. Colloquium Mathematicum, 2017. CODEN CQMAAQ. ISSN 0010-1354 (print), 1730-6302 (electronic).

Borwein:2013:CLB

[BSM13] Jonathan M. Borwein, Matthew Skerritt, and Christopher Maitland. Computation of a lower bound to Giuga’s primality conjecture. Integers, 13(??):??, ???? 2013. CODEN INTEHN.

ISSN 1867-0652 (print), 1867-0660 (electronic). URL http:

//docserver.carma.newcastle.edu.au/1511/.

Borwein:2013:NCR

[BST13] J. M. Borwein, B. Sims, and M. K. Tam. Norm convergence of realistic projection and reflection methods. ArXiv e-

prints, December 2013. URL http://adsabs.harvard.edu/abs/
2013arXiv1312.7323B;http://docserver.carma.newcastle.
edu.au/1493/.

Borwein:2015:NCR

[BST15] Jonathan M. Borwein, Brailey Sims, and Matthew K. Tam. Norm convergence of realistic projection and reflection methods. Optimization, 64(1):161–178, January 2015. CODEN OPTZDQ. ISSN 0233-1934, 0323-3898. URL http://docserver.carma.

newcastle.edu.au/1493/.

Borwein:2015:DSU

[BSV15] J. M. Borwein, A. Straub, and C. Vignat. Densities of short uniform random walks in higher dimensions. ArXiv e-

prints, August 2015. URL http://adsabs.harvard.edu/abs/
2015arXiv150804729B;http://docserver.carma.newcastle.
edu.au/1699/.

Borwein:2016:DSU

[BSV16] Jonathan M. Borwein, Armin Straub, and Christophe Vignat. Densities of short uniform random walks in higher dimensions. Journal of Mathematical Analysis and Applications, 437(1):668–707, May 2016. CODEN JMANAK. ISSN 0022-247X (print), 1096-0813 (electronic). URL http://docserver.

carma.newcastle.edu.au/1699/;http://www.sciencedirect.
com/science/article/pii/S0022247X16000421.

Borwein:1982:SSI

[BSW82] J. M. Borwein, G. P. H. Styan, and H. Wolkowicz. Solutions: Some inequalities involving statistical expressions (L. V. Foster). SIAM Review, 24(3):340–342, ???? 1982. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Borwein:2013:TSF

[BSW13] Jonathan M. Borwein, Armin Straub, and James Wan. Three-step and four-step random walk integrals. Experimental Mathematics, 22(1):1–14, 2013. CODEN ???? ISSN 1058-6458 (print), 1944-

950X (electronic). URL http://docserver.carma.newcastle.
edu.au/763/.

Borwein:2011:DSU

[BSWZ11] Jonathan M. Borwein, A. Straub, J. Wan, and Wadim Zudilin. Densities of short uniform random walks. ArXiv e-prints, March

2011. URL http://adsabs.harvard.edu/abs/2011arXiv1103.
2995B;http://docserver.carma.newcastle.edu.au/1388/.

Borwein:2012:DSU

[BSWZ12] Jonathan M. Borwein, Armin Straub, James Wan, and Wadim Zudilin. Densities of short uniform random walks. Canadian Journal of Mathematics = Journal canadien de mathématiques, 64(5):961–990, 2012. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL http://arxiv.org/

abs/1103.2995;http://docserver.carma.newcastle.edu.au/

1388/. With an appendix by Don Zagier.

Borwein:1983:PSA

[BSZ^{+}83] J. Borwein, Gabor J. Szekely, Andras Zempleni, Michael Barr, Edmund Butler, Allen J. Schwenk, and Thomas Q. Sibley. Problems and solutions: Advanced problems: 6430-6435. American Mathematical Monthly, 90(6):402–403, 1983. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Borwein:2013:NTR

[BSZ13] Jonathan M. Borwein, Igor E. Shparlinski, and Wadim Zudilin, editors. Number theory and related fields: in memory of Alf van der Poorten, volume 43 of Springer proceedings in mathematics and statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-6641-5. ISSN 2194-1009. LCCN QA241

.N86755 2013. URL http://link.springer.com/book/10.
1007/978-1-4614-6642-0;http://www.loc.gov/catdir/enhancements/
fy1404/2013936399-d.html;http://www.loc.gov/catdir/enhancements/
fy1404/2013936399-t.html.

Borwein:1984:SCS

[BT84] J. M. Borwein and D. W. Tingley. On supportless convex sets. Report DALTR 84-04, Department of Mathematics, Dalhousie University, Halifax, NS, Canada, 1984. 13 pp.

Borwein:1985:SCS

[BT85] J. M. Borwein and D. W. Tingley. On supportless convex sets. Proceedings of the American Mathematical Society, 94(3):471– 476, 1985. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-

6826 (electronic). URL http://docserver.carma.newcastle.
edu.au/1619/.

Borwein:1992:STS

[BT92] J. M. Borwein and M. Théra. Sandwich theorems for semicontinuous operators. Canadian mathematical bulletin = Bulletin canadien de mathématiques, 35(4):463–474, 1992. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic).

Borwein:2013:CDRa

[BT13a] J. M. Borwein and M. K. Tam. A cyclic Douglas–Rachford iteration scheme. ArXiv e-prints, March 2013. URL http://adsabs.

harvard.edu/abs/2013arXiv1303.1859B;http://docserver.
carma.newcastle.edu.au/1469/.

Borwein:2013:CDRb

[BT13b] J. M. Borwein and M. K. Tam. The cyclic Douglas– Rachford method for inconsistent feasibility problems. ArXiv

e-prints, October 2013. URL http://adsabs.harvard.edu/abs/
2013arXiv1310.2195B.

Borwein:2014:RMIb

[BT14a] J. M. Borwein and M. K. Tam. Reflection methods for inverse problems with application to protein conformation determination. ArXiv e-prints, August 2014. URL http://adsabs.

harvard.edu/abs/2014arXiv1408.4213B;http://docserver.
carma.newcastle.edu.au/1678/.

Borwein:2014:RMIa

[BT14b] J. M. Borwein and M. K. Tam. Reflection methods for inverse problems with applications to protein conformation determination. In CIMPA school Generalized Nash Equilibrium Problems, Bilevel programming and MPEC New Delhi, India, December 2012, page ?? ????, ????, 2014. URL http://docserver.carma.

newcastle.edu.au/1678/.

Borwein:2014:CDR

[BT14c] Jonathan M. Borwein and Matthew K. Tam. A cyclic Douglas– Rachford iteration scheme. Journal of Optimization Theory and Applications, 160(1):1–29, 2014. CODEN JOTABN. ISSN 0022-3239 (print), 1573-2878 (electronic). URL http://arxiv.org/

abs/1303.1859;http://docserver.carma.newcastle.edu.au/
1469/.

Borwein:2015:CDR

[BT15] Jonathan M. Borwein and Matthew K. Tam. The cyclic Douglas– Rachford method for inconsistent feasibility problems. Journal of Nonlinear and Convex Analysis, 16(4):573–584, April 2015. ISSN 1345-4773 (print), 1880-5221 (electronic). URL

http://docserver.carma.newcastle.edu.au/1496/;http://
www.ybook.co.jp/online2/jncav16-4.html.

Borwein:2017:RMI

[BT17] Jonathan M. Borwein and Matthew K. Tam. Reflection methods for inverse problems with applications to protein conformation determination. In Generalized Nash equilibrium problems, bilevel programming and MPEC, Forum Interdiscip. Math., pages 83– 100. Springer, Singapore, 2017.

Ben-Tal:1988:DAM

[BTBT88] A. Ben-Tal, J. M. Borwein, and M. Teboulle. A dual approach to multidimensional L_{p} spectral estimation problems. SIAM Journal on Control and Optimization, 26(4):985–996, 1988. CODEN SJCODE. ISSN 0363-0129 (print), 1095-7138 (electronic).

Borwein:1995:NCC

[BTZ95] Jonathan M. Borwein, Jay S. Treiman, and Qiji J. Zhu. Necessary conditions for constrained optimization problems with semicontinuous and continuous data. Report, Department of Mathematics & Statistics, Simon Fraser University, Burnaby, BC V5A 156, Canada, September 1995. URL http://docserver.carma.

newcastle.edu.au/116/2/95_051-Borwein-Treiman-Zhu.pdf.

Borwein:1997:SAR

[BTZ97] Jonathan M. Borwein, Jay Treiman, and Qiji Zhu. Sensitivity analysis in reflexive Banach spaces. Preprint., August 1997.

Borwein:1998:NCC

[BTZ98] Jonathan M. Borwein, Jay S. Treiman, and Qiji J. Zhu. Necessary conditions for constrained optimization problems with semicontinuous and continuous data. Transactions of the American Mathematical Society, 350(6):2409–2429, 1998. CODEN TAM-TAM. ISSN 0002-9947 (print), 1088-6850 (electronic). URL

http://docserver.carma.newcastle.edu.au/116/.

Borwein:1999:PSVa

[BTZ99] Jonathan M. Borwein, Jay S. Treiman, and Qiji J. Zhu. Partially smooth variational principles and applications. Nonlinear Analysis, Theory, Methods and Applications, 35(8):1031–1059, 1999. CODEN NOANDD. ISSN 0362-546x (print), 1873-5215 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
182/;http://www.sciencedirect.com/science/article/pii/
S0362546X98001138.

Borwein:1993:DKK

[BV93a] J. M. Borwein and J. Vanderwerff. Dual Kadec–Klee norms and the relationships between Wijsman, slice and Mosco convergence. ArXiv Mathematics e-prints, February 1993. URL

http://adsabs.harvard.edu/abs/1993math......2212B.

Borwein:1993:FAS

[BV93b] Jon Borwein and Jon Vanderwerff. Further arguments for slice convergence in nonreflexive spaces. Report, Department of Mathematics & Statistics, Simon Fraser University, Burnaby, BC V5A 156, Canada, November 6, 1993. 18 pp. URL http://docserver.

carma.newcastle.edu.au/62.

Borwein:1994:BSA

[BV94a] J. M. Borwein and J. D. Vanderwerff. Banach spaces that admit support sets. Report, Department of Mathematics & Statistics, Simon Fraser University, Burnaby, BC V5A 156, Canada, July 10,

1994. 6 pp. URL http://docserver.carma.newcastle.edu.au/
87.

Borwein:1994:CFS

[BV94b] Jon Borwein and Jon Vanderwerff. Convex functions on “sequentially reflexive” Banach spaces. Report, Department of Mathematics & Statistics, Simon Fraser University, Burnaby, BC V5A 156, Canada, September 14, 1994. 11 pp.

URL http://docserver.carma.newcastle.edu.au/115/2/95_
050-Borwein-Vanderwerff.pdf.

Borwein:1994:DKK

[BV94c] Jon Borwein and Jon Vanderwerff. Dual Kadec–Klee norms and the relationships between Wijsman, slice, and Mosco convergence. Michigan Mathematical Journal, 41(2):371–387, 1994. CODEN MIMJA5. ISSN 0026-2285 (print), 1945-2365 (electronic).

Borwein:1994:FAS

[BV94d] Jon Borwein and Jon Vanderwerff. Further arguments for slice convergence in nonreflexive spaces. Set-Valued Analysis, 2(4):529– 544, 1994. CODEN SVANEG. ISSN 0927-6947 (print), 1572-

932x (electronic). URL http://docserver.carma.newcastle.
edu.au/62/.

Borwein:1995:CLR

[BV95a] Jonathan M. Borwein and Jon D. Vanderwerff. Convergence of Lipschitz regularizations of convex functions. Journal of Functional Analysis, 128(1):139–162, 1995. CODEN JFUAAW. ISSN 0022-1236 (print), 1096-0783 (electronic). URL

http://docserver.carma.newcastle.edu.au/1540/;http://
www.sciencedirect.com/science/article/pii/S0022123685710269.

Borwein:1995:EUC

[BV95b] Jonathan M. Borwein and Jon D. Vanderwerff. Epigraphical and uniform convergence of convex functions. Report, Department of Mathematics & Statistics, Simon Fraser University, Burnaby, BC V5A 156, Canada, March 17, 1995. 19 pp. URL http://

docserver.carma.newcastle.edu.au/94/.

Borwein:1995:SRSa

[BV95c] Jonathan M. Borwein and Jon D. Vanderwerff. A survey on renorming and set convergence. Report, Department of Mathematics & Statistics, Simon Fraser University, Burnaby, BC V5A 156, Canada, March 25, 1995. 19 pp. URL http://docserver.

carma.newcastle.edu.au/97. Published in [BV95d].

Borwein:1995:SRSb

[BV95d] Jonathan M. Borwein and Jon D. Vanderwerff. A survey on renorming and set convergence. Topological Methods in Nonlinear Analysis, 5(2):211–228, 1995. ISSN 1230-3429. URL http://

docserver.carma.newcastle.edu.au/97/;https://www.tmna.
ncu.pl/static/files/v05n2-01.pdf.

Borwein:1996:BSAa

[BV96a] J. M. Borwein and J. D. Vanderwerff. Banach spaces that admit support sets. Proceedings of the American Mathematical Society, 124(3):751–755, 1996. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL http://docserver.carma.

newcastle.edu.au/87/.

Borwein:1996:BSAb

[BV96b] J. M. Borwein and J. D. Vanderwerff. Banach spaces that admit support sets. Proceedings of the American Mathematical Society, 124(3):751–756, March 1996. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL http:

//docserver.carma.newcastle.edu.au/87/;http://www.ams.
org/journals/proc/1996-124-03/S0002-9939-96-03122-X/.

Borwein:1996:EUC

[BV96c] Jonathan M. Borwein and Jon D. Vanderwerff. Epigraphical and uniform convergence of convex functions. Transactions of the American Mathematical Society, 348(4):1617–1631, April 1996. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850

(electronic). URL http://docserver.carma.newcastle.edu.
au/94/;https://www.jstor.org/stable/2155155.

Borwein:1997:CFB

[BV97] Jon Borwein and Jon Vanderwerff. Convex functions on Banach spaces not containing l_{1}. Canadian mathematical bulletin = Bulletin canadien de mathématiques, 40(1):10–18, 1997. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287 (electronic). URL

http://docserver.carma.newcastle.edu.au/1532/.

Borwein:199x:SCF

[BV9x] Jonathan M. Borwein and Jon D. Vanderwerff. A survey of convex functions and sequential convergence. Technical report, Faculty of Computer Science, Dalhousie University, Halifax, NS, B3H2W5

Canada, 199x. 1–10 pp. URL https://carma.newcastle.edu.
au/jon/Preprints/Books/CUP/CUPold/convsurvey.pdf.

Borwein:2000:CBC

[BV00a] J. M. Borwein and J. D. Vanderwerff. On the continuity of biconjugate convex functions. Report, Department of Mathematics, Simon Fraser University, Burnaby, BC V5A 1S6, Canada, 2000. 10

pp. URL http://docserver.carma.newcastle.edu.au/230/.

Published in [BV02].

Borwein:2000:CFL

[BV00b] Jonathan M. Borwein and Jon D. Vanderwerff. Convex functions of Legendre type in general Banach spaces. Report, Department of Mathematics, Simon Fraser University, Burnaby, BC V5A 1S6, Canada, 2000. 14 pp. URL http://docserver.carma.

newcastle.edu.au/228/. Published in [BV01].

Borwein:2001:CFL

[BV01] Jonathan M. Borwein and Jon D. Vanderwerff. Convex functions of Legendre type in general Banach spaces. Journal of Convex Analysis, 8(2):569–581, 2001. ISSN 0944-6532 (print), 2363-

6394 (electronic). URL http://docserver.carma.newcastle.
edu.au/228/;http://www.heldermann.de/JCA/JCA08/JCA082/
jca08030.htm.

Borwein:2002:CBC

[BV02] J. M. Borwein and J. D. Vanderwerff. On the continuity of biconjugate convex functions. Proceedings of the American Mathematical Society, 130(6):1797–1803, June 2002. CO-DEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
230/;https://www.jstor.org/stable/2699775.

Borwein:2004:CCS

[BV04] Jonathan M. Borwein and Jon D. Vanderwerff. Constructible convex sets. Set-Valued Analysis, 12(1–2):61–77, March/June 2004. CODEN SVANEG. ISSN 0927-6947 (print), 1572-932x

(electronic). URL http://docserver.carma.newcastle.edu.
au/38/.

Borwein:2009:DCF

[BV09] Jonathan Borwein and Jon Vanderwerff. Differentiability of conjugate functions and perturbed minimization principles. Journal of Convex Analysis, 16(3–4):707–711, 2009. ISSN 0944-6532 (print), 2363-6394 (electronic). URL http://docserver.

carma.newcastle.edu.au/383/;http://www.heldermann.de/
JCA/JCA16/JCA163/jca16042.htm.

Borwein:2010:FLF

[BV10a] Jonathan M. Borwein and Jon Vanderwerff. Fréchet–Legendre functions and reflexive Banach spaces. Journal of Convex Analysis, 17(3–4):915–924, 2010. ISSN 0944-6532 (print),

2363-6394 (electronic). URL http://www.heldermann.de/JCA/

JCA17/JCA173/jca17057.htm. Special volume in honour of Hedi Attouch.

Borwein:2010:CFC

[BV10b] Jonathan M. Borwein and Jon D. Vanderwerff. Convex functions: constructions, characterizations and counterexamples, volume 109 of Encyclopedia of Mathematics and its Applications. Cambridge University Press, Cambridge, UK, 2010. ISBN 0-521-85005-3. x +

521 pp. URL http://ebooks.cambridge.org/ebook.jsf?bid=
CBO9781139087322.

Borwein:2012:CUC

[BV12] Jonathan M. Borwein and Jon Vanderwerff. Constructions of uniformly convex functions. Canadian mathematical bulletin = Bulletin canadien de mathématiques, 55(4):697–707, December 2012. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic). URL http://docserver.carma.newcastle.edu.
au/778/.

Bradshaw:2024:DIV

[BV24] Zachary P. Bradshaw and Christophe Vignat. Dubious identities: a visit to the Borwein zoo. American Mathematical Monthly, 131 (5):400–416, 2024. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See [BB92a].

Borwein:2014:NFI

[BvdPSZ14] Jonathan M. Borwein, Alfred Jacobus van der Poorten, Jeffrey Outlaw Shallit, and Wadim Zudilin. Neverending Fractions: an Introduction to Continued Fractions, volume 23 of Australian Mathematical Society lecture series. Cambridge University Press, Cambridge, UK, 2014. ISBN 0-521-18649-8. x + 212 pp. LCCN QA295 .B667 2014. URL http://docserver.

carma.newcastle.edu.au/1722/;http://ebooks.cambridge.
org/ebook.jsf?bid=CBO9780511902659.

Ponomarenko:2021:YAMi

[bVP21] Edited by Vadim Ponomarenko. 100 years ago this month in The American Mathematical Monthly. American Mathematical Monthly, 128(9):855, 2021. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Borwein:2001:LLC

[BVW01] J. M. Borwein, J. Vanderwerff, and Xianfu Wang. Local Lipschitzconstant functions and maximal subdifferentials. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada,

2001. 28 pp. URL http://docserver.carma.newcastle.edu.
au/55/2/01_158-Borwein-Vanderwerff-Wang.pdf. Published

in [BVW03].

Borwein:2003:LLC

[BVW03] J. M. Borwein, J. Vanderwerff, and Xianfu Wang. Local Lipschitzconstant functions and maximal subdifferentials. Set-Valued Analysis, 11(1):37–67, 2003. CODEN SVANEG. ISSN 0927-6947 (print), 1572-932X (electronic). URL http://link.springer.

com/article/10.1023/A%3A1021975622768.

Borwein:1979:COA

[BW79a] J. M. Borwein and H. Wolkowicz. Characterizations of optimality for the abstract convex program. Research Report 19, Department of Mathematics, Dalhousie University, Halifax, NS, Canada, July 1979. 54 pp.

Borwein:1979:COC

[BW79b] J. M. Borwein and H. Wolkowicz. Characterizations of optimality without constraint qualification for the abstract convex program.

Research Report 14, Department of Mathematics, Dalhousie University, Halifax, NS, Canada, June 1979. 59 pp.

Borwein:1981:CCP

[BW81a] J. M. Borwein and H. Wolkowicz. Cone-convex programming, stability and affine constraints. In Schaible and Ziemba [SZ81], pages 379–397. ISBN 0-12-621120-5. LCCN QA402.5 .G45. Invited paper.

Borwein:1981:RAC

[BW81b] Jon Borwein and Henry Wolkowicz. Regularizing the abstract convex program. Journal of Mathematical Analysis and Applications, 83(2):495–530, 1981. CODEN JMANAK. ISSN 0022-247X (print), 1096-0813 (electronic). URL http://docserver.

carma.newcastle.edu.au/1645/;http://www.sciencedirect.
com/science/article/pii/0022247X81901384.

Borwein:1981:COA

[BW81c] Jon M. Borwein and Henry Wolkowicz. Characterization of optimality for the abstract convex program with finite dimensional range. Journal of the Australian Mathematical Society. Series A: Pure Mathematics and Statistics, 30(4): 390–411, April 1981. CODEN JAMADS. ISSN 0263-6115 (print), 2396-8192 (electronic). URL http://docserver.carma.

newcastle.edu.au/1649/;http://journals.cambridge.org/
action/displayAbstract?fromPage=online&aid=4896312.

Borwein:1981:FRC

[BW81d] Jon M. Borwein and Henry Wolkowicz. Facial reduction for a cone-convex programming problem. Journal of the Australian Mathematical Society. Series A: Pure Mathematics and Statistics, 30(3):369–380, February 1981. CODEN JA-MADS. ISSN 0263-6115 (print), 2396-8192 (electronic). URL

http://docserver.carma.newcastle.edu.au/1650/;http://
journals.cambridge.org/action/displayAbstract?fromPage=
online&aid=5441184.

Borwein:1982:COCa

[BW82a] J. M. Borwein and H. Wolkowicz. Characterizations of optimality without constraint qualification for the abstract convex program. Mathematical Programming Study, 19:77–100, 1982. CODEN MP-STDF. ISSN 0303-3929. Optimality and stability in mathematical programming.

Borwein:1982:COCb

[BW82b] J. M. Borwein and H. Wolkowicz. Characterizations of optimality without constraint qualification for the abstract convex program. In Optimality and Stability in Mathematical Programming, page ?? Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. URL http://link.springer.

com/chapter/10.1007/BFb0120983.

Borwein:1986:SCQ

[BW86] J. M. Borwein and H. Wolkowicz. A simple constraint qualification in infinite-dimensional programming. Mathematical Programming, 35(1):83–96, 1986. CODEN MHPGA4. ISSN 0025-5610 (print), 1436-4646 (electronic). URL http://docserver.

carma.newcastle.edu.au/1606/;http://link.springer.com/
article/10.1007/BF01589443.

Borwein:1995:DDF

[BW95a] J. M. Borwein and Xianfu Wang. Distinct differentiable functions may share the same Clarke subdifferential at all points. Report, Centre for Computer-assisted Research Mathematics and its Applications (CARMA), School of Mathematical and Physical Sciences, University of Newcastle, Callaghan, NSW 2308, Australia, June 29, 1995. 10 pp. URL https://docserver.carma.

newcastle.edu.au/112/2/95_047-Borwein-Wang.pdf. Pub-

lished in [BW97a].

Borwein:1995:SRR

[BW95b] J. M. Borwein and E. Wong. A survey of results relating to Giuga’s conjecture on primality. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, May 8, 1995. 21 pp. Published in [BW97b].

Borwein:1997:DDF

[BW97a] J. M. Borwein and Xianfu Wang. Distinct differentiable functions may share the same Clarke subdifferential at all points. Proceedings of the American Mathematical Society, 125(3):807– 813, 1997. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-

6826 (electronic). URL http://docserver.carma.newcastle.
edu.au/112/.

Borwein:1997:SRR

[BW97b] J. M. Borwein and E. Wong. A survey of results relating to Giuga’s conjecture on primality. In L. Vinet, editor, Advances in mathematical sciences: CRM’s 25 years (Montreal, PQ, 1994), volume 11 of CRM Proc. Lecture Notes, pages 13–27. American Mathematical Society, Providence, RI, USA, 1997. URL

http://docserver.carma.newcastle.edu.au/101/.

Borwein:1998:CMV

[BW98a] J. M. Borwein and Xianfu Wang. The converse of the mean value theorem may fail generically. American Mathematical Monthly, 105(9):847–848, November 1998. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL

http://docserver.carma.newcastle.edu.au/1530/;https://
www.jstor.org/stable/2589216.

Borwein:1998:LFM

[BW98b] Jonathan M. Borwein and Xianfu Wang. Lipschitz functions with maximal Clarke subdifferentials are generic. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, September 11, 1998. URL http://docserver.carma.

newcastle.edu.au/206/. Published in [BW00].

Borwein:1999:STC

[BW99] J. M. Borwein and X. Wang. Subdifferentiability of typical continuous functions. Report, Centre for Experimental and Constructive Mathematics (CECM) at Simon Fraser University (SFU), Burnaby, BC V5A 1S6, Canada, August 13, 1999. 10 pp. URL

http://docserver.carma.newcastle.edu.au/241/. Published

in [BW01].

Borwein:2000:LFM

[BW00] Jonathan M. Borwein and Xianfu Wang. Lipschitz functions with maximal Clarke subdifferentials are generic. Proceedings of the American Mathematical Society, 128(11):3221–3229, 2000. CO-DEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic).

URL http://docserver.carma.newcastle.edu.au/206/.

Borwein:2001:STC

[BW01] J. M. Borwein and X. Wang. Subdifferentiability of typical continuous functions. Nonlinear Analysis Forum, 6(1):49– 58, 2001. ISSN 1226-7228. URL http://docserver.carma.

newcastle.edu.au/241/;http://prof.ks.ac.kr/bslee/naf/
table/vol-0601/v0601-05a.PDF. Nonlinear analysis and its ap-

plications (St. John’s, NF, 1999).

Borwein:2003:LFL

[BW03] J. M. Borwein and Xianfu Wang. Lipschitz functions on the line with prescribed Hölder subdifferentials. Advanced Studies in Contemporary Mathematics (Kyungshang)., 7(2):93–117, 2003. ISSN

1229-3067. URL http://docserver.carma.newcastle.edu.au/
1526/.

Borwein:2005:CMF

[BW05a] Jonathan M. Borwein and Xianfu Wang. Cone-monotone functions: differentiability and continuity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 57(5):961–982, 2005. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Borwein:2005:LFM

[BW05b] Jonathan M. Borwein and Xianfu Wang. Lipschitz functions with maximal Clarke subdifferentials are staunch. Bulletin of the Australian Mathematical Society, 72(3):491–496, 2005. CODEN ALNBAB. ISSN 0004-9727 (print), 1755-1633 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
302/;http://journals.cambridge.org/action/displayAbstract?
fromPage=online&aid=4829356.

Borwein:2006:AMO

[BW06] Jonathan M. Borwein and H. Wiersma. Acyclic monotone operators. Atlantic Analysis Days, January 20–21, Dalhousie University, Halifax, NS, Canada., January 20, 2006.

Borwein:2007:ADMa

[BW07] Jonathan Borwein and Herre Wiersma. Asplund decomposition of monotone operators. SIAM Journal on Optimization, 18(3): 946–960, 2007. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-

7189 (electronic). URL http://docserver.carma.newcastle.
edu.au/321/.

Borwein:1997:MIM

[BWB97] Jonathan M. Borwein, Carolyn Watters, and Ephraim J. Borowski, editors. The MathResource interactive math dictionary: version 1.0. The math visualizing program Newton and Leibniz didn’t have. [An essential tool for high school, college and university students, teachers, scientists, engineers, researchers and anyone else who uses or studies mathematics]. MathResources, Inc., Halifax, NS, Canada, 1997. ISBN 3-540-14650-4. ???? pp.

LCCN ???? URL http://www.mathresources.com.

Bauschke:2010:BWD

[BWY10] Heinz H. Bauschke, Xianfu Wang, and Liangjin Yao. On Borwein–Wiersma decompositions of monotone linear relations. SIAM Journal on Optimization, 20(5):2636–2652, 2010. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

Borwein:1984:ANV

[BY84] J. M. Borwein and D. T. Yost. Absolute norms on vector lattices. Proceedings of the Edinburgh Mathematical Society (2), 27(2):215–222, June 1984. CODEN PEMSA3. ISSN 0013-0915 (print), 1464-3839 (electronic). URL http://docserver.carma.

newcastle.edu.au/1624/;http://journals.cambridge.org/
action/displayAbstract?fromPage=online&aid=3076132.

Borwein:2006:QCT

[BY06] Jonathan Borwein and Peter (Liquin) Ye. Quadratic convergence of the tanh–sinh quadrature rule. Preprint., February 2006. URL

http://docserver.carma.newcastle.edu.au/342/.

Borwein:2012:LTI

[BY12a] J. M. Borwein and L. Yao. Legendre-type integrands and convex integral functions. ArXiv e-prints, August 2012. URL http://

adsabs.harvard.edu/abs/2012arXiv1208.5217B;http://docserver.
carma.newcastle.edu.au/1455/.

Borwein:2012:MSM

[BY12b] J. M. Borwein and L. Yao. Maximality of the sum of a maximally monotone linear relation and a maximally monotone operator. ArXiv e-prints, December 2012. URL http://adsabs.

harvard.edu/abs/2012arXiv1212.4266B;http://docserver.
carma.newcastle.edu.au/1483/.

Borwein:2012:RPMa

[BY12c] J. M. Borwein and L. Yao. Recent progress on monotone operator theory. ArXiv e-prints, October 2012. URL http://adsabs.

harvard.edu/abs/2012arXiv1210.3401B;http://docserver.
carma.newcastle.edu.au/1457/.

Borwein:2012:SRC

[BY12d] J. M. Borwein and L. Yao. Some results on the convexity of the closure of the domain of a maximally monotone operator. ArXiv

e-prints, May 2012. URL http://adsabs.harvard.edu/abs/
2012arXiv1205.4482B;http://docserver.carma.newcastle.
edu.au/1373/.

Borwein:2012:STMa

[BY12e] J. M. Borwein and L. Yao. Structure theory for maximally monotone operators with points of continuity. ArXiv e-prints, March

2012. URL http://adsabs.harvard.edu/abs/2012arXiv1203.
1101B;http://docserver.carma.newcastle.edu.au/1375/.

Borwein:2012:MMO

[BY12f] Jonathan M. Borwein and Liangjin Yao. Maximally monotone operators of negative infimum type are of dense type: the proof revisited. Preprint., June 2012.

Borwein:2013:STMa

[BY13a] J. M. Borwein and L. Yao. Sum theorems for maximally monotone operators of type (FPV). ArXiv e-prints, May

2013. URL http://adsabs.harvard.edu/abs/2013arXiv1305.
6691B;http://docserver.carma.newcastle.edu.au/1503/.

Borwein:2013:MSM

[BY13b] Jonathan M. Borwein and Liangjin Yao. Maximality of the sum of a maximally monotone linear relation and a maximally monotone operator. Set-Valued and Variational Analysis, 21(4):603–616, 2013. ISSN 1877-0533 (print), 1877-0541

(electronic). URL http://arxiv.org/abs/1212.4266;http://
docserver.carma.newcastle.edu.au/1483/.

Borwein:2013:STMb

[BY13c] Jonathan M. Borwein and Liangjin Yao. Structure theory for maximally monotone operators with points of continuity. Journal of Optimization Theory and Applications, 157(1):1–24, 2013. CODEN JOTABN. ISSN 0022-3239 (print), 1573-2878

(electronic). URL http://arxiv.org/abs/1203.1101;http://
docserver.carma.newcastle.edu.au/1375/.

Borwein:2014:LTI

[BY14a] Jonathan M. Borwein and Liangjin Yao. Legendre-type integrands and convex integral functions. Journal of Convex Analysis, 21(1): 261–288, 2014. ISSN 0944-6532 (print), 2363-6394 (electronic).

URL http://arxiv.org/abs/1208.5217;http://docserver.
carma.newcastle.edu.au/1455/;http://www.heldermann.de/
JCA/JCA21/JCA211/jca21015.htm.

Borwein:2014:SRC

[BY14b] Jonathan M. Borwein and Liangjin Yao. Some results on the convexity of the closure of the domain of a maximally monotone operator. Optimization Letters, 8(1):237–246, 2014. ISSN 1862-4472 (print), 1862-4480 (electronic). URL http://arxiv.org/

abs/1205.4482;http://docserver.carma.newcastle.edu.au/
1373/.

Borwein:2014:STMa

[BY14c] Jonathan M. Borwein and Liangjin Yao. Sum theorems for maximally monotone operators of type (FPV). Journal of the Australian Mathematical Society, 97(1):1–26, 2014. CODEN JAU-MAX. ISSN 1446-7887 (print), 1446-8107 (electronic). URL

http://arxiv.org/abs/1305.6691;http://docserver.carma.
newcastle.edu.au/1503/;http://journals.cambridge.org/
action/displayAbstract?fromPage=online&aid=9306990.

Borwein:2015:RPM

[BY15] Jonathan M. Borwein and Liangjin Yao. Recent progress on monotone operator theory. In Reich and Zaslavski [RZ15], pages 51–81. ISBN 1-4704-1480-5 (paperback). LCCN QA329 .I54 2015.

URL http://docserver.carma.newcastle.edu.au/1457/.

Borwein:1986:FMT

[BZ86] J. M. Borwein and D. Zhuang. On Fan’s minimax theorem. Mathematical Programming, 34(2):232–234, 1986. CODEN MH-PGA4. ISSN 0025-5610 (print), 1436-4646 (electronic). URL

http://docserver.carma.newcastle.edu.au/1608/;http://
link.springer.com/article/10.1007/BF01580587.

Borwein:1987:EIE

[BZ87] J. M. Borwein and I. J. Zucker. Elliptic integral evaluation of the gamma function at values with small rational denominators. Report, Department of Mathematics, Statistics and Computing Science, Dalhousie University and Department of Physics, King’s College, Halifax, NS B3H 3J5, Canada and London WC2, UK, September 2, 1987. 13 pp.

Borwein:1988:VNS

[BZ88] J. M. Borwein and D. M. Zhuang. Verifiable necessary and sufficient conditions for openness and regularity of set-valued and single-valued maps. Journal of Mathematical Analysis and Applications, 134(2):441–459, 1988. CODEN JMANAK. ISSN 0022-247X (print), 1096-0813 (electronic). URL http://docserver.

carma.newcastle.edu.au/1594/;http://www.sciencedirect.
com/science/article/pii/0022247X88900340.

Borwein:1991:SEC

[BZ91] J. M. Borwein and D. M. Zhuang. Super efficiency in convex vector optimization. Zeitschrift für Operations Research. Mathematical Methods of Operations Research, 35(3):175–184, 1991. CODEN

ZMMRFZ. ISSN 0340-9422. URL http://link.springer.com/
article/10.1007/BF01415905.

Borwein:1992:FEG

[BZ92] J. M. Borwein and I. J. Zucker. Fast evaluation of the gamma function for small rational fractions using complete elliptic integrals of the first kind. IMA Journal of Numerical Analysis, 12 (4):519–526, 1992. CODEN IJNADN. ISSN 0272-4979 (print), 1464-3642 (electronic).

Borwein:1993:SEV

[BZ93] J. M. Borwein and D. Zhuang. Super efficiency in vector optimization. Transactions of the American Mathematical Society, 338 (1):105–122, 1993. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic).

Borwein:1994:CPP

[BZ94a] Jonathan M. Borwein and J. Zhu. Control problems with perturbations in non-reflexive space. 33rd CDCIEEE Meetings Orlando, FL, USA., December 16, 1994.

Borwein:1994:VAN

[BZ94b] Jonathan M. Borwein and Qiji J. Zhu. Variational analysis in nonreflexive spaces and applications to control problems with L^{1} perturbations. Report, Department of Mathematics & Statistics, Simon Fraser University, Burnaby, BC V5A 156, Canada, May 2,

1994. 40 pp. URL http://docserver.carma.newcastle.edu.
au/78.

Borwein:1995:VSV

[BZ95] Jonathan M. Borwein and Qiji J. Zhu. Viscosity solutions and viscosity subderivatives in smooth Banach spaces with applications to metric regularity. Report, Department of Mathematics & Statistics, Simon Fraser University, Burnaby, BC V5A 156, Canada, March 18, 1995. 40 pp. URL http://docserver.carma.

newcastle.edu.au/80.

Borwein:1996:VSV

[BZ96] Jonathan M. Borwein and Qiji J. Zhu. Viscosity solutions and viscosity subderivatives in smooth Banach spaces with applications to metric regularity. SIAM Journal on Control and Optimization, 34(5):1568–1591, 1996. CODEN SJCODE. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http://docserver.carma.

newcastle.edu.au/80/.

Borwein:1997:VAN

[BZ97] Jonathan M. Borwein and Qiji J. Zhu. Variational analysis in nonreflexive spaces and applications to control problems with L^{1} perturbations. Nonlinear Analysis, Theory, Methods and Applications, 28(5):889–915, 1997. CODEN NOANDD. ISSN 0362-546x (print), 1873-5215 (electronic).

URL http://docserver.carma.newcastle.edu.au/78/;http:
//www.sciencedirect.com/science/article/pii/0362546X9500186Y.

Borwein:1998:LCE

[BZ98] Jonathan M. Borwein and Qiji J. Zhu. Limiting convex examples for nonconvex subdifferential calculus. Journal of Convex Analysis, 5(2):221–235, 1998. ISSN 0944-6532 (print), 2363-6394 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
191/;http://www.heldermann-verlag.de/jca/jca05/j163.pdf.

Borwein:1999:MFAa

[BZ99a] Jonathan M. Borwein and Qiji J. Zhu. Multifunctional and functional analytic techniques in nonsmooth analysis. Report, Department of Mathematics, Simon Fraser University, Burnaby, BC V5A 1S6, Canada, November 2, 1999. 97 pp. URL

http://docserver.carma.newcastle.edu.au/49.

Borwein:1999:MFAb

[BZ99b] Jonathan M. Borwein and Qiji J. Zhu. Multifunctional and functional analytic techniques in nonsmooth analysis. In F. H. Clarke and R. J. Stern, editors, Nonlinear analysis, differential equations and control (Montreal, QC, 1998), volume 528 of NATO Sci. Ser. C Math. Phys. Sci., pages 61–157. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands,

1999. URL http://docserver.carma.newcastle.edu.au/49/.

Borwein:1999:SSC

[BZ99c] Jonathan M. Borwein and Qiji J. Zhu. A survey of subdifferential calculus with applications. Nonlinear Analysis, Theory, Methods and Applications, 38(6):687–773, 1999. CO-DEN NOANDD. ISSN 0362-546x (print), 1873-5215 (elec-

tronic). URL http://docserver.carma.newcastle.edu.au/
199/;http://www.sciencedirect.com/science/article/pii/

S0362546X98001424. Dedicated to Francis Clarke on the occasion of his fiftieth birthday and the twenty-fifth birthday of the Clarke generalized gradient. See addendum [BZ02a].

Borwein:2002:ASS

[BZ02a] Jonathan M. Borwein and Qiji J. Zhu. Addendum: “A survey of subdifferential calculus with applications” [Nonlinear Anal. 38 (1999), no. 6, Ser. A: Theory Methods, 687–773; (2000j:49024)]. Nonlinear Analysis, Theory, Methods and Applications, 49(2, Ser. A: Theory Methods):295–296, 2002. CODEN NOANDD. ISSN 0362-546x (print), 1873-5215 (electronic). See [BZ99c].

Borwein:2002:SSC

[BZ02b] Jonathan M. Borwein and Qiji J. Zhu. A survey of subdifferential calculus with applications: [nonlinear anal. tma 38 (1999) 687–773]. Nonlinear Analysis, Theory, Methods and Applications, 49(2):295–296, 2002. CODEN NOANDD. ISSN 0362-546x (print), 1873-5215 (electronic). URL http://docserver.

carma.newcastle.edu.au/199/;http://www.sciencedirect.
com/science/article/pii/S0362546X01000992.

Borwein:2005:TVA

[BZ05] Jonathan M. Borwein and Qiji J. Zhu. Techniques of Variational Analysis. CMS Books in Mathematics/Ouvrages de Mathématiques de la SMC, 20. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. ISBN 0-387-24298-8. vi + 362 pp. URL http://docserver.carma.

newcastle.edu.au/1735/.

Borwein:2006:VMC

[BZ06] Jonathan M. Borwein and Qiji J. Zhu. Variational methods in convex analysis. Journal of Global Optimization, 35(2):197–213, 2006. CODEN JGOPEO. ISSN 0925-5001 (print), 1573-2916

(electronic). URL http://docserver.carma.newcastle.edu.

au/259/. Special issue in honour of Alex Rubinov’s 65th birthday.

Borwein:2011:MHM

[BZ11] Jonathan M. Borwein and Wadim Zudilin. Math honours: Multiple zeta values. Classroom notes, Centre for Computer Assisted RMA, University of Newcastle, Callaghan, NSW 2308, Australia,

October 19, 2011. 65 pp. URL http://carma.newcastle.edu.
au/MZVs/mzv.pdf.

Borwein:2013:VMP

[BZ13] Jonathan M. Borwein and Qiji J. Zhu. Variational methods in the presence of symmetry. Advances in Nonlinear Analysis, 2(3): 271–307, 2013. ISSN 2191-9496 (print), 2191-950X (electronic).

URL http://docserver.carma.newcastle.edu.au/1514/.

Borwein:2016:VAL

[BZ16] J. M. Borwein and Qiji (Jim) Zhu. A variational approach to Lagrange multipliers. Journal of Optimization Theory and Applications, 171(3):727–756, December 2016. CODEN JOTABN. ISSN 0022-3239 (print), 1573-2878 (electronic). URL http://link.

springer.com/article/10.1007/s10957-015-0756-2. Special

issue of JOTA on Nondifferentiable Optimization and Nonsmooth Analysis, dedicated to Vladimir Demyanov.

Bailey:2020:I

[BZ20a] David H. Bailey and Qiji J. Zhu. Introduction. In Bailey et al. [BBB^{+}20], pages 229–231. ISBN 3-030-36567-0 (print), 3-030-36568-9 (e-book). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Borwein:2020:EMF

[BZ20b] Jonathan M. Borwein and Qiji J. Zhu. Entropy maximization in finance. In Bailey et al. [BBB^{+}20], pages 275–295. ISBN 3-030-36567-0 (print), 3-030-36568-9 (e-book). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Borwein:2008:ECE

[BZB08] J. M. Borwein, I. J. Zucker, and J. Boersma. The evaluation of character Euler double sums. The Ramanujan Journal, 15(3): 377–405, 2008. CODEN RAJOF9. ISSN 1382-4090 (print), 1572-

9303 (electronic). URL http://docserver.carma.newcastle.
edu.au/255/;http://link.springer.com/article/10.1007/
s11139-007-9083-z.

Calude:2016:HFC

[Cal16] Cristian S. Calude. The Human Face of Computing, volume 9 of Advances in computer science and engineering. Texts. Imperial College Press, London, UK, 2016. ISBN 1-78326-643-0 (hardcover), 1-78326-645-7 (e-book). xvi + 432 pp.

LCCN QA76.9.C66. URL http://www.worldscientific.com/
worldscibooks/10.1142/p992.

Campbell:2016:RUN

[Cam16] Alexander Campbell. Research uncovers new sources of financial model risk: Past performance of financial models is no guarantee of future success, two forthcoming papers suggest. Risk.net, ??

(??):??, May 20, 2016. URL http://www.risk.net/risk/news/
2458806/research-uncovers-new-sources-of-financial-model-risk.

Cass:1999:BRP

[Cas99] Peter Cass. Book review: Pi and the AGM. Mathematical Gazette, 83(497):334–335, July 1999. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.jstor.org/

stable/3619084.

Calude:2020:RQC

[CC20a] Cristian S. Calude and Elena Calude. The road to quantum computational supremacy. In Bailey et al. [BBB^{+}20], pages 349–367. ISBN 3-030-36567-0 (print), 3-030-36568-9 (e-book). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Chan:2020:RWT

[CC20b] Eunice Y. S. Chan and Robert M. Corless. A random walk through experimental mathematics. In Bailey et al. [BBB^{+}20], pages 203–226. ISBN 3-030-36567-0 (print), 3-030-36568-9 (ebook). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Calkin:2021:WNM

[CDH^{+}21] Neil J. Calkin, Killian Davis, Evan Haithcock, Catherine M. Kenyon, and Sylvia Wu. What Newton might have known: Experimental mathematics in the classroom. American Mathematical Monthly, 128(9):845–855, 2021. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Campbell:2020:HP

[CDS20] John M. Campbell, Jacopo D’Aurizio, and Jonathan Sondow. Hypergeometry of the parbelos. American Mathematical Monthly, 127(1):23–32, 2020. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Chieu:2018:SCS

[CFG^{+}18] N. H. Chieu, J. W. Feng, W. Gao, G. Li, and D. Wu. SOS-Convex semialgebraic programs and its applications to robust optimization: A tractable class of nonsmooth convex optimization. Set-Valued and Variational Analysis, 26(2):305–326, June 2018. CODEN ???? ISSN 1877-0533 (print), 1877-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s11228-017-0456-1.

Cotrina:2018:EPE

[CG18] John Cotrina and Yboon García. Equilibrium problems: Existence results and applications. Set-Valued and Variational Analysis, 26(1):159–177, March 2018. CODEN ???? ISSN 1877-0533 (print), 1877-0541 (electronic). URL http://link.springer.

com/article/10.1007/s11228-017-0451-6.

Cohen:1995:AAA

[CGM95] G. (Gérard) Cohen, Marc Giusti, and Teo Mora, editors. Applied algebra, algebraic algorithms, and error-correcting codes: 11th international symposium, AAECC-11, Paris, France, July 1995: proceedings, volume 948 of Lecture notes in computer science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1995. ISBN 3-540-60114-7 (softcover).

LCCN QA268 .A35 1995. URL http://www.loc.gov/catdir/
enhancements/fy0815/95021560-d.html.

Chamberland:2003:BBF

[Cha03] Marc Chamberland. Binary BBP-formulae for logarithms and generalized Gaussian–Mersenne primes. Journal of Integer Sequences, 6(3):Article 03.3.7, 10, 2003. ISSN 1530-7638.

Chapman:2016:LEJ

[Cha16] Scott T. Chapman. A letter from the Editor: Jonathan M. Borwein (1951–2016). American Mathematical Monthly, 123(9): 847–848, November 2016. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL https://www.jstor.org/

stable/10.4169/amer.math.monthly.123.9.847.

Cook:1992:PSS

[CJKB92] Matthew Cook, Walther Janous, Marcin E. Kuczma, and David Borwein. Problems and solutions: Solutions of elementary problems: E3388. American Mathematical Monthly, 99(1):69–70, January 1992. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Stu90].

Cohn:2016:SPP

[CKM^{+}16] Henry Cohn, Abhinav Kumar, Stephen D. Miller, Danylo Radchenko, and Maryna Viazovska. The sphere packing problem in dimension 24. arxiv.org, ??(??):1–12, March 21, 2016. URL

http://arxiv.org/abs/1603.06518.

Casazza:2015:M

[CKR15] Peter G. Casazza, Steven G. (Steven George) Krantz, and Randi D. Ruden, editors. I, mathematician. Spectrum series. Mathematical Association of America, Washington, DC, USA, 2015. ISBN 0-88385-585-2 (print), 1-61444-521-4 (e-book). xiii + 273 pp. LCCN QA28 .I22 2015.

Cohen:2015:BRM

[Coh15] Marion Cohen. Book review: Mathematicians on Creativity. Edited by Peter B. Borwein, Peter Liljedahl, and Helen Zhai, The Mathematical Association of America, Washington DC, 2014, xviii + 199 pp, ISBN 978-0-88385-574-4, $30.00. American Mathematical Monthly, 122(6):613–616, June/July 2015. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.
math.monthly.122.6.613.

Combettes:2018:PFP

[Com18] Patrick L. Combettes. Perspective functions: Properties, constructions, and examples. Set-Valued and Variational Analysis, 26(2):247–264, June 2018. CODEN ???? ISSN 1877-0533 (print), 1877-0541 (electronic). URL http://link.springer.

com/article/10.1007/s11228-017-0407-x.

Cosgrave:2017:EGB

[Cos17] John Cosgrave. Extension of Gauss’ binomial coefficient congruence. Lecture at Trinity College, Dublin, Ireland., 2017.

Crisci:2020:SPB

[CPRZ20] Serena Crisci, Federica Porta, Valeria Ruggiero, and Luca Zanni. Spectral properties of Barzilai–Borwein rules in solving singly linearly constrained optimization problems subject to lower and upper bounds. SIAM Journal on Optimization, 30(2):1300–1326, ???? 2020. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

Crandall:2004:BIJ

[Cra04] Richard E. Crandall. On a Bessel-integral of J. Borwein. Report, Reed College, Portland, OR, USA, August 2004. 2 pp. URL

http://people.reed.edu/~crandall/papers/borweinJ.pdf.

Crandall:2012:GTB

[Cra12] Richard Crandall. The googol-th bit of the Erdős–Borwein constant. Integers, 12(5):811–840, 2012. CODEN INTEHN. ISSN 1867-0652 (print), 1867-0660 (electronic).

Chamberland:2021:ALE

[CS21] Marc Chamberland and Armin Straub. Apéry limits: Experiments and proofs. American Mathematical Monthly, 128(9):811– 824, 2021. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Cvijovic:2010:PBB

[Cvi10] Djurdje Cvijović. Proof of the Borwein–Broadhurst conjecture for a dilogarithmic integral arising in quantum field theory. arxiv.org, ??(??):??, October 31, 2010. URL http://arxiv.org/abs/1011.

0195.

Campbell:2016:LEJ

[CW16] Christiane Rousseau Eddy Campbell and Graham P. Wright. Letter to the Editor: Jonathan Borwein and the Canadian Mathematical Society. Canadian Mathematical Society Notes, 48(6):21, December 2016. ISSN 0045-5164. URL http://cms.math.ca/

notes/v48/n6/Notesv48n6.pdf.

Chen:2021:BBG

[CZX21] Yannan Chen, Xinzhen Zhang, and Yanwei Xu. A Barzilai– Borwein gradient algorithm for spatio-temporal Internet traffic data completion via tensor triple decomposition. Journal of Scientific Computing, 88(3):65:1–65:24, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s10915-021-01574-0.

Dai:2015:PBB

[DABY15] Yu-Hong Dai, Mehiddin Al-Baali, and Xiaoqi Yang. A positive Barzilai–Borwein-like stepsize and an extension for symmetric linear systems. In Numerical analysis and optimization, volume 134 of Springer Proc. Math. Stat., pages 59–75. Springer, Cham, 2015.

Demir:1988:PSSa

[DAK88] Huseyn Demir, Gene Arnold, and Vaclav Konecny. Problems and solutions: Solutions of elementary problems: E3164. American Mathematical Monthly, 95(7):660–661, August/September 1988. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [DNG^{+}86].

Deutsch:1988:PSSa

[DBCB88] Emeric Deutsch, Jonathan M. Borwein, Paul Cull, and G. Behrendt. Problems and solutions: Solutions of elementary problems:

E3159. American Mathematical Monthly, 95(6):557–558, June/ July 1988. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [DNG^{+}86].

Davis:2015:MSS

[DD15] Ernest Davis and Philip J. Davis, editors. Mathematics, Substance, and Surmise: Views on the Meaning and Ontology of Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2015. ISBN 3-319-21472-1, 3-319-21473-X (e-book). LCCN QA8.4 .M38 2015.

Deville:199x:ADT

[Dev9x] Robert Deville. Autour du théorème de Borwein et Preiss. (French) [On the theorem of Borwein and Preiss]. In Séminaire d’Initiation à l’Analyse, volume 95 of Publ. Math. Univ. Pierre et Marie Curie, pages Exp. No. 25, 7. Univ. Paris VI, Paris, 199x.

Devlin:2017:LJB

[Dev17] Keith Devlin. The legacy of Jonathan Borwein. Web blog.,

September 20, 2017. URL http://devlinsangle.blogspot.
com/2017/09/the-legacy-of-jonathan-borwein.html.

Devlin:2020:HML

[Dev20] Keith Devlin. How mathematicians learned to stop worrying and love the computer. In Bailey et al. [BBB^{+}20], pages 133–139. ISBN 3-030-36567-0 (print), 3-030-36568-9 (e-book). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Dai:2005:PBB

[DF05] Yu-Hong Dai and Roger Fletcher. Projected Barzilai–Borwein methods for large-scale box-constrained quadratic programming. Numerische Mathematik, 100(1):21–47, 2005. CODEN NUMMA7. ISSN 0029-599x (print), 0945-3245 (electronic).

Dinh:2020:CRS

[DGLV20] Nguyen Dinh, Miguel A. Goberna, Marco A. López, and Michel Volle. Characterizations of robust and stable duality for linearly perturbed uncertain optimization problems. In Bailey et al. [BBB^{+}20], pages 43–74. ISBN 3-030-36567-0 (print), 3-030-36568-9 (e-book). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Dai:2006:CBB

[DHSZ06] Yu-Hong Dai, William W. Hager, Klaus Schittkowski, and Hongchao Zhang. The cyclic Barzilai–Borwein method for unconstrained optimization. IMA Journal of Numerical Analysis, 26(3):604–627, 2006. CODEN IJNADN. ISSN 0272-4979 (print), 1464-3642 (electronic).

Dilcher:2020:NIB

[Dil20] Karl Dilcher. Nonlinear identities for Bernoulli and Euler polynomials. In Bailey et al. [BBB^{+}20], pages 369–376. ISBN 3-030-36567-0 (print), 3-030-36568-9 (e-book). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Dilcher:2021:ACC

[Dil21] Karl Dilcher. Analytic continuations of character and alternating Tornheim zeta functions. American Mathematical Monthly, 128 (9):780–795, 2021. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Dai:2016:BBC

[DK16] YuHong Dai and CaiXia Kou. A Barzilai–Borwein conjugate gradient method. Science China. Mathematics, 59(8):1511–1524, July 2016. ISSN 1674-7283 (print), 1869-1862 (electronic).

Dai:2002:LCB

[DL02] Yu-Hong Dai and Li-Zhi Liao. R-linear convergence of the Barzilai and Borwein gradient method. IMA Journal of Numerical Analysis, 22(1):1–10, January 2002. CODEN IJNADN. ISSN 0272-4979 (print), 1464-3642 (electronic). URL http:

//www3.oup.co.uk/imanum/hdb/Volume_22/Issue_01/220001.
sgm.abs.html;http://www3.oup.co.uk/imanum/hdb/Volume_
22/Issue_01/pdf/220001.pdf.

Dai:2005:ABB

[DLL05] Yu-Hong Dai, Li-Zhi Liao, and Duan Li. An analysis of the Barzilai and Borwein gradient method for unsymmetric linear equations. In Optimization and control with applications, volume 96 of Appl. Optim., pages 183–211. Springer, New York, 2005.

DiazMillan:2020:CAR

[DLR20] Reinier Díaz Millán, Scott B. Lindstrom, and Vera Roshchina. Comparing averaged relaxed cutters and projection methods: Theory and examples. In Bailey et al. [BBB^{+}20], pages 75–98. ISBN 3-030-36567-0 (print), 3-030-36568-9 (e-book). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Deutsch:1986:PSE

[DNG^{+}86] Emeric Deutsch, L. I. Nicolaescu, Ira Gessel, Paul Monsky, Clark Kimberling, and Hüseyin Demir. Problems and solutions: Elementary problems: E3159–E3164. American Mathematical Monthly, 93(7):565–566, August/September 1986. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [DAK88, DBCB88, GC88, KC89, Mon89, NJS88].

Daniilidis:2018:PAD

[DP18] Aris Daniilidis and Colin Petitjean. A partial answer to the Demyanov–Ryabova conjecture. Set-Valued and Variational Analysis, 26(1):143–157, March 2018. CODEN ???? ISSN 1877-0533 (print), 1877-0541 (electronic). URL http://link.springer.

com/article/10.1007/s11228-017-0439-2.

dePrado:2021:FST

[dPB21] Marcos López de Prado and David H. Bailey. The false strategy theorem: A financial application of experimental mathematics. American Mathematical Monthly, 128(9):825–831, 2021. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Dolgov:2020:PCI

[DS20] Sergey Dolgov and Dmitry Savostyanov. Parallel cross interpolation for high-precision calculation of high-dimensional integrals. Computer Physics Communications, 246(??):Article 106869, January 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465519302565.

Eberhard:2008:SOC

[EB08] A. C. Eberhard and J. M. Borwein. Second order cones for maximal monotone operators via representative functions. Set-Valued Analysis, 16(2–3):157–184, 2008. CODEN SVANEG. ISSN 0927-6947 (print), 1572-932X (electronic). URL http://docserver.

carma.newcastle.edu.au/348/;http://link.springer.com/
article/10.1007/s11228-008-0075-y.

Engquist:2001:MUB

[ES01] Björn Engquist and Wilfried Schmid, editors. Mathematics unlimited: 2001 and beyond. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-66913-2. xv + 1237 pp. LCCN QA7 .M32423 2001.

Erdos:1986:PSS

[EWM86] P. Erdős, J. B. Wilker, and L. E. Mattics. Problems and solutions: Solutions of elementary problems: E3000. American Mathematical Monthly, 93(7):571–572, August/September 1986. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [ANO^{+}83].

Fabian:1989:STL

[Fab89] Marián Fabián. Subdifferentiability and trustworthiness in the light of a new variational principle of Borwein and Preiss. Acta Univ. Carolin. Math. Phys., 30(2):51–56, 1989. CODEN AUMMBZ. ISSN 0001-7140. 17th Winter School on Abstract Analysis (Srní, 1989).

Faltings:1996:MM

[Fal96] Gerd Faltings. Moderne Mathematik. Verständliche Forschung. Spectrum Akademischer Verlag, Heidelberg, Germany, 1996. ISBN 3-8274-0025-2. 183 pp. LCCN ????

Ferris:1991:WTP

[Fer91] Timothy Ferris, editor. The World Treasury of Physics, Astronomy, and Mathematics. Little, Brown and Co., Boston, MA, USA, 1991. ISBN 0-316-28129-8. xv + 859 pp. LCCN QC71 .W67 1991. With a foreword by Clifton Fadiman, general editor.

Finch:1995:MBB

[Fin95] Steven Finch. The miraculous Bailey–Borwein–Plouffe pi algorithm. Recent URLs redirect to an unrelated site, but the one given here worked on 26-Apr-2011., October 1, 1995. URL

http://replay.web.archive.org/20020917121814/;http://
www.mathsoft.com/ASOLVE/plouffe/plouffe.html.

Foster:2000:IBE

[FK00] William Foster and Ilia Krasikov. An improvement of a Borwein– Erdélyi–Kós result. Methods and Applications of Analysis, 7(4): 605–614, 2000. ISSN 1073-2772 (print), 1945-0001 (electronic).

Fletcher:2005:BBM

[Fle05] Roger Fletcher. On the Barzilai–Borwein method. In Optimization and control with applications, volume 96 of Appl. Optim., pages 235–256. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005.

Fang:2015:DSF

[FN15] Xiaowei Fang and Qin Ni. A direct search frame-based adaptive Barzilai–Borwein method. Journal of Computational Mathematics, 33(2):179–190, 2015. CODEN JCMMEB. ISSN 0254-9409 (print), 1991-7139 (electronic).

Ganz:2014:DES

[Gan14] Reinhard E. Ganz. The decimal expansion of π is not statistically random. Experimental Mathematics, 23(2):99–104, 2014. CODEN ???? ISSN 1058-6458 (print), 1944-950X (electronic). See the reproduction of results, and reanalysis, in [BBBR16], that reveals a flaw in the statistical analysis in this paper: Ganz used only a single blocksize in sampling digits, and that blocksize produces anomalous statistics.

Ganz:2017:RRC

[Gan17] Reinhard E. Ganz. Reply to “Reproducibility in Computational Science: A Case Study: Randomness of the Digits of Pi” [Bailey et al. 17]. Experimental Mathematics, 26(3):306– 307, 2017. CODEN ???? ISSN 1058-6458 (print), 1944-950X

(electronic). URL http://www.tandfonline.com/doi/full/10.
1080/10586458.2016.1173413. See [BBBR17].

Gessel:1988:PSSb

[GC88] Ira Gessel and David Callan. Problems and solutions: Solutions of elementary problems: E3161. American Mathematical Monthly, 95(8):770–771, October 1988. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [DNG^{+}86].

Gao:2015:BBL

[GDT15] Huan Gao, Yu-Hong Dai, and Xiao-Jiao Tong. Barzilai–Borweinlike methods for the extreme eigenvalue problem. J. Ind. Manag. Optim., 11(3):999–1019, 2015. ISSN 1547-5816 (print), 1553-166X (electronic).

Georgiev:2005:PBP

[Geo05] Pando Gr. Georgiev. Parametric Borwein–Preiss variational principle and applications. Proceedings of the American Mathematical Society, 133(11):3211–3225, November 2005. CODEN PAM-YAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL

https://www.jstor.org/stable/4097574.

Gourevitch:2007:CBS

[GG07] Boris Gourévitch and Jesús Guillera Goyanes. Construction of binomial sums for π and polylogarithmic constants inspired by BBP formulas. Applied Mathematics E-Notes, 7:237–246,

2007. ISSN 1607-2510. URL http://www.math.nthu.edu.tw/
%7Eamen/2007/061028-2.pdf.

Garvan:2001:SCN

[GI01] Frank (Frank G.) Garvan and Mourad Ismail, editors. Symbolic Computation, Number Theory, Special Functions, Physics, and Combinatorics, volume 4 of Developments in mathematics. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 1-4020-0101-0. x + 283 pp. LCCN QA295 .S86 2001.

Giladi:2018:RCD

[Gil18] Ohad Giladi. A remark on the convergence of the Douglas– Rachford iteration in a non-convex setting. Set-Valued and Variational Analysis, 26(2):207–225, June 2018. CODEN ???? ISSN 1877-0533 (print), 1877-0541 (electronic). URL http://link.

springer.com/article/10.1007/s11228-018-0477-4.

Garcia-Lirola:2018:MRN

[GLR18] Luis García-Lirola and Matías Raja. Maps with the Radon– Nikodým property. Set-Valued and Variational Analysis, 26(1): 77–93, March 2018. CODEN ???? ISSN 1877-0533 (print), 1877-

0541 (electronic). URL http://link.springer.com/article/
10.1007/s11228-017-0428-5.

Galicki:2016:CAB

[GN16] Aleksander Galicki and André Nies. A computational approach to the Borwein–Ditor theorem. In Arnold Beckmann, Laurent Bienvenu, and Natasša Jonoska, editors, Pursuit of the Universal: 12th Conference on Computability in Europe, CiE 2016, Paris, France, June 27–July 1, 2016, Proceedings, volume 9709 of Lecture Notes in Computer Science, pages 99–104. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016.

Goos:2020:CBF

[Goo20] Merrilyn Goos. Crossing boundaries: Fostering collaboration between mathematics educators and mathematicians in initial teacher education programmes. In Bailey et al. [BBB^{+}20], pages 141–148. ISBN 3-030-36567-0 (print), 3-030-36568-9 (e-book). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Galvin:1997:PSP

[GRM^{+}97] Fred Galvin, Franz Rothe, Wen-Xiu Ma, Dan Sachelarie, Vlad Sachelarie, Yury J. Ionin, Robin R. Lewis, Joseph Rosenblatt, Jonathan M. Borwein, and C. G. Pinner. Problems and solutions: Problems: 10599–10605. American Mathematical Monthly, 104 (6):566–567, 1997. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Grippo:2002:NGT

[GS02] L. Grippo and M. Sciandrone. Nonmonotone globalization techniques for the Barzilai–Borwein gradient method. Computational optimization and applications, 23(2):143–169, 2002. CODEN CPPPEF. ISSN 0926-6003 (print), 1573-2894 (electronic).

Gold:2008:POD

[GS08] Bonnie Gold and Roger A. Simons, editors. Proof and Other Dilemmas: Mathematics and Philosophy. Spectrum series. Mathematical Association of America, Washington, DC, USA, 2008. ISBN 0-88385-567-4. xxxii + 346 pp.

LCCN QA8.4 .P755 2008. URL http://www.loc.gov/catdir/
enhancements/fy0916/2008922718-b.html;http://www.loc.
gov/catdir/enhancements/fy0916/2008922718-d.html;http:
//www.loc.gov/catdir/enhancements/fy0916/2008922718-t.
html.

Guillera:2008:EPS

[Gui08] Jesús Guillera. Easy proofs of some Borwein algorithms for π. American Mathematical Monthly, 115(9):850–854, November 2008. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL https://www.jstor.org/stable/27642614.

Guillera:2016:NPB

[Gui16] Jesús Guillera. New proofs of Borwein-type algorithms for pi. Integral Transforms and Special Functions, 27(10):775–782, June 2016. CODEN ???? ISSN 1065-2469 (print), 1476-8291 (electronic).

Guillera:2017:SRB

[Gui17] Jesús Guillera. Self-replication and Borwein-like algorithms. The Ramanujan Journal, August 2017. ISSN 1382-4090 (print), 1572-9303 (electronic).

Hong:2009:NBC

[HC09] Shao Fang Hong and Wei Cao. Notes on the Borwein–Choi conjecture of Littlewood cyclotomic polynomials. Acta Math. Sin. (Engl. Ser.), 25(1):65–76, 2009. ISSN 1439-7617 (print), 1439-8516 (electronic).

Hu:2007:IBB

[HD07] Yi-Qing Hu and Yu-Hong Dai. Inexact Barzilai–Borwein method for saddle point problems. Numerical Linear Algebra with Applications, 14(4):299–317, 2007. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Hanna:2012:PPM

[Hd12] Gila Hanna and Michael de Villiers, editors. Proof and proving in mathematics education: the 19th ICMI study, volume 15 of New ICMI study series. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 94-007-2128-5, 94-007-2129-3 (e-book). xii + 475 pp. LCCN QA9.54

.P766 2012. URL http://www.springerlink.com/content/
978-94-007-2129-6.

Higham:2015:PCA

[HDG^{+}15] Nicholas J. Higham, Mark R. Dennis, Paul Glendinning, Paul A. Martin, Fadil Santosa, and Jared Tanner, editors. The Princeton Companion to Applied Mathematics. Princeton University Press, Princeton, NJ, USA, 2015. ISBN 0-691-15039-7 (hardcover). xvii + 3 + 994 + 16 pp. LCCN QA155 .P75 2015. URL http:

//press.princeton.edu/titles/10592.html.

Huang:2021:EBB

[HDL21] Ya-Kui Huang, Yu-Hong Dai, and Xin-Wei Liu. Equipping the Barzilai–Borwein method with the two dimensional quadratic termination property. SIAM Journal on Optimization, 31(4):3068– 3096, ???? 2021. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

Hersh:2005:BRM

[HF05] Reuben Hersh and Roger Frye. Book review: Mathematics by Experiment: Plausible Reasoning in the 21st Century. SIAM Review, 47(4):832–833, December 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https://www.jstor.

org/stable/20453726.

Hirschhorn:1993:CAJ

[HGB93] Michael Hirschhorn, Frank Garvan, and Jon Borwein. Cubic analogues of the Jacobian theta function θ(z, q). Canadian Journal of Mathematics = Journal canadien de mathématiques, 45 (4):673–694, ???? 1993. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). URL http://docserver.carma.

newcastle.edu.au/1557/.

Hirschhorn:2017:CTF

[Hir17] Michael D. Hirschhorn. The cubic theta-function analogues of Borwein, Borwein and Garvan. In The Power of q, pages 185– 203. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017.

Huang:2015:BBT

[HL15a] Yakui Huang and Hongwei Liu. A Barzilai–Borwein type method for minimizing composite functions. Numerical Algorithms, 69 (4):819–838, 2015. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Huang:2015:RCP

[HL15b] Yakui Huang and Hongwei Liu. On the rate of convergence of projected Barzilai–Borwein methods. Optimization Methods and Software, 30(4):880–892, 2015. ISSN 1026-7670 (print), 1055-6788 (electronic).

Huang:2016:SPC

[HLY16] Yakui Huang, Hongwei Liu, and Tengteng Yu. Smoothing projected cyclic Barzilai–Borwein method for stochastic linear complementarity problems. International Journal of Computer Mathematics, 93(7):1188–1199, 2016. CODEN IJCMAT. ISSN 0020-7160.

Huang:2014:BBT

[HLZ14] Yakui Huang, Hongwei Liu, and Sha Zhou. A Barzilai–Borwein type method for stochastic linear complementarity problems. Numerical Algorithms, 67(3):477–489, 2014. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Huang:2015:EMP

[HLZ15a] Yakui Huang, Hongwei Liu, and Sha Zhou. An efficient monotone projected Barzilai–Borwein method for nonnegative matrix factorization. Applied Mathematics Letters, 45:12–17, 2015. CO-DEN AMLEEL. ISSN 0893-9659 (print), 1873-5452 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0893965915000087.

Huang:2015:QRP

[HLZ15b] Yakui Huang, Hongwei Liu, and Shuisheng Zhou. Quadratic regularization projected Barzilai–Borwein method for nonnegative matrix factorization. Data Min. Knowl. Discov., 29(6):1665–1684, 2015. ISSN 1384-5810 (print), 1573-756X (electronic).

Hussain:2020:MTS

[HMM20] Mumtaz Hussain, Seyyed Hassan Mahboubi, and Abolfazl Seyed Motahari. Metrical theory for small linear forms and applications to interference alignment. In Bailey et al. [BBB^{+}20], pages 377– 393. ISBN 3-030-36567-0 (print), 3-030-36568-9 (e-book). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Han:2009:APB

[HNP10] Lixing Han, Michael Neumann, and Upendra Prasad. Alternating projected Barzilai–Borwein methods for nonnegative matrix factorization. Electronic Transactions on Numerical Analysis, 36: 54–82, 2009/10. ISSN 1068-9613 (print), 1097-4067 (electronic).

URL http://etna.mcs.kent.edu/volumes/2001-2010/vol36/
abstract.php?vol=36&pages=54-82.

Hoare:2005:BRM

[Hoa05] Graham Hoare. Book review: Mathematics by Experiment, by Jonathan M. Borwein, David H. Bailey. Plausible Reasoning in the 21st Century. Mathematical Gazette, 89(514):143–144, March 2005. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.jstor.org/stable/3620673.

Holmes:2020:MEC

[Hol20] Kathryn Holmes. Mathematics education in the computational age: Challenges and opportunities. In Bailey et al. [BBB^{+}20], pages 149–152. ISBN 3-030-36567-0 (print), 3-030-36568-9 (ebook). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Honor:1985:DTR

[Hon85] R. B. Honor. Density and transitivity results on l^{∞} and l^{1}. Journal of the London Mathematical Society (series 2), S2-32(3):521– 527, December 1985. CODEN JLMSAK. ISSN 0024-6107 (print),

1469-7750 (electronic). URL http://jlms.oxfordjournals.
org/content/s2-32/3/521.full.pdf+html. The ‘author’ is an

acronym for the family names of seven people: Heydar Radjavi, Jon Borwein, Don Hadwin, Richard O’Brien, Eric Nordgren, Mehmet Orhon, and Peter Rose.

Howlett:2014:BRC

[How14] Phil Howlett. Book review: Convex functions: Constructions, characterizations and counterexamples, by Jonathan M. Borwein and Jon D. Vanderwerff. Australian Mathematical Society Gazette, 41(3):173–177, July 2014. ISSN 0311-0729 (print),

1326-2297 (electronic). URL http://www.austms.org.au/Publ/
Gazette/2014/Jul14/BkRev.pdf.

Hong:2014:MSI

[HY14] Hoon Hong and Chee-Keng Yap, editors. Mathematical Software — ICMS 2014: 4th International Conference, Seoul, South Korea, August 5–9, 2014. Proceedings, volume 8592 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-662-44198-5, 3-662-44199-3 (e-book). LCCN QA76.95.

Han:2009:MBB

[HYG09] Le Han, Gao Hang Yu, and Lü Tai Guan. Multivariate Barzilai– Borwein method and its application in elastic registration of medical image. Acta Scientiarum Naturalium Universitatis Sunyatseni. Zhongshan Daxue Xuebao. Ziran Kexue Ban, 48(3):8–12, 21, 2009. CODEN CHTHAJ. ISSN 0529-6579.

IEEE:2008:HIS

[IEE08] IEEE, editor. HPCS 2006: [20th International Symposium on High-Performance Computing in an Advanced Collaborative Environment]; 20th anniversary of HPCS; hosted by ACEnet at Memorial University of Newfoundland, [St. John’s], May 14th– 17th 2006; [co-hosted with OSCAR 2006]. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2008. ISBN 0-7695-2582-2, 0-7695-3200-4. LCCN QA76.88 .I5858 2006.

Ioffe:1992:ONA

[IMR92] Aleksandr Davidovich Ioffe, Moshe Marcus, and Simeon Reich, editors. Optimization and nonlinear analysis: [proceedings of a Binational Workshop on Optimization and Nonlinear Analysis, held at Technion City, Haifa, 21–27 March 1990], volume 244 of Pitman research notes in mathematics series. Longman Scientific and Technical, Harlow, Essex, UK, 1992. ISBN 0-582-08065-7, 0-470-21943-2. ISSN 0269-3674. LCCN QA402.5 .O6424 1991.

Iannazzo:2017:RBB

[IP17] Bruno Iannazzo and Margherita Porcelli. The Riemannian Barzilai–Borwein method with nonmonotone line search and the matrix geometric mean computation. IMA Journal of Numerical Analysis, 38(1):495–517, April 2017. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic).

Iannazzo:2018:RBB

[IP18] Bruno Iannazzo and Margherita Porcelli. The Riemannian Barzilai–Borwein method with nonmonotone line search and the matrix geometric mean computation. IMA Journal of Numerical Analysis, 38(1):495–517, January 25, 2018. CODEN IJNADH. ISSN 0272-4979 (print), 1464-3642 (electronic). URL http:

//academic.oup.com/imajna/article/38/1/495/3573895.

Jackson:2009:JER

[Jac09] Allyn Jackson. John Ewing retires from the AMS. Notices of the American Mathematical Society, 56(1):48–51, January 2009. CODEN AMNOAN. ISSN 0002-9920 (print),

1088-9477 (electronic). URL https://www.ams.org/journals/
notices/200901/200901-full-issue.pdf. With contributions

by Jonathan Borwein and Jane Kister. Borwein’s is a sidebar on page 49, with the title Some Reflections on John Ewing.

Jalilian:2024:TVM

[Jal24] Hamed Jalilian. Total variation method based on modified Barzilai–Borwein algorithm to noise reduction in MRI images. The Journal of Supercomputing, 80(1):601–619, January 2024. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (elec-

tronic). URL https://link.springer.com/article/10.1007/
s11227-023-05500-z.

Jungic:2021:EMM

[JB21] Veselin Jungić and Andrijana Burazin. On experimental mathematics and mathematics education. American Mathematical Monthly, 128(9):832–844, 2021. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Jiang:2013:FBB

[JD13] Bo Jiang and Yu-Hong Dai. Feasible Barzilai–Borwein-like methods for extreme symmetric eigenvalue problems. Optimization Methods and Software, 28(4):756–784, 2013. ISSN 1026-7670 (print), 1055-6788 (electronic).

Jungic:2020:DVM

[JJ20] Damir Jungić and Veselin Jungić. Dynamic visual models: Ancient ideas and new technologies. In Bailey et al. [BBB^{+}20], pages 189–201. ISBN 3-030-36567-0 (print), 3-030-36568-9 (e-book). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Jimenez:2003:NLP

[JN03] B. Jiménez and V. Novo. A notion of local proper efficiency in the Borwein sense in vector optimisation. The ANZIAM Journal, 45 (1):75–89, July 2003. CODEN AJNOA2. ISSN 1446-1811 (print),

1446-8735 (electronic). URL https://www.cambridge.org/
core/journals/anziam-journal/article/notion-of-local-proper-efficiency-in-the-borwein-sense-in-vector-optimisation/
54FFD574CEC237456161E9A9A39D8D3B.

James:2014:STR

[JWDS^{+}14] D. James, N. Wilkins-Diehr, V. Stodden, D. Colbry, C. Rosales, M. Fahey, J. Shi, R. F. Silva, K. Lee, R. Roskies, L. Loewe, S. Lindsey, R. Kooper, L. Barba, D. Bailey, J. Borwein, O. Corcho, E. Deelman, M. Dietze, B. Gilbert, J. Harkes, S. Keele, P. Kumar, J. Lee, E. Linke, R. Marciano, L. Marini, C. Mattman, D. Mattson, K. McHenry, R. McLay, S. Miguez, B. Minsker, M. Perez-Hernandez, D. Ryan, M. Rynge, I. Santana-Perez, M. Satyanarayanan, G. St.Clair, K. Webster, E. Hovig, D. S. Katz, S. Kay, G. Sandve, D. Skinner, G. Allen, J. Cazes, K. W. Cho, J. Fonseca, L. Hwang, L. Koesterke, P. Patel, L. Pouchard, E. Seidel, and I. Suriarachchi. Standing together for reproducibility in largescale computing: Report on reproducibility@XSEDE. ArXiv e-

prints, December 2014. URL http://adsabs.harvard.edu/abs/
2014arXiv1412.5557J.

Jiang:2012:BBB

[JY12] Jianlin Jiang and Xiaoming Yuan. A Barzilai–Borwein-based heuristic algorithm for locating multiple facilities with regional demand. Computational optimization and applications, 51(3): 1275–1295, 2012. CODEN CPPPEF. ISSN 0926-6003 (print), 1573-2894 (electronic).

Kennedy:2015:CAM

[KAA^{+}15] Stephen F. Kennedy, Donald J. Albers, Gerald L. Alexanderson, Della Dumbaugh, Frank A. Farris, Deanna B. Haunsperger, and Paul Zorn, editors. A Century of Advancing Mathematics. Mathematical Association of America, Washington, DC, USA, 2015. ISBN 0-88385-588-7 (hardcover), 1-61444-522-2 (e-book). xi + 423 pp. LCCN QA26 .C46 2015.

Kimberling:1989:PSS

[KC89] Clark Kimberling and David Callan. Problems and solutions: Solutions of elementary problems: E3163. American Mathematical Monthly, 96(1):57, January 1989. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [DNG^{+}86].

Kisilevsky:2004:NTP

[KG04] Hershy Kisilevsky and Eyal Z. (Eyal Zvi) Goren, editors. Number theory. Papers from the 7th conference of the Canadian Number Theory Association, University of Montreal, Montreal, QC, Canada, May 19–25, 2002, volume 36 of CRM proceedings and lecture notes. American Mathematical Society, Providence, RI, USA, 2004. ISBN 0-8218-3331-6. LCCN QA241 .N8645 2004.

Kimberling:1986:PSS

[KJ86] Clark Kimberling and A. A. Jagers. Problems and solutions: Solutions of elementary problems: E2998. American Mathematical Monthly, 93(9):733–734, November 1986. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [ANO^{+}83].

Krejic:2016:BBM

[KJR16] Nataša Krejić, Nataša Krklec Jerinkić, and Sanja Rapajić. Barzilai–Borwein method with variable sample size for stochastic linear complementarity problems. Optimization, 65(2):479– 499, 2016. CODEN OPTZDQ. ISSN 0233-1934, 0323-3898. URL

http://arxiv.org/abs/1312.7323.

Kortenkamp:2016:JMB

[KMT16] Ulrich Kortenkamp, John Monaghan, and Luc Trouche. Jonathan M. Borwein (1951–2016): exploring, experiencing and experimenting in mathematics — an inspiring journey in mathematics. Educational Studies in Mathematics, 93(2):131–136, October 2016. CODEN EDSMAN. ISSN 0013-1954 (print), 1573-0816

(electronic). URL http://link.springer.com/accesspage/
article/10.1007/s10649-016-9729-0.

Kitazume:2000:BIV

[KMY00] Masaaki Kitazume, Masahiko Miyamoto, and Hiromichi Yamada. Borwein identity and vertex operator algebras. Journal of Number Theory, 82(1):100–108, 2000. CODEN JNUTA9. ISSN 0022-314X (print), 1096-1658 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022314X99924824.

Kellar:2003:DCM

[KMZ^{+}03] Melanie Kellar, Bonnie MacKay, Rui Zhang, Carolyn Watters, David Kaufman, and Jonathan Borwein. Dynamic composition of math lessons. Journal of Educational Technology & Society, 6(4):100–111, ???? 2003. ISSN 1176-3647 (print), 1436-

4522 (electronic). URL http://docserver.carma.newcastle.

edu.au/1525/;http://ifets.ieee.org/periodical/6_4/10.
pdf;https://www.jstor.org/stable/jeductechsoci.6.4.100.

Kohlenbach:2001:QVT

[Koh01] Ulrich Kohlenbach. A quantitative version of a theorem due to Borwein–Reich–Shafrir. Numerical Functional Analysis and Optimization, 22(5–6):641–656, 2001. CODEN NFADOL. ISSN 0163-0563 (print), 1532-2467 (electronic).

Komatsu:2000:IDA

[Kom00] Takao Komatsu. On inhomogeneous Diophantine approximation and the Borweins’ algorithm. Sūrikaisekikenkyūsho Kōkyūroku, 1160:155–162, 2000. Analytic number theory and related topics (Japanese) (Kyoto, 1999).

Komatsu:2002:IDA

[Kom02] Takao Komatsu. On inhomogeneous Diophantine approximation and the Borweins’ algorithm. II. In Analytic number theory (Beijing/Kyoto, 1999), volume 6 of Dev. Math., pages 223–242. Kluwer Acad. Publ., Dordrecht, 2002.

Komatsu:2004:IDA

[Kom04] Takao Komatsu. On inhomogeneous Diophantine approximation and the Borweins’ algorithm. Far East J. Math. Sci. (FJMS), 12 (2):203–224, 2004. ISSN 0972-0871. URL http://www.pphmj.

com/abstract/137.htm.

Kruger:2016:BPV

[KPS16] A. Y. Kruger, S. Plubtieng, and T. Seangwattana. Borwein– Preiss variational principle revisited. Journal of Mathematical Analysis and Applications, 435(2):1183–1193, 2016. CO-DEN JMANAK. ISSN 0022-247X (print), 1096-0813 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022247X15010343.

Kruger:2017:BPV

[KPS17] A. Y. Kruger, S. Plubtieng, and T. Seangwattana. Borwein–Preiss vector variational principle. Positivity, January 2017. CODEN ???? ISSN 1385-1292 (print), 1572-9281 (electronic).

Kruger:2018:AIT

[Kru18] Alexander Y. Kruger. About intrinsic transversality of pairs of sets. Set-Valued and Variational Analysis, 26(1):111–142, March 2018. CODEN ???? ISSN 1877-0533 (print), 1877-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s11228-017-0446-3.

Knuth:1990:PSSa

[KWK^{+}90a] D. E. Knuth, Jungmin Woo, Hans Kappus, J. M. Borwein, and P. B. Borwein. Problems and solutions: Solutions of elementary problems: E3335. American Mathematical Monthly, 97(10):927, 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL 2-8&origin=MSN;http://links.jstor.org/
sici?sici=0002-9890(199012)97:10<927:E>2.0.CO.

Knuth:1990:PSSb

[KWK^{+}90b] D. E. Knuth, Jungmin Woo, Hans Kappus, J. M. Borwein, and P. B. Borwein. Problems and solutions: Solutions of elementary problems: E3335. American Mathematical Monthly, 97(10):927, 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL 2-8&origin=MSN;http://links.jstor.org/
sici?sici=0002-9890(199012)97:10<927:E>2.0.CO.

Knuth:1990:PSSc

[KWK^{+}90c] D. E. Knuth, Jungmin Woo, Hans Kappus, J. M. Borwein, and P. B. Borwein. Problems and solutions: Solutions of elementary problems: E3335. American Mathematical Monthly, 97(10):927, December 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Rud89].

LaCruz:2009:EBB

[La 09] William La Cruz. Extension of the Barzilai–Borwein method for quadratic forms in finite Euclidean spaces. Numerical Functional Analysis and Optimization, 30(3–4):306–321, 2009. CO-DEN NFADOL. ISSN 0163-0563 (print), 1532-2467 (electronic).

Lassonde:2018:USF

[Las18] Marc Lassonde. Upper semismooth functions and the subdifferential determination property. Set-Valued and Variational Analysis, 26(1):95–109, March 2018. CODEN ???? ISSN 1877-0533 (print), 1877-0541 (electronic). URL http://link.springer.

com/article/10.1007/s11228-017-0467-y.

Li:2015:SNB

[Li15] Xiangli Li. Smoothing nonmonotone Barzilai–Borwein gradient method and its application to stochastic linear complementarity problems. Mathematical Problems in Engineering, 2015: 425351, 2015. ISSN 1024-123X. URL http://www.hindawi.com/

journals/mpe/2015/425351/.

Liu:2000:BCT

[Liu00] Zhi-Guo Liu. The Borweins’ cubic theta function identity and some cubic modular identities of Ramanujan. The Ramanujan Journal, 4(1):43–50, 2000. CODEN RAJOF9. ISSN 1382-4090 (print), 1572-9303 (electronic).

Liu:2001:SES

[Liu01] Zhi-Guo Liu. Some Eisenstein series identities associated with the Borwein functions. In Garvan and Ismail [GI01], pages 147–169. ISBN 1-4020-0101-0. LCCN QA295 .S86 2001.

Lewis:2001:BCT

[LL01] Richard Lewis and Zhi-Guo Liu. The Borweins’ cubic theta functions and q-elliptic functions. In Garvan and Ismail [GI01], pages 133–145. ISBN 1-4020-0101-0. LCCN QA295 .S86 2001.

Lin:2009:PPM

[lL09] Fou lai Lin, editor. Proof and Proving in Mathematics Education: ICMI Study 19 Conference Proceedings = shu xue zheng ming yu lun zheng de jiao yu xue. Department of Mathematics, National Taiwan Normal University, Taipei, Taiwan, 2009. ISBN 986-01-8210-8. LCCN QA109; QA11.A1; QA9.54. 2 volumes.

Liu:2013:MSB

[LL13] Hongwei Liu and Xiangli Li. Modified subspace Barzilai–Borwein gradient method for non-negative matrix factorization. Computational optimization and applications, 55(1):173–196, 2013. CO-DEN CPPPEF. ISSN 0926-6003 (print), 1573-2894 (electronic).

Limber:1995:DRF

[LLC^{+}95] M. A. Limber, M. N. Limber, A. Celler, J. S. Barney, and J. M. Borwein. Direct reconstruction of functional parameters for dynamic SPECT. IEEE Transactions on Nuclear Science, 42(4):1249–1256, August 1995. CODEN IRNSAM. ISSN 0018-9499 (print), 1558-1578 (electronic). URL http://adsabs.

harvard.edu/abs/1995ITNS...42.1249L;http://docserver.
carma.newcastle.edu.au/1535/.

Li:2011:FSM

[LLS11] Xiangli Li, Hongwei Liu, and Xiaojun Sun. Feasible smooth method based on Barzilai–Borwein method for stochastic linear complementarity problem. Numerical Algorithms, 57(2):207–215, 2011. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Lauster:2018:SCM

[LLT18] Florian Lauster, D. Russell Luke, and Matthew K. Tam. Symbolic computation with monotone operators. Set-Valued and Variational Analysis, 26(2):353–368, June 2018. CODEN ???? ISSN 1877-0533 (print), 1877-0541 (electronic). URL http://link.

springer.com/article/10.1007/s11228-017-0418-7.

Lord:1990:BRD

[Lor90] Nick Lord. Book review: A Dictionary of Real Numbers. Mathematical Gazette, 74(470):395–396, December 1990. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.jstor.org/stable/3618159.

Lorentzen:2008:CDR

[Lor08] Lisa Lorentzen. Convergence and divergence of the Ramanujan AGM fraction. The Ramanujan Journal, 16(1):83–95, May 2008. ISSN 1382-4090 (print), 1572-9303 (electronic).

Lord:2009:BRE

[Lor09] Nick Lord. Book review: Experimental Mathematics in Action. Mathematical Gazette, 93(528):564–566, November 2009. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.jstor.org/stable/40378594.

Lindqvist:2001:PSS

[LPB01] Peter Lindqvist, Jaak Peetre, and Jonathan M. Borwein. Problems and solutions: Solutions: Generalized trigonometric functions: 10744. American Mathematical Monthly, 108(5):473–474, 2001. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Li:2000:GEV

[LS00] Yongxin Li and Shuzhong Shi. A generalization of Ekeland’s ϵvariational principle and its Borwein–Preiss smooth variant. Journal of Mathematical Analysis and Applications, 246(1):308–319, 2000. CODEN JMANAK. ISSN 0022-247X (print), 1096-0813 (electronic).

Lupas:2002:SBF

[Lup02] Alexandru Lupaş. Some BBP-functions. Univ. Beograd. Publ. Elektrotehn. Fak. Ser. Mat., 13:26–29 (2003), 2002. ISSN 0353-8893 (print), 2406-0852 (electronic).

Li:2018:NAB

[LW18] Ting Li and Zhong Wan. New adaptive Barzilai–Borwein step size and its application in solving large-scale optimization problems. The ANZIAM Journal, 61(1):76–98, December 2018. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (electronic).

Li:2019:NAB

[LW19] Ting Li and Zhong Wan. New adaptive Barzilai–Borwein step size and its application in solving large-scale optimization problems. The ANZIAM Journal, 61(1):76–98, January 2019. CODEN AJNOA2. ISSN 1446-1811 (print), 1446-8735 (elec-

tronic). URL https://www.cambridge.org/core/journals/
anziam-journal/article/new-adaptive-barzilaiborwein-step-size-and-its-application-in-solving-largescale-optimization-problems/
ACD9681C2B8C9F20A84A62658C3DEAB8.

Li:2018:NOC

[LY18] An Li and Jane J. Ye. Necessary optimality conditions for implicit control systems with applications to control of differential algebraic equations. Set-Valued and Variational Analysis, 26(1):179– 203, March 2018. CODEN ???? ISSN 1877-0533 (print), 1877-

0541 (electronic). URL http://link.springer.com/article/
10.1007/s11228-017-0444-5.

Li:2021:DSP

[LY21] Pei Heng Li and Hee Yong Youn. Distributed stochastic principal component analysis using stabilized Barzilai–Borwein stepsize for data compression with WSN. The Journal of Supercomputing, 77(10):11032–11051, October 2021. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL https://

link.springer.com/article/10.1007/s11227-021-03707-6.

Li:2014:PBB

[LZ14] Xiang Li Li and Sha Zhou. Projected Barzilai–Borwein method for a class of stochastic linear complementarity problems. Acta Math. Appl. Sin., 37(2):278–285, 2014. ISSN 0254-3079.

Marcos:1991:CPB

[Mar91] Raydan M. Marcos. Convergence properties of the Barzilai and Borwein gradient method. Thesis (ph.d.), Rice University, Houston, TX, USA, 1991. 60 pp. URL http://search.proquest.

com/docview/303929682.

Merca:2015:DII

[Mer15] Mircea Merca. A double inequality involving Erdős–Borwein constants. Miskolc Math. Notes, 16(1):277–281, 2015. ISSN 1787-2405.

Michalos:2003:BTE

[Mic03] Alex C. Michalos, editor. The best teacher I ever had: personal reports from highly productive scholars. Althouse Press, London, ON, Canada, 2003. ISBN 0-920354-53-X. 294 pp. LCCN LA2321 .B47 2003.

Miller:1989:GRB

[Mil89] Harry I. Miller. Generalization of a result of Borwein and Ditor. Proceedings of the American Mathematical Society, 105(4):889– 893, April 1989. CODEN PAMYAR. ISSN 0002-9939 (print),

1088-6826 (electronic). URL https://www.jstor.org/stable/
2047048.

Miller:1989:FER

[Mil90] Harry I. Miller. A further extension of a result of Borwein and Ditor. Real Anal. Exchange, 15(2):713–723, 1989/90. ISSN 0147-1937 (print), 1930-1219 (electronic).

Monsky:1989:PSS

[Mon89] Paul Monsky. Problems and solutions: Solutions of elementary problems: E3162. American Mathematical Monthly, 96(3):258– 259, March 1989. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [DNG^{+}86].

Moors:2018:NCS

[Moo18] Warren B. Moors. Nearly Chebyshev sets are almost convex. Set-Valued and Variational Analysis, 26(1):67–76, March 2018. CODEN ???? ISSN 1877-0533 (print), 1877-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s11228-017-0445-4.

Momeni:2018:NCG

[MP18] M. Momeni and M. R. Peyghami. A new conjugate gradient algorithm with cubic Barzilai–Borwein stepsize for unconstrained optimization. Optimization Methods and Software, ??(??):1–15, January 2018. ISSN 1055-6788.

Morovati:2016:BBM

[MPB16] Vahid Morovati, Latif Pourkarimi, and Hadi Basirzadeh. Barzilai and Borwein’s method for multiobjective optimization problems.

Numerical Algorithms, 72(3):539–604, July 2016. CODEN NUA-LEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//link.springer.com/article/10.1007/s11075-015-0058-7.

Molina:1996:PBB

[MR96] Brigida Molina and Marcos Raydan. Preconditioned Barzilai– Borwein method for the numerical solution of partial differential equations. Numerical Algorithms, 13(1–2):45–60, 1996. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Musev:2011:QJB

[MR11] Boil Musev and Nadezhda Ribarska. On a question of J. Borwein and H. Wiersma. Proceedings of the American Mathematical Society, 139(5):1707–1716, 2011. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic).

Monaghan:2016:TMI

[MTB16] John Monaghan, Luc Trouche, and Jonathan M. Borwein. Tools and Mathematics: Instruments for Learning, volume 110 of Mathematics Education Library. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016. ISBN 3-319-02395-0, 3-319-02396-9 (e-book). ISSN 0924-4921 (print), 2214-983X (electronic). xxi + 483 pp. LCCN QA11.2 .M65 2016. URL

http://docserver.carma.newcastle.edu.au/1717/;http://
sfx.ethz.ch/sfx_locater?sid=ALEPH:EBI01\%26genre=book\
%26isbn=9783319023960;http://swbplus.bsz-bw.de/bsz470387262cov.
htm.

Marechal:1998:APR

[MTCB98] P. Marechal, D. Togane, A. Celler, and J. M. Borwein. Assessment of the performance of reconstruction processes for computed tomography. In 1998 IEEE Nuclear Science Symposium. Conference Record. Toronto, ON, Canada, November 8–14, 1998, volume 2, pages 1353–1357. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISSN

1082-3654. URL http://docserver.carma.newcastle.edu.au/
1529/.

Marechal:1999:CSA

[MTCB99] P. Marechal, D. Togane, A. Celler, and J. M. Borwein. Computation and stability analysis for regularized tomographic reconstructions. IEEE Transactions on Nuclear Science, 46:2177– 2184, December 1999. CODEN IRNSAM. ISSN 0018-9499 (print),

1558-1578 (electronic). URL http://adsabs.harvard.edu/abs/
1999ITNS...46.2177M.

Miller:2012:MBD

[MW12] Harry I. Miller and Leila Miller-Van Wieren. More on the Borwein–Ditor theorem. Sarajevo J. Math., 8(21)(2):367–369, 2012. ISSN 1840-0655 (print), 2233-1964 (electronic).

Marchant:2016:PJM

[MW16] T. R. Marchant and G. A. Willis. Professor Jonathan M. Borwein. Journal of the Australian Mathematical Society, 101(03): 289, November 2016. CODEN JAUMAX. ISSN 0004-9735 (print), 2059-9234 (electronic).

Nosratipour:2017:ANG

[NFB17a] Hadi Nosratipour, Omid Solaymani Fard, and Akbar Hashemi Borzabadi. An adaptive nonmonotone global Barzilai–Borwein gradient method for unconstrained optimization. Optimization, 66(4):641–655, February 2017. CODEN OPTZDQ. ISSN 0233-1934, 0323-3898.

Nosratipour:2017:OCV

[NFB17b] Hadi Nosratipour, Omid Solaymani Fard, and Akbar Hashemi Borzabadi. Optimal control of viscous Burgers equation via an adaptive nonmonotone Barzilai–Borwein gradient method. International Journal of Computer Mathematics, pages 1–16, June 2017. CODEN IJCMAT. ISSN 0020-7160.

Nimbran:2015:TSA

[Nim15] Amrik Singh Nimbran. Taylor series for arctan and BBP-type formulas for π. The Mathematics Student, 84(3–4):39–52, 2015. CODEN MTHSBH. ISSN 0025-5742.

Nicolaescu:1988:PSSb

[NJS88] L. I. Nicolaescu, A. A. Jagers, and Robert E. Shafer. Problems and solutions: Solutions of elementary problems: E3160. American Mathematical Monthly, 95(8):769–770, October 1988. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [DNG^{+}86].

Novinger:1986:PSS

[NOL86] Phil Novinger, Daniel Oberlin, and Nick Lord. Problems and solutions: Solutions of elementary problems: E2996. American Mathematical Monthly, 93(4):303–304, April 1986. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [ANO^{+}83].

Narushima:2009:EBB

[NWY09] Yasushi Narushima, Takahiko Wakamatsu, and Hiroshi Yabe. Extended Barzilai–Borwein method for minimizing a strictly convex quadratic function. In Nonlinear analysis and convex analysis, pages 141–150. Yokohama Publ., Yokohama, 2009.

Narushima:2010:EBB

[NWY10] Yasushi Narushima, Takahiko Wakamatsu, and Hiroshi Yabe. Extended Barzilai–Borwein method for unconstrained minimization problems. Pac. J. Optim., 6(3):591–613, 2010. ISSN 1348-9151 (print), 1349-8169 (electronic).

Octavio:1996:IT

[OBB^{+}96] Alfredo Octavio, Jonathan Borwein, Peter B. Borwein, Roland Girgensohn, and S. Parnes. The “indexed” theorem. The Mathematical Intelligencer, 18(4):9–18, September 1996. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Odlyzko:2011:BRE

[Odl11] Andrew Odlyzko. Book review: Experimental Mathematics in Action, by D. H. Bailey, J. M. Borwein, N. J. Calkin, R. Girgensohn, D. R. Luke, and V. H. Moll. A K Peters, Wellesley, MA, 2007. xii + 322 pp., ISBN 978-1-56881-271-7. $65. American Mathematical Monthly, 118(10):946–951, December 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (elec-

tronic). URL https://www.jstor.org/stable/10.4169/amer.
math.monthly.118.10.946;https://www.jstor.org/stable/
pdfplus/10.4169/amer.math.monthly.118.10.946.pdf.

Osburn:2005:RCB

[Osb05] Robert Osburn. A remark on a conjecture of Borwein and Choi. Proceedings of the American Mathematical Society, 133(10):2903– 2909, October 2005. CODEN PAMYAR. ISSN 0002-9939 (print),

1088-6826 (electronic). URL https://www.jstor.org/stable/
4097904.

Pospisil:2018:PBB

[PD18] Lukáš Pospíšil and Zdeněk Dostál. The projected Barzilai– Borwein method with fall-back for strictly convex QCQP problems with separable constraints. Mathematics and Computers in Simulation, 145:79–89, March 2018. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Pearl:2007:OPT

[Pea07] Elliott Pearl, editor. Open problems in topology II. Elsevier, Amsterdam, The Netherlands, 2007. ISBN 0-444-52208-5, 0-08-047529-9 (e-book). 763 pp. LCCN QA611 .O562 2007.

Piantadosi:2012:CME

[PHB12] Julia Piantadosi, Phil Howlett, and Jonathan Borwein. Copulas with maximum entropy. Optimization Letters, 6(1):99–125, 2012. ISSN 1862-4472 (print), 1862-4480 (electronic). URL http://

docserver.carma.newcastle.edu.au/1394/.

Piantadosi:2013:MSS

[PHB13] Julia Piantadosi, Phil Howlett, and Jonathan Borwein. Modelling and simulation of seasonal rainfall. Report, MODSIM/ASOR, Adelaide, SA, Australia, December 2013. ?? pp. URL http:

//docserver.carma.newcastle.edu.au/1453/.

Piantadosi:2014:MSS

[PHB14] Julia Piantadosi, Phil Howlett, and Jonathan Borwein. Modelling and simulation of seasonal rainfall using checkerboard copulas of maximum entropy. Submitted eMathematics and Computers in Simulation MATCOM, May 2014., 2014. URL http:

//docserver.carma.newcastle.edu.au/1694/.

Piantadosi:2012:MEM

[PHBH12] Julia Piantadosi, Phil Howlett, Jonathan Borwein, and John Henstridge. Maximum entropy methods for generating simulated rainfall. Numerical Algebra, Control and Optimization, 2(2):233–256, ???? 2012. CODEN ???? ISSN 2155-3289 (print), 2155-3297

(electronic). URL http://aimsciences.org/article/doi/10.
3934/naco.2012.2.233.

Piantadosi:2013:GSR

[PHBH13] Julia Piantadosi, Phil Howlett, Jonathan Borwein, and John Henstridge. Generation of simulated rainfall data at different time-scales. In Bean [Bea13], page ?? LCCN ???? URL

http://docserver.carma.newcastle.edu.au/1381/.

Phillips:2020:MEI

[PL20] Collin Phillips and Fu Ken Ly. Mathematics education for indigenous students in preparation for engineering and information technologies. In Bailey et al. [BBB^{+}20], pages 153–169. ISBN 3-030-36567-0 (print), 3-030-36568-9 (e-book). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Pospisil:2013:OAB

[Pos13] Lukáš Pospíšil. An optimal algorithm with Barzilai–Borwein steplength and superrelaxation for QPQC problem. In Programs and algorithms of numerical mathematics 16, pages 155– 161. Acad. Sci. Czech Repub. Inst. Math., Prague, 2013.

Pilehrood:2011:ABB

[PP11] Kh. Hessami Pilehrood and T. Hessami Pilehrood. A qanalogue of the Bailey–Borwein–Bradley identity. Journal of Symbolic Computation, 46(6):699–711, 2011. CODEN JSYCEH. ISSN 0747-7171 (print), 1095-855x (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0747717111000186.

Phillips:1992:SMS

[PR92] Fred Young Phillips and John James Rousseau, editors. Systems and Management Science by Extremal Methods: Research Honoring Abraham Charnes at Age 70. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 1-4613-6599-6 (print), 1-4615-3600-6 (e-book). xxv + 580 pp. LCCN HD30.23.

Petrov:2014:BBM

[PT14] Miroslav S. Petrov and Todor D. Todorov. Barzilai–Borwein method for a nonlocal elliptic problem. Mat. Bilten, 38(2):23– 30, 2014. ISSN 0351-336X.

Platt:2020:IBB

[PT20] Dave Platt and Tim Trudgian. Improved bounds on Brun’s constant. In Bailey et al. [BBB^{+}20], pages 395–406. ISBN 3-030-36567-0 (print), 3-030-36568-9 (e-book). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Qiu:2007:LCL

[QR07] J. H. Qiu and S. Rolewicz. Local completeness of locally pseudoconvex spaces and Borwein–Preiss variational principle. Studia Mathematica, 183(2):99–115, 2007. CODEN SMATAZ. ISSN 0039-3223 (print), 1730-6337 (electronic).

Qiu:2014:NAB

[QYX14] Yuanying Qiu, Jianlei Yan, and Fanyong Xu. Nonmonotone adaptive Barzilai–Borwein gradient algorithm for compressed sensing. Abstract and Applied Analysis, 2014:410104, 2014. ISSN 1085-3375 (print), 1687-0409 (electronic).

Rowe:2005:EDC

[R^{+}05] Kerry Rowe et al. Engines of Discovery: The 21st Century Revolution. The Long Range Plan for HPC in Canada. NRC Press, ????, August 2005. ???? pp.

Raydan:1993:BBC

[Ray93] Marcos Raydan. On the Barzilai and Borwein choice of steplength for the gradient method. IMA Journal of Numerical Analysis, 13 (3):321–326, 1993. CODEN IJNADN. ISSN 0272-4979 (print), 1464-3642 (electronic).

Raydan:1997:BBG

[Ray97] Marcos Raydan. The Barzilai and Borwein gradient method for the large scale unconstrained minimization problem. SIAM Journal on Optimization, 7(1):26–33, 1997. CODEN SJOPE8. ISSN 1052-6234 (print), 1095-7189 (electronic).

Reisner:2002:NTB

[Rei02] Shlomo Reisner. A note on a theorem of Borwein, Borwein, Fee and Girgensohn. Mathematical Inequalities & Applications, 5(4): 753–754, 2002. ISSN 1331-4343 (print), 1848-9966 (electronic).

Robin:2006:BRP

[Rob06] Anthony C. Robin. Book review: Pi: A Source Book. Mathematical Gazette, 90(518):375–376, July 2006. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL

https://www.jstor.org/stable/40378657.

Rockafellar:2020:RUD

[Roc20] R. Tyrrell Rockafellar. Risk and utility in the duality framework of convex analysis. In Bailey et al. [BBB^{+}20], pages 21–42. ISBN 3-030-36567-0 (print), 3-030-36568-9 (e-book). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Rajkovic:2009:GBC

[RP09] Predrag M. Rajković and Marko D. Petković. Generalized Borwein conjecture and partitions of natural numbers. Funct. Anal. Approx. Comput., 1(2):47–56, 2009. ISSN 1821-410X (print), 2406-1573 (electronic).

Raydan:2002:RSD

[RS02] Marcos Raydan and Benar F. Svaiter. Relaxed steepest descent and Cauchy–Barzilai–Borwein method. Computational optimization and applications, 21(2):155–167, 2002. CODEN CPPPEF. ISSN 0926-6003 (print), 1573-2894 (electronic).

Richter:1993:PSP

[RSP^{+}93] R. Bruce Richter, Josef Siran, Carl Pomerance, Kiran S. Kedlaya, Michael Golomb, John Sarli, David Borwein, Jonathan Borwein, Ignacy I. Kotlarski, and Horst Alzer. Problems and solutions: Problems: 10330–10337. American Mathematical Monthly, 100 (8):796–798, 1993. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Rudin:1989:PSE

[Rud89] Walter Rudin. Problems and solutions: Elementary problems: E3325. American Mathematical Monthly, 96(5):445, May 1989. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Reich:2015:IPO

[RZ15] Simeon Reich and Alexander J. Zaslavski, editors. Infinite products of operators and their applications: a research workshop of the Israel Science Foundation: May 21–24, 2012, Haifa, Israel: Israel mathematical conference proceedings, volume 636 of Contemporary mathematics. American Mathematical Society, Providence, RI, USA, 2015. ISBN 1-4704-1480-5 (paperback). LCCN

QA329 .I54 2015. URL http://www.ams.org/books/conm/636/.

Reich:2018:CMU

[RZ18] Simeon Reich and Alexander J. Zaslavski. Contractive mappings on unbounded sets. Set-Valued and Variational Analysis, 26(1): 27–47, March 2018. CODEN ???? ISSN 1877-0533 (print), 1877-

0541 (electronic). URL http://link.springer.com/article/
10.1007/s11228-017-0430-y.

Schoenberg:1987:APS

[SB87] I. J. Schoenberg and Peter B. Borwein. Advanced problems and solutions: Solutions: 6501. American Mathematical Monthly, 94 (4):390, April 1987. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Sch85].

Stodden:2013:SDR

[SBB13] Victoria Stodden, Jonathan Borwein, and David H. Bailey. “Setting the default to reproducible” in computational science research. SIAM News, 46(5):4–6, June 2013. ISSN 0036-1437.

URL https://sinews2.siam.org/DetailsPage/tabid/607/

ArticleID/351/%E2%80%9CSetting-the-Default-to-Reproducible%
E2%80%9D-in-Computational-Science-Research.aspx.

Singh:1984:ATS

[SBW84] S. P. Singh, J. W. H. Burry, and B. Watson, editors. Approximation Theory and Spline Functions. NATO Advanced Study Institute held at Memorial University of Newfoundland during August 22–September 2, 1983, volume 136 of NATO ASI Series, Series C: Mathematical and Physical Sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1984. ISBN 94-009-6466-8, 94-009-6468-4. ISSN 1389-2185. LCCN ????

Schoenberg:1985:PSA

[Sch85] I. J. Schoenberg. Problems and solutions: Advanced problems: 6501. American Mathematical Monthly, 92(8):595, October 1985. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [SB87, SH87].

Scherzer:2015:HMM

[Sch15] Otmar Scherzer, editor. Handbook of Mathematical Methods in Imaging. SpringerReference, New York, NY, USA, second edition, 2015. ISBN 1-4939-0789-1 (set), 1-4939-0790-5 (e-book). xviii + 2178 (3 volumes) pp. LCCN RC78.7.D53 H358 2015.

Sopyla:2015:SGD

[SD15] Krzysztof Sopyła and Pawe Drozda. Stochastic gradient descent with Barzilai–Borwein update step for SVM. Information Sciences, 316:218–233, 2015. CODEN ISIJBC. ISSN 0020-0255 (print), 1872-6291 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0020025515002467.

Nature-Inspired Algorithms for Large Scale Global Optimization.

Seife:2001:PKE

[Sei01] Charles Seife. Pi keeps ’em guessing. Science, 203(5530):??, July 27, 2001. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/news/2001/07/
pi-keeps-em-guessing.

Selin:2016:EHS

[Sel16] Helaine Selin. Encyclopaedia of the History of Science, Technology, and Medicine in non-Western cultures. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2016. ISBN 94-007-7746-9 (set), 94-007-7747-7 (e-book), 94-007-7748-5 (bundle). ???? pp. LCCN ????

URL http://link.springer.com/referencework/10.1007/

978-94-007-7747-7;http://www.loc.gov/catdir/enhancements/
fy1608/2015957805-b.html;http://www.loc.gov/catdir/enhancements/
fy1608/2015957805-d.html;http://www.loc.gov/catdir/enhancements/
fy1608/2015957805-t.html.

Schurger:1987:PSS

[SH87] Klaus Schurger and Ellen Hertz. Problems and solutions: Solutions of advanced problems: 6502. American Mathematical Monthly, 94(5):470–471, May 1987. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Sch85].

Shallit:2005:BRM

[Sha05] Jeffrey Shallit. Book review: Mathematics by Experiment, Jonathan Borwein and David Bailey, A K Peters, 2003, 288 pages, $45.00, ISBN 1-56881-211-6. Experimentation in Mathematics, Jonathan Borwein, David Bailey, and Roland Girgensohn, A K Peters, 2004, 357 pages, $49.00, ISBN 1-56881-136-5. Notices of the American Mathematical Society, 52(8):863–865, September 2005. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-

9477 (electronic). URL http://www.ams.org/notices/200508/
rev-shallit.pdf.

Shareef:2016:NCG

[SI16] Salah Shareef and Alaa Ibrahim. A new conjugate gradient for unconstrained optimization based on step size of Barzilai and Borwein. Science Journal of University of Zakho, 4(1):104–114, June 2016. ISSN 2410-7549 (print), 2414-6943 (electronic).

Simons:2018:QMM

[Sim18] Stephen Simons. Quasidense monotone multifunctions. Set-Valued and Variational Analysis, 26(1):5–26, March 2018. CODEN ???? ISSN 1877-0533 (print), 1877-0541 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s11228-017-0434-7.

Sole:1995:A

[Sol95] Patrick Sole. D_{4}, E_{6}, E_{8} and the AGM. In Cohen et al. [CGM95], pages 448–455. ISBN 3-540-60114-7 (softcover). LCCN QA268

.A35 1995. URL http://www.loc.gov/catdir/enhancements/
fy0815/95021560-d.html.

Solomon:2015:DS

[Sol15] Benjamin T. Solomon. ET, the deafening silence. Huffington Post, ??(??):??, August 10, 2015. URL http://www.huffingtonpost.

com/benjamin-t-solomon/et-the-deafening-silence_b_7608462.

html. See [?].

Sturm:1990:PSE

[Stu90] Jacob Sturm. Problems and solutions: Elementary problems: E3384. American Mathematical Monthly, 97(5):427, May 1990. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Sidoli:2014:ATB

[SV14] Nathan Sidoli and Glen Van Brummelen, editors. From Alexandria, Through Baghdad: Surveys and Studies in the Ancient Greek and Medieval Islamic Mathematical Sciences in Honor of J. L. Berggren. SpringerLink: Bücher. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-642-36735-6 (hardcover), 3-642-36736-4 (ebook). xv + 583 pp. LCCN QA21 .F76 2014. URL http://

scans.hebis.de/HEBCGI/show.pl?33313183_aub.html;http:
//scans.hebis.de/HEBCGI/show.pl?33313183_toc.html.

Skerritt:2020:EPA

[SV20] Matthew P. Skerritt and Paul Vrbik. Extending the PSLQ algorithm to algebraic integer relations. In Bailey et al. [BBB^{+}20], pages 407–421. ISBN 3-030-36567-0 (print), 3-030-36568-9 (ebook). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Stoica:2021:ICC

[SW21] George Stoica and Yousef Wardat. An inequality can change everything …. American Mathematical Monthly, 128(9):810, 2021. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Swedroe:2017:SGF

[Swe17] Larry Swedroe. Swedroe: Grading the forecasters. Web

blog., July 28, 2017. URL http://www.etf.com/sections/
index-investor-corner/swedroe-grading-forecasters.

Schaible:1981:CCP

[SZ81] Siegfried Schaible and William T. Ziemba, editors. Generalized Concavity in Optimization and Economics. NATO conference, Academic Press, New York, NY, USA, 1981. ISBN 0-12-621120-5. LCCN QA402.5 .G45. Papers presented at the proceedings of the NATO Advance Study Institute held at the University of British Columbia, Vancouver, Canada, August 4–15, 1980.

Sendov:2014:SBA

[SZ14] Hristo S. Sendov and Ričardas Zitikis. The shape of the Borwein– Affleck–Girgensohn function generated by completely monotone and Bernstein functions. Journal of Optimization Theory and Applications, 160(1):67–89, 2014. CODEN JOTABN. ISSN 0022-3239 (print), 1573-2878 (electronic).

Straub:2020:SWA

[SZ20] Armin Straub and Wadim Zudilin. Short walk adventures. In Bailey et al. [BBB^{+}20], pages 423–439. ISBN 3-030-36567-0 (print), 3-030-36568-9 (e-book). ISSN 2194-1009 (print), 2194-1017 (electronic). LCCN ????

Shallit:1986:PSS

[SZUM86] J. D. Shallit, Karel Zikan, University of South Alabama Problem Group, and Z. A. Malzak. Problems and solutions: Solutions of elementary problems: E2999. American Mathematical Monthly, 93(5):402–403, May 1986. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [ANO^{+}83].

Tijs:1980:SGC

[TB80] S. H. Tijs and J. M. Borwein. Some generalizations of Carathéodory’s theorem via barycentres, with application to mathematical programming. Canadian mathematical bulletin = Bulletin canadien de mathématiques, 23(3):339–346, September 1980. CODEN CMBUA3. ISSN 0008-4395 (print), 1496-4287

(electronic). URL http://docserver.carma.newcastle.edu.
au/1656/.

Theodorescu:2000:PSS

[TB00] Radu Theodorescu and Jonathan M. Borwein. Problems and solutions: Solutions: Moments of the Poisson distribution: 10738. American Mathematical Monthly, 107(7):659, August/September 2000. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972

(electronic). URL https://www.jstor.org/stable/2589135.

Thangadurai:2002:NCB

[Tha02] R. Thangadurai. A note on a conjecture of Borwein and Choi. Arch. Math. (Basel), 78(5):386–396, 2002. CODEN ACVMAL. ISSN 0003-889x (print), 1420-8938 (electronic).

Thera:2016:HSH

[Thé16] Michel Théra. Homo sapiens, homo ludens. Journal of Optimization Theory and Applications, 172:1–8, December 29, 2016. CO-DEN JOTABN. ISSN 0022-3239 (print), 1573-2878 (electronic). Paper in memory of Jonathan Borwein.

Takahashi:1997:IAC

[TK97] Daisuke Takahashi and Yasumasa Kanada. Improvement of the algorithms for π calculation: the Gauss–Legendre algorithm and the Borwein’s quartically convergent algorithm. Transactions of the Information Processing Society of Japan, 38(11):2406–2409, 1997. CODEN JSGRD5. ISSN 0387-5806.

Todd:2003:CAN

[Tod03] Michael J. Todd. Convex Analysis and Nonlinear Optimization: Theory and Examples. Jonathan M. Borwein and Adrian S. Lewis, Springer, New York, 2000 [book review]. International Journal of Robust and Nonlinear Control, 13(1):92–93, 2003. ISSN 1049-8923 (print), 1099-1239 (electronic).

Tressider:2013:EDV

[Tre13] Virginia Tressider. The explainer: from déjà vu to why the sky is blue, and other conundrums. CSIRO Publishing, Collingwood, VI, Australia, 2013. ISBN 1-4863-0050-2 (paperback). xiv + 334

pp. URL http://www.publish.csiro.au/pid/7165.htm.

Thera:2013:MRR

[TSB13] Michel Thera, Brailey Sims, and Jonathan M. Borwein. In memoriam: Robert R. Phelps (1926–2013). Math Drudge, February 17,

2013. URL https://experimentalmath.info/blog/2013/02/
in-memoriam-robert-r-phelps/.

Ullman:2021:PSi

[UVW^{+}21] Daniel H. Ullman, Daniel J. Velleman, Douglas B. West, Paul Bracken, Ezra A. Brown, Zachary Franco, George Gilbert, László Lipták, Rick Luttmann, Hosam Mahmoud, Frank B. Miles, Lenhard Ng, Rajesh Pereira, Kenneth Stolarsky, Richard Stong, Stan Wagon, Lawrence Washington, and Li Zhou. Problems and solutions. American Mathematical Monthly, 128(9):856–864, 2021. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Viazovska:2016:SPP

[Via16] Maryna Viazovska. The sphere packing problem in dimension 8. arxiv.org, ??(??):1–22, March 14, 2016. URL http://arxiv.org/

abs/1603.04246.

Virosztek:2014:PBB

[Vir14] Dániel Virosztek. On a problem of Bauschke and Borwein. arxiv.org, ??(??):??, December 5, 2014. URL http://arxiv.org/

abs/1412.2602.

Warnaar:2001:GBC

[War01] S. Ole Warnaar. The generalized Borwein conjecture. I. The Burge transform. In q-series with applications to combinatorics, number theory, and physics (Urbana, IL, 2000), volume 291 of Contemp. Math., pages 243–267. Amer. Math. Soc., Providence, RI, 2001.

Warnaar:2003:GBC

[War03] S. Ole Warnaar. The generalized Borwein conjecture. II. Refined q-trinomial coefficients. Discrete Mathematics, 272(2–3):215–258, 2003. CODEN DSMHA4. ISSN 0012-365x (print), 1872-681x

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0012365X03000475.

Ward:1987:NCF

[WB87] D. E. Ward and J. M. Borwein. Nonsmooth calculus in finite dimensions. SIAM Journal on Control and Optimization, 25(5): 1312–1340, September 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http://docserver.carma.

newcastle.edu.au/1601/.

Wei:2015:SBT

[Wei15] Chuanan Wei. Several BBP-type formulas for π. Integral Transforms Spec. Funct., 26(5):315–324, 2015. ISSN 1065-2469 (print), 1476-8291 (electronic).

Wuppuluri:2017:STL

[WG17] Shyam Wuppuluri and Giancarlo Ghirardi. Space, Time, and the Limits of Human Understanding. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2017. ISBN 3-319-44417-4, 3-319-44418-2 (e-book). xxiv + 530 pp. LCCN BD161.

Wimp:1988:BRP

[Wim88] Jet Wimp. Book review: Pi and the AGM: A Study in Analytic Number Theory and Computational Complexity. SIAM Review, 30(3):530–533, September 1988. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL https://www.jstor.

org/stable/2030735.

Wang:2007:PBB

[WM07] Yanfei Wang and Shiqian Ma. Projected Barzilai–Borwein method for large-scale nonnegative image restoration. Inverse Problems in Science and Engineering, 15(6):559–583, 2007. CODEN IPSECR. ISSN 1741-5977 (print), 1741-5985 (electronic).

Wang:2015:BBS

[WSdSY15] Liumei Wang, Wenyu Sun, Raimundo J. B. de Sampaio, and Jinyun Yuan. A Barzilai and Borwein scaling conjugate gradient method for unconstrained optimization problems. Applied Mathematics and Computation, 262:136–144, 2015. CO-DEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0096300315005044.

Wu:2016:BBL

[WSL16] Lei Wu, Zhe Sun, and Dong-Hui Li. A Barzilai–Borwein-like iterative half thresholding algorithm for the L_{1/2} regularized problem. Journal of Scientific Computing, 67(2):581–601, May 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/
s10915-015-0094-4;http://link.springer.com/content/pdf/
10.1007/s10915-015-0094-4.pdf. See [BB88a] for the original

Barzilai–Borwein work.

Xie:2011:SSC

[XC11] Zhipeng Xie and Songcan Chen. SCIHTBB: Sparsity constrained iterative hard thresholding with Barzilai–Borwein step size. Neurocomputing, 74(17):3663–3676, 2011. CODEN NRCGEO. ISSN 0925-2312 (print), 1872-8286 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0925231211003857.

Xiao:2008:SBB

[XH08] Yunhai Xiao and Qingjie Hu. Subspace Barzilai–Borwein gradient method for large-scale bound constrained optimization. Applied Mathematics and Optimization, 58(2):275–290, 2008. CODEN AMOMBN. ISSN 0095-4616 (print), 1432-0606 (electronic).

Xiao:2012:MCG

[XSW12] Yunhai Xiao, Huina Song, and Zhiguo Wang. A modified conjugate gradient algorithm with cyclic Barzilai–Borwein steplength for unconstrained optimization. Journal of Computational and Applied Mathematics, 236(13):3101–3110, 2012. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S037704271200043X.

Xiao:2014:NBB

[XWQ14] Yunhai Xiao, Soon-Yi Wu, and Liqun Qi. Nonmonotone Barzilai– Borwein gradient algorithm for ℓ_{1}-regularized nonsmooth minimization in compressive sensing. Journal of Scientific Computing, 61(1):17–41, 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

Xia:2012:ESI

[XY12] Ernest X. W. Xia and Olivia X. M. Yao. Eisenstein series identities involving the Borweins’ cubic theta functions. Journal of Applied Mathematics, 2012:181264, 2012. ISSN 1110-757X (print), 1687-

0042 (electronic). URL http://www.hindawi.com/journals/
jam/2012/181264/.

Yang:1994:EBP

[Yan94] Xin Min Yang. The equivalence of Benson proper efficient solutions and Borwein proper efficient solutions. Math. Appl. (Wuhan), 7(2):246–247, 1994. ISSN 1001-9847.

Yongxin:2000:GEV

[YS00] Li Yongxin and Shi Shuzhong. A generalization of Ekeland’s ϵvariational principle and its Borwein–Preiss smooth variant. Journal of Mathematical Analysis and Applications, 246(1):308–319, 2000. CODEN JMANAK. ISSN 0022-247X (print), 1096-0813

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022247X00968136.

Yuan:2012:BBG

[YW12] Gonglin Yuan and Zengxin Wei. The Barzilai and Borwein gradient method with nonmonotone line search for nonsmooth convex optimization problems. Math. Model. Anal., 17(2):203–216, 2012. ISSN 1392-6292 (print), 1648-3510 (electronic).

Zaharescu:2006:BCA

[Zah06] Alexandru Zaharescu. Borwein’s conjecture on average over arithmetic progressions. The Ramanujan Journal, 11(1):95–102, 2006. CODEN RAJOF9. ISSN 1382-4090 (print), 1572-9303 (electronic).

Zalinescu:1986:LEJ

[Zăl86] Constantin Zălinescu. Letter to the editor: on J. M. Borwein’s paper: “Adjoint process duality” [Math. Oper. Res. 8 (1983), no.

3, 403–434; MR0716121 (85h:90092)]. Mathematics of Operations Research, 11(4):692–698, November 1986. CODEN MOREDQ. ISSN 0364-765X (print), 1526-5471 (electronic). URL https://

www.jstor.org/stable/3690014. See [Bor83a].

Zeilberger:2005:SSM

[Zei05] Doron Zeilberger. Serving a silicon master: Book reviews: Mathematics by Experiment: Plausible Reasoning in the 21st Century. Jonathan Borwein and David Bailey, x + 288 pp. A K Peters, 2004. $45. Experimentation in Mathematics: Computational Paths to Discovery. Jonathan Borwein, David Bailey and Roland Girgensohn. + 357 pp. A Peters, 2004. $49. American Scientist, 93(2):182–183, April 2005. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL https:

//www.jstor.org/stable/27858557.

Zhang:2006:PPA

[ZH06] Hongchao Zhang and William W. Hager. PACBB: a projected adaptive cyclic Barzilai–Borwein method for box constrained optimization. In Multiscale optimization methods and applications, volume 82 of Nonconvex Optim. Appl., pages 387–392. Springer, New York, 2006.

Zhao:2010:CBB

[Zha10] Jianqiang Zhao. On a conjecture of Borwein, Bradley and Broadhurst. Journal für die reine und angewandte Mathematik, 639: 223–233, 2010. CODEN JRMAA8. ISSN 0075-4102 (print), 1435-5345 (electronic).

Zhang:2013:NFB

[Zha13] Wenlong Zhang. New formulae of BBP-type with different moduli. Journal of Mathematical Analysis and Applications, 398(1):46–60, 2013. CODEN JMANAK. ISSN 0022-247X (print), 1096-0813 (electronic).

Zhou:2012:EBP

[Zho12] Zhi Ang Zhou. Equivalence of the Borwein properly efficient solution and the Benson properly efficient solution of a set-valued optimization problem. Math. Pract. Theory, 42(1):247–250, 2012. ISSN 1000-0984.

Zhuang:1991:BCC

[Zhu91] D. M. Zhuang. Bases of convex cones and Borwein’s proper efficiency. Journal of Optimization Theory and Applications, 71 (3):613–620, 1991. CODEN JOTABN. ISSN 0022-3239 (print), 1573-2878 (electronic).

Zhang:2022:BBR

[ZL22] Yaling Zhang and Hongwei Liu. A Barzilai and Borwein regularization feasible direction algorithm for convex nonlinear SOC programming with linear constraints. Journal of Computational and Applied Mathematics, 401(??):1–15, February 2022. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0377042721004039.

Zhang:2012:BTF

[ZS12] Wenlong Zhang and He-Shan Song. BBP-type formulae modulo 5. Utilitas Mathematica, 89:107–112, 2012. CODEN UTMADA. ISSN 0315-3681.

Zhang:2010:NFB

[ZSQ10] Yan Zhang, Wenyu Sun, and Liqun Qi. A nonmonotone filter Barzilai–Borwein method for optimization. Asia-Pac. J. Oper. Res., 27(1):55–69, 2010. ISSN 0217-5959 (print), 1793-7019 (electronic).

Zhou:2016:NSM

[ZSZ16] QunYan Zhou, WenYu Sun, and HongChao Zhang. A new simple model trust-region method with generalized Barzilai–Borwein parameter for large-scale optimization. Science China. Mathematics, 59(11):2265–2280, July 2016. ISSN 1674-7283 (print), 1869-1862 (electronic).

Zhang:2014:NFB

[ZZ14] Wenlong Zhang and Ying Zhang. New formulae of BBP-type modulo 6. Quaest. Math., 37(3):393–400, 2014. ISSN 1607-3606 (print), 1727-933X (electronic).